A Complete Bibliography of Publications in Scientometrics: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

23 March 2022 Version 1.53

Title word cross-reference

#Psychology [VSK18].

+ [RCN^{+}14]. 013 [dCPF14]. 10 [TBW^{+}12]. 20 [LGH^{+}14]. 3 [HZQ^{+}17, ZLLD19]. 9 [vdBBdK16]. ^{2} [TSMTDLCH11]. a [San12c]. C^{3} [PPM^{+}17]. ch [FMPP10]. χ [LFBI19]. d [DCS12]. f [FESD11]. g [ADV13, ACHVH10, BB16, LF12a, SS14]. H [BLA16, Egg10e, Egg11e, Egg11f, Lun19, MHLGHV14, MHC^{+}15, MZE19, ZYZ14, Abb13, ADV13, uARA19, ACHVH10, AA19, Ask18, AA16, AMI18, BL10, BIL15, BI18a, BK11, BBL17a, BBL17b, BW10, BSMD11, BL18, CG14, CF18, DGDG11, Egg10d, Egg11b, Egg11c, Egg13a, Egg13b, Egg14b, ER19a, FGMM12, FMM13a, GP13, GBB15, GQAM19, Glä10, Glä12, HAA14, HBA19, HSW10, HK12, Jac12, Laz10, LF12a, LZGQ13, MR13, MJHG13, MPH19, MKHB15a, MKHB15b, OBG11, PDAN19, Pra10e, Pra17a, Pra19c, Pra19d, RAA18, Rya16, Saa10, San12c, SCGZR16, SS10b, Sch15b, SS14, SJOC18, TH19, Tur16, YR10, YAC10, ZGY16, ZGJ18, dS18, dSD18b, dSD18c]. H_{2} [VB12]. h_{α} [Hir19b, LBO19, Hir19a]. H_{l} [ZYZ14]. h_{m} [TH19]. h_{men} [Ask18]. h_{T} [Pra10b, San12c]. hg [ACHVH10, FM11b]. hw [BIL15]. j [Tod11]. N [Par14c, BASL16]. P [dFVDU^{+}19, AA10, Pra10a, Pra11b, Pra19c]. ϕ [Cab13]. π_{v} [Vin10b]. q [AdAdAM10, CS11a]. R [San12c]. S [LW10]. φ [SOBM16]. z [ZCL14]. zp [ZP16].

-0967-y [dCPF14]. -Classics [MHLGHV14, MHC^{+}15]. -core [YR10]. -cores [SCGZR16]. -curve [LW10]. -exponential [AdAdAM10]. -grams [BASL16]. -Index [SJOC18, uARA19, ACHVH10, AA19, AA10, AMI18, BL10, BIL15, BI18a, BK11, BBL17a, BBL17b, Cab13, CF18, DCS12, DGDG11, Egg10d, Egg11b, Egg11c, Egg13a, Egg13b, ER19a, FMPP10, FM11b, FGMM12, GQAM19, Glä10, HAA14, HSW10, Jac12, Laz10, LFBI19, LF12a, MJHG13, MKHB15a, MKHB15b, OBG11, PDAN19, Pra10a, Pra10e, Pra17a, Pra19c, RAA18, Rya16, Saa10, SOBM16, Sch15b, TH19, Tod11, Tur16, Vin10b, YAC10, ZP16, dS18, dSD18b, dSD18c, BLA16, BB16, Egg11e, Egg11f, FMM13a, Lun19, MHLGHV14, MZE19, PPM^{+}17, ZYZ14]. -Index-based [ZGJ18]. -indices [ACHVH10, MPH19, Pra11b]. -like [BW10]. -rank [BIL15]. -related [ZGY16]. -score [ZCL14, dFVDU^{+}19]. -sequence [Egg10e]. -similarity [SS10b]. -Tuple [Par14c]. -Type [Abb13, Egg11c, Egg14b, Glä12, LZGQ13, MR13, Pra19d]. -Value [FESD11]. -year [LGH^{+}14, TBW^{+}12, vdBBdK16].

//doi.org/10.1007/s11192 [Glä18]. //doi.org/10.1007/s11192-016-2142-8 [Glä18].

10.1007/s11192 [Ano18b, dCPF14]. 10.1007/s11192- [dCPF14]. 10.1007/s11192-017-2490-z [Ano18b]. 10.1007/s11192-017-2506-8 [Ano18b]. 109 [Glä18]. 10th [Ano10]. 11 [KGB^{+}18]. 12th [KÖG12, LLRG10]. 13th [KG13, OING12]. 14th [Ano15, GGH^{+}14]. 15th [SSAG16]. 1970s [NBR^{+}11]. 1980s [Sot12]. 1997-2017 [LRY18]. 1o/oo [Kos18a]. 1st [BMM14].

2 [LGD11]. 20-year [ML18]. 2000-2015 [GVS17]. 2003-2012 [Liu16]. 2009 [Ano11]. 2010 [GGG^{+}11, JG14, VB12]. 2011 [Ano12c, GGG^{+}12, OL11, Pra12a]. 2012 [BMM14]. 2013 [Ano14, Ano15]. 2015 [KL16, SSAG16, Tsa15]. 2017 [Ano17a]. 2019 [Dan19]. 2020 [EC16]. 20th [LHG16a]. 21st [ACORC10, MCL^{+}11, Sot12]. 22-year [KKT^{+}18]. 28 [BB19].

30-year [MDDG17].

4.0 [DDS^{+}19a].

5 [VFA10]. 5-top [TBB^{+}16]. 50-year [PFL19].

6 [VFA10]. 6th [OA10a, OA10b, OA10c].

7 [BPHL16]. 7th [KÖG12].

87 [OL11]. 88 [Pra12a]. 8th [KG13].

97 [dCPF14]. 983 [Ho16].

= [Rou19].

A&HCI [WF18]. A. [Fie15b]. ability [Cab13, HH19, Rou12a, SOBM16, Sch12b]. above [GRSFV11]. abridged [MAA^{+}11]. abroad [PD10, RCdJ^{+}14, SL17]. ABS [WLZ^{+}15]. Absolute [MKHB13a, MKHB13b]. absorptive [MJC12]. abstract [SI17, ZXT^{+}19, Zon19]. abstracts [KD18, LY16a, Pau10, RM18]. Abt [Bur12]. abuse [VASNU^{+}19]. academia [GB14a, HO19, HYC15, Lei16, OK13, Oze12a, Oze12b, BKG16, BA15]. Academic [CHM15, CXZ19, DY18, DC17, Har16d, LCY14, MCCU16, NZL^{+}19, SM14, TCR10, Tor13, YC10, ZW17b, ZL15b, VHD^{+}16, AAH10, ADD11d, ADM14, ADR16a, ADD18b, ACFL11, ADM19, AI17, AÇA^{+}14, ÁCCG^{+}15, BHB13, BI18a, BCT19, Bre13, CMUdF15, Car16, CÖT16a, CHC13, CB18, CR18, DTM^{+}13, DRMMC19, DMV10, Doc13, DRS14, DJWS11, DGF17, Dya14, Egg11d, Eld19, FYC15, FFL16, FSC14, FDVZ16, pGSyW^{+}19, GRSS16, GRSFV14a, GRSFV18, GRSFV19b, GVS17, GB16, GB17a, GB17b, Gus19, HAL11, HK19, HJL18, HC14a, HC15a, HC16a, HB18b, HFW^{+}14, HG13, Hud17, IT11, JKPL18, JBMR11, JL18b, JKSK15, Keg15, KBT14, Ken18, KG10b, KJW^{+}17, KM12, KR17b, LCFC14, LM15, LDG17, LLX^{+}18, LZCZ18, MPY^{+}13, MR15, MM18, MM19, MARMSG19, MMOMALC16, MMOMLC18b, MR18b, MT12a, MHKB16, OVJM17, OKCPS17, OMOR13b]. academic [OMR14, OMMMTLC17, Ort15, dFPYdCL12, PNVCB18, PTMT10, PTMT11, PPI17, RG12, Rha17, RPDCRVRP15, RD13, SBT18, SHB14, SS14, SHS15, SZAS16, Shi11, SC10, SPdSM16, Sku19, SM17, TSRGG17, Tor14, VB12, WM19, WHLP16, WYB^{+}17, WBX^{+}17, WZ19b, XWL19, XZFD19, YQX10, YS13, YZ17, Yur15, Yur16a, ZZL^{+}10, ZMW^{+}18, ZZL19, ZZLS19, ZL15a, ZGL^{+}17, ZB15, ZYS16, Zhu17, dSNV18, dSD18b, dSD18c, vPD13, HA17a, HA17b, Har19b, HHBB18, HOB17, HB17b, Saf19, The18a, The18e]. academically [PL18]. academicians [DBO^{+}18, KKE13, YGW^{+}15]. academics [Ben15, BC13b, COS11b, HM15a, HM15b, ILP13, KM12, MS13, Mix18, San18, TV17, Yur18b, Yur18c, ZW17a]. Academy [CMPD19, Fan15a, YGW^{+}15, ZCZ^{+}16, KHH18, CD16a, HSBW10, WS13b]. accelerated [VGPdlC^{+}17]. Acceleration [Kra10, San12c]. acceptable [KB11b]. acceptance [AND19, HTL15, KBAK17, Oos15]. accepted [BHDI18, TC11, TC13]. acces [GBDG19]. Access [AC12, GOPG13, SMLHCP17, Wra16a, AOFU10, AChO19, Asa19, Ase10, BND11, Cop19a, DGGBDG17, EdS19, Fuk17, GFC18, GWA14, Hen19, Hen20, JN15, JMM18, KPJT14, Laa14, LL16, LP18a, MA19, OMLC15, PB18, PP16, SBD^{+}19, Sni16, SE18, yTwTlW19, VZAMG19, WLMF15a, WLF15, WLMF15b, WZCC19, YXW18, ZW18a, ZW14, Zhu17, MSB18, PROG19, Wra16b, ZW19]. accessibility [RWG^{+}15]. accident [OMJLVSN19]. acclimation [MSP^{+}15]. accommodate [DdS19b, HAA14]. accomplishments [Hei13]. according [AYS16, BS15a, HV18b, OMLC14, PYW18, SOBM16, SYDW19, Waa13]. account [GTMRE^{+}16, GTMRE^{+}19, Hir10, LIdMAM11, NF13, ZL18a]. accounting [ADR13, BKL15, SP12b, Tol11]. accreditation [STCRPA18]. Accumulation [Pet18a]. Accuracy [Cab11, DRS18, KT15, ÁBMB17, Don17, Lop10, MVS10, Ort18, PLWS14]. Accurate [Mue16, KB10, MCR^{+}12]. accurately [MB10a]. achievement [NASR11]. Achilles [HLSC18]. acknowledgement [MK18, PHDC16, RJ14, WS11]. acknowledgements [Mor16, Rig13]. acknowledgment [SMAABJ11]. acknowledgments [ÁBMB17, McC18]. acquisition [GSM^{+}16, PYK13]. across [ADM14, ADD14b, BKG16, BM12b, Cha16, CZPR17, DMM13, DGDGSV15, GRSFV14a, GZ14a, GNS^{+}15, HL13, Hur17, KR17b, LNMQRR15, MS18a, ÖS17, PPI17, SC10, Sug11, TCB16]. Act [TR14]. Active [HWQ^{+}18, WJCC19]. activism [ML16]. activities [AAH10, ADD17b, GGG14, GRSFV14a, GL15, HAL11, HMCD^{+}19, HR15, KM12, LZZ^{+}12, LM13b, MT15, Ort17, PS13, TR14, WS10, ZS17]. activity [BGM17, CP12b, GZ11, HHK^{+}12, HHZ14, HIG^{+}17, Kaz15, LMdBG16, MB10b, NTM^{+}18, PRDG17, Rou19, TBB^{+}16, WS13a, YYS^{+}10, ZZ14]. actor [RTP17]. actual [LL16, WZCC19]. acupuncture [FZZ^{+}12b]. adaptable [SSG^{+}18]. adaptive [VHH16]. added [MMSS11, TSG13]. added-value [TSG13]. Adding [Wal16, HO19]. additional [CC14]. additive [LNRSRBB18, TTC17]. address [AKB^{+}10, Kra13, MAGAM13, SRW18]. addresses [MSA13]. adequate [SLISC17]. adjectives [Lei16, OK13]. Adjusted [Doc12, OCM^{+}12]. Administration [CDD15, CAS16, QA18]. administrative [CYW^{+}11]. administrators [ALH15, LZCZ18]. adopting [AD16]. adoption [HPBI^{+}14]. advanced [Cho12, RTP17, SHR^{+}10]. advancement [ADR14a, HEH17, Lar12]. advantage [DGGBDG17, GK18, GXC^{+}19, Mik17, Rou19, SGY15, SE18, WLMF15a, WLMF15b, WG11]. advantages [DLL^{+}16b, MSDJ19, The18d, Wan16]. adverbs [Lei16, OK13]. adversely [ZHG16]. advertised [SA16]. advertising [FZQ17]. advice [FB10, LMKG19]. advisee [DRMMC19, LTK^{+}18]. advisor [DRMMC19, HB18b, LTK^{+}18]. advisor-advisee [LTK^{+}18]. Aedes [VACCAJ18]. aegypti [VACCAJ18]. affairs [dJC15]. affect [Asu19, CMUdF15, CRMdMA15, GYZ15, HC16a, JJS^{+}12, LCZ17, Sch10b, WG11, Yur16a]. affecting [ACFL11, ACMP13, CC13, IMHG12, qJnShPL17, TAA16]. affects [ZHG16]. affiliated [CAV^{+}19, YP19]. Affiliation [FD14, IFT^{+}18, VO17]. affiliations [BBP14, CLB13, HL17, SGM^{+}16, TA14a, ZSCR^{+}18]. affinity [CRBRG^{+}18, VYL17]. Africa [Pou10, AGHL14, Asu19, Ban18, CdJD15, Fed13, ILP11, ILP13, LMM15, MOA16a, Mêg13a, Mêg13c, OM11, PMJF19, Pou10, PP11, Pou12, PH14, SdJDD19, SS10c, Soo10a, Soo10b, Soo11a, Soo14a, Soo17, Soo18, Soo19, ZL18a]. African [Bos10, vWBS^{+}16, CG15b, KB18, KCM19, Mat12, Mat13, Med15, MPF18, PM18, PB12, QA18, TP11, ZCKZ16]. after [BL11b, HM18, IJF16, KGB^{+}18, Kra10, Mix18, Soo10b, WHW^{+}19]. against [CFS18, Fed13]. Age [Soó14b, SST^{+}16, ADD18b, Abt12, Bjø19, Bur12, CB11, Egg10c, GB17a, JKJL14, MZ14, Par14a, RAA18, San12b]. age-independent [San12b]. Age-sensitive [Soó14b]. ageing [GZ17, ZG17d]. agencies [HH18]. Agency [EBR16, CFP14, MÁB18]. agenda [HS16b, Jon10]. agendas [HS16b]. Agent [Ahr17, NH11, BGBS18, KM11, KPRT16, KTRP17, RT17, WG11]. Agent-based [Ahr17, NH11, BGBS18, KPRT16, KTRP17, RT17, WG11]. ages [WYB^{+}17]. agglomeration [BS13b]. Aggregated [Bha11, XLR15]. aggregates [ACORC11]. aggregating [GRSFVdMA14, dZLwC^{+}15]. aggregation [DR10b, Yu15]. Aggregative [YKLK14]. aging [Bou11, WZFD19, ZG17e]. agony [GGG16b]. agree [BD16a, BGJ^{+}16]. agricultural [DKS18, Fuk19, KHJ^{+}12, LLL12, PPI17, WAT16, ZL18b]. Agriculture [CP12b, BBJS16, SK16, ZZY13, KSB11]. Agrifood [BS11]. AHFS [TKA17]. AIB [WS13b]. aid [HN16]. aided [Glä15]. aiding [Che12]. AIDS [Lja16, PP11, QA18, WT14]. air [BSS15]. al [Ho16, OL11, Pra12a, Pra16b, dM10]. al. [BD16b]. Albuquerque [Bas14]. Alex [PT17]. algorithm [AT18, AA18, Cop19b, Kim19b, PLT14, Ric15, SH19]. algorithmic [Oli15a]. algorithms [CL17b, GSM^{+}16, JSZ13]. all-elements-sleeping-beauties [LY12, Li14]. Allegation [BH18a]. allergy [VSVR15]. allied [WOW13]. allocate [Osó18]. allocation [ACP12, BZ17, HBT16, KD14, WFZD19b, ZM16]. allocations [Gal11]. Allometric [CB16, DLL^{+}17]. allow [BM14a, KR17a]. allowed [Laa14]. Almeida [dCPF14]. almost [BBL17a]. almost-exact [BBL17a]. Alone [Mus12, NPT^{+}15]. along [BHPVdPMR18, HG13]. alphabetical [LF14a]. alphabetization [Yur16a]. also [BC13b, Ley15a, vR17]. alter [DWGL16]. Alternative [Bor15a, hCcTmWH15, FMP17a, FGMM12, FZZ^{+}11, HBA19, KTRP17, LBA19, MBT16, RCETS19, VB12, VRF12, WWC19, XWL19, YG18, ZCW14]. Alternatives [Ley12, OBG11]. altmetric [CGG19, HBS^{+}19, HP18, Ort18, GGG16b, HB17a, MSC18]. Altmetric.com [Ort18]. Altmetrics [ENST16, RCETS19, Bor15a, Eld19, GG15a, Ham14, HIG^{+}17, HPBI^{+}14, Hau16, HV18b, LSE^{+}18, MS18b, NCG^{+}19, PP18, PKL^{+}16, SBA^{+}19, ST14b, Yu17, ZCW14, ZWZ^{+}19, ZWX22, BB17b, ZW17c].

Alzheimer [CWJC14, SHL15]. ambidexterity [GLÚGML16, HWLL14]. ambidextrous [DY18]. ambiguity [GKF17]. ambiguous [Cop19a]. ambivalences [LWB16]. America [BB19, BMTA15, CR14, CRLMLM17, CS19, MC10, MSB18, SPB18]. American [LV12, ZGL^{+}17, AR18, ALYZ15, Ano18c, CSR^{+}18, CRZGVQMA15, CRZGVQdMA16, CRLMRPA17, GWBSVWB13, MHC14, OMLC15, PEPUT15, RPDCRVRP15, RG18, SLGO17, SS14, TC11, TC13, WMT^{+}12, WXLL12, ZGCRVQ18]. ammunition [ACMP13]. among [BHKP11, BPJ^{+}14, BPHL16, CGKB18, CG17, Cho12, DCS12, DCY^{+}17, FK16, FFL16, GGG14, Hal13, HF19, HM15c, HLL14, IBL13, JX13, Kim10, KFKS15, Ley11b, LZ14, LWB18b, MC13, MHM^{+}13, MK19, OO12, Ort18, PML^{+}17, PSY^{+}19, QDY14, SFBS17, SS10c, SM16b, SVCFI14, SN10, WTM^{+}16, YZB18, Yur18a, dCdAMB19]. amongst [Hud17]. amount [GRSFV19b, LAL15]. Amsterdam [Ano11]. anaerobic [ZLT^{+}14]. anaesthesiology [CAV^{+}19]. anaesthetists [OO12]. analgesia [KB11b]. analgesic [Kis11a]. analogy [FM11c, Sma10]. analyses [ÁBV^{+}14, BB15, CWL10, CHL15, CLHH10, IS16, Kra10, Kra16, KPY16, Lee10b, Pra14a, Sch16, TCC17, Wal16]. Analysing [BSPL19, Fan18, MAGSTRC15, OI17, RY14]. Analysis [ÁCCG^{+}15, BPVM11, CHY16, CXZ19, EKR19, FMPP10, Gom19, Ioa06, IJF16, ICC16, JG12, KM18a, KKLP17, KBZS15, MVS10, MYN^{+}15, PV15, Pen19, RC13a, RBC^{+}10, SIS17, YLL15a, YL12, YST12, AHUR11, ADV10, ADD11c, ADS17a, ADS17b, ADD19b, Agu12, APPS15, AG13, AR18, AW10, ABGS14, AYS16, AW11, ABL17, ACD14, Ase10, ANA18, BB10, BL10, Bai18, BY13, BHPVdPMR18, BLdlCV17, BHKP11, BPJ^{+}14, BVB13, BHS14, BK11, Bar17b, Bas11, BMM17, BFGVV^{+}18, BFS17, BL11a, Bel13, Ben11, BGMB16, Bha16, BP11, BBP14, BS17, BMTA15, BWD10, BHJD12, Bor15a, Bor16, BYY17, BHH18, BL13, BZBLP16, ByLbH16, BWbH^{+}18, Bue15, BS19, ÇAAÇ15, CD16a, CD17, CHWL12, CZW13, CD14, CVD14, CD16b, CB15, CLL^{+}17, CHC17, CZPR17, CL11, Che11, CBF13, CWJC14, CJY^{+}15, CLSW19]. analysis [CLW^{+}19, CY13, CGG19, CRZGVQMA15, CS11b, CL17a, CYK^{+}11, CBKL13, Chu14, Cle16, CFM15, CG15b, CPF18, CdMCdMMdP17, CvLvR11, CV14, COS11b, CNC18, CAS16, DSG^{+}15, DRMMC19, Dan14, Das16, DMM13, DVB14, DVB15, DGPL15, DC17, DB19, DXL^{+}12, DGGBDG17, DB16, ES16a, EMSH16, EDEH16, EBK16, Eld19, EW15, Ell18, EC16, ÉMS^{+}13, ENST16, EN17, ET15, ES18, Etz13a, FJ11, FZQ17, Fan18, FMU16, FZZ17, FFR^{+}17, FS12, FMP17a, FFR16, FRPP17, FP18, FdSdO17, FB16, FRF^{+}19, FM11a, FM12, Fra17, FSLR10, FLB19, FZZ^{+}11, FZZ^{+}12b, FLH14, FH16, Fuk14, Fuk17, FK17, FK18, FSO11, FSOS12, GW15a, pGDTP12, GGG14, GG14, GCLcG15, pGSyW^{+}19, GREL14, Gar15, Geo17, GE11, GM12, GW10a, Glä12, GHT17, GNVQdMAG11, GTGABAG15, GTAG18, GdOdAG^{+}13, GWA14, GK19, GFK^{+}18, GBGB13]. analysis [GB12, GHA^{+}15, HG10, HAL11, HT11, Hal13, Ham11, HM18, HH17a, HSAK18, HB18a, HB18b, HLSC18, HMCD^{+}19, HL13, Ho14, HH15c, HH17c, HLC17, Hol10, Hos11, Hou17, HYC18, HTL15, HC15c, HC17, HHGZ11, HHDL13, HZ17, HWQ^{+}18, HWS18, HYYL12, HLY14, HC16b, HFC11, HZD^{+}15, HYC15, HHA^{+}16, HH17e, HH18, HYS18, Hud16, HOB17, HLLT14, IWK18, IA19, IQT^{+}19, IH14, IFH15, JYW11, JC12, JPZ14, JBC19, JDG14, JMM19, KGNB11, KG10a, Ke13, KZ13, KW15, KLPP16, Kim10, Kim14, KGL^{+}14, KPL19, KHK13, KB18, KGB^{+}18, KGZML^{+}19, KS17, KG16, KM16, KM18c, LP18a, LGR17, Lam12, LBCO19, LGL10, LKP11, LVGV^{+}11, LGPC18, LLL12, LKS^{+}14, LKSK15, Lee15, LJS16, LK17, LPMK17, LL19, LZZ^{+}12, LZZ^{+}13, Ley13a, Ley13b, LKY17, LYWSV13, LGZ^{+}13, LJKG15, LSS15, LSS16, LQW17, LWL17, LWR^{+}17, LP18b, LRS^{+}18, LRY18]. analysis [LYW19, LC18, LCFC14, LW15, Lin12, LCY14, Lin18, LD16, LW10, LSY11, LHW12, LFLG14, LXL15, LWM^{+}15, LTGH15, LYLD15, Liu16, LXL16, LLW^{+}16, LCLX16, LM16, LZC17, LTK^{+}18, LLH19, LdSdFFNM17, LDVSGDR16, LDVSGD19, LJMF15, LA19, LCIADG19, LAW14, LDZ17, LWW^{+}11, LPZ17, MSYW12, MH16a, ML18, MC13, Mad15, MnaeR^{+}15, Mag14a, MM14a, MB19, MWH14, MBL18, MJC14, MDDG17, MGMY^{+}18, MCvFP16, Meh07, MCCU16, MF14, MdFdA^{+}14, MTT15, MSB18, MK19, Moe17, MPH16, MD18, MdNS^{+}19, MT12a, MDFGAM14, MEG15, MFF^{+}16, MHM12b, Mue18, MRR17, MAGSTRC15, MKF14, NTM^{+}18, NRAW17, NSH^{+}11, NSKO15, NFH12, NP11, NHLL17, NBR^{+}11, NPP^{+}12, NHY^{+}14, NT17, OGOPPR17, OZK11, OMOR13a, OMLC14, OA10c, Ort17, OMJLVSN19, OE15, PFL19, PFDL17, PJL19, Pan14, PY14, PLJ18, PHBN^{+}15, PSB^{+}17, PLW^{+}15, PR10, PCRMCB^{+}18, PRSB16, PKL^{+}16]. analysis [PhD18, PQG14, PFPCM^{+}19, PMN16, PHS12, QRJ^{+}17, QL12, RG15, RMCM13, RM18, RAS15, RR17, RGGBV16, RJ14, RCJ18, RFGBMA13, RBBG18, RRRGVD19, SR15, Saf19, SC18, SFBS17, SH15a, SZMS17, SB15, SMCC18, SBB16, SHS15, SD13, SK17, SC10, SFNO12, SWH14a, SWH14b, SDS14a, SLD^{+}17, SL12a, Sku19, SM15, SY16b, SHL15, SZD16, Soo10a, Soo17, Soo18, SÁV18, SZ15, SDEB16, SCLC15, SZ12, tScL13, SG16, SX16, SZAJS14, TG16, TFH14, TUCR15, TCH^{+}15, TA18, TA11, TK16, The18a, The19a, TBMM18, THFBdMA18, TW16, TP11, TSMTDLCH11, Tsa11, yT11, yT15, yTnL17, TT13, TCT^{+}13, TE18, UHAR12, VH17, VMM15, VSVR15, VRF12, VASNU^{+}19, VuHL10, VYL17, VO17, VT10, VSK18, WRM17, WvEvL^{+}11b, WYH10, WS11, WLLL12, WLHZ13, WZLZ13, WL14, WLR^{+}14, WLD^{+}14, WCB^{+}15, WZW15]. analysis [WTM^{+}16, WXZ^{+}16, WH16, WDS16, WMH^{+}17, WC18, WRC^{+}19, WHW^{+}19, WL18, WP17, WOW10, WOW13, WY12, WK15, WKK16, WW11, WLH^{+}17, WHH^{+}18, XZZ15, Xie15, XGCK19, XM13, YDZ10, YDJ12, mYqS15, YLL15b, YLC18, YMSQ10, YPH10, YWC12, YCL^{+}13a, YCL^{+}13b, YZW^{+}17, YQW13, YCK11, YK12, YPK13, Yos13, YLH^{+}17, YLJ^{+}17, YWY10, YYLW14, Yur18b, ZCW14, ZAJ19, ZL18b, ZLYF14, ZYSS14, ZLG^{+}15, ZXH10, ZJLG10, ZG12a, ZLT^{+}14, ZZ14, jZhLY15, ZLL^{+}15, ZXM^{+}16, ZG17c, ZWW^{+}18, ZWZ^{+}19, ZWX22, ZS11, ZC14, ZZ16, ZyZZ^{+}11, ZZZ^{+}12, ZyZZ^{+}14, ZZZ^{+}14, ZWW^{+}15, ZWW^{+}16, ZGL^{+}17, ZG10, ZXLX14, ZZP^{+}14b, ZP15, ZZ15, ZHZY19, ZG13c, ZY15, ZH17, ZLN^{+}13, ZSY^{+}13, ZCMVQS11, dART^{+}17, dJC15, dPSS18, dPdCAdMC^{+}16, ddMS15, KK15, vLvWW16, vWBS^{+}16, vdBSS18, JG14, Mix18, Puu10]. analysts [Lyk18]. analytic [CZV10, Pet18a]. analytical [Ano16c, CMPD19]. analytics [BLdlCV17, RTP17]. analyze [CB16]. analyzed [PY19]. Analyzing [CC10a, GZ17, HH15a, KY17, KKK^{+}14, MP15, PJY17, Sah16, SHK14, TSRGCCJC14, WCL14, YZB18, ZFY^{+}17, BCML19, BHL^{+}10, CIK^{+}18, GD16, HB17b, Ma12, SCL12, VuHL10, MR18a]. anatomist [ÁBV^{+}14]. anatomy [Soó11b]. anchoring [CdMCdMMdP17]. anchors [AA18]. and-independent [WFH^{+}16]. Andras [Pen19]. anesthesia [WCK^{+}12]. anesthesiology [PFL19, TBW^{+}12]. angel [Zit12]. Angewandte [BD10b]. angiogenesis [EG18]. anglophone [Ben19]. animal [GMJ^{+}17, KKBW17]. animation [KdBBK15]. Anna [MB13]. Annals [ZGL^{+}17]. Anne [Ben11]. Anne-Wil [Ben11]. anniversary [BS13a]. annual [GRBBS17, HAA14, Rya16, SA16]. annually [BYY17]. Anomaly [HD17, CLLH15]. Antarctic [FH16, HLY14]. Antarctica [JPZ14, SRP13]. Antecedents [KH17]. anticancer [BMM17, Xie15]. anticipating [Ano16b]. anticipation [Ley15a]. antioxidants [AG13]. Antisemitism [TG18a]. Any [Fan13a, MR15]. anything [Mou15a]. Aparna [Pra16b]. apart [Kor19]. apartheid [Soo10b]. APC [PB18]. apparent [PZ17]. appear [CT15a]. appearance [BWbH^{+}18, DLM15]. appearing [HYYL12, HLY14]. applicability [TYYW16]. applicable [SA16]. applicant [YK14]. applicants [ACT18, HN16, PJY17]. Application [San12a, WTM^{+}16, CMM17, DMM17, Ell18, GRSFVdMA14, GTD14, GSMT10, GWA14, GC10, ILP13, JKJL14, JL18a, JL19, KKK^{+}14, Lam12, LZFW15, LGD11, LF12a, Mag14b, Rya16, San12d, SDS14b, VEJC^{+}18a, WYvE11, YK14, YLL10, ZAJ19, ZK19, ZKC^{+}16, ZWW^{+}16]. applications [AAS^{+}19, CFP14, CC16, KM17b, Lor14, MT12b, NT17, SPS14, TTC17, YK15, ZRY^{+}12]. applied [uHBLKH19, BFMRM19, HRH10, HSXL14, KHK13, LGL10, MDG10, SM16a, VB12, dACdFC18, ddMS15, Pen19]. apply [EC16]. Applying [BG17, KLL14, Saa10, ZC14, Mad15, NF13, TE18]. appointing [BN14]. appointment [Hos11]. appointments [GG15b]. appraisal [GY12, VPM16]. approach [AF15a, AF15b, APR19, AS18a, Ama18b, AOd15, Ano17c, AMMT16, AJCACRdMA16, BCML19, BVOL18, BASL16, BT18b, BT18c, BGM17, BM14a, BCC^{+}17, BDC^{+}12, CMT19, CLL^{+}17, CSC14, CFdC^{+}14, CMVP16, DLL^{+}16b, DLL^{+}16a, DFS15, DNAH15, DQ11, DB19, DGDGSV15, DPF^{+}16, DCY^{+}17, FMP17a, FVVSGM^{+}18, FGMM12, GRTPMLAJ19, GGS17, GSMT10, GGP14, GG15b, HT11, Har13a, HSAK18, HB18b, Ho16, HR15, HWQ^{+}18, JBMR11, JYM^{+}16, JC19, JDLIV14, KHS^{+}15, KMP^{+}11b, KCP12, KLPP16, KD14, KPRT16, KTRP17, LLC^{+}17, Lam12, LGL10, LMKG19, LPC17, LGPC18, LYGQ12, Li18, LLGW13, LdZwC^{+}17, LL13b, LS19b, MXZ18, MB19, MRGT13, MHTB17, MH14, Moh12, MGGdP17, MRN14, MAGBBM13, NJ10, Nic14, dBONM^{+}19, OA10a, Par14b, Pra10c, Pra10d, RAM18, RPAMR19, RD13, SH19, SP12a, SMAABJ11, SMCC18, SLK12, tScL13, TSG13, Tor14, VuHL10]. approach [VG14, WNS13, WWP14, WDP11, WA18a, XCS^{+}16, YLL15a, YHL19, YSY^{+}13, YKCK13, ZK19, ZG17e, ZLL^{+}17, ZHZY19, dSTL18]. approaches [ANOdFC12, CA19, CLB13, KKK^{+}14, MB16a, Moe10, RNF19, SN10, VBG^{+}17, WvE13, WC18]. approaching [OI17]. appropriability [SZ18]. appropriate [ADC12, MCR^{+}12, Moe10]. appropriation [GAPP18]. approximate [TW10]. aquaculture [DMM13, NFH12]. aquatic [LH14, LAdAMJ17]. Arab [Ibr18, SZAJS14, TBT19]. Arabia [SZMS17, SZMS18, SFM16, SLK12]. arbitrage [YKCK13]. archive [ZW18a]. archived [BC17]. Archives [CRLMLM17, MKYM^{+}17, The18a]. archiving [Laa14, LP18a]. Archivos [CRLMLM17]. Arctic [HYYL12]. ARDL [ILP13]. Area [CDCK13, BLdlCV17, Egg13c, GdOdAG^{+}13, HTHB11, KHVGA^{+}16, Kha13a, Kha13b, LOMLPA^{+}17, MCvFP16, MGMW14, OA10b, OA10c, Pin15, RGdCMM17, San12c, SP12b, SÁV18, UBTS16, PQG14]. areal [Sch12a]. areas [FK17, GGS14, GWB11, HHK^{+}12, HHZ14, LCC12, LSR13, MNdF16, MGC19, Pet18a, RCdJ^{+}14, SH15b, FK18]. arena [WLY14]. Argentina [CRFM^{+}12, MC13, MMAHS10]. Argentinean [CRMdMA15]. Argentinian [GZM15]. arise [GRSFV14a]. arisen [Har16d]. arithmetically [Hag10b]. Armenia [GHA^{+}15]. art [Ano16b, DG16, ENST16, LWB16, LMR16, PFDL17, RMdO17, UMdSV12, Wad16, Wad17]. artefact [Van12]. Article [BS15b, JN11, PB18, ABMRVZ14, Avk13, BS15a, BBSS16c, CGPT15, CP16, DB16, EGR13, dCPF14, FESD11, GBDG19, GdA14, KCT^{+}17, LLH^{+}16, LDZ17, MT13a, MSC18, NG16, QA19, SP14, The19a, WLMF15a, WLMF15b, XGL^{+}19, LX19, Zit12, dW15, AZSA14]. article-level [LLH^{+}16]. Articles [Bha11, AYS14, Ase10, BIH18, Cab14, CC11a, CSS^{+}16, CVD14, Cav15b, Cha14, Cha18c, Cha19b, CHY13, Che18a, Che20, CHL10, CH12, DBO^{+}18, DGD19, ESB15, Eld19, Emm19, dCPF14, FH16, FI16, GRSFV12a, GdA14, HII^{+}18, HBS^{+}19, Ho14, HV18b, HYYL12, HLY14, HW12, HP10, IWK18, JN15, Jam17, JNA18, JDG14, KPJ16, KY17, Kos18b, LL16, LY16a, LS16, LHM^{+}11, LTGH15, LHTL18, LNMQRR15, dNMVQL16, MS18a, MBR^{+}13, MASM14, MASM16, MZ14, NJM18, ND16, PNS^{+}10, PFL19, PB18, PTMT11, Ric17, RPK16, RPK18, RMH14, SGM^{+}16, SCGZR16, SM17, SML16, TA11, The18c, Tom17, Tom18, VVN16, VH17, WLF15, WFG16, WFS16, WK17, WLC17, WBX18, WJCC19, WWC19, YHL19, YJ11, ZW18b, dAG13, dW15]. articles-how [GdA14]. artifacts [NSC13]. Artificial [CGG^{+}17, AAS^{+}19, CC10a, ILB11, LVSL18, OMCP17, ZWZ^{+}19, ZWX22]. artificially [Ama15]. arts [LXH^{+}18, HH15c]. ARWU [Saf19, Doc12, JJS^{+}12]. arXiv [BI14]. ASAC [CST11]. ascertaining [BMP^{+}14]. ascription [Pei19]. ASEAN [HHK^{+}12, KRR14, NP11, PS15, RS12, Vin12b]. ashes [Har16d]. Asia [Chu14, HTHB11, KP12a, KTT11, SK14c, WCK^{+}12, WCB^{+}15, WHW^{+}19, WLZ^{+}15]. Asian [Bar17a, CDCK13, HIC12, KHJ^{+}12, KHK13, Moe16b, PP18, WG12]. Aspect [IA19, GHvdB12]. aspects [Bal12, BN10, Ell18, ZG17d]. Asperger [LLP^{+}16]. aspirant [WF17]. assassins [CPRSFVG19]. assertion [BSK15]. assess [ADS12, ADC12, ACD15, ANOdFC12, Doc11a, Doc11b, HRH10, MBA13, RGGBV16, Sch16]. assessed [Fed13, FCWH11, LLG14]. Assessing [ADS10a, BM11, CGG^{+}17, CYT^{+}12, DFG^{+}18, Gál17, GK18, GPN10, GG13, KLP12, KHVGA^{+}16, LTM12, Lee10a, LSM^{+}15, MARMSG19, MGLZ10, MHM^{+}13, McC14, MT13a, MSH16, PY14, PLBZ18, Sch11a, mYqS15, BAB13, BLS15, BSMD11, FMP17a, Ham14, HBA19, ILB13, LNRSRBB18, NCG^{+}19, Tol12, YWL16]. Assessment [BBDS^{+}14, BP11, CLD13, LSC10, ADS10b, ADV10, ADD11b, ADD11a, AD11b, ACD13, ADD14a, AD16, ACC^{+}16, BD16a, BYR13, BKL15, BR11, BGJ^{+}16, BS16, CTL^{+}19, CFSSP16, CG18a, DJWS11, FAI^{+}18, FE14, FE16a, FE16b, GTD14, GZ18, GRG12, GG15b, GMM16, GAGT15, Hal14, HYYR14, ILP11, LPC17, Lyk18, MWH14, McC18, MYN^{+}15, MCR^{+}12, MTA^{+}18, PS15, Pin15, Pou10, Pra11a, Pra18e, RNB19, Sal17, SZMS18, SH15b, SBT18, SBSR19, Siv16b, SDP^{+}19, TAB13, VG11b, WXZ^{+}16, WAT16, ZK19, ZZL^{+}10, ZC16, dCPRP18, DPF^{+}16]. Assessments [TR14, Vin17, vLCCMV13]. assets [NSMMDB19]. assignee [CF14, SWCH14]. assignees [HCS^{+}15, LZZ^{+}13]. Assigning [CFP14]. assignment [Don17]. assisted [VAJCC17]. associated [CHC17, Hen18, KBAK17, MRGT13, SFR^{+}19, TR14, XGL^{+}19, Zhu17]. Association [GWBSVWB13, ZGL^{+}17, CGKB18, CdMCdMMdP17, LYS^{+}17, LVHS^{+}15, YWS18, ZYNZ18]. associations [CG17]. assortative [PR10]. assumptions [Lin18, OCCSM11]. Astrobiology [GG13, Coc18, TA15]. astronomers [BGÖ^{+}13]. astronomical [Tri10]. Astronomy [Wil15, CH13a, CH14, GT17, RLW14]. Astrophysics [BI14, MASM14, MASM16, CH13a, CH14, HL15, TC11, TC13, VYL17]. asymmetry [CRBRG^{+}18]. Atapuerca [LRA14]. Atlas [Ley11a]. Atmospheric [BND11, LWR^{+}17, BHJD12]. attached [Han15]. attachment [SL16]. attain [Yur18a]. attempt [BH16c]. Attention [MSC18, BH18a, Bru10, GRSFV18, HV18b, MR18b, WLM15, WLMF15b, XBD^{+}18, XDB^{+}19, WLMF15a]. attract [CFG^{+}14, ZLTY18]. attracting [ZS18]. attractive [AKB12]. attribute [BD13]. attributes [LZ14, NG16, TG17]. attributions [Mix18]. auctioned [LL10]. Aude [DF15]. August [LT10b]. Australia [GB14b, MWH14]. Australian [GB17b, HG10, LXH^{+}18, MR15, MS13, SM14, VG17, WOW13, YWG14]. Austria [GRG12]. Author [BG12, FD14, GHvdB12, KT15, Leb12, SKCK14, WD13, Yur18a, AT17, Abt17, Ama18a, AOd15, AA10, Asu19, Aus13, AP14, Ber18, BKRG13, ByLbH16, BWbH^{+}18, Bur14, CL11, CIK^{+}18, DAMC15, DRS18, DdS19b, DWGL16, Dor17, DGDGSV15, EGR13, Emm19, Gál17, GRSFV15, GRSFV16b, GRSFVCP19, GD16, GSOLHO19, HYF^{+}17, Har12, wH15, HS17, JX13, Kim18, KK18, Kim19a, Kim19b, KKOS19, LP18a, Li18, LH12, MY16, MM14c, MSL11, MJHG13, MPH19, MRR17, OZK11, Pei19, PR15, PPM^{+}17, Pra19b, QDY14, Rou12a, SGSS17, Sch12b, Sch16, SGY15, SMM15, SE18, TA14a, TDG17, TH19, TC13, Whi15, WSL14, Wil15, WLPH14, YWW17, ZLW19, ZS11, ZC14, ZW14, ZZLS19, ZGL^{+}17, Wra16a, Wra16b]. author- [DdS19b]. author-citation [ZZLS19]. author-editor [GRSFV15]. author-inventors [MM14c]. author-level [WSL14, Wil15]. author-paper-citation [Pra19b]. Author-Pay [Wra16a, Wra16b]. author-pays [SGY15, SE18]. Author-weighted [Yur18a]. authored [KO18, LF12a, Osó18]. authoring [KM15a]. authority [CYT^{+}12]. Authors [GRSFV16a, Har17, ADR13, Ask18, Aus13, uHBK19, BD12b, CBWJ18, Cha18c, Cha19b, CXWW18, FM17, FE14, FE16a, FE16b, GHT16, GBHT16, GW10b, Har16b, HC16a, HL15, HV18b, HJM^{+}13, HL17, Hud16, JMM19, LPC17, LXL15, Liu16, LXL16, MHM^{+}13, MS14, PEPUT15, QA18, QL12, SOBM16, SLGO17, San12a, San12d, San12c, TBW^{+}12, VO17, WW15, WQY12, WBX18]. Authorship [Fer14, RKZK18, Abb11, AF18, Aus14b, AP14, BHB13, BM12b, CMUdF15, CVC^{+}15, Cha19c, CRFM^{+}12, CHL10, CB11, CFK16, Fie15b, Fie15a, Fie15c, FDVZ16, Gal11, GLS16, GA18, GTGABAG15, Hag10b, Hen16, Hen18, HLW19, ICC16, JCK11, JX13, KY16, Kim10, KGG15, KPSL12, LH12, LF14a, LF14b, LX15, Med18, MCL^{+}11, MdNS^{+}19, NPT^{+}15, OVJM17, OROMAA16, OKK14, PFL19, PYL16, PROGMA10, PG14b, PB16, Pra11b, Pri15, Pri16b, RGGBV16, RP17b, RPK18, SP14, SL14, SSN19, SJ19, Tol11, UHAR12, VuHL10, VO17, Wal16, YB14, YSD11, dSF13]. authorships [BL15, CHM15, FK16, KB18, Med18]. automated [Gom19, LBW17, ST14c, DPF^{+}16]. Automatic [ABM19, KBT14, MXZ18, vEWNB10, ES16b, GMJ^{+}17, MSA13]. automatically [KKOS19, LSS15, LSS16, MÁB18, TL18]. automation [MAGAM13]. automatizing [Lam12]. automobile [WLN^{+}14]. automotive [CD16b]. autonomous [VT10]. autonomy [CRMPRS18]. autoregressive [ILP13]. Availability [BC17, ABMSSP16, ABSF^{+}19, CLD13, LL16, LP18a]. average [ADV11, Egg10c, GBDG19, YY16]. avoid [LMKG19]. Avoiding [The17b]. awakeners [Fan19]. awakens [EG18]. award [MTA^{+}18, Fie15b]. awardees [Han11]. awarding [HN16]. awards [CT15b, JL18b, MTU17, ZL15a]. aware [CU16]. away [CS19, SS16]. axiomatics [Que10]. Azerbaijan [GHA^{+}15].

B [PAL13]. Back [MMOMALC16, CG18b, SMF18]. background [CT15b, Ye14]. backward [CKPY19, CLHH10, YST12]. bad [HS16a]. badly [WP18]. Baidu [CD18]. Baike [CD18]. Bakare [Ano18b]. balance [Ley18, MHM^{+}13, Pra19a]. balanced [WLZ^{+}19]. Balassa [Rou19]. ball [Cha19a]. Baltic [ZKD11]. bang [MB10a]. Bangladesh [HMK^{+}12, MR10]. bankruptcy [dPdCAdMC^{+}16]. bans [CRBRG^{+}18]. bar [RCJ18, Ano17a, Ano19, The17c]. Bar-Ilan [Ano17a, Ano19, The17c]. Barsotti [Pra17a]. base [FRPP17, GLÚGML16, Ham11, HMK^{+}12, MKF14, RY14, SdJDD19, WOW10, ZG17c, ZLH^{+}15]. Based [CXZ19, DLL^{+}16b, AAH10, AF15a, AJSN18, ACD11, Ahr17, uARA19, AÇA^{+}14, ACHVH10, ÁCCG^{+}15, Ama18a, ADD^{+}15, AW11, AGLNRR14, AMMT16, AYS^{+}13, Ask18, AMI18, BK10, BKL15, Bel17, BGBS18, BT18b, BT18c, BSBG18, BM14b, BH16c, BYY17, Boy17b, BN14, BDC^{+}12, Cab11, CHWL12, CHL15, Cha18c, CZPR17, CCLL14, CC12b, CYH13, CjZZ^{+}19, CLSW19, CHL10, CL16, De 13, DMB17, DdS19b, Egg14b, ÉMS^{+}13, Eto13, EHK12, FZZ17, FSSPG^{+}15, FE16a, FE16b, FMM14, FA10, Gal11, pGSyW^{+}19, GTD14, HA19, HK19, HYF^{+}17, HBA19, HGH17a, HB18b, HLC17, HH10, HP18, HJM^{+}13, Hos11, HTL15, HSPY15, HZL^{+}17, HH17e, HYS18, IA19, JSZ13, JC19, JH16, JDLIV14, JS15, Jun12, KMD^{+}18, KM15a, KWM^{+}18, KT15, KPJ16, Kha13a, Kha13b, KHR^{+}19, KP12b, KD14, KL17, Kis11b, KB13, KB18, KJES16, KW17, KPRT16]. based [KTRP17, KK17, KR17a, LGR17, LVSL18, LSC10, LS17a, LPMK17, LL19, LZZ^{+}12, LZZ^{+}13, LCR13, LKR14, LZ14, LYWSV13, LYGQ12, LGZ^{+}13, LZFW15, LSS15, LSS16, LPL16, Li18, LP18b, LRY18, LSY11, LCD^{+}14, LGH^{+}14, LWM^{+}15, LM16, LdZwC^{+}17, LL13b, LDZ17, MLOY18, MVS10, MR13, MGLZ10, MB13, MHTB17, MCR^{+}12, MM15a, MD18, Moo15, MKHB13b, NH11, NA12, PFDL17, PYK12, PJY17, PHL17, PMJF19, PROG19, PRDG17, PZ17, PPM^{+}17, RAA18, RRL16, RT17, RFGBMA13, RPK16, RP17a, RPK17, SOBM16, SBSR19, SS10b, Sch10a, SZAS16, SK17, SLK12, SKCK14, Shu17, SV19, SPdSM16, Smi12, SM15, SH18, SDP^{+}19, TW10, TLSH14, TA18, TSMTDLCH11, Van14, VASNU^{+}19, VO17, VT10, WRM17, WYAY12, WLLL12, WZLZ13, WCL14, WLY14, WDS16, WK17, WLC17, WG11, Won13, WLH^{+}17, WZ19b, WKHS19, XGY^{+}16, Yan14, mYqS15, YLL15a, YQX10, YWC12]. based [YS13, YZW^{+}17, YHL19, YLL10, YCK11, YLH^{+}17, YL10, YYL10, YHL^{+}18, ZPG^{+}14, ZYZ14, ZZL^{+}10, ZJLG10, ZCL14, ZZPG14, ZCL15, ZGY16, ZLW16, ZFY^{+}17, ZG17d, ZG17e, ZQH^{+}17, Zha17, ZWW^{+}18, ZSC18, ZZW19a, ZLLL19, ZyZZ^{+}11, ZGL^{+}17, ZM16, ZZZC16, ZGJ18, ZHZY19, ZG13c, ZWHH13, dZLwC^{+}15, ZLF18, dS17a, dPSS18, vEW17, vLvWW16, vRvLV11, DLL^{+}16a, AF15b, KCP11, YK11, YPK13]. baseline [GT18]. basemap [SK12]. bases [Sch12a]. Basic [HSXL14, ÁBDFB19, BFMRM19, BBL17a, DDS^{+}19a, DWGL16, wH15, NJ10, RRLNAG15, WZ17, vL12, ZKC^{+}16]. basic-clinical [RRLNAG15]. basins [Vil10]. basis [BMP^{+}14, GRSFV12a, GRSFV^{+}12b, Kar12, MG12, RNB19, Sch15a, Wu18]. Basu [Pra16b]. batches [CT15a]. battery [HLLT14]. Bayesian [ILB11, RBC^{+}10]. Bayh [LM10, LM13a, TR14]. BCE [Cav15a]. be [ADD10, ACD13, BM14a, BG18, DdlPPL^{+}19, GGH^{+}10, HL15, Ley15a, Ley15b, LCFC14, Lin10, MS16a, MBA13, MDG10, Ort16, Pei19, PS16b, SLISC17, SSZL18, Sot12, The18a, dSBC17]. beamplots [HBA19]. beauties [CL16, DRCG17, DW18, Fan18, HH17b, HH17d, HY19, LY12, Li14, LY16b, ON12, ZXZ17, ZLF18, vR17, TVA17, vRW18]. beauty [Fan19, WRM17, EG18, GR16, SSZL18]. because [Cop19a]. become [Bai18, GW10a, LM11]. becoming [Cav15a, HAL11, Kra16]. beer [Lor10]. before [IJF16, MSYW12, WHW^{+}19]. begets [ALvH19]. beginning [Bar11, MCL^{+}11]. behave [BDF^{+}17, YWL16]. behaving [WP18]. behavior [ADD19a, ADD19b, BKRG13, CSS^{+}16, DY18, GTGABAG15, HFC11, KY17, Kra10, LAS14, LZR14, RMA12, San12a, San12d, Sch10b, Won19, WHH^{+}18]. behavioral [BCT19, CdMCdMMdP17]. behaviors [KC12]. behaviour [LP18a, MS14, VNA16, WG10]. behaviours [GBMB10, GGS14]. Behind [Zit11, Emm19, GFK^{+}18, JKMS17, Oli15a, ÖS17, Pan14, RY14, RMA12, ZLL19]. Beijing [KZSZ19]. being [GRSFV19a, Har19a, Tor13, WLM15]. Belgium [CAV^{+}19]. beliefs [OBG11]. believe [BBV10]. belong [Koz15]. belonging [MRGT13]. Benchmarking [CRZGVQdMA16, GRSFV^{+}13, MvdH13, OCJB15, CFL12, GRSFV^{+}14c, GT18, RS12, SSN19]. benchmarks [BG18, ZZL^{+}10]. beneficial [QZZD18]. Benefit [ZZW14, BBJS16, PPI17, WHLP16, Wra16a, ZZW16]. benefits [SMM^{+}19]. benefitting [Lor14]. benevolent [DH13b]. Benford [AYS14, AYS16, CC11a, cTnHwH17]. Benin [Mêg13b, Mêg13c]. benthic [dCdSNB15]. Berlin [GGG^{+}11]. Bertocchi [BD16b]. Bertoli [Pra17a]. Bertoli-Barsotti [Pra17a]. Best [GRSFV^{+}14c, Cav15b, CFG^{+}14, DJWS11, FLM16, GRSFV19a, MT15, SR16]. Best-in-class [GRSFV^{+}14c]. better [ADR19, AC13, Fan13a, FFL16, GRSFV16b, GZGAC16, GZGAC17, HC16a, HK12, IA19, Lyk18, MS18a, MDFGAM14, SK16, Sch10b, WWC19]. Betteridge [CP16]. between [ADS11, AZKR13, ABMSSP16, AMFLH15, AND19, BCML19, BB10, BL10, Bak17, BD12b, BB19, BSS15, BBL17b, BS15b, BL15, BZBLP16, BL17b, CKPY19, Cha13, Cha18b, CC10b, CS11a, CSC14, Che18a, Che20, CIK^{+}18, Cle16, CdMCdMMdP17, DFG^{+}18, DC15b, Dya14, EMSH16, ERW12, Egg10d, Egg13b, ER19a, EMH^{+}10, Fin11, Fin15, FM11c, Fuk14, FI16, pGDTP12, GRSFV18, GD16, GZ11, GB14b, GdOdAG^{+}13, GP18b, GBMA14, GMSZ18, HJL18, Har14b, HN16, wHwH11, HC12, HCLC14, ILB11, ILBG14, JG12, JG14, qJnShPL17, JK10b, JKSK15, KZSZ19, KLP17, KA13, Kim10, KHJ^{+}12, KTLD16, KG10b, KJ14, KM12, LAL15, LBGBdMA13, LV11, LL10, Lee10b, LH12, LHW16, LJC^{+}15, LLP^{+}16, LXH^{+}18, ML18, MVS10, MM16, MR15, MRGT18, MSL11, MA19, MGMY^{+}18, MDG10, MPM18, MSC18, NA18, APFR^{+}13, OMA15, dFPYdCL12]. between [PCR18, PJY17, Pen19, PRRC15, PROG19, Pra17a, RGLE16, RGCM14, RSGFV14, RPK16, RP17b, RPK17, RPK18, San12c, Sch13b, SH15c, SZ18, TB19a, TLSH14, TA11, The16, The18c, TYWZ12, Tod11, VTY17, WV13, WvEvL^{+}11a, WWP14, WWH^{+}17, WWP17, Wol15, XXL^{+}17, XGCK19, YSM^{+}19, YWS18, YYDH12, YWW17, YST12, ZYG15, ZFY^{+}17, ZSCR^{+}18, ZYNZ18, ZG12b, ZG13b, ZW19, ZRY^{+}12, ZL15b, ZB15, ZH17, dJC15, dW15, dSTL18]. between-department [PRRC15]. Betweenness [LWB18a, GLM11]. Beyond [ATM16, McC18, SK12, CRMdMA15, HH15b, KB11b, KR17a, Mik17, OKCPS17, Pra10b, Ran09, RPAMR19]. bi [BB19, GZ11]. bi-directional [GZ11]. bi-regional [BB19]. Bias [CFS18, ADR16a, CGSS13, CdMCdMMdP17, DD18, GRSFV16a, HE16, KKdBK12, KK13, LRZ13, MRLW15, Mou15b, Mue18, SC13, TN19, WL18]. biases [CdMCdMMdP17, MBT16, Mes11]. Bibliographic [TZG15, ATJ16, BC13a, DRS18, DCS12, FS12, GD16, GKF17, Gus19, HA19, HD17, HWL11, HH10, HC14b, hHC15, KKLP17, KWS17, Liu17, LJMF15, MB19, MGLZ10, dCCMAW16a, dCCMAW16b, OMAT19, SKCK14, SL12a, Soó14b, dFVDU^{+}19, VO17, YSY^{+}13, dCPRP18]. Bibliographical [LKP11, CHL15]. bibliographically [Sch12a]. Bibliometric [APPS15, AMFLH15, Ano18a, BY13, BSS15, BGJ^{+}16, CFM18, CVD14, CIK^{+}18, CdMCdMMdP17, CZ18, DSG^{+}15, DCM16, EDEH16, FP18, FM11a, FZZ^{+}11, GALR16, HL15, HZD^{+}15, Kaz14, LLL12, LJKG15, LHC16, LLP^{+}16, LA19, LWW^{+}11, LPZ17, MM14a, MM14b, MC12, MKP16, NBR^{+}11, OMJLVSN19, RNF19, SH19, SÁV18, SZAJS14, TW10, Ter17, yTnL17, TE18, VFA10, vLvWW16, ADS12, APT13, AG13, AAB^{+}13, AEFP16, APR19, ATCCAAB19, AATBPAB15b, ABGS14, AYS16, Ama18b, AC13, ACC^{+}16, AMMT16, Asu19, BMM14, Bar17b, BK15, BL11a, Bel13, BLS15, BMTA15, Bor16, BG17, BTL19, BZBLP16, CHWL12, CZW13, CLHH10, hCyL12, CWJC14, CJY^{+}15, CRZGVQMA15, CH12, CC11b, CG15b, CvLvR11, CNC18, CAS16, DMB17, DVB15, DSM11, DSH^{+}10, DXL^{+}12, EBR16, EBK16, EW15, Ell18, FZQ17, Fed13, FS12]. bibliometric [FPS14, FFR16, FdSdO17, FRF^{+}19, Fra10, FM11c, FMM15a, FMM15b, Fra17, FA10, FZZ^{+}12b, FH13, FLH14, FH16, GW15a, GG14, GREL14, Gar15, GMJ^{+}17, GW10a, Glä12, GM13, GTC16, GZM15, GGW11, GG15b, GGG16a, GFK^{+}18, GBGB13, GGW^{+}13, GWG17, HH17a, HHK^{+}12, HHZ14, HTHB11, HPBI^{+}14, Hei13, Ho13a, Ho14, HH15c, Ho16, HH17c, HSBW10, HdSV16, HC15b, HC15c, HC17, HC16b, HFC11, HW12, HSX^{+}15, HHA^{+}16, HZL^{+}17, ILB11, IQT^{+}19, IH14, IFH15, JPZ14, JDG14, JPZ^{+}10, Kaz15, KHVGA^{+}16, Kim14, KKBW17, Kis11a, Kis11b, KGZML^{+}19, Kor19, KVC15, KBL15, Kra10, KvES11, LBCO19, LKP11, LNRSRBB18, LWR^{+}17, Lin12, LZH^{+}12, LLGW13, LZH^{+}13, LFLG14, LTGH15, LYLD15, LLW^{+}16, LM16, LTK^{+}19, LdSdFFNM17, LSL15, Lyk18, MH16a, ML18, MdBdP^{+}19, MWH14, MB13, MHTB17, MDDG17, MSL11, MYN^{+}15, MGMW14, MS14, MH14]. bibliometric [MSH16, MD18, MdNS^{+}19, MEG15, MRN14, NFH12, NHY^{+}14, NT17, Ort17, PFDL17, PJL19, PYL16, PHBN^{+}15, PS15, PLW^{+}15, PEFP13, PFPCM^{+}19, Pra11a, Pra14a, Pra14b, Pra14c, Pra17c, Pra18a, PHS12, RG15, RC13a, RMdO17, RAM18, RGCM14, RPNC13, Saa10, SC18, SRGMF15, SZMS17, SH15b, SMCC18, Sch16, SFR^{+}19, SAR19, SSN19, Sku19, SM15, SS15, SCLC15, SG16, Suo14, TG16, TG18a, TFH14, TCC17, TCH^{+}15, TA15, TFJD14, TM12, TA11, TW16, Tod11, TP11, TA17, Tsa11, Tsa15, yT11, yT15, TT13, VMM15, VZAMG19, VEJC^{+}18a, VSVR15, VSK18, WvEvL^{+}11a, WYvE11, WYH10, WLHZ13, WL14, WCB^{+}15, WXZ^{+}16, WH16, WC18, WHW^{+}19, WL18, WSL14, Wil15, WOW10, WOW13, XM13, YWL16, YLC18, YYDH12, YCL^{+}13a, YJ11, ZL18a, ZXH10, ZJLG10, ZLT^{+}14, ZLL^{+}15, ZXM^{+}16, ZWW^{+}18, ZY19, ZWW^{+}15, ZWW^{+}16, ZZY13, ZXLX14]. bibliometric [ZZ15, ZG13c, ZHMX14, ZH17, ZLN^{+}13, dART^{+}17, dJC15, dMALIM14, dPdCAdMC^{+}16, ddMS15, vEWNB10, vEW10, vLCCMV13, vR12, vWBS^{+}16, BD16a, Glä18, WWC19]. Bibliometric-enhanced [Ano18a, CFM18, SH19]. bibliometricly [LD16]. Bibliometrics [Ano15, DGWZ13, Glä15, MMOMLC18b, PB12, AD11a, ADD11a, ADR19, Agu12, ABMRVZ14, ABRVZ15, BH17b, BM19, BJIB16, CLD13, CYW^{+}11, CS19, DPF^{+}16, GWP16, JVM17, KHS^{+}15, KDFL14, LWB16, Lin10, LSE^{+}18, Mad15, MB16a, MS15a, NRAW17, OCJB15, QDY14, RTP17, TE18, UMdSV12, Vin10a, WV13, WTM^{+}16, Whi18, mYqS15, ZAJ19, Zit15]. Bibliometrics-aided [Glä15]. bibliometry [GY12]. bidimensional [TSMTDLCH11]. Bidirectional [GP18b]. Big [Hal14, HH15b, HSPY15, Pra14a, Ano16b, HIG^{+}17, HBS^{+}19, LPB14, MB10a, NT17, Par14a, RNM18, Sko14, YWG14, ZVC11, HZ17, JK19, ZLH^{+}15, SBSU15]. bilateral [BKSS15, GW15b]. billion [Jac18]. Binary [Aus14a, Par15]. Bio [JAAA18, LM13b]. bio-scientists [LM13b]. Biochemistry [CAGL15, TN19]. biodiversity [CFdC^{+}14, aSS17]. bioenergy [LGH^{+}14]. biofuels [GD17]. Biographical [IWK18, SBB16]. biographies [Ken18]. bioinformatics [KJS14, LLG14, SK13, SHK14]. Biological [Kim14, ACAGD^{+}17, NPT^{+}15, RCdJ^{+}14]. Biology [CAGL15, RBF^{+}10, CWJBT10, LABL13, RASP13, SKY17, TN19, ZW18b, ZL17]. biomass [LGH^{+}14, ZKC^{+}16]. biomass-based [LGH^{+}14]. Biomédica [STCRPA18]. Biomedical [CD18, KKT^{+}18, ABMRVZ14, ABRVZ15, ÁBDFB19, AGLNRR14, BAC13, BvdB14, BN10, Bre10, BN14, CXWW18, CL17b, FA10, FAA13, HEH17, JK10a, JDH12, JH16, Kis11b, Lan13, LT10b, LS15, MV19, OCJB15, TTC17, VHG^{+}15, ZLG^{+}15]. bionanoscience [RM10]. biopharmaceutical [AJCACRdMA16, GSPLVG^{+}18, ZT18, ZT19]. biophysics [Var11]. biosciences [JTZ14]. BIOTA [CFdC^{+}14]. biotechnological [RR17]. Biotechnology [MJC14, BY13, CG11, CdSPdM13, EES13, GZ11, HCLC14, PS15, PS13, ZLH^{+}16]. biped [LW10]. bird [DF15]. birth [GR16]. black [GP15, Sku19]. Blaise [Ano14]. blanket [ZLT^{+}14]. blind [SB17]. block [Har19b]. blockbuster [BMR12]. blockchain [FRF^{+}19]. blockmodel [XLZ^{+}18]. Blockmodeling [CRFM^{+}12]. blogs [HC16a]. BMC [Asa19]. BMJ [DGD19]. board [Kim10, MHM^{+}13, TO18, YP19, ZL18b]. boards [GCGP10]. bodies [BM12b]. Bogazici [SSAG16]. Bonded [KM16]. Bonded-communities [KM16]. Bonitz [BS13a]. bonobo [Hir19b, LBO19]. Book [HCDT16, HFC11, LDG17, Ye14, Ben11, GTC16, GGP14, HH17a, JMM19, MRGT18, TSRGCCJC14, TSRGG17, VE14, WF18, ZZZC16]. bookmarks [SMM15]. Bookmetrix [ENA19]. books [ENA19, GTMRE^{+}16, GTMRE^{+}19, OE15, Sch14a, Sch17b, TSRGCCJC14, yTmShL16]. booming [AvLS14]. boost [ES16a]. boosted [WBH^{+}12]. boosted-trees [WBH^{+}12]. Boosting [GKS18, LS19b]. Border [PSY^{+}19, CKB^{+}14]. borders [HL13].

Börner [Ley11a]. Bornmann [Hir19b, Dan19, Ste17]. both [HM18]. boundaries [BMR12, Fie15b, MGMY^{+}18]. boundary [Cop19a]. bounded [McC14]. bounds [KD19]. box [Sku19]. BR [ZKC^{+}16]. BR-CI [ZKC^{+}16]. brain [MT12a, Vel12, WMW^{+}13]. branch [PL18]. branches [PPE14]. brasiliense [GVGSEPRC15]. Brazil [dCPF14, CdSPdM13, FWFM18, FKRS14, GdOdAG^{+}13, GdA14, Han15, HRH10, HdSV16, RGdCMM17, RCdJ^{+}14, dSSdMAF14, SSdOS17, dAG13, dSF13]. Brazilian [BSG17, CFdC^{+}14, CDdS^{+}12, DRMMC19, FP18, HR11, HG17, LML11, LSM^{+}15, dNMVQL16, MCL^{+}11, MdNS^{+}19, OCM^{+}12, PB18, PPK^{+}16, RCdJ^{+}14, SLD^{+}17, SRP13, WV13, dSNV18, dART^{+}17]. breadth [GPN14, HR15, YMSQ10]. Breakthrough [PLWS14, FM11c, GKS18, Paj15, WFH^{+}16]. breakthroughs [RNB19, WT14, WT15]. BRIC [AATBPAB15a, WLY14, YYDH12]. brick [GRSS16]. BRICS [BPHL16, Fin15, FB16, SK18, WW12, YQW13]. Bridge [WMT^{+}12, RCJ18, XG18]. bridges [ACD14]. bridging [AChO19, XG18]. brief [Bha18, LT10b, MBR^{+}13]. brimstone [WRM17]. broad [LLYC14]. brokerage [PW13, WhCL10]. brokers [Fie15a]. browsing [KWS17]. BT [KL17]. bubble [FM11c, ZWW^{+}16]. budget [Tod14]. Building [HSK18, WJCC19, ÁRS17, KR17a, McC18, SK16, xShLY^{+}15, VHD^{+}16]. buildings [CNC18]. builds [XG18]. burden [HEH17]. burdensome [dS18]. Business [ALvH19, CDD15, ELP11, HM15b, KG16, SRGMF15, WS13b, ZWW^{+}18, AL12, AHP17, BSS15, Bha16, CLLZ15, CIK^{+}18, Cle16, CAS16, DTM^{+}13, EBD15, FPS14, HLSC18, KJ13, LQW17, LHM^{+}11, LDVSGDR16, LDVSGD19, ML10, PHV17, PHS12, RBBG18, TE18, WHC^{+}13, WLZ^{+}15, HM15a, HSBW10]. butterfly [WRM17]. BWM [Sal17, SDP^{+}19]. byline [ADR13, MSL11, YWW17].

C [EDEH16]. CAD [AA18]. Calculating [BWdMA17, Sch13a]. calculation [CvLB10, FMM13a, MBSB17, Pra12d, RKT^{+}15]. calculus [MGT14]. calibrated [KPRT16]. California [HFW^{+}14]. call [Sug11]. calls [RJ14]. Calophyllum [GVGSEPRC15]. came [CMO11]. camel [GAGT15]. campaigns [Par14b]. Campanario [Egg14a, Ano18b]. Campbell [Par14c]. campuses [PL18]. Can [Bru10, HL15, Kis11a, LRC19, Ley15a, Ley15b, LCFC14, MS16a, MBA13, MBT16, MDG10, Pei19, SL17, SS10c, SLH18, The18a, WFZD19a, ACD13, ALH15, BG18, ENA19, GRSFV14a, Ley18, LM15, LF17, MS18a, PS16b, PLWS14, SAPR18, SLISC17, WhCL10, WM17, WWC19, Yur18a, HC16a]. Canada [EES13, HAL11, Lun19, TBS15]. Canadian [GHS18]. cancer [BL17b, KGG15, LR12b, LRWS16, WKHS19]. candidate [ST14c]. canonical [LW15]. Can’t [Lyk18]. Cao [PT17]. capabilities [FLZ17, FFR^{+}17, LGR17, Wu14]. capability [LKS^{+}14, Liu17, vWBS^{+}16]. capacity [KS17, MJC12, VHD^{+}16]. capital [BDC^{+}12, DR10a, EN17, FMU16, GB14a, KMS16, MARMSG19, PS13, TO18, ZW17a]. capital-based [BDC^{+}12]. capitals [SWH14a, SWH14b]. capture [CGG19]. Capturing [APYS13, TB19b, TW18, CMRC15]. Carayannis [Par14c]. carbon [CWL10, EGUB12]. card [BI12b, BI12a]. cardiology [Sch17a]. Cardiovascular [BPTG10, SZMS17, CdJD15]. care [CYW^{+}11, YLY^{+}14]. Career [ADR14a, DH13a, CT15b, CMT18, Egg13a, GHA^{+}16, HAA14, HCL14, KKS16, Lee19a, LD16, VO17, XA15, ZG12a]. Careers [Kos15, Waa13, CHM15, Emm19, KK13, QZL^{+}17, Yur16a, vdBS16]. Caribbean [BB19, CR14, MSB18]. Carlo [Sch16]. cars [Jun12]. cascades [HBDL18]. Case [GS15, ADV11, ACD13, AW10, Ama18a, Ard12, Asa19, ÁRM13, BCML19, BHB13, BI10b, BI14, BIH17, BSG17, BPGGdMA12, BR11, BSK15, BKSS15, BD10b, BSMD11, BND11, Bor18, BPVM11, BvdB14, BJIB16, CWJBT10, CSS^{+}16, Car16, CGG^{+}17, CFP14, CO10, CWL10, CJC13, CAGL15, CLO18, CLLH15, CD18, Chi14, CRFM^{+}12, CHL10, CP12b, CC13, CYK^{+}11, CdSPdM13, DRS18, DC15a, DXL^{+}18, Dya17a, Dya17b, EMSS16, ET15, EES13, Fin11, FI16, pGSyW^{+}19, GRSFV^{+}12b, GD11, GL15, GSTD11, GT18, GRG12, GG13, GZ14a, GLM11, HH15a, Hen18, HMCL16, HL18, HNG19, HTL15, HCLC14, HZQ^{+}17, HYS18, HW10, ILB11, IBL13, IF13, ILP11, JNA18, JKPL18, Jun12, KLM16, KKCG18, KM15b, KM15c, KM15d, KL17, Kim18, Kim19a, Kli16, KK19, KWW15, Kra19, KTT11, LR12a, LLL12, LSCK12, LYS^{+}17, LT10a]. case [LNRSRBB18, LW15, LM15, cSL10, LABL13, LL13b, LSE^{+}18, ML16, Mag14a, MRGT13, MMOMLC18b, MB13, MHFB17, MBA13, Mêg13a, MK18, MYN^{+}15, Mes11, MNdF16, Moe16a, MZE19, MCB15, NSKO15, NJ10, NBR^{+}11, OMOR13b, OMLC15, OMA15, Ort11, OBG11, PY14, PHL17, PLJ18, PSB^{+}17, PRRC15, Pet18a, Pol16b, PPI17, PB17b, PLG19, RHMH17, RM10, RASP13, RMA12, RGTSLCH14, RPP18, RF19, RNM18, SCGZSL^{+}13, SND19, SIS17, SW19a, SHR^{+}10, SG10, Sch13a, SS10c, SJ10, Shi11, SV19, SK11, Soó14b, Sot12, SK14a, SGY15, SE18, SL16, THAL15, TUCR15, The18a, TG18b, Tsa15, VNA16, VRF12, VLV14, WS11, WLR^{+}14, WLN^{+}14, WLD^{+}14, WPW^{+}14, WT14, WY12, XBD^{+}18, YIK^{+}10, YCL^{+}13b, YSY^{+}13, YKCK13, YKLK14, YCK11, Yos13, ZVC11, ZZPG14, ZSC18, ZW18b, Zha10, ZC14, ZLLL19, ZGL^{+}17]. case [LX19, ZLLD19, dSD18a, vLvWW16, vdBS16]. case-control [RF19]. cases [FR11, WD13]. CATAR [TT13]. catch [JK19, WRV14]. Catching [ÖS17, RY14]. categories [BBJS16, CC14, Cam17, Cam18, DGDG13, DGGBDG17, Eld19, GRSFV^{+}14c, Har13c, KB18, LCR13, QRJ^{+}17, RSGFV14, WWH^{+}17, ZCL14]. categorisation [JDH12]. categorization [AZKR13, ANZ15, TA18]. categorize [HK19]. category [CWH11, hHSL19, THFBdMA18, Yu17]. causal [BKRG13]. causality [ILBG14, LLCL11]. caused [Bor15b]. causes [MB15]. cautionary [BI18a, CF18, DSH^{+}10, dSD18b, dSD18c]. cautions [TCH^{+}15]. CBkCI [Ye14]. CC [AAH10]. CC-index [AAH10]. CE [Cav15a]. CE/ [Cav15a]. CEFIPRA [BKSS15]. cell [ÁRS17, BL17b, LZZ^{+}13, dNMVQL16, SYLC17, Suo14, YKCK13, ZZP^{+}14a, ZS11, BHDI18, SRW^{+}15]. cells [AW11, CHY16, CYK^{+}11, HLL14, HDC13, HYC15, LAHH15, Pra14c, WLR^{+}14, YCK11]. centage [Sch15c]. Center [Das16, Fie15c, You14, ZTRH18]. centers [EGUB12, KZSZ19]. Central [BBSS16b, BBSS16c, DGDG11, GBHT16, Egg13c, Pra16b, BBSS16a, GVS17, JMM19, LGR17, WCB^{+}15, WHW^{+}19]. centrality [Abb13, BHB13, BHA15, DLGP16, FLB11, FK17, FK18, GLM11, KJES16, RPK17, SP12b, ZLF^{+}14, ZRL18]. centre [SS10c]. centre-periphery [SS10c]. Centres [MEG15, MDFGAM14, MCB15]. century [ACORC10, BCC^{+}17, GZ18, MCL^{+}11, SCGZR16, Sot12, TW16, Won19]. CERIF [ISR11]. certain [ZG17a]. Certainty [AD13, SBK19]. certification [Van12]. Cext [ZLF^{+}14]. Cext-N [ZLF^{+}14]. CH [KWW15]. chain [BLdlCV17, NSH^{+}11, QDK19, SDT15]. chair [MB16b]. challenge [BGG^{+}17, MSA13, SÁV18]. Challenges [AJdMA10, Ano16b, BH16a, CRAJdMACÁ15, DG16, FS12, GTC16, HAG^{+}16, Hau16, ZLL^{+}15]. champion [MM14b]. chance [CYH13]. Change [CRLMLM17, MB16a, ARE^{+}18, AM18, ÁRM13, BP11, BHM16, CÖT15, CHY13, CL17a, CIL^{+}16, HYS18, HLLT14, JVM17, LWB18b, LDZ17, MB10a, MHFB17, MHTB17, PG14a, SK16, SY16a, Soo17]. Changes [Asa19, IFT^{+}18, AvLS14, Cam10, GKS16, Han11, HC12, ÖS17, RF19, SV19, Sob11, YS14]. Changing [Chi15, EOS12, ADD11b, HLSW18, MT13b, OFP16]. channel [KCT^{+}17]. character [ATJ16, Kol12]. Characterisation [GB17a, LFBI19, CPF13, RBF^{+}10]. characterises [ÁBDFB19]. Characteristic [Egg10a, BYY17, Glä10]. Characteristics [BBS17, FCWH11, Fuk17, LLG14, OE15, SLGO17, SM15, WTG15, Zha10, ZLH^{+}16, BIH18, BSvEK13, Fuk14, FK17, FK18, GFC18, GW15a, GKK15, GA18, GW17, HdSV16, HN16, JCK11, KSB11, LL10, LPMK17, LVHS^{+}15, Lin12, MM14b, MJHG13, MC12, MTA^{+}18, PY19, PLA10, RMH14, San13, SCL12, TSRGCCJC14, WZFD19, WSL14, YZB18, YC12, Yos13]. Characterization [GSPLVG^{+}18, PHDC16, AdAdAM10, DRCG17, ER19b, FP18, HMI19, VAJCC17]. characterizations [Que11]. characterize [ACHVH10, MK18, UBTS16, Vin10b, ZYF^{+}17]. Characterizing [JKN19, MHC^{+}15, PLA10, PAL13, YWY10, MHLGHV14, SLXD15]. charge [PB18]. charges [BS15b]. Charles [GR16, MHFB17]. Charting [Pen19]. Chatman [GTAG18]. Chaves [dCPF14]. chemical [FLH14, GVGSEPRC15, Kaz15, Tom17, Tom18]. Chemie [BD10b]. Chemistry [BND11, BHJD12, RHGKD16, Tom17, BHB13, BSMD11, CGG19, GW10a, KA13, Lee10b, LRZ13, TA17, Tom18]. chemists [Kos15, ZCKZ16]. Chengyu [PT17]. CHI [MTA^{+}18]. chicken [CMO11]. Chief [Glä14]. child [VASNU^{+}19]. Chile [MC13]. chimpanzee [Hir19b, LBO19]. China [Liu16, Shu17, BFHS18, BSB12, BL17b, CG11, CJC13, CXpHqZ15, CXWW18, CRR14, CHL10, DGWZ13, FCWH11, FH13, FLH14, GW15a, GGR11, GP18a, GW10b, HWL11, HHGZ11, HHDL13, HGH17b, HSXL14, HSX^{+}15, HYS18, qJnShPL17, KZSZ19, KHJ^{+}12, LZZ^{+}12, LZ14, LYGQ12, LGZ^{+}13, LL15, LCZ17, LSY11, LHW12, LLG14, LF14a, LXL15, LG15, LTGH15, LJC^{+}15, LXL16, LPZ17, Ma12, ML18, MM14b, MHM^{+}12a, MT13b, Pan14, PYL16, PP18, QDK19, QMSM^{+}19, SMY15, SH15c, SL16, TS11a, TS11b, TYWZ12, WXLL12, WXW^{+}13, Wan16, WWP17, WZ19a, WY12, Wu14, XCS^{+}16, YDZ10, mYqS15, YCL^{+}13b, YHC^{+}15, YGW^{+}15, YXW18, YSND17, YSD11, YYS^{+}10, YHL^{+}18, ZSY14, Zha14, ZGL14, ZZ14, ZZP^{+}14a, ZYG15, ZG17b, ZLL^{+}17, ZW17a, ZRL18, ZY19, ZW14, ZLH^{+}16, Zha18, ZW19, ZyZZ^{+}11, ZZZ^{+}12, ZM16, ZG10, ZZY13, ZT14, ZL15b, ZB15, ZSY^{+}13, dS17a]. China-a [YYS^{+}10]. China-related [LTGH15]. Chinese [ALYZ15, BHS14, CZ18, DQ11, Fan15a, FYC15, FL16, FZZ^{+}12a, GGG14, pGSyW^{+}19, GXC^{+}19, GC10, HYF^{+}17, HZD^{+}15, LYWSV13, LCWY12, LLX^{+}18, LZL10, MLC14, NQ14, SLG10, SW19a, SLH18, SDS14b, kWhHRkS10, WXLL12, WMW^{+}13, WHC^{+}13, WLN^{+}14, WRV14, WWP14, WPW^{+}14, WHL^{+}15, WDS16, XA15, YMSQ10, YGW^{+}15, Ye14, YSND17, ZYSS14, ZCZ^{+}16, ZZ11, ZHMX14, ZL17]. chip [VMM15]. choice [MRLW15, Mou15b, PROG19, Sil13, Wad18, Wan13]. choices [XA15]. cholinesterase [CWJC14]. Chornobyl [MHKB16]. chosen [FSLR10]. Christmas [Har13b, Har14a]. Chuang [HW12]. CI [ZKC^{+}16]. CIBER [MCB15]. circle [LM15]. circles [TA14b]. circular [CZ18, NT17, San12c]. circulation [Vel12]. circumstances [GKS18]. CIS [Kar12]. citability [ADD16]. citable [Cam17]. Citance [KMD^{+}18]. Citance-based [KMD^{+}18]. Citation [ANZ15, BI10b, BRS^{+}16, CC14, CjZZ^{+}19, CH12, CL17b, FLH14, FH16, HG10, Ho13b, Ho14, HH15c, HTL15, HW12, HZL^{+}17, hHSL19, HOB17, IH14, IFH15, Jac12, JPT13, KWW15, LBGBdMA13, LvI10, Li14, LLG14, Lop10, LL12, MH16a, MSB18, MCR^{+}12, NSH^{+}11, PD10, PRDG17, QRJ^{+}17, RMA12, SLG10, SD13, Sil13, SSS^{+}11, SL12b, Sot10, SDEB16, SDS14b, SC13, SZ12, TFH14, TYWZ12, VASNU^{+}19, Wan13, Ye14, YJ11, YYLW14, Zyc10, dSD18a, vEW17, AD13, ADS16, AD16, ACP12, AA18, uARA19, Ama16, ALH15, ACD14, Ase10, Ask18, BL17a, BVZV16, Bal12, BAB13, BK11, Bel17, Ben11, BASL16, BBL17b, BOS14, BP11, BL15, BYY17, BL18, BYY18, BW19, BW20, BC13b, Bou11, BT15, BGJB16, BS16, Brz15, ByLbH16, BWbH^{+}18, Bur14, CTL^{+}19, CD16a, CD17, Cam11, CSS^{+}16, CPY13, Car16]. citation [CD16b, Cha13, CHL15, CHC17, CLHH10, CL11, CY13, Chi15, CG17, CG18a, CGG19, CS11b, CA19, CL16, Cop19a, Dan14, Das16, DD18, DC15b, Don17, DGDG11, DGDGSV15, DGGBDG17, DB16, DGF17, EMSH16, EG18, EMH^{+}10, ÉMS^{+}13, ES18, Eto13, EHK12, Fan19, FFR^{+}17, FFR16, Fia11, FA10, FZZ^{+}12b, Gál17, pGDTP12, GD11, GD16, GK18, GZ11, GSTD11, GTD14, GTC16, GHT16, GT18, GSKM17, GAE15, GXC^{+}19, GBDG19, GTGABAG15, GTAG18, GGP14, GGS14, GS12, Hag10b, HT11, Ham11, Har13d, HSAK18, HB18a, HSK18, HLC17, HT19, Hou17, HYC18, HHGZ11, HRC12, HWS18, hHC15, HC16c, HW10, Hur17, IA19, IL14a, ILGZ^{+}14, JBC19, JDH12, JDLIV14, KLP17, KPJ16, KZ13, KY16, Kim10, KCK14, KJES16, KPJT14, Kos18b, LV11, LGPC18, LZB10, Lee15, LK17, LS17b, Lee19b, LT10a, Ley13a]. citation [LWB18a, LWB18b, LPL16, LQW17, LWL17, LC18, LG10, Lin18, LSY11, LC12, LCD^{+}14, LABL13, LWT16, LDZ17, LS19b, MY16, MHLGHV14, MB15, MBT16, Meh07, MDG10, Mik17, MRS^{+}16, MPH19, MT13b, MKHB13b, MKF14, NSKO15, OKCPS17, OZK11, OM11, Ort16, PNS^{+}10, PKR15, PYW18, PW17, PLJ18, PAL13, PW13, PRA16a, Pra12d, Pra17a, Pra18c, Pra19b, PML14, QL12, RAA18, RGGBV16, RCJ18, RPAMR19, RKT^{+}15, Rod17, RPK16, RP17a, RPK17, RPK18, RMH14, RC13b, San12a, San12c, San13, SBT18, SG10, SvLVA19, Sch13b, SA17, Shu17, SL12a, Sma10, Sma11, SBK19, Smi12, SI17, Soo17, SGY15, SH18, SE18, Ste19, SX16, SD18, TB19a, TK10, TA14a, TA18, TA14b, The16, The18a, The18d, The18f, The19a, TN19, TDG17, Tol11, TSRGCCJC14, TSRGG17, TC13, yT15]. citation [yTmShL16, VYL17, dFVDU^{+}19, WM19, WvE13, WZX11, WYAY12, WQY12, WZLZ13, WTG15, WLMF15a, WLMF15b, WWH^{+}17, WZCC19, WST14, WDN17, Wra16a, WW11, Wu13, Xie15, YWS18, YMSQ10, YHL19, YST12, Yos13, YLH^{+}17, YP19, You14, YWY10, YL10, YYL10, ZJLG10, ZFY^{+}17, ZG17d, ZG17e, ZG17c, ZYNZ18, ZS11, ZC14, ZZLS19, ZGJ18, Zit15, ZLF18, ZXT^{+}19, Zon19, Zuc10, dS17a, dS17b, vRvLV11, vWWtH14, McC18, MPS^{+}18, YLL^{+}15c, LV12]. citation-analysis [KZ13]. citation-anchors [AA18]. Citation-based [HTL15, MCR^{+}12, PRDG17, vEW17, Ask18, Bel17, GTD14, HLC17, LPL16, MKHB13b, RPK16, RP17a, RPK17, Shu17, Smi12, SH18, ZG17d, ZG17e, dS17a, vRvLV11]. citation-lexical [LZB10]. citation-related [San13]. citation-words [Zit15]. Citations [ADD10, Abt17, BI10a, BHKP11, Bha11, Ort15, Ran09, SBK19, Tom17, ANOdFC12, AL12, AC13, AdAdAM10, Asu19, Bak17, BIH17, BPJ^{+}14, BM14b, BL15, BL18, BHH18, BRS^{+}16, ByLbH16, BWbH^{+}18, BNV11, CC11a, CC14, Cam14, Cam17, CC10a, CKB^{+}14, CvLB10, DdS19b, Egg10d, Egg14a, EG16, Fan13a, Fan13b, Fan18, FMS17, FE14, FE16a, FE16b, FMM15b, FMM16, FAA13, GP15, GPN14, GW17, GSOLHO19, GBMA14, GMSZ18, HM18, Har12, Har17, HII^{+}18, HPS19, HPKS18, HNG19, HdSV16, HQY^{+}18, IBL13, JN11, JYW11, JMM19, JH16, JW18, Kol12, Kra17, Leb12, LY16a, LYWSV13, LRZ13, LH12, LABL13, LSE^{+}18, MSdBC16, MGLZ10, MBR^{+}13, MA19, Mes11, MDG10, ML10, MGC19, MPM18, MdNS^{+}19, NG16, NA18, NPP^{+}12, PJY17, PKL^{+}16, PB17b, PZ17, PMN16, Pra19g, QA19, RMA12]. citations [RW11, RP17b, RBBG18, San12d, San12c, Sch12a, SGG^{+}14, SAR19, SRW^{+}15, SRW18, Sni16, SMM15, TAA16, TB19a, TK16, TK17, The18e, Tom18, VTY17, VT10, Wad18, WRV14, WZFD19, WW11, XXL^{+}17, XGCK19, XGL^{+}19, YSM^{+}19, YQX10, YK14, YK15, YSY^{+}13, YST12, YS14, ZCL14, Zha17, ZXLX14, Zit11, dW15, MMOMLC18b]. cite [COS11b, Har12, LT10b, dSM17]. Cited [Bha11, Ioa06, LCS^{+}16, WLZ^{+}19, BB15, Bor16, BWdMA17, BBS17, BL18, BC13b, COS11b, DC19, DW18, FH16, GRSFV19a, Ham11, Har16b, Har19a, HBS^{+}19, HAJ12, Ho13b, HH17c, HLW19, JNA18, Kor18, KPS12, KN15, Ley12, Li16, LHW16, LDZ17, MXZ18, MMOMLC18a, MHC^{+}15, MB16a, MHTB17, MBA13, MBTKA14, MC12, NvLvR10, PLA10, PAL13, PCRMCB^{+}18, Per10, RH18, SCGZR16, TKA17, The19a, The19b, VG17, WYY11, WBX18, WFZD19a, Whi15, XHA^{+}19, YY14, YHL19, YSY^{+}13, ZG17c, ZWW^{+}18, ZW18b, dSD17, dSBC17, vR17]. citedness [MHFB17, PKSG12]. CitedReferencesExplorer [TMLB16]. Citer [AW10, Egg13b, FMM14, LAW14]. citer-based [FMM14]. citer-success-index [FMM14]. citers [FMPP10, Fra17]. CiteScore [dSM17]. CiteSeer [Fia11]. CiteULike [SMM15]. cities [DSH^{+}10, qJnShPL17, Li19, Pra18a]. Citing [GBMB10, CCM^{+}11, EO14, FSSPG^{+}15, HBDL18, Kra10, PYW18, WBX18, ZYZ14, Zit11]. citing-side [Zit11]. citizen [LWB16]. CitNetExplorer [WLH^{+}17, vEW17]. city [ISL18, qJnShPL17, xShLY^{+}15, dCdAMB19, HQY^{+}18, LOMLPA^{+}17]. CIVETS [YQW13]. Civil [LOPAGS19, JJR10, Kaz14, KPJT14, RAA18]. claims [CRLMRPA17]. Class [SFM16, BM13a, BM13b, BRS^{+}16, GRSFV^{+}14c, HF19, LP12]. class-selective [BRS^{+}16]. classes [AD13, CFL12, GTD14, TDG17]. Classic [Ho14, The19a]. classical [Cam11, Egg13b, Geo17, GN19, MBL18]. classics [CL16, MHLGHV14, MHLGHV14, MHC^{+}15]. Classification [LKR14, NA12, BBJS16, Bor18, Cha18b, De 16a, DLMX15, GNVQdMAG11, Gom19, GSE^{+}18, HL18, KBT14, Kos16b, LBW17, LHCH18, LS19a, MYP19, PY14, TZG15, VDV16, VPM16, ZJLG10, ZLW16, ZZD^{+}18]. classifications [DC15a, Ley15b, LKY17, RGCM14]. classified [FZQ17]. classifier [MXZ18]. classifiers [GR14]. Classifying [GDP16, Li18]. cleaning [XHA^{+}19]. Climate [BP11, AM18, BHM16, JVM17, MHFB17, MHTB17, SK16]. Clinical [KHK13, AAB^{+}13, ÁBDFB19, CHC17, IPIU13, LNK^{+}14a, LNK^{+}14b, OCM^{+}12, APFR^{+}13, RRLNAG15, SZMS18, TK16]. ClinicalTrials.gov [TK16]. clique [Keg15]. Close [Fie15a]. Closed [BWD12]. closer [DMV10, FFL16]. closure [PW13, WhCL10, WBX^{+}17]. cloud [DGF17, HH17e, HC19]. clues [Cab11, CKB^{+}14]. clumping [ZZPG14]. Cluster [AF15a, ILB13, KLCS14, KGL^{+}14, Mad15, SH15a, YLJ^{+}17, AF15b]. Cluster-based [AF15a, AF15b]. clustered [LBW17]. Clustering [KBT15, WK17, YPNS14, CLL^{+}17, GN19, JC12, Kim19b, KKOS19, KGZML^{+}19, LCS^{+}16, LGD11, LGD12, MLT^{+}15, TZG15, TG18b, TT13, VYL17, WZLZ13, ZLLD19, ZWL^{+}18, vEW17]. clusters [AGHL14, Cho12, GT11, GT17, KLCS14, KW17, Bor16]. Cluttered [OK13, Lei16]. CNPq [PPK^{+}16]. Co [AW11, Fie15b, Hol10, OGOPPR17, OKK14, RFGBMA13, SSN19, Whi15, ACD14, Aus13, Aus14b, BHB13, BAB13, BHS14, BP11, BL15, BHH18, ByLbH16, BWbH^{+}18, CHM15, CMUdF15, CD16b, CHL15, CL11, CJC13, CLSW19, CY13, CRFM^{+}12, CHL10, CFK16, Dan14, DAMC15, DVB14, DC15b, EGR13, Eto13, Fan19, FK16, FZZ17, FFR^{+}17, FFR16, Fie15a, Fie15c, FDVZ16, GW15a, GGR11, pGDTP12, GLS16, GA18, GTGABAG15, HD17, Hen16, Hen18, HYC18, HHGZ11, HHDL13, HLW19, hHC15, ICC16, JCK11, JX13, qJnShPL17, KM15a, KO19, KY16, KTLD16, KB18, KGG15, KPSL12, LGL10, LSS15, LSS16, LQW17, LH12, LHW12, LC12, LX15, LM16, LCIADG19, MJHG13, Med18, MYP19, MKF14, OVJM17, OZK11, OROMAA16, PY14, PYL16, PROGMA10, PB16, PLG19, PML14, QDY14, RAS15, RGGBV16, RPAMR19, RP17b, RBBG18, Rou12a]. co [Sch12a, Sch12b, SL14, SD13, SJ19, Sma10, SL10, Tol11, UHAR12, VuHL10, VO17, WZX11, WLLL12, WZLZ13, WCL14, WZW15, WS13a, Xie15, YSM^{+}19, YLL15b, YWC12, YHL19, YB14, YSD11, ZLL^{+}17, ZS11, ZC14, ZGL^{+}17, ZLH^{+}15, ZZW^{+}19b, ZSY^{+}13, ZLF18, ZCMVQS11, dSF13, vdPR18]. co-author [Aus13, DAMC15, EGR13, MJHG13, OZK11, Rou12a, Sch12b, ZGL^{+}17]. co-authoring [KM15a]. co-authors [Aus13]. Co-authorship [Fie15b, OKK14, SSN19, Aus14b, BHB13, CMUdF15, CRFM^{+}12, CHL10, CFK16, Fie15a, Fie15c, FDVZ16, GLS16, GA18, GTGABAG15, Hen16, Hen18, HLW19, ICC16, JCK11, JX13, KY16, KGG15, KPSL12, LH12, LX15, Med18, OVJM17, OROMAA16, PYL16, PROGMA10, PB16, RGGBV16, RP17b, SL14, SJ19, Tol11, UHAR12, VuHL10, VO17, YB14, YSD11, dSF13]. co-authorships [BL15, CHM15, FK16, KB18, Med18]. co-citation [ACD14, BAB13, BP11, ByLbH16, BWbH^{+}18, CD16b, CHL15, CL11, CY13, DC15b, Eto13, Fan19, FFR^{+}17, FFR16, pGDTP12, GTGABAG15, HYC18, HHGZ11, hHC15, LQW17, LC12, MKF14, RPAMR19, SD13, Sma10, WZX11, WZLZ13, Xie15, YHL19, ZS11, ZC14, ZLF18]. co-citations [BL15, RBBG18, Sch12a, YSM^{+}19]. Co-cited [Whi15, YHL19]. co-classification [MYP19, PY14]. co-evolution [HD17, vdPR18]. co-existing [PLG19]. Co-inlinking [Hol10]. co-invention [CJC13]. co-inventor [ZZW^{+}19b]. co-inventorships [GGR11]. co-link [LGL10]. co-occurrence [BHH18, LSS15, LSS16, LM16, LCIADG19, QDY14, SL10, ZLH^{+}15]. co-opinion [YSM^{+}19]. co-patent [qJnShPL17]. co-publication [BHS14]. co-publications [WS13a]. co-publishing [KTLD16, PML14]. co-term [ZCMVQS11]. Co-word [AW11, OGOPPR17, RFGBMA13, CLSW19, Dan14, DVB14, FZZ17, GW15a, HHDL13, KO19, LHW12, RAS15, WLLL12, WCL14, WZW15, Xie15, YLL15b, YWC12, ZLL^{+}17, ZSY^{+}13]. coactivity [CHY16]. coastal [dCdSNB15]. coauthors [Aus14a, BLA16, Bou14b]. coauthorship [BLdlCV17, CÖT16a, FB16, GPN10, Hir10, KD14, LL15, XXL^{+}17, Xie19, YDZ10]. Cobbold [MKP16]. COCI [HPS19]. cocitation [Zhe19]. code [HF19, LPB14]. codes [BM12b]. coefficient [Rou11, SS10a]. coevolution [XXL^{+}17]. cognition [SMCC18]. Cognitive [PSY^{+}19, Bru10, CdMCdMMdP17, GSE^{+}18, ML13, Oze12b, RGLE16, xShLY^{+}15, TW10, VYL17, BD12b]. Cogno [JAAA18]. Coherence [LKSK15, RM10]. coherent [YLH^{+}17]. cohesion [AP16, TCC17]. Cohesive [CH14, Keg15]. cohort [CIL^{+}16, SST^{+}16, ZXT^{+}19]. cohorts [CÖT16b, ÖS17, TCB16]. coin [dOM16]. collaborate [IDKF17, SLXD15]. collaborating [LBGBdMA13, WXW^{+}13]. Collaboration [KM15b, KMFD12, LSCK12, LCC12, Mor19, Ort11, Oze12a, PR10, PR14, RPGM16, RHGKD16, TCB16, WLR^{+}14, WHL^{+}15, AAH10, AHUR11, ADS10a, ADM14, ADD19a, ADD19b, AER^{+}14, ACFL11, AGHL14, APR19, Ama16, Ard12, Asu19, Avk13, BFMRM19, BH10, BB19, BKG16, BPGGdMA12, BKSS15, BAC13, Bos10, BPHL16, Bre13, BFM^{+}14, BMD^{+}18, CGC18, CÖT15, CH13a, CH14, CBF13, CF14, CG17, CdJD15, CL17a, Cho12, CB16, CdSPdM13, CQB16, EES13, ELP11, FYC15, FUR10, FK16, FKM^{+}15, FFL16, Fin15, FB16, GGG14, Gau17, GD11, GZGAC16, GZGAC17, GHS18, GKK15, GPN10, GRG12, GLM11, HSL^{+}14, HTHB11, HG17, HdSV16, HEH18, wHwH11, HCL14, HDC13, HYC15, IBL13, JK19, JvGH10, Jar10, JCK11, JC12, KM15a, KLM16, KA13, KA17, KLL14, KD19, KKBW17, KB11c, Kos16a, KCM19, KdBBK15, KGG15, KJ14, KM16]. collaboration [KR17b, Lan13, LZZ^{+}12, LZZ^{+}13, LT16, LLLL18, LCWY12, Lia11, LZC17, LNK^{+}14a, LNK^{+}14b, LRA14, LXH^{+}18, MLC14, ML18, MM16, MM14a, MOA16a, MR15, MCLL17, MLVJ12, Mêg13a, MGB16, NHLL17, NA18, NQ14, OVJM17, OA10b, OA10c, OE15, OPGW^{+}13, ONB17, Oze12b, PCR18, PY19, PML^{+}17, PS15, PJB^{+}12, PPI17, PH14, Pra13, RVFEdlM10, RRLNAG15, RRSA18, RPDCRVRP15, RPK16, RF19, SFBS17, SH15b, SS10a, SIR^{+}14, SIS17, SS10c, SL13, SC10, SLK12, SB14, Soo10b, Soo14a, Soo17, Soo19, THAL15, TS11a, TS11b, TA11, TP11, TA17, VuHL10, WWL17, WXLL12, WXW^{+}13, WLN^{+}14, WTG15, WHLP16, WYB^{+}17, WY19, WDL17, WOW10, WD13, YYP17, YB14, YHL^{+}18, ZHG16, ZCKZ16, ZLYF14, ZYSS14, ZG17b, ZW17a, ZRL18, ZSCR^{+}18, ZML19, ZG11, ZZW14, ZLH^{+}16, ZZ16, ZZW16, ZLLL19, ZZZ^{+}12, ZyZZ^{+}14, ZZY13, ZT14]. collaboration [ZL15b, ZB15, ZZFD18a, vdPR18, vPD13]. Collaborations [KM18c, ADS12, ADD18a, CGKB18, GLS16, GR14, qJnShPL17, KJ13, KRR14, LMM15, MSdBC16, MGGdP17, OI17, PSB^{+}17, QMSM^{+}19, SW19a, SLXD15, TUCR15, WWP14, WWP17, WBX^{+}17, ZCZ^{+}16]. Collaborative [TA15, dSFSF15, AhOL14, AOd15, BGAAM15, FSO11, GGR11, GYZ15, GSPLVG^{+}18, LT10a, LG15, OM11, Rou11, SS10a, UCH19, WZ19a, ZS17, ZLF^{+}14, ZS11, ZTRH18]. collaborator [KJW^{+}17, QZL^{+}17]. collaborators [DCS12, GLD16, Moe16b, QA18]. collapse [WHW^{+}19]. collected [ÁBMB17]. collection [ACAGD^{+}17, Ben12, GBM^{+}16, Jac18, Liu19, MSB18]. collections [SK13]. Collective [ZMW^{+}18, Cha17b, MAA^{+}11, WFZD19b, WG11]. college [PHS12, Tei11, WM19]. colleges [GN17, PNVCB18]. collegiality [FMU16]. COLLNET [KÖG12, KG13]. Colombia [ACAGD^{+}17, MJC14]. Colombian [BDC^{+}12]. colon [vWWtH14]. combat [KM17b]. Combination [YAC10, AD16, ACC^{+}16]. combine [ZZPG14]. combined [ByLbH16, CRZGVQMA15, KKK^{+}14, MCLL17, MG12]. Combining [Ano15, LJMF15, MS15a, WMH^{+}17, FDVZ16, HQY^{+}18, KPY16, MM15b]. Comment [Ano18b, BGJ^{+}16, Shu17, Ste17, BD16b, Hir19b, JW18, Lor10, OL11, Pra12a, Zit12]. commentary [Sch18b, Pat18, Wu18]. Comments [Ano12a, BI18a, Bur12, CH12, Egg14a, Egg14b, dCPF14, Ho13a, LFLG14, Liu16, PA12, Pra11c, Pra17a, Rou11, Rou12a, BB17a, BWD10, BWD12, BHJD12, GdA14, HW12, LM13a, LZH^{+}13, MS16a, Pra16b, Rod16, Zon19, vL12]. Commerce [Glä18, LHC16, CYH13, SD13, Tsa15, YLL15a]. commercial [MT15, yTwTlW19]. commercialization [CK14, GP18a, WHLP16, WS13a, YC10]. commitment [LBRR19, YC10]. committees [Cha19a, SA12]. common [Jar16, Kos14, Tod11, Van10, dSTL18]. commonly [Hal13]. commons [dM10]. Communicating [Ano11]. Communication [Pen19, RCETS19, Bai18, BHKP11, CLLZ15, CD18, CyPP12, CP12c, CBKL13, FLB11, GRSFV18, GBHT16, HT14, KLPP16, LJS16, LYS^{+}17, MY16, NSC13, PZ17, RCCM14, RHGKD16, SFNO12, SST^{+}16, SX16, XDB^{+}19, YP19]. communications [GD17]. communicative [Sug11]. communist [JKMS17, Pra18d]. communities [AAH10, APT13, BMZ^{+}17, CZPR17, CyPP12, HGH17a, HLW19, hHSL19, KMFD12, KM16, MGGdP17, SPdSM16, Sob11, TA11, URU10a, YDJ12, ZGL^{+}17]. Community [Bos10, ELP11, ZGL^{+}17, BSG17, BHJD12, BH15, CQB16, De 13, Gar15, GZM15, HPBI^{+}14, KKLP17, LML11, LZFW15, LAdAMJ17, MCLL17, MHKB16, RPGM10, RG18, SK16, Vel12, WY19, ZB12, dCdSNB15, Pou10]. companies [CC10a, LLC^{+}17, WZX11]. company [HR15, LJ10]. Comparative [AS18b, BL10, CHC17, GHA^{+}15, JMM19, MC13, MK19, QDY14, SMY15, ddMS15, Ano17c, BPTG10, ÇAAÇ15, hCyL12, FM11a, FSOS12, GGS17, GdOdAG^{+}13, GW10b, HG17, HC12, hHC15, HC16c, KZSZ19, KHJ^{+}12, KB18, KJ14, LRS^{+}18, Liu13, LXL15, Liu16, LXL16, LZL10, MSDJ19, Moe17, MPH16, NSKO15, NASR11, RCdJ^{+}14, SH15c, SWCH14, Tom18, Wan16, WW11, YQW13, ZPG^{+}14, ZYG15, ZP15]. compare [ADS10a, BL18, Har19b, RD13]. compared [ADV11, BG18, BTL19, Med15, Ole12, PDAN19, Sot12]. Comparing [ABILO10, Ama15, CGZ10, CGG19, DGDG13, Mik10, Sin18, SJOC18, SVCFI14, SH18, BB17b, HDW^{+}15, KW17, MR10, Mêg13c, NSC13, PP18, SvLVA19, Tsa11, Gus19, Sak19]. Comparison [BLS15, CG18a, CA19, ES18, FH13, KKCG18, MH16a, MBR^{+}13, MGC19, MKHB13b, OCM^{+}12, PPK^{+}16, SG10, SGG^{+}14, TYWZ12, VBG^{+}17, YK15, ZCL14, ADD11a, AZKR13, ACORC10, AYS14, Ask18, BSBG18, Cha13, CHY13, Che18a, Che20, CA12, Fra10, GRSFV12a, GTMRE^{+}19, GTD14, GB14b, GZ14b, GW15b, GSE^{+}18, HA16, HIG^{+}17, HHBB18, HWL11, KCP12, KZC16, KHK13, LV11, LKY17, LLH^{+}16, LJC^{+}15, MMOMLC18a, MBT16, ML10, OMAT19, Ort18, PRRC16, PROG19, PPE14, Sch11a, SSS^{+}11, SM17, Sot10, SN10, TBS15, yT11, yTmShL16, Vin12b, WvE13, Wil15, XTZ15, YY14, YYDH12, YWW17, Yur15, ZW19, vPD13]. Comparisons [MS18a, Wu14, BM14a, CSO17, FLZ17, KKL14, Moe16a, Tod11, VHH16]. compatability [MS18b]. compensation [HFW^{+}14]. competences [CC12b, SB19]. competencies [LLC^{+}17, WF17]. competency [Lee12]. Competing [PLG19]. Competition [GRSFV18, HBT16, SFNO12, YPK13]. Competitive [Zin16, ACD11, ADD14b, Fan11, LF12b, Rou19, RCCM14, VGPdlC^{+}17, WLD^{+}14, ZZPG14]. Competitiveness [ZKC^{+}16, DCGZ^{+}12, DNAH15, DQ11, MLOY18, WMT^{+}12]. competitors [HF19]. compilations [HH10]. complement [DGDG11, RNM18]. complementarities [LWB16]. complementary [Dan14, FZZ^{+}11, PLG19, Sah16, VG14]. complements [dFVDU^{+}19]. complete [ACD13]. completely [MSC18]. completeness [ÁBMB17, CIL^{+}16]. Completing [MM15a, AA16]. completing- [AA16]. Complex [ÁCCG^{+}15, KPRT16, LCIADG19, ANOdFC12, KRP19, Pen19, RP17a, ZG13c, ZWHH13, Zin16]. complexity [GZ17, LRC19, Mag14b, WZS12]. compliance [Jam17]. component [TG18b, YZW^{+}17]. components [Ley11b]. composers [Geo17, GN19]. Composite [BBSS16b, BBSS16c, BR11, MvdH13, PDAN19, Pra16b, tScL13, BBSS16a]. composition [Kim10, PRSB16, Pir19]. compound [AP14]. compounds [ATCCAAB19, ZXH10]. comprehension [IA19]. Comprehensive [GQAM19, JDG14, LZGQ13, Moe16a, SL12a, AAV13, CCLL14, GRTPMLAJ19, SRGMF15, SK17, TAA16, WMT^{+}12, ZS11, YLL^{+}15c]. Comprehensiveness [yTwTlW19]. comprising [APFR^{+}13]. Computation [PT17, Cam14, Egg14a]. Computer [CHM15, CVC14, FAA13, GZM15, HP10, SLD^{+}17, SUP15, ATK17, dSAEE15, BCML19, BI10b, BSG17, Cab13, Cav15b, Cha18a, DRCG17, ES16b, ERW12, FM17, Fie15b, Fra10, FSOS12, GLS16, GKV11, ILB11, ILB13, IBL13, IQT^{+}19, LL13a, LRS^{+}18, LSM^{+}15, LHBC18, LSE^{+}18, Mue18, QRJ^{+}17, RMCM13, SB19, The17c, USPO15, ZY15, vEW10]. Computing [KHH18, VHH16, CVC^{+}15, CLLH15, GSM^{+}16, Ho16, HH17e, LVSL18, LLGW13, NH11, ZW11]. concentrated [Pir19]. Concentration [Ioa06, Jon10]. concept [FRdA16, GZ17, HRB^{+}13, HRB^{+}14, LWM^{+}15, LM15, MHLGHV14, McC14]. concepts [ANOdFC12, CNPG17, MG12, NSC13, PPE14, SB19]. conceptual [BM19, Car16, CLMN19, GZJ^{+}15, MAGSTRC15, PCRMCB^{+}18]. conceptualisation [Bor15b]. conceptualization [San12c]. concern [Glä18]. concerning [LSL15, SM17]. condensed [TH19]. Condorcet [MOA16b]. conduct [KM18c]. conducted [Bjø19]. Conference [Ano10, Ano15, GGH^{+}14, HSBW10, KÖG12, KG13, LHG16a, LLRG10, OING12, SSAG16, ADM19, ATK17, BSG17, Bar11, Bar17b, KKV^{+}13, LMKG19, LYS^{+}17, Lee19b, LRS^{+}18, MF14, MAA17, SFNO12, SHK14, WBX^{+}17, BI10b]. conference-related [Lee19b]. conferences [dSAEE15, ERW12, ICC16, LMKG19, MGLZ10, SA12, SDP^{+}19, WBX^{+}17, Web16, WS13b]. Confidence [Bor17, Ste17]. configuration [Ley11b]. confirmation [CdMCdMMdP17]. confirmatory [GRSFV16a]. conflict [JPZ^{+}10]. conflicting [GRSFVdMA14]. Conflicts [LS15]. conformity [EO14]. confused [AM18]. confusing [GG15a]. confusion [Sch15a, Wu18]. congruence [MSC18]. connect [Vil10]. connected [BSKB17, SK16]. Connecting [RG18, Hsi11]. connection [Cha18b]. connections [BL15, Tu19]. connectivity [KGL^{+}14, dFPYdCL12]. cons [Dem18]. consequences [HBT16, KH17, MG12, Hal14]. conservation [CWJBT10, CLD13]. consideration [KM18b]. Considerations [BN14, BH17b, GRG12, Hei13, PHDC16, dSSdMAF14]. considering [LWL17, WLMF15a, WLMF15b, WZFD19]. consistency [KKL14]. consistency-driven [KKL14]. consistent [De 16a]. consolidated [FRdA16]. consolidation [GAPP18]. consortial [TÜ10]. constant [HT19, SS18]. constraints [BGBS18]. Constructing [CLHH10, LMdBG16, HFL14, LSS15, LSS16]. Construction [dCPRP18, JYM^{+}16, xShLY^{+}15, WDS16, YHL19]. consumption [BZBLP16]. contain [Har15b]. containing [LHTL18]. contemporary [ZRL18]. content [BHPVdPMR18, BWD10, BHJD12, DGPL15, FSLR10, FA10, GPN10, Hol10, LRY18, LDZ17, MS18a, QA19, RKZK18, SHL15, TA18, WOW10, WHH^{+}18, YPK13, ZAJ19]. content-based [FA10, LRY18, LDZ17, TA18]. contents [Boy17b, ESB15]. contested [BH16a]. Context [DB16, Yu17, AD18, Ano18c, BIH17, BASL16, BHH18, Cam18, CA12, DCM16, DGGBDG17, Eto13, GVGSEPRC15, GTAG18, HII^{+}18, HSAK18, JW18, Kaz14, Kaz15, KPJ16, KP12b, LCD^{+}14, MWH14, RGTSLCH14, RC13b, SV19, Sma11, SH18, VDV16, Vel12]. context-based [Eto13]. contexts [Cha13, SMAABJ11, Sma10]. Contextual [LPC17]. Contextualization [KWS17]. continent [CRLMRPA17]. contingent [Vel12]. continuance [CHM15]. continue [BL11b, dSBC17]. Contradicting [WDL17]. contrarian [JVM17]. contribute [OM11]. contributed [Cha14, Cha18c, Cha19b]. Contribution [DMM13, Fan15a, AG13, BMM14, DWGL16, EGR13, JX13, KB13, Lar12, LKW^{+}16, TG18b, VG14, YHC^{+}15, YWW17, Zha14]. Contributions [KKE13, ALYZ15, Ano11, CH15, GK14, LLYC14, PL18, TBW^{+}12, ZL17]. contributor [HO19]. contributors [WW15]. control [EdS19, RF19, ZZ14]. controlled [CIL^{+}16, KCT^{+}17]. controversial [Cop19a]. controversy [JVM17]. conundrum [ACS18]. convergence [BFMRM19, GMM16, JKC15, KCK14, KL17, LLW13]. convergent [BTL19]. converging [BNV11, JAAA18]. conversion [AA16]. convolutional [LHCH18]. Cooperation [KRP19, UMK14, BHS14, IJF16, JJR10, RPGM10, SC18, ZG10, dPSS18]. Cooperative [ZZW19a, Cha17a, Xie19]. coordinates [GGG16a]. cope [Bue15]. Copiello [Zon19]. Coping [GBSZL15, Ano16b]. copying [GBSZL15]. Copyright [Jam17, LL16, LP18a]. Core [BH16a, BH16b, Cho12, Jac18, KLM16, Liu19, MSB18, Vin19, Aus13, Aus14a, Bou14b, CRR14, CS11b, Col17, Egg11f, GT11, Glä12, GT12, GT17, HLL14, HYYL12, HLY14, KMP11a, SD13, WQY12, WF17, XCS^{+}16, YZW^{+}17, YR10, Zel12, ZGY16, CKPY19, ICC16]. Core-periphery [Cho12, KLM16, Zel12]. CORE-ranked [ICC16]. coreness [PY14]. cores [SCGZR16]. corporate [CC10b, Che11, CSC12, TE18, vPD13]. corporations [PYK13]. correct [FMM16]. correcting [Wra10, dSD18a]. Correction [APPF18, BT18a, BT18b, BW20, Che20, Dya17a, FK18, Hen20, JL19, Kim19a, MM19, RV18a, ZT19, ZWX22, ZTP18a, Ano16a, Egg13a]. Corrective [DdS19b]. correlate [The18f]. correlated [HCL14, SMM15, Tor13]. correlates [Lun19]. Correlation [ABMSSP16, EMH^{+}10, wHwH11, SM16b, LHW16, APFR^{+}13, QDY14, WvEvL^{+}11a, XGCK19]. Correlations [AND19, WV13, AAB^{+}13, MK19, The16, TC13]. correlative [WF17]. correspond [WM19]. correspondence [MSL11]. corresponding [FMS17, Han11, MSL11]. cosmology [FG15, MB10a]. costs [FPS14, HH15b, Hsi11]. Could [HO19, ADD10, CLD13, Sot12]. Council [RGdCMM17]. Count [BL18, Rig13, WFS16, Yu17]. Counting [AL12, ML10, Ber18, Hag10b, MAGBBM13, Osó18, ZZZ^{+}14]. countrie [GK18, Mue16, ZHG16]. countries [AvLS14, BKZ^{+}16, Bar17a, BSvEK13, BSK15, BGAAM15, Bos10, BPHL16, CRMPRS18, Cho12, CMM17, DMM13, DGWZ13, DB19, EN17, Fan12, Fin15, FB16, FKRS14, Gan12, GCGP10, GTMRE^{+}16, GTMRE^{+}19, GAVZAB12, GRG12, GdOdAG^{+}13, GZ14a, HV18a, HH18, Ibr18, IPIU13, JL18a, JL19, JKMS17, Kar12, KMP^{+}11b, KHJ^{+}12, KHK13, KBL15, KWW15, KEP^{+}18, KBZS15, LBGBdMA13, LRC19, LA19, Med15, MSB18, Moe16b, MZE19, NJ10, NP11, ÖS17, OM11, PCR18, PP18, Pir19, PL18, Pra10d, Pra18d, RV18b, RLW14, RPK17, San13, SS10c, SK18, SAR19, SZAJS14, TA14b, TE18, TBT19, UMK14, Vil10, Vin12b, WLDW12, WXW^{+}13, WLY14, WTM^{+}16, WM17, WF18, XTZ15, YYDH12, ZGCRVQ18, ZG11, dPSS18, vZ13, RV18a]. countries/territories [XTZ15]. Country [BL17a, GNVQdMAG11, Pie18, ADD16, AvLS14, ACT18, BFMRM19, Bas10, Ben15, BB15, CHY13, CR14, DAYY18, DC17, Fuk17, GZ14b, HSL^{+}14, IFT^{+}18, KK19, LCY14, MCCU16, Mes11, OA10a, OA10b, PR14, Sch11a, SM12, SAR19, Sot12, SWCH14, Vel12, ZZZ^{+}14]. country-level [KK19, PR14]. Country-specific [Pie18]. counts [ATK17, Bas11, BH16c, BYY17, BL18, EHK12, FA10, Hal14, JX13, Kos18b, MvdH13, MBT16, MPH19, SBK19, SI17, SWCH14, TB19a, The16, The17a, The18f, Tol11, dFVDU^{+}19, YP19, ZXT^{+}19, Zon19]. coupled [Sch12a]. coupling [CHL15, GD16, HA19, HC14b, hHC15, Liu17, Soó14b, TZG15, YSY^{+}13, ZLW16]. cover [WLM15]. Coverage [HPBI^{+}14, MMOMLC18a, CHC17, Fan13a, GRTPMLAJ19, Har14b, HB17b, LvI10, MS12, MF14, MPH16, SL12a, VBTK19]. covered [RV18a, RV18b, Shi14]. CPC [YLL^{+}15c]. crammed [CS19]. CRC [Tom17]. create [HSPY15, OMMMTLC17]. created [CÖT16b, Kon12]. Creating [BMZ^{+}17, Car16, YG18]. creation [EES13, Fuk16, GP18b, Hu11, LM13b, RPGM16, URU10b, ZS17, ZZW19a]. Creative [Hei13, LHLH19, Won13, GY12]. Creativity [Sol06, FR11]. credibility [Fra17, HO19, LWIB16]. Credit [CXZ19, Tol11, BZ17, Ber18, Hag10b, KD14, LF12a, MdBdP^{+}19, MCB15, TBW^{+}12, WSC16, WFZD19b, dPdCAdMC^{+}16]. credits [Osó18]. CRExplorer [TMLB16, TBMM18]. criminal [Wal15]. criminology [Wal15]. CRIS [vLvWW16]. crisis [FRdA16, HC15c, HC17, KKT^{+}18, MMAHS10]. criteria [ND16, OCM^{+}12, PRDG17, Siv16b]. criterion [Fan19, ZLF18]. Critical [KB11a, Moe17]. criticality [LWB18b]. critically [Bas10]. Criticism [FM11b, Jac12]. critique [PA12, VB12, vL12]. Croatia [BPVM11, Bra12a]. Croatian [JZL10]. Crohn [Kos14]. Cronin [Ano14]. crops [JBC19]. Cross [Ben15, LDVSGD19, MS18b, AHUR11, ADD11c, BFGVV^{+}18, BN10, BL11b, CGKB18, COS11a, CHY13, CC11b, CKB^{+}14, DC17, Fie15a, Fie15c, GY12, GZ14b, HA16, Hos11, qJnShPL17, LRY18, LCY14, OO12, SMAABJ11, VEJC^{+}18b, WQY12, YWS18, ZCW14, ZJLG10, ZG17d, ZG17e]. cross-citation [WQY12, ZJLG10]. cross-city [qJnShPL17]. Cross-country [Ben15, CHY13, DC17, GZ14b, LCY14]. cross-disciplinary [BFGVV^{+}18, CC11b, Fie15a, Fie15c, GY12, HA16, YWS18, ZCW14, ZG17d, ZG17e]. cross-discipline [COS11a]. cross-domain [LRY18]. cross-generic [SMAABJ11]. Cross-metric [MS18b]. Cross-national [LDVSGD19]. cross-sectional [BN10, BL11b, qJnShPL17, OO12, VEJC^{+}18b]. cross-time [AHUR11, ADD11c]. cross-topic [CGKB18]. Crossing [BBCP14]. Crossref [HPS19, Har19b, Ort18]. crown [WvEvL^{+}11b]. crucial [Cam12, Ley15b]. Cruz [STCRPA18]. cryptographic [Pal15]. crystal [Cha19a, HWLL14]. crystals [SLG10]. CSCW [CPF18]. CSE [AT17]. CSIC [Ort11]. CSS [BYY17, BG17, GT18]. CSSCI [WDS16]. Cuban [AJdMA10, AJCACRdMA16, CRAJdMACÁ15, CRZGVQdMA16, GRTPMLAJ19, GSPLVG^{+}18, PCR18, RPK16]. cue [QA19]. cue-terms [QA19]. cult [Etz13a]. cultivation [MSP^{+}15]. culture [HH10, LDVSGDR16]. cultures [MHC14, SM12]. culturomics [CMT19]. cumulative [BD12a, San12d, WG11]. currency [LH12]. Current [YLY^{+}14, CRAJdMACÁ15, FAI^{+}18, GTMRE^{+}16, Kaz15, KTT11, MSP^{+}15, SK18, TG16, ZSY14, ZWW^{+}15, ZWW^{+}16]. curricula [BCT19]. curve [ER19a, GSKM17, LW10]. curves [CV15, Li14, MdFdA^{+}14]. curvilinear [BHA15]. custom [TE18]. custom-made [TE18]. customer [Tsa11]. customized [LKS^{+}14]. cut [CÖT16a]. cutting [MBSB17]. cyber [AAS^{+}19, JC19]. cybermetric [OMOR13a]. cycle [Bou11, Jun12, KKS16, QJZ^{+}14, ZAJ19, vdPR18]. cycles [Fie15c, KG16, LWIB16, Ley15b]. Czech [PV15, VFA10, Van14]. Czekanowski [ST14a].

D [KWW15, DCM16, HZQ^{+}17, ZLLD19]. D-A-CH-L-L [KWW15]. D-index [DCM16]. dactylifera} [AAG14]. Daegu [KCP12]. Daiichi [KNK^{+}19]. dance [HH15c]. Dancing [YZ17]. dangers [ACD11]. Danish [Hen18, IL14a]. dark [OMLC15]. Darwin [MB14]. Data [ASW18, Com15, DLL^{+}16b, DLL^{+}16a, EBD15, GBM^{+}16, HSPY15, MRR17, RBC^{+}10, Siv16a, ZC16, AR18, ABMSSP16, ABSF^{+}19, AS18a, ÁBMB17, AYS14, Ano16b, Ano16d, Ano17c, BOS14, Bor16, Boy17a, CB15, CyPP12, CQB16, CNC18, DG16, De 17, DRS18, DCS12, Don17, DGF17, EDEH16, FS12, Fia11, FMM13a, FDVZ16, Fuk16, GSM^{+}16, GTC16, GGS17, Gom19, GKF17, HK19, HT11, HLE10, Har13d, HIG^{+}17, HBS^{+}19, HB17a, HB18a, Hau16, HWL11, HL13, HWS18, ISR11, qJnShPL17, JJS^{+}12, JS15, KZC16, KL17, KK18, KKOS19, KPRT16, LPMK17, Ley13b, LKR14, Li16, LPB14, LGD11, LSY11, LTK^{+}18, Liu19, LJMF15, LWT16, LSL15, MK18, MNdF16, MM17a, MM17b, MOO17, Mix18, MKYM^{+}17, dCCMAW16b, MS16b, Mou15a, NT17, Osw10, Pan14, PK14, Par14a, PW17, PHDC16, PRSB16, PKL^{+}16, Pra14a]. data [RC13a, RJ14, RNM18, SBA^{+}19, SGG^{+}14, SFR^{+}19, SYDW19, Sko14, SHK14, xShLY^{+}15, Tsa11, XYW^{+}17, YLSW16, Yu17, YHL^{+}18, Yur15, ZYS16, dPdCAdMC^{+}16, vLvWW16, SBSU15, HZ17, Ort18, ZLH^{+}15, dCCMAW16a]. data-different [Ano17c, GGS17]. Data-driven [EBD15]. Data-mining [Com15]. Database [CYW^{+}11, BMM14, BB15, CHWL12, Cav15b, CWH11, Cop19a, DGGBDG17, EDEH16, FMM13a, GY12, GK18, KT15, LVSL18, McC14, MS12, dANR15, PFDL17, PIB18, Rod16, SA11, SA12, SM15, TE18, VO17, kWhHRkS10, Ye14, ZWW^{+}18, ZCMVQS11, dART^{+}17]. databases [ATJ16, FMM15a, Gus19, Jac12, OMAT19, Pau10, Rod16, SSN19, yTwTlW19, ZYX^{+}14]. dataset [GT17, MXZ18, BH16a, BH16b]. datasets [BYY18, CKPY19]. date [AAG14]. David [MHFB17, Par14c]. dawn [Etz13a]. Day [BHDI18]. DBLP [Kim18, Kim19a, SHK14]. de-globalization [MHM^{+}12a]. DEA [HL13, ZK19, ZG17f]. dealt [GVS17]. death [BS13a]. debate [HRH10]. Debunking [Pol16a]. decade [BHPVdPMR18, DFS15, GKB^{+}19, HMCD^{+}19, KVC15, MAA17, SA11, SA12, The18b]. decade-project [GKB^{+}19]. Decades [Nar12, FPS14, Fra14, FZZ^{+}11, FZZ^{+}12a, FZZ^{+}12b, Li19, LTK^{+}19, RG15, TG18b, WHH^{+}18, ZAJ19]. decentralization [LOMLPA^{+}17]. decided [CKCK10]. decimal [Cam14, Egg14a]. decision [CdMCdMMdP17, LS17a, LD16, Ora17, vdBSS18]. decision-making [vdBSS18]. decisions [BD13, Fox17, GWG17, Med15, MG12, PR15, RSGFV18]. decline [HH17a, LvI10, LM10]. declining [Hor18, THFBdMA18]. decomposing [CRZGVQdMA16]. decomposition [HM18, IL14b, MYP19, XCS^{+}16]. deconcentration [MGM^{+}17]. Deconstructing [Hag10a]. deconstruction [LAL15]. decrease [Egg10c]. Decreasing [IS16]. deductive [SP12a]. Deep [HII^{+}18, SH19]. deepest [CÖT16a]. DeepPatent [LHCH18]. defenses [Pol16a]. define [AD14]. Defining [PHS12, BGJB16]. Definition [Cha19c, CPY13, Hei13]. definitions [HLSC18, Mar11]. degree [ADS11, CMdT14, JKPL18, LGZ^{+}13, RPK17, SS10a, WLF15]. Déjà [GREL14]. delay [LHW16, PS10, SRW^{+}15, ZFY^{+}17, ZYF^{+}17]. Delayed [EG18, BYY18, Li14, SSZL18]. deliberation [KP12b]. Delineating [CGV12, MNdF16]. Delineation [LZB10, MAA^{+}11, TT13, Zit15]. deliver [ADD17a]. delivery [KPY16]. DEM-related [PLW^{+}15]. demand [BR12]. Demise [KO18]. democracy [HRH10]. demographic [Ort15]. demographical [Han11]. demographics [ZG12a]. demonstrate [SLISC17]. demonstration [CL16, Ye14]. Dempster [WLPH14]. Denmark [SCGZSL^{+}13]. denominator [dSTL18]. densities [Sch12a]. dental [CVD14, KG10a, VVN16]. Dentistry [GdOdAG^{+}13, VEJC^{+}18b, VEJC^{+}18a, HG17]. Department [KSB11, PRRC15, Yur15]. departmental [BKL15, MKHB15a, MKHB15b, ZSCR^{+}18]. departments [ACRC17, CKT17, CAV^{+}19, GRSFV12a, Kaz14, Kaz15, KS17, Laz10, MHC14, PRA16a, WM19, WCK^{+}12]. depend [Bas10, Cop19a, MSP^{+}15]. dependence [Egg11c, LG10, SL16]. dependences [DCS12]. Dependencies [CRMPRS18, Hau16, KGSS16, LWIB16]. dependency [BMZ^{+}17, KB11a]. dependent [Pra19f, WFH^{+}16, ZMW^{+}18]. depicting [MMOMLC18b]. depth [GE11, IS16, PR10, ZG10]. Derek [Ano12c, Ano14, Ano17a, Dan19, KL16, CXZ19]. derivation [Egg11c]. derivative [ESB15, Fan18, KS18]. derived [Sch13a, WC18]. descendants [CLJH12]. describe [HR15, MB10a]. describes [AdAdAM10]. Describing [OA10a, BL17b, ZRY^{+}12]. description [Mag14b, Nii17, PHDC16]. descriptions [VH17]. descriptive [ABMRVZ14, ByLbH16, Liu17]. DESIDOC [LM19]. Design [EO14, CIL^{+}16, Hei19, HNG19, MG12, RNF19, SDS14b]. Designing [BBSS16a, BBSS16b, HNG19, MH16b, Pra16b, BBSS16c]. destination [Iwa17, PCR18]. destruction [Won13]. detailed [MSA13]. details [Hud16]. detect [MVS10]. detected [WJCC19]. Detecting [BK11, CH15, EBK16, HE16, HC14b, LWM^{+}15, PROGMA10, SK13, TL18, WLC17, YLSW16, YK12, GT12, hHC15, HC16c, LZFW15, RCN^{+}14, WWH^{+}17, ZHZY19, ZZW^{+}19b]. Detection [IMSK14, AA18, Boy17a, CLLH15, HD17, Jar16, KGY^{+}17, YK12, ZSC18]. determinant [TN19]. Determinants [AGLNRR14, FVVSGM^{+}18, Fuk19, KM17b, BBCP14, BS17, BC13b, CLW^{+}19, HWS18, JCK11, PNS^{+}10, PS16a, Pie18, PR14, Rha17, XA15]. determination [EO14]. determine [Ber18, GPN14, VEJC^{+}18a, YZB18]. determines [JC11]. determining [BGJB16, PR15, SA16, YB14]. develop [JDH12]. developed [EDEH16, KBL15, ZCW14]. developing [BH16b, CP14, GZ18, KMP^{+}11b, KTT11, PL18, Sko14, Sot12, WM17]. Development [Bos10, CFP14, GMJ^{+}17, Pou10, Ye14, YLL10, YLH^{+}17, ZLLL19, dSdSSB16, Bai18, Che12, CYH13, CdJD15, CH15, DMM13, Geo17, HHZ14, HMCL16, JH16, Kar12, KPRT16, KvES11, Lan13, Lee10b, LLL12, LWB16, LW10, LCLX16, cSL10, MGT14, MB19, MSB18, MK19, MRN14, QMSM^{+}19, QL12, RRBA10, RT17, SdJDD19, SHS15, SP12b, TS11b, US10, WT14, WG12, ZSC18, ZyZZ^{+}14, ZH17, dPSS18, RGdCMM17]. Developmental [PNVCB18, JKC15]. developments [APYS13, GTMRE^{+}16, HTHB11, HYC18, KC15, LGH^{+}14, SvLVA19]. device [CdJD15, LJ10, SdJDD19]. devices [FSOS12, SHR^{+}10]. devil [Zit12]. diabetes [SZMS18, SZAJS14]. diachronic [Col18, Ing12, SM16a, SMAABJ11, YS14]. diachronous [BL13, ZG17d]. diadromous [NBR^{+}11]. diaspora [Bas13]. Did [MM16, SBB16]. differ [BWD10, BWD12, BHJD12, HC14a, LT10a, VVN16]. Difference [YXW18, BSFCC15, FK16, MM16, Men18, WHW^{+}19, YY16, vWWtH14]. Differences [The18c, Bas13, Ben15, BG18, DVB14, GD16, HAA14, HFW^{+}14, HT14, HDC13, hHSL19, KFKS15, KR17b, LVGV^{+}11, Lee10b, LB12, LWL17, LIdMAM11, MRGT18, MR18a, NF13, PFL19, PYH16, PJY17, PS16b, SK14a, VTY17, WTM^{+}16, YZB18, Zha18, vAvdWvdB12, vdBS16, LNMQRR15]. Different [LL10, AS18a, Ano17c, BPGGdMA12, CMM17, CvLB10, Dor17, DGDG13, GD11, GGS17, GGG16a, HH19, HGH17b, HP10, IPIU13, JN15, JX13, KZC16, KWW15, LCS^{+}16, Mar11, MGC19, Mor19, MSC18, PRRC16, Puu10, PML14, QRJ^{+}17, Rya16, San12a, Sch15a, Sch13a, Sch16, Sot10, WvE13, Wu18, ZWZ^{+}19, ZWX22, dPdCAdMC^{+}16]. Differentiation [RRL16, Geo17, Li18]. differently [AF18]. difficult [GW10a, QZZ17]. Difficulty [Fox17, KBAK17]. diffuses [GZ14a]. diffusing [CMRC15]. Diffusion [DT16, ACMP13, FMPP10, Gál17, pGDTP12, GG12, GZ17, HL13, HLL14, Hu11, JC19, LJC^{+}15, LL13b, LLHN17, MGMY^{+}18, Oze12a, SHR^{+}10, SS15, WSC16, WG10, Wu13, YC10, YWY10, ZXZ17, ZLW19, ZY15]. Digital [CZV10, CRR14, KY17, SB19, ZVC11, BMZ^{+}17, Fan15b, GB17a, HRH10, Kim18, Kim19a, LHW12, MZ14, TCC17, WFG16, XCS^{+}16, ZZ11, ZWL^{+}18, BC17]. DII [kWhHRkS10]. dilemma [PG12, RJ14]. dimension [ADS12, BG12, GSKM17, Li15, SZZC18, Van10]. dimensional [HS16b, JC19, KdBBK15, LYGQ12, LP18b, Li19, NJ10, Pra14c, Pra17c, Pra18e, Sch12a]. Dimensionless [Pra18c]. dimensions [Bha16, BCC^{+}17, FFL16, GD17, JL14, MR18a, Pal15, Rod17, Wal16, Har19b, Bor18, HL18]. diode [hHC15]. direct [AR18, BL15, HSK18, MvdH13, SLISC17]. directed [ZZFD18a]. directional [GZ11]. dirty [FMM13a]. disambiguating [AOd15]. disambiguation [AT17, CLB13, GHvdB12, HYF^{+}17, HYYR14, Kim18, KK18, Kim19a, Kim19b, KKOS19, MRR17, PLT14, Sch16, SKCK14, TW10, WBH^{+}12, WD13, WLPH14, ZYX^{+}14, ZWL^{+}18]. disappearing [Fan12]. disaster [KNK^{+}19, MHKB16]. disasters [RF19]. Disciplinarity [Rod17, BTNS14]. Disciplinary [DAYY18, GD17, HT14, PYH16, BFGVV^{+}18, CC11b, DFG^{+}18, Ell18, Fie15b, Fie15a, Fie15c, GY12, Gau17, GD16, GTD14, HAA14, HA16, Li17, LWL17, LIdMAM11, MGMY^{+}18, MKP16, VTY17, VO17, YWS18, YYDH12, ZCW14, ZZL^{+}10, ZG17d, ZG17e, ZSCR^{+}18]. discipline [ADS17a, ADS17b, COS11a, CAGL15, CBKL13, CST11, GAPP18, GDP16, LG10, LSE^{+}18, NLCC17, SSS^{+}11, WDS16, WZ19b, Zhu17]. disciplines [ADD14b, AS18a, AC13, BHKP11, BKG16, BM12b, Cha16, CG18a, CGG19, Fan12, Fan13b, Hud16, Hud17, MMOMLC18b, Mor19, PHS12, PML14, QL12, SC10, Soo18, TCB16, WV13, WL14, WJD15, XG18, ZY15]. discontinued [PhD18]. Discontinuities [LWB18b]. Discouraging [PG14b]. discourse [CRLMRPA17, SMAABJ11, Sma10]. discover [Cam17, ILB11]. discoverer [MHFB17]. discoveries [LW15]. Discovering [BAB13, CGKB18, DCS12, HZ17, KLP17, Ma12, RPAMR19, VGPdlC^{+}17].

Discovery [BK10, WYAY12, WTM^{+}16, Arb11, Cha17b, CYH13, CRLMRPA17, De 17, KGL^{+}14, KGY^{+}17, Kos14, LP18b, Pra14a, VG14, WKHS19]. discrepancies [RGCM14]. discrete [BD12a]. discrimination [MM18, MM19, SH18]. discriminations [KK13]. discriminatory [OBG11]. Discussing [TBT19, vL12]. Discussion [Ano12a, Ano12b, HB15, EMSH16, KKCG18, CG15a]. Disease [HEH17, Kos14, CWJC14, DMM13, SZMS17, SHL15, TZ15]. Diseases [VACCAJ18, AP16, FR11, MK19]. Disentangling [CD16b]. disorders [GWBSVWB13]. Disparities [HJM^{+}13, PHBN^{+}15, TN19]. disparity [Ley18]. dispersed [Pir19]. dispersion [ADS16, CM18]. display [CWL10, HWLL14, JYW11, LWL17]. dispute [ZXM^{+}16]. Disruptive [BT19a]. disseminated [LHLH19]. Dissemination [PQG14, ESB15, FRdA16, Fuk16, LZR14, NLCC17, RPGM16]. dissertation [MS16b]. dissertations [AH11, Ban18, Hag10a, KKCG18, KHH18, KKE13, ZSY^{+}13]. dissertators [KH17]. Distance [GPN14, ACFL11, APT13, FZZ17, LDVSGD19, LXH^{+}18, PML^{+}17, SS16, vPD13]. distances [RGLE16]. distinction [SVS18, WWH^{+}17]. distinctions [Men18]. distinguish [Ley18]. Distinguishing [LY16b]. distortion [ADS10b]. distributed [ILP13, OROMAA16]. Distribution [Cam10, Shi11, ADS16, AdAdAM10, BS15a, BBP14, Bou11, DGDG11, Egg10b, Egg11a, wHwH16, KCT^{+}17, LXWC17, Lin11, LM13b, NQ14, OKK14, PYW18, WZS12, WLD^{+}14, WZ19b, YMSQ10, YL10, ZLF^{+}14, ZFY^{+}17, ZZW14, ZZW16, vZ13]. distributions [BCJ^{+}17, Brz15, Glä10, Glä13, RC13b, Sch13b, Sch15c, yTmShL16, cTnHwH17, YR10]. Divergence [LLW13]. diversification [ADD17b, ADD18b, ADD18a, CJW10, CSC12, CC12b, HMCL16, Luo12]. Diversified [FYC15, CLLL10, LHWW14]. Diversity [DKS18, Ley18, RM10, WJD15, YS14, AJSN18, CP12a, CKPY19, Gau17, GS12, GZJ^{+}15, IS16, KWM^{+}18, Ley15b, LWB18a, Lia11, LAdAMJ17, MKP16, PLWS14, aSTS17, SK11, Wu13, YST12, ZT18, ZSCR^{+}18, ZT19, ZRY^{+}12]. divide [ADR16b]. divisible [dS17b]. DMAIC [QJZ^{+}14]. Do [ADD17a, ATK17, BD16a, BDF^{+}17, CFG^{+}14, CT15a, CÖT15, CP16, CB11, FMM16, GRSFV19a, HC14a, HPKS18, LABL13, MBTKA14, MZ14, MTT15, Mix18, MDFGAM14, NCG^{+}19, OMMMTLC17, San18, SRW18, Soo17, The18d, TSG13, WM19, WPCG13, WS13a, YWL16, ZLTY18, vRW18, AD14, ADS17a, ADS17b, ALH15, AAV13, Bar17b, BLA16, BWD10, BWD12, BHJD12, BWdMA17, BCZ12, Chi14, COS11b, GE11, GYZ15, GWG17, Har19b, HV18b, IDKF17, Joh18, LVSL18, LNMQRR15, Lyk18, MS18a, Mar11, Med18, OM11, RGCM14, Saf13, SBB16, Sch13a, STCRPA18, TB19a, TG17, VVN16, WLY14, WRV14, YP19, ZW18a, dSBC17, BGJ^{+}16]. Dobránszki [BI18a, CF18]. Doctor [SDS19]. Doctoral [ZSY^{+}13, DY18, FCTV12, Hag10a, HB18b, HR11, RPSA17, SK17, SJOC18, Var12, ZTP18a, ZTP18b]. doctorate [Han11, HMCD^{+}19, RCdJ^{+}14]. doctorates [SH15a]. Document [Don17, Har13c, ACD14, CCM^{+}11, GT11, GT12, GT17, HYC18, KJES16, RP17b, Whi18, Xie15, YSM^{+}19]. documents [BHM16, pGDTP12, Glä12, HYF^{+}17, HB17a, HFL14, IBL13, Kra19, LHBC18, MVS10, MMOMLC18a, MGC19, RKZK18, TB19a, WPCG13, YPNS14, ZGY16]. Dodo [DF15]. Does [Ama18a, Bas10, BSMD11, BH17a, Kor18, KR17a, LT10a, LXH^{+}18, MSP^{+}15, Men18, Par15, PPI17, PMN16, SP14, SB17, SLISC17, SZZC18, The18e, TV17, TÜ10, WLM15, WHLP16, WG11, WHZ14, YY16, Yur16a, dS17a, BSvEK13, BSFCC15, BTL19, BT19b, DWGL16, Hag10a, HTHB11, JJS^{+}12, KKV^{+}13, LCZ17, Lin10, Lyk18, MB10a, MR18a, MPF18, NG16, PRA16a, Rig13, RJ14, TG18a, Zuc10, Shu17]. DOI [Ano18b, dCPF14, FMM15a, HPS19, XHA^{+}19, ZHL19, ZLL19]. DOI-to-DOI [HPS19]. doi.org [Glä18]. doing [dCPF14, GdA14, dAG13]. Dole [TR14, LM10, LM13a]. domain [ÁBDFB19, BS19, CNPG17, CZV10, CDM18, GQAM19, GN17, HWQ^{+}18, HQY^{+}18, JYM^{+}16, KW15, KZC16, LSS15, LSS16, LRY18, MYP19, NRAW17, NSKO15, SZ18, THFBdMA18, yT15, WKK16, WS13b, ZLH^{+}15, ZH17]. domain-oriented [LSS15, LSS16]. domains [CAS16, GRSFV14a, KMP11a, LAS14, LNMQRR15, ZZ11]. domestic [HWL11, PY19, PML14, SL17, TUCR15]. Dominance [Van10, GG19, IB15, Moe16b, Pei19]. dominated [uHBK19]. domination [WZ17]. Don’t [Mou15a]. door [SL13]. doors [RRLNAG15]. dormitory [vR17]. Dotter [GR16]. down [MAA17]. Download [DX17, kWhHRkS10]. downloading [KY17]. downloads [GBMA14, HBA19, JN11, SGG^{+}14, VTY17]. downs [SRP13]. Dr. [Zon19]. drain [MT12a, WMW^{+}13]. drawer [Pau10]. drawn [WT15]. driven [EBD15, KKLP18, KKL14, VPM16]. drivers [DR10b]. drives [HM15c]. drivetrain [NSKO15]. Driving [CHC13, CL17b]. dropout [XZFD19]. droughts [Bue15]. drug [HFC11, KPY16]. drugs [KB11b, KB12, WT14]. drugs.com [TKA17]. dual [LS17a]. due [dCCMAW16a, dCCMAW16b, Tol11]. Duplicate [LL13a]. durability [CvLvR11]. Duration [HS17]. Durfee [Pra10b]. during [ADD11b, CZW13, CJY^{+}15, GG14, IPIU13, IJF16, JPZ14, Lee10b, MGT14, Man15, NHY^{+}14, Sot10, SK14a, SZAJS14, TBW^{+}12, ULFRU^{+}14, WLY14, WCB^{+}15, WXZ^{+}16, YK14, ZLN^{+}13]. dusk [Etz13a]. Dutch [HM15a, Fra14, HM15b, vLvWW16]. dye [WLR^{+}14, ZZP^{+}14a]. dye-sensitized [WLR^{+}14, ZZP^{+}14a]. Dynamic [BZ17, KCK14, LG15, WLN^{+}14, ZY15, dSAEE15, AHP17, CLHH10, FFR^{+}17, Hsi11, HC16b, KO19, KJW^{+}17, Ley13b, LRY18, PTMT11, RC13b, SH15a, SvLVA19, SWH14a, SWH14b, SG16, WCL14, YPK13, ZG12a, ZGY16, ZLL^{+}17, ZG12b, ZZW19a, ZWL^{+}18, Zin16]. dynamically [BYY18]. Dynamics [RPGM10, WY19, AHUR11, ÁRS17, BGSvdB11, BvdB14, Cam12, DC14, FKM^{+}15, FR11, HMK^{+}12, HKWC15, KLM16, Kli16, KPSL12, Lam12, LMM15, LK17, LWIB16, LAHH15, LG16, LDG17, MT13b, MD18, RHGKD16, Sch17b, SHK14, TS11a, VLV14, WT14, YB14, ZG13b].

E-commerce [YLL15a, CYH13, Tsa15]. e-government [KMP^{+}11b, SZZC18]. E-learning [THFBdMA18, CL11]. e-mail [SRW18]. e-research [JDLIV14]. each [CR14, OM11]. earlier [LS17b, vL12]. Early [CLSW19, HZL^{+}17, The18f, WT15, BDE11, BS16, CGZ10, CMT18, DRCG17, Lee19a, LD16, MHTB17, TK17, The18e]. earn [San18]. Earth [Mik10, JS15, NHY^{+}14]. earthquake [Ho13a, LZH^{+}12, LZH^{+}13]. earthworm [XZZ15]. ease [Arb11, Har16a, Har16c]. East [GNS^{+}15, KHK13, Moe16b, TA11, WCK^{+}12]. Eastern [LGR17, SZAJS14, Cav16, GVS17, JMM19, KBL15, Paj15, TA14b]. eating [GWBSVWB13]. Ebola [Oli15a]. Ecodesign [LJJ^{+}16]. ecological [CBWJ18, NTM^{+}18, PNS^{+}10, PPI17, XM13]. ecologists [Med18, PML^{+}17, RRL16]. ecology [ASPF^{+}16, BF17, Fox17, JKJL14, Leb12, LABL13, Med18, PLA10, PAL13, PS10, PB12, dCdSNB15]. econometric [KG16]. Economic [Gan12, vB13, ACFL11, BB10, HT11, Hud17, ILP13, ILBG14, KZ13, KKBW17, KKT^{+}18, LRC19, Lee10a, LLCL11, LKW^{+}16, MMAHS10, NA18, OMR14, PR15, SY16b, SX16, TB19b]. Economical [MC15]. Economics [BMTA15, GRSS16, HM15a, HM15b, Hen18, LT10a, McC14, ZWW^{+}18, ACRC17, BS15c, BW19, CGZ10, CKT17, CÖT16b, Etz13b, FMU16, GRSFV12a, GW17, GZJ^{+}15, HM15c, Kor19, Kos18b, KG16, KO18, KRR14, KKE13, MTU17, Mix18, NvLvR10, ÖS17, Osw10, PRRC15, PPI17, SND19, SL14, Var11, ZL18b, ZZY19, ZYT^{+}16, BOS14, CMT18, BW20]. economies [AATBPAB15a, CP14, KTT11, PN15, Sch14c, Sko14, WG12, WF17, ZP15]. Economists [WP18, BD10a, CMUdF15, Fra14, GVS17, HM15c, Kor18, Pra10a]. economy [CZ18, Etz13b, GG19, KG16, LZ14, NP11, NT17, dART^{+}17]. ecosystem [LK17, LH14, ZG17a]. Ecosystems [Ioa06]. ecstasy [GGG16b]. Edge [Pac19, Fie15a, KGY^{+}17, MBSB17]. Edinburgh [KCP12]. Edited [OE15, TSRGCCJC14]. Edition [BD10b]. Editor [BI18a, BD10b, Egg10e, GRSFV15, GRSFV16b, Pra16b, Pra17a, Pra18c, Pra19c, Pra19d, SGSS17, Sch10b, THB18, BBSS16c, Bor15b, Glä14, Lor10, Pra11c, Ric17]. Editor-in-Chief [Glä14]. Editorial [Ano16a, Ano18b, BEMS17, Bra10, BSG10, Bra12b, Gar14, GSB18, KP12a, MHM12b, PHV17, RSGFV18, SRW^{+}15, TO18, XWL19, CIL^{+}16, Fox17, GCGP10, HE16, Kim10, LHW16, MRLW15, MHM^{+}13, Mou15b, MFF^{+}16, Ora17, PR15, YP19, ZL18b, vLCCMV13]. editorials [WvBvE11, Waa13]. editormetrics [MPF18]. editors [BPVM11, YP19]. editorship [BB10]. Edler [dCPF14]. eds [Pen19]. educated [Yur18b]. Education [BSFW10, CD17, PQG14, PSY^{+}19, Saf19, ACD11, ADD14b, ADR16b, CGK^{+}14, CC12a, CDdS^{+}12, DVB14, DVB15, FWFM18, GGG^{+}12, HIC12, KY17, Kwi18, LKSK15, LAS14, LVHS^{+}15, MSdBC16, Med15, NSMMDB19, PSZ15, Pin15, PFPCM^{+}19, QJZ^{+}14, SB19, VZAMG19, VPM16, VAJCC17, Vin12b, WDP11]. Educational [AZSA14, AZSA16, BKY^{+}15, Moo15, CT15b, KKCG18, LdSdFFNM17, RPSA17, SV19, TCT^{+}13]. EEE [Pra11a]. EERQI [Moo15]. Effect [CCM^{+}11, EES13, MSdBC16, ADD16, ADD18a, AvLS14, BHDI18, BND11, BM12a, Bor12, Boy17a, CC14, Cam14, CLJH12, CR14, CGSS13, CdMCdMMdP17, DEC15, Egg14a, FMM13a, GRSFV18, GLD16, GK18, Gom19, GMSZ18, Hir10, HNG19, Ibr18, IT11, LS17a, LM10, LM13a, LABL13, MR15, MB16b, NSKO15, QZL^{+}17, RG15, RCJ18, RP17b, RMH14, SGSS17, SW19a, SWH14a, SWH14b, SI17, SL16, TBT19, Van14, Var12, WZ19a, XZFD19, YGW^{+}15, ZCL14, ZCL15, ZW17a, dSD18a, vRvLV11]. Effective [PHL17, Ben11, YLL10, ZS18]. Effectiveness [Cle16, IT11, MPY^{+}13, RMCM13, TCT^{+}13, YSM^{+}19, ZWZ^{+}19, ZWX22]. Effects [Bra12a, GRBBS17, GP18a, IF13, JJR10, PJB^{+}12, PROG19, Puu10, ADD11b, ADD18b, AML17, BHA15, BS15c, Bou14b, BS13b, CGK^{+}14, CC10a, CJW10, CSC12, CHC13, CU16, CvLB10, CvLvR11, CR18, DC15a, ET15, HAJ12, HBT16, JCK11, KPS12, LKW^{+}16, MBSB17, MGB16, MSH16, OCCSM11, PS13, QDK19, Saf19, SFM16, SST^{+}16, WHL^{+}15, WW11, YC10, You14, ZCZ^{+}16]. Efficiency [GMM16, RBC^{+}10, WDP11, Bas14, Cle16, CNC18, GZ14b, HT11, LCIADG19, Pra10d, QDK19, Rha17, SR16, UHAR12, ZK19, ZG17f]. Efficient [CLB13, BCJ^{+}17]. efficiently [BWdMA17]. effort [Cha16, LAL15]. egg [CMO11]. ego [SC13, ZCZ^{+}16]. ego-network [ZCZ^{+}16]. Egyptian [EDEH16, Eld19]. Eigenfactor^{T M} [YAC10]. eight [AH11, KEP^{+}18, LL16]. either [BWD10]. election [Par14b]. electric [HSLP14, YSY^{+}13]. electrical [SFNO12]. electrochemical [WLH^{+}17]. electrolyte [CYK^{+}11]. electromagnetic [ARK^{+}15]. Electromobility [Zha18]. electron [FSOS12]. Electronic [Glä18, LHC16, Ban18, JYW11, SD13, SFNO12, TÜ10]. electronic-paper [JYW11]. electronics [KCK14, Lee10b]. element [JS15]. elements [LY12, Li14]. Elenara [dCPF14]. Elfreda [GTAG18]. Elias [Par14c]. elite [FMM14, PLA10, PAL13, Saa10, Vin17, Vin19, WMW^{+}13, YZ17, Yur18b]. elitism [PG14a]. else [Har16b]. Elsevier [SGY15, SE18]. EM-index [BT18b, BT17, BT18a, BT18c]. Embedded [KGG15, WZ19a, vWBS^{+}16]. embeddedness [LLHN17]. embedding [KWM^{+}18, LHCH18, ZZL19]. embraced [WHC^{+}13]. emerged [Bru10]. Emergence [CRLMRPA17, CQB16, BFS17, BSPL19, CNPG18, FVVSGM^{+}18, KN15, MB13, MLT^{+}14, Moe16b, SYP10, SKY17]. emergency [FCFG17]. Emergent [CR18, CNPG17, Kli16, LRA14, MGMW14, SHR^{+}10]. Emerging [BSK15, GAPP18, HYC18, JYW11, STCRPA18, TZ15, US10, AUS12, AATBPAB15a, ALYZ15, AJCACRdMA16, APYS13, AYS^{+}13, ÁRM13, BGM17, BRS^{+}16, CMRC15, CLMN19, Cha19a, CWL10, CRMPRS18, CS11b, CDCK13, ÉMS^{+}13, GT12, GALR16, GWB11, HSPY15, IMSK14, JKJL14, KO19, Kha13a, Kha13b, KW15, KC15, KdBBK15, KK17, LQW17, MCvFP16, OHT10, OH19, PN15, RNM18, SBSR19, SJ10, SD13, SZD16, SÁV18, SG16, WP17, WG12, WF17, ZZPG14, ZRL18, ZG11, ZLLD19, HZL^{+}17]. EMF [ARK^{+}15]. eminence [Kor18]. emission [CWL10]. emitting [hHC15]. emphasis [BN10, PH14]. Empirical [Cam11, HFL14, ZW17a, Avk13, BYY17, Che12, CKCK10, CG17, FMU16, FWFM18, Fuk14, HB17a, HB18a, Jun12, KPRT16, LF14b, MRR17, Nii17, Par15, RD13, SBT18, SL12a, Soo14a, VPM16, Wad16, Wad17, WvEvL^{+}11b, WvE13, WLZ^{+}15, YWL16, YQX10, ZT18, ZT19, ZM16, LCZ17]. empirics [JPT13]. employee [CC10a]. empowering [RG18]. enabled [KPY16]. encompassing [PKR15]. encountered [CGK^{+}14]. encouraging [JX13]. encyclicals [RAM18]. Encyclopedia [Tom18]. encyclopedias [Tom18]. end [Bar11, Glä13, HRC13, LM10, LM13a]. ending [BA15]. endocrinology [LPZ17]. endodontics [AMK13]. endogenous [YP19]. endure [Kor18]. enemy [KFB18]. Energy [CNC18, KJ14, Pra11a, Pra12b, FLZ17, FAI^{+}18, GBMB10, HH15a, HYS18, ILGZ^{+}14, JK19, LO12, LG15, LLW13, MSdBC16, Pra12d, Pri16a, RFGBMA13, SPB18, SCGZSL^{+}13, WLH^{+}17, YG18, ZYG15, ZKC^{+}16, ZG17c]. Enfermeria [Kra19]. engagement [DTM^{+}13, PMJF19, WHLP16, WKK16]. engine [MMOMALC16, vdBBdK16]. Engineering [SH15c, WMT^{+}12, ATK17, Fer14, FLH14, Gar15, GF17, GWA14, HAL11, Kaz14, Kaz15, KPJT14, KKV^{+}13, cSL10, MGB16, MPM18, PPK^{+}16, PLG19, RAA18, RPNC13, SH19, SBT18, SFNO12, Sin18, TAB13, TABA16, ULFRU^{+}14, VCC12, Zha14, ZW17a, dJC15, dSdSSB16]. engineers [Soo11a]. engines [Gus19]. England [DD18, WL18]. English [BL11b, GAVZAB12, LRZ13, LHTL18, LNMQRR15, MASM14, MASM16, SW19b]. English-medium [MASM14, MASM16]. enhanced [Ano18a, CFM18, SH19, WS11]. enhancement [GGP14]. enhances [Sot10]. Enhancing [SBD^{+}19, DCGZ^{+}12, Pra10b]. enlarged [KKBW17]. enology [AATBPAB15a]. enough [Liu19, SYDW19]. enquiry [MGMW14]. Enriching [CRLMRPA10]. enrollment [BR12]. ensemble [WLZ^{+}19]. enter [LM15]. enterprise [LCWY12]. enterprise-university [LCWY12]. enterprises [ZLLL19]. entire [GSKM17]. entities [Ken18, KKL14, SZZC18]. entity [Saf13]. Entrepreneurial [SMCC18, YC10, ZG17a]. entrepreneurship [KM18b, LLYC14, MLT^{+}14, SMF18, Sku19, Tei11, ZSY14]. Entropy [Pra11a, BFGVV^{+}18, CC12b, LAL15, LZFW15, Par14b, Pra12d, ZQH^{+}17]. entropy-based [CC12b, LZFW15, ZQH^{+}17]. Entry [AYS^{+}13]. Envelopment [RBC^{+}10, HT11, HL13, PRSB16]. environment [AGLNRR14, HJL18, RCCM14, WRM17, ZZW19a]. Environmental [MWH14, Wil15, JKJL14, Lop10, MH16a, PLA10, PAL13, PPI17, RHMH17, RF19, SL14]. epigenetic [OGRMOP19]. epistemic [APT13]. epistemological [OBG11]. eponym [ARE^{+}18]. eponyms [Cab14]. Equal [TBW^{+}12, BT19b, CÖT16b]. equality [Yur18a]. equally [Hag10b, PPI17]. equals [SVCFI14]. equations [GP13]. equitably [Hag10b]. equivalence [TW10]. equivalent [AD13, YK15]. era [Cho12, VB12]. ergonomics [LMR16]. erosion [ZDZ^{+}15]. Erratum [AF15b, ADS17a, BI12a, BBSS16b, BM13a, Doc11a, FE16a, GZGAC17, Har16a, HM15a, HRB^{+}14, JG14, Kha13a, KPS12, LV12, LSS16, LXL16, LNK^{+}14a, MASM16, MM17a, dCCMAW16a, SWH14b, Wad17, WLMF15a, ZZW16]. error [ET15, Wu18]. Errors [FMM15a, EdS19, LZR14, LD16, XHA^{+}19, ZHL19]. Erythronium [CRLMRPA17]. Erythronium/Vanadium [CRLMRPA17]. ESA [EBR16]. ESCI [HZL^{+}17]. ESI [CPF13, WFZD19a]. essence [Vin10a]. Essential [FCWH11, Har15b, CWH11, ZWW^{+}18]. Essentials [TKA17]. ESSS [GGH^{+}10, GGG^{+}11, GGG^{+}12]. established [BKSS15, HR11]. Establishing [DCY^{+}17, SFM16, RC13b]. Establishment [ZZL^{+}10]. esteem [BCT19]. Estimate [LL12, De 13, Wra10]. Estimating [SOBM16, vdBBdK16, MD12, OMAMMLC15]. estimation [Cha17a]. estimators [VHH16]. ethanol [dSFSF15]. ethical [JX13, YB14]. ethics [BM12b, LP18a]. ethnic [LKW^{+}16]. ethnicity [HJM^{+}13, LRWS16, RLW14]. ethnicized [DFS14]. ethno [Han15]. ethnobiology [CSS^{+}16]. ethnobotany [RMA12]. ethnographic [KHR^{+}19]. ETRI [SWH14b, SWH14a]. ETSI [XWL19]. Etzkowitz [LM13a]. EU [BM11, GRSFV12a, HHK^{+}12, HLSW18, Mor19, SCGZSL^{+}13, WWP17]. EU27 [TFJD14]. Eugene [BHL18, Ano17b, BI18b, Bha18, BHH18, Che18b, GSB18, GA18, Jac18, LWB18a, Pra18b, RH18]. Europe [ACFL11, ACF^{+}17, BKZ^{+}16, BB19, BS11, Bre10, CGZ10, CKT17, EBR16, FFL16, GBM^{+}16, KKBW17, LGR17, LAS14, MK19, Paj15, PCR18, UMK14]. European [ACORC10, ACT18, BS17, CMM17, DSM11, EBR16, FM12, GTMRE^{+}16, GTMRE^{+}19, GGH^{+}10, GGG^{+}11, GVS17, HV18a, HRC13, JMM18, JMM19, Kaz14, KHK13, KBL15, KEP^{+}18, LV11, LVHS^{+}15, MM16, MBA13, Moo15, OA10c, OPGW^{+}13, PSZ15, PQG14, Pir19, SK16, TA11, TA14b, WDP11]. eutrophic [GCLcG15]. eutrophication [CH12, HW12, YJ11]. evaluate [AC13, BT17, BT18a, BT18b, BT18c, BM14b, CMdT14, FMS17, FMPP10, MH16b, MOO17, NJ10, Pra10d, RNF19, Ric15, VB12, YLL^{+}15c, YAC10, ZMW^{+}18]. evaluated [ADV10, LCFC14]. Evaluating [AAH10, ARK^{+}15, AD11a, DBO^{+}18, FMM13b, HDW^{+}15, KDFL14, Kim18, KTRP17, MCL13, Pra12c, SK16, SZAS16, ST14b, Zha17, Abb11, BW10, FCFG17, FM11d, FGMM12, JBMR11, KCP11, PRRC16, Pra11c, Pra13, QDK19, WST14, XWL19, YSM^{+}19, Glä13, Kim19a]. Evaluation [uARA19, AA19, BKZ^{+}16, BB15, LdZwC^{+}17, RAA18, RG12, ZW19, ÁBV^{+}14, AKB^{+}10, Ano16d, ASW18, BD16a, BMM14, BBDS^{+}14, BBSS16a, BBSS16b, BBSS16c, Ben12, BGJ^{+}16, Bha11, BM19, BDE11, Bou14a, CCLL14, CFdC^{+}14, DQ11, DGWZ13, DB19, ES16b, FCCMTRVR18, FLM16, pGSyW^{+}19, GQAM19, GTMRE^{+}19, GNHT18, HGH17a, HH15b, HAJ12, ISR11, KT15, Kaz14, Kim19b, KKL14, KPS12, KR17a, KKE13, LMKG19, LR12b, Ley13a, LYGQ12, LAW14, MWDC12, MLOY18, MB16a, MBSB17, MM17b, MZE19, NZL^{+}19, OCCSM11, Paj15, PYK13, PYL16, PP18, Pra10c, Pra14c, Pra16b, Pra17c, PKSG12, RWG^{+}15, Sch17b, SLISC17, xShLY^{+}15, Suo14, The12, VFA10, WV13, WMT^{+}12, Wan13, Yan14, ZW17b, ZW17c, dZLwC^{+}15, MM17a]. Evaluations [Eto13, BD10b, BG18]. Evaluative [Pen19, BH17a]. evaluators [CFP14, RGLE16]. evaluees [RGLE16]. Even [JL19, JL18a]. Event [GGG^{+}12, Ort18, Ano15]. events [BFM^{+}14, JK10a]. ever [ADD10, GVS17]. ever-increasing [GVS17]. Evidence [MT15, WOW10, ADS17a, ADS17b, ADD19b, AAB^{+}13, AZSA14, AZSA16, BHA15, BCT19, BGAAM15, BM11, Brz15, CSC13, CRMdMA15, CK14, CB11, DRMMC19, DRS14, FWFM18, FDVZ16, Fuk19, GRBBS17, GP18a, HV18a, HWLL14, IS16, JC12, qJnShPL17, KGG15, KEP^{+}18, LY16a, LCZ17, LJ10, LG15, LSR13, MM18, MM19, PNS^{+}10, Rha17, SL14, Soo14a, SK12, SC13, TK16, TN19, VG17, Wad16, Wad17, Wad18, Wal16, WHL^{+}15, Wra16a, Wra16b, YG18, YGW^{+}15, ZG17b, ZW17a, ZLH^{+}16, Zin16, ZXT^{+}19, ÁBDFB19, Bai18, MR15, MZ14, MS13, PIB18, PL17, SZZC18, TKA17, WRV14]. evidenced [Hur17]. evidences [AMK13, ZG13b, dJC15, GYZ15, ZG17a]. evidential [LdZwC^{+}17, dZLwC^{+}15]. Evolution [BSKB17, CHL15, FM17, HEH18, Hu11, IPIU13, KTLD16, TA17, YLJ^{+}17, BSG17, BKG16, CV15, CVC^{+}15, CZV10, CAGL15, CW17, CU16, CB16, Coc18, CZ18, DVMS17, FMP17b, Fox17, FM11c, Geo17, GHA^{+}16, GZJ^{+}15, HD17, Hou17, wH15, HZQ^{+}17, KM15a, KLCS14, Ley13b, Li15, LL15, LQW17, LHM^{+}11, Lja16, LG15, LX15, LYLD15, LRA14, LLW13, Ma12, MHKB16, MAGSTRC15, Oli15b, OMLC14, QZZD18, RBBG18, SdJDD19, SFBS17, SB15, SIR^{+}14, SJ19, SHK14, SZ18, SYLC17, TM12, TTC17, UCH19, WCL14, WSH16, WLH^{+}17, XBD^{+}18, Yoo15, ZGY16, ZLL^{+}17, ZY19, ZyZZ^{+}11, ZGL^{+}17, LX19, vdPR18]. Evolutionary [AHUR11, ÁRM13, BFM^{+}14, HKWC15, Li17, RSGFV14, ZLYF14, ZYSS14, BSPL19, BS19, CPRSFVG19, Cha17a, CH13b, Coc18, HY19, LZC17, LABL13, LJJ^{+}16, NRAW17, ZCZ^{+}16, ZZ16, ZZP^{+}14b]. evolutions [RPP18]. Evolving [AP16, BMM17, CBF13, Pri16a, CF14, Fan15b, SH15a, TCR10, YK11, ZYS16, ZB12]. ex [CA19]. ex-post [CA19]. exact [BBL17a, YGD17]. Examination [Lun19, Bre10, FK16, Nii17, Oli15b, PW17, TA14b, YWS18, ZY15]. examine [CS11a, CL11, CLW^{+}19]. Examiner [ACT18, BN14, Wad18, YK14, YK15]. examiners [Bre13, PJY17]. Examining [BHB13, Cha18c, HW10, JC19, KK19, SvLVA19, WZCC19, WZ19a, CRMPRS18, DFS14, Lia11]. example [BB15, GRSFV14a, Kol12, PB12]. examples [Hei13, Mus12, Whi15]. Excellence [MSH16, MKHB15a, BPGGdMA12, BR11, BCHH17, BWdMA17, DTM^{+}13, DH13b, DMV10, EGUB12, HIC12, MKHB13a, MKHB13b, MKHB15b, TW18, YGD17, Bor16]. exchange [CYK^{+}11, GD16]. exchanges [BPHL16]. execution [CIL^{+}16]. exegesis [Pra12d]. exemplars [MLT^{+}15]. Exemplary [Kra10]. exercise [AD16, BD16a, MWH14, BGJ^{+}16]. Exercises [CLD13, ADS10b, ADD11b, ADD11a, ACD13, ADD14a, SBT18]. Exergy [Pra11a, NPP^{+}12, Pra12d]. exhaustive [HIG^{+}17, HC19]. exist [HTHB11]. existing [AAV13, PLG19]. expand [GTC16, RCN^{+}14]. Expanded [FLH14, Ho13b, IH14, LLG14, TFH14, IFH15, VASNU^{+}19, WF18, FH16, MH16a]. Expanded-based [VASNU^{+}19]. expansion [GZGAC16, GZGAC17, dWZD14]. expectancy [BL13]. expectations [uHBLKH19, PD10]. Expected [Pra19b, BG18, Pra12d, RNB19]. expenditures [DR10a, PS13]. expensive [Hal14]. experience [Bou14a, HJL18, HIC12, VPM16]. Experiences [ZCKZ16, TCB16, kWhHRkS10, KR17a]. experiment [Sni16, ZS18]. Experimental [CA12, MBSB17, ZGY16]. Experimentally [uHBLKH19]. Experimenting [Cab13]. experiments [BL17b, PRSB16, JNA18]. expert [BH15, KR17a, ZLW16]. expert-based [KR17a]. expertise [BAB13, Lyk18]. experts [LZB10, LXDL13, LdZwC^{+}17, Moe16a]. explain [SS10c]. explained [PS16b]. Explaining [BCHH17]. explanation [BKRG13, CPRSFVG19, KK13, SM17]. explicit [Har13a]. Explicitly [Hsi11]. Exploitation [FLB19, CC13]. exploiting [CGKB18, QA19]. Exploration [Bel13, Hou17, Rai19, ACF^{+}17, BFMRM19, CFG^{+}14, JDG14, NvLvR10, PR14, WBX18, XYHD18, YS13]. explorative [LPMK17]. Exploratory [DGF17, KD18, MDFGAM14, WY19, XDB^{+}19]. Explore [CSC14, Che11, CC12b, CyPP12, CU16, LLC^{+}17, Li15, QZZ17, UCH19]. explored [PKL^{+}16]. Exploring [Cha14, hCcTmWH15, CAGL15, CST11, FD14, GGW^{+}13, HLL14, HSWC13, HCLC14, HC19, KJW^{+}17, KVC15, LP12, LS17a, LQW17, LC18, LSE^{+}18, ML16, MVS10, Moo15, MKF14, SND19, VRF12, XCS^{+}16, ZNB^{+}17, dSTL18, BTNS14, CYH13, GLÚGML16, ILB13, JCK11, LML11, LAHH15, dNMVQL16, Pin15, Saa10, SDT15, WF17, CLMN19, MDG10, RGCM14]. exponent [cTnHwH17]. exponential [AdAdAM10]. exportation [Rod17]. Expression [Glä18, ISR11]. extended [KS18, PKL^{+}16]. Extending [LAW14, Sko14, HWQ^{+}18]. extension [BM15, ISR11]. extensions [GQAM19, Lee10a, RAA18]. extent [ADD17b, SL14]. external [Bre13, BN14, CFG^{+}14, CLJH12, CHC13, MGB16, SPdSM16]. Extracting [Cab14, FMP17a, YLH18]. extraction [NJM18, VBG^{+}17, XLZ^{+}18]. extraneous [Ora17]. extreme [Bue15, Mag14a]. eyes [OMAT19, dSM17].

F [Par14c, Pen19]. F1000 [MT13a]. fabric [TW16]. Facebook [LdSdFFNM17]. faceted [DLMX15, PPM^{+}17]. facets [NSC13]. facilitate [WFH^{+}16]. Facilitating [LP18b]. facilitators [CDCK13]. facilities [Hal13, Hal14, SSN19]. facilitymetrics [Hal13]. facing [BGBS18]. fact [AM18]. Factor [ATM16, KCP11, ADD10, ABMSSP16, ASPF^{+}16, AA16, BL10, Ben12, BM12a, Bro13, BCZ12, CCM^{+}11, DGDG13, Egg10b, Egg11a, Har12, wHwH16, Ing12, Jac12, Ley12, LBA19, MRS^{+}16, APFR^{+}13, Ora17, PKR15, PA12, PMN16, Rou12b, Saf13, VEJC^{+}18b, Van12, VSS12, XLR15, YWS18, YLSW16, YYL10, Yur18a, Zit11, Zit12, BND11, EMH^{+}10, GAVZAB12, KM17a, KR17a, Vin12a]. Factors [ACFL11, ACMP13, CGC18, GNHT18, MOA16a, PR15, SFR^{+}19, TAA16, Bal12, BK10, Cam10, CC11a, Cam11, Cam14, CHC17, CHC13, CRMdMA15, DdS19b, DX17, ES16a, FSSPG^{+}15, Gan12, GRG12, Hen18, IMHG12, IMH13, qJnShPL17, Kos14, Lee19b, LWL17, LABL13, MBSB17, MD12, Pac19, PML^{+}17, PG12, Smi12, VHG^{+}15, WYAY12, XGL^{+}19, ZNB^{+}17, Zhu17, Zyc10, vWWtH14, BW10, Bor17, Cam12, Ste17, vL12, Egg14a]. facts [GM13]. faculties [BDE11]. Faculty [OCJB15, SRF16, ADR14b, ALH15, CSO17, CR18, FLB11, Kaz15, Lun19, MCL13, OBG11, SIS17, SDS19, The18c, WM19, ZYS16]. fading [CS19]. fads [YWG14]. fails [Mue18]. failure [ABL17]. Fairly [LF12a]. Fake [AM18, LL13a]. fall [HH17a]. fallen [Pan14]. falling [ÖS17, RY14]. false [Pen19, Pra14a]. Fame [CMT19, MRLW15]. families [DMM17, Mar11, NF13]. family [CIK^{+}18, HLSC18, LL12, NSKO15]. Fang [PT17]. far [BWD12, JKMS17, SS16]. Farewell [Gar14]. farm [KKBW17]. farming [AATBPAB15b, dSNV18]. fast [Kim19b, WRV14]. fate [CGG^{+}17]. FDH [GRSS16]. feasibility [MVS10]. Feature [LTGH15, GZGAC16, GZGAC17, SH19]. featured [WLM15]. Features [PT17, FA10, GYZ15, LT10b, TMLB16, WYY11, ZG17a]. fecundity [HB18b]. federal [CLO18]. feedback [MR15]. feelings [DLM15]. fees [CFG^{+}14, GRSFVCP19]. fellows [ZTP18a, ZTP18b]. fellowships [RGdCMM17]. Female [LM11, BL11b, CVC^{+}15, DH13a, Keg15, MB10b, SMF18]. fences [Mix18]. Fermilab [PRSB16]. Few [KCU19, CB18, FR11, LLH19]. fidelity [Car16]. Field [ACP12, Bor18, HB18a, HL18, TSMTDLCH11, ADV11, ADD17b, ACMP13, ACF^{+}17, uARA19, AL12, AW11, AA16, BH10, BK10, BFS17, BK15, BM13a, BM13b, BAC13, BTL19, BM11, BNV11, CWL10, CHY16, DL16b, FRdA16, FM12, FR11, FI16, GZ14a, GYZ15, GLM11, HH15a, Ho14, HH15c, HDC13, HC14b, HCLC14, hHC15, HH18, JYW11, JKJL14, JOGC17, KY17, KB11b, KPJT14, KK17, KJ13, KJ14, LZB10, Leb12, LC18, LHM^{+}11, LHW12, LX15, LJJ^{+}16, Lyk18, MLC14, MB19, MLT^{+}14, ML10, MÁB18, PNVCB18, PPK^{+}16, QZZ17, RBF^{+}10, RPGM10, RBBG18, SZZC18, SHK14, Tei11, TO18, The18d, The19b, TBMM18, VEJC^{+}18b, VEJC^{+}18a, WYvE11, WKHS19, XDB^{+}19, YLL15b, YP19, ZLYF14, ZCL14, ZKC^{+}16, ZG17a, ZG17c, ZWW^{+}18, ZT18, ZSC18, Zha10, ZS11, ZL17, ZT19]. Field- [HB18a]. field-normalized [BTL19, WYvE11]. field-standardized [ADV11]. fields [ARK^{+}15, AKB12, ACORC11, AH11, BCML19, BAB13, BFMRM19, BBDS^{+}14, BBJS16, BD12b, CSS^{+}16, Cha18b, Coc18, CRLMRPA10, DLL^{+}17, DGDG13, DGDGSV15, GRSFV^{+}12b, GALR16, GGG16a, GF11, Han11, HG17, HCS^{+}15, ILGZ^{+}14, JN15, JKPL18, JvGH10, JKJL14, JK10b, KM15a, KLP17, KCU19, KR17b, Lor14, NH14, OHT10, APFR^{+}13, PRRC16, PSY^{+}19, RGCM14, RD13, The17a, The18d, Var11, VT10, VLV14, WNS13, WP17, YS14, Yur18a, Zit15, dPdCAdMC^{+}16]. Fifteen [KGB^{+}18]. fifty [MGT14]. file [Pau10]. file-drawer [Pau10]. files [CC16]. filing [HV18a]. filings [Hei19]. Filling [TSRGG17]. film [HWLL14, YCK11]. filter [YSY^{+}13]. Finance [Avk13, CdMCdMMdP17, MZ14, VT10, FUR10]. Financial [HF19, Dem18, HR11, HC15c, HC17, KM17b]. financing [CFM15]. finches [MB14]. find [The18e]. Finding [LXDL13, WZ19b, KBAK17]. findings [Hei13, The17b]. finds [TK17]. fine [UBTS16]. fine-grained [UBTS16]. fingerprints [HYF^{+}17, TW10]. fingerprints-based [HYF^{+}17]. Finnish [Puu10]. firm [CC10a, CC12b, Hun12, KLCS14, LPMK17, LCZ17]. firms [AYS^{+}13, GRBBS17, JYW11, LZ14, LYWSV13, MDDG17, PS13, WPCG13, WRV14, WY12]. First [De 16a, Bar11, BS13a, BH16c, Bor18, Cam18, CÖT16a, CMO11, CMM17, CA12, GRSFVMB12, GBM^{+}16, GA18, GZ18, GKB^{+}19, GNHT18, Hal13, Hei19, HL18, HL17, LS17b, Li14, LHG16b, MM14c, MBSB17, NvLvR10, SCGZR16, Sch14b, SL10, The19a, TW16]. first- [MBSB17]. first-quartile [LHG16b]. fish [KJES16, NBR^{+}11, dSNV18]. Fisheries [CP12b]. fit [RPP18]. fitness [Ke13]. fits [BVZV16]. fitted [vZ13]. five [Bar11, Cam11, CAS16, DMB17, GTMRE^{+}16, Moe17, PSB^{+}17, RG15, SMAABJ11, WF18, WOW13, ZCKZ16]. five-year [Cam11, DMB17]. Fixed [YLC18, DR10a, SvLVA19]. Fixed-income [YLC18]. flag [HH19]. Flanders [VDV16]. flash [Li14]. flavors [BPGGdMA12]. flawed [Pat18, Pol16b, Sch18a, Sch18b]. Flesch [Har16a, Har16c]. flexible [GG15b]. flow [AHP17, KJES16, SZZC18, yT15]. flows [Ama16, CKB^{+}14, HH13, HH15a, HSAK18, Iwa17]. foci [OHT10]. Focus [ZSY14, BH17a, CXpHqZ15, GK14, Gau17, TA17]. Folkman [EG18]. Follow [Tod14, CP16, CB19, MAA18]. Follow-up [Tod14, CB19, MAA18]. followee [YZB18]. follower [YZB18]. follower-followee [YZB18]. following [CÖT15, SRF16]. Food [CP12b, TCH^{+}15, VSVR15, ZZY13, ACF^{+}17]. footprint [CGV12]. footprints [GGW11, WFG16]. forecast [Mue16, VEJC^{+}18b]. forecasted [Tsa15]. Forecasting [Ano11, KL17, KvES11, LW10, SB14, BGM17, MdFdA^{+}14, YLH^{+}17]. foreign [Bre13, GXC^{+}19, Jar10, Kim10, Mor19, Vel12, WY12]. foreseeable [VZAMG19]. Foresight [SB15, SL10]. forest [GR14]. Forestry [CP12b, BMP^{+}14, VB12]. Foreword [GS15]. formal [BHA15, BI18a, CGC18, LWM^{+}15, dSD18b, dSD18c]. formalized [HO19]. formation [CGC18, DR10a, LSC10, LOMLPA^{+}17, TCR10, XBD^{+}18]. Formational [KD19]. former [IJF16, JJR10, KBL15]. formulation [GRSFV17b, Van10]. Fortune [WZX11]. Forty [PHBN^{+}15]. forum [Bre10]. forward [CP12a, CKPY19, LS17b, YK14, YK15]. fostering [Fan11, LF12b]. found [KA17]. Foundation [Nic14, ZM16, Fed13]. foundational [HR17]. Foundations [CAS16, DVMS17]. Four [WHH^{+}18, ZAJ19, AER^{+}14, BPGGdMA12, DWGL16, EGU10, GGS14, GGG16a, HFW^{+}14, LY16a, NSC13, RLW14, TG18b, WW11, The18d]. FP [VFA10]. FP-5 [VFA10]. FP-6 [VFA10]. FP6 [BM11]. FP7 [HHK^{+}12]. fracking [LJKG15]. fractal [GSKM17]. Fractional [Pri16b, Gal11, MGT14, Pra11b]. fragmentation [BK15, PZ17]. Fragmented [FEHC19]. frames [BTNS14]. Framework [BM11, LVHS^{+}15, OA10a, OA10b, OA10c, OPGW^{+}13, Car16, Egg10a, FB16, GSM^{+}16, GWA14, HK19, KHVGA^{+}16, LP18b, Moo15, MKHB15b, dSSdMAF14, UBTS16, UCH19, ZRY^{+}12, ZYX^{+}14, ZWL^{+}18, MKHB15a]. frameworks [WH16]. France [BKSS15, CAV^{+}19, vRvLV11]. Franceschini [Pra11c]. fraud [KM17b]. free [APPS15, BBP14, CYT^{+}12, DW18, Jac18].

French [BR11, BZBLP16]. frequencies [Ben19]. Frequency [HWQ^{+}18, HW10, PNS^{+}10, PYW18, RNB19, SI17, YST12, Yos13]. frequently [MHTB17]. fresh [JBMR11]. freshwater [dSNV18]. friend [WZ19b]. Frontier [WKHS19, GC10, HRB^{+}13, HRB^{+}14, Pac19]. fronts [CL16, HC14b, HCLC14, hHC15, HC16c, LCLX16, ML16, Sch12a, US10]. fruit [TFJD14]. fruits [BH10]. FTA [SB15]. Fuel [KJ14, CHY16, CYK^{+}11, HLL14, HYC15, Suo14]. Fukushima [KNK^{+}19]. full [BWbH^{+}18, Cab14, CB19, HII^{+}18, JN15, Jam17, MM18, MM19, MBR^{+}13, McC14, PROG19, RCdJ^{+}14, SH19, SK13, kWhHRkS10]. full-text [Cab14, CB19, HII^{+}18, JN15, Jam17, McC14, SH19, SK13, kWhHRkS10]. fully [JMM18]. function [BBL17a, ER12, HH17e, YCK11]. Functional [GKK15, LHBC18, LAdAMJ17, Egg10b, Soó11b]. functions [ER19b, KGG15, Lin18]. fund [BCJ^{+}17, pGSyW^{+}19, HH10]. fundamental [Vin12a]. Funded [ZT14, ÁBDFB19, ABL17, Com15, Zha10, ZLTY18, ZLLL19]. funders [QA18]. Funding [ÁBMB17, BGR19, KWM^{+}18, MÁB18, WS11, XTZ15, ZM16, ACD11, Ano16d, Asu19, Bas14, CHC13, ES16a, Fan11, GTMRE^{+}19, GBM^{+}16, HH18, IF13, KT15, LVGV^{+}11, LXWC17, LF12b, MK18, Mor16, Mor19, PHDC16, Rig13, RJ14, SV19, TBS15, Van14, WLDW12, WJD15, YWS18, ZYNZ18, ZYT^{+}16, Zin16]. funds [ACP12, Shi11]. Further [Que11]. fusion [ABM19, KHS^{+}15, XYW^{+}17, ZSC18]. Future [HK12, Ano16b, BT15, CWJC14, DG16, FAI^{+}18, FdSdO17, GP13, Ley15a, Par14a, PMN16, RBBG18, SA17, WWC19, YLY^{+}14, ZWW^{+}15, ZWW^{+}16, ZTC15, dCdSNB15]. fuzzy [WYAY12, ZAJ19].

G [Par14c, BPHL16]. G-7 [BPHL16]. G7 [YYDH12, ZG11]. G9 [HH18]. gain [Vel12]. gains [ADD17a]. Galton [GGW11]. Game [SHB14, ADD11b, BGBS18, Cha17b, GRSFV15, Xie19, ZG17f]. games [RSGFV14]. gamesmanship [SC13]. Gangan [BBSS16c]. gap [BCHH17, Dya14, MCL^{+}11, PL17, SH15a, TSRGG17]. gaps [APR19, Hei19, RHMH17, RMdO17]. Garfield [Ano17b, GSB18, BI18b, Ben12, Bha18, BHL18, BHH18, Che18b, GA18, Jac18, LWB18a, McC18, Pra18b, RH18, SS18, The18b, Vin12a, Whi18]. gas [SHS15, ZLH^{+}15]. gastroenterology [DWGL16, MDFGAM14]. Gastroenterology/ [MDFGAM14]. gatekeepers [NSC13]. Gatekeeping [MPF18]. gatherings [GALR16]. gauging [CGG^{+}17]. GenBank [CQB16]. Gender [ADR16a, APR19, CSO17, GBB15, HFW^{+}14, KKdBK12, KK13, Men18, SK14a, SK14b, vAvdWvdB12, vdBS16, ADS17a, ADS17b, ADD18b, ADD19b, BHB13, BVB13, BBVO10, BN10, CFS18, DVB14, FK16, GHS18, Hei19, HAJ12, HJM^{+}13, KFKS15, KPS12, LSL15, MS16a, MS18a, MM18, MM19, MHM^{+}13, MR18a, MCL^{+}11, OKK14, PFL19, PHBN^{+}15, PMJF19, PL17, PZ17, PB12, Puu10, SM15, SM17, TN19, VO17, CÖT16b, Kra17, Zhu17]. gender-based [PZ17]. gender-inclusive [GHS18]. gender-specific [PB12]. gendered [AF18, DFS14, The18d]. gene [BL17b, Per18]. genealogy [DRMMC19, SZ12, dSNV18]. General [Coc18, SJ10, ByLbH16, FG15, NG16, RCN^{+}14, RWG^{+}15, SFR^{+}19, YWW17, ZRY^{+}12, GBM^{+}16]. generalising [The17b]. generalization [KRP19]. generalized [AD13, LAL15, MD12, vZ13]. generated [Ama15, BM14a, TL18]. Generating [ADM19, KKOS19, LBW17, OH19]. Generation [NSMMDB19, CFSSP16, HB18b, KNK^{+}19, LL19, MnaeR^{+}15, Tri10, dSFSF15]. generational [SH15a]. generations [FE16a, FE16b]. generator [YLH^{+}17]. generic [SMAABJ11]. genesis [FK17, FK18]. genetic [cSL10, dANR15, SZ12]. genetically [JBC19]. genetics [TN19]. genius [GGW11]. genomics [CQB16, LZB10]. Genres [PT17, Sug11]. genuine [Paj15]. Geographers [ZGL^{+}17]. Geographic [ACRC17, BF17, CRMdMA15, LZ14, LAHH15, PML^{+}17]. Geographical [APT13, MT12a, Wu13, LLW^{+}16, MT12b, PLWS14, SMAABJ11, jZhLY15]. geographically [OROMAA16]. Geographies [CD17]. geography [ACF^{+}17, GGG16a, LG16, MGM^{+}17, WL14]. geohistoriometric [FVVSGM^{+}18]. geometric [Ber18, ER19a]. geometrical [LZGQ13]. geometrically [Hag10b]. geometry [GSKM17]. Georgia [GHA^{+}15]. German [Ama16, AML17, BHS14, BSvEK13, BR11, BH16a, BH16b, Cle16, KWW15, MSH16, ÖS17, PKSG12, TE18]. German-speaking [BSvEK13, KWW15, ÖS17, TE18]. Germany [BB15, Bor16, BBS17, Chi14, Chi15, CDD15, Cle16, HL17, SCGZSL^{+}13, SHB14, SS10c, Zha18, ZL15b, ZB15, vRvLV11]. get [uHBLKH19, PS13, RPP18]. gets [EC16]. getting [HA17b]. gh [Gal11]. gh-index [Gal11]. Ghana [Mêg13c, ONB17]. ghost [OMMMTLC17, HQY^{+}18]. giant [MR10, MMOMALC16, ZG11]. giants [GN17, QZL^{+}17]. given [GRSFV11, MCB15, TBW^{+}12]. Glänzel [Pen19]. Global [AATBPAB15a, AAG14, BJIB16, CSR^{+}18, FZZ^{+}12a, JBC19, LCR13, LH14, LFLG14, LLW^{+}16, NHY^{+}14, OMCP17, PLW^{+}15, SFNO12, THFBdMA18, VSVR15, YCL^{+}13a, jZhLY15, ZYT^{+}16, ZZ15, ZLN^{+}13, AM18, Ano18c, BSK15, BS11, Boy17a, ÇAAÇ15, CZW13, CH15, CKB^{+}14, EGUB12, GVGSEPRC15, GS12, HSL^{+}14, He13, Ho16, HKWC15, HSLP14, HH18, JCCC13, KSSB13, KHA17, LTM12, LLGW13, LZC17, LSR13, Lor10, LL12, LWW^{+}11, MLY^{+}14, MGM^{+}17, MM14a, MHFB17, MT13b, OGRMOP19, OMJLVSN19, RY14, RG18, Saf13, SB15, SM16b, SY16b, SZD16, Sot10, WW12, WLY14, WHW^{+}19, Zel12, ZLG^{+}15, ZXM^{+}16, Bai18]. Globalization [MHM^{+}12a, Paj15, AD18, Cho12, KPSL12]. GM [JL18a, JL19]. GMF [ZXM^{+}16]. GMF-related [ZXM^{+}16]. go [KJES16, KR17a]. Going [Pra10b]. Gold [GOPG13, DGGBDG17]. Golden [Ber18]. Golden-ratio [Ber18]. good [JVM17, MDG10, NG16, Tu19]. Google [MM17a, AT17, Agu12, AL12, BI10a, EGU10, Fra10, Gus19, Har13d, Har14b, HA16, Har19b, HWQ^{+}18, JN15, MMOMLC18a, MMOMLC18b, Mik10, ML10, MM17b, MOO17, OMCP17, OMLC14, OMLC15, OMAMMLC15, OMAT19, Ort15, SX16, TK17, Var12, Wil15, ZLLL19, dS18, dWZD14].

Gothenburg [SV19]. governance [MPF18, PHV17, QA18, TE18, VDV16]. Government [CP12b, CHC13, Chu14, Com15, Das16, IL14b, KMP^{+}11b, LZZ^{+}12, LPL14, PYL16, SZZC18, YPH10, VG17]. government-funded [Com15]. governmental [Med15]. governments [HSWC13]. GPS [WLHZ13]. grade [AD16]. graduate [CÖT16b, SLD^{+}17, ZTP18a, ZTP18b]. graduates [CGZ10, HR11]. grained [UBTS16]. grammatical [TL18]. grams [ATJ16, BASL16]. Grand [Ano16b, DG16, Hau16]. Granger [LLCL11]. grant [CFP14, GNHT18, HRB^{+}13, HRB^{+}14, HBT16, KCU19, Shi11, YPH10, Zha10, vdBSS18]. grant-funded [Zha10]. granted [RGdCMM17]. granting [ZYF^{+}17]. Grants [HN16, CLSW19, EC16, PPK^{+}16, aSTS17, WLF15, YHL^{+}18]. Grants-in-aid [HN16]. granularity [GT18, ZZZC16]. granulomatosis [ARE^{+}18]. grape [ATCCAAB19]. Graph [JSZ13, BH18b, KM15a, SKCK14, ZLZ19, Zyc10]. Graph-based [JSZ13]. graphene [AYS^{+}13, EGUB12, Kli16, KPY16, LWW^{+}11, YKLK14]. graphical [wHwH16, LLH^{+}16, Mag14b]. graphs [Fie15c, HC14a, ICC16]. gratitude [SMAABJ11]. gravity [HT19, ZG17b]. great [CT15a, CMO11, EW15]. greater [WLM15]. Greece [KKT^{+}18]. Greek [Kaz14, KDFL14, Kaz15]. Green [Laa14, Suo14, ZW18a]. Greetings [Glä14, Har13b, Har14a]. Grey [PJL19, WST14, ZW18a]. Grim [Jac12]. GRIs [Sha12]. ground [MCvFP16]. grounded [MdBdP^{+}19]. group [CGC18, KBT15, KB11a, Kon12, LF14b, MKHB13a, MKHB13b, NPT^{+}15, Wal16, ZZY19, vR12]. group- [Wal16]. grouping [LLHN17]. groups [Cam18, CGC18, GNHT18, KR17b, LKW^{+}16, MSC18, PROGMA10, RGLE16, RGGBV16, RBC^{+}10, SYDW19, VHG^{+}15, YSD11, ZC14, ddMS15]. grow [YWG14]. Growing [ML18, YHC^{+}15, Zha14, dCPF14, FK17, FK18, GdA14, NLCC17, SYLC17, WM17, dAG13]. Growth [BDC^{+}12, CR14, dANR15, RRSA18, SAPR18, ULFRU^{+}14, WWL17, AH11, CV15, Cha17a, Che11, CSC12, FCTV12, GB16, ILP13, ILBG14, JL18a, JL19, JSZIZ13, KM15c, KLCS14, LvI10, MS12, NQ14, OMAT19, RRBA10, SLG10, San12d, SY16b, Suo14, TS11b, TA11, WG10, Won19]. guanxi [LM15]. guarantee [WLM15]. guarantors [MAGBBM13]. guide [Ben11]. guided [Moe10]. guidelines [BM12b, JSZIZ13]. guiding [Sch14a]. guilds [RRL16]. Guimarães [dCPF14]. Gulf [Moe16b]. GVC [LM16].

H [CSO17]. H-Index [CSO17]. H7N9 [TZ15]. habitats [dCdSNB15]. Handbook [Tom17]. handicap [Bor14]. handling [Ano16b, SGSS17]. handover [ML18]. HantaVirus [KM16]. happened [SS18]. happiness [DLM15, KN15]. hard [ACD13, BDF^{+}17]. Harmonic [Hag10b, Ber18, JX13, Pra11b]. harmonisation [Ano16d]. Harvard [GD11]. Harzing [Ben11]. hashing [AT18]. Hawaii [GG13]. Hawthorne [Bor12]. hazard [FGP13, MV19]. haze [YHC^{+}15]. HCI [MAA17]. HCVDBegy [EDEH16]. heading [AW11, KGB^{+}18]. Headings [LCS^{+}16, GWBSVWB13]. heads [Sch10b]. Health [CYW^{+}11, CRZGVQMA15, Das16, Har15b, Wil15, AGLNRR14, BJIB16, BSKB17, CRZGVQdMA16, FEHC19, GMJ^{+}17, HRH10, JDG14, KB18, MC10, MC13, MT12b, NLCC17, OROMAA16, OA10b, OA10c, PLBZ18, PEFP13, SFBS17, WLF15, WOW13, XG18, YLY^{+}14, ZCMVQS11, ABL17]. Healthcare [HFW^{+}14, FdSdO17]. Hebrew [LKP11]. hedges [uHBLKH19]. heel [HLSC18]. held [Ano11, Ano15]. Helices [Par14c]. Helix [Chu14, CP14, IL14b, KP12a, KP12b, KPSL12, LPL14, MGMW14, Par14b, Par14c, Sko14, ZZP^{+}14a, HMK^{+}12, KZSZ19, KCP12, KHJ^{+}12, Par14a, PM18, SLK12, Yoo15, YYP17]. help [YG18]. helps [FK17, FK18]. Hemingway [MC15]. Henk [Pen19]. hepatitis [SAPR18, EDEH16]. Hepatology [MDFGAM14]. her [GR16]. herbal [BMM17]. heterodisciplinarity [Sch15c]. Heterogeneity [BAC13, Hau16]. heterogeneous [ADD^{+}15, Ano16d, HD17, Pra19b, RD13, ZZLS19]. Heuristics [BM19]. hexmap [BSBG18]. hexmap-based [BSBG18]. hIa [Rya16, HAA14, Rya16]. hidden [AHP17]. hierarchical [BCML19, JL18b, LBW17, LGD12, TZG15, WHS19, ZLZ19]. hierarchy [IB15, MYP19, XCS^{+}16]. High [AMK13, CWH11, Glä13, Jar16, Kwi18, Pra19g, BWD10, BPHL16, Che15, ERW12, GRSFV14b, GBMB10, Hen19, Hen20, JK19, MSdBC16, NH14, PMN16, RKZK18, SPB18, WZ19b, ZZ14]. High-end [Glä13]. high-energy [GBMB10, JK19]. High-impact [CWH11, BWD10, Che15, WZ19b]. High-level [AMK13]. high-profile [Hen19, Hen20]. high-quality [ERW12, GRSFV14b]. High-ranked [Pra19g]. high-speed [ZZ14]. high-technology [BPHL16]. Higher [BSFW10, CD17, PQG14, Saf19, ACD11, ADD11d, ADD14b, ADR16b, ADD17a, BLA16, CC12a, DVB14, DVB15, EdS19, FWFM18, HV18b, HIC12, Kwi18, LAS14, LVHS^{+}15, LT10a, MSdBC16, Med15, MZ14, NSMMDB19, PSZ15, Pin15, PFPCM^{+}19, QJZ^{+}14, San18, SL17, TK16, The17b, VPM16, VAJCC17, WDP11, WS13a]. highest [CFG^{+}14, GRSFVMB12, GRSFV19a]. Highly [DC19, FH16, HH17c, NvLvR10, PCRMCB^{+}18, VH17, dSD17, BB15, Bor16, BWdMA17, BBS17, BL18, Ham11, HBS^{+}19, Ley12, Li16, LHW16, LDZ17, MMOMLC18a, MHC^{+}15, MBA13, MBTKA14, MC12, PLA10, PAL13, Per10, TKA17, The19b, VVN16, WYY11, WFZD19a, YY14, ZG17c, ZWW^{+}18, ZW18b, ZS11]. highly-cited [BB15, BL18, MMOMLC18a, ZG17c]. hinder [LXH^{+}18, SB17]. hindrance [YZ17]. hiring [CR18, ZYS16]. Hirsch [Rou12a, Egg11f, Que10, Que11, Ran09, San12c, Sch12b, Sch15c, YWL16]. Hirsch-core [Egg11f]. Hirsch-inspired [Sch15c]. Hirsch-type [Rou12a, Sch12b, YWL16]. Hirschian [Sch10a]. Hirschman [Rou18]. Historical [GK19, EBK16, Hou17, Lee10b, PHBN^{+}15, WLHZ13]. historiography [Col18, CL16, Kol12, Oli15a]. history [Col17, Hei13, HT19, Kol12, Kra16, Mus12, PR15, RCJ18, Smi12, Soó14b, WMXZ14]. HIV [Lja16, PP11, QA18, WT14]. HIV/AIDS [Lja16, PP11, QA18, WT14]. Hochberg [dM10]. Hokkaido [GY12]. hold [TÜ10]. holders [GNHT18]. holding [MB16b, VHD^{+}16]. hole [GP15]. Home [CGSS13, DD18, Fan15a]. homepages [KBT15]. homonyms [Har15b]. homophily [PZ17, WDN17]. Hong [Liu16, Hor18, LYGQ12, LXL15, LXL16, ML18, Pra18a]. honorific [PG14b]. Horizon [EC16]. horns [RJ14]. hospitality [ZTC15]. hosting [AR18]. hostnames [Lin11]. hot [BYY18, CW17]. hotspots [ZDZ^{+}15]. house [EDEH16]. HRI [Bar11]. Hsinchu [Hu11, Hun12]. https [Glä18]. hub [SK16]. Human [FMU16, WHH^{+}18, BSG17, Bar17b, BL17b, Coc18, GRSFV14a, JKN19, TYWZ12, WL14]. Humanities [EOS12, HH15c, Lin10, OE15, VE14, BDF^{+}17, Cha13, hCcTmWH15, CRMdMA15, DL16b, GSE^{+}18, Ham14, HYYL12, HLY14, KEP^{+}18, LL15, Lin18, LXH^{+}18, MLC14, MRGT13, Pol16a, RG12, SL12a, Siv16b, SMY15, TCC17, VO17, Whi15, Wra16a, Wra16b, YQX10]. Hungarian [PV15, SND19, SK11]. Hungary [SK12]. hurdle [BBCP14]. Hybrid [pGDTP12, LGD11, Abb13, BM13a, BM13b, CLL^{+}17, CFM15, GT17, HZQ^{+}17, Jun12, LZB10, LGD12, MBSB17, Mou16, SE18, SDP^{+}19, TSG13, TG18b, ZZP^{+}14a, ZHZY19, ZYX^{+}14, Zit15]. hydrates [SHS15]. hydrogen [MnaeR^{+}15]. hype [CV15, Jun12]. hype-type [CV15]. hypergraph [LGS18]. hypergraphs [TCR10]. hypothesis [BMR12, EG18, Sch15a, Wu18, LBGBMA10].

I-UGR [RGTSLCH14]. i/o [CL16]. I3 [Ley12, BTL19]. I3* [LBA19]. Iberian [CPF13]. Ibero [PEPUT15, SLGO17]. iCE [Pra10c, Pra10d]. Iceberg [LBGBMA10]. ICIS [CST11]. ICM [Dan14]. ICT [BM11, GGG14, SK14c]. iCX [NPP^{+}12]. ID [Ama18a, KT15]. idea [LL19, WHS19]. ideas [Cam12, LHLH19, Yur16b]. Identification [AA16, CZPR17, HLW19, LP10, PYK13, QA19, RBF^{+}10, YL10, Ano17c, ABM19, AP14, BGG^{+}17, CLSW19, CYK^{+}11, DDS^{+}19a, CNC18, DRCG17, DXL^{+}18, EO14, GGS17, HB18b, HY19, HW10, IFT^{+}18, LKS^{+}14, LHBC18, MXZ18, MSA13, MAGAM13, PNVCB18, PCRMCB^{+}18, RKZK18, WLZ^{+}19, WT15, WFH^{+}16, YZW^{+}17, YKLK14, vEWNB10]. identified [WRM17]. Identifiers [BC17]. identify [AUS12, CLL^{+}17, GALR16, HPKS18, JKJL14, LL13b, MP15, MÁB18, RGGBV16, SHR^{+}10, SHB14, SD13, SÁV18, WMH^{+}17, ZG17a, ZLLD19]. Identifying [AKB12, AHP17, BHPVdPMR18, BYY18, CGV12, Cha19a, CLLZ15, CXpHqZ15, CDCK13, DLGP16, HT18, HQY^{+}18, Hud17, JS15, Ken18, MM14c, OROMAA16, PYK12, PLJ18, Shi14, SHL15, SD18, TBMM18, WLD^{+}14, WP17, YK11, YPK13, Zel12, ZZFD18a, APYS13, CJW10, CF14, CL17b, DW18, Fan19, FAA13, GWB11, KL17, PROGMA10, RHMH17, RMdO17, SJ10, SKM15, SDEB16]. Identity [OVJM17, AP14, ML13, SRF16]. IEEE [KY17, SFNO12]. if [DDR17, Sch14a]. IFQ [TSMTDLCH11]. Igniting [HR17]. ignorance [GRSFVCP19]. II [GTMRE^{+}19, HAJ12, HH17c, KPS12, LF12b, TC13, ZG17d]. Ilan [Ano17a, Ano19, The17c]. illustrate [vWBS^{+}16]. illustrated [HSPY15, Osw10]. illustration [IB15, LJC^{+}15, RR17]. imbalanced [KK18, MXZ18]. iMetrics [ML13]. immediacy [kWhHRkS10]. Impact [ATM16, BW10, BND11, Bor17, BDE11, Bre13, Bro13, BB17b, Cam12, DPF^{+}16, EMH^{+}10, wHwH16, KM17a, KPJT14, MWDC12, MS14, SBSR19, SAR19, SK17, Smi12, Ste17, TUCR15, Van12, Vin12a, ZW17c, vL12, ADD10, ADV11, ADR19, AZKR13, AER^{+}14, ATK17, ACP12, AhOL14, AW10, ABMSSP16, ASPF^{+}16, AdAdAM10, Avk13, AMI18, BL10, Bal12, Ban18, Bar17a, Bar17b, Ben12, BT17, BT18a, BT18b, BT18c, BAC13, BGAAM15, BWD10, BM12a, BM14a, Bor15b, BHM16, BH17a, BH17b, BYY18, BH18b, BTL19, BW19, BW20, BBVO10, BD13, BZBLP16, BT15, BMZ^{+}17, BCZ12, CTL^{+}19, CFSSP16, Cam10, CC11a, CCM^{+}11, Cam11, Cam14, CV15, Car16, Cha18a, Che15, Che18b, CG17, CWH11, CFM15, CM18, CR14, CFdC^{+}14, CG15b, CIL^{+}16, DL16a, DL16b, DNAH15, Din14]. impact [DGDG13, Egg10b, Egg11a, Egg14a, EW15, Emm19, ENST16, ENA19, FMP17a, FSSPG^{+}15, FRdA16, FESD11, FZZ^{+}12b, GRSFVMB12, GPN14, GF17, GD11, GLD16, GK18, GZ11, GT18, GOPG13, GKS16, Ham14, Har12, HH13, HH15b, HMCL16, HNG19, HEH18, wHwH11, ILP11, Ing12, IL14a, ILGZ^{+}14, IPIU13, Jac12, JDLIV14, Ke13, KY16, KK18, KJES16, KKK^{+}14, KFB18, KTT11, LVGV^{+}11, LRC19, Ley12, LO12, LBA19, LTG12, LRS^{+}18, LCFC14, LJ10, LF14a, LABL13, LZCZ18, LNK^{+}14a, LNK^{+}14b, LAW14, LDZ17, MDG10, MMAHS10, MRS^{+}16, MTU17, MT13a, Mor19, Mue18, MD12, NHLL17, NF19, NPP^{+}12, APFR^{+}13, OM11, PKR15, PYH16, PJB^{+}12, PLBZ18, PA12, PHV17, PRA16a, PP16, PPI17, PMN16, PG12, PML14, Ric17, Rig13, RPDCRVRP15, RP17a, RPK18, Rou12b, SBT18, SvLVA19, Sch16, SSN19, SM14]. impact [SY16b, Sot10, The12, TK16, The17b, The17a, Tol12, TSRGG17, TR14, VEJC^{+}18b, VSS12, Van10, Vel12, Veu10, VG11a, VG11b, Vin10b, Vin17, WM19, WvE13, WYAY12, Wan13, WTG15, WLM15, WJCC19, WST14, WZ19b, XLR15, YSM^{+}19, YWS18, YLSW16, YXW18, YYL10, YYLW14, Yur18a, ZS17, ZHG16, ZG13a, ZGL14, Zha17, ZNB^{+}17, ZMW^{+}18, ZYNZ18, Zha10, ZW17b, ZZZC16, ZY15, Zit11, Zit12, ZXT^{+}19, Zon19, Zyc10, vLvWW16, vR12, vdBS16, BSFCC15, GAVZAB12, KR17a]. impact-citations-exergy [NPP^{+}12]. impacted [DdS19b]. impacts [Ama16, Fuk19, JDH12, JH16, Ley13a, Lia11, WS11]. Imperfect [Bro13]. implementation [Car16, NLCC17, QJZ^{+}14]. implications [AP14, Bas11, BM15, CT15b, DKS18, GHS18, GKK15, HF19, PW17, RRBA10, SM16b, Sko14, SVS18]. implicit [uHBLKH19]. Importance [dOM16, ADR13, BMM17, BLS15, Com15, RGTSLCH14]. Important [LLLL18, Bue15, QA19, ZL15a, ZZFD18a]. importation [Rod17]. impossibility [Osó18]. improve [DQ11, Lia11, PLWS14, RRLNAG15]. Improved [TDG17, AR18, DH13a, IFT^{+}18, WLZ^{+}19]. improvement [KLP12, dANR15, RMCM13, TZG15, ZYZ14, ZJLG10]. Improvements [CHL10]. Improving [BR11, FZZ17, FT19, FDVZ16, GNVQdMAG11, ZCL15, ZYF^{+}17]. In-depth [ZG10, GE11, IS16, PR10]. in-house [EDEH16]. in-licensing [LYWSV13].

Inaccuracies [ABRVZ15, Har15a]. inaccuracy [ABMRVZ14]. inaugurational [GGG^{+}11]. inbreeding [IT11, SM14]. incentive [CLD13, HS16a]. incentives [MCL13, WDL17]. Incidence [SL14]. include [HLC17]. inclusive [GHS18, HM15a, HM15b, SW19b, Sug11]. inclusiveness [Cop19a]. income [YLC18]. incomplete [dCCMAW16a, dCCMAW16b]. inconsistencies [MS18b]. Inconsistent [FGP13]. incorporation [McC14]. increase [ALH15, Egg10c, PP16, SLH18]. increased [Mik17]. increases [BH18a, Fan13a, KCT^{+}17]. increasing [GG19, GVS17, Hen18, Sch14c]. increments [LBGBdMA13]. incubators [DVMS17]. Independent [IH14, LW15, Abt12, Bur12, FH13, PDAN19, San12b, WFH^{+}16]. Index [BBSS16b, CH12, Egg10c, FD14, FH16, GWBSVWB13, HPS19, Ho13b, HW12, HZL^{+}17, IH14, IFH15, KB11b, MWDC12, MH16a, MPS^{+}18, SKM15, SJOC18, TFH14, TYWZ12, VASNU^{+}19, ZKC^{+}16, ZGJ18, Abt12, uARA19, AvLS14, ACHVH10, AA19, Ask18, AA10, AA16, AMI18, BL10, BIL15, BI18a, BK11, BBSS16c, BLA16, BB16, BBL17a, BBL17b, BT18b, BT18c, BSMD11, Bur12, Cab13, CP12a, CG14, CF18, DCM16, DCS12, Don17, DGDG11, DW18, Egg10d, Egg11b, Egg11c, Egg11e, Egg11f, Egg13a, Egg13b, ER19a, FMPP10, FM11b, FGMM12, FMM13a, FMM13b, FMM14, Gal11, GP13, GBB15, GQAM19, Glä10, GTC16, HAA14, HBA19, HSW10, Hir05, Hir10, Hir19a, HK12, Jac12, KPJ16, KB12, KCP11, Laz10, LFBI19, LF12a, Lun19, MHLGHV14, MJHG13, MT13b, MZE19, MKHB15a, MKHB15b, NASR11]. index [OBG11, PD10, PDAN19, PPM^{+}17, Pra10b, Pra10a, Pra10e, Pra16b, Pra17a, Pra19c, Que10, Que11, RAA18, RZ12, Ran09, RG12, Rou12a, Rya16, Saa10, SOBM16, Sah16, SAPR18, San12b, San12c, Sch12b, ST14a, Sch15b, SS14, SL12a, Ste19, TH19, Tod11, TSMTDLCH11, TSRGCCJC14, Tur16, VB12, dFVDU^{+}19, Vin10b, kWhHRkS10, XWL19, XGY^{+}16, YAC10, ZYZ14, ZLF^{+}14, ZP16, dS18, dSD18b, dSD18c, vdBBdK16, CG15a, BBSS16a, BH16c, CSO17, FLH14, HB15, Ho14, HH15c, HWQ^{+}18, LvI10, LLG14, LL12, McC18, MSB18, PDAN19, SLG10, Sot10, SDS14b, Ye14, YJ11, AAH10, BT17, BT18a, LLX^{+}18, MH16a, Tom17]. indexed [BI10a, Bas10, CR14, IWK18, JZL10, MMA18, ZLH^{+}16]. indexes [AD13, ADV13, AC13, CFSSP16, CRBRG^{+}18, DGDG11, Fed13, GGP14, TA14a, BI10b]. indexing [Cop19a, FMM15a, GWBSVWB13, IFT^{+}18]. India [BBSS16b, AG13, Bas13, BBSS16a, BBSS16c, BSB12, BSK15, BKSS15, CO10, KGNB11, KG10a, KM15b, KM15c, KM15d, KGG15, LM19, MR10, MH15, MH16b, MM14b, NPP^{+}12, Pan14, Pra16b, Pra17c, RASP13, SR15, SUP15]. Indian [BMM14, Bas13, EKR19, GK14, GGG^{+}12, GKV11, KM18c, LR12b, MR10, SLG10, SJ19]. indicate [MPF18]. indicated [Ley15b]. Indicator [APLHF18, Sol06, CNPG18, CC11b, DQ11, Doc12, FMPP10, FMM14, IL14a, Kis11a, Kis11b, KB13, LML11, LPL14, LBA19, LMdBG16, MWDC12, Moe16a, MKHB13b, PLWS14, Pra19g, Sch15c, SPdSM16, VG11a, VG11b, Vin10b, WYvE11, WvEvL^{+}11b, XWL19, ZQH^{+}17, APPF18, XLR15]. Indicators [AOFU10, BvdB14, FCWH11, Vin10a, AAB^{+}13, AYS16, ACC^{+}16, Ano16c, AAV13, BVZV16, Bas14, BFHS18, BR11, BBL17b, BW10, BG17, BTL19, CG18a, CGG19, CRZGVQMA15, CWH11, De 16b, DCM16, DMM17, DSM11, Egg14b, EHK12, Fan15a, dCPF14, FE16a, FE16b, Fra10, FH13, GSM^{+}16, GSTD11, GM13, GALR16, GdA14, GWB11, GHvdB12, HL15, HRC12, JPZ^{+}10, Kar12, KP12b, Kos16b, KBZS15, LVSL18, LO12, LWB16, LPL16, LLP^{+}16, MS18a, MR10, MR13, MBT16, MGMW14, MdFdA^{+}14, MK19, Mue16, MKP16, NTM^{+}18, NF13, OMR14, Pol16b, Pra12b, Pra13, Pra14a, Pra18c, Pra19d, Pra19e, RC13a, RM10, RCETS19, RPNC13, SG10, Sch13a, SAR19, Sug11, SN10, The16, The17a, The19b, TW18, Vin12a, WM19, WvEvL^{+}11a, WvE13, WC18, WJCC19, WSL14, Wil15, WFH^{+}16, WF17, YWL16]. indicators [dAG13, dOM16, vR12, Ano10, Har15b, LHG16a, Moo15, ZWW^{+}18, Pra17a]. indicators/approaches [WC18]. indices [Abb11, AAH10, ACHVH10, Ask18, Bal12, Egg11c, Egg13c, EMH^{+}10, Glä12, GKB^{+}19, ILB11, KKBW17, Koz15, LZGQ13, MPH19, MKP16, Pra11b, Pra14b, TABA16, Tod11, Van10, WWC19, WOW10, WZ17, YWL16, ZWZ^{+}19, ZWX22, Zyc10, ddMS15]. indifference [AM18]. indigenous [LM15, WY12]. indirect [Egg11e, Egg11f, FE14, FE16a, FE16b, JH16, VB12]. Individual [Oze12b, AD11b, ADD17b, BBCP14, BSvEK13, BM14b, Bur14, CHY16, CvLB10, CvLvR11, FGMM12, Fuk14, GHT16, GSKM17, HAA14, Hir05, Hir10, Hir19a, HS16b, MBA13, PRRC15, RC13a, RCN^{+}14, Rya16, San12a, San12d, San13, SSS^{+}11, TH19, Wal16, WJD15, ZS17, ddMS15, BC13b]. individual-level [Wal16]. individuals [GHA^{+}16, HSWC13, WXW^{+}13]. Indonesian [WAT16]. Industrial [Tom18, uHBK19, FK16, KLCS14, Kon12, KFKS15, LLHN17, ZWW^{+}15]. industrial-organizational [FK16]. industrial/organizational [uHBK19, KFKS15]. industrialised [SS10c]. industries [Cha14, HSWC13, KCP12, KL17, WF17, Wu14]. Industry [HYC15, ZyZZ^{+}11, ADS12, AJCACRdMA16, CSC12, CHC13, CSC13, CKCK10, CK14, Chu14, DMM13, FYC15, FLZ17, GGG14, GSPLVG^{+}18, HF19, HWLL14, IL14b, JCCC13, LZZ^{+}12, LZZ^{+}13, LPL14, LJ10, LLW13, OMA15, RVFEdlM10, SR16, Sha12, SN10, TM12, TYYW16, VRF12, WLN^{+}14, WHLP16, WA18a, WS13a, WF17, ZL18a, ZZP^{+}14a, ZW17a, ZLLD19, ZLH^{+}15, DDS^{+}19a]. Industry-academia [HYC15]. inefficiencies [PS16a]. Inefficiency [ADD14a]. Inequality [GHS18, Kor19, LXWC17, Pra17b, ZM16]. inertia [WHZ14]. infectious [KM11, TZ15]. infer [Mou15a]. inferring [HRB^{+}13, HRB^{+}14]. inflation [DWGL16, dSD18a]. Influence [AJSN18, Cha16, FMM15b, IL14a, ILGZ^{+}14, KY16, MC10, MdNS^{+}19, TBB^{+}16, YY14, Avk13, BPGGdMA12, BGAAM15, BGJB16, CD17, CFG^{+}14, CM18, CB11, DNAH15, Egg11b, FYC15, GFC18, GSM^{+}16, GTAG18, GBMA14, HWLL14, ILP13, KJW^{+}17, LWL17, LXWC17, OPGW^{+}13, PW13, Per18, PRA16a, Pra19e, RW11, SP14, SND19, Sch13a, SLISC17, SK18, SZZC18, TM12, The18d, The19b, WTG15, WZCC19, WW11, ZW17a, ZW14, ZW19, ZZLS19, vWWtH14]. influenced [CNPG17]. Influences [ZZZ^{+}14, CLLZ15, CC12b, KGSS16, ZT18, ZT19]. influencing [BK10, BKSS15, BMR12, CGC18, DX17, ES16a, GNHT18, MOA16a, VEJC^{+}18a, VHG^{+}15, WYAY12]. Influential [HBS^{+}19, GRSFV12a, GRSFV^{+}12b, LDVSGD19, RH18, SKM15, TBMM18]. influenza [TZ15]. Info [JAAA18]. infomap [VYL17]. inform [CRBRG^{+}18]. informatics [BCML19, JK10a]. Information [ABMSSP16, Ano15, Ano18a, Ard12, CXpHqZ15, Glä18, GAPP18, hHSL19, KW15, LM19, LHC16, ML13, NJM18, PEFP13, PB17b, RPK17, YSM^{+}19, ZSY^{+}13, AF15a, AF15b, AW10, ASW18, Bai18, BH16b, BG12, BKRG13, BS16, Bue15, CFM18, CHL15, Cha18c, Cha19b, CYT^{+}12, CRFM^{+}12, CST11, Das16, DXL^{+}18, Dor17, Fia11, FMP17b, FLB19, GP15, GRSFV^{+}13, GRSFV19b, GKK15, Glä15, GTGABAG15, GWA14, GWG17, HD17, HG17, HYC18, HC15b, HHGZ11, HHDL13, HC12, KZSZ19, KHS^{+}15, Kha13a, Kha13b, KZC16, KWS17, KW17, KGG15, LL16, LY16a, LT16, LPL14, LYW19, LC18, Lin12, LD16, LWM^{+}15, LLW^{+}16, LTK^{+}19, LJMF15, Lun19, MP15, MBA13, MS15a, MS15b, MT12b, MMSS11, MM15b, NSKO15, Oli15b, OGOPPR17, OI17, PK14, PYK13, PQG14, PPE14, PEPUT15, Pin15, PFPCM^{+}19, PHS12, QZZ17, SMLHCP17, SH19]. information [SM16a, SBB16, SGG^{+}14, Sha12, SKCK14, SFNO12, Soo19, SMM15, SB19, SN10, tScL13, The17c, yT11, yT15, yTmShL16, TT13, VDV16, WW15, WYvE11, Whi10, Wol15, XGY^{+}16, YDZ10, YL12, Ye14, ZLYF14, ZSC18, Zha10, Zhe19, dZLwC^{+}15, ZYS16, dSF13, CWH11, RPK18]. information-seeking [BKRG13]. informational [EO14]. informative [GMSZ18]. informatization [ZY19]. informed [AD11a, BD16a, BGJ^{+}16, LD16, MCR^{+}12, RSGFV18]. Informetrics [Ano15, KÖG12, KG13, RM18, SSAG16, AER^{+}14, BI10a, CP14, GLM11, KP12a, Wol15, ZRY^{+}12, GGH^{+}14, LLRG10, OING12, Pen19]. informing [TKA17]. infrastructure [JC19, KM11, LRA14, SK14c, WK15]. infrastructures [JDLIV14, MT15]. infringement [Jam17, PYK12]. inherited [AF18]. inhibiting [IMH13]. Inhibitors [WT14, CWJC14]. initial [BD10b, FRdA16]. Initiative [Bor16, MSH16]. inlinking [Hol10]. inner [CGPT15]. Inno [CC12a]. Innovating [URU10a]. Innovation [ALvH19, CS19, JCCC13, Koz15, UMdSV12, dSFSF15, Ahr17, APPF18, APLHF18, Ano16b, Ano17d, AMMT16, ÁRM13, BD12a, BS19, CSC12, CHC13, CLkS11, CyPP12, CP14, CL17b, CKB^{+}14, DLL^{+}16a, DMM13, De 16a, DdlPPL^{+}19, FLZ17, FWFM18, Fuk16, GRBBS17, GL15, GWS15, GP18a, GZ14b, HMCD^{+}19, KZSZ19, KRP19, KP12a, KVC15, KPY16, Lee12, LPMK17, LAHH15, LJ10, LJC^{+}15, LYLD15, LLHN17, MvdH13, MDDG17, MCCU16, MT15, PM18, RVFEdlM10, RS12, RP17a, RBBG18, SLK12, Sil13, SN10, SG16, TLSH14, TVA17, TP11, WhCL10, WLY14, WY19, WZ19a, Wu14, WHZ14, YKCK13, Yoo15, YLJ^{+}17, ZK19, ZZP^{+}14a, ZQH^{+}17, ZG17a, ZT18, ZT19, ZG13c, ZB12]. innovation-related [LPMK17]. innovations [BMZ^{+}17, HMK^{+}12, MBP19, WY12]. innovative [CP12b, HR15, JC11, LL13b, MvdH13, TW18]. inophyllum [GVGSEPRC15]. Input [PSB^{+}17, ADS10b, KHK13, Pra18e]. Input-output [PSB^{+}17, KHK13, Pra18e]. insight [MM14c, RGTSLCH14, SZ12, WHC^{+}13]. insights [AMMT16, Bou14a, Fuk16, HH10, HZL^{+}17, KHH18, MKYM^{+}17, SWH14a, SWH14b, YCPS17]. Inspecting [HLSC18]. inspiration [CPV14]. inspired [Sch15c, aSS17]. instant [ZC14]. instead [BL18, LNMQRR15, RTP17]. institute [Kon12, PLJ18, SWH14a, SWH14b, CWH11, GG13, TA15]. Institutes [ABL17, GGG14, KM15d, KZ13]. Institution [GY12, HYYR14, TSMTDLCH11, WL18, BKSS15, CMdT14, CXpHqZ15, DD18, GLD16, HSL^{+}14, PKSG12, SSS^{+}11, WLD^{+}14]. institution-specific [CXpHqZ15]. institution-topic [WLD^{+}14]. Institutional [SY16a, ADR13, Ban18, BHS14, BGBS18, BL15, BG18, BCC^{+}17, CGC18, CHY16, Doc12, DC15a, DC17, Fan15a, GTD14, GW15b, HL17, Kos16a, Kra13, LJC^{+}15, MOO17, OROMAA16, PY19, RCCM14, SIS17, SRW18, SZ15, TABA16, TW16, TSRGG17, YZB18, You14]. institutional-level [BHS14]. Institutionalization [MT12b]. institutions [ADV13, ALYZ15, Bor14, BL15, BWdMA17, CH13a, CXpHqZ15, CRZGVQdMA16, DB19, Fan15a, FFL16, FMM13b, GJ11, GHA^{+}16, JKSK15, KMS16, KM18c, LSCK12, LAS14, LVHS^{+}15, LLLL18, LLX^{+}18, LOMLPA^{+}17, MSdBC16, NPP^{+}12, NSMMDB19, PSZ15, QJZ^{+}14, RPGM16, SBB16, SUP15, UBTS16, VB12, VAJCC17, WXW^{+}13, WDP11, WLZ^{+}15, YP19, dMALIM14]. Instituto [STCRPA18]. instrument [EO14, GM12, HS16b, HYS18, ILP11, MARMSG19]. instrumental [FAA13]. instrumentation [LR12a]. instruments [LAHH15, Pou12]. Insularity [CSS^{+}16, LTM12, MGMY^{+}18]. Insurance [CYW^{+}11]. intangible [NSMMDB19]. integer [Van10]. integer-valued [Van10]. Integrase [WT14]. Integrated [MWDC12, BH16b, DXL^{+}18, KJS14, LGPC18, LBA19, LP18b, Wal15, ZZW^{+}19b, dACdFC18]. Integration [YWC12, Ano16b, Ano16d, BCML19, DLL^{+}16a, DG16, Keg15, KS18, Siv16a, ZC16]. integration-state [DG16]. integrative [Dan14, FCCMTRVR18, Kim19b]. integrity [Var11, YSND17]. Intellectual [BMM14, JK10a, dNMVQL16, ZS11, ACD14, AMMT16, BDC^{+}12, CO10, CL11, CLSW19, Ham11, HC19, KJS14, Ley11b, Ley15a, LWM^{+}15, LYLD15, LG16, Ma12, Oli15b, PNVCB18, PHS12, RAS15, SDT15, TCC17, TO18, YLL15b, ZG17c]. Intelligence [BB17b, SÁV18, AAS^{+}19, ILB11, OMCP17, WLD^{+}14, ZZPG14, ZWZ^{+}19, ZWX22]. intelligent [ES16b]. intelligible [Moe16a]. intended [SFM16]. intensity [ADD17b, uARA19, BGAAM15, Lia11, RAA18, WM17]. intensive [EN17]. Inter [JOGC17, KGL^{+}14, Saf19, ZCZ^{+}16, ACFL11, Cab11, LAL15, MY16, SL16, VDV16, WZ19a]. Inter-cluster [KGL^{+}14]. Inter-organizational [ZCZ^{+}16, VDV16]. Inter-ranking [Saf19]. Inter-rater [JOGC17]. inter-regional [ACFL11, SL16, WZ19a]. inter-researcher [Cab11]. inter-specialty [MY16]. inter-textual [LAL15]. interaction [BHA15, BSG17, Bar17b, pGDTP12, LZB10, ZC14]. Interactions [dCdAMB19, CÖT16a, HSWC13, KLP17, Mor16, RRBA10, ZB12]. Interactive [LKR14, BND11, DGF17, PC18]. interbasin [ZLL^{+}15]. intercontinental [KCM19]. intercultural [CY13]. interdependencies [DFG^{+}18]. Interdisciplinarity [Ham11, KM15a, MGMY^{+}18, WNS13, ZSCR^{+}18, BP11, CMdT14, Cha18a, Cha18c, GG13, JKPL18, JL14, LWB18a, Ley18, dCCMAW16a, dCCMAW16b, MRN14, MKP16, Nic14, PLWS14, RM10, Rod17, RPNC13, SDEB16, Sug11, WWH^{+}17, XGY^{+}16]. Interdisciplinary [JAAA18, XGY^{+}16, XDB^{+}19, ADD17a, BM14a, BTNS14, CAGL15, CY13, DXL^{+}18, HZ17, HC12, JOGC17, KLP17, KCU19, KH17, LBRR19, LSY11, LX15, Rai19, SZZC18, Sma10, SW19b, TA15, VNA16, VLV14, WZ19b, XBD^{+}18, YPH10]. Interest [Cop19b, Zin16, BH15, KJW^{+}17, LS15, SA17]. interested [BHJD12]. interests [JPZ^{+}10]. interface [WT15]. Interfaces [dJC15]. Interfield [Yur15]. interinstitutional [LLLL18]. interlinking [LAS14, OROMAA16]. Interlocking [BB10, TO18]. intermediaries [LP10]. intermediarity [PY14]. Internal [Dya17b, CFG^{+}14, DC14, SSN19, Dya17a]. International [AGHL14, Ano10, Ano15, BD10b, COS11a, DMM17, FSOS12, GGH^{+}14, HSL^{+}14, HSBW10, KB11c, KÖG12, KG13, KRR14, LHG16a, LLRG10, LKR14, Mêg13a, NHLL17, NA12, OING12, PML14, SSAG16, WXW^{+}13, WK15, WS13b, Yur16b, ZPG^{+}14, ZG11, ZZZ^{+}12, ZyZZ^{+}14, AChO19, APR19, Ama18a, ATM16, BLdlCV17, BFMRM19, Bha16, BGAAM15, BN10, Bre10, CWJBT10, CH13a, CH14, Cha17a, CJC13, Cho12, CB16, DCGZ^{+}12, Eld19, FLZ17, FPS14, FB16, FS10, GCGP10, Gau17, GE11, GRG12, HH13, HTHB11, HWL11, JK19, JL18b, KA13, KA17, KM15c, Kaz15, KY16, Kim10, KKBW17, KPSL12, LML11, LHM^{+}11, LM15, LDVSGDR16, LDVSGD19, LNK^{+}14a, LNK^{+}14b, MH14, MdNS^{+}19, Mor19, NA18, NQ14, PML^{+}17, PSB^{+}17, PJB^{+}12, Per10, PS16b, PL18, Pra13, QMSM^{+}19, QA18, RRSA18, RGTSLCH14, RPK16, SH15b, SS10c, Sch11a, SB14]. international [tScL13, TUCR15, TA11, TW18, TA17, WWL17, WTG15, Xie15, YY14, YHL^{+}18, ZHG16, ZYSS14, ZG17b, ZRL18, ZW19, ZL15a, ZG10, WTM^{+}16, Wad16, Wad17]. internationalisation [LVSL18]. internationality [CBWJ18]. Internationalization [Dya14, QJZ^{+}14, VE14, ADS11, AChO19, AZSA14, AZSA16, MT12b, PMJF19, Siv16b]. internationally [Ama18b]. Internet [MCvFP16, CP12b, LYS^{+}17, NT17, PW13, RWG^{+}15, YLL15b]. interoperability [DLL^{+}16b, DCY^{+}17, ZC16]. interplay [Zit15]. interpret [dMALIM14]. interpretation [CV15, LZGQ13, SBK19, VuHL10]. interpretative [BVOL18]. Interpreting [Sma11, The16, XA15]. interregional [TS11b]. intersectoral [TV17]. interval [Egg13c]. intervals [Bor17, Ste17]. interventional [GR16]. interview [Par14a, Par14c]. intra [CGC18, SIS17, WZ19a]. intra- [WZ19a]. intra-institutional [CGC18, SIS17]. intramural [KSB11]. Intriguing [Pri15]. intrinsic [Wu18]. Introducing [Hal13, XLR15, MM15a]. Introduction [BI10a, PC14, DG16, RG18]. introspective [ZTC15]. intrusion [ZSC18]. intuitive [LLH^{+}16]. invariance [KRP19]. invariant [DGDGSV15]. invasion [PB12]. invasions [MB19]. invented [CD14, Lee10b]. inventing [GW10b]. Invention [YCK11, CJC13]. Inventions [AAV13, Hsi11, HDW^{+}15, LHWW14, RR17, vR17]. inventive [LZZ^{+}12, LS19a]. Inventor [AD18, BT15, vPD13, GW10b, PLT14, SWCH14, ZLW19, ZZW^{+}19b]. inventor-authors [GW10b]. inventors [FMP17a, KB13, LZZ^{+}13, MM14c, PLT14]. inventorships [GGR11]. inventory [HS16b]. investigación [CRLMLM17]. investigate [FSLR10, WS11]. Investigating [AAS^{+}19, Boy17a, GD11, KJS14, KPL19, VTY17, WHLP16]. Investigation [WLF15, ATK17, Avk13, CT15b, CG17, HB17a, HP18, LHM^{+}11, LF14b, Lor14, LZCZ18, MdBdP^{+}19, PZ17, SBT18, Sma11, SHK14, WMW^{+}13, WWP14, YR10, Yu17, ZZY13]. investigators [HR11, KKLP18, Ste19]. investment [PHL17]. invisible [GN17, PNVCB18]. involvement [MB10b]. involvements [TG16]. involving [ZCMVQS11, dPdCAdMC^{+}16]. IoT [MCvFP16, YLL15b]. IP [TTC17]. IPC [LKR14, NA12]. IPCC [BP11, JVM17]. IPM [yT11]. IPR [Hei19]. IR [KMD^{+}18, Zit15]. IR-bibliometrics [Zit15]. Iran [FGP13, Moe16b, RHMH17, Sot10, Sot12]. Iranian [AKB^{+}10, Ase10, BVOL18, FJ11, FGP13, IMHG12, IMH13, Moh12, NA12, NH14, Sot10]. Ireland [OO12]. iris [ABGS14]. iron [HLLT14]. irreconcilable [DTM^{+}13]. irrigation [Bue15]. iSEER [ES16b]. ISI [CVD14, FGP13, Har13c, IMHG12, IMH13, Mik10]. Islamic [SH15b]. Islamophobia [TG18a]. Israel [BD10a, YGD17]. ISSI [SSAG16]. Issue [PC14, LHG16a, RG18, GSB18]. issues [Ano16b, GS15, KMP^{+}11b, McC18, PB17b, SLGO17, SHR^{+}10, Waa13, YB14]. Istanbul [SSAG16]. Italian [BGJ^{+}16, ACD13, ADR16b, AD16, BD16a, BKG16, CD14, DRS18, FDVZ16, GE11, GMM16]. Italy [ADV11, ADS17a, ADS17b, ADD19b, ACS18, Fan13a, MM18, MM19, MT12a, SBT18]. items [Cam17, Chi14, CB19, Egg11b, KKS^{+}17, San12d]. iterative [KKOS19, MM15a]. ITRI [SWH14b, SWH14a]. Ivory [HAL11].

J [GZ18, Par14c]. J. [Egg14a, Zit12]. Jaccardized [ST14a]. JAERI [YIK^{+}10]. Jaime [BI18a]. Japan [Fuk19, FI16, HL17, HWLL14, KT15, Kon12, LW10, LPZ17, MK19, PP18, aSTS17, SN10, Won13, YK15]. Japanese [Hos11, HN16, IF13, IS16, MRN14, Sak19, Shi11, SK17, YST12, Yos13]. JASIST [yT11]. JCR [AYS16, DGDG13, GSMT10]. JCR(R) [AYS14]. Jewish [KB13]. JIF [Cam17, YY16]. JIF-Plots [Cam17]. Jing [PT17]. Job [BS15c, CÖT16b, FLB11, SRF16]. JOD [yT11]. John [Ano12c, Ano14, Ano17a, CXZ19, Dan19, KL16, Fie15b]. joining [MM16]. Joint [ZYNZ18, CL17a, Har16b, XBD^{+}18, XDB^{+}19]. joint-authors [Har16b]. Jorge [dCPF14]. Journal [BI12b, DD18, GNVQdMAG11, Ioa06, JNA18, LL16, RCJ18, RPGM10, The12, TT13, WL18, ZJLG10, ZXT^{+}19, ADD10, AD16, ACP12, Asa19, Bai18, Bal12, BKL15, BWD10, BND11, BWD12, BM12a, Bor12, BH15, BH18b, BM12b, BF17, CWJBT10, CCM^{+}11, Cam11, Cam14, Cav15b, CGK^{+}14, CMO11, CHY13, CGG19, CM18, CRLMLM17, CP16, CIL^{+}16, DLGP16, DAYY18, DdS19b, Don17, DX17, Egg14a, EMH^{+}10, EBD15, GFC18, GSTD11, GBDG19, GSOLHO19, GB16, GBMA14, HT11, HHGZ11, Ing12, Jam17, JKSK15, KBAK17, Kis11b, KB18, KPRT16, Kra19, Laa14, LP18a, Lee15, LS16, Ley12, LBW17, LWB18a, LBA19, LHM^{+}11, LAW14, LS19b, MWDC12, MSYW12, MS18a, MHM^{+}13, MA19, MPF18, MCR^{+}12, MPH16, MKF14, MD12, NvLvR10, APFR^{+}13, PJB^{+}12, PA12, PW13, PROG19, PHV17, Pra10c, Pra12c]. journal [RGLE16, RM18, RAS15, RPP18, RTP17, Rou12b, SOBM16, SND19, SvLVA19, Sch14b, SA16, SM17, SI17, SMM15, SL10, TABA16, TA11, yTnL17, ULFRU^{+}14, VEJC^{+}18b, VEJC^{+}18a, VHH16, Van12, VT10, VG11b, WM19, kWhHRkS10, WLM15, WJCC19, WZCC19, WHH^{+}18, XWL19, XLR15, YHL19, YWZ^{+}17, ZLW16, Zha17, ZGL^{+}17, Zit11, Zit12, vR12, ANZ15, BI12a, BW10, BND11, Bor17, CD16a, Cam12, CC14, CGG^{+}17, EMH^{+}10, hHSL19, Jac12, KM17a, MWDC12, Ste17, vL12, CMPD19, FUR10, LM19, SL14, LX19]. journal-based [DdS19b, ZJLG10]. journal-ranking [YWZ^{+}17]. Journals [Glä18, KB11b, LHC16, AZKR13, ANZ15, AR18, ABMRVZ14, ABRVZ15, ABMSSP16, ABSF^{+}19, ASPF^{+}16, AYS14, ATM16, Ase10, AND19, BB10, BL10, uHBK19, BHKP11, Bas10, Bha11, BS15b, BMP^{+}14, BH16c, BPVM11, CWJBT10, CBWJ18, CC11a, Cam17, CHWL12, CVD14, Cha19b, Cha19c, CCLL14, Che15, CRR14, CG18a, CP12c, CR14, Dan14, DL16a, DL16b, DWGL16, DGDG13, Dya14, ERW12, EdS19, FCFG17, FSSPG^{+}15, Fox17, FE14, Fra10, FM11a, FMM15b, Fra17, FLH14, Fuk17, GRTPMLAJ19, GCGP10, GRSS16, GRSFV11, GRSFVMB12, GRSFV18, GSM^{+}16, GSMT10, GWA14, GBHT16, GB17a, GOPG13, HG10, HK19, HM15c, HP18, HYYL12, HLY14, ILB11, IB15, JZL10, JMM18, JMM19, KM17a, KY17, Kim10, KB12, KCP11, Kos18b, Kra10, KKdBK12, KR17a, LL16, LPC17, LY16a, LWB18b]. journals [LRZ13, LC18, LCFC14, LXL15, Liu16, LXL16, LHG16b, LDG17, Lop10, LA19, MPY^{+}13, MCL^{+}11, MZ14, MGC19, MdNS^{+}19, MHM12b, Oos15, PB17a, PFL19, PS10, PW13, PHV17, PhD18, PP16, PR15, Pol16b, QL12, RZ12, RCETS19, RWG^{+}15, Rod17, RP17b, RPK18, RTP17, SM16a, SP12a, SLGO17, San13, SIR^{+}14, SG10, SGG^{+}14, SS10b, Sch10a, Sch13b, SFR^{+}19, SRW^{+}15, SGY15, SE18, xShLY^{+}15, STCRPA18, TBW^{+}12, TA14b, The18c, TÜ10, yT11, yTmShL16, VVN16, Vin10b, Vin19, WP18, WLF15, WF18, WOW13, WAT16, WLZ^{+}15, YWW17, YHL19, YAC10, YL10, YWZ^{+}17, ZL18b, ZW19, ZXLX14, dSdSSB16, Cam18, LV12]. journey [GGW11]. JoVE [XYHD18, JNA18]. JSTOR [McC14]. Judah [EG18]. Judit [Ano17a, Ano19, The17c]. July [Ano15, SSAG16, LT10b, Tsa15]. junctures [KS18]. June [SSAG16, LLYC14]. junior [MS13]. junk [PG14a]. jury [MOA16b]. just [FM11c]. justice [Wal15]. justification [Egg13c].

K-core [XCS^{+}16]. Karenina [MB13]. Karlovcec [Rod16]. Katy [Ley11a]. KCI [KCP11]. KCI-based [KCP11]. Keeling [MHFB17]. Key [GRG12, ZLW19, BGR19, DLGP16, HLLT14, SDS14a, ZZY19, ZyZZ^{+}11]. key-route [HLLT14]. KeyGraph [CYH13]. KeyGraph-based [CYH13].

Keyword [Ben19, ADM19, BM13a, BM13b, BHH18, CHL15, HWQ^{+}18, HQY^{+}18, KC12, KW15, MM15a, SI17, SL10, VRF12, WNS13, WC18, WP17, XCS^{+}16, XBD^{+}18, YC12, YLL10]. keyword-based [YLL10]. Keywords [ZLH^{+}15, CXpHqZ15, Li18, LCIADG19, MÁB18, YLH18, ZWHH13]. kids [Har19b]. kill [PLT14]. Kirk [Tom18]. knockdown [BL17b]. know [Ley11a]. Knowledge [ACAGD^{+}17, ALvH19, BHA15, CD18, Fuk16, JvGH10, KNK^{+}19, LGZ^{+}13, Lja16, Moh12, OM11, SLK12, yT15, WZLZ13, AChO19, AEFP16, Ama18b, AHP17, ÁRS17, AC12, BCML19, BFGVV^{+}18, BMR12, CMRC15, CLLZ15, CU16, CH13b, CRLMRPA10, CRLMRPA17, CKB^{+}14, DT16, DCY^{+}17, EES13, FMP17a, FRPP17, Fin11, FKRS14, FVVSGM^{+}18, FK17, FK18, Gál17, GGR11, GG12, GLÚGML16, GN17, GZ14a, GP18b, GZ17, HH13, HH15a, HSAK18, HRH10, HL13, HLL14, HMK^{+}12, HEH17, Hu11, HFL14, HWLL14, HDW^{+}15, Hur17, IA19, Iwa17, JC19, KM12, Lar12, LSC10, LK17, Ley11b, LZ14, LCS^{+}16, LC18, LP10, LSY11, LM13b, LJC^{+}15, LL13b, MLY^{+}14, MY16, MGMY^{+}18, MKF14, NTM^{+}18, NP11, Oze12a, Pet18a, PPE14, PSY^{+}19, Rod17, RPGM16, RCCM14, SK16, SdJDD19, SMF18, Sch12a, SZZC18, SD13, SWH14a, SWH14b, SK13, Soo19]. knowledge [SL10, STCRPA18, SX16, SD18, SZ18, SK14c, TLSH14, TO18, UCH19, US10, URU10a, URU10b, WPW^{+}14, WDS16, WY19, WA18a, WK15, Wu13, WKHS19, WS13b, YG18, YC10, YS13, YKCK13, YC12, YLL10, YWY10, ZS17, ZSC18, ZLW19, ZZ11, ZZW19a, ZY15, ZLH^{+}15, ZH17, dSF13, dPdCAdMC^{+}16, vdPR18, Ard12, Har13c]. Knowledge-based [SLK12, LZ14, TLSH14]. knowledge-building [ÁRS17]. Knowledge-transfer [WZLZ13, KM12]. Kohsetsushi [Fuk19]. Kong [Liu16, Hor18, LYGQ12, LXL15, LXL16, ML18, Pra18a]. Kor [KCP11]. Kor-Factor [KCP11]. Korea [SWH14b, CP12b, CC13, CL17a, CH13b, FKRS14, HP18, HWLL14, JKSK15, KHJ^{+}12, Kim14, KTLD16, KPSL12, LPZ17, Par14b, PYL16, PHL17, PP18, PY19, SP12b, SL13, SWH14a, Won13, Wu14, Yoo15]. Korea-based [HP18]. Korean [AhOL14, AChO19, CK14, CyPP12, Kim10, KP12b, KLL14, KCP11, KM12, Lee10b, LSCK12, Lee12, PY14, PLJ18, Sha12, YPH10, YL12, YYP17]. Korean-invented [Lee10b]. KOS [WKHS19]. Kosetsushi [Fuk16]. Kosovo [JJR10]. Kuhn [MB10a]. Kuhnian [MB13].

L [KWW15]. lab [VMM15]. lab-on-a-chip [VMM15]. labeled [KKOS19]. labelling [GT12, KW17]. labels [MT13a]. labor [ADS10b, CGZ10, EN17]. laboratorie [JL14]. laboratories [CLO18]. labour [TV17]. labs [CFM15, RPNC13]. lag [ILP13, NSH^{+}11]. lagging [JKMS17]. lake [Vil10, mYqS15]. lakes [GCLcG15]. landfill [RNF19]. Landmark [ZXLX14, WT15]. Lando [Pra17a]. landscape [BSK15, KJS14, LL19, Pri16a, Rai19, SK18, SHL15, ZYS16]. landscapes [KD18, OGOPPR17]. Language [BF17, Bar17a, BS15a, BWD12, CG18b, DGD19, GXC^{+}19, GBMA14, LNMQRR15, MSYW12, MBT16, SW19b, VBTK19, vRvLV11]. languages [Fuk17, OMLC14]. Lanka [MR10]. laparoscopy [CZW13]. Large [ACC^{+}16, SYDW19, SS15, BS17, BYY18, CQB16, FEHC19, GKK15, Hal14, IF13, Kon12, KTRP17, LLH^{+}16, LGD12, LRA14, MSdBC16, NF19, PIB18, Pra14a, RC13a, SSN19, ZZL19, ZYX^{+}14]. Large-scale [ACC^{+}16, SS15, BYY18, FEHC19, GKK15, IF13, KTRP17, LLH^{+}16, LGD12, LRA14, NF19, ZZL19]. largest [KT15]. lasso [VHH16]. last [JC11, MGT14, OHT10, RBBG18]. late [BCC^{+}17]. latecomers [JK19]. Latent [MVS10, LPL16, WLC17, WP17, WDN17]. later [BWD10, HL15, Sni16, The18f]. latest [WLC17]. Latin [AR18, Ano18c, BB19, BMTA15, CSR^{+}18, CRZGVQMA15, CRZGVQdMA16, CR14, CRLMLM17, CS19, MC10, MSB18, OMLC15, RPDCRVRP15, RG18, SPB18, ZGCRVQ18]. latter [ADD10]. launched [GBM^{+}16, GGH^{+}10]. laureate [CÖT16a]. laureates [Ama18a, CÖT15, CT15b, CZPR17, Fie15a, SBB16, YGD17, ZZY19]. Law [TÜ10, Aus13, ANFF16, LAL15, Lin11, MS13, RPK18, SND19, SM14, cTnHwH17, YGD17, AYS14, AYS16, CC11a, CP16, WHH^{+}18]. law-the [LAL15]. laws [BS15a, Brz15, MH16a]. layers [ZWL^{+}18]. LCD [YKCK13]. LDA [WKHS19]. lead [ER19b, MZ14, RJ14, Tu19, WS13a]. leader [LF14b]. leaders [Bue15, Cam18, JK19]. Leadership [TG16, BFHS18, CRZGVQdMA16, Hal13, Hir19a, KB10, LZCZ18, MT13b, QMSM^{+}19, TG17, TFJD14]. Leading [CMRC15, Asa19, Cam18, CV15, FI16, IMSK14, NPP^{+}12, Osw10, RWG^{+}15, RPGM16, WLZ^{+}15]. leads [JSZIZ13]. lean [FdSdO17]. learn [Li16, LF17, WRV14]. learned [LYWSV13]. Learning [PSY^{+}19, BK10, CL11, CLJH12, CG18b, DAMC15, FA10, HII^{+}18, HSAK18, KMD^{+}18, Ken18, KK18, KK17, NZL^{+}19, RG15, SM12, THFBdMA18, WPCG13, WG11, ZML19, AMFLH15]. least [Cha16, dS17b]. leave [HA17a]. led [Kon12]. legal [SMLHCP17]. legend [MB14]. Leibniz [Ama18a]. Leiden [BG18, DF15, LB12]. Leinster [MKP16]. leishmaniasis [SFBS17]. length [ADC12, Egg13a, GMSZ18, HAA14, XGCK19]. lengthening [BL13]. lens [BM19, CS19, LJJ^{+}16]. lense [WB15]. Less [QZZ17, AC13, CVC14, MS18a]. lessons [QA18, WT15]. Letter [BBSS16c, Bor15b, Egg10e, Lor10, Pra11c, Pra16b, Pra17a, Pra18c, Pra19c, Pra19d, Ric17, HL18, BI18a]. letters [THB18, LHW16]. level [AD11b, AvLS14, AMK13, BPJ^{+}14, BFMRM19, BHS14, BBSS16a, BBSS16b, BBSS16c, CA19, CB11, Fuk19, GRSFV^{+}13, GHT17, GC10, GW15b, HSL^{+}14, Har13a, Keg15, KK19, KS17, LWIB16, LCS^{+}16, LXDL13, LLH^{+}16, MOO17, Mue16, NF13, PC18, PR14, Pra16b, RGLE16, SBA^{+}19, TABA16, TSRGG17, UBTS16, Wal16, WSC16, WSL14, Wil15, XTZ15, Yan14, Yur15, ZCL14, Zit15]. levels [AHUR11, CHY16, CvLB10, DAYY18, EO14, GBMA14, SMM15, WS13a, ZWZ^{+}19, ZWX22]. Leveraging [AChO19]. Lewison [Ano18b]. Lexical [GHT17, LZB10, MAA^{+}11, RNM18, TG18b]. Leydesdorff [Hir19b, Pra12a, LM13a, Par14a, WvEvL^{+}11a]. libraries [Kim18, Kim19a, OMR14, ZWL^{+}18]. Library [ABMSSP16, Ard12, CXpHqZ15, GAPP18, hHSL19, LM19, RPK18, ZSY^{+}13, AW10, Ben12, CHL15, Cha18c, Cha19b, CRFM^{+}12, DXL^{+}18, FMP17b, HHGZ11, HHDL13, HC12, LYW19, LC18, Lin12, LHW12, LWM^{+}15, Lun19, MH15, QZZ17, SMM15, TT13, WW15, WYvE11, XCS^{+}16, YDZ10, YL12, Zha10, ZZ11, ZYS16, KY17, PB17b]. licensed [WPW^{+}14]. licensing [LYWSV13, LCZ17]. lies [Cop19a]. Life [CRV12, ADR13, BHS14, BDF^{+}17, BM14b, Bou11, BL13, Cha18a, CIL^{+}16, FI16, GK14, Jon10, KKS16, KHR^{+}19, Ley15b, OHT10, OH19, WLF15, YGD17, ZAJ19, ZM16, vdPR18]. life-cycle [Bou11, ZAJ19, vdPR18]. life-cycles [Ley15b]. lifecycle [HG13]. lifespan [SSS^{+}11]. lifetime [CHM15]. light [hHC15, JVM17, LP18a, WM17, YLH^{+}17]. light-emitting [hHC15]. like [BW10, BDF^{+}17, BL18]. limitations [Ran09]. Limited [GP13, YGD17]. limnology [CHWL12]. linear [Bak17, ER19b, Hos11, LW15, ZCL15]. lines [BL17b, ZCMVQS11]. Linguistic [PT17, BASL16, Gan12, IA19, vdBSS18]. linguistics [Bar17a, SM16a]. link [CU16, GR14, HGH17a, KD19, LGL10, Lin10, ST14c, VRF12, WS10, YS13, YWZ^{+}17, ZGJ18]. link-based [YS13]. linkage [FLZ17, Lee12, cSL10, TYYW16, ZL18a, dSNV18]. linkages [CHC13, FI16, PK14]. Linked [ST14c, ACAGD^{+}17, EdS19, HHBB18]. Linking [MOA16b, BSC^{+}17, LF17, OMOR13b, RPNC13]. links [BB10, BL15, GZ11, HH15a, HMI19, MS15b, TM12]. liposome [ZZ15]. liquid [HWLL14, SLG10]. LIS [Cha18c, Cha19b, AZKR13, ANZ15, ALYZ15, Asu19, Cha18c, Cha19b, HSL^{+}14, Ma12, MH15, MOA16a, WDS16]. LISA [FMP17b]. list [ATJ16, HO19]. Listing [DWGL16]. lists [BKL15, YWW17, ZSCR^{+}18]. literacy [PEFP13, PPE14, PEPUT15, Pin15, PFPCM^{+}19, SB19]. Literature [Ioa06, Kos14, LCD^{+}14, ACD14, BMM14, BZ17, Bha16, CLD13, Cav15a, Che12, CXWW18, Col17, CGSS13, CvLvR11, CNC18, FA10, FAA13, GTC16, GAPP18, Ham11, HHZ14, HH15a, HIG^{+}17, HTL15, HZD^{+}15, IWK18, JKN19, JDH12, KGZML^{+}19, KM11, LL13a, LY12, LJKG15, LCIADG19, LLYC14, ML16, MnaeR^{+}15, MdBdP^{+}19, MB16a, MHTB17, Mue18, NTM^{+}18, dANR15, NFH12, NT17, OGOPPR17, PYH16, PN15, PLGC18, PLG19, QZZD18, QL12, RPAMR19, SR15, SLG10, SAPR18, SZ12, TAA16, TM12, TYWZ12, TZ15, yTnL17, VACCAJ18, VSK18, WW15, WMXZ14, WST14, YSND17, ZL18a, ZG13a, ZLL^{+}15, ZGY16, ZG17d, ZG17e, ZDZ^{+}15, dSD18a, dCdAMB19]. Literature-related [Kos14]. literatures [LLH^{+}16, VG14]. lithium [HLLT14]. litigated [SCL12]. litigation [CLW^{+}19, SCL12]. LitStoryTeller [PC18]. little [YWG14]. Liu [Ho16]. LOCA [YIK^{+}10]. local [CRLMLM17, GSM^{+}16, HGH17a, MC10, MR10]. locality [AT18]. locally [Sch14c]. locations [HP10]. Loet [Par14a]. log [Bak17, Hos11]. log-linear [Hos11]. logic [Par15]. logical [Wu18]. logically [Pat18, Sch18a, Sch18b]. logistic [Par15]. logistics [LMR16, dCdAMB19]. lone [KO18]. Long [LT16, SVS18, ADS17a, ADS17b, JC11, Kos18b, WJCC19]. long-run [Kos18b]. Long-term [SVS18, WJCC19]. Longitudinal [BL11b, AP16, BN10, FRPP17, Har13d, Har14b, HA16, HKWC15, PW13, PR10, PQG14, THAL15, TCC17, VEJC^{+}18b, WP17, WS13b, YMSQ10, dWZD14, vdBS16, vdBBdK16]. look [CAV^{+}19, DMV10, Hei19, HL17, SA11, SA12, SMF18, SL10]. Looking [FLM16, Rig13, Sug11, dCdSNB15]. Lorenz [ER19a]. lose [BH17a]. loss [KKS^{+}17]. lost [DF15]. Lotka [ER12, WZS12]. Lotkaian [Egg10a]. Love [ZG13b]. low [BWD10, Pri16a]. low- [BWD10]. low-energy [Pri16a]. lower [ADD11d]. LSH [AT18]. LTE [WLD^{+}14]. Lucio [Pra17a]. luminarie [WM17]. luminous [WM17]. lunches [BBP14]. Lung [LRWS16]. Lutz [Dan19].

M [Egg14a, Fie15b]. M&As [LJ10]. MA [XA15]. MACA [ByLbH16]. Macau [Liu16, LXL15, LXL16]. machine [BK10, CG18b, DAMC15, FA10, HII^{+}18, HSAK18, KMD^{+}18, KK18, KK17, NZL^{+}19]. machine-learning [HII^{+}18, HSAK18]. macro [CBF13, MdFdA^{+}14, Mue16, NF13, TW16, GC10]. macro-indicators [MdFdA^{+}14]. macro-level [Mue16]. macro-trends [TW16]. macroinvertebrates [LAdAMJ17]. made [CC11b, TE18]. made-to-measure [CC11b]. Madison [LSL15, MS16a]. Madrid [HMCD^{+}19]. magazine [LT10b]. Magnetic [LCLX16]. mail [SRW18]. main [BH17a, HdSV16, HLLT14, KS18, LLH19, ML16, MCL^{+}11, YKLK14, ZCMVQS11]. Mainland [LYGQ12, Liu16, CXWW18, LF14a, LXL15, LXL16, ML18]. mainstream [Fan11, LF12b]. maintain [ADS17a, ADS17b]. Maisano [Pra11c]. Major [AEFP16, qJnShPL17, AZSA14, AZSA16, ACAGD^{+}17, CT15b, DWGL16, EGU10, FJ11, FMM15b, KCU19, Med15, MTU17, MÁB18, PFL19, SM16b]. make [BSFCC15, Hag10a, KKV^{+}13, MM16, Men18, YY16]. maker [CW17]. makers [JSZIZ13]. makes [CWJBT10, CMO11, LS17b, NG16, vWWtH14]. Making [CZV10, Kor18, CdMCdMMdP17, pGDTP12, LD16, vdBSS18, CPV14]. Malaysia [KJ13, KJ14, LNK^{+}14a, LNK^{+}14b]. Malaysian [ANA18, LKW^{+}16, TAB13, TABA16, TUCR15]. Male [MB10b, BL11b, Keg15, Kra17]. mammography [THAL15]. managed [HAL11]. Management [CD16a, DLL^{+}16b, ELP11, AEFP16, AL12, ACD14, uHBLKH19, Ben12, BH16b, CJY^{+}15, CST11, CGSS13, CAS16, DCY^{+}17, EBD15, FFR^{+}17, FFR16, GRBBS17, GZ14a, HMCD^{+}19, HNG19, KLP12, KW15, Kos16a, KJ13, Lee15, LPMK17, LGZ^{+}13, LXL15, Liu16, LXL16, LZCZ18, MH15, ML10, MGGdP17, NG16, Oze12a, PMJF19, PHV17, RPGM10, RPDCRVRP15, RPGM16, RP17b, RPP18, RTP17, SDT15, TO18, Tsa11, VT10, WLZ^{+}15, ZAJ19, ZYSS14, DLL^{+}16a, RPK17, SL14]. managers [LTG12]. managing [APPF18, APLHF18]. Mandelbrot [Aus14b]. Manfred [BS13a]. Manhattan [WM19]. manifesto [DF15]. manipulating [uHBLKH19]. manipulation [BK11, YLSW16, YYL10]. manpower [Bas14]. manufacturer [WHZ14]. Manufacturing [FMM15b, FM11a, FM12, LNRSRBB18, MDDG17, TTC17]. manuscript [BM12a, GRSFV19b, KBAK17, SGSS17, SA17]. manuscripts [Ama15, BWD10, BD10b, BWD12, BHJD12, CGG^{+}17, CGK^{+}14, GRSFV14b, Sch11b]. many [Hag10a, HB17a, HB18a, JL14, LL13a, Oos15]. map [BSBG18, HZL^{+}17, KWM^{+}18, Pra10d, WvBvE11, YS13, YLL10]. map-based [HZL^{+}17]. Mapping [AATBPAB15b, ATYL17, BBJS16, CG11, CF14, CP14, FAI^{+}18, FMP17b, GGR11, GTGABAG15, GJ11, GKV11, HMK^{+}12, HGH17b, KGNB11, KG10a, KS18, KKBW17, KGZML^{+}19, Kos16a, KJ13, LM19, LYS^{+}17, LKY17, LG16, LMR16, MdBdP^{+}19, MSB18, NFH12, NLCC17, NA12, Par14b, RAS15, SS10b, SL10, SZ18, SK14c, Tei11, VYL17, WY12, YDZ10, YLL15b, ZZ11, ZL15a, ZB12, dPSS18, BD12b, Boy17b, CIK^{+}18, DC15b, DLMX15, GN19, HLSW18, KD18, KO19, KMP11a, LGD12, LJJ^{+}16, MLT^{+}15, Moh12, MAGSTRC15, Oli15a, OMCP17, SMF18, SBSU15, Soó11b, SK12, SDEB16, SB19, USPO15, ZLH^{+}15, vEWNB10, vEW10, KN15]. mappings [LCS^{+}16]. MapReduce [HAG^{+}16]. Maps [Sma10, dSAEE15, IS16, LCR13, LKR14, LBW17, MLY^{+}14, Sma11, TW10, YPK13]. March [BMM14]. marine [dCdSNB15]. markers [dBONM^{+}19]. market [CO10, CSC13, CSC14, CLW^{+}19, CFM15, FT19, GRSFV14a, MR18b, NF13]. market-oriented [CFM15]. Marketing [CMPD19, BHPVdPMR18, MAGSTRC15, Saa10, VT10, WZW15]. markets [CMRC15, FS10, PN15]. Markov [AHP17]. Martin [LM13a]. mass [GALR16, KB11a, SKM15]. Massacrator© [PLT14]. master [Jac18]. masters [BR12]. match [ATJ16, AT18, RGLE16]. matching [ZC16]. material [vLCCMV13]. materials [KSSB13, LLLL18]. MATH [CB15]. Mathematical [Egg11c, FSSPG^{+}15, BD13, BFM^{+}14, KM18c, SIR^{+}14, Zuc10]. mathematicians [DRS14, ZCKZ16]. Mathematics [BL11a, uARA19, AA16, BD12b, Bra12a, GLS16, GQAM19, HG17, LD16, SL12b, ZT14]. Matrices [RRBA10]. matrix [HH17e, YHL19]. matter [HSBW10, LVSL18, Mar11, NG16, TH19, WHZ14]. matters [Har16b, Yu17]. Matthew [PD10, YGW^{+}15]. Maturity [RMdO17, KHVGA^{+}16, LNRSRBB18]. may [BBJS16, Cop19a, SA17]. me [KB12]. me-too [KB12]. mean [Laz10, TG17, ZCL14, ZCL15]. mean-based [ZCL14, ZCL15]. meaningful [BB15]. meaningfully [BM14b]. means [SYDW19, dSNV18]. measure [ADS10b, AD14, AW10, Ama18b, BT17, BT18a, CNPG17, CLHH10, CC12b, CC11b, CB16, EGU10, Har16a, Har16c, HS16b, LPL16, Liu17, Lun19, Pei19, PMN16, RZ12, RKT^{+}15, Saf13, SS10a, Sch10a, TB19a, YWZ^{+}17, ZLF^{+}14, ZLZ19, ZTRH18, Zhe19, dSFSF15]. measured [BM14a, GM12, GK18, JZL10, VSS12]. Measurement [Rha17, ASW18, BTL19, ET15, LTG12, MARMSG19, Osw10, PPM^{+}17, RBC^{+}10]. measurements [HT19, dCCMAW16a, dCCMAW16b]. Measures [ER19b, Moe10, SA16, Abb13, BLS15, Cab11, CFSSP16, CM18, DLGP16, DGDGSV15, FSSPG^{+}15, GRSFV^{+}13, GPN14, GKK15, GLM11, Hal13, HK12, Kim10, KJES16, LWB18a, MvdH13, MKHB13a, MKHB13b, RPNC13, SS16]. Measuring [ADR13, ALYZ15, BMD^{+}18, Car16, CKT17, CJW10, Cla15, Das16, DL16a, DL16b, DNAH15, Din14, EGR13, GRTPMLAJ19, HT11, HH13, HIG^{+}17, HWL11, HCS^{+}15, Hun12, JDLIV14, JK10b, KP12b, KB18, KS17, LZ14, MG12, RGLE16, RVFEdlM10, SR16, Sch13b, SV19, SN10, tScL13, Var11, Wal15, WWH^{+}17, WRC^{+}19, ZZLS19, ZZZC16, ZHMX14, ZL17, ZG17f, ADV13, ATM16, ACAGD^{+}17, BHM16, ENST16, FESD11, GSMT10, HSAK18, KLL14, KB10, LPL14, Nic14, PB17b, PB12, RW11, Sch15c, SPS14, Sni16, SML16, TSRGG17, Wra18, ZQH^{+}17, ZNB^{+}17]. mechanical [JC12]. Mechanism [PG14a, Gál17, San12a, San12d, ZZW14, ZZW16, ZZFD16]. mechanisms [CFM15, Hos11]. Medal [Ano12c, Ano14, Ano17a, CXZ19, Dan19, Fie15b, KL16]. Media [AMMT16, CHC17, CD18, CyPP12, CP12c, CV14, ENST16, GW15a, HIG^{+}17, HMI19, Kha13a, Kha13b, NTM^{+}18, RCCM14, SMM^{+}19, VSK18, WLMF15a, WLMF15b, WFG16, ZG13a, ZW17b]. Media-based [AMMT16, Kha13a, Kha13b]. median [Egg10c]. mediating [EMSH16]. Médica [CRLMLM17]. Médica/Archives [CRLMLM17]. Medical [AKB^{+}10, CRLMLM17, ESB15, GWBSVWB13, Soo10a, ZGCRVQ18, BCML19, CdJD15, CMM17, ES18, FJ11, FI16, HE16, KM15b, KM15c, KM15d, KDFL14, KGB^{+}18, LJ10, MdNS^{+}19, Pau10, PDAN19, PKSG12, RWG^{+}15, SMAABJ11, SFR^{+}19, TKA17, VVN16, VNA16, YWW17, ZG13a, LCS^{+}16]. medicalized [SSdOS17]. medicinal [SSdOS17]. Medicine [KHK13, BMM17, BGMB16, Dan14, Fan15b, FCFG17, FZZ^{+}11, FZZ^{+}12a, HFW^{+}14, HZD^{+}15, LNK^{+}14a, LNK^{+}14b, MLY^{+}14, OCM^{+}12, STCRPA18, Fie15a, OCJB15, DD18, WL18]. medium [MASM14, MASM16]. Medline [ZCMVQS11]. Meeting [KÖG12, KG13]. meets [Glä15, Lei16, ZG12a]. mega [EdS19, MA19]. mega-journal [MA19]. megajournals [Hen19, Hen20]. Mekong [SCLC15]. mellitus [SZAJS14]. member [Lia11, MM16, WWP17, YP19, ZL18b, ZZW14, ZZW16]. members [ALH15, BHJD12, KS17, MHM^{+}13, OBG11, RGLE16]. membership [XLZ^{+}18]. membrane [CYK^{+}11, DSG^{+}15, ZPG^{+}14]. meme [ZXZ17]. memes [AF18, SD18]. Memetic [HGH17a]. Memorial [GSB18]. Memorias [STCRPA18]. men [HC14a, HAJ12, KPS12]. Mendel [CvLvR11]. Mendeley [ATK17, Ask18, BI14, Eld19, KCT^{+}17, The17a, The18c, The18f]. mental [BJIB16]. mentee [CZPR17]. mention [PYW18]. mentioned [BHM16, HB17a]. mentions [ST14c]. mentor [CZPR17]. mentoring [CZPR17, DY18]. Merck [Tom17]. merger [PYK13, SRF16]. mergers [DEC15]. merging [SKCK14]. merit [ADR14b]. MeSH [LCS^{+}16]. Meso [TABA16, Zit15, CvLB10]. Meso-level [TABA16, Zit15]. mesoscopic [VuHL10]. messaging [ZC14]. Meta [GBGB13, Bor15a, EBD15, Mue18, VHH16, XGCK19]. Meta-analysis [GBGB13, Bor15a, Mue18, XGCK19]. meta-ranking [EBD15, VHH16]. metabolism [LPZ17]. metadata [BSG17, ByLbH16, QA19]. metaknowledge [ZG17a]. metal [WH16]. metal-organic [WH16]. metaphors [BSBG18]. method [ADR19, AHP17, Bel17, BM15, BG17, BYY17, ByLbH16, CCLL14, DXL^{+}18, DLMX15, FZZ17, HSPY15, HZQ^{+}17, IMSK14, KGY^{+}17, KKOS19, KB10, KKL14, LJS16, LNRSRBB18, LZFW15, LSS15, LSS16, MMOMLC18b, MM15a, Osó18, Osw10, OI17, RNB19, SR16, SSG^{+}18, SCL12, Var11, WBH^{+}12, WCL14, XHA^{+}19, ZCL14, ZCL15, ZZD^{+}18, ZZP^{+}14b, ZLLD19, dZLwC^{+}15]. Methodi [PKR15, dCPRP18]. Methodological [AAB^{+}13, PB17b, CT15b, GMJ^{+}17, LGL10, OBG11]. Methodology [Bha11, BM13a, BM13b, CKB^{+}14, EMSS16, FCCMTRVR18, FMP17a, HL13, MD12, PKR15, RCN^{+}14, TSMTDLCH11, Tsa11, Tsa15, dCPRP18]. Methods [OMAMMLC15, SJ10, AUS12, BM14b, Boy17b, CMM17, CvLB10, Din14, ES18, GT17, HY19, ILB13, JTZ14, JPZ^{+}10, LZB10, Lor14, MM14c, MSH16, OROMAA16, SB15, YYDH12, YWC12, ZZP^{+}14a, ZCL15, ZZZ^{+}14, ZGJ18, ZHMX14, Zit15, dCdSNB15]. metric [LLH^{+}16, MS18b, PQG14, PDAN19, PPM^{+}17, RZ12, RG18, ZCW14, ZC16, dSM17, dS18, dSdSSB16]. Metrics [PMJF19, Pol16b, VB12, AD16, AP16, ASW18, Bor15a, Bou14a, hCcTmWH15, DdS19b, EMSH16, EG16, GG15a, HCS^{+}15, MBL18, ML13, MCR^{+}12, Pra12c, Pra18b, Shu17, TYYW16, WMXZ14, WWC19, dS17a, OMLC14]. metrics- [Shu17, dS17a]. Metrics-based [PMJF19]. metropolitan [LSR13]. Mexican [CRLMRPA10, LOMLPA^{+}17, LOPAGS19, RHGKD16, VAJCC17]. Mexico [AD18, CRLMRPA17, FVVSGM^{+}18, LBCO19, LOMLPA^{+}17, MC13, USPO15]. Michael [GZ18, KL16]. micro [CBF13, CB11, ZWW^{+}16]. micro-level [CB11]. micro-macro [CBF13]. micro/nano [ZWW^{+}16]. micro/nano-bubble [ZWW^{+}16]. microbial [MYN^{+}15]. microbiology [dSTL18]. microbiome [Coc18]. microplastics [dSTL18]. microRNA [MM14a]. Microsoft [Har16d, HA17a, HA17b, Har19b, HHBB18, HOB17, HB17b, The18a, The18e]. Middle [Cav16, GNS^{+}15, SZAJS14]. might [uHBLKH19, Har15b]. migration [BD12b, Dya17a, Dya17b, MAP13, MH14]. milestones [AUS12, CL17b]. military [ACMP13, LS17a]. millennium [Kra10]. million [Cab13, Jac18]. Mimicry [Bor11]. minds [CT15a]. mineralogy [FVVSGM^{+}18]. Mining [CLLH15, Fia11, SBA^{+}19, WYY11, ZPC17, Com15, EDEH16, GRBBS17, GWS15, HD17, JS15, Lor14, MVS10, Moh12, PYK13, PC18, RHMH17, RNF19, SK13, Tsa11, WRM17, WNS13, WA18a, YLL15a, YLSW16, ZLG^{+}15, ZLW19, ZZZC16, ZHZY19, Mad15]. Ministry [CP12b]. miracle [RT17]. mirror [DFS14, VHD^{+}16]. mis [LZR14, CL11, Glä18, LHC16]. misattributions [Kra17]. Misconduct [ANA18, BH18a, ZG13a]. misinterpretations [Sch15a, Wu18]. Missing [Kra19]. mission [aSTS17, WHC^{+}13]. mission-oriented [aSTS17]. Mistakes [CGK^{+}14, ANA18]. Mistaking [Etz13a]. Misunderstanding [Har13c]. mix [Sch15a, Wu18]. mix-up [Sch15a, Wu18]. Mixed [GWB11, VEJC^{+}18b, XLZ^{+}18]. Mixed-indicators [GWB11]. mixed-membership [XLZ^{+}18]. mixing [PR10]. Mladenić [Rod16]. MNCs [SL17]. mobile [Ama18b, LK17, PFPCM^{+}19, KK15]. mobilities [YY14]. mobility [AD18, ACRC17, Ama18a, BD12b, BS15c, COS11a, CRBRG^{+}18, CR18, DRS14, FSO11, FSOS12, KA17, TV17, WD13]. mock [Pra10e]. modal [GF11]. modalities [GBDG19, RCdJ^{+}14]. mode [KM12]. Model [LSC10, ACD15, AHP17, Aus14b, Bas14, BSS15, BBL17b, BGBS18, Bou11, BD13, BS13b, Cha17a, DFS15, Egg10c, FCFG17, GG12, GLS16, GZGAC16, GZGAC17, GSM^{+}16, GHT16, GWB11, Hei19, Hos11, HWQ^{+}18, HWS18, IL14b, ISR11, Ke13, KHK13, KPRT16, LWB18b, LQW17, LW10, LGD11, LGS18, MH16b, MB10a, MB13, NG16, Nic14, NZL^{+}19, OMOR13a, Par15, Pra17a, RT17, SHR^{+}10, SS10c, SKCK14, SGY15, SDP^{+}19, VEJC^{+}18b, VEJC^{+}18a, WYAY12, WLZ^{+}19, WHS19, Xie19, YLH18, YKLK14, Yoo15, YL10, YWZ^{+}17, ZFY^{+}17, ZG17b, ZYF^{+}17, ZMW^{+}18, ZSC18, ZLZ19, ZLLD19, ZG17f]. Modeling [Ano11, BR12, BGSvdB11, Cha17b, GC10, KHS^{+}15, LJC^{+}15, SLG10, WFH^{+}16, WG10, XXL^{+}17, ZG12b, AUS12, BVOL18, DGPL15, FS12, FMP17b, GDP16, GF11, HFL14, LKSK15, MHM12b, SLXD15, WZCC19, YPNS14, ZYNZ18]. Modelling [GAE15, ZKC^{+}16, AF15a, AF15b, BGM17, KTRP17, MFF^{+}16, MKF14]. Models [Jon10, CLLH15, CB16, DdlPPL^{+}19, FLM^{+}19, FA10, FAA13, HII^{+}18, HV18a, JL18a, KK13, LDVSGD19, LWT16, MMSS11, MM15b, NH11, Par15, PROG19, RMdO17, Wal15, JL19]. moderating [LLHN17, WZ19a, YC10, ZW17a]. moderation [WY19, ZQH^{+}17]. modern [CRLMRPA17, FVVSGM^{+}18]. modes [JCK11]. modification [LF12a]. Modified [SS10a, cTnHwH17, ByLbH16, CLHH10, JBC19, Rou11]. modularity [TK10]. Moed [Pen19]. Molecular [CAGL15, RBF^{+}10, Jon10, dBONM^{+}19, TN19]. Momentum [GNHT18]. monetary [ZS18]. money [ACD13, Cop18]. monitor [CWL10]. monitored [OHT10]. monitoring [BSC^{+}17, EMSS16, GM12, TW16, YLJ^{+}17]. monograph [Sni16, VO17]. monographic [GTC16, SLGO17]. monographs [Ham11, ZC16]. monopoly [Fan11, LF12b]. Monte [Sch16]. MOOC [GZ17]. Moravcsik [GZ18]. Moroccan [Bou14a]. Morocco [BDE11, Med15]. morphology [WMH^{+}17]. Most [Ioa06, Fan12, GdOdAG^{+}13, Kor18, Ley12, LDVSGD19, MHTB17, PLA10, PAL13, Pra10a, RV18a, RV18b, SCGZR16, SAR19]. Most-Cited [Ioa06, SCGZR16]. most-highly [Ley12]. motion [Lja16]. motivate [MCL13]. motivations [HL13, LNMQRR15]. motives [ONB17]. motorcycle [OMJLVSN19]. move [RM18]. moving [DC19, Yoo15]. MS [CCLL14]. much [uHBK19, BSFW10, GRSFV16b, LVSL18, OM11, Pra10b, Sch13a, Vil10]. Multi [GF11, LLX^{+}18, MLT^{+}15, XYW^{+}17, AHUR11, BPJ^{+}14, Ber18, BD13, DLMX15, GHvdB12, HS16b, wH15, JC19, KL17, KO18, LWIB16, LGD11, LF12a, MXZ18, NPT^{+}15, NH11, Osó18, PC18, PPM^{+}17, Pra18e, RKT^{+}15, RPK18, SP14, TH19, VRF12, You14, ZZZC16]. multi-agent [NH11]. multi-aspect [GHvdB12]. multi-attribute [BD13]. multi-author [Ber18, wH15, TH19]. multi-authored [KO18, LF12a, Osó18]. multi-authorship [NPT^{+}15, RPK18, SP14]. multi-classifier [MXZ18]. multi-dimensional [HS16b, JC19, Pra18e]. multi-faceted [DLMX15, PPM^{+}17]. multi-granularity [ZZZC16]. multi-industry [VRF12]. multi-institutional [You14]. multi-level [BPJ^{+}14, LWIB16, PC18]. multi-levels [AHUR11]. Multi-modal [GF11]. Multi-source [XYW^{+}17]. multi-stage [RKT^{+}15]. multi-technology [KL17]. Multi-view [MLT^{+}15, LGD11]. Multi-views [LLX^{+}18]. multicountry [WDP11]. multidimensional [GRSFV12a, GRSFV^{+}12b, SSG^{+}18, TSRGG17, Tor14]. multidisciplinarity [ÁBV^{+}14, CBWJ18, MRGT18, SVS18]. multidisciplinary [ADD18a, BH10, DGGBDG17, GTD14, KB18, Lee15, MMOMLC18a, ZS11]. multifaceted [SP12a]. Multilevel [SC10, BS17, OMOR13a]. multimedia [TH13]. multimodality [Xie19]. multinational [HEH18]. multinationals [WY12]. Multiparametric [VAJCC17]. Multiple [BI18a, Yos13, dSD18b, dSD18c, Abb11, FDVZ16, Gal11, Har15b, Hir10, HL17, LdZwC^{+}17, MR18a, Pra11b, SGM^{+}16, Sch11b, SK12, TG16, YG18, YWZ^{+}17, ZWL^{+}18]. multiple-link [YWZ^{+}17]. multiple-part [Sch11b]. multipolar [Veu10]. Multivariate [dPdCAdMC^{+}16, ACD15, CLL^{+}17, FCFG17, OA10a, Tod11, VHG^{+}15, dCdSNB15]. municipal [Hol10, YCL^{+}13b]. museum [ACAGD^{+}17]. music [Geo17, GN19]. musical [KCP12]. must [DDR17, Sch14a]. Mutual [KW17, KZSZ19, LPL14, SN10, Zhe19]. mutually [QZZD18, YWZ^{+}17]. my [LNMQRR15, Per18].

n [ATJ16, ZLF^{+}14]. n-grams [ATJ16]. Name [GKF17, SHB14, AT17, ADD16, ARE^{+}18, CLMN19, HYF^{+}17, HYYR14, Kim18, KK18, Kim19a, Kim19b, KKOS19, McC18, MRR17, Sch16, SKCK14, TW10, WBH^{+}12, WD13, ZYX^{+}14, ZWL^{+}18]. named [Ken18]. names [AOd15, AP14, DRS18, FGP13, SSdOS17]. Nano [BS15a, SK14a, GHT17, KPY16, LG15, WNS13, ZT18, ZT19, JAAA18]. Nano-Bio-Info-Cogno [JAAA18]. nano-biopharmaceutical [ZT18, ZT19]. nano-bubble [ZWW^{+}16]. nano-enabled [KPY16]. nano-energy [LG15]. nano-level [GHT17]. nanobiopharmaceuticals [ZG11]. nanocellulose [MNdF16]. nanofiltration [ZPG^{+}14]. Nanomaterials [dSSdMAF14]. nanomedicine [BAB13]. nanoparticles [LCLX16]. Nanoscience [GB14b, Ley13a, YWY10, BYR13, CRMPRS18, HMCL16, KGNB11, MDG10, Moh12, QZZD18, SW19a, SK14b, SDEB16, ZG13b]. nanostructures [ÁRM13]. nanotechnological [ÁRM13]. nanotechnologies [Fin11, SJ10]. Nanotechnology [PLG19, YWY10, APYS13, BYR13, BK10, BSB12, CRMPRS18, CC13, DT16, EGUB12, FSAB10, GHS18, GB14b, GW10b, GS12, JKPL18, KGSS16, KGNB11, Ley13a, LJC^{+}15, LZC17, MAA^{+}11, MHM^{+}12a, Men18, MdFdA^{+}14, OZK11, OPGW^{+}13, OI17, SYP10, SK14b, SDEB16, TBS15, TS11a, TS11b, Ter17, WS11, WXLL12, ZG12b, ZG13b, ZyZZ^{+}14]. nanotube [CWL10]. nanotubes [EGUB12]. Narin [Pra19e]. NASA [GG13, TA15]. National [ADS10b, ADD11b, ADD11a, AD11b, ACD13, ABL17, CLD13, CYW^{+}11, KA17, KM15c, Kaz14, Kim14, Nic14, Zin16, vWBS^{+}16, ADV11, ADD11c, ADD14a, AD16, AÇA^{+}14, BBDS^{+}14, Bas14, BBSS16a, BBSS16b, BBSS16c, ÇAAÇ15, Fed13, FKM^{+}15, GE11, GTD14, HMCL16, Hos11, JK19, Kaz15, KR17a, KPSL12, Li17, LDVSGDR16, LDVSGD19, MBP19, MWH14, OA10a, PY14, Pra16b, RGTSLCH14, RLW14, dSSdMAF14, SBT18, SvLVA19, Shi11, SK17, SN10, SG16, WMW^{+}13, Das16, KPSL12, RGdCMM17]. National-scale [AD11b, ADD11c]. nationalist [AC12]. nationalities [GCGP10]. Nations [KRR14, GG19, GHA^{+}16, GNS^{+}15, HLL14, KJ14, NASR11, Pac19, Pra17b, Pra18e]. native [GSM^{+}16]. Natural [KHK13, PTMT10, BM14b, Cha13, Dya14, Egg11d, HAL11, Pau10, RPK16, SVS18, SW19b, XTZ15, ZS18, ZM16]. Nature [CG15a, BBP14, CLD13, CC10b, GGW^{+}13, HZ17, Lee15, VZAMG19, Vel12, ENA19, HB15, BBP14, BHDI18, CIL^{+}16, DAYY18, Emm19, GRSFV18, KA17, LHW16, LLX^{+}18, SRW^{+}15, WvBvE11, Waa13, WMXZ14]. Navigating [KPY16]. NBIC [JAAA18]. NDGM [JL19, JL18a]. near [KH17, MM15a, WRM17]. near-term [KH17]. nearby [SS16]. necessity [BS13b]. need [Koz15, Lin10, Shu17, dS17a]. needs [Ano16b]. needs-state [Ano16b]. Nefrologica [Kra19]. Negative [Fan12, Bal12, GGW^{+}13, Par14b]. Neglected [VSS12]. negotiation [WK15]. neighbouring [Vil10]. Neophilia [PB17a]. Nepal [Gau17, GB12]. nephrology [APFR^{+}13]. nepotism [ADR14b]. nest [HA17a]. Nested [dS17b, HAL11]. Netherlands [CAV^{+}19]. Network [CB15, GG12, KMP11a, NQ14, OA10b, PW13, SFBS17, WhCL10, WWP17, AJSN18, AP16, BB10, BHB13, BHA15, BAB13, BHKP11, BPJ^{+}14, BD12b, BFS17, BSBG18, Bor16, BFM^{+}14, Bue15, BSKB17, CD17, CD14, CH13a, CH14, CLL^{+}17, CZPR17, CC10a, CY13, CS11b, CL17a, CYK^{+}11, CU16, DDS^{+}19a, DT16, DAMC15, DLGP16, DCY^{+}17, ÉMS^{+}13, ET15, EES13, FMP17a, FR11, FSC14, FDVZ16, GLS16, GKK15, Glä12, GBHT16, GP18b, GLM11, HKWC15, HZ17, HLW19, HSK17, HH17e, HW10, JBC19, JC19, KLPP16, KP12b, KCK14, KLCS14, KD14, KTLD16, KKK^{+}14, Kos16a, KM16, KM18c, LPC17, LSC10, LSCK12, Lee15, LJS16, LK17, LYS^{+}17, LZFW15, LZR14, LW15, LSY11, LX15, LM16, LZC17, LCIADG19, Mad15, MB19, MJHG13, MCvFP16, MPH19, MGGdP17, NRAW17, NSKO15, OMA15, OA10c, PR10, RM10, RRSA18, RKT^{+}15, RPGM10, RPGM16, RTP17]. network [SBA^{+}19, SHL15, SÁV18, tScL13, SL16, SYLC17, Suo14, UHAR12, VYL17, WLR^{+}14, WY19, WK15, WKK16, YDZ10, YLC18, YPH10, YLL^{+}15c, YLH18, YCK11, YLH^{+}17, YSD11, ZYSS14, ZLF^{+}14, ZCZ^{+}16, ZLL^{+}17, ZRL18, ZSC18, ZLW19, ZZW14, ZZW16, ZZW19a, ZyZZ^{+}14, ZGL^{+}17, ZGJ18, ZHZY19, ZG13c, ZWHH13, BHL^{+}10, BRS^{+}16, The18a]. network-based [AJSN18, CZPR17, KD14]. network-level [SBA^{+}19]. network-related [ZLL^{+}17]. networking [DLMX15, IQT^{+}19, MDFGAM14, MCB15, WZ19b]. Networks [ÁCCG^{+}15, dSFSF15, AHUR11, Abb13, ATYL17, ANOdFC12, AOd15, ÁRS17, BLdlCV17, BC13a, BHL^{+}10, CGC18, Cha17a, CBF13, CF14, CdJD15, CRFM^{+}12, CK14, CIK^{+}18, CQB16, CKB^{+}14, CR18, FB10, FK17, FK18, FSO11, GLS16, GA18, GAE15, GTGABAG15, GF11, GSPLVG^{+}18, HSL^{+}14, HD17, HWL11, HSK18, HRC12, Hur17, ILB11, JvGH10, JBC19, KLM16, KLP17, KKLP17, KO19, Keg15, KD19, KJES16, KKK^{+}14, KdBBK15, KGG15, KJ14, KR17b, LMM15, LGL10, LGPC18, Ley11b, LWB18a, LHCH18, LLLL18, Lia11, LXDL13, LGD12, LG15, LJC^{+}15, LJMF15, MY16, Med18, OZK11, Ort11, PYL16, PROGMA10, Pra19b, QDY14, QA18, RPP18, RF19, RBC^{+}10, SMF18, SP12b, SSN19, SJ19, SD18, THAL15, TK10, VuHL10, VAJCC17, WZX11, WCL14, WZW15, WY19, WZ19a, WDN17, XCS^{+}16, Xie19, YDJ12, YG18, YC12, YK11, ZLYF14]. networks [ZZ16, ZZLS19, ZZFD18a, ZZW^{+}19b, Zin16, dPSS18, vdPR18]. netzwerk [BHL^{+}10]. Neumann [Fie15b]. neural [CC10a, LHCH18, MPH19, SYLC17, VAJCC17]. neurobiology [Bru10]. neurological [HFW^{+}14]. neurology [IPIU13]. Neuroscience [XG18, AA19, Bru10, BJIB16, KGZML^{+}19]. Neurosciences [HdSV16, APFR^{+}13]. never [BA15]. never-ending [BA15]. NEViewer [WCL14]. newcomers [SB17]. newly [ZZPG14]. NewQIS [GKB^{+}19]. News [LT10b, AM18, ZG13a]. newspaper [Fan13a]. Newton [HT19]. next [HB18b, YQW13]. NHST [Pat18, Sch18b, Pat18, Sch18a]. NIEs [WG12]. night [WM17]. night-light [WM17]. NIH [ZYNZ18]. NIHR [GWG17]. Nitrogen [GG14]. No [BBP14, MRLW15, TN19, Wra16a, AAB^{+}13, Wra16b, ZLF18]. NOAA [Bel13]. Nobel [BOS14, Bjø19, CÖT15, CT15b, CÖT16a, CMT18, CZPR17, Cha17b, Fie15a, GW10a, Har13d, HR17, MSYW12, SBB16, TA17, ZZY19]. Nobelists [ZXLX14]. node [dSAEE15, SKCK14, ZLF^{+}14]. nodes [ZLW19]. noise [BSFW10]. nominees [GW10a]. Non [LRZ13, ACD11, ADD14b, CGV12, Cha18c, Chi14, GSM^{+}16, GAVZAB12, GW10b, HWS18, HC19, LL10, LBA19, LCFC14, LHTL18, Moe16a, MT13a, NvLvR10, OVJM17, Sak19, SRW18, Shi14, SCL12, VE14, ZS18, ZG13a]. non-auctioned [LL10]. non-citation [HWS18]. non-competitive [ACD11, ADD14b]. Non-English [LRZ13, LHTL18]. non-English-speaking [GAVZAB12]. non-exhaustive [HC19]. non-experts [Moe16a]. non-institutional [SRW18]. non-inventing [GW10b]. non-journal [NvLvR10]. non-LIS [Cha18c]. non-litigated [SCL12]. non-local [GSM^{+}16]. non-medical [ZG13a]. non-monetary [ZS18]. non-parametric [LBA19]. non-patent [CGV12, LCFC14, Shi14]. non-peer [VE14]. non-peer-reviewed [Sak19]. non-source [Chi14]. non-standard [MT13a]. non-university [OVJM17]. none [Fan13a]. noninferiority [DD18]. nonlinear [CC10a, CC10b, WFZD19b, XZFD19]. nonparametric [CMVP16]. normal [Wra18]. Normalisation [BW19, GSTD11, MSA13, BW20]. normalised [The19b]. normalization [CA19, HB18a, JJS^{+}12, PYL16, Pra12b, PG12, ZCL14, ZCL15, Zit11]. normalize [BH16c]. normalized [ACP12, BYY17, BL18, BYY18, BTL19, DGDG13, MZE19, WYvE11, WvE13]. Normalizing [BB16, MM17b, ATYL17, MM17a]. normative [EO14]. norms [Sot10, YB14]. north [ADR16b, KHK13, Med15, PY19, SL13, ZCKZ16]. Northeast [KHJ^{+}12, CdSPdM13]. northern [MWH14, LMM15]. Norway [KG10b]. Norwegian [ACP12, HS16a, San18]. notation [Cav15a]. Note [AZSA16, DR10a, HC17, MSDJ19, ST14a, Wra18]. notes [LLH19]. notions [DFS14]. novel [FE16a, FE16b, HIG^{+}17, HSAK18, HB18b, JDH12, Li15, Li18, Mag14b, MMOMLC18b, SH19, SZ12, WOW10, YHL^{+}18, ZZD^{+}18, ZLLD19, ZWL^{+}18]. novelty [GM12, LS19a, WRM17]. Novosibirsk [MLVJ12]. NPR [LCFC14]. NRAEs [CLD13]. NSFC [YHL^{+}18]. nuclear [HYS18, KNK^{+}19, Pri15, Pri16a, Pri16b, YIK^{+}10]. nucleation [San12a, San12d]. Null [Sch15a, Wu18]. Number [HBDL18, ACORC11, Bas10, BPVM11, CCM^{+}11, CR14, Egg10d, FM17, GA18, GMSZ18, HHBB18, IBL13, JN11, KKS^{+}17, Lev15, MBR^{+}13, PKR15, Sch10b, TAA16, ULFRU^{+}14, XGL^{+}19, YHL19]. numbers [Abt17, Fan18, LABL13, TC13]. numerical [Ric15]. numerology [Etz13a]. numerous [Sch15a, Wu18]. nurses [BL11b]. nursing [DBO^{+}18, WOW13].

o [CL16]. OA [Ase10, BS15b, SGY15, Zhu17]. Obituary [Ano17b, Ano19, Cze11]. Object [BC17, CYW^{+}11]. objective [KB10, WM19]. Objectivity [SM17, Fed13]. obliteration [McC14]. obscure [RMA12, The17b]. observation [Tod14]. observational [DGD19]. observations [KHR^{+}19]. observatories [SÁV18]. observed [Pra19b]. obsolescence [SMY15, SML16, WZCC19]. Obstacles [Wad16, Wad17]. obtained [Cam11]. occur [LS17b]. occurrence [BHH18, LSS15, LSS16, LM16, LCIADG19, QDY14, SL10, ZLH^{+}15]. Ocean [Bel13, SH15c, WTM^{+}16, Vil10]. Oceanography [SP12a]. October [Ano11]. odyssey [HC15a]. oeuvres [RCN^{+}14]. off [FRPP17, LHLH19]. Office [ACT18, Bel13]. offices [Wad16, Wad17, Wad18]. Offshore [SL17]. often [BHM16]. OIC [KJ14]. oil [FP18, ZLH^{+}15]. OLAP [LJMF15]. old [FMM16, HA17a, MM16]. older [LS16]. OLED [hHC15, HC14b]. Olle [Ano12c]. omitted [FMM15b, FMM16]. Oncology [BCC^{+}17, RLW14, SMM^{+}19, SG10, YSD11]. Oncotarget [GFK^{+}18]. One [BVZV16, hHSL19, dB12, BM12a, HC16a, HA17a, HB18b, Ley18, SRW18, Vin12a, dW15, BHDI18]. ones [ADD11d, Har16b]. Online [FZQ17, GP15, KCT^{+}17, CYT^{+}12, CyPP12, HP18, HV18b, LTG12, Mou16, OMA15, VSK18, WST14, ZZZC16]. only [dB12, EdS19, Fan13b]. Ontology [DLL^{+}16b, LSC10, DLL^{+}16a]. ontology-based [LSC10, DLL^{+}16a]. Open [FL16, GOPG13, JN15, LP18a, LPB14, Wra16a, Wra16b, AOFU10, Asa19, Ase10, BND11, BWD12, Cop19a, DGGBDG17, EdS19, Fuk17, GFC18, GRSFV17a, GBDG19, GWA14, HAG^{+}16, HPS19, Hen19, Hen20, JMM18, KPJT14, KVC15, Laa14, LL16, MA19, Mik17, OMLC15, PB18, PTMT11, PP16, Sni16, SE18, yTwTlW19, VZAMG19, WLMF15a, WLF15, WLMF15b, WZCC19, YXW18, ZW18a, ZW14, ZW17c, Zhu17, MSB18, PROG19, ZW19]. open-access [YXW18]. OpenCitations [HPS19]. Opening [RRLNAG15, Sku19]. openings [CÖT16b]. openness [Cop19a, DLL^{+}16b]. operational [Lee10a]. operations [ZAJ19]. operator [Yu15]. ophthalmology [ON12]. Opinion [GM13, Bue15, Fan11, LF12b, Sob11, YSM^{+}19]. opinions [BH15]. Opportunities [ZLL^{+}15, KPL19, MP15, TBMM18, WPCG13, WMH^{+}17, YK12]. opportunity [KGL^{+}14, KGY^{+}17, LKS^{+}14, LL19]. Opthof [Hir19b, Pra12a, WvEvL^{+}11a]. Optimal [LGD12, Luo12, PRSB16, GRSFV19b, Lev15, SY16a]. optimization [HEH18, KCP11, RJ14, ZCL15]. optimization-based [ZCL15]. optimize [BD13]. optoelectronic [SHR^{+}10]. optoelectronics [Lam12]. OR/MS [CCLL14]. Orange [HFC11]. ORCIDs [YCPS17]. order [ADR13, CA12, CB11, Egg11a, GA18, JX13, MBSB17, Ora17, Pra13, Pra19d, The19a, TSG13, Wal16]. Ordinatio [PKR15, dCPRP18]. organic [AATBPAB15b, hHC15, WH16, ZXH10]. organisation [BL15]. organisational [CGC18]. organisations [GBM^{+}16]. Organization [Har13a, CP12b, DCY^{+}17, HEH18, HFL14, KdBBK15, Ley11b, YC12, dSF13]. Organizational [GLÚGML16, VHG^{+}15, uHBK19, BK15, BSvEK13, BS13b, FK16, GRSFV14a, GSE^{+}18, KFKS15, LVHS^{+}15, LZ14, RGCM14, VDV16, ZCZ^{+}16]. Organizations [LOPAGS19, Ano16c, BvdB14, CLMN19, MK18, WKK16]. organized [LWB18b]. organizing [dSAEE15]. orientation [BH17a, Won19]. Orientations [VNA16]. oriented [CMRC15, CFM15, FYC15, LSS15, LSS16, aSTS17, YZW^{+}17, ZSC18]. Origin [MLT^{+}14, MB14, PCR18, Vel12]. original [DBO^{+}18, DF15, MBR^{+}13, PFL19]. origins [AMMT16, RLW14]. ornithology [YWG14]. orphans [Jon10]. orthopaedic [SdJDD19]. Oswaldo [STCRPA18]. other [ASPF^{+}16, BHKP11, BBJS16, BBL17b, BTL19, ES16a, Egg11c, EMH^{+}10, dCPF14, GK18, GdOdAG^{+}13, GdA14, MRGT18, Mor19, OM11, Pra17a, SM17, SN10, The16, Tod11, Zhu17, dAG13]. Othmer [Tom18]. outcome [VNA16, YIK^{+}10]. outcomes [BF17, ZYNZ18]. outdated [Van12]. outlier [Glä13, KGY^{+}17, YK12]. outliers [GPSM18, Pra14b]. outlines [GBM^{+}16]. output [AAG14, ACHVH10, AJdMA10, AJCACRdMA16, ALvH19, BYR13, BBVO10, BM11, CRZGVQMA15, DB19, EN17, EGUB12, FMPP10, FM11d, FGMM12, FLB19, GRTPMLAJ19, pGSyW^{+}19, GCGP10, GRSFV^{+}12b, GK14, GSKM17, GYZ15, GHA^{+}16, GKV11, GB12, GHA^{+}15, Hir05, Hir10, HB17b, ILP13, ILBG14, JL18a, JL19, JKSK15, KG10a, KKBW17, KHK13, KBZS15, LLCL11, LKW^{+}16, LGS18, MM14a, MBL18, Med15, MT15, MGB16, MMA18, MAGBBM13, Mue16, NP11, OM11, Pan14, PSZ15, PS16a, PSB^{+}17, PL17, Pra11c, Pra18e, RASP13, San12c, SB15, SW19a, SK18, SY16b, SZAJS14, TLSH14, Van14, WLDW12, WTM^{+}16, WM17, YCL^{+}13b, Yur15, ZGCRVQ18, ZHG16, ZL18b, ZZ11, ZM16, ZP16, vLvWW16]. output/growth [JL18a, JL19]. outputs [ACC^{+}16, APYS13, BDE11, COS11b, GP18a, HH10, ILP11, Sal17, YCPS17, vZ13]. outstanding [PLWS14, QZL^{+}17]. over-citation [BL17a]. over-competitive [Fan11, LF12b]. Overall [GRSFV11]. overcome [MBT16]. overconfidence [CdMCdMMdP17]. overestimating [LXWC17]. overhaul [Rou12b]. Overlapping [XLZ^{+}18, HGH17a]. overlaps [HC19]. overlay [HZL^{+}17, LKR14]. Overlaying [YDJ12]. overshadow [Gus19]. overshadowed [HSXL14]. Overturning [OCCSM11]. Overview [MPS^{+}18, OGRMOP19, RGdCMM17, ABGS14, BMP^{+}14, DMB17, DGPL15, Gau17, LG16, MB16a, OMJLVSN19, WvE13, ZB15]. own [BPVM11, LNMQRR15, Pra11d]. own-toward [Pra11d]. owned [CD14]. oxidative [WH12]. OzCHI [MAA17].

P2V [ZZL19]. Pacific [HTHB11, HIC12, WLZ^{+}15]. PACIS [CST11]. package [HBA19]. Pact [KBL15]. page [KBT14]. PageRank [CCLL14, PPM^{+}17, RR17, Yan14, ZMW^{+}18]. pages [BIL15]. pair [OKK14]. paired [VHH16]. pairs [HC16a, YK15]. pairwise [KKL14, RKT^{+}15]. Pakistan [BHB13, BY13, BYR13, FAI^{+}18, MR10]. palm [AAG14]. pan [Li14]. pandemic [PP11]. panel [CSC13, CSC14, HWS18, RGLE16, ZG17b]. panels [GWG17]. Paper [Ano12a, Ano12b, BH18b, BL13, BB17b, CRR14, Egg10d, Egg14a, GM13, GBMA14, HA19, HO19, HYC15, JYW11, LCS^{+}16, OL11, PYW18, PC18, PLWS14, Pra11c, Pra16b, Pra17a, Pra19b, RCJ18, WLM15, WSC16, XGCK19, ZZL^{+}10, ZZL19, ZZD^{+}18, ZZZ^{+}12, dS17b]. paper-reviewer [ZZD^{+}18]. Papers [Ioa06, MSYW12, ATK17, AA18, ANOdFC12, Ano10, Ano15, Ano18b, ABM19, ANA18, ALvH19, AND19, BHDI18, BSMD11, Bor16, BWdMA17, BL18, BYY18, BHH18, BT19a, BB17b, CGG19, CWH11, DLGP16, DX17, Egg11d, Egg11f, FMS17, FE14, FE16a, FE16b, FMM14, Fuk17, GRSFV19a, GF17, GXC^{+}19, Hag10a, HM18, Har16b, HB17a, HLC17, wH15, Hud16, ILGZ^{+}14, IMSK14, JTZ14, KO18, LL13a, Lee19b, LHG16a, LT10b, LS15, Ley12, LPL16, LRY18, LRZ13, LHW16, LF12a, Liu13, LF17, LNK^{+}14a, LNK^{+}14b, MC10, MHC^{+}15, MC12, MAGBBM13, MTA^{+}18, MAA18, NPT^{+}15, NA18, Ort16, PKR15, PJB^{+}12, PTMT10, PCRMCB^{+}18, Per10, PD10, PLWS14, Pra12a, RMA12, RG12, Sak19, SSAG16, San12a, San12d, SRW18, SL10, SCLC15, The19b, TC11, TC13, ULFRU^{+}14, WYY11, WFH^{+}16, XXL^{+}17, YY14, YLL15a, YXW18, YYLW14, ZYZ14]. papers [ZG17c, ZWW^{+}18, ZMW^{+}18, ZLTY18, ZLLL19, ZXLX14, dSD17, dSBC17]. para [Pet18a]. para-scientific [Pet18a]. paradigm [BS19, FE14, MB10a, MB13, PLGC18]. paradigms [PLG19]. paradise [DF15]. paradox [BLA16, HRC13]. paragraph [KWM^{+}18]. parallel [Ho16, LLGW13, ZG17f]. parameter [DW18, MBSB17, MZE19, Pra14b]. parameter-free [DW18]. parameters [CC13]. parametric [LBA19, Van10]. paratext [SMAABJ11]. Pareto [Aus14b, vZ13]. Park [Hu11, Hun12]. Parkinson [Kos14]. parks [DVMS17, FLM16, MTT15, vWBS^{+}16]. part [GTMRE^{+}19, Sch11b, ZG17d, ZG17e, HAJ12, KPS12, LF12b, PKSG12, TC13]. Participation [HMCD^{+}19, LOPAGS19, VCC12, GRSFV17a, KP12b, OA10a, OA10b, The18d, YZB18]. Participations [LVHS^{+}15]. particles [CRLMRPA10]. particularly [vRvLV11]. partner [AhOL14, ZT18, ZT19]. partners [CLO18, SK18, YHL^{+}18]. partnership [Cab13, CLJH12, dBONM^{+}19, Rou12a, SOBM16, Sch12b, Soo10a]. parts [WHZ14]. past [BHPVdPMR18, FAI^{+}18, Li19, MMOMALC16, OFP16, Par14a, SMF18, ZTC15, dCdSNB15]. Patent [ACT18, BVZV16, BRS^{+}16, CS11b, HY19, LZL10, Mar11, NF13, NA12, OI17, QZZ17, SPS14, SCL12, WZX11, YLL^{+}15c, YYS^{+}10, ZZ14, ATYL17, AAV13, Bak17, BAB13, BK10, BM13a, BM13b, BM15, CGV12, CD16b, CWL10, Cha17a, Cha18b, CC10a, CLHH10, CJW10, CC10b, Che11, CC12b, CF14, CLW^{+}19, CKCK10, CMM17, CKB^{+}14, DMM17, ÉMS^{+}13, FM12, FLB19, Fuk16, GGR11, GM12, Gom19, GP18a, HM18, HV18a, HFL14, HSWC13, HYC15, HSK17, HW10, Hur17, JYW11, JBC19, qJnShPL17, KM17b, KCK14, KL17, KB13, KK17, LLC^{+}17, LKS^{+}14, LK17, LPMK17, LS17b, LL19, LZZ^{+}12, LZZ^{+}13, LKR14, Ley15b, LKY17, LYWSV13, LHCH18, LCFC14, LS19a, MP15, MVS10, Mes11, MDG10, Moe10, MG12, MM15a, MYP19, NSKO15, Nii17, Ort11, PYK12, PYK13, PY14, PJY17, PLJ18, RKT^{+}15, Sch11a, Shi14, SWCH14, SYLC17, Wad16]. patent [Wad17, Wad18, WhCL10, WPCG13, WRV14, WLD^{+}14, WRC^{+}19, WY12, YK14, YK15, YSY^{+}13, YK11, YK12, YPK13, Yos13, YS14, YLH^{+}17, YLJ^{+}17, ZGL14, ZYF^{+}17, ZyZZ^{+}11, ZZZ^{+}12, ZyZZ^{+}14, ZZZ^{+}14, ZZP^{+}14b, ZHZY19, dPSS18]. patent-related [KB13]. patented [ACMP13, HDW^{+}15]. patenting [FLZ17, FWFM18, Fin11, GGG14, GS12, KG10b, KTT11, LM10, LSR13, MB10b, Men18, PS13, dSSdMAF14, SWH14a, SWH14b, WS10]. Patents [CSC13, LAHH15, MdFdA^{+}14, AHP17, BD12a, BK10, BM13a, BM13b, BT15, CPV14, CD14, CG11, CS11a, CHY13, CSC14, CK14, CL17a, CYK^{+}11, Com15, FS10, FI16, GM12, GZ11, HFC11, HC19, JS15, KB13, KKK^{+}14, KPY16, LGR17, LVSL18, LL10, Lee10b, LS17a, Liu13, LZC17, NA12, Ort11, QZZD18, RR17, SHB14, Shi14, SCL12, SD18, TB19b, WRM17, WLY14, YG18, YCK11, YST12, Yos13, ZPG^{+}14, ZFY^{+}17, ZQH^{+}17, ZYNZ18, vR17, vPD13]. Path [EMSH16, Yur18b, HLLT14, KGSS16, KS18, LGPC18, LLH19, LL13b, Pen19, SL16, Won13, YKLK14]. path-dependencies [KGSS16]. pathogens [TYWZ12]. paths [ÁRM13, FSLR10, GHA^{+}16, KS18, ML16]. pathway [WG12]. pathways [HZQ^{+}17, MP15, ZZP^{+}14b]. PATSTAT [LVSL18]. pattern [Che18a, Che20, GKF17, GZ17, HD17, KC12, LZZ^{+}12, SSS^{+}11, WLR^{+}14, ZZPG14, ZYT^{+}16, dSF13]. Patterns [BKRG13, LL15, Siv16b, ADM14, AHP17, AML17, BASL16, Bre13, BGJB16, CÖT15, CLLZ15, CJC13, CF14, Chi15, CRAJdMACÁ15, CH13b, CRLMRPA10, DX17, ESB15, EOS12, ELP11, FYC15, FMP17a, GHS18, HJM^{+}13, HC16b, HYS18, Hur17, IA19, JC19, JS15, KCK14, KK19, KdBBK15, KEP^{+}18, LSCK12, LZZ^{+}13, Li17, LC18, LCC12, LDG17, NHLL17, Ole12, OKCPS17, ÖS17, Ort11, ONB17, Oze12b, OKK14, PAL13, PR10, PSY^{+}19, RRBA10, RRSA18, RP17b, RPK18, RHGKD16, Sak19, SS10c, SV19, Soo10a, Soo19, TS11a, TP11, UMK14, US10, VuHL10, WLN^{+}14, WFS16, WYB^{+}17, YL12, ZY15, Med18]. Paula [Etz13b]. pay [Mik17, TV17, Wra16a, Wra16b]. pay-walls [Mik17]. pays [SGY15, SE18, XG18]. peace [PFDL17]. pediatrics [ABSF^{+}19, MRS^{+}16]. Peer [ADV10, ADR19, Fan11, LF12b, PS10, AD11a, ADD11a, ACD13, ANFF16, AND19, BD16a, BFS17, BGJ^{+}16, BGBS18, BSFCC15, BCJ^{+}17, BS15c, Bor12, BHJD12, BH15, BF17, CGG^{+}17, CPRSFVG19, Cop18, DGD19, DFS14, Fed13, FLM^{+}19, FT19, GRSFV17a, GRSFV17b, GPSM18, GWG17, HG10, HH19, HS17, JOGC17, JPZ^{+}10, KBAK17, KPRT16, KTRP17, MPY^{+}13, MS18a, Moo15, MFF^{+}16, MKHB13b, ND16, Ort17, PG14a, PTMT11, RMCM13, Rig13, RCJ18, RT17, Sak19, SB17, SBM17, SC13, VE14, WV13, WvEvL^{+}11a, ZS18, dZLwC^{+}15]. peer-review [ACD13, GRSFV17b, JPZ^{+}10, KBAK17, KPRT16, Ort17]. peer-reviewed [BSFCC15, MS18a, Moo15, Sak19]. peer-to-peer [BCJ^{+}17]. peers [GW10b]. Peirce [PMN16]. penalty [LHTL18]. Pennants [Whi18]. people [Lor10]. perceived [VSS12]. percentage [Pra19g, Sch15c]. percentage-valued [Sch15c]. percentile [Pra12b, YY16]. percentiles [BM14b, Sch13a]. perception [AGLNRR14]. perception-based [AGLNRR14]. Perceptions [JX13, BTNS14, ZCKZ16]. perfect [Osó18]. Perfectionism [SKM15]. perform [BTL19, MDFGAM14]. Performance [LYGQ12, Pra18d, AJSN18, ARK^{+}15, ADS10a, ADS10b, ADV10, ACD11, ADD11c, AD11b, ADS11, ADC12, ADV13, ADR14a, ADD14b, ACD15, AÇA^{+}14, ACORC10, AS18b, BKZ^{+}16, BBCP14, BHB13, BHA15, BBSS16a, BBSS16b, BBSS16c, BGÖ^{+}13, BW10, BS17, CGZ10, CMT18, CLHH10, CC10b, CSC12, CC12a, CC12b, CHC13, hCcTmWH15, CRZGVQdMA16, CRMPRS18, CC11b, CR18, CAS16, DR10b, DLM15, DGWZ13, Doc11a, Doc11b, DC17, DGDGSV15, ES18, EES13, EHK12, FZZ^{+}12a, GRSFV^{+}13, Glä13, GTD14, GW10b, GC10, GW15b, GNS^{+}15, GNHT18, Hal14, HH15b, HH10, HC16b, HSWC13, HDC13, HCS^{+}15, HDW^{+}15, ILB13, Ibr18, IL14a, JKMS17, KA13, Kaz14, KDFL14, KLL14, Kim19b, Lee10a, LSCK12, Lee19a, LL12, LZCZ18, MvdH13, MLVJ12, OCJB15, APFR^{+}13, Ort17, OCCSM11, PPK^{+}16, PPM^{+}17, Pra10d, Pra11d, Pra12c, Pra12d, Pra16b, PB12, RRLNAG15, RGGBV16]. performance [Ric15, Rod16, RPK16, RPK17, RD13, Sah16, SND19, Sch13a, SPS14, SZZC18, SDS14a, SV19, SPdSM16, SK11, SVCFI14, THAL15, Tu19, TBT19, VB12, Van14, VASNU^{+}19, VHG^{+}15, VAJCC17, WYvE11, WYH10, WLC17, WH12, YLY^{+}14, ZCL15, ZW17a, ZT18, ZT19, ZTRH18, ZHMX14, Zyc10, vLCCMV13, vR12, vdBS16]. performance-based [ACD11, HH10, SV19, Van14]. performances [CFM15]. performers [FLM16, Kwi18]. performing [GBM^{+}16]. period [ADC12, ABGS14, BGÖ^{+}13, BMP^{+}14, CJY^{+}15, IPIU13, KKE13, Lee10b, RV18a, RV18b, SAPR18, SZAJS14, TBW^{+}12, ULFRU^{+}14, ZZ15]. periodical [GTC16]. periodicals [SFNO12]. periodontists [BMM14]. periphery [Cho12, FKRS14, KLM16, SS10c, Zel12]. perish [Ben11, GVS17, NPT^{+}15]. permanence [LAL15]. permeability [BTNS14]. Persian [Moe16b]. persistence [CNPG17]. Persistent [LS16, ÖS17]. persisting [vAvdWvdB12]. person [BB17b]. Personal [WHS19, LWL17, MBA13, McC18]. personalities [Har15b]. personalities-the [Har15b]. Personality [BB17b]. Personalizing [BB17b]. Perspective [NH14, AAG14, BYR13, BAB13, BKRG13, BBVO10, BS11, BZBLP16, BSKB17, BSPL19, CL11, CB11, FFR^{+}17, GZ18, GZM15, HMK^{+}12, HSXL14, HYC15, HS17, KSSB13, KZSZ19, KPJ16, KMP^{+}11b, KvES11, KM18c, Li15, Li19, Lin18, LGH^{+}14, LZL10, MS18a, MCLL17, MB16a, MGMW14, MMAHS10, NA18, PM18, PMJF19, Pra19a, SRF16, The12, WHC^{+}13, WBX18, WS13b, XTZ15, XBD^{+}18, YQW13, YYS^{+}10, YHL^{+}18, ZCL15, ZCZ^{+}16, ZW17b, ZW17c, ZZZ^{+}12, ZyZZ^{+}14, ZZZ^{+}14]. perspectives [Ano16b, CLW^{+}19, DG16, HH18, LAHH15, SS16, SRP13, Wan16, WW11]. Persson [Ano12c]. pertaining [ADM19, SHS15]. pervasive [ZW11]. perverse [CLD13]. Ph.D. [SR16, SK17]. pharmaceutical [CC10a, CSC13, PR14]. pharmaceuticals [Mes11]. pharmacoeconomic [GK19]. pharmacological [GVGSEPRC15]. pharmacology [BL10, DGWZ13]. pharmacology/pharmacy [DGWZ13]. Pharmacy [RGdCMM17, DGWZ13]. phase [KGY^{+}17]. Phases [Suo14]. PhD [BN10, Bre10, BL11b, Bre13, BN14, CGZ10, Hag10a, KKCG18, KKE13, Lar12]. PhDs [ÖS17, Sin18]. phenolic [ATCCAAB19]. phenomena [RPP18, WW11]. phenomenon [HW10, KK13, LY12, WZS12, ZWHH13, vAvdWvdB12]. Pheonix [AAG14]. pheromone [RASP13]. Philippines [NA14, Vin12b]. Philosophy [Wil15, WB15, APPS15, MGMY^{+}18, Pet18a, Pol16b, Wra14]. phoenix [HA17b, Har16d, HA17a]. phosphate [HLLT14]. phosphorus [GCLcG15]. photovoltaic [JCCC13, LLC^{+}17, LAHH15, Wu14, dPSS18]. Physica [BHDI18]. Physical [SDS19, DLM15, ZW18a, LHW16]. physicists [Dya17a, Dya17b, ZCKZ16]. Physics [BND11, Mat13, CRLMRPA10, GBMB10, GW10a, JK19, KMS16, KM17a, LRZ13, MSYW12, MSdBC16, Pri15, Pri16a, Pri16b, SPB18, SJ19, TH19, Web16, ZZY19, ZXLX14, ZL15b, dACdFC18, BHJD12, Tom17, ZXT^{+}19]. Physiological [LV12]. physiology [PV15, Fie15a]. picture [AAV13]. pilot [AYS^{+}13, Cha18b, FLM16]. piloting [JH16]. Pinski [Pra19e]. pitfalls [SMM^{+}19]. place [Asu19, Ben12, Hud17, Pra10e]. places [Cam14, Egg14a]. plagiarism [GREL14, Jar16, JTZ14, ZLZ19]. planning [KKK^{+}14, MT12b, PHL17, YLL10, YK11, YPK13, dCdSNB15]. plant [Jon10, KNK^{+}19, PS15, dSD18a]. plants [SSdOS17]. plate [MB13]. platforms [OMOR13b, SMM^{+}19]. play [Chi14, KCU19, WLY14, Zuc10]. players [BSB12, SK11]. playground [Bai18]. Playing [SHB14]. plea [Har13a]. PLOS [dW15, BHDI18]. plot [Pra19f]. Plots [BH18b, Cam17, Cam17]. PlumX [Ort18, TSRGG17]. plus [HB18b]. PNAS [DAYY18]. pneumococcal [CSR^{+}18]. point [Har17, HTHB11, MB16a].

Poland [KKL14, KK19, KR17a]. Policies [ABSF^{+}19, HSX^{+}15, Laa14, LOMLPA^{+}17, Shu17, dS17a]. Policy [BHM16, HYS18, Ahr17, Ano16b, BDE11, Bou14a, BS19, CRBRG^{+}18, CIL^{+}16, CZ18, DLL^{+}16a, DSM11, Dem18, DKS18, EdS19, GFC18, GHS18, GP18a, GKS16, HB17a, HRH10, HSXL14, ILP11, JSZIZ13, KK19, LKW^{+}16, MC13, MS16b, Pou12, SM16b, VPM16, Vin18, Waa13, WF17, ZL18b, ZCZ^{+}16, ZY19, ZB12]. policy-makers [JSZIZ13]. policy-related [BHM16, HB17a]. policymaking [FLB19]. Polish [Kol12, Kos15]. Politécnica [HMCD^{+}19]. Political [CA18, Hen18, Chi14, Chi15, Gan12, KP12b, MOA16b, NvLvR10, Sak19]. pollution [LWR^{+}17, YYS^{+}10]. polyangiitis [ARE^{+}18]. polymer [CYK^{+}11, Pra14c]. polyrepresentation [AF15a, AF15b]. poor [GHS18]. Popular [TH13, RMA12]. popularity [Aus14b, HF19]. populated [Ort15]. population [LSR13, Ort15, SM15]. portfolio [CFM15, LLC^{+}17, LKY17, SRGMF15, dART^{+}17, dCPRP18]. portfolios [Luo12, RJ14, Sch11a, SK12, TTC17, ZP15]. portrait [Jac18, TM12]. Portugal [BCC^{+}17]. position [CRR14, GP18b, MSL11, MT13b, Pei19, Puu10, dMALIM14]. Positioning [URU10b, CH15, FM11a, Koz15]. positions [CLL^{+}17, Pac19, Sch17a]. Positive [Bal12, Fan13b]. possible [ADS10a, DKS18, Ley15a, VB12, ZHL19]. possibly [XGL^{+}19]. Post [BIH17, BTNS14, JW18, CA19, JKMS17, LA19, Pra18d, ZTP18a, ZTP18b]. post-communist [JKMS17, Pra18d]. post-doctoral [ZTP18a, ZTP18b]. Post-interdisciplinary [BTNS14]. post-Soviet [LA19]. postdoctoral [BBVO10]. posteriori [GSTD11]. postmodern [HLE10]. postulate [DGGBDG17]. potential [CF14, DGDGSV15, FMM13b, JX13, JMM18, LS19b, MP15, MKP16, PLJ18, SA17, SL12a, TCH^{+}15, ZQH^{+}17]. Power [Brz15, Sch17a, AJSN18, BS15a, CNPG17, DXL^{+}12, FMP17a, GRSFVCP19, GGP14, GYZ15, KNK^{+}19, Lee19b, Lin11, Pra19e, RPK18, SCGZSL^{+}13, SLISC17]. power-law [Lin11]. power-weakness [Pra19e]. powers [SH18]. pp [Ho16]. PPP [SZD16]. practicable [BM14a]. practice [BMZ^{+}17, Bur14, Din14, DJWS11, ER12, FE14, Kha13a, Kha13b, LCFC14, Paj15, PW17, TKA17, Zhu17]. practices [BM12b, DY18, GZM15, JX13, YB14]. pragmatic [Har13a]. pragmatics [Ing12, LX19]. Prathap [BBSS16c, BB17a]. precaution [SCL12]. precautions [YSND17]. precise [CPY13]. predatory [Fra17, LMKG19, PIB18, WP18]. predict [Bar17b, BF17, FA10, GP13, GW10a, GHA^{+}16, Kis11a, NTM^{+}18, PMN16, Rig13, SA17, WhCL10, WM17, WFZD19a, WWC19]. predictability [KKS^{+}17]. predicted [HL15, HK12]. Predicting [AMI18, BGJB16, FLB11, JKPL18, JL18a, Kos18b, Lee19a, MJHG13, MPH19, MKHB15a, MKHB15b, BSK18, SI17, SLXD15, JL19]. Prediction [CXZ19, ÉMS^{+}13, Bou11, CGC18, CU16, DAMC15, DXL^{+}18, GZGAC16, GZGAC17, GR14, KD19, KJES16, LD16, NZL^{+}19, PLWS14, PLGC18, YYLW14, ZGJ18]. Predictive [Lee19b]. predictor [BT15]. predicts [ADR19, Fox17]. Preface [GW16, GMS10, Tsa12, CFM18, GSB18, MS15a]. prefer [vRW18]. preferable [ADD10]. preference [ANZ15, Ric15, SC10, WAT16, ZLLL19]. preferences [GRSFVdMA14, LF14b, OBG11, SW19b]. preferential [SL16, ZZFD16]. preferred [HBA19]. prefix [Han15]. preliminary [AR18, Har13d, Sma11, Wra14]. premier [LHM^{+}11]. premium [Cop19a]. preprint [The18a]. pres [GBDG19]. Prescribed [BM12b]. presence [Doc12, OMR14, RAM18, SK14b, ZCW14]. present [Par14a, TG16, ZWW^{+}16]. presentation [HO19]. presentations [Ano11]. presented [CWH11]. President [AM18]. presidential [Par14b]. press [BH17b]. Presses [MRGT18]. pressures [BGBS18]. prestige [FSSPG^{+}15, GRSFV11, GRSFV12a, GRSFV^{+}12b, SZZC18, YWZ^{+}17]. prestigious [LSE^{+}18]. prevalence [RV18a, RV18b, ZW18a]. Price [AH11, Ano12c, Ano14, Ano17a, CXZ19, Dan19, Egg10c, KL16, BS15b, SA16, De 13, Wra10]. primarily [Cha19b]. primary [ZYNZ18]. primordial [Cab18]. PRIN [Zin16]. prince [ZLF18, vRW18, EG18, SSZL18]. princes [TVA17]. Principal [CC13, KKLP18, Ste19]. principle [Bor14, Cha16, MB13]. principles [BH16a]. print [Mou16]. print-online [Mou16]. printed [KCK14]. printing [HZQ^{+}17, LKP11, LP18b, ZLLD19]. prior [LYWSV13, Wad16, Wad17]. priori [GSTD11]. priorities [RHMH17]. Priority [ND16, HHK^{+}12, LGR17, LVSL18, NH14, SAPR18]. private [ADS10a, BD12a, Com15, Mor16, dMALIM14]. Prize [BOS14, CMT18, CÖT15, MSYW12, SBB16, Bjø19, GW10a, Har13d, HR17, MSYW12, TA17, YGD17]. prize-winning [SBB16]. prizes [GW10a]. pro [GHS18]. pro-poor [GHS18]. probabilistic [JC12]. probabilities [MRGT13]. Probability [RNB19, CMT18, EC16, GRSFV19a, LHG16b, WLD^{+}14]. probe [XGL^{+}19]. Probing [YR10]. probiotics [VGPdlC^{+}17]. Problem [Soó14b, GKF17, Har15b, HO19, Pau10, TA14a, ZZPG14]. problematic [HH19, Lin18]. Problems [Egg11d, GRSFV17a, HT18]. procedure [Bre10, IB15, MNdF16, YK14]. procedures [BH15, HH19]. Proceedings [BI10b, Bar17b, HSBW10, ILGZ^{+}14, MF14, SFNO12]. process [BND11, Bur14, FT19, HS17, KBAK17, KdBBK15, LZB10, LWIB16, LZC17, LOMLPA^{+}17, MGM^{+}17, MDDG17, MHM12b, OH19, WZCC19, WG10, ZYF^{+}17, Zuc10]. processes [Ano17d, BH16a, BN10, Cha17b, CyPP12, Ley15a, YL10, dCdAMB19]. processing [BS15b, PB18]. Produce [CHY13, BWdMA17]. producers [NSC13, SKM15]. produces [The17b]. producing [HBA19]. Product [MDDG17, ADV10, Fuk19]. product-level [Fuk19]. Production [RPSA17, ACF^{+}17, AATBPAB15a, ÁBV^{+}14, ACAGD^{+}17, BVB13, BSFCC15, Bra12a, BDC^{+}12, CSR^{+}18, Cav16, CVC14, CH13b, CRLMRPA10, CFdC^{+}14, CDdS^{+}12, Das16, ES16a, EW15, EGU10, FP18, dCPF14, Fin11, FKRS14, FM12, FZZ^{+}12b, GGR11, GSMT10, GC10, GdA14, HR11, IBL13, IL14a, JCCC13, JvGH10, MC10, MD18, NJ10, NH14, OM11, PPK^{+}16, PEPUT15, PFPCM^{+}19, Pra10d, RGdCMM17, RCdJ^{+}14, Sch14c, Sch17b, Soo18, TBS15, Tol12, YC10, YZ17, dAG13, dJC15]. productive [ADD11d, BL11b, GdOdAG^{+}13, HAJ12, KPS12, MS13, RV18a, RV18b, SAR19]. Productivity [BT19b, DRS14, KHK13, ADD11d, ADR13, AD14, AER^{+}14, ACS18, ACRC17, ÁCCG^{+}15, AA10, Bar17a, Bas10, Bas14, Ben15, BS15c, BA15, CMUdF15, CVC14, CÖT16a, CKT17, CC10a, CHY13, CHL10, CÖT16b, CMVP16, DKS18, DJWS11, Fuk14, Gan12, GNS^{+}15, GKS18, HH15b, HCL14, Ibr18, IMHG12, IMH13, IJF16, JSZIZ13, JZL10, KFB18, Kwi18, KKT^{+}18, LVGV^{+}11, LPC17, LRC19, LLCL11, LT16, LD16, Lor10, Lun19, Mat13, MR18a, MB16b, NA14, OO12, dFPYdCL12, PRRC15, PPK^{+}16, Pra18e, PKSG12, Puu10, RGdCMM17, SC18, SH15b, Sin18, SJOC18, SM14, SDS19, Soo10a, Soo14a, Sot10, SK14a, SK14b, Tor13, Tor14, TBT19, VO17, Vin12b, Wal15, Wal16, WLZ^{+}15, YGD17, ZGL14, dOM16, vAvdWvdB12]. products [ADD14a]. Professional [LWB16, BM12b, CB11, DMB17, FMS17, HMI19, JX13, Puu10, RRL16, TKA17, WOW13]. professional-based [DMB17]. professionals [GAPP18]. professions [DMB17]. professor [MM18, MM19]. professorial [GG15b]. professors [Hos11, LVGV^{+}11, PRA16a, ZTP18a, ZTP18b]. Profile [LGH^{+}14, Car16, EBR16, GK14, Hen19, Hen20, LSM^{+}15, MJC14, Oli15b, RGdCMM17, SLD^{+}17, VH17, dSF13]. Profiles [VO17, DFG^{+}18, GT18, Li17, Pir19, SM12, SPdSM16, VAJCC17]. profiling [CLkS11, MJC12, PMJF19, TDG17]. Profitability [CKCK10, CC10a]. Program [CFdC^{+}14, SA12, TBB^{+}16, YPH10, vEW10, WTM^{+}16]. programme [SBSR19, OA10c]. Programmes [BM11, OA10a, OA10b, OPGW^{+}13]. programs [BR12, HIC12, IF13, PYL16, SLD^{+}17, SJOC18, Yur18c, LVHS^{+}15]. Progress [Ho16, LLGW13, DSG^{+}15, GKB^{+}19, MAA18, Nar12, RR17]. Progressive [San12d, MV19, San12a]. project [BH16a, GBM^{+}16, GKB^{+}19, KM15a, Luo12, dZLwC^{+}15, Zin16]. projects [ABL17, ALvH19, BM11, LdZwC^{+}17, SR16, TCT^{+}13, Zin16]. prolific [PRA16a, Pra10a]. prologue [ZTC15]. Prominent [CH13a]. promise [HM15a, HM15b]. promising [MAGBBM13, WF17, YKCK13, ÖS17]. promote [MTT15]. promoted [SDS19]. Promoting [aSTS17, aSS17, WBX^{+}17]. promotion [Bor16, GCGP10, Hos11, Per18, YZ17]. promotions [MM18, MM19]. Proof [HM15b, HM15a]. propagating [WG10]. propensity [CJC13, CFdC^{+}14, MD12, EC16]. properly [Koz15]. Properties [vR12, Ama15, Coc18, FSSPG^{+}15, Glä10, Gom19, Saa10, SBA^{+}19, Zit11]. property [ACD14, CO10, PNVCB18, YCK11]. property-function [YCK11]. proportions [The18c]. Proposal [OMOR13a, VEJC^{+}18a, BCJ^{+}17, BM14b, BM14a, BSK18, HP10, MARMSG19, Mou16]. Proposals [FM11d, Pra11c, BB15, HRB^{+}13, HRB^{+}14]. proposed [Bre10, FT19, KHVGA^{+}16, LWT16, PKR15]. proposing [DCM16]. proposition [Sah16]. Pros [Dem18]. Prospective [RWG^{+}15]. prospects [DH13a, ZWW^{+}15, ZWW^{+}16]. protection [CO10]. proteomics [TFH14]. Protocol [CIL^{+}16]. proton [CYK^{+}11]. prototype [Moo15]. provide [AAV13, GRSFV19b, TSG13]. provided [LvI10, MCR^{+}12, SM16a]. providers [Ort18]. province [GC10, JJR10]. province-level [GC10]. Proving [MCvFP16]. provisional [CC16]. Proximity [FFL16, SL16, Fie15b, Fie15a, LC12]. proxy [ADD10]. PSTR [CSC14]. psychiatric [LR12a, MCL^{+}11]. Psychiatry [MDFGAM14, BL10, SS14]. Psychological [GWBSVWB13, LA19, VSK18]. psychologists [DJWS11, FK16, Kor18]. psychology [uHBK19, uHBLKH19, BVB13, BSvEK13, Ben19, Bru10, GBB15, GBGB13, HH17a, HH17b, HH17d, KB11c, KFKS15, Kra10, KvES11, KWW15, Kra16, NvLvR10, SLGO17, Vin12b, ZTP18b, ZTP18a]. Public [CRZGVQMA15, Mor16, SÁV18, Wil15, ADS10a, ABMSSP16, ABSF^{+}19, AC12, BR12, BHJD12, CRZGVQdMA16, CFM15, CAS16, CZ18, FSAB10, Fuk19, GSMT10, GMM16, HRH10, Hun12, Kon12, LSCK12, LOMLPA^{+}17, MC10, MC13, OROMAA16, QA18, SMLHCP17, SFBS17, SWH14a, SWH14b, TBS15, WDP11, dJC15]. Public-private [Mor16, ADS10a]. Publication [DMB17, DC15b, FM12, GZ11, GZM15, GBDG19, KSSB13, KEP^{+}18, MLY^{+}14, MHTB17, Ole12, Soo14a, TG17, WB15, ADC12, ADD16, Abt12, ALH15, AML17, Asu19, BHS14, Bas11, BB19, Ber18, BSK15, BND11, BHL18, BC13b, BZBLP16, Boy17b, Bur12, Bur14, Chi15, CRAJdMACÁ15, CA19, CH15, DGF17, ESB15, Egg11e, Egg11f, EOS12, Fuk17, pGSyW^{+}19, GRSFVCP19, GHT16, GBHT16, Hag10b, Hal14, HS16a, HK12, HJM^{+}13, Hou17, HB17b, IFT^{+}18, JKMS17, KFB18, KK19, KPRT16, KTRP17, KWW15, LvI10, Lee10a, MSYW12, MM14a, MR15, MB14, Mat13, MR18a, MS14, MNdF16, MT13b, MHC14, Mue18, OVJM17, OCCSM11, PKR15, PAL13, Puu10, RAA18, RCN^{+}14, Sak19, San12b, San12c, SND19, SBT18, SFR^{+}19, SV19, SPdSM16, SW19b, SSG^{+}18, Tur16, TBB^{+}16, WWH^{+}17, WT15, Won19, YGD17, YY14]. publication [YDJ12, YL12, ZL18b, ZYF^{+}17, ZP15, ZTP18a, ZTP18b]. publication-citation [Bur14]. publication-level [CA19]. Publications [CAV^{+}19, HH15c, IFH15, ATJ16, ADS16, AD16, ADR19, AC13, Asu19, BI14, BSG17, Bas13, BBP14, BSMD11, Bor15b, Bor18, BHL18, BPVM11, BC17, BL17b, CGV12, CA18, CXWW18, CRMdMA15, CPF18, CB11, Das16, DWGL16, Dor17, DKS18, EDEH16, Fan19, FM17, Gau17, GMJ^{+}17, GD11, GZ11, GHT17, GALR16, GB14b, GSE^{+}18, GB12, HC14a, HHBB18, He13, HPKS18, HL18, HH17c, HH19, HR17, HLW19, IPIU13, IQT^{+}19, IH14, JG12, JG14, Kim14, Kis11b, KPJT14, KWW15, Kra16, KPY16, LL13a, LCC12, LF14a, LXL15, Liu16, LXL16, LOPAGS19, Lor10, LABL13, LA19, LPZ17, MH16a, MC10, MVS10, MAA^{+}11, MH16b, Man15, Mik17, MCB15, NvLvR10, OFP16, Osó18, PY19, PV15, PROG19, Pra19g, Pri16a, QRJ^{+}17, RRRGVD19, SH19, SA11, SBD^{+}19, Sch12a, Sch17a, SLISC17, SZAS16]. publications [SK18, Shi14, SLH18, SM15, Soo11a, Soo17, Soo18, Sot12, SD18, TBMM18, Tri10, VSS12, VCC12, VE14, VG17, Vin17, WP18, WCK^{+}12, WMXZ14, WTG15, WFZD19a, WFZD19b, WZFD19, WL18, WS10, WS13a, XYHD18, YWL16, YP19, ZCW14, ZKD11, ZPG^{+}14, ZYNZ18, ZW18a, ZWZ^{+}19, ZZY19, ZLH^{+}16, ZZFD16, KK15, vEW17, Wra16a, ZWX22]. publicity [Fan13a]. Publish [GVS17, NPT^{+}15, Ben11, BPVM11, BC13b, BL11b, DR10b, LNMQRR15, San18, YP19]. published [ABMRVZ14, ABMSSP16, ABSF^{+}19, uHBLKH19, BWD10, BSMD11, BM12a, BHH18, BT19a, CVD14, CHL10, CH12, DGD19, EdS19, Fuk17, HW12, LL16, LRY18, LHM^{+}11, LNK^{+}14a, LNK^{+}14b, PD10, RM18, San13, TB19a, TA11, VVN16, WLF15, WM17, YXW18, YJ11, ZG13a]. publisher [Asa19, GFC18, TSRGCCJC14]. publishers [GRSFVCP19, Laa14, MRGT18, SLGO17]. Publishing [SW19b, ASPF^{+}16, Ben15, CKT17, CRR14, CRLMLM17, FEHC19, GRSFV14a, GVS17, HE16, KTLD16, KHA17, KG10b, LHG16b, MRGT13, Mêg13a, Mêg13b, Mêg13c, Mik17, Mou16, PB18, PIB18, Puu10, PML14, SC10, ZL15b, ZB15, Zhu17]. Publons [Ort17]. PubMed [BC17, EDEH16, HZD^{+}15, IFT^{+}18, BMM14]. Purpose [SJ10]. purposes [Ano16d, BI18a, dSD18b, dSD18c]. pursuit [DTM^{+}13].

Q [GLM11]. Q-measures [GLM11]. Quadruple [LPL14, Par14b]. Quadruple-Helix [LPL14]. Qual [CC12a]. Qualifying [HH15b, LHLH19]. Qualitative [ÁBV^{+}14, MnaeR^{+}15, Pen19, AA19, BFHS18, dCPF14, GdA14, HNG19, KBZS15, MAGSTRC15, Wan16, dAG13, dCPRP18]. qualitatively [Rya16]. qualities [dB12]. Quality [BH17b, BB17b, CL13, Moo15, PDAN19, Sal17, Abb11, ADD10, AR18, AAB^{+}13, Ano16b, ANFF16, ASW18, BKZ^{+}16, BKL15, BR11, BS15b, BH17a, CGPT15, CKCK10, Cla15, DLL^{+}16b, ERW12, EdS19, GRSFV14b, GKB^{+}19, HJL18, Hau16, HS17, KM15d, KB11a, KKL14, Lia11, LCY14, Lyk18, MS18a, MGLZ10, MZ14, MOO17, MPM18, NCG^{+}19, PS16a, Pra11d, Pra12d, Pra19g, RZ12, Ric17, SP14, SZMS18, SL17, Sch16, SA16, SDP^{+}19, VEJC^{+}18a, Wal15, Wal16, WhCL10, WXZ^{+}16, WAT16, WS10, ZYZ14]. quantification [Cha18a, UCH19]. quantify [Hir05, Hir10, Hir19a]. Quantifying [Arb11, ANFF16, CBWJ18, MHKB16, OFP16, SGSS17, ZP16, AAH10, Cab14, GSKM17, SS10a]. quantitation [WOW10]. Quantitative [FSO11, GB17b, Pen19, ÁBV^{+}14, AA19, BFHS18, CLO18, DRMMC19, FSOS12, GF17, GB12, HLSC18, HWS18, HYYL12, HLY14, ISR11, KZC16, KBZS15, LWL17, LDVSGDR16, MPY^{+}13, MnaeR^{+}15, MRGT13, MFF^{+}16, Oli15b, SFNO12, WWP14, Wan16, WBX18, YKCK13, Yos13, dCdSNB15]. quantitatively [Rya16]. Quantity [dB12, GF17, GKB^{+}19, KM15d, Pra11d, Pra12d, TYWZ12, Wal15, Wal16, WS10]. quantophrenia [Etz13a]. quantum [FG15]. quartile [GRSFVMB12, LHG16b, MGC19]. Quasity [Pra11d, Pra12d]. Quebec [OBG11, LVGV^{+}11]. query [MAA^{+}11, RNM18]. quest [YSM^{+}19]. question [DdlPPL^{+}19]. Questionable [Che15].

Questioning [FCCMTRVR18]. questionnaire [BN14]. questionnaire-based [BN14]. questions [CP16]. quickly [TC11, TC13]. Quintuple [YYP17]. quo [Pol16a].

R [HBA19]. R&D [ZT19, AChO19, BGAAM15, CO10, CHY13, CC13, DR10a, FLB19, GC10, HMK^{+}12, HFC11, Hun12, qJnShPL17, JSZIZ13, KKK^{+}14, LLC^{+}17, LPMK17, LdZwC^{+}17, Luo12, MT15, PY14, Par15, PHL17, PS13, Pir19, QDK19, SL17, TYYW16, TW18, Van14, WT14, mYqS15, YPH10, YLL10, YK11, YPK13, ZT18, ZG17f]. R&I [Ano16b]. rabies [SR15]. race [BSB12, HJM^{+}13, MT13b]. Radiation [SMM^{+}19, Hal13]. radio [HW10]. Radiology [LV11, GR16, KBAK17]. radius [San12c]. railway [ZZ14]. Rainbow [SSG^{+}18]. Ramanujacharyulu [Pra19e]. Randić [Egg10e]. random [Egg11b, GR14, HJM^{+}13]. randomized [KCT^{+}17]. Rank [GNVQdMAG11, PG12, ADD18b, ADD^{+}15, BHA15, BIL15, CB11, Egg11a, PKR15, RGCM14, SP12a, Sch14a, SYDW19, Wu13, YR10, AMFLH15]. rank-order [Egg11a]. ranked [ICC16, KHK13, Pra19g]. Ranking [dSAEE15, BVOL18, BD10a, Bor14, GRSFV^{+}12b, IB15, KM15d, Laz10, SKM15, WZFD19, WLZ^{+}15, ZZFD16, AOFU10, AÇA^{+}14, Aus13, BIL15, BBSS16a, BBSS16b, BBSS16c, Bel17, BBV10, ÇAAÇ15, CC11b, DH13b, DMV10, Doc11a, Doc11b, Doc12, Doc13, DC14, DC15a, DEC15, DB16, EBD15, FL16, FGP13, GRSS16, GRSFV11, GSTD11, HG10, HM15a, HM15b, JBMR11, JSZ13, JJS^{+}12, KR17a, LB12, Li18, LRS^{+}18, LdZwC^{+}17, LIdMAM11, NASR11, OCM^{+}12, PB17a, Pra12b, Pra16b, QRJ^{+}17, RPGM16, Saf19, SND19, SAR19, SSG^{+}18, TSMTDLCH11, VHH16, VPM16, XLR15, YWZ^{+}17, BG18, FCCMTRVR18]. Rankings [MHC14, Saf19, ADS10b, ADV10, ADD11a, ABILO10, BR11, BGR19, BSFW10, BG17, BD13, hCyL12, CPF13, DCGZ^{+}12, DMV10, Fra14, Fre14, GE11, Joh18, KHA17, KKV^{+}13, MSdBC16, Mat12, Moe17, Pie18, PS16b, RGTSLCH14, RGCM14, RTP17, Saf13, SM16b, TYYW16, ZZZ^{+}14, dMALIM14, vRvLV11]. ranks [ADD11d, ADM14, JX13]. rapidly [YK11]. rare [AP16, FR11, JS15, MK19]. rate [CLHH10, EdS19, LvI10, MRS^{+}16, Rou18, SBT18, Sch10b]. rater [JOGC17]. rates [ACP12, DD18, KBAK17, Lee19b, LABL13, MB15, RMH14, Sch11b, SL12b, TC13]. rather [AL12]. rating [ILP11, KKS^{+}17, LR12a]. ratio [Ber18, LHW16, Pra19e, YZB18, YR10, Yur18a]. Rational [Sol06]. rationale [Egg13b]. rationality [McC14]. ratios [BL17a, MV19, XTZ15]. RC [AAH10]. RC-index [AAH10]. RCN [KM16]. Re [CS11a, CLW^{+}19, PW17]. Re-examine [CS11a, CLW^{+}19]. re-use [PW17]. reach [SL17]. reaction [MHKB16]. Readability [LY16a, HC16a]. reader [ATK17, The17a]. readers [KY17, The18b, The18f]. readership [Ask18, Eld19, HPKS18, SGG^{+}14]. readership-based [Ask18]. reading [GBMB10, Har16a, Har16c, OKCPS17]. reads [CB19, Eld19]. ready [HA17a]. real [Ama15, LABL13, PIB18]. reality [HHA^{+}16]. realized [PD10]. really [BSMD11, Cam18, Joh18, Mar11, MS13, Par15, Rig13, Saf13]. realm [CRMdMA15]. realms [SM17]. realtimely [WWX13]. reasoning [LdZwC^{+}17, dZLwC^{+}15]. Reasons [HTHB11, PZ17, WBX18, HV18b]. reassessment [HIC12]. Rebuttal [Ho16]. rec [LFBI19]. rec-index [LFBI19]. receive [AC13, Fan13b, GRSFV16b, HV18b, SRW18]. received [Asu19, Egg10d, XGL^{+}19]. Receiving [Sha12, WPCG13]. Reception [Dan14, CÖT15, MHFB17]. Recipient [Dan19]. recipients [PPK^{+}16]. reciprocity [DFS14]. reclassifications [WSH16]. Recognition [BRS^{+}16, ABGS14, BCT19, BYY18, CMT18, EG18, Li14, SSZL18]. recognized [KHA17, Sot10]. recombinant [LL19]. recombination [ZHZY19]. recombine [Ley18]. recommend [YKCK13]. recommendation [CjZZ^{+}19, EMSH16, HA19, KC15, KJW^{+}17, WHS19, ZML19, ZZD^{+}18]. recommendations [DJWS11]. recommender [AI17]. Recommending [GR14]. Reconsidering [DGGBDG17]. record [Gau17, SGN15]. records [BC13b, CCM^{+}11, DCS12, Jac18]. recruiting [Fox17]. recruitment [ADR16a]. recursive [WYvE11]. red [BHL^{+}10]. rediscovery [SSZL18]. reduce [KKS^{+}17]. reduction [ZZ14]. redundancy [IL14b]. redundant [Cop19b]. Redux [Cab18]. REF [Hud16]. referee [BWD10, GRSFV16b]. Reference [BHL18, CFL12, BTNS14, Cab18, CH15, Egg10c, EHK12, GNVQdMAG11, Kra10, LTG12, LCFC14, MH16b, MB14, Rod17, Sch10a, Sch13b, Yur18a, ZSCR^{+}18, MHTB17]. reference-based [Sch10a]. Referenced [WB15, Hou17]. References [QL12, ATJ16, CGV12, CPV14, COS11b, FK17, FK18, GXC^{+}19, HHBB18, Jac18, JKJL14, KWW15, KN15, LCS^{+}16, Lin11, LH12, MB16a, PYW18, RWG^{+}15, Shi14, The19a, ULFRU^{+}14, WPCG13, XHA^{+}19]. referencing [BHH18, LZR14, YL10]. referential [LPL16, WLC17]. referring [RTP17]. refined [CCLL14]. reflect [ATK17]. reflected [Bor16, JK10a, MSL11, Mor16, SLG10]. reflecting [MAA17, Soó14b]. Reflection [GY12, Bar11]. Reflections [CF18, FMS17, Bha18, Har19a, Jac12]. reforms [AZSA14, AZSA16]. Refrain [AD16]. refusals [Wad18]. regarding [ABSF^{+}19, uHBLKH19]. regime [CRLMLM17]. regimes [Ley13b]. region [ADS17a, ADS17b, CdSPdM13, LT10a, MLVJ12, WLZ^{+}15, Moe16b]. Regional [Jar10, SP12b, TS11b, ADS12, ACFL11, BBSS16a, BBSS16b, BBSS16c, BB19, BBP14, DCM16, HWL11, HDC13, MC10, MvdH13, Pra16b, QDK19, QA18, RVFEdlM10, RG18, STCRPA18, SL16, TP11, VO17, WM17, WZ19a, ZK19]. regionalization [SSdOS17]. regions [AZSA14, AZSA16, BPGGdMA12, HGH17b, RY14, SW19a, ZG17f]. registered [BL11b, vLvWW16]. registering [The18c]. Registers [PT17]. regression [CSC13, CSC14, LRS^{+}18, Par15, QRJ^{+}17, Yos13, YYLW14]. regularity [FM11d, Pra11c]. regulatory [dSSdMAF14]. Rehabilitation [HHA^{+}16, DMB17, MKYM^{+}17]. reinforced [YWZ^{+}17]. rejected [BWD10, CGG^{+}17]. rejecting [ZQH^{+}17]. Rejection [Sch11b, Sch10b]. rejections [CGK^{+}14, Oos15]. Rejoinder [dSD18c, LO12]. relate [BSvEK13, LABL13, MR18a]. related [ASPF^{+}16, BGM17, BHM16, CSR^{+}18, CSC12, CC12b, CH12, HLE10, HHK^{+}12, HB17a, HYYL12, HLY14, HW12, KB13, Kos14, LPMK17, Lee19b, LTGH15, NHY^{+}14, Ort17, PHS12, RPAMR19, San13, SGN15, xShLY^{+}15, SCLC15, tScL13, TABA16, WL14, YJ11, ZGY16, ZXM^{+}16, ZLL^{+}17, ZWW^{+}16, PLW^{+}15]. relatedness [ADD17b, Hsi11, JK10b, LLW13, ZYG15, ZG17c]. relating [Hos11]. Relation [CMT18, HN16, Bak17, DKS18, Egg10b, Egg10d, Egg13b, ER19a, Glä12, GBDG19, HJL18, HBS^{+}19, JN11, KKBW17, MSL11, MPM18, Nii17, SRW^{+}15, WZ17, ZG12b, vR12]. Relations [SÁV18, APT13, CY13, Fin11, IL14b, JAAA18, Ley11b, LPL14, LWB18b, PCR18, SGSS17, YSM^{+}19]. Relationship [ERW12, GBMA14, IBL13, NA18, dFPYdCL12, ADS11, BHB13, BSS15, BBL17b, BS15b, BAC13, CKPY19, CS11a, CSC14, Cop19a, EMSH16, Fuk14, GD16, GP18b, GMSZ18, HM18, KA13, KG10b, KM12, LAL15, LH12, LTK^{+}18, MA19, Med15, MDG10, NP11, OMR14, Ort11, Ort16, Pra17a, QZZD18, RPK16, RP17b, RPK17, RPK18, RHGKD16, Saf19, San12c, TLSH14, Tsa11, VZAMG19, WQY12, Wol15, YWW17, YST12, ZLF^{+}14, ZSCR^{+}18, dW15]. Relationships [BZBLP16, Cam17, CZPR17, CC10b, Chu14, DRMMC19, HCLC14, ILB11, JKSK15, KPSL12, OROMAA16, SN10, VTY17, YPH10]. Relative [Dor17, JKJL14, MJC12, RZ12, AvLS14, BLS15, DGWZ13, MV19, Pra19b, SAPR18, SK18, SDS14a, VG11b]. Relatives [ADR14b]. relativity [FG15]. released [ZKD11]. relevance [Bel17, GD17, HM15c, LDVSGDR16, SLGO17, Whi10, Whi15]. relevant [BM13a, BM13b, Gál17, LP18b, MOA16b, MNdF16, PKR15, ZPG^{+}14]. reliabilities [dZLwC^{+}15]. Reliability [Ort18, YYL10, Bor18, EO14, HL18, JOGC17]. Reliability-based [YYL10]. reliable [KB10]. reliably [SYDW19]. remarks [CB18]. remember [Sch14b]. Remembering [BS13a]. remote [ZLN^{+}13]. removal [Egg11b]. Renewable [SCGZSL^{+}13, RFGBMA13]. renewal [LS17a]. repeat [Rou18]. repeated [CÖT16a, Oos15]. repetition [SI17]. replaceable [Bro13]. Replication [JTZ14]. Replies [LZH^{+}13]. Reply [BD16b, BB17a, Pet18b, BH17b, DF15, LM13a, WvEvL^{+}11a, MS16a]. Report [BP11, AP14, BI12a, BI12b, GGG^{+}12, LWB16, LTGH15, ANZ15, CC14]. reporting [CIL^{+}16]. reports [DGD19, GRSFV16b, GRBBS17, JVM17, MBR^{+}13, SSN19, vdBSS18, hHSL19, Jac12, Wad16, Wad17]. repositories [AOFU10, CMRC15, CQB16, DCM16, Fan15a, OMLC15, RCCM14]. repository [Ban18, SZ15]. representation [GCGP10, GT11, GT17, wHwH16, KO19, Mag14b, WW12]. representations [LCS^{+}16]. representative [ADM19, ARE^{+}18, DGWZ13, ZZFD18b]. representing [LGS18]. reprise [DL16a]. Reproducibility [Doc13]. reproductive [BGMB16]. Republic [VFA10, Van14]. republics [IJF16, JJR10]. reputable [KKV^{+}13]. Reputation [GPSM18, CB18, HM15c, dFVDU^{+}19, XWL19]. reputational [Saf19]. reputations [OMMMTLC17]. Requirement [YZW^{+}17]. Requirement-oriented [YZW^{+}17]. requirements [GVS17]. requiring [HO19]. Rescaling [Sch15b]. Research [ADD11d, Ano16c, ALvH19, Bar17a, BGÖ^{+}13, BS16, CLD13, CHWL12, CYW^{+}11, CLkS11, CXZ19, Cop19b, DVB14, DVB15, DVMS17, DPF^{+}16, EG16, FWFM18, GW15a, pGSyW^{+}19, Glä18, HMCL16, HAJ12, HH10, KPS12, KJ14, KR17b, LLCL11, Lin12, LHC16, LZCZ18, MH15, MLVJ12, Moo15, NA14, OO12, dBONM^{+}19, ONB17, PP18, PM18, PKL^{+}16, PL18, PH14, PKSG12, RASP13, SK18, SDS19, The18a, Tsa11, VASNU^{+}19, VPM16, Vin12b, WYH10, WQY12, WLLL12, WH16, YCL^{+}13b, YSND17, YSD11, ZLG^{+}15, ZKC^{+}16, ZTRH18, dZLwC^{+}15, ZDZ^{+}15, ZCMVQS11, Abb11, AJSN18, ARK^{+}15, ADS10a, ADS10b, ADV10, ACD11, AD11a, ADV11, ADD11b, ADD11a, ADD11c, AD11b, ADS11, ADS12, ADC12, ADR13, ACD13, ADV13, AD14, ADD14a, ADM14, ADD14b, ACD15, AD16, ADD17a, ADD17b, ADD18b, ADD18a, AER^{+}14, ACP12, AA18, AhOL14, AW10, ACS18, AEFP16]. research [AKB12, APR19, ACRC17, ABMSSP16, ABSF^{+}19, ATCCAAB19, AATBPAB15b, AMFLH15, AAG14, ÁBDFB19, ÁBV^{+}14, AP16, ALYZ15, AC13, AH11, ABL17, ACC^{+}16, Ano16b, AGLNRR14, APYS13, Asu19, AS18b, ASW18, BD16a, BHB13, BHA15, BY13, BYR13, uHBK19, BHPVdPMR18, BLdlCV17, BFMRM19, BBDS^{+}14, BBJS16, Bas13, BBSS16a, BBSS16b, BBSS16c, BMM17, BFGVV^{+}18, BFS17, BKY^{+}15, BKG16, BPGGdMA12, Ben15, BGJ^{+}16, BH16b, BGM17, Bjø19, BMP^{+}14, BPTG10, BS15c, BS17, BCHH17, BAC13, BGAAM15, BM14a, Bor15a, BHM16, Bor16, BG18, BM19, BS11, Bos10, BDE11, Bou14a, BSK18, BvdB14, BCC^{+}17, BFM^{+}14, BJIB16, BNV11, CFSSP16, CMRC15, Cam12, CV15, CZW13, CRV12, CFL12, CGZ10, Car16, CVC14, CVC^{+}15, CGC18, CFP14, CHL15, CHC17, CZV10, CWJC14, CJY^{+}15, CXpHqZ15, CW17, CLO18]. research [CLSW19, CLLH15, CY13, CG18a, CRMdMA15, CRMPRS18, CDCK13, CC13, CL17a, CFS18, CDD15, Cle16, CFM15, CB16, Coc18, CFdC^{+}14, CH15, CL16, CL17b, CÖT16b, CB19, CV14, CIL^{+}16, CZ18, CL13, DSG^{+}15, DLL^{+}16a, DMM13, DFS15, DR10b, DSM11, DNAH15, DLM15, DGWZ13, Doc11a, Doc11b, DC17, DB19, DXL^{+}12, DWGL16, DKS18, DCY^{+}17, DRS14, DJWS11, ES16a, EBR16, EBK16, EC16, ENST16, ES18, FYC15, FL16, Fan11, FAI^{+}18, Fed13, FPS14, FFR16, FRPP17, FdSdO17, FMPP10, FM11d, FGMM12, FMM13b, FSLR10, FZZ^{+}11, FCWH11, FH13, FLH14, Fuk19, Fuk14, GGG14, GG14, GCLcG15, GRSFV^{+}12b, GRSFV^{+}13, GPN14, Gar15, GY12, GD11, GL15, GKK15, GTD14, GBM^{+}16, GZ18, GVGSEPRC15, GALR16, GTGABAG15, GB14b, GGG16a, GW10b, GW15b, GMM16, GHA^{+}16, GNS^{+}15, GR14, GWB11, GKV11, GAGT15]. research [GKS16, GKS18, GNHT18, HA19, Ham14, Han11, HJL18, Har13a, HHK^{+}12, HH13, HHZ14, HTHB11, HMCD^{+}19, HMI19, Hir05, Hir10, Ho13a, Ho13b, Ho16, HV18b, HRB^{+}13, HRB^{+}14, HS16b, HN16, HKWC15, HBT16, HTL15, HEH18, HC15b, HC15c, HC17, HSLP14, HZ17, HGH17b, HLSW18, HLW19, HC16b, HSXL14, HC14b, HYYR14, hHC15, HC16c, HH18, Hun12, HP10, ILB13, IBL13, Ibr18, IF13, ILP11, ILP13, ILBG14, IL14a, ILGZ^{+}14, ISL18, IQT^{+}19, IMHG12, IMH13, ISR11, Iwa17, JG12, JG14, JKJL14, Jar10, JL18a, JL19, JCK11, JC12, JPZ14, JSZIZ13, JDH12, JH16, Jon10, JDLIV14, JKSK15, KLP12, KGSS16, KGNB11, KLP17, KKLP17, KA13, KO19, KG10a, KM15b, KM15c, Kaz14, KDFL14, Kaz15, KPJ16, KHVGA^{+}16, KBT15, KB11a, KZ13, Kha13a, Kha13b, KCU19, KY16, KLL14, KHH18]. research [KKBW17, KHK13, KB18, KB10, Kli16, KGB^{+}18, Kon12, KJW^{+}17, KSB11, Kos16a, Kos16b, KM11, KS17, KVC15, KK13, KGG15, KJ13, KRR14, KM16, KM18c, Kwi18, KPSL12, KM12, KKT^{+}18, LP18a, Lam12, Lan13, LKP11, LVGV^{+}11, LBRR19, Lee10a, LLL12, LSCK12, LKSK15, Lee19a, LT10a, Lev15, LT10b, LR12b, LYGQ12, LGZ^{+}13, Li17, LXWC17, LRY18, LCWY12, Lia11, LH14, Lja16, LCY14, Lin18, LZH^{+}12, LCC12, LF12b, LLGW13, LZH^{+}13, LFLG14, LLG14, LGH^{+}14, LG15, LX15, LYLD15, LLW^{+}16, LG16, LCLX16, LM16, LF17, cSL10, LdSdFFNM17, LDVSGDR16, LNMQRR15, Lor14, LL12, LRA14, LJJ^{+}16, Lun19, LLYC14, LMdBG16, LWW^{+}11, MLY^{+}14, ML16, MLOY18, dNMVQL16, MC10, MC13, MBP19, MR10, MM14a, MOA16a, MWH14, MR15, MBL18, MARMSG19, MJC12, MHFB17, MHTB17, MBT16, MSP^{+}15, MR18a, Med15, MCvFP16, MHM^{+}12a, MK18]. research [MASM14, MASM16, MCCU16, MLT^{+}14, MNdF16, ML13, MM17a, MM17b, MOO17, MTT15, MGB16, MGC19, MB16b, MS13, Mix18, MK19, MMA18, MHC14, MGGdP17, MPM18, MV19, MS16b, MCB15, MÁB18, Mor19, MRN14, MAGBBM13, MKHB13a, MKHB13b, MKHB15b, Mue16, MT12b, MAGSTRC15, Mus12, NSH^{+}11, NFH12, NHLL17, NBR^{+}11, NPP^{+}12, NHY^{+}14, NQ14, NCG^{+}19, OCJB15, OHT10, Oli15a, OGRMOP19, OM11, OA10c, Ort16, OMJLVSN19, Osw10, PFL19, Paj15, Pal15, Pan14, PK14, PW17, PLJ18, PSZ15, PS16a, PSB^{+}17, PMJF19, PLBZ18, PW13, PLW^{+}15, PR10, PROGMA10, PRRC16, Per18, PRSB16, PR14, PPI17, PL17, PRDG17, PSY^{+}19, PP11, Pou12, Pra11c, Pra14c, Pra16b, Pra17c, Pra18d, Pra18e, PB12, QA19, QZZ17, QA18, RG15, RHMH17, RGLE16, RM18, RRLNAG15, RGGBV16, Rha17, Rig13, RJ14]. research [Rod16, RNB19, RPNC13, RPDCRVRP15, RPGM16, RBBG18, RBC^{+}10, RHGKD16, RG18, SR15, Sah16, SM16a, SA11, SCGZSL^{+}13, SZMS17, SZMS18, SB15, SS10a, SIS17, SW19a, SHR^{+}10, Sch12a, SM12, SHS15, SLISC17, SYP10, SP12b, SD13, Shi11, aSTS17, SC10, SWH14a, SWH14b, Shu17, SV19, SSN19, SUP15, SBSU15, Sin18, Siv16b, SRF16, Smi12, SJOC18, SM14, SY16b, SHL15, SZD16, Soo10a, SK11, SK12, SÁV18, SS15, SRP13, SZ15, SDEB16, SCLC15, SH15c, SZ18, Suo14, SZAJS14, THAL15, TABA16, TFH14, TUCR15, TS11b, TCH^{+}15, TA15, TFJD14, TO18, TK16, The17b, TKA17, TN19, TBMM18, TP11, Tol12, TH13, Tsa15, TCT^{+}13, TE18, TBT19, USPO15, UBTS16, UCH19, US10, UMdSV12, VMM15, VNA16, VDV16, VB12, Van14, VSVR15, VHG^{+}15, VT10, Vin18, VLV14, WS11, WLDW12]. research [Wan13, WLHZ13, WNS13, WWX13, WCL14, WCB^{+}15, WHL^{+}15, WLF15, WZW15, WXZ^{+}16, Wan16, WM17, WWH^{+}17, WHS19, WHW^{+}19, WP17, WH12, WDL17, WFH^{+}16, Won19, WJD15, WLZ^{+}15, WLH^{+}17, XZZ15, XCS^{+}16, Xie15, XM13, XBD^{+}18, YLL15a, YZB18, YWS18, YIK^{+}10, YC10, YPH10, YL12, YCL^{+}13a, YHC^{+}15, YLY^{+}14, YZ17, YYP17, YB14, Yu15, YHL^{+}18, ZYSS14, ZSY14, ZXH10, ZJLG10, ZVC11, ZLT^{+}14, jZhLY15, ZXM^{+}16, ZFY^{+}17, ZG17b, ZLL^{+}17, ZW17a, ZG17a, Zha10, ZS11, ZZ11, ZW11, ZZW14, ZLH^{+}16, ZZW16, Zha18, ZWW^{+}15, ZWW^{+}16, ZM16, LX19, ZT14, ZZ15, ZG13c, ZHMX14, ZLH^{+}15, ZH17, ZLN^{+}13, ZTC15, ZL17, dS17a, dSNV18, dMALIM14, dPdCAdMC^{+}16, vLCCMV13, vR12, vWBS^{+}16, vZ13, vdBS16, Bel13, BH16a, BH16b, MKHB15a, CRLMLM17]. Research-driven [VPM16]. research-group [MKHB13b]. research-orientation [Won19]. research-oriented [CMRC15, FYC15]. Researcher [KKS16, KFB18, WSC16, AML17, BB15, Cab11, Egg11e, Egg11f, Egg13b, GG13, HJL18, JCK11, KA17, LYS^{+}17, LWL17, LNMQRR15, APFR^{+}13, RRLNAG15, RJ14, YY14]. ResearcherID [Jac18]. Researchers [dSF13, AAH10, AI17, ACHVH10, AGLNRR14, BHB13, BSvEK13, BLS15, BM14b, BBS17, BH18b, CGKB18, COS11a, CC11b, CMVP16, DH13a, Dem18, DC19, ES16b, FM12, FSOS12, HAL11, HBA19, IDKF17, ILP11, JX13, JSZ13, Keg15, KFKS15, KCM19, Kra17, Lev15, LM11, LRWS16, LSM^{+}15, LOMLPA^{+}17, MBA13, MBTKA14, Med15, MCL13, OCM^{+}12, PCR18, PPK^{+}16, PDAN19, RCN^{+}14, RGdCMM17, RLW14, SFBS17, Soo14a, The18c, TH19, VG11a, Vil10, WF18, Wil15, YCPS17, ZW18a, ZP16, ddMS15]. researches [HG17]. ResearchGate [AT17, CB18, Cop19b, CB19, Jam17, OMMMTLC17, TK17, YZB18]. Resilience [XM13, QA18]. resolve [PG12]. resource [BGBS18, LWIB16, LL13b, MRR17, PPI17, SL14]. resource-based [LL13b]. resources [CWH11, DC17, DLMX15, FYC15, KKBW17, LCY14, RJ14, SBD^{+}19, Tom17, mYqS15, YQX10]. respect [Bas14]. Response [BBSS16c, GdA14, HL18, Hir19b, HW12, Sch18b, Zon19, CG15a, CF18, LSL15]. responses [MSP^{+}15]. responsible [Ben11]. rest [GB14b]. restless [GR16]. restrict [CLD13]. restricted [ZML19]. restrictions [LL16, LP18a]. restructuring [SRGMF15]. resulting [Ano11]. results [Ano16c, Ano17c, Bel17, Doc13, Fan12, Fan13b, GBM^{+}16, GGS17, GKB^{+}19, GGW^{+}13, HFW^{+}14, ISR11, MC13, MM14c, MKHB15a, MKHB15b, SGN15, VFA10, Van14, VBG^{+}17, WM19]. Rete [BHL^{+}10]. Rete-netzwerk-red [BHL^{+}10]. rethink [Shu17, dS17a]. Rethinking [LSL15, Wra10]. retirement [BDE11]. Retracted [CXWW18, ANA18, BIH18, dSD17, dSBC17, AZSA14]. Retraction [AZSA16, He13, HC17, RV18a, BIH17, JW18, RV18b]. Retractions [RV18b, WA18b, DdS19b, EKR19, RV18a]. Retrieval [Ano15, Ano18a, AF15a, CFM18, Glä15, KMD^{+}18, KHS^{+}15, LCD^{+}14, MS15a, MS15b, MM15b, OI17, SH19, Whi15, Whi18, Wol15, YSM^{+}19, Zit15, AF15b]. retrieve [MNdF16]. retroactive [OMAT19]. retrospective [CMPD19, CIL^{+}16, SFR^{+}19, ZXT^{+}19]. return [Vel12, Yur18a]. returnee [Sin18]. returnee-PhDs [Sin18]. Returns [CDD15, Sch14c]. reunification [AML17]. reuse [JKN19]. Reuters [Har15b, PA12]. reveal [TG18a]. revealed [LZR14, MSDJ19, Rou19, WAT16]. Revealing [OMA15]. revenue [Asa19]. Reverse [KKV^{+}13]. Reverse-engineering [KKV^{+}13]. Review [FE14, HAG^{+}16, LHW16, MFF^{+}16, Pen19, ADV10, AD11a, ADD11a, ACD13, ADR19, AATBPAB15b, AND19, BD16a, BFS17, BK15, BGJ^{+}16, BGBS18, BCJ^{+}17, Bor12, BHJD12, BH15, BM12b, BF17, CGG^{+}17, CPRSFVG19, CCLL14, Che12, Che18b, Cop18, CZ18, DGD19, DFS14, Fan11, Fan15b, FZQ17, Fed13, FLM^{+}19, FCCMTRVR18, dCPF14, Fox17, FT19, GRSFV17a, GRSFV17b, GGG^{+}11, GPSM18, GKB^{+}19, GdA14, GWG17, Hal13, HPS19, HLC17, HH19, HS17, JKN19, JPZ^{+}10, KBAK17, KHVGA^{+}16, KC15, KVC15, KPRT16, KTRP17, LRY18, LF12b, LTK^{+}19, LLYC14, MPY^{+}13, MPS^{+}18, MKHB13b, NG16, ND16, Ort17, PG14a, PS10, PLBZ18, PTMT11, RMCM13, Rha17, Rig13, RCJ18, RT17, RBBG18, SRGMF15, SB17, SZD16, SBM17, SS15, SC13, SZ18, TAA16, TB19a, WvEvL^{+}11a, WLHZ13, WSL14, YLC18, Yu15, ZL18a, ZLL^{+}15]. review [ZZZC16, dZLwC^{+}15, ZDZ^{+}15, Zuc10, dAG13, dCdAMB19, vdBSS18]. reviewed [BSFCC15, MS18a, Moo15, Sak19, VE14]. Reviewer [Bar17b, SA17, BHJD12, DFS14, GRSFVdMA14, Ort17, ZZD^{+}18, Zuc10, dM10]. reviewers [ANFF16, uHBLKH19, Fox17, GRSFV16a, KBAK17, RSGFV18, Sch10b, ZS18]. Reviewing [NRAW17, BWD12]. reviews [GGP14, HCDT16, HH17a, JOGC17, LDG17, WF18]. revision [RCJ18, vL12]. revisit [Hu11]. revisited [AH11, Egg11a, GGW11, wHwH16, LBGBMA10, Lei16, LBA19, MKHB15b, Pra19e]. Revisiting [Bou11, Pra19c, SM12, Sni16, WWL17]. revitalized [MS15b]. revolutions [De 17, LWB18b]. rewards [Shu17, ZS18, dS17a]. RFID [HW10, SZAS16]. rhetorical [BASL16, RM18]. rheumatology [PCRMCB^{+}18]. rice [MRN14]. rich [RPP18]. richer [RPP18]. Ricker [BH17b]. right [Hei19]. rights [CO10, YB14]. rigor [HNG19]. Rise [KO18, FRF^{+}19, Hen16, KB11c, Meh07]. rising [BFHS18, Cho12, DAMC15, NZL^{+}19, ZZW^{+}19b]. risk [KFB18, LP18b, MHFB17, MYN^{+}15, dPdCAdMC^{+}16]. risk-smoothing [KFB18]. risks [MV19]. river [VLV14, WXZ^{+}16, SCLC15]. roadmapping [ZZPG14, ZZP^{+}14a]. robot [Bar17b, LW10]. robotics [Coc18, FSOS12, GL15, KK17, MK18]. Robust [Fre14, ZYX^{+}14, RNB19]. ROC [LD16]. Role [BKSS15, Cha18a, ACFL11, Asu19, Chi14, Col17, Glä10, Glä12, GT18, GPSM18, HV18a, HLC17, KCU19, KCM19, Ley15b, Li18, LCWY12, LOMLPA^{+}17, LLHN17, MHM^{+}12a, OFP16, PTMT11, RPGM10, RPDCRVRP15, RC13b, RCCM14, Sma10, TBS15, WXLL12, WHLP16, WBX^{+}17, Zuc10, vLCCMV13, vWBS^{+}16, vdPR18]. roles [CLL^{+}17, EMSH16, GAPP18, HLSW18, LZCZ18, ONB17, PML^{+}17, RTP17, STCRPA18, WLY14]. Romania [SDS14a]. Romanian [ATM16, GLS16, VPM16]. roots [BSPL19, EBK16, Hou17, TM12]. rough [WYAY12, SDP^{+}19]. route [HLLT14]. routine [LPL14]. RPYS [BHL18, MB14, WB15, CL16]. RSI [AvLS14]. RTRS [AI17]. rule [LdZwC^{+}17, dZLwC^{+}15]. rule-based [dZLwC^{+}15]. rules [ADD11b, CGKB18]. Ruling [WDN17]. run [Kos18b]. Rural [ZY19, WL14].

Russia [Dya17a, Dya17b, GKS16, KGSS16, LM11, MBL18, Ole12, Ter17, ZK19]. Russian [AS18b, GKS18, MMA18, MPS^{+}18, PHBN^{+}15, PD10, TBB^{+}16].

s [Tei11]. S&T [SH15b, BKSS15, De 16b, FKRS14, GB12, ZLLL19]. s11192- [dCPF14]. s11192-016-2142-8 [Glä18]. s11192-017-2490-z [Ano18b]. s11192-017-2506-8 [Ano18b]. SADC [Bos10]. safety [YIK^{+}10]. Saharan [OM11, ZL18a]. salaries [San18]. salient [ALYZ15]. salt [mYqS15]. Same [Ano17c, GGS17, ADR14b, BM12a, Che18a, Che20, GCGP10, GRSFV14a, LS19a, YWL16, dOM16]. sample [HJM^{+}13, Rya16, SYDW19]. SAO [CYK^{+}11, PYK12, WMH^{+}17, WRC^{+}19, YZW^{+}17, YK11, YPK13]. SAO-based [YK11, YPK13]. SAOx [KPL19]. Sapere [DF15]. SARS [KM11, TZ15]. Saudi [SZMS17, SZMS18, SFM16, SLK12]. save [EMSH16]. SBIR [ALvH19, ABL17]. Scale [Pra19f, ADD11c, AD11b, ACC^{+}16, BS17, BYY18, CCM^{+}11, CNPG17, CJW10, CDD15, CQB16, Egg11c, FEHC19, GG14, GKK15, IF13, KRP19, KKS^{+}17, KTRP17, LLH^{+}16, LGD12, LRA14, NF19, Sch14c, SS15, ZZL19, BS13b]. scale-dependence [Egg11c]. Scale-dependent [Pra19f]. scale-invariance [KRP19]. scales [BYY17, Egg10a, Glä10, LR12a, RP17a]. Scaling [LSR13, DLL^{+}17, RPK16, RP17b, RPK17]. Scandinavia [Siv16a]. scapegoat [Zit12]. scatter [Pra19f]. scattering [BG12, BKRG13]. scenario [PB18]. schemas [LS19a]. schemes [DLMX15, TZG15]. Scholar [DLM15, MMOMLC18b, MM17a, OMLC14, GRSFV18, NZL^{+}19, Puu10, SW19b, WYB^{+}17, Agu12, AL12, BI10a, EGU10, Fra10, Gus19, Har13d, Har14b, HA16, Har19b, JN15, MMOMLC18a, Mik10, ML10, MM17b, MOO17, OMCP17, OMLC15, OMAMMLC15, OMAT19, Ort15, SX16, TK17, Wil15, dS18, dWZD14]. Scholarly [CTL^{+}19, MRGT13, Mik17, MTU17, Pen19, SMAABJ11, ADR19, ATK17, AR18, BKRG13, BHL^{+}10, BPVM11, Cav15a, CYT^{+}12, hCcTmWH15, Che18b, Che18a, Che20, COS11b, Das16, DL16b, Din14, EW15, GRTPMLAJ19, GRSFV18, GTMRE^{+}16, GTMRE^{+}19, HH13, HII^{+}18, HT14, JK10a, JZL10, Ke13, KHR^{+}19, KPJT14, Laa14, LYS^{+}17, LTG12, LXDL13, LTK^{+}18, MH16b, MRGT18, MMSS11, MM15b, NSC13, RW11, RPP18, SBD^{+}19, SA16, SST^{+}16, SX16, TA11, Tom17, Tom18, UCH19, VHG^{+}15, Wal15, Wal16, WFG16, WFS16, XWL19, XDB^{+}19]. scholars [AAH10, Bai18, BT18c, CFG^{+}14, Cha19a, CP12c, CP16, Fra10, RC13a, Saa10, SGM^{+}16, Tol12, YMSQ10, ZYSS14, ZNB^{+}17, ZZFD18a, ZL17, BT18b]. scholarship [Cha19a, FL16]. scholarships [BSFCC15]. scholastic [GSM^{+}16]. School [GGH^{+}10, GGG^{+}11, Kim14]. schools [Cle16, KDFL14, MOA16a, MS13, SM14, Tol12, URU10b]. Schubert [Pen19, Egg10d]. SCI [TYWZ12, CHWL12, SZAS16, Shi14, Sot12, WF18]. SCIE [AKB^{+}10, ZLH^{+}16]. SciELO [MSB18, dART^{+}17]. Science [ACF^{+}17, ABMSSP16, Ano10, Ano11, Ard12, BSBG18, Bou14a, CHM15, Cha17b, CXpHqZ15, Che18a, Che20, CD18, CH12, FCWH11, FI16, GAPP18, Hal14, Har15b, HH15b, Hen18, Ho13b, Hu11, HW12, hHSL19, Hun12, IH14, Jac18, KGSS16, LBCO19, LR12a, LHG16a, Liu19, MH16a, MSB18, MPS^{+}18, MMSS11, MM15b, Nic14, PSY^{+}19, RPK18, Sch17b, SLD^{+}17, Soo10b, TFH14, The18a, TYWZ12, WLDW12, Wil15, WF17, WA18b, ZWW^{+}18, ZSY^{+}13, AF15a, AF15b, ADD17a, ATK17, Ahr17, AR18, AW10, ACORC11, AS18a, dSAEE15, APPF18, APLHF18, Ama16, Ano17c, Ano17d, Ano18c, AGLNRR14, BCML19, BSC^{+}17, BI10b, BHPVdPMR18, BFMRM19, Bas14, BFHS18, Ben12, BSK15, BH16a, BH16b, BCJ^{+}17, BGSvdB11, Bor11, BDE11, BS13b, Cam12, CA18, Cav15b, CVC14, CHL15, Cha18c, Cha19b, Cha19c, Chi14, Chi15]. science [CRFM^{+}12, CRBRG^{+}18, CWH11, CG15b, CvLvR11, CDdS^{+}12, DVB14, DVB15, DY18, DRCG17, DVMS17, DXL^{+}18, DGDG13, DLMX15, DW18, EKR19, Etz13b, Fan18, FM17, FLM16, Fie15b, Fie15c, FMP17b, Fra10, Fra17, FEHC19, pGDTP12, GLS16, GG19, GGS17, GXC^{+}19, GKB^{+}19, GB17b, GKV11, GKS16, HG10, HK19, HLE10, HM18, Hei13, HWL11, HG17, HK12, HYC18, HHGZ11, HHDL13, HC12, HSX^{+}15, Hud17, HP10, ILB11, ILB13, IBL13, IS16, JVM17, JC12, JMM19, Jon10, KSSB13, KZSZ19, Kar12, KKLP18, KG10a, KLPP16, Kim10, KK19, Kra16, LL13a, LTM12, LY16a, Lei16, Lev15, LT16, Ley11a, Ley11b, LCR13, LY16b, LRS^{+}18, LLLL18, LYW19, LC18, LPB14, LSM^{+}15, Lin12, LWM^{+}15, cSL10, LOMLPA^{+}17, Lop10, LHBC18, Lun19, LSE^{+}18, MCLL17, MOA16b, MGMY^{+}18, MDG10, MS12, MTT15, MKYM^{+}17, MS16b]. science [MAGSTRC15, NG16, NA14, NvLvR10, NHY^{+}14, NLCC17, NH14, OHT10, Oli15a, OGOPPR17, OA10a, PLA10, PAL13, PHBN^{+}15, Pau10, PZ17, Pou10, Pra18b, QRJ^{+}17, QZZ17, RHMH17, RY14, RMCM13, RRBA10, RRSA18, RT17, RS12, RF19, RG18, Sak19, SH15a, SH15b, SBSR19, SS10b, SYP10, aSTS17, Sil13, SUP15, Sin18, Sma10, Sma11, Soo17, Soo18, Soó11b, SK12, Soó14b, SK14a, SMM15, TLSH14, TW16, yT11, yT15, TT13, USPO15, UMK14, US10, VH17, VNA16, Veu10, Vin18, Waa13, WW12, WWL17, WW15, WYvE11, WG11, Whi10, WAT16, WT15, WG10, WG12, Won13, WB15, Wra18, XTZ15, XGY^{+}16, YDZ10, YL12, YYDH12, YGW^{+}15, YYP17, You14, Zel12, ZVC11, ZZPG14, ZW18a, Zha10, ZG12b, ZG13b, ZYT^{+}16, ZM16, ZG10, ZY15, ZYS16, ZB12, dSF13, AT18, AZKR13, ÁBMB17, AL12]. Science [BI10b, BBDS^{+}14, BBJS16, Cam18, DBO^{+}18, DNAH15, Don17, DGGBDG17, Fra10, FLH14, FGP13, GRSFV^{+}14c, GY12, Gau17, GK18, GBM^{+}16, GZM15, HA16, Har19b, HB17a, HHBB18, HBA19, Ho14, HSPY15, IWK18, IFH15, Jac12, JK19, JZL10, LvI10, LCR13, LRY18, LLG14, LL12, MdBdP^{+}19, MMOMLC18a, MB15, MF14, Mik10, ML10, MMA18, MPH16, OMAT19, PHDC16, PB17b, RAM18, RCETS19, RPK18, SLG10, Sot10, SDS14b, VASNU^{+}19, VBTK19, WFS16, Wil15, XHA^{+}19, YJ11, ZCL14, ZHL19, ZLL19, dWZD14, CMPD19, DAYY18, Emm19, FH16, KA17, LHW16, MH16a, SRW^{+}15, WvBvE11, Waa13]. science-based [HK19, SBSR19, Won13]. science-related [HLE10]. Science-technology-industry [WF17]. science/technology [HM18]. Sciences [EOS12, Har13c, Kim14, KHK13, Mik10, OE15, VE14, ADR13, ACD13, Abt17, ATM16, BHS14, BDF^{+}17, BM14b, BTNS14, BN14, BFM^{+}14, CRV12, Cha18a, CMT19, CMO11, Cha13, hCcTmWH15, Chi14, CRMdMA15, CIL^{+}16, Dya14, EGU10, FI16, GK14, GSE^{+}18, HAL11, Hen16, Hen18, HYYL12, HLY14, wH15, JKJL14, JL18b, KEP^{+}18, Ley13b, Ley15a, LL15, Lin18, LXH^{+}18, MLC14, MH16b, MWH14, MRGT13, MGMY^{+}18, NJ10, OH19, Paj15, PFDL17, PEFP13, Pol16a, RG12, RPK16, RCdJ^{+}14, SND19, SIS17, Sin18, SL12a, Siv16b, SM17, SMY15, SVS18, SW19b, SST^{+}16, TA14b, VO17, WTG15, WLF15, Wra16a, Wra16b, XTZ15, YGD17, YQX10, YLY^{+}14, dSD18a, Fan15a, YGW^{+}15, ZCZ^{+}16]. Scientific [Ard12, AJCACRdMA16, BVB13, BBVO10, BPHL16, CWH11, CdSPdM13, EN17, FLZ17, FKM^{+}15, Fin15, FB16, FZZ^{+}12b, GWP16, HRH10, HG17, Ioa06, JMM18, JKMS17, KHA17, LF14b, cSL10, MR10, MM16, Man15, Mêg13b, Mêg13c, NP11, NT17, PCR18, Pen19, PFPCM^{+}19, Pra17b, SC18, SL13, Sol06, Soo11a, Soo19, WCK^{+}12, WSC16, WYB^{+}17, ZKD11, ZYG15, ZG17c, AAH10, AHUR11, ADR14a, AZKR13, ACFL11, ACF^{+}17, AhOL14, ACORC10, ATCCAAB19, AATBPAB15a, AATBPAB15b, AMFLH15, AM18, AS18a, ACHVH10, ÁBV^{+}14, AYS14, Ama18a, ANOdFC12, Ama15, AdAdAM10, ABL17, AF18, ACAGD^{+}17, Arb11, AJdMA10, ABM19, ALvH19, Aus14a, ÁRS17, AMI18, BKZ^{+}16, BBCP14, BZ17, BSG17, Bas10, Bas14, BSvEK13, BB19, Bha11, BT17, BT18a, BT18b, BT18c, BSFCC15, BW10, BHJD12, BH15, BH17a, BH18a, Bra12a, BM11]. scientific [BMZ^{+}17, BMD^{+}18, BDC^{+}12, CMRC15, CGV12, CPV14, CC11a, CV15, CSS^{+}16, CSR^{+}18, Cav16, Cha14, Cha18b, CLLZ15, CZV10, CDM18, CRR14, CG17, CG18a, CRZGVQMA15, CRAJdMACÁ15, CRZGVQdMA16, CRBRG^{+}18, Cho12, CFM15, CRLMRPA17, CFdC^{+}14, CPF18, CB11, CvLvR11, CDdS^{+}12, CNC18, Dan14, DFG^{+}18, De 17, De 13, DCS12, DQ11, DLL^{+}17, DGDGSV15, DLMX15, ES16a, ES16b, ET15, EGUB12, ELP11, FK16, FCTV12, FFL16, FMS17, dCPF14, Fin11, FVVSGM^{+}18, FESD11, FMPP10, FM11a, FR11, FI16, GRTPMLAJ19, Gan12, GCGP10, GRSFV^{+}14c, GK14, GLS16, Gau17, GLD16, GZGAC16, GZGAC17, GN17, GK18, GPN10, GHT17, GSKM17, GD17, GSMT10, GAPP18, GDP16, GJ11, GF11, GYZ15, GdA14, GNHT18, GHA^{+}15, Hal13, Hal14, HM18, HH13, HHZ14, HH15a, HIG^{+}17, HSAK18, HB17a]. scientific [HTHB11, He13, HT18, HR11, Hir05, Hir10, Hir19a, HP18, Hor18, HN16, HLSW18, HDC13, HCLC14, Ibr18, IWK18, IA19, IT11, ILP11, ILP13, IPIU13, ISR11, IJF16, JvGH10, JYM^{+}16, JC19, JDLIV14, JJR10, KM15a, KMS16, KA17, KMP11a, KTLD16, KG10b, KKL14, KJW^{+}17, Kor18, KPRT16, KTRP17, KMFD12, KBZS15, KPSL12, LL13a, LMM15, LvI10, LRC19, LML11, LS16, LKW^{+}16, LWB18b, LY12, LJKG15, LLLL18, LCC12, LGD12, LXL15, LJC^{+}15, Liu16, LXL16, LNMQRR15, LOPAGS19, Lor10, LRA14, LHBC18, LGS18, LPZ17, ML18, dNMVQL16, MS18a, MVS10, MGM^{+}17, MCLL17, MGLZ10, MB14, Med15, Mêg13a, MLT^{+}15, MMAHS10, MAP13, MH14, Moe16b, MGGdP17, MD18, MAGBBM13, MHM12b, MHKB16, Mue18, NTM^{+}18, NJM18, dANR15, ND16, NJ10, NHLL17, NH14, OPGW^{+}13, PB17a, Pac19, PKR15, PYH16, PLA10]. scientific [PAL13, PRRC16, PROG19, Pet18a, PC18, PEPUT15, PP16, PL18, PRDG17, PLGC18, PLG19, Pra10d, Pra13, QZZD18, QL12, Rai19, RMA12, RMdO17, Ric17, RGdCMM17, RPAMR19, RPGM10, RCdJ^{+}14, RF19, RCCM14, RRRGVD19, SGM^{+}16, SdJDD19, Sal17, San13, SH15a, SH15b, SND19, SIR^{+}14, SW19a, SS10c, Sch14c, SM12, SZZC18, SB14, Sob11, Soo10b, SVCFI14, SK14a, SK14b, SDP^{+}19, SZ12, SLXD15, SML16, SD18, TB19a, TBS15, TS11a, TM12, THFBdMA18, TCB16, Tu19, TBT19, Van10, Var11, VACCAJ18, Vel12, VG14, VT10, Vin18, WXLL12, WXW^{+}13, WMXZ14, Wan16, WM17, WWP17, WBX^{+}17, WFZD19b, WZFD19, WT14, WS10, Won19, Wra10, WD13, WJD15, XXL^{+}17, XG18, XYHD18, Yan14, YPNS14, YC12, YAC10, YL10, YYLW14, ZGCRVQ18, ZCW14, ZHG16, ZCKZ16, ZCZ^{+}16, ZFY^{+}17, ZNB^{+}17, ZRL18, ZMW^{+}18, ZZW14]. scientific [ZZW16, ZZZ^{+}12, ZM16, ZZFD16, ZZFD18a, Zit15, ZP16, dSD18a, dSNV18, dAG13, dOM16, vAvdWvdB12, PDAN19]. scientific/technological [CV15]. scientifically [Sot12]. Scientist [LT10b, LZR14, ADS11, ADS16, ADD19b, FM11d, Hir19b, LBO19, MB16b, Pra11c, The17c, WWX13]. scientists [ADS17a, ADS17b, ADD17b, ADD19a, AKB12, Ama18b, Bas13, BGBS18, BT17, BBVO10, BCZ12, Cab13, CvLvR11, Dya17a, Dya17b, FB10, FSLR10, FI16, GK14, HCL14, Lee19a, LM13b, MBA13, OBG11, PLA10, PAL13, QZL^{+}17, Ric15, SKM15, Sot10, TLSH14, Vel12, WV13, WXLL12, ZZFD18b, BT18a]. sciento [UBTS16]. sciento-text [UBTS16]. ScientoBASE [GSM^{+}16]. scientodiversity [aSS17]. Scientometric [Bha16, CPF18, FJ11, GCLcG15, GK14, GZ18, GGG^{+}12, GKS16, ILP11, JPZ^{+}10, Kra16, Pal15, Pou12, SR15, SDS14a, SBSU15, xShLY^{+}15, TAB13, USPO15, AUS12, AAG14, AJCACRdMA16, BYR13, BMM17, BSvEK13, BGMB16, BPTG10, BCC^{+}17, Cam17, Che18b, CFM15, Coc18, CST11, CA12, CV14, De 17, FAI^{+}18, FG15, FK16, Glä10, Glä13, GdOdAG^{+}13, GAGT15, HLE10, HSLP14, Jac18, Kar12, KGNB11, KG10a, KC15, KGB^{+}18, LPC17, Ley13b, LHM^{+}11, LG16, LCLX16, Mag14a, MMAHS10, MTT15, NTM^{+}18, dBONM^{+}19, Ort17, PK14, Pou10, RRRGVD19, SMLHCP17, SB15, Soo10a, Soo17, SZ18, TCT^{+}13, VAJCC17, Vin17, WS13b, XZZ15, YYS^{+}10, jZhLY15, dSTL18, KK15]. scientometrician [The17c]. Scientometrics [Ano15, GS15, GGH^{+}10, GGG^{+}11, GGH^{+}14, HRH10, KÖG12, KG13, LLRG10, MS15b, OING12, PP11, SSAG16, SBM17, ZVC11, ZPC17, Aus13, Bor15a, BH17a, CP14, DBO^{+}18, DF15, Fan15b, Glä15, GHT17, Har13a, KMP^{+}11b, KP12a, LZFW15, MSDJ19, MM15b, RAS15, Smi12, Vin10a, Vin12a, Vin19, WQY12, XYW^{+}17, YQW13, Yu15, ZG12a, ZZ16, vB13, BT19a, dCPF14, Glä18, OL11, Pra12a, CBF13, GP18b, RM18, RAS15, SOBM16, Sch14b, TG18b]. ScientoPy [RRRGVD19]. SCIgen [LL13a]. SCImago [GNVQdMAG11]. Scope [KSB11, CJW10, CLLL10, DBO^{+}18, KVC15, SDEB16]. scoping [FLM^{+}19]. SCOPUS [PFDL17, AT18, AZKR13, AYS14, Ama18a, BI10a, BI14, BBDS^{+}14, Brz15, CRAJdMACÁ15, CRMdMA15, EGU10, FMM16, HA16, Har19b, HIG^{+}17, KT15, Kra19, MdBdP^{+}19, Man15, MMOMLC18a, Mêg13b, MAP13, MMA18, MPH16, dANR15, OMAT19, PhD18, VBTK19]. score [CFdC^{+}14, CB18, GRSFV11, GGG16b, KPY16, MD12, dFVDU^{+}19, ZCL14, CB19, MSC18]. scores [ALH15, Bar17b, BYY17, BYY18, CSO17, Doc12, Egg10a, Fox17, Glä10, Pra19b, OMMMTLC17]. scrutiny [Ano16c]. SCSs [DLMX15]. Search [Har16d, HG13, Wad16, Wad17, APYS13, Bel17, BM15, DGF17, Gus19, HGH17a, HSPY15, JDH12, Jun12, KM18a, KK17, LL19, MMOMALC16, Med18, MM15a, Mue18, MM15b, RNM18, Saf13, SGN15, ST14c, VG14, ZLL19, vdBBdK16]. search-term [RNM18]. Searching [BNV11, Eto13, Hsi11, SH19, Wad16, Wad17, YSND17]. Second [Pra13, Pra19d, CA12, GA18, MBSB17, TSG13, dSFSF15]. second-generation [dSFSF15]. Second-order [Pra19d, CA12, MBSB17, TSG13]. secondary [SB19]. secrets [ZLL19]. section [HP10, JTZ14]. sectional [BN10, BL11b, qJnShPL17, OO12, VEJC^{+}18b]. Sections [HA19, KKCG18, The19a]. Sections-based [HA19]. sector [Com15, GGG14, KHJ^{+}12, SMLHCP17, SBT18, ZM16, dMALIM14]. Sectoral [CdJD15, GL15, Lan13, Lee10b, Fuk16, SdJDD19]. sectors [Hur17, PY14, SN10]. securities [YLC18]. security [AAS^{+}19, HC16b, MBP19, Oli15b, PFDL17]. sediment [NHY^{+}14]. sediment-related [NHY^{+}14]. SEE [KBZS15]. Seed [RCN^{+}14]. seeing [Cha19a]. seeking [BKRG13]. seeks [FB10]. seemingly [GRSFV14a]. seen [Man15, Mêg13b, WSH16]. select [FMM14, Moe10, PKR15, SP12a]. Selected [Ano10, Ano11, Ano15, SSAG16, AATBPAB15a, BKZ^{+}16, EN17, JL18a, JL19, KY17, MR10, WG12, ZP15]. selecting [ADD14a, BM13a, BM13b, MÁB18]. selection [AhOL14, ACRC17, BA15, Cha19a, Egg11d, GRSFV14b, KC12, LMKG19, LdZwC^{+}17, PTMT10, RNM18, YK14, dZLwC^{+}15]. selections [WM19]. Selective [OMOR13b, BRS^{+}16, LRY18]. Selectivity [KB11b, KB12, PS10]. Self [BCT19, CvLB10, LV11, MRS^{+}16, dSAEE15, BK11, CPY13, EG18, Gál17, GREL14, GSOLHO19, Har12, HEH18, KdBBK15, Laa14, LP18a, Leb12, LWB18b, LH12, SBT18, SAR19, Tod14, WG10, YK14, dS17b, LV12]. self-archiving [Laa14, LP18a]. Self-citation [LV11, MRS^{+}16, BK11, CPY13, EG18, Gál17, SBT18, dS17b, LV12]. Self-citations [CvLB10, GSOLHO19, Har12, Leb12, LH12, SAR19]. Self-esteem [BCT19]. self-observation [Tod14]. self-organization [HEH18, KdBBK15]. self-organized [LWB18b]. self-organizing [dSAEE15]. self-plagiarism [GREL14]. self-propagating [WG10]. self-selection [YK14]. self-symbolizing [BCT19]. Semantic [CDM18, HYF^{+}17, MVS10, PK14, ZLZ19, BCC^{+}17, EMSS16, FZZ17, GM12, GGW^{+}13, HH15a, HQY^{+}18, KWM^{+}18, MM15a, PYK12, RPAMR19, Var11, VG14, WLLL12, WK17, WMH^{+}17, WRC^{+}19, WP17, WK15, YK11, YK12, ZZPG14, ZZP^{+}14a, ZSC18]. semantic-based [WLLL12, ZSC18]. semi [CMVP16, CLB13, GMJ^{+}17, ZLLD19]. semi-automatic [GMJ^{+}17]. semi-nonparametric [CMVP16]. semi-supervised [CLB13, ZLLD19]. semiautomatic [MM15a]. semiconductor [CSC12, CKCK10]. seminal [CH15, HPKS18]. sending [Har13b]. Senegal [Mêg13c]. senior [MS13]. seniority [Zhu17]. sense [CPV14, CZV10, pGDTP12, Kor18]. sensing [ZLN^{+}13]. sensitive [AT18, DB16, Soó14b]. sensitivity [ADV10]. sensitized [WLR^{+}14, ZZP^{+}14a]. sensor [KHR^{+}19, PR10]. sensor-based [KHR^{+}19]. sentence [Mag14b]. Sentences [Sch14b, TL18]. sentiment [IA19]. sentiments [Sma11]. Seoul [Kim14]. separation [KM12]. September [KGB^{+}18]. Sequence [BYY17, Egg10e]. sequences [Pra11a, XLR15]. Sequential [Ora17]. Serbia [IH14, IFH15]. Serbian [SIR^{+}14, SIS17]. series [ADM19, BSG17, Bas11, BOS14, FCTV12, MD18, Pen19, TSRGCCJC14, XGY^{+}16]. series-based [MD18]. service [Jac18, KK17, ZG13c]. services [Dor17, MMSS11, YLY^{+}14]. SERVQUAL [SDP^{+}19]. set [RC13a, SvLVA19, Vin17, WYAY12]. sets [BM13a, BM13b, Boy17b, MRR17, Pra14a, SSG^{+}18, ZAJ19]. setting [HS16b]. settings [MBSB17, OVJM17]. seven [ILGZ^{+}14, NBR^{+}11]. several [BM12a, BTL19]. Severe [vRvLV11]. Sex [LVGV^{+}11, LRWS16, RLW14, SM17]. sex/gender [SM17]. sexual [VASNU^{+}19]. Shafer [WLPH14]. Shanghai [Doc11a, BBV10, DMV10, Doc11b, Doc13, DC14, DC15a, DEC15, FCCMTRVR18, KZSZ19]. shapes [Etz13b, KK19]. Shaping [LSY11, OA10c, AAV13]. Shapley [Tol12]. share [ADV10, MGC19]. shared [Vil10]. shareholder [ML16]. sharing [GPN10, Hag10b, LF12a, PW17, PSY^{+}19]. shift [BWdMA17, BS19, MB13]. shifted [ER12]. shifts [CDM18, Li16, PLGC18, RY14, VGPdlC^{+}17]. shoreline [Pra19f]. Short [Fie15c]. Should [BBV10, BM14a, LIdMAM11]. shoulders [Lar12, MMOMALC16, QZL^{+}17]. show [Joh18]. showing [BL11b, Egg10c]. shrimp [DMM13]. SICRIS [Rod16]. side [OMLC15, Zit11]. sides [dOM16]. SIF [XLR15, XLR15]. SIF- [XLR15]. SIF-indicator [XLR15]. signal [MZ14, WPCG13]. signals [YK12]. significance [Cam14, Egg14a, PRDG17, Sch15a, Wu18, Ye14]. significant [LHWW14]. significantly [CNPG17, Pan14, Yur16a]. silicon [YCK11]. silicon-based [YCK11]. Silva [BI18a, CF18]. similarities [BL17b, CA12, MBSB17, Moe10, MM15a, PYK12, TSG13]. similarity [Cab11, Geo17, GN19, GHvdB12, MVS10, MBSB17, MG12, RKZK18, RPAMR19, RKT^{+}15, RC13b, SS10b, Sch10a, Sch13b, ST14a, TL18, WK17, WRC^{+}19, Zhe19, ZRY^{+}12, ZZFD16]. similarity-preferential [ZZFD16]. simple [AC13, BBL17b, BM14a, CV15, GGP14, Hal14, HR15, Pen19, Pra17a]. simulated [Ley15a]. Simulating [Ano17d, HBT16]. simulation [Ahr17, DGPL15, FLM^{+}19, GDP16, IL14b, Mue18, MKF14, PG14a, WXZ^{+}16, WG11, CGG^{+}17]. Simulations [Sob11, Sch16]. simultaneous [LW15]. sin [DF15]. since [KBL15, ML18, MH15]. Singapore [Hor18, PP18, Pra18a]. Single [Pra14b, BSMD11, BH18b, BL17b, Egg10d, Egg11e, Egg11f, Har16b, HBA19, NPT^{+}15, SB17, TBMM18, Tur16, YWL16]. single- [NPT^{+}15]. Sino [YYP17]. Site [LGL10, Ort15]. sites [WZ19b]. situated [SMCC18]. situation [CRAJdMACÁ15, PYW18, SMY15]. six [Abt17, Fox17, GRG12, SM16b, WWL17, Yur18c]. Sixty [AER^{+}14]. Sixty-four [AER^{+}14]. Size [LAS14, Aus14a, BVZV16, BT15, CC10a, De 13, DC15a, KB11a, LCZ17, LSR13, OMAMMLC15, PDAN19, RP17a, SY16a, SYDW19, Vin17, Wra10, vdBBdK16]. size-independent [PDAN19]. sizeable [GP15]. sizes [Gus19, LABL13, RP17b]. SJR [GNVQdMAG11]. skewness [ACORC11]. skills [SB19]. Sky [CZV10, ZVC11]. skyline [Pra19f]. skyline-shoreline [Pra19f]. sleep [GR16]. Sleeping [DRCG17, HH17d, ON12, ZXZ17, ZLF18, vR17, CL16, DW18, Fan18, Fan19, HH17b, HY19, LY12, Li14, LY16b, EG18, GR16, SSZL18, TVA17, vRW18]. sliding [HC14b]. slippery [Paj15]. Sloan [CZV10, ZVC11]. slope [Paj15]. Slovak [CFP14]. Slovenia [BBDS^{+}14, KLM16]. Slovenian [KKLP17, KMFD12]. Slow [MHFB17]. sludge [ZLT^{+}14]. Small [ALvH19, ZWHH13, Fie15c, GS12, HW10, LGL10, UMK14, ZGL14]. Small-world [ZWHH13]. Smart [ISL18, Li19, xShLY^{+}15]. SME [YKCK13]. SMEs [LKS^{+}14]. smooth [CSC14]. smoothing [KFB18]. SNA [MB19]. SNIP [Moe16a]. snub [TA17]. Social [AMMT16, CGG^{+}17, EOS12, GB14a, Har13c, Kha13a, Kha13b, KHR^{+}19, LL12, MB19, MHC^{+}15, OE15, RTP17, SDS14b, VE14, WZW15, YB14, ATM16, BD12a, BDF^{+}17, BTNS14, Cab11, CSO17, Cha13, CLL^{+}17, Cha19c, hCcTmWH15, CLLH15, CD18, Chi14, CRMdMA15, CyPP12, CV14, DGDG13, DCY^{+}17, Dya14, ENST16, EGU10, FMP17a, FSC14, GW15a, GD17, GXC^{+}19, GF11, GSE^{+}18, HG10, HIG^{+}17, Hen16, Hen18, Ho14, HZ17, HYYL12, HLY14, Hud17, JYM^{+}16, JL18b, JMM19, KC12, KLPP16, KP12b, KEP^{+}18, KM18c, Lev15, LZFW15, LL15, Lin18, LZC17, LXH^{+}18, MLC14, MH16b, MRGT13, MARMSG19, MBTKA14, MCvFP16, MGGdP17, NTM^{+}18, OMOR13b, Ort15, OBG11, Paj15, PFDL17, PLA10, Pau10, PLBZ18, PEFP13, Pol16a, RG12, SMM^{+}19, SND19, SYP10, Sin18, SL12a, Siv16b, SRF16, SJOC18, SM17, SMY15, Soo17, SVS18, SW19b]. social [SST^{+}16, TA14b, VO17, WLMF15a, WLMF15b, WFG16, WKK16, Wra16a, Wra16b, WZ19b, XTZ15, YQX10, ZHG16, ZW17a, ZW17b, The18a, MH16a]. socialization [BK15]. socially [SMCC18]. societal [BM14a, BHM16, BH17a, JH16]. Societies [CGG^{+}17]. Society [GGH^{+}14, LV12, LOPAGS19, SSAG16, SÁV18, yTmShL16]. socio [KMP^{+}11b, MMAHS10, ML13, XM13]. socio-cognitive [ML13]. socio-ecological [XM13]. socio-economic [MMAHS10]. socio-technical [KMP^{+}11b]. socioeconomic [CRZGVQMA15, PML^{+}17]. socioeconomics [BF17]. sociologists [ACS18, Kor18]. sociology [Kor19, MHC14, Var11]. Söderlund [LSL15, MS16a]. soft [WYAY12]. Software [HPS19, RRRGVD19, vEW10, CL16, Fer14, Gar15, GF17, GWA14, ICC16, KLCS14, MPM18, PYH16, VCC12, ZW17c, dSdSSB16]. soil [NA14, ZDZ^{+}15]. solar [DXL^{+}12, FLZ17, HDC13, JCCC13, LZZ^{+}13, LLW13, MnaeR^{+}15, Pra14c, WLR^{+}14, Wu14, YKCK13, YCK11, ZZP^{+}14a, ZYG15, ZG17c]. solid [LFLG14, YCL^{+}13a, YCL^{+}13b]. solidarity [dART^{+}17]. Solla [Ano12c, Ano14, Ano17a, CXZ19, Dan19, KL16]. solo [Emm19]. solution [PP16, ZZPG14, ZSC18, ZZW^{+}19b]. solution-oriented [ZSC18]. solutions [CTL^{+}19, HT18, ZHL19]. solve [HO19]. Some [Bas13, Cam12, Har19a, San13, Whi10, WZ17, BBL17a, Bou14a, Fan13b, GA18, HH10, HV18b, Kar12, McC18, Med15, OCCSM11, OBG11, dSSdMAF14, SVS18, ZG13b, Zit11, dSBC17, ddMS15, vL12, BG18, CRMdMA15, CB11]. Somes [CRAJdMACÁ15]. something [MSC18]. sore [dSM17]. sorites [APPS15]. soundness [EdS19]. soundness-only [EdS19]. Source [WvE13, Chi14, GRSFVCP19, GTC16, Har13d, HWLL14, Liu19, MBA13, PK14, SPS14, SLISC17, XYW^{+}17, ZW17c]. sourced [LSE^{+}18]. Sources [CPV14, HZL^{+}17, LWR^{+}17, LS19a, Ano16d, BG12, BHM16, DRS18, Egg11b, FDVZ16, HPBI^{+}14, HWL11, JN15, LXWC17, Mor19, YLJ^{+}17]. South [Kim14, Soo10a, Soo14a, Won13, ADR16b, Bai18, Ban18, Bos10, CdJD15, CP12b, CyPP12, Fed13, FKRS14, HP18, ILP11, ILP13, JKSK15, KP12b, KB18, KCM19, KPSL12, LPZ17, MOA16a, Mat12, Mat13, Moe16b, Par14b, PP18, PM18, PP11, Pou12, PB12, SdJDD19, SS10c, SP12b, Soo10b, Soo11a, Soo17, Soo19, SK14c, Yoo15, ZCKZ16, vWBS^{+}16]. South-East [Moe16b]. Southeast [BKZ^{+}16, Bar17a]. southern [ZCKZ16, Bos10, Pou10]. Soviet [LA19]. Space [LSC10, ATYL17, BMZ^{+}17, EBR16, GZGAC16, GZGAC17, GHA^{+}16, Hol10, HQY^{+}18, KMS16, KM17a, LM13b, WPW^{+}14, YS13, VDV16]. Spain [BPTG10, BGAAM15, MM14c, MB10b, MDFGAM14, RPSA17, SCGZSL^{+}13]. Spanish [Ard12, BR12, DCGZ^{+}12, EGU10, FCTV12, GRSFV^{+}12b, GSMT10, ILB13, IBL13, LNMQRR15, MRGT13, MRGT18, MBT16, MCB15, OGOPPR17, OMOR13b, PQG14, RGTSLCH14]. spans [MXZ18, WLZ^{+}19]. sparking [HR17]. sparse [KO19]. Spatial [QDK19, ZML19, SdJDD19, Wu13]. Spatial-temporal [ZML19]. speaking [BSvEK13, CFG^{+}14, GAVZAB12, KWW15, ÖS17, TE18]. Special [LHG16a, PC14, Ano15, FR11, GS15, MH16b, RG18, SLGO17]. Specialization [ADD17b, MRGT18, Wra14, AvLS14, KPY16, LIdMAM11, PS16a, ZGCRVQ18]. specialized [CLLL10]. specialties [De 13, HFW^{+}14, Sot12, WWL17, Wra10]. specialty [MY16, ML13, Oli15a]. Species [Soó14b, MSP^{+}15, NBR^{+}11]. Specific [Kol12, CFSSP16, CXpHqZ15, MB15, MKHB13a, MKHB13b, Pie18, PB12]. specificities [ACT18]. specimen [ACAGD^{+}17]. Spectroscopy [WB15, BHL18, CH15, Hou17, MB14, Ter17, MHTB17]. spectrum [MRGT13]. speed [GPN14, SFR^{+}19, You14, ZZ14]. spell [JC11]. spend [PS13]. sphere [RG18]. spillover [QDK19, tScL13]. spillovers [ALvH19, Hur17, PHL17, QDK19, WZ19a]. spin [FRPP17]. spin-off [FRPP17]. splitting [dSAEE15, SKCK14]. sponsored [Che15, MK18]. sponsors [MÁB18]. sponsorship [WS11]. sport [WTG15]. spot [CW17]. spotlights [HGH17b]. spread [Bue15]. Spring [Ibr18, TBT19]. Springer [Che18a, Che20, ENA19, SGY15]. square [Pra10b]. Sri [MR10]. SSCI [SZAS16, WF18]. SSH [SVS18]. SSME [HC15b]. stability [BMD^{+}18, CC14, CFK16, HSW10]. staff [BD10b, KS17]. stage [BND11, Fan19, JKC15, LKS^{+}14, LHLH19, RKT^{+}15, VO17, WT15, WDP11]. stages [HCL14]. stakeholders [dCdAMB19]. stand [GE11]. standard [Cav15a, MT13a]. standardisation [Ano16d]. Standardization [Kra13, TA14a, BH16a, CLkS11, GRBBS17, HP10]. standardize [DDR17]. standardized [ADV11]. standards [BH16b, Ley13a, RNF19]. Standing [QZL^{+}17, BGR19]. standpoint [UMK14]. stands [TBS15]. star [Aus14a, DAMC15, KO18, NZL^{+}19]. stardom [ADS17a, ADS17b, FMU16]. stars [Cho12, HL15, ZZW^{+}19b]. State [FdSdO17, UMdSV12, Ano16b, DG16, ENST16, LWB16, LMR16, PFDL17, RMdO17]. state-of-art [PFDL17]. state-of-the-art [ENST16, LWB16, RMdO17]. stated [ANZ15, WAT16]. statements [LS15, Ric15]. States [ZKD11, Ley15a, MM16, WWP17, CC16, GK18, Lun19, PFL19, SDS19, TBS15, TYWZ12, WMW^{+}13, YK15, Yur18b, Yur18c]. static [MB10a, RC13b, WDN17]. Statistical [JKSK15, SGN15, CG14, CG18b, DMM17, ES16a, Geo17, GE11, Hos11, LKY17, SPB18]. statistically [LB12]. statisticians [FDVZ16]. Statistics [Ley13b, BBL17a, CBWJ18, DFS15, FMM15b, GHT16, LLC^{+}17, LG10, RC13b, TW18, VVN16, vB13, Das16]. status [CHWL12, Chu14, GW15a, JL18b, LGZ^{+}13, Lin12, MSP^{+}15, PJB^{+}12, Pol16a, ZSY14, ZLG^{+}15, Zuc10]. staying [CNPG17]. stem [AW11, ÁRS17, dNMVQL16, ZS11, FB10, TN19]. step [EC16]. Stephan [Etz13b]. stepping [Van12]. stepping-stone [Van12]. steps [De 16a, Ley13a]. stepwise [YYLW14]. STI [DKS18]. Still [Wra16b, Dya14, JKMS17, Pat18, Sch18a]. stimulation [PYL16]. Stirling [Rou18]. STM [SVS18]. stochastic [ACD15, VEJC^{+}18a, XLZ^{+}18, YKLK14]. stock [GRSFV14a]. Stocktaking [BB19]. stone [Van12]. Stopped [LWT16]. storage [WLH^{+}17]. story [GFK^{+}18, LS19a]. storytelling [PC18]. stratagem [KM18a]. Strategic [CLJH12, FFR^{+}17, FFR16, GRSFV^{+}14c, Kos16a, Lee10b, RPGM10, RPP18]. strategies [AYS^{+}13, DMM17, GRSFV17b, HV18a, HG13, HBT16, HSPY15, JKN19, KFB18, LLL12, PYK13, VGPdlC^{+}17, WHL^{+}15, DCGZ^{+}12]. strategy [ACD14, APYS13, BSC^{+}17, GSOLHO19, KK17, LZL10, ZLLL19, GRSFV17b]. stratification [Pra19f, SST^{+}16]. Streamlining [MKYM^{+}17]. strength [CMUdF15, UBTS16]. strengthen [SÁV18]. strengthened [KPSL12]. Strengthening [Lee12]. strengths [LKY17]. stress [MSP^{+}15, SRF16, WH12]. strike [RF19]. Striking [BL17b]. stroke [MKYM^{+}17]. strong [CMUdF15, XZFD19]. Structural [FRPP17, HRC12, LLHN17, MYP19, PY19, AJSN18, BVOL18, BM15, CLLZ15, Col17, KLPP16, LBRR19, Ley11b, PW13, XA15, YC12, Zha18]. Structuration [Ley11b]. Structure [HSK17, KC12, KM11, LX15, RBBG18, SJ19, Vin18, ACD14, Asa19, BASL16, BGSvdB11, CD16b, CLMN19, CL11, CLSW19, CY13, CL17a, DT16, DMM13, De 13, EES13, ELP11, Fan15b, FUR10, HAL11, HBDL18, HC19, HP10, JvGH10, JK10a, KM17a, KLCS14, KTLD16, LMM15, LWM^{+}15, LYLD15, LHBC18, Ma12, dNMVQL16, MM14a, MSC18, NQ14, OZK11, Oze12a, PHS12, QDY14, RAS15, RGCM14, SIR^{+}14, SDT15, SK13, SL10, SG16, SYLC17, TW10, TL18, VYL17, WhCL10, Wan16, WDS16, WWP17, YLL15b, YYDH12, YYP17, Zel12, ZS11]. structured [LLC^{+}17, MKYM^{+}17, RRBA10]. structures [Ano17c, CR18, CFK16, DAYY18, FDVZ16, GRSFV14a, GGS17, KMFD12, KGG15, LG16, MCB15, Oze12b, SK18, SK12, XLZ^{+}18]. student [HB18b, XA15]. students [Lar12, The18c, ZTP18a, ZTP18b]. studied [MS16a, MSP^{+}15]. Studies [GS15, WS13b, DC15b, GTC16, Ham11, HSBW10, HNG19, KHR^{+}19, KKdBK12, KN15, LJS16, LP18b, Li19, LPB14, LAdAMJ17, LSL15, MS16a, MS18a, MSDJ19, MPF18, MYN^{+}15, MS14, MPM18, dBONM^{+}19, PK14, RM10, RCETS19, RPNC13, Sil13, SM15, SM17, SSZL18, SG16, TB19a, TVA17, yTnL17, WL14, XA15, ZLG^{+}15, dSdSSB16]. Study [Iwa17, MLC14, Xie15, ZW14, AUS12, ANZ15, AW10, ACS18, AEFP16, ABMRVZ14, ATCCAAB19, AMFLH15, Ama18a, AP16, AGLNRR14, AJdMA10, AYS^{+}13, Asu19, BCML19, Ban18, BI14, BIH17, BSvEK13, BPGGdMA12, BR11, BSS15, BSK15, BKSS15, BPTG10, BD10b, BSMD11, BND11, Bor18, BPVM11, BvdB14, BN10, BL11b, BRS^{+}16, BJIB16, CWJBT10, CCM^{+}11, Cam11, Cam17, COS11a, Car16, Cha18b, CZV10, CYW^{+}11, hCyL12, Che12, CLO18, CKCK10, CD18, CY13, Chi14, CRFM^{+}12, CHL10, CYK^{+}11, CH12, CIK^{+}18, Coc18, CST11, Col18, CB19, CQB16, CIL^{+}16, CNC18, DD18, DGD19, DRS18, DXL^{+}18, FG15, FLM16, FPS14, FRF^{+}19, FEHC19, GRSFV^{+}12b, GF17, GLS16, GMJ^{+}17, GD11, GN17, Glä18, GRG12, GG13, GdOdAG^{+}13, GB17b, GW10b, GZ14a, GBGB13, GLM11, Har13d, Har14b, HHK^{+}12, HHZ14]. study [HH15a, Hen18, HL18, HG17, Ho13a, HLL14, HNG19, HdSV16, Hos11, HL17, HTL15, HC15b, HFL14, HSLP14, HC12, HW12, HSX^{+}15, hHC15, HC16c, HZL^{+}17, HZQ^{+}17, HYS18, HW10, ILB11, IBL13, IF13, IMHG12, IMH13, Jun12, KBAK17, KZSZ19, KLM16, KM15b, KM15c, KM15d, Kaz15, KY17, KHJ^{+}12, KLCS14, KCP11, KBL15, KWW15, KTT11, KJ14, KM18c, KKT^{+}18, Laa14, LL16, LM19, LKP11, LT10a, LQW17, Lin11, LHC16, LZH^{+}12, LZH^{+}13, Liu13, Liu19, LLP^{+}16, LABL13, LSE^{+}18, MPY^{+}13, ML16, MM14a, MB19, MB13, MHFB17, MSL11, MYN^{+}15, MHC14, MZE19, MCB15, MKP16, MM15b, NSKO15, NASR11, NJ10, NF19, Nii17, NBR^{+}11, NT17, OO12, APFR^{+}13, OMOR13b, Ort11, Ort15, PYW18, PSB^{+}17, PW13, PIB18, PHV17, PEFP13, PB17b, PLG19, QDY14, RHMH17, RASP13, RMdO17, RCdJ^{+}14, RF19, RG18]. study [SMLHCP17, SM16a, SPB18, SCGZSL^{+}13, SCGZR16, Sch13a, SJ10, SFR^{+}19, SL12b, SMY15, SW19b, SST^{+}16, SH15c, SWCH14, TG18a, TUCR15, TCC17, TA15, TA18, TFJD14, Tod14, Tom18, TA17, Tsa15, yT15, UMdSV12, Var11, VHG^{+}15, VAJCC17, WXLL12, WLDW12, WLR^{+}14, WLN^{+}14, WLY14, WLD^{+}14, WPW^{+}14, WDS16, WFS16, WF18, WAT16, WY12, Wra14, XGY^{+}16, XYW^{+}17, XBD^{+}18, XDB^{+}19, YWL16, YIK^{+}10, YQX10, YCL^{+}13b, YKCK13, YKLK14, YJ11, ZPG^{+}14, ZVC11, ZZPG14, ZYG15, ZGY16, ZLW16, ZG17d, ZG17e, ZW17a, ZT18, ZSC18, ZT19, ZY19, Zha10, ZC14, ZLLL19, ZGL^{+}17, ZM16, ZG13c, ZXT^{+}19, dSD18a, dWZD14, dPdCAdMC^{+}16, vLvWW16, vdBS16, vdBBdK16, PD10]. Studying [MAP13, vdBSS18, BGSvdB11, BM19, GWS15]. sturgeon [JG12, JG14]. styles [AS18a, MHC14]. stylistic [Hud16]. sub [ACORC11, HHZ14, ILGZ^{+}14, OM11, XLR15, ZL18a]. sub-areas [HHZ14]. sub-fields [ACORC11, ILGZ^{+}14]. sub-impact [XLR15]. sub-Saharan [OM11, ZL18a]. subcontinent [MR10]. subdiscipline [SL12b]. subdisciplines [TC13]. subdivision [LXWC17]. subdomains [APPS15]. Subfield [Bou14b, hHSL19, QA18, TT13, ZY15]. subfields [QRJ^{+}17, Sot10]. subgroups [CH14, Keg15]. Subject [LJS16, LCS^{+}16, AZKR13, AW11, BBJS16, Cam17, Cam18, CWH11, DGDG13, DGGBDG17, GRSFV^{+}14c, GNVQdMAG11, GGS14, GWBSVWB13, HH18, JN15, KB18, MB15, Pin15, RSGFV14, SP12a, TZG15, ZJLG10, ZCL14, ZLW16]. subject-classification [TZG15, ZLW16]. Subject-method [LJS16]. subject-specific [MB15]. subjects [CHL15]. submission [ADD14a, AND19, BHDI18]. submitted [BHJD12, Hud16]. subscription [PROG19, SA16]. subscription-based [PROG19]. subsequent [SRW^{+}15, vWWtH14]. subset [ADM19, FMM14]. subsets [Vin19]. Substance [WST14]. substitute [Ber18]. substitution [MV19]. succeed [EC16]. success [AGLNRR14, BSvEK13, BSK18, CT15b, FGMM12, FMM13b, FMM14, GZGAC16, GZGAC17, HV18b, HK12, JPT13, KCU19, Kis11a, KWW15, WWC19, Yur18b]. success-index [FGMM12, FMM13b]. Successful [KJES16, CZPR17, GG19, KM17b, Pou12]. successfully [LHLH19, MBTKA14, PMN16]. succession [CIK^{+}18]. successive [Par15]. Succinct [GMSZ18]. such [CdMCdMMdP17]. suffer [GRSFV16a]. suffice [Ama18a]. sufficiently [dSD18a]. sugarcane [dBONM^{+}19]. suggested [Osw10]. suggestions [RMCM13]. suitability [ADV13]. sum [Doc12, LWT16]. summarization [JKN19, KMD^{+}18]. Summer [GGH^{+}10, GGG^{+}11, GGG^{+}12]. superficial [vWWtH14]. supernetwork [ZZW19a]. supervised [CLB13, Ken18, ZML19, ZLLD19]. supervising [KKE13]. supervision [dSNV18]. supplemental [GP15]. supplements [Che15]. Supply [SDT15, BLdlCV17, NSH^{+}11]. support [Dem18, FB10, GWG17, HR11, MT15, PG14a, SFBS17, SBSR19, Zhu17]. supported [Bor16, KKL14, WLF15]. supportive [PC18]. surname [Kis11b, KB13]. surname-based [Kis11b, KB13]. surrogate [SBD^{+}19]. Survey [CZV10, BN14, CTL^{+}19, hCyL12, HM15c, HJM^{+}13, LJMF15, NJM18, NH11, RRRGVD19, SGM^{+}16, SK17, vEW10, ZVC11]. Surviving [ZXM^{+}16]. susceptibility [dOM16]. Sustainability [SE18]. sustainable [HHZ14, ILGZ^{+}14, QL12, Sot12, ZH17]. sustained [WJCC19]. swan [ZZY19]. swans [JSZ13]. Sweden [BL11b, BN14]. Swedish [BN10, Bre10, SM15]. swing [Mix18]. symbiotic [Wol15]. symbolizing [BCT19]. synchronic [YS14]. synchronous [ZG17e]. synchrotron [Hal13]. syndrome [LLP^{+}16, CvLvR11]. Synergy [KM12, LZ14, LPL14]. synonym [GKF17]. syntactic [VG14, WZS12]. syntactic-semantic [VG14]. synthesizing [Hsi11]. synthetic [LMdBG16, SKY17]. System [Ano11, ADR16b, AI17, APPF18, APLHF18, Ama16, BR12, BH16a, BH16b, BCJ^{+}17, Che12, CC12a, CH15, ES16b, FKM^{+}15, FT19, GRBBS17, GMJ^{+}17, HS16a, HFL14, JK19, Jar16, KGSS16, KPRT16, KPSL12, LBW17, LLW^{+}16, MBP19, MMAHS10, Mou16, OROMAA16, OMOR13b, PJL19, PC18, RVFEdlM10, RG12, RP17a, SH15a, SN10, WWH^{+}17, Yoo15, ZK19, ZZL^{+}10, ZLW16, ZQH^{+}17, Zuc10, HFW^{+}14]. Systematic [MF14, GWA14, HSPY15, PLBZ18, Rha17, SMCC18, ZL18a, dCdAMB19]. Systems [Glä18, LHC16, UMdSV12, dSFSF15, ACD11, ADV11, ADD14b, ÁRM13, ASW18, Bue15, ÇAAÇ15, Cha18b, CST11, Doc11b, Doc12, DCY^{+}17, FM11c, FM12, Fuk16, GL15, GWA14, HC15b, KRP19, Kha13a, Kha13b, KC15, KPRT16, KTRP17, Kwi18, LL16, LRS^{+}18, LYLD15, LTK^{+}19, MT12b, MMSS11, NH11, OA10a, OCCSM11, PK14, PHS12, RP17a, SGG^{+}14, SV19, TP11, yTwTlW19, ZLYF14, ZZ14, ZRY^{+}12, ZL17, RPK17, Doc11a].

tables [HC14a]. tabulation [Hos11]. tag [DGF17]. tags [KC12]. tail [Glä10, Glä13, YR10]. tail-core [YR10]. Tailor [Gal11]. Taiwan [Liu16, SWH14b, Cha14, CYW^{+}11, CSC12, hCcTmWH15, CS11b, HIC12, Hu11, HWLL14, KLP12, LLL12, LYGQ12, Lin12, LXL15, LXL16, MC12, PP18, SWH14a, TCT^{+}13, Wu14]. Taiwanese [LCC12, Sha12]. take [Hag10a, KKV^{+}13, LHLH19]. take-off [LHLH19]. takes [Hir10]. Taking [GTMRE^{+}16, GTMRE^{+}19]. tale [DSH^{+}10, Pra18a]. talent [MGGdP17]. tales [Jac12]. talk [Bru10]. tapered [San12c]. target [DC19, HYS18]. target-policy [HYS18]. targeting [WF17]. taxonomy [De 16b, JC12, LSS15, LSS16, Rod17]. teach [SAPR18]. teacher [Tod14]. teaching [DH13b, GMM16, WHLP16]. team [BT15, LPMK17, PRSB16, TCR10, XWL19, ZZW14, ZZW16]. teams [ADD17a, CÖT16a, MARMSG19, PRDG17]. Tech [ZPC17, Suo14]. TechMining [PC14, VGPdlC^{+}17]. Technical [AP14, MBP19, CNPG17, CNPG18, HCLC14, KMP^{+}11b, SHS15, VG14]. technique [JDH12, JH16, LW10]. techniques [ABM19, CNC18, DAMC15, Lee10a, MVS10, NZL^{+}19, RHMH17, YLSW16]. techno [ÁRS17, PLG19]. techno-scientific [ÁRS17, PLG19]. Technological [CLLL10, HLLT14, LGR17, LZZ^{+}13, RGdCMM17, SS16, WSH16, AAV13, BM13a, BM13b, BM11, CV15, CWL10, CJW10, CSC12, CC12b, Com15, DQ11, EES13, FLZ17, GS12, HR15, HDC13, HCS^{+}15, JK10b, KCK14, KS18, Kli16, LLC^{+}17, Lee10b, LKS^{+}14, LZ14, LAHH15, LNRSRBB18, LCFC14, LP10, MYP19, PYK12, RR17, RBF^{+}10, SPS14, SFNO12, SÁV18, SD18, WLD^{+}14, WY12, Won13, Wu14, YZW^{+}17, YK11, YK12, YPK13, ZHG16, ZQH^{+}17, ZZZ^{+}12, ZHZY19, vRW18]. Technologies [SJ10, AUS12, ÁRS17, Bai18, CS11b, CMM17, DDS^{+}19a, EMSS16, ÉMS^{+}13, HLL14, HSPY15, JAAA18, KCK14, RNM18, SBSR19, WF17, YKCK13, ZyZZ^{+}11, ZLLD19]. Technology [Ano10, BM15, CK14, HH17e, JKC15, KHK13, LKS^{+}14, LL19, LHG16a, MEG15, NASR11, SK14a, SL10, ACMP13, Ahr17, ATYL17, APPF18, APLHF18, Ano17d, Ano18c, APYS13, AYS^{+}13, BFHS18, BKY^{+}15, BGM17, BDE11, Bou14a, BPHL16, BRS^{+}16, BSPL19, CGV12, CD16b, CWL10, CLJH12, CYK^{+}11, DSG^{+}15, EMSS16, Etz13a, FM12, FK17, FK18, pGDTP12, GM12, HM18, HF19, Har14a, HTL15, HSWC13, HSX^{+}15, HHA^{+}16, HZQ^{+}17, Hur17, JYW11, JC12, JBC19, KZSZ19, KKCG18, KW15, KGL^{+}14, KL17, KGY^{+}17, KS18, KPL19, KKK^{+}14, KPY16, LLL12, Lee15, LS17a, LM13a, Ley15b, LYWSV13, Li15, LCZ17, Liu13, cSL10, LLW13, LLHN17, MP15, MTT15, Moh12, MRN14, NH14, OA10a, Ort11, PY14, PJY17, PHL17, PLJ18, RG15, RRBA10, RS12, RG18, SH15b, SZAS16, tScL13, SL16, Suo14, TR14, US10, WLN^{+}14, WLD^{+}14, WPW^{+}14, WMH^{+}17, WZ19a]. technology [WT15, WG10, WG12, WS13a, Won13, WF17, YSY^{+}13, YLH^{+}17, ZPG^{+}14, ZZPG14, ZZ14, ZZP^{+}14a, ZFY^{+}17, ZG12b, ZG13b, ZZP^{+}14b, ZZW^{+}19b, dPSS18, KK15, vdPR18, ACF^{+}17, Mad15, LM19, Tom18]. technology-class [HF19]. Technology-function [HH17e]. Technology-industry [CK14]. technology-relatedness [LLW13]. tectonics [MB13]. Teixeira [CF18]. Teixera [BI18a]. telescopes [Tri10]. tell [BSG17, CGPT15, ENA19]. Temporal [BIH18, wH15, CU16, HCLC14, SZ15, ZML19]. temporary [COS11a]. Ten [FS12, DGWZ13]. Tendencies [EGUB12, Fan18, FdSdO17, PNVCB18]. tendency [CjZZ^{+}19]. Tenure [Yur18c, Mix18]. terahertz [Liu13]. term [KJES16, KH17, LT16, LSS15, LSS16, RNM18, SVS18, WJCC19, ZZPG14, ZCMVQS11, vEWNB10, HWQ^{+}18]. term-document [KJES16]. terminology [GG15a]. Terms [GWBSVWB13, BS15a, GK19, Ley13b, LZ14, Li18, QA19, RS12, XGY^{+}16]. Territorial [DdlPPL^{+}19]. territories [XTZ15]. territory [FP18, LJJ^{+}16]. terrorism [Mag14a]. test [Har13d, LLCL11, Wal15]. Testing [HM18, LB12, Bor18, Glä10, HL18, PLT14, SC13]. tests [Sch15a, SGN15, Whi10, Wu18]. text [BWbH^{+}18, Cab14, CB19, GRBBS17, HII^{+}18, HT18, JKN19, JN15, Jam17, JS15, LAL15, MXZ18, MVS10, McC14, Moh12, PYK13, PC18, RHMH17, RNF19, SH19, SK13, SH18, TA18, UBTS16, WRM17, kWhHRkS10, WLZ^{+}19, WA18a, YLL15a, YLH18, ZLG^{+}15, ZHZY19, PT17]. textbooks [Kor18]. textual [LAL15, Moe10, MG12]. their [AAH10, ADS11, ADS17a, ADS17b, ALH15, Aus13, BPGGdMA12, BPVM11, BL11b, Cam18, CGK^{+}14, CÖT15, CLO18, DCS12, DX17, Fan15a, FMS17, FK17, FK18, Glä12, GW10b, GWG17, HJL18, HC14a, Har17, HR15, KGSS16, KKE13, LVHS^{+}15, LRWS16, LYWSV13, LIdMAM11, MSYW12, MG12, MCB15, NA18, OMR14, Ort11, PRA16a, PML14, RPK17, RCCM14, SBB16, SYDW19, Sot10, TVA17, Tri10, VSS12, VBG^{+}17, Vel12, WHC^{+}13, WLY14, WLC17, WS10, YST12, YP19, YCPS17, ZHG16, ZCL15, ZW18a, ZXLX14]. Thelwall [KL16]. them [Gus19, Koz15, MCL13, VH17]. thematic [Ano17c, GGS17, OGOPPR17, OA10b, OA10c, RFGBMA13, SK18, UBTS16, XLZ^{+}18]. Theme [WLH^{+}17, SZ18, TA17]. themed [jZhLY15]. themes [ALYZ15, KW15, TA17, ZW14]. theorem [MOA16b]. theoretic [GRSFV^{+}13]. Theoretical [Egg13c, SPB18, BBL17b, BSPL19, Hei13, KD18, Pra17a, dACdFC18]. Theories [KLPP16, GTAG18, Sch15a, TG16, Wu18]. theorisation [Bor15b]. theorization [Won19]. Theory [ER12, Bur14, CBKL13, FPS14, FE14, JPT13, KM18b, LL13b, MB10a, Oos15, PJL19, Pra11d, RPSA17, SRF16, tScL13, Whi10, Whi15, WLPH14, ZAJ19, ZLH^{+}15]. therapy [LL12, SDS19]. there [BFMRM19, BS19, Dya14, FK16, KFKS15, MR15, MCL^{+}11, Mou15b, APFR^{+}13, PL17, PZ17, Pra10e, RY14, Sch14c, Wu18, vR17]. thermodynamic [ES18, FM11c, Pra19a]. Thesaurus [Boy17b]. Thesaurus-based [Boy17b]. theses [Ban18, ESB15, FCTV12, KHH18, LGZ^{+}13, RPSA17, Var12]. thin [HWLL14, YCK11]. Things [MCvFP16, YLL15b, YWG14, NT17]. think [BCZ12, MC15]. thinking [XM13]. Third [BP11, WHC^{+}13]. Thirty [SRP13, dCdSNB15]. Thomas [MB10a]. Thomson [Har15b, PA12]. those [SRW18, VVN16]. thought [Ben12, URU10b]. thoughts [GM13]. threat [PIB18]. Three [FE16a, FE16b, KdBBK15, WC18, Bai18, Bor15a, EO14, EMH^{+}10, FPS14, Fra14, FZZ^{+}11, FZZ^{+}12a, FZZ^{+}12b, HCS^{+}15, HC16c, Li15, LP18b, Pra14c, Pra17c, Sch12a, Sch10b, TA17, yT11, Wal15, WOW10, YWC12, YWW17, ZYNZ18, ZG11]. three-dimension [Li15]. Three-dimensional [KdBBK15, LP18b, Pra14c, Pra17c, Sch12a]. three-journal [Bai18]. threshold [CSC13, GRSFV11, LHLH19, SS10c]. throughout [LJJ^{+}16]. ties [CMUdF15, KA17]. till [BMM14]. Time [Bas11, BOS14, CU16, FCTV12, Fin11, ILBG14, MD18, AHUR11, ADD11c, ASPF^{+}16, Cam10, CGPT15, Cle16, Egg10c, FSLR10, HC15a, HCDT16, Har16a, Har16c, HB18a, IMSK14, Jar16, JPT13, MFF^{+}16, RCJ18, SC10, Tod14, Wan13, WFH^{+}16, ZFY^{+}17, ZNB^{+}17]. Time-aware [CU16]. time-dependent [WFH^{+}16]. time-normalization [HB18a]. Time-varying [ILBG14]. Times [BSFW10, Har19a, SGSS17, Saf19]. timing [MTU17]. tip [Pou10]. tissue [Zha14]. Title [GW17, ADD16, BS15a, GMSZ18, JN11, NG16, RMH14, ST14c]. Titles [MASM14, MASM16, ABMRVZ14, ABRVZ15, CP16, GZJ^{+}15, Har15a, Hud16, SM16a, LX19]. Tobin [CS11a]. today [GN17]. together [Mus12]. tolerance [MSP^{+}15]. Tommaso [Pra17a]. Too [BSFW10, Oos15, KB12]. tool [ADS12, Cam17, CFL12, CL16, FCCMTRVR18, GRSFV17b, JKJL14, PB12, RRRGVD19, SYLC17, WS11, ZZPG14, BHL^{+}10]. toolbox [PC18]. toolkit [TT13]. tools [MCR^{+}12]. Top [KB11b, KB12, KHK13, LPMK17, ADS16, ADS17a, ADS17b, ADD19a, ADD19b, ACRC17, uHBK19, CMVP16, FH13, GRSFV12a, GKS18, Ho13b, Keg15, KHA17, Kos18a, Kos18b, Kwi18, LL16, Ley12, LSM^{+}15, LSE^{+}18, MCL13, SM16b, SLD^{+}17, SUP15, SK11, TBB^{+}16, WW15, WWP14]. top-10 [Ley12]. top-25 [Ley12]. top-cited [Ho13b]. top-level [Keg15]. Topic [ADD^{+}15, BGG^{+}17, MLOY18, MCLL17, Yan14, Boy17a, CGKB18, Cam12, DFS15, DXL^{+}18, FMP17b, HFL14, Kli16, LKSK15, LJS16, Li18, LXDL13, MHKB16, Nic14, QZZD18, Sch14c, SHK14, SYLC17, yTmShL16, VBG^{+}17, VYL17, WLD^{+}14, WHS19, XCS^{+}16, YPNS14, ZMW^{+}18, ZLLD19]. Topic-based [ADD^{+}15, Yan14]. topic-dependent [ZMW^{+}18]. topic-level [LXDL13, Yan14]. Topical [BL15, Cab11, CGC18, LYW19, MP15, ZMW^{+}18]. Topics [PN15, BHPVdPMR18, CYH13, GT11, GT12, GT17, HGH17a, HQY^{+}18, KWS17, LM19, Lam12, MCLL17, OHT10, OH19, PPI17, RBF^{+}10, RRRGVD19, The17b, TH13, WCL14, WLD^{+}14, XGY^{+}16, YDJ12]. Topological [AOd15, Ama15, KM17a]. Topological-collaborative [AOd15]. topology [KJS14]. Toronto [OCJB15]. total [Egg10d, Rou12b]. Totalized [Pra18e]. tourism [FRdA16, jZhLY15, ZTC15]. tourism-themed [jZhLY15]. Tower [HAL11]. TR [AÇA^{+}14]. trace [HCS^{+}15, HZQ^{+}17, ZZP^{+}14b]. Tracing [ÁRS17, JDH12, JH16, MB14, WWX13, WA18a, ZZY19]. track [Ama18a]. Tracking [CMM17, HFC11, MAA18, SKY17, TK10, WFG16, YCPS17, MH14, SD18]. trade [tScL13, WK15, Yur16b]. trade-related [tScL13]. trademark [Hei19]. trading [ZY15]. traditional [FZZ^{+}12a, HK12, HZD^{+}15, NTM^{+}18]. traffic [Jun12]. tragedy [dM10]. trainee [OO12]. training [KK18, SK17]. traits [CC10b, RRL16]. trajectories [AAV13, HY19, KS18, KPY16, Ley13b, Won13]. trajectory [Cha18a, NPP^{+}12, ZTP18a, ZTP18b]. transaction [FPS14, HSK17, SL16]. transatlantic [BCHH17]. Transfer [LM13b, Ama18b, BFGVV^{+}18, Etz13a, GN17, HL13, KM12, LM13a, Ort11, PLJ18, TR14, WZLZ13, WA18a]. transferral [Hos11]. transferred [WPCG13]. transfers [CMM17, ZLL^{+}15]. transformation [CW17, JKN19]. transforms [LW15]. transience [CHM15]. transistor [HWLL14]. Transition [HLE10, Par14c, CSC14, Fan19, IMSK14, JK19, MB10a, PN15, SFNO12]. transitional [CP14, Sko14]. transitionary [FK17, FK18]. transitions [KKS16]. translation [DC15b]. translational [MLY^{+}14, RRLNAG15, VNA16]. transliteration [FGP13]. transmission [ZS17]. Transnational [FS10, GS12]. transparency [SMLHCP17]. transparent [GGP14, dSM17]. transport [BSS15, LG16, NRAW17]. transportation [dSdSSB16]. Travel [CRBRG^{+}18]. treated [ADS10b]. treatment [DSG^{+}15, YYS^{+}10]. Tree [MSP^{+}15]. trees [GVGSEPRC15, WBH^{+}12]. trench [MM18, MM19]. Trend [BSC^{+}17, LNK^{+}14a, LNK^{+}14b, SHS15, UHAR12, AP14, CH12, EMSS16, FAI^{+}18, HW12, Kra16, LH14, LWW^{+}11, NPT^{+}15, dANR15, RRRGVD19, SP14, SZAS16, SB14, VSVR15, WH16, YHC^{+}15, YJ11, YLH^{+}17, ZLG^{+}15, Zha14]. TrendNets [KO19]. Trends [Fra14, GVGSEPRC15, LLYC14, MMA18, OHT10, TCH^{+}15, WH12, AEFP16, AATBPAB15a, ÁBV^{+}14, AW11, AS18b, BKY^{+}15, BMP^{+}14, BRS^{+}16, CHWL12, CZW13, Cav16, CWL10, CLLH15, CYK^{+}11, CAS16, DVB14, DVMS17, DC15b, DJWS11, Fer14, GFC18, GB16, Han11, HSL^{+}14, HYC18, KSSB13, Kar12, KO19, KC15, KvES11, LQW17, LFLG14, LF14a, LLW^{+}16, LCLX16, LL12, MLY^{+}14, MR10, MH15, NHY^{+}14, OGRMOP19, PJL19, PLW^{+}15, Pri15, RG15, RBBG18, SK18, SZD16, Soo10a, Soo19, SVS18, SB19, TYWZ12, TZ15, TW16, Tsa11, Tsa15, VASNU^{+}19, Veu10, WWX13, WHS19, Xie15, YCL^{+}13a, YLY^{+}14, YWG14, YK11, YPK13, ZXH10, ZG17a, ZWW^{+}15, ZWW^{+}16, ZLN^{+}13, ZDZ^{+}15]. Triadic [Mes11, TB19b, WBX^{+}17]. trial [KCT^{+}17]. trials [AAB^{+}13, TK16]. triangular [GHT16]. tribology [EBK16]. tribute [GZ18, LWB18a]. TRIF [PG12]. trilateral [Wad16, Wad17, Wad18]. Triple [Chu14, CP14, IL14b, KZSZ19, KP12a, KP12b, KHJ^{+}12, KPSL12, LPL14, MGMW14, Par14b, Par14c, Sko14, ZZP^{+}14a, HMK^{+}12, KCP12, Par14a, PM18, SLK12, Yoo15]. Triple-Helix [LPL14, SLK12]. TRIZ [ZZPG14, ZZP^{+}14a]. Tropical [VSS12, GVGSEPRC15, STCRPA18]. troubles [Hen19, Hen20]. true [Har16b]. Trump [AM18]. trust [ACT18]. trustworthiness [WAT16]. truth [Mue18]. Tsallis [AdAdAM10]. Tucker [LGD11]. Tucker-2 [LGD11]. tumor [EG18]. Tuple [Par14c]. Turing [Fie15b]. Turkey [SSAG16, CA18, DBO^{+}18, DT16, KKCG18, KGZML^{+}19, KJ14, OMA15]. Turkic [SC18]. Turkish [AÇA^{+}14, BGÖ^{+}13, Dem18, Gar15, KKE13, Oze12a]. turn [ACORC10, BCC^{+}17, KN15]. turnaround [RCJ18]. turnover [Yur18c]. tweeted [VH17, ZW18b]. tweeters [HBS^{+}19]. tweeting [Ort16]. tweets [AM18, LF17, Sni16, VH17, XYHD18, dW15]. twentieth [BCC^{+}17]. Twenty [CG18b, TW16, WOW13, CAS16, GZ18, SCGZR16]. Twenty-first [TW16, GZ18, SCGZR16]. Twenty-five [WOW13, CAS16]. Twitter [BH16c, BH18a, CP12b, HB18a, HT14, KP12b, LYS^{+}17, Ort16, SBA^{+}19, SLH18, VH17, Web16]. Twitter-based [KP12b]. Two [AC13, Har19b, KGY^{+}17, LPL16, LTK^{+}19, AAH10, APPS15, BCML19, BSB12, BND11, BG18, Cam11, CD14, DSH^{+}10, EC16, FKRS14, HC16a, hHSL19, JvGH10, KLP17, KY17, KJ14, LKS^{+}14, LCS^{+}16, LYGQ12, Li19, LS19a, Lor14, Moe16a, NJ10, PFL19, PRRC16, PPE14, Pra18a, SGG^{+}14, Sch15a, Sch14a, Sch17b, Sch10b, SYDW19, cTnHwH17, WDP11, Wu18, dOM16]. two-dimensional [LYGQ12, Li19, NJ10]. two-exponent [cTnHwH17]. Two-phase [KGY^{+}17]. two-stage [BND11, LKS^{+}14, WDP11]. two-step [EC16]. two-year [Cam11]. Type [Abb13, Asu19, BHH18, BPVM11, CCM^{+}11, CV15, Don17, Egg11c, Egg14b, ER19b, Glä12, JN11, LZGQ13, MR13, Pra19d, Rou12a, SZMS18, Sch12b, YWL16, Yu17]. Types [VG17, XHA^{+}19, GD11, HP10, KBT14, KWW15, MK18, Puu10, RP17b, Soo17, SH18, TSRGCCJC14]. typical [SVS18, WYY11]. typically [AC13]. typology [SK11].

U.S. [ACORC10, ABL17, ALvH19, uHBK19, BRS^{+}16, CSC13, CLO18, CKCK10, Han11, KSB11, OMR14, PSB^{+}17, RRL16, Shi14]. UASB [ZLT^{+}14]. ubiquitous [ZW11]. UGR [RGTSLCH14]. UK [COS11b, Fan13a, HL17, HLSW18, Hud16, MBT16, MTT15, ZZY13]. Ullmann [Tom18]. ultimate [Oos15]. un-cited [ZG17c]. unavailable [Mou15a]. Unbalanced [ZHG16, HDC13]. unbiased [MD12]. uncertain [ZZW19a]. uncertainty [KHS^{+}15, dCCMAW16a, dCCMAW16b]. uncited [HLW19, YY14, YSY^{+}13]. uncitedness [Egg10b, NF19]. Uncovering [DMV10, Lee15, MY16]. Under-cited [RH18, DW18]. under-citedness [MHFB17]. Underestimating [LXWC17]. undergraduate [Yur18c]. underlying [CL13, HC16b, QL12]. undermining [Fra17]. understand [You14]. Understanding [Ano11, LTK^{+}18, LDVSGDR16, LDZ17, MGGdP17, PTMT11, XBD^{+}18, ZYS16, BM19, FK17, FK18, MPM18, OMR14, PCR18, THAL15]. undertake [BL11b]. undifferentiated [Kwi18]. undisclosed [Cop19b]. unequivocally [PL17]. unhistorical [Kra16]. uniform [ADS10b]. uninformed [RSGFV18]. unintended [SFM16]. union [ACORC10, MM16]. unions [MdBdP^{+}19]. uniqueness [yTwTlW19]. unit [dS17b]. United [SDS19, CC16, GK18, Lun19, PFL19, TBS15, TYWZ12, WMW^{+}13, YK15, Yur18b, Yur18c]. units [Lee10a, LWIB16, PRRC16]. universal [CRLMLM17, GGW11]. Universality [EHK12]. universally [SA16]. universe [KWS17, MB10a]. Universidad [HMCD^{+}19]. Universitie [LBRR19, AChO19, AC12, CFL12]. Universities [AKB^{+}10, BBSS16b, BBSS16c, OMAT19, Saf19, ADR14a, ADD14b, AhOL14, AChO19, AZSA14, AZSA16, AÇA^{+}14, APLHF18, AS18b, BVOL18, BPJ^{+}14, BR11, Ben15, BS17, BG18, BZBLP16, BDC^{+}12, CD14, CPF13, DCGZ^{+}12, DH13b, DQ11, Doc13, DC17, EGU10, FJ11, FYC15, FL16, FCCMTRVR18, FH13, GRSFV^{+}12b, GE11, GSMT10, GP18a, GMM16, GKS18, Hos11, HN16, HSWC13, JBMR11, KY16, KLL14, KHK13, KHA17, KB18, KKE13, LP12, Lee12, LYGQ12, LCWY12, LIdMAM11, LZL10, LXH^{+}18, LSE^{+}18, Mat12, Mat13, MT15, PS16a, PM18, PSB^{+}17, PQG14, Pra16b, Pra18d, RGTSLCH14, RGCM14, SY16a, Sha12, SM16b, TAB13, TW18, TR14, TBB^{+}16, WMT^{+}12, WMW^{+}13, WHC^{+}13, WWP14, WHL^{+}15, YZB18, ZCKZ16, ZHMX14, BBSS16a, APPF18]. University [CD14, Che15, GY12, Joh18, Kim14, MRGT18, Saf19, SSAG16, SFM16, TYYW16, ZL18a, ADS10a, ADD11c, ADS12, ADR14b, ACF^{+}17, ABILO10, ACP12, ALH15, ALvH19, Ban18, Bas11, BPGGdMA12, BR11, BR12, BGR19, BG17, BS13b, ÇAAÇ15, hCyL12, CHC13, CK14, Chu14, Cla15, CAV^{+}19, DH13a, Doc11a, Doc11b, Doc12, DC15a, DEC15, FYC15, FGP13, GGG14, GRSFV^{+}13, HB17b, HDW^{+}15, IL14b, KHA17, KS17, LVGV^{+}11, Laz10, LZZ^{+}12, LWIB16, LM10, LPL14, LCWY12, MR15, MOO17, Moe17, OVJM17, OMOR13a, OMOR13b, OMA15, PMJF19, Pie18, PS16b, PRA16a, Puu10, RVFEdlM10, RGTSLCH14, RGCM14, Saf13, SR16, SDS14a, SN10, TM12, Tod14, WHL^{+}15, WHLP16, WA18a, WS13a, ZW17a, vLvWW16, vRvLV11, GD11, GG13, GGG^{+}12, HFW^{+}14, OCJB15, RCCM14, SV19, WS10]. university-affiliated [CAV^{+}19]. University-industry [TYYW16, ZL18a, ADS12, FYC15, RVFEdlM10, SR16, TM12, WHLP16, WA18a]. university-industry-government [CHC13, Chu14, IL14b, LZZ^{+}12, LPL14]. university-invented [CD14]. University-owned [CD14]. university-research [GGG14]. University-sponsored [Che15]. Unravelling [ATCCAAB19]. unrelated [CSC12, CC12b, GRSFV14a]. Unseen [WW12]. Unsupervised [WLPH14]. unusable [Pen19]. unusual [The18d]. Unveiling [AMMT16, CKB^{+}14, Keg15]. unwarranted [Pol16a]. update [HM15c]. updated [APYS13, CCLL14, TB19a]. Updating [Rou12b]. upflow [ZLT^{+}14]. upon [CC12b]. ups [SRP13]. URAP [AÇA^{+}14]. URAP-TR [AÇA^{+}14]. urban [JDG14, LSR13]. urbanization [ZLL^{+}17]. URL [Lin11]. Urquhart [TÜ10]. USA [CGZ10, Dya17a, Dya17b, KKCG18, MT13b, TFJD14, The18d, WMW^{+}13, YXW18, ZYG15, ZW19]. usability [WAT16]. Usage [Che18a, GG15a, GGS14, MBL18, WMXZ14, WFS16, BM19, CG17, CG18a, CGG19, GKF17, HF19, KCT^{+}17, MA19, MF14, SG10, WLMF15a, WLMF15b, ZLTY18, Che20]. Use [AT17, AT18, EGU10, GWS15, MS16b, SMM^{+}19, SS14, AL12, BI18a, BW10, BWD12, BL18, CP12b, CH15, CP16, CF18, Das16, GBM^{+}16, GWG17, HC14a, Hun12, JKSK15, KD18, KPL19, KSB11, Kos16b, LS17a, LWB16, LRY18, MBTKA14, MM17a, MM17b, PYH16, PW17, PHDC16, Rod16, RNB19, SB15, Sch15c, TDG17, TÜ10, Web16, You14, dSD18b, dSD18c, dCdSNB15]. used [AD16, BD12b, BG17, Fuk17, Hal13, MCR^{+}12, SLISC17, The18a]. useful [Agu12, Hsi11, MBA13, QZZ17, Sch14a, The17a, dS18]. Usefulness [Tur16, BT19b, YK14, ZHG16]. user [Ell18, Jun12, VH17, YZB18, Yu17, ZC14]. users [PLJ18, Tri10, WZ19b]. Using [ATJ16, ANOdFC12, BWbH^{+}18, CWL10, CL11, Che11, CC12b, CyPP12, DAMC15, FSLR10, FA10, GT11, GT12, GT17, Ham14, ILB11, MK18, MOO17, RHMH17, RR17, RGGBV16, SCGZR16, Sch16, YLL^{+}15c, Zhe19, dSAEE15, BMM17, BASL16, BHL^{+}10, BYY18, BS16, BNV11, BB17b, CWJBT10, CBWJ18, Cam11, Cam17, Car16, CSO17, CZPR17, CC10a, CLHH10, CRZGVQMA15, CQB16, DGPL15, Doc11a, Doc11b, DGF17, EDEH16, Fan18, FMP17b, FH13, FH16, GGR11, GRSS16, GPN14, GY12, GSKM17, GBHT16, GHA^{+}16, GR14, GHvdB12, Hal14, HHZ14, HII^{+}18, HSAK18, HB17a, HB18a, HD17, HFL14, HWS18, HLW19, HC14b, hHC15, HCS^{+}15, HHA^{+}16, HC19, IA19, IMSK14, IB15, KMD^{+}18, KLP17, KWM^{+}18, KDFL14, KP12a, KLPP16, KP12b, KKK^{+}14, Kol12, KN15, Laz10, LKSK15, LFLG14, LJJ^{+}16, MY16, MSDJ19, MJHG13, MdFdA^{+}14, MT13a]. using [MKHB15a, MKHB15b, MYP19, dBONM^{+}19, OROMAA16, PYK12, PYK13, PY14, Par15, PLJ18, PP18, Pra10a, PMN16, Pra18e, RJ14, Rod16, RBC^{+}10, SGM^{+}16, SR16, Sal17, SKCK14, Sma11, SJOC18, SYLC17, TCC17, TT13, TE18, VHH16, kWhHRkS10, WJCC19, WHS19, WLPH14, XCS^{+}16, YCL^{+}13a, YK11, YK12, YPK13, YLJ^{+}17, YYLW14, Yur15, ZLZ19, ZZLS19, ZLLD19, ZHMX14, ZWL^{+}18, ZP16, dCPRP18, vEW17]. USPC [WSH16]. USPTO [LKR14, LZC17, WLY14]. USSR [WHW^{+}19]. utilisation [KM18a]. utility [CRBRG^{+}18, HV18a, Hei19, Shi14]. utilization [HDW^{+}15, YQX10].

vaccines [CSR^{+}18]. Validating [LTG12, RPNC13, YK14]. validation [EO14, LZB10, Rya16, TZG15, ZJLG10]. validity [BD10b, Bor18, BTL19, GP13, HL18, JOGC17, Nii17, TYYW16]. valuable [dSM17]. Value [FESD11, Bak17, BD12a, BK10, BZBLP16, CO10, CSC13, CSC14, CLW^{+}19, Cop18, FMS17, FSAB10, LS17a, LS16, LPL16, LD16, Mes11, MMSS11, NF13, PJY17, Pra12d, QDK19, Sch11a, SA16, SDS14b, TB19b, TSG13, THB18, YLL^{+}15c]. value-added [MMSS11]. valued [CFSSP16, Sch15c, Van10]. values [CL13, Tol12]. Vanadium [CRLMRPA17]. Vanclay [Jac12, PG12, Zit12]. Variability [ADD14b, PRRC15, vdBBdK16]. variables [ASPF^{+}16, Sch15c, VEJC^{+}18a]. variance [HAL11]. variant [BIL15]. variants [uARA19, AA19, LWT16, RAA18]. Variation [ADM14, AKB^{+}10, BF17, Fan18, SL12b]. variety [Ley15b, Ley18, TYWZ12]. various [HCL14, LCC12]. vary [WYB^{+}17]. varying [ILBG14]. Vector [CFSSP16, LSC10, SML16, VACCAJ18]. Vector-Space [LSC10]. Vector-valued [CFSSP16]. vegetable [TFJD14]. vehicle [HSLP14, YSY^{+}13, ZZ14]. velocity [GPN14]. Venezuela [CRV12]. venue [SBT18]. verification [Par15]. vernacular [SSdOS17]. versions [BI18a, BM12a, Moe16a, dSD18b, dSD18c]. versus [ADD10, ADD17b, ADR19, BD16a, BGJ^{+}16, BSBG18, BWD12, Cam11, Cam17, GF17, GG15a, GGS14, GWBSVWB13, Hsi11, HBDL18, JX13, KM15c, LO12, MBR^{+}13, Pir19, Ran09, SMLHCP17, SL17, SKM15, SVS18, TK17, WAT16, dWZD14]. vertically [Kwi18]. via [AUS12, ABM19, DLGP16, GSM^{+}16, Iwa17, KO19, LGD11, LTK^{+}18, NZL^{+}19, Xie15, ZZFD18a]. vibration [ZZ14]. vibration-reduction [ZZ14]. video [JNA18, XYHD18, ZXT^{+}19, Zon19]. Vienna [Ano15, GGG^{+}12]. Vietnam [Man15, NHLL17]. view [BBDS^{+}14, Har17, HTHB11, Lee10b, LGD11, MB10a, MB16a, MLT^{+}15, VACCAJ18, ZW17b, ZTC15]. viewed [Ard12, WB15]. Viewing [PPE14, Pin15]. viewpoint [YK14]. views [LLX^{+}18, dW15]. violating [WZ17]. virtual [HHA^{+}16]. virus [EDEH16]. Visibility [DCGZ^{+}12, Kim10, ADD16, AR18, ATM16, CRMdMA15, CP12c, Dor17, EMSH16, EGR13, GSOLHO19, HP18, LP12, LML11, MGM^{+}17, PROG19, PQG14, SLGO17, SLH18]. visible [LM11, Tei11, VVN16]. vision [ERW12, FSOS12]. Visual [LYW19, CZV10, NH11, PC18]. Visualization [Liu13, Sch12a, CL16, GY12, HZ17, KZC16, Kor19, KS17, KdBBK15, Ley15a, PJL19, SZD16, SB19, SYLC17, WKHS19, Xie15, YWC12, ZCMVQS11]. visualizations [BSBG18, DGF17]. Visualized [JNA18, LW15]. Visualizing [ACD14, BHH18, Fan15b, GN19, Li19, LYLD15, LM16, NSC13, WPW^{+}14, ZW11, ZH17, ALYZ15, BHL^{+}10, BH18b, KMP11a, Ley11a, PROGMA10]. vitae [BCT19]. vitality [WLC17]. viticulture [AATBPAB15a]. vocabularies [JYM^{+}16]. vol [Ho16]. volatile [ZXH10]. volume [WM17]. volumes [Sch14b]. voluntary [BDE11, SMLHCP17]. VOSviewer [vEW10, vEW17]. voting [JYM^{+}16, MOA16b]. vouchers [ACAGD^{+}17]. VQR [AD16]. vs [GSTD11, MBL18]. vu [GREL14].

wait [PR15]. wake [KNK^{+}19, SSZL18]. walking [LW10]. wall [GRSS16]. walls [Mik17]. Waltman [Pra12a, OL11]. War [HH17c, Bra12a, JJR10]. warfare [MM18, MM19]. warming [AM18, MHFB17]. warning [Har15b]. wars [IJF16]. Warsaw [KBL15]. waste [ACD13, CJY^{+}15, LFLG14, RJ14, YCL^{+}13a, YCL^{+}13b]. wastewater [ZWW^{+}15]. Watch [RV18a, RV18b]. Water [WYH10, CWH11, DSG^{+}15, HC16b, WXZ^{+}16, YYS^{+}10, ZLL^{+}15]. watershed [GG14]. way [KLL14, MCR^{+}12, Moe10, RGGBV16]. ways [Har13b, PRRC16, Sch13a]. Weak [PNS^{+}10, CO10, MS15b]. weak-links [MS15b]. weakest [Lin10]. weakness [Pra19e]. weaknesses [LKY17]. wealth [Hor18, Kor19, PL18, Pra17b]. weapons [ACMP13]. weather [AM18]. Web [Che20, Har13c, Mik10, AYS^{+}13, BIL15, BPJ^{+}14, CL16, GRSFVdMA14, GWS15, KBT14, LP12, LAS14, Lor14, MBTKA14, Moo15, OMR14, SMY15, SK14b, TCH^{+}15, YQX10, YS13, AT18, AZKR13, ÁBMB17, AL12, Ard12, BI10b, BI18b, BBDS^{+}14, BBJS16, Cam18, Che18a, CP12c, DBO^{+}18, DNAH15, Don17, DGGBDG17, Fra10, FGP13, GRSFV^{+}14c, GY12, Gau17, GK18, HA16, Har19b, HB17a, HHBB18, HBA19, Hol10, HSPY15, IWK18, Jac12, Jac18, JZL10, LCR13, LRY18, Liu19, MdBdP^{+}19, MMOMLC18a, MB15, MF14, ML10, MSB18, MMA18, MPH16, OMAT19, PHDC16, RAM18, RCETS19, VRF12, VBTK19, WFS16, Wil15, XHA^{+}19, ZCL14, ZHL19, ZLL19, dWZD14]. web-based [AYS^{+}13, CL16, Moo15]. web-citations [BPJ^{+}14]. Web-of-Science [LCR13]. webometric [AOFU10, Agu12, Fan15a, Hol10, The12]. Webometrics [KÖG12, KG13, CP14, KP12a, KLPP16, MOA16b, Lor14]. website [KBT15, MBA13]. websites [OMOR13b, PQG14]. Wechat [ZW17b]. week [BHDI18]. Wegener [ARE^{+}18]. weight [Pra19e]. Weighted [Abb11, Abb13, ACC^{+}16, CGKB18, EG16, FSSPG^{+}15, LGPC18, LSS15, LSS16, Yur18a, Zha17, Zyc10]. weighting [AW11, CTL^{+}19, MBSB17]. weights [ACP12, DQ11, NF13]. welfare [GMJ^{+}17, KKBW17]. well [BTL19, KHA17, OCM^{+}12, Tor13, ZCW14, ZHG16]. well-adjusted [OCM^{+}12]. well-being [Tor13]. West [Mêg13a, Mêg13c, Ole12]. Western [Bai18, Geo17, SDS19]. Whatever [SS18]. Where [GE11, KGB^{+}18, TBS15, YP19, ZG12a, ZW18a, AC13, Cop19a, Glä15, Koz15, Laa14, Tol11]. Which [ADR19, BG18, Chi14, GG19, MHTB17, MT15, TK17, WWC19, ADC12, Bjø19, CL13, DGDGSV15, GRSFVMB12, KM18c, WLF15, XGL^{+}19, CMO11]. white [JSZ13]. Who [CC16, Eld19, EC16, HH17b, JSZIZ13, Kwi18, Sak19, XG18, XYHD18, YHL^{+}18, Zhu17, AC12, BC13b, BL11b, FB10, GRSFV16a, San18, VH17]. whom [IDKF17]. wide [CCM^{+}11]. wide-scale [CCM^{+}11]. wider [JDH12]. widespread [Sch15a, Wu18]. Wikipedia [PB17b, SLISC17]. Wil [Ben11]. Will [CGPT15, APPS15, JC11, PG12, RRLNAG15]. willingness [PR15]. Wind [SLISC17, GYZ15, SCGZSL^{+}13]. window [Cam11, HC14b, SL13, Wan13]. windows [CLHH10, HC16c]. wine [ATCCAAB19]. wings [HA17b]. winners [Fie15b, GW10a, Har13d, HR17, MSYW12]. winning [CMT18, MTA^{+}18, SBB16]. wins [Ano12c, Ano14, Ano17a, KL16]. withdrawal [QA18]. Within [PRRC15, ADD14b, ACFL11, APT13, Ama16, BKG16, BL15, CGV12, CyPP12, DLGP16, HFW^{+}14, LC18, LMR16, MWH14, PYW18, Pra17b, RPGM10, Shi14, The19a, UMK14, ZRY^{+}12, ZB12]. Within- [PRRC15]. without [HLSW18, KKS^{+}17, Lyk18, WST14]. woke [HH17b]. Wolfgang [Pen19]. Women [FB10, PFDL17, DVB15, GK14, HC14a, HMCD^{+}19, HAJ12, IMHG12, IMH13, KPS12, KK13, NH14, yTnL17, VCC12, ZCMVQS11]. won [MSYW12]. word [AW11, CDM18, CLSW19, Dan14, DVB14, FZZ17, GW15a, HHDL13, KD18, KO19, LSS15, LSS16, LHCH18, LHW12, OGOPPR17, RAS15, RFGBMA13, WZS12, WLLL12, WCL14, WZW15, Xie15, YLL15b, YWC12, ZLL^{+}17, ZSY^{+}13]. Word2Vec [HWQ^{+}18]. words [Zit15]. Work [MHC^{+}15, WKK16, ARK^{+}15, BSMD11, CSO17, Ho14, Lyk18, MAA18, NCG^{+}19, Par15, RH18, SBB16, SJOC18, Tor13]. Workbench [BHL^{+}10]. workflows [GWP16, MFF^{+}16]. working [BM14b, PCR18, PRRC16, Tod14]. works [Ho13b, KMS16, MM14b, MHTB17, ZZFD18b]. workshop [Ano11, Ano15]. World [GAGT15, Saf19, SFM16, ADV11, AZSA14, AZSA16, ACRC17, BMM14, BPJ^{+}14, BPGGdMA12, hCyL12, Doc13, GZ18, GB14b, GS12, HHZ14, HYYL12, HLY14, HW10, JBMR11, Kor19, LP12, LJJ^{+}16, Mat12, Moe17, Osw10, Pie18, RY14, SC18, SUP15, SH15c, TYYW16, Veu10, ZXH10, ZGL14, ZWHH13, HH17c, PP11, SH15b]. world-class [LP12]. world-leading [Osw10]. Worldwide [WW15, XZZ15, Hal13, LVSL18]. worse [HS16a]. Worsening [Pau10]. worst [SR16]. WoS [Che18a, Che20, GK18, BI10a, FMM16, MÁB18, SOBM16]. Wray [Pet18b]. Writing [AS18a, HC15a, HC16a, JKN19, Lei16, MC15, OK13]. written [BHJD12, Har16b, MSYW12, ZXLX14]. wrong [Cha17b]. wronged [vRvLV11].

X [Sch15c, Saf13]. X-centage [Sch15c].

Year [BT18c, Egg14b, MR13, MHTB17, BHL18, Cam11, CH15, DMB17, HA17a, Hou17, KKT^{+}18, LGH^{+}14, ML18, MB14, MDDG17, PKR15, PFL19, TBW^{+}12, vdBBdK16, BT18b]. Year-based [Egg14b, MR13]. yearly [Fan18]. Years [WB15, AER^{+}14, AAG14, AP16, Bar11, CAS16, CG18b, DWGL16, EBR16, HLSC18, JDG14, KGB^{+}18, MGT14, OHT10, PHBN^{+}15, RBBG18, SMF18, Sni16, SRP13, The18c, Tod14, VACCAJ18, WLC17, WOW13, WS13b, YWG14, ZLG^{+}15, dCdSNB15]. Yes [MRLW15, Mou15b]. Yesterday [GN17]. young [CSS^{+}16, HL15, KY16, Kos15, QZL^{+}17]. younger [vRW18]. Yugoslav [IJF16]. Yugoslavia [IJF16, JJR10].

z [Ano18b]. Zealand [HH10]. zealots [CPRSFVG19]. Zentralblatt [CB15]. Zero [NF19]. zeros [HB18a]. Zipf [Aus14b, ANFF16, ER19b, LAL15]. ZnO [ÁRM13]. zone [ABM19]. Zoonoses [VSS12]. zoonotic [HKWC15].

References

Assimakis:2010:NAP

[AA10] N. Assimakis and M. Adam. A new author’s productivity index: p-index. Scientometrics, 85(2):415–427, November 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0255-z.

Ayaz:2016:ICF

[AA16] Samreen Ayaz and Muhammad Tanvir Afzal. Identification of conversion factor for completing-h index for the field of mathematics. Scientometrics, 109(3):1511–1524, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2122-z.

Ahmad:2018:CAC

[AA18] Riaz Ahmad and Muhammad Tanvir Afzal. CAD: an algorithm for citation-anchors detection in research papers. Scientometrics, 117(3):1405–1423, December 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2920-6.

Ameer:2019:EIQ

[AA19] Madiha Ameer and Muhammad Tanvir Afzal. Evaluation of h-index and its qualitative and quantitative variants in neuroscience. Scientometrics, 121(2):653–673, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03209-6.

Akcan:2013:MQC

[AAB^{+}13] Derya Akcan, Susanna Axelsson, Christina Bergh, Thomas Davidson, and Måns Rosén. Methodological quality in clinical trials and bibliometric indicators: no evidence of correlations. Scientometrics, 96(1):297–303, July 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0949-0.

Alhaider:2014:GRO

[AAG14] Ibrahim Alhaider, Mueen K. K. Ahmed, and B. M. Gupta. Global research output on date palm (Pheonix dactylifera): a 12 years scientometric perspective. Scientometrics, 98(1): 157–171, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-0991-y.

Abbasi:2010:ESB

[AAH10] Alireza Abbasi, Jörn Altmann, and Junseok Hwang. Evaluating scholars based on their academic collaboration activities: two indices, the RC-index and the CC-index, for quantifying collaboration activities of researchers and scientific communities. Scientometrics, 83(1):1–13, April 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0139-2.

Abbas:2019:IAA

[AAS^{+}19] Naveed Naeem Abbas, Tanveer Ahmed, Syed Habib Ullah Shah, Muhammad Omar, and Han Woo Park. Investigating the applications of artificial intelligence in cyber security. Scientometrics, 121(2):1189–1211, November 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03222-9.

Aleixandre:2015:GTS

[AATBPAB15a] Jose Luis Aleixandre, Jose Luis Aleixandre-Tudó, Máxima Bolaños-Pizarro, and Rafael Aleixandre-Benavent. Global trends in scientific production in enology and viticulture in selected emerging economies (BRIC). Scientometrics, 103 (2):649–668, May 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1543-4.

Aleixandre:2015:MSR

[AATBPAB15b] José Luis Aleixandre, José Luis Aleixandre-Tudó, Máxima Bolaños-Pizarro, and Rafael Aleixandre-Benavent. Mapping the scientific research in organic farming: a bibliometric review. Scientometrics, 105(1):295–309, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1677-4.

Arts:2013:IST

[AAV13] Sam Arts, Francesco Paolo Appio, and Bart Van Looy. Inventions shaping technological trajectories: do existing patent indicators provide a comprehensive picture? Scientometrics, 97(2):397–419, November 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1045-1.

Abbas:2011:WIE

[Abb11] Ash Mohammad Abbas. Weighted indices for evaluating the quality of research with multiple authorship. Scientometrics, 88(1):107–131, July 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
011-0389-7.

Abbasi:2013:THC

[Abb13] Alireza Abbasi. h-type hybrid centrality measures for weighted networks. Scientometrics, 96(2):633–640, August 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-0959-y.

Alvarez-Bornstein:2019:WCF

[ÁBDFB19] Belén Álvarez-Bornstein, Adrián A. Díaz-Faes, and María Bordons. What characterises funded biomedical research? Evidence from a basic and a clinical domain. Scientometrics, 119(2):805–825, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03066-3.

Alvarez-Betancourt:2014:OIR

[ABGS14] Yuniol Alvarez-Betancourt and Miguel Garcia-Silvente. An overview of iris recognition: a bibliometric analysis of the period 2000–2012. Scientometrics, 101(3):2003–2033, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1336-1.

Aguillo:2010:CUR

[ABILO10] Isidro F. Aguillo, Judit Bar-Ilan, Mark Levene, and José Luis Ortega. Comparing university rankings. Scientometrics, 85(1):243–256, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
010-0190-z.

Andersen:2017:FSR

[ABL17] Martin S. Andersen, Jeremy W. Bray, and Albert N. Link. On the failure of scientific research: an analysis of SBIR projects funded by the U.S. National Institutes of Health. Scientometrics, 112(1):431–442, July 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Asadi:2019:AZI

[ABM19] Nasrin Asadi, Kambiz Badie, and Maryam Tayefeh Mahmoudi. Automatic zone identification in scientific papers via fusion techniques. Scientometrics, 119(2):845–862, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03060-9.

Alvarez-Bornstein:2017:FAW

[ÁBMB17] Belén Álvarez-Bornstein, Fernanda Morillo, and María Bordons. Funding acknowledgments in the Web of Science: completeness and accuracy of collected data. Scientometrics, 112(3):1793–1812, September 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2453-4.

Aleixandre-Benavent:2014:DSI

[ABMRVZ14] Rafael Aleixandre-Benavent, Vicent Montalt-Resurecció, and Juan Carlos Valderrama-Zurián. A descriptive study of inaccuracy in article titles on bibliometrics published in biomedical journals. Scientometrics, 101(1):781–791, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1296-5.

Aleixandre-Benavent:2016:CBI

[ABMSSP16] Rafael Aleixandre-Benavent, Luz María Moreno-Solano, Antonia Ferrer Sapena, and Enrique Alfonso Sánchez Pérez. Correlation between impact factor and public availability of published research data in information science and library science journals. Scientometrics, 107(1):1–13, April 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-016-1868-7.

Aleixandre-Benavent:2015:ITB

[ABRVZ15] Rafael Aleixandre-Benavent, Vicent Montalt Resurrecció, and Juan Carlos Valderrama-Zurián. Inaccuracies in titles on bibliometrics in biomedical journals. Scientometrics, 103(1):331–332, April 2015. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-014-1504-3.

Aleixandre-Benavent:2019:PRP

[ABSF^{+}19] Rafael Aleixandre-Benavent, Antonia Ferrer Sapena, Silvia Coronado Ferrer, Fernanda Peset, and Alicia García García. Policies regarding public availability of published research data in pediatrics journals. Scientometrics, 118(2): 439–451, February 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2978-1.

Abt:2012:PII

[Abt12] Helmut A. Abt. A publication index that is independent of age. Scientometrics, 91(3):863–868, June 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0525-4. See comments [Sch12b].

Abt:2017:CAN

[Abt17] Helmut A. Abt. Citations and author numbers in six sciences. Scientometrics, 111(3):1861–1867, June 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Alvarez:2014:QQA

[ÁBV^{+}14] Pablo Álvarez, Houria Boulaiz, Celia Vélez, Fernando Rodríguez-Serrano, Raul Ortiz, Consolación Melguizo, Esmeralda Carrillo, Antonio Martínez-Amat, and Jose Prados. Qualitative and quantitative analyses of anatomists’ research: evaluation of multidisciplinarity and trends in scientific production. Scientometrics, 98(1):447–456, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1006-8.

Azagra-Caro:2012:AUP

[AC12] Joaquín M. Azagra-Caro. Access to universities’ public knowledge: who’s more nationalist? Scientometrics, 91 (3):671–691, June 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0629-5.

Anania:2013:TSN

[AC13] Giovanni Anania and Annarosa Caruso. Two simple new bibliometric indexes to better evaluate research in disciplines where publications typically receive less citations. Scientometrics, 96(2):617–631, August 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0951-6.

Alasehir:2014:UTN

[AÇA^{+}14] Oguzhan Alasehir, Murat Perit Çakir, Cengiz Acartürk, Nazife Baykal, and Ural Akbulut. URAP-TR: a national ranking for Turkish universities based on academic performance. Scientometrics, 101(1):159–178, October 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1333-4.

Arbelaez-Cortes:2017:KLM

[ACAGD^{+}17] Enrique Arbeláez-Cortés, Andrés R. Acosta-Galvis, Carlos DoNascimiento, Diana Espitia-Reina, Arturo González-Alvarado, and Claudia A. Medina. Knowledge linked to museum specimen vouchers: measuring scientific production from a major biological collection in Colombia.

Scientometrics, 112(3):1323–1341, September 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2461-4.

Anfossi:2016:LSA

[ACC^{+}16] Alberto Anfossi, Alberto Ciolfi, Filippo Costa, Giorgio Parisi, and Sergio Benedetto. Large-scale assessment of research outputs through a weighted combination of bibliometric indicators. Scientometrics, 107(2):671–683, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1882-9.

Alvarez:2015:AAP

[ÁCCG^{+}15] R. Álvarez, E. Cahué, J. Clemente-Gallardo, A. Ferrer, D. Íñiguez, X. Mellado, A. Rivero, G. Ruiz, F. Sanz, E. Serrano, A. Tarancón, and Y. Vergara. Analysis of academic productivity based on complex networks. Scientometrics, 104(3):651–672, September 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1627-1.

Abramo:2011:DPB

[ACD11] Giovanni Abramo, Tindaro Cicero, and Ciriaco Andrea D’Angelo. The dangers of performance-based research funding in non-competitive higher education systems. Scientometrics, 87(3):641–654, June 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
011-0355-4.

Abramo:2013:NPR

[ACD13] Giovanni Abramo, Tindaro Cicero, and Ciriaco Andrea D’Angelo. National peer-review research assessment exercises for the hard sciences can be a complete waste of money: the Italian case. Scientometrics, 95(1):311–324, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0875-6.

Appio:2014:VSB

[ACD14] Francesco Paolo Appio, Fabrizio Cesaroni, and Alberto Di Minin. Visualizing the structure and bridges of the intellectual property management and strategy literature: a document co-citation analysis. Scientometrics, 101(1):623–661, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1329-0.

Abramo:2015:MSM

[ACD15] Giovanni Abramo, Corrado Costa, and Ciriaco Andrea D’Angelo. A multivariate stochastic model to assess research performance. Scientometrics, 102(2):1755–1772, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1474-5.

Acosta:2017:GUS

[ACF^{+}17] Manuel Acosta, Daniel Coronado, Esther Ferrándiz, M. Dolores León, and Pedro J. Moreno. The geography of university scientific production in Europe: an exploration in the field of Food Science and Technology. Scientometrics, 112 (1):215–240, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Acosta:2011:FAI

[ACFL11] Manuel Acosta, Daniel Coronado, Esther Ferrándiz, and M. Dolores León. Factors affecting inter-regional academic scientific collaboration within Europe: the role of economic distance. Scientometrics, 87(1):63–74, April 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0305-6.

Ahn:2019:LBU

[AChO19] JongWuk Ahn, Hyundo Choi, and Dong hyun Oh. Leveraging bridging universities to access international knowledge: Korean universities’ R&D internationalization. Scientometrics, 120(2):519–537, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03138-4.

Alonso:2010:INI

[ACHVH10] S. Alonso, F. J. Cabrerizo, E. Herrera-Viedma, and F. Herrera. hg-index: a new index to characterize the scientific output of researchers based on the h- and gindices. Scientometrics, 82(2):391–400, February 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0047-5.

Acosta:2013:FAD

[ACMP13] Manuel Acosta, Daniel Coronado, Rosario Marín, and Pedro Prats. Factors affecting the diffusion of patented military technology in the field of weapons and ammunition. Scientometrics, 94(1):1–22, January 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0857-8.

Albarran:2010:CSP

[ACORC10] Pedro Albarrán, Juan A. Crespo, Ignacio Ortuño, and Javier Ruiz-Castillo. A comparison of the scientific performance of the U.S. and the European union at the turn of the 21st century. Scientometrics, 85(1):329–344, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0223-7.

Albarran:2011:SSS

[ACORC11] Pedro Albarrán, Juan A. Crespo, Ignacio Ortuño, and Javier Ruiz-Castillo. The skewness of science in 219 subfields and a number of aggregates. Scientometrics, 88(2): 385–397, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0407-9.

Ahlgren:2012:FNC

[ACP12] Per Ahlgren, Cristian Colliander, and Olle Persson. Field normalized citation rates, field normalized journal impact and Norwegian weights for allocation of university research funds. Scientometrics, 92(3):767–780, September 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0632-x.

Albarran:2017:GMR

[ACRC17] Pedro Albarrán, Raquel Carrasco, and Javier Ruiz-Castillo. Geographic mobility and research productivity in a selection of top world economics departments. Scientometrics, 111(1):241–265, April 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-017-2245-x.

Akbaritabar:2018:CRP

[ACS18] Aliakbar Akbaritabar, Niccolò Casnici, and Flaminio Squazzoni. The conundrum of research productivity: a study on sociologists in Italy. Scientometrics, 114(3):859–882, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2606-5.

Azagra-Caro:2018:ETA

[ACT18] Joaquín M. Azagra-Caro and Elena M. Tur. Examiner trust in applicants to the European Patent Office: country specificities. Scientometrics, 117(3):1319–1348, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2894-4;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2894-
4.pdf.

Abramo:2011:ERI

[AD11a] Giovanni Abramo and Ciriaco Andrea D’Angelo. Evaluating research: from informed peer review to bibliometrics. Scientometrics, 87(3):499–514, June 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0352-7.

Abramo:2011:NSR

[AD11b] Giovanni Abramo and Ciriaco Andrea D’Angelo. Nationalscale research performance assessment at the individual level. Scientometrics, 86(2):347–364, February 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0297-2.

Abatemarco:2013:CEC

[AD13] Antonio Abatemarco and Roberto Dell’Anno. Certainty equivalent citation: generalized classes of citation indexes. Scientometrics, 94(1):263–271, January 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0758-x.

Abramo:2014:HDY

[AD14] Giovanni Abramo and Ciriaco Andrea D’Angelo. How do you define and measure research productivity? Scientometrics, 101(2):1129–1144, November 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1269-8.

Abramo:2016:RAC

[AD16] Giovanni Abramo and Ciriaco Andrea D’Angelo. Refrain from adopting the combination of citation and journal metrics to grade publications, as used in the Italian national research assessment exercise (VQR 2011–2014). Scientometrics, 109(3):2053–2065, December 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2153-5.

Aboites:2018:IMM

[AD18] Jaime Aboites and Claudia Díaz. Inventors’ mobility in Mexico in the context of globalization. Scientometrics, 115 (3):1443–1461, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2645-6.

Anastasiadis:2010:TED

[AdAdAM10] Aristoklis D. Anastasiadis, Marcelo P. de Albuquerque, Marcio P. de Albuquerque, and Diogo B. Mussi. Tsallis qexponential describes the distribution of scientific citations — a new characterization of the impact. Scientometrics, 83 (1):205–218, April 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0023-0.

Abramo:2012:WAL

[ADC12] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Tindaro Cicero. What is the appropriate length of the publication period over which to assess research performance? Scientometrics, 93(3):1005–1017, December 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0714-9.

Abramo:2010:CVJ

[ADD10] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. Citations versus journal impact factor as proxy of quality: could the latter ever be preferable? Scientometrics, 84(3):821–833, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0200-1.

Abramo:2011:NRAb

[ADD11a] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. National research assessment exercises: a comparison of peer review and bibliometrics rankings. Scientometrics, 89(3):929–941, December 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0459-x.

Abramo:2011:NRAa

[ADD11b] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. National research assessment exercises: the effects of changing the rules of the game during the game. Scientometrics, 88(1):229–238, July 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0373-2.

Abramo:2011:NSC

[ADD11c] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. A national-scale cross-time analysis of university research performance. Scientometrics, 87(2):399–413, May 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0319-0.

Abramo:2011:RPH

[ADD11d] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. Research productivity: Are higher academic ranks more productive than lower ones? Scientometrics, 88(3): 915–928, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0426-6.

Abramo:2014:ISP

[ADD14a] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. Inefficiency in selecting products for submission to national research assessment exercises. Scientometrics, 98 (3):2069–2086, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1177-3.

Abramo:2014:VRP

[ADD14b] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. Variability of research performance across disciplines within universities in non-competitive higher education systems. Scientometrics, 98(2):777–795, February 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1088-3.

Amjad:2015:TBH

[ADD^{+}15] Tehmina Amjad, Ying Ding, Ali Daud, Jian Xu, and Vincent Malic. Topic-based heterogeneous rank. Scientometrics, 104(1):313–334, July 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
015-1601-y.

Abramo:2016:ECN

[ADD16] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. The effect of a country’s name in the title of a publication on its visibility and citability. Scientometrics, 109(3):1895–1909, December 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2120-1.

Abramo:2017:DIR

[ADD17a] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. Do interdisciplinary research teams deliver higher gains to science? Scientometrics, 111(1):317–336, April 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-017-2253-x.

Abramo:2017:SVD

[ADD17b] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. Specialization versus diversification in research activities: the extent, intensity and relatedness of field diversification by individual scientists. Scientometrics, 112(3):1403– 1418, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2426-7.

Abramo:2018:EMC

[ADD18a] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. The effect of multidisciplinary collaborations on research diversification. Scientometrics, 116(1):423–433, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2746-2.

Abramo:2018:EGA

[ADD18b] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. The effects of gender, age and academic rank on research diversification. Scientometrics, 114(2):373–387, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2529-1.

Abramo:2019:CBT

[ADD19a] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. The collaboration behavior of top scientists. Scientometrics, 118(1):215–232, January 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2970-9.

Abramo:2019:GAT

[ADD19b] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Flavia Di Costa. A gender analysis of top scientists’ collaboration behavior: evidence from Italy. Scientometrics, 120(2): 405–418, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03136-6.

Abramo:2014:VRC

[ADM14] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Gianluca Murgia. Variation in research collaboration patterns across academic ranks. Scientometrics, 98(3):2275–2294, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1185-3.

Adhikari:2019:GRK

[ADM19] Agniv Adhikari, Paramita Das, and Abhik Mukherjee. Generating a representative keyword subset pertaining to an academic conference series. Scientometrics, 119(2):749–770, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03068-1.

Abramo:2013:MIR

[ADR13] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Francesco Rosati. Measuring institutional research productivity for the life sciences: the importance of accounting for the order of authors in the byline. Scientometrics, 97(3):779–795, December 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1013-9.

Abramo:2014:CAS

[ADR14a] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Francesco Rosati. Career advancement and scientific performance in universities. Scientometrics, 98(2):891–907, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-1075-8.

Abramo:2014:RSU

[ADR14b] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Francesco Rosati. Relatives in the same university faculty: nepotism or merit? Scientometrics, 101(1):737–749, October 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1273-z.

Abramo:2016:GBA

[ADR16a] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Francesco Rosati. Gender bias in academic recruitment. Scientometrics, 106(1):119–141, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1783-3.

Abramo:2016:NSD

[ADR16b] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Francesco Rosati. The north–south divide in the Italian higher education system. Scientometrics, 109(3):2093–2117, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2141-9.

Abramo:2019:PRV

[ADR19] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Emanuela Reale. Peer review versus bibliometrics: Which method better predicts the scholarly impact of publications? Scientometrics, 121(1):537–554, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03184-y.

Abramo:2010:APP

[ADS10a] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Marco Solazzi. Assessing public-private research collaboration: is it possible to compare university performance? Scientometrics, 84(1):173–197, July 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
009-0104-0.

Abramo:2010:NRA

[ADS10b] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Marco Solazzi. National research assessment exercises: a measure of the distortion of performance rankings when labor input is treated as uniform. Scientometrics, 84(3):605–619, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0164-1.

Abramo:2011:RBS

[ADS11] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Marco Solazzi. The relationship between scientists’ research performance and the degree of internationalization of their research. Scientometrics, 86(3):629–643, March 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0284-7.

Abramo:2012:BTA

[ADS12] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Marco Solazzi. A bibliometric tool to assess the regional dimension of university-industry research collaborations. Scientometrics, 91(3):955–975, June 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0577-5.

Abramo:2016:DCD

[ADS16] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Anastasiia Soldatenkova. The dispersion of the citation distribution of top scientists’ publications. Scientometrics, 109(3):1711–1724, December 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2143-7.

Abramo:2017:EHL

[ADS17a] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Anastasiia Soldatenkova. Erratum to: How long do top scientists maintain their stardom? An analysis by region, gender and discipline: evidence from Italy. Scientometrics, 113 (3):1825, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2255-8;

http://link.springer.com/content/pdf/10.1007/s11192-
017-2255-8.pdf. See [ADS17a].

Abramo:2017:HLD

[ADS17b] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Anastasiia Soldatenkova. How long do top scientists maintain their stardom? An analysis by region, gender and discipline: evidence from Italy. Scientometrics, 110(2):867–877, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2193-x. See

erratum [ADS17a].

Abramo:2010:PRR

[ADV10] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Fulvio Viel. Peer review research assessment: a sensitivity analysis of performance rankings to the share of research product evaluated. Scientometrics, 85(3):705–720, December 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0238-0.

Abramo:2011:FSA

[ADV11] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Fulvio Viel. The field-standardized average impact of national research systems compared to world average: the case of Italy. Scientometrics, 88(2):599–615, August 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0406-x.

Abramo:2013:SIM

[ADV13] Giovanni Abramo, Ciriaco Andrea D’Angelo, and Fulvio Viel. The suitability of h and g indexes for measuring the research performance of institutions. Scientometrics, 97(3): 555–570, December 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1026-4.

Akhavan:2016:MTK

[AEFP16] Peyman Akhavan, Nader Ale Ebrahim, Mahdieh A. Fetrati, and Amir Pezeshkan. Major trends in knowledge management research: a bibliometric study. Scientometrics, 107 (3):1249–1264, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1938-x.

Abrizah:2014:SFY

[AER^{+}14] A. Abrizah, Mohammadamin Erfanmanesh, Vala Ali Rohani, Mike Thelwall, Jonathan M. Levitt, and Fereshteh Didegah. Sixty-four years of informetrics research: productivity, impact and collaboration. Scientometrics, 101(1): 569–585, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1390-8.

Abbasi:2015:CBP

[AF15a] Muhammad Kamran Abbasi and Ingo Frommholz. Clusterbased polyrepresentation as science modelling approach for information retrieval. Scientometrics, 102(3):2301–2322, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-014-1478-1. See erratum

[AF15b].

Abbasi:2015:ECB

[AF15b] Muhammad Kamran Abbasi and Ingo Frommholz. Erratum to: Cluster-based polyrepresentation as science modelling approach for information retrieval. Scientometrics, 103(3):1151–1152, June 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-015-

1590-x.pdf. See [AF15a].

Araujo:2018:SMI

[AF18] Tanya Araújo and Elsa Fontainha. Are scientific memes inherited differently from gendered authorship? Scientometrics, 117(2):953–972, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2903-7.

Ahmed:2013:ICA

[AG13] K. K. Mueen Ahmed and B. M. Gupta. India’s contribution on antioxidants: a bibliometric analysis, 2001– 10. Scientometrics, 94(2):741–754, February 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0700-2.

Adams:2014:ICC

[AGHL14] Jonathan Adams, Karen Gurney, Daniel Hook, and Loet Leydesdorff. International collaboration clusters in Africa. Scientometrics, 98(1):547–556, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1060-2.

Antonio-Garcia:2014:DSB

[AGLNRR14] M. Teresa Antonio-García, Irene López-Navarro, and Jesús Rey-Rocha. Determinants of success for biomedical researchers: a perception-based study in a health science research environment. Scientometrics, 101(3):1747–1779, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1376-6.

Aguillo:2012:GSU

[Agu12] Isidro F. Aguillo. Is Google Scholar useful for bibliometrics? A webometric analysis. Scientometrics, 91(2):343– 351, May 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0582-8.

Andersen:2011:PRG

[AH11] Jens Peter Andersen and Björn Hammarfelt. Price revisited: on the growth of dissertations in eight research fields. Scientometrics, 88(2):371–383, August 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0408-8.

Ahn:2014:SIP

[AhOL14] Jongwuk Ahn, Dong hyun Oh, and Jeong-Dong Lee. The scientific impact and partner selection in collaborative research at Korean universities. Scientometrics, 100(1):173– 188, July 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1201-7.

An:2017:IDK

[AHP17] Yoonjung An, Mintak Han, and Yongtae Park. Identifying dynamic knowledge flow patterns of business method patents with a hidden Markov model. Scientometrics, 113(2):783–802, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
017-2514-8.

Ahrweiler:2017:ABS

[Ahr17] Petra Ahrweiler. Agent-based simulation for science, technology, and innovation policy. Scientometrics, 110(1): 391–415, January 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-016-
2105-0.

Abbasi:2011:EDS

[AHUR11] Alireza Abbasi, Liaquat Hossain, Shahadat Uddin, and Kim J. R. Rasmussen. Evolutionary dynamics of scientific collaboration networks: multi-levels and cross-time analysis. Scientometrics, 89(2):687–710, November 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0463-1.

Alam:2017:RRS

[AI17] Mohammad Mahbub Alam and Maizatul Akmar Ismail. RTRS: a recommender system for academic researchers. Scientometrics, 113(3):1325–1348, December 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2548-y.

Arencibia-Jorge:2016:SOE

[AJCACRdMA16] Ricardo Arencibia-Jorge, Elena Corera-Alvarez, Zaida Chinchilla-Rodríguez, and Félix de Moya-Anegón. Scientific output of the emerging Cuban biopharmaceutical industry: a scientometric approach. Scientometrics, 108(3):1621–1636, September 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2023-1.

Arencibia-Jorge:2010:CSC

[AJdMA10] Ricardo Arencibia-Jorge and Felix de Moya-Anegón. Challenges in the study of Cuban scientific output. Scientometrics, 83(3):723–737, June 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
009-0150-7.

Abbasi:2018:INB

[AJSN18] Alireza Abbasi, Mahdi Jalili, and Abolghasem Sadeghi-Niaraki. Influence of network-based structural and power diversity on research performance. Scientometrics, 117(1): 579–590, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2879-3.

Aminpour:2010:IMU

[AKB^{+}10] Farzaneh Aminpour, Payam Kabiri, Mohammad Ali Boroumand, Abbas Ali Keshtkar, and Seyed Shamsoddin Hejazi. Iranian medical universities in SCIE: evaluation of address variation. Scientometrics, 85(1):53–63, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0248-y.

Akritidis:2012:IAR

[AKB12] Leonidas Akritidis, Dimitrios Katsaros, and Panayiotis Bozanis. Identifying attractive research fields for new scientists. Scientometrics, 91(3):869–894, June 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0646-4.

Amara:2012:CCF

[AL12] Nabil Amara and Réjean Landry. Counting citations in the field of business and management: why use Google Scholar rather than the Web of Science. Scientometrics, 93(3): 553–581, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0729-2.

Amara:2015:WCU

[ALH15] Nabil Amara, Réjean Landry, and Norrin Halilem. What can university administrators do to increase the publication and citation scores of their faculty members? Scientometrics, 103(2):489–530, May 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
015-1537-2.

Audretsch:2019:KBK

[ALvH19] David B. Audretsch, Albert N. Link, and Martijn van Hasselt. Knowledge begets knowledge: university knowledge spillovers and the output of scientific papers from U.S. Small Business Innovation Research (SBIR) projects. Scientometrics, 121(3):1367–1383, December 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03260-3.

An:2015:MVC

[ALYZ15] Lu An, Xia Lin, Chuanming Yu, and Xinwen Zhang. Measuring and visualizing the contributions of Chinese and American LIS research institutions to emerging themes and salient themes. Scientometrics, 105(3):1605–1634, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1640-4.

Allen:2018:FNI

[AM18] David E. Allen and Michael McAleer. Fake news and indifference to scientific fact: President Trump’s confused tweets on global warming, climate change and weather. Scientometrics, 117(1):625–629, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2847-y.

Amancio:2015:CTP

[Ama15] Diego Raphael Amancio. Comparing the topological properties of real and artificially generated scientific manuscripts. Scientometrics, 105(3):1763–1779, December 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1637-z.

Aman:2016:HCI

[Ama16]  Valeria Aman.  How collaboration impacts citation

flows within the German science system. Scientometrics, 109(3):2195–2216, December 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2092-1.

Aman:2018:DSA

[Ama18a] Valeria Aman. Does the Scopus author ID suffice to track scientific international mobility? A case study based on Leibniz laureates. Scientometrics, 117(2):705–720, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2895-3.

Aman:2018:NBA

[Ama18b] Valeria Aman. A new bibliometric approach to measure knowledge transfer of internationally mobile scientists. Scientometrics, 117(1):227–247, October 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2864-x.

Alejo-Machado:2015:BSS

[AMFLH15] Oscar J. Alejo-Machado, Juan Manuel Fernández-Luna, and Juan F. Huete. Bibliometric study of the scientific research on “Learning to Rank” between 2000 and 2013. Scientometrics, 102(2):1669–1686, February 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1467-4.

Ayaz:2018:PSI

[AMI18] Samreen Ayaz, Nayyer Masood, and Muhammad Arshad Islam. Predicting scientific impact based on hindex. Scientometrics, 114(3):993–1010, March 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2618-1.

Asgary:2013:HLE

[AMK13] Saeed Asgary, Leili Mehrdad, and Sanam Kheirieh. Highlevel evidences in endodontics. Scientometrics, 94(3): 955–962, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0742-5.

Archambault:2017:ERG

[AML17] Antoine Archambault, Philippe Mongeon, and Vincent Larivière. On the effects of the reunification on German researchers’ publication patterns. Scientometrics, 111(1): 337–347, April 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-017-
2254-9.

Appio:2016:UIO

[AMMT16] Francesco Paolo Appio, Antonella Martini, Silvia Massa, and Stefania Testa. Unveiling the intellectual origins of social media-based innovation: insights from a bibliometric approach. Scientometrics, 108(1):355–388, July 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1955-9.

Aspura:2018:AMR

[ANA18] M. K. Yanti Idaya Aspura, A. Noorhidawati, and A. Abrizah. An analysis of Malaysian retracted papers: Misconduct or mistakes? Scientometrics, 115(3):1315–1328, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2720-z.

Ausloos:2019:CBS

[AND19] Marcel Ausloos, Olgica Nedic, and Aleksandar Dekanski. Correlations between submission and acceptance of papers in peer review journals. Scientometrics, 119(1):279–302, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03026-x;   http://
link.springer.com/content/pdf/10.1007/s11192-019-
03026-x.pdf.

Ausloos:2016:QQP

[ANFF16] Marcel Ausloos, Olgica Nedic, Agata Fronczak, and Piotr Fronczak. Quantifying the quality of peer reviewers through Zipf’s law. Scientometrics, 106(1):347–368, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1704-5.

Anonymous:2010:SPI

[Ano10] Anonymous. Selected papers of the 10th International Conference on Science and Technology Indicators. Scientometrics, 82(3):457–458, March 2010. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-010-0170-3.

Anonymous:2011:SCR

[Ano11] Anonymous. Selected contributions, resulting from presentations at the workshop “Modeling Science–Understanding, Forecasting and Communicating the Science System,” held in Amsterdam, October 6–9, 2009. Scientometrics, 89 (1):345, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-011-0461-
3.pdf.

Anonymous:2012:CDP

[Ano12a] Anonymous. Comments on the discussion paper. Scientometrics, 92(2):239, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-012-
0815-5.pdf.

Anonymous:2012:DP

[Ano12b] Anonymous. Discussion paper. Scientometrics, 92(2):209, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-012-0814-6.pdf.

Anonymous:2012:OPW

[Ano12c] Anonymous. Olle Persson wins the 2011 Derek John de Solla Price Medal. Scientometrics, 90(2):327–330, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0499-2.

Anonymous:2014:BCW

[Ano14] Anonymous. Blaise Cronin wins the 2013 Derek John de Solla Price Medal. Scientometrics, 98(1):5–10, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1169-3.

Anonymous:2015:SPW

[Ano15] Anonymous. Selected papers of the Workshop “Combining Bibliometrics and Information Retrieval” held as special event at the 14th International Conference of Scientometrics and Informetrics, Vienna, 15 July 2013. Scientometrics, 102(3):2189, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-015-1532-
7.pdf.

Anonymous:2016:EC

[Ano16a] Anonymous. Editorial correction. Scientometrics, 108(3): 1673, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-016-2076-
1.pdf.

Anonymous:2016:GCD

[Ano16b] Anonymous. Grand challenges in data integration for research and innovation (R&I) policy: handling big data, coping with quality issues and anticipating new policy needsstate of the art and future perspectives. Scientometrics, 108(1):389, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-016-1915-
4.pdf.

Anonymous:2016:ROU

[Ano16c] Anonymous. Research organizations under scrutiny new indicators and analytical results. Scientometrics, 109(3): 2159, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-016-2129-
5.pdf.

Anonymous:2016:TSH

[Ano16d] Anonymous. Towards standardisation, harmonisation and integration of data from heterogeneous sources for funding and evaluation purposes. Scientometrics, 106(2):819, February 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1812-2;   http://link.
springer.com/content/pdf/10.1007/s11192-015-1812-
2.pdf.

Anonymous:2017:JBI

[Ano17a] Anonymous. Judit Bar-Ilan wins the 2017 Derek John de Solla Price Medal. Scientometrics, 113(3):1233–1234, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2552-2.

Anonymous:2017:O

[Ano17b] Anonymous. Obituary: Eugene Garfield (1925–2017). Scientometrics, 111(1):1, April 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-017-
2352-8.pdf.

Anonymous:2017:SDD

[Ano17c] Anonymous. Same data-different results? Towards a comparative approach to the identification of thematic structures in science. Scientometrics, 111(2):979, May 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-017-2295-0.pdf.

Anonymous:2017:SPS

[Ano17d] Anonymous. Simulating the processes of science, technology, and innovation. Scientometrics, 110(1):385, January 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s11192-016-2103-2.pdf.

Anonymous:2018:BEI

[Ano18a] Anonymous. Bibliometric-enhanced information retrieval. Scientometrics, 116(2):1223, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2862-z;   http://link.springer.com/
content/pdf/10.1007/s11192-018-2862-z.pdf.

Anonymous:2018:ECP

[Ano18b] Anonymous. Editorial comment on the papers of Bakare and Lewison (DOI: 10.1007/s11192-017-2490-z), and Campanario (DOI: 10.1007/s11192-017-2506-8). Scientometrics, 114(2):779, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2642-9;   https://link.springer.com/content/
pdf/10.1007/s11192-018-2642-9.pdf. See [BL17a, Cam17].

Anonymous:2018:LAS

[Ano18c] Anonymous. Latin American science and technology in a global context. Scientometrics, 115(3):1437, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2756-0;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2756-
0.pdf.

Anonymous:2019:OJB

[Ano19] Anonymous. Obituary: Judit Bar-Ilan (1958–2019). Scientometrics, 121(2):601, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03237-2;   http://link.springer.com/content/
pdf/10.1007/s11192-019-03237-2.pdf.

Amancio:2012:UCN

[ANOdFC12] D. R. Amancio, M. G. V. Nunes, O. N. Oliveira, Jr., and L. da F. Costa. Using complex networks concepts to assess approaches for citations in scientific papers. Scientometrics, 91(3):827–842, June 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0630-z.

Abrizah:2015:LJC

[ANZ15] A. Abrizah, A. Noorhidawati, and A. N. Zainab. LIS journals categorization in the Journal Citation Report: a stated preference study. Scientometrics, 102(2):1083–1099, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1492-3.

Amancio:2015:TCA

[AOd15] Diego R. Amancio, Osvaldo N. Oliveira, Jr., and Luciano da F. Costa. Topological-collaborative approach for disambiguating authors’ names in collaborative networks. Scientometrics, 102(1):465–485, January 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1381-9.

Aguillo:2010:IWR

[AOFU10] Isidro F. Aguillo, José L. Ortega, Mario Fernández, and Ana M. Utrilla. Indicators for a webometric ranking of open access repositories. Scientometrics, 82(3):477–486, March 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0183-y.

Avila-Poveda:2014:TRT

[AP14] Omar Hernando Avila-Poveda. Technical report: the trend of author compound names and its implications for authorship identity identification. Scientometrics, 101(1):833–846, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1359-7.

Amat:2016:ECM

[AP16] Carlos B. Amat and François Perruchas. Evolving cohesion metrics of a research network on rare diseases: a longitudinal study over 14 years. Scientometrics, 108(1):41–56, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1952-z.

Oliveira:2013:TCB

[APFR^{+}13] Eduardo Araujo Oliveira, Roberto Peicots-Filho, Daniella Reis Martelli, Isabel Gomes Quirino, Maria Christina Lopes Oliveira, Mariana Guerra Duarte, Sergio Veloso Pinheiro, Enrico Antonio Colosimo, Ana Cristina Simões e Silva, and Hercílio Martelli-Júnior. Is there a correlation between journal impact factor and researchers’ performance? A study comprising the fields of clinical nephrology and neurosciences. Scientometrics, 97(2):149–160, November 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-0992-x.

Amador:2018:ISM

[APLHF18] Soleidy Rivero Amador, Maidelyn Díaz Pérez, María José López-Huertas, and Reinaldo Javier Rodríguez Font. Indicator system for managing science, technology and innovation in universities. Scientometrics, 115(3):1575–1587, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-018-2721-y. See correc-

tion [APPF18].

Amador:2018:CIS

[APPF18] Soleidy Rivero Amador, Maidelyn Díaz Pérez, María José López-Huertas Pérez, and Reinaldo Javier Rodríguez Font. Correction to: Indicator system for managing science, technology and innovation in universities. Scientometrics, 115 (3):1589, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2794-7;
http://link.springer.com/content/pdf/10.1007/s11192-
018-2794-7.pdf. See [APLHF18].

Ahlgren:2015:BAT

[APPS15] Per Ahlgren, Peter Pagin, Olle Persson, and Maria Svedberg. Bibliometric analysis of two subdomains in philosophy: free will and sorites. Scientometrics, 103(1):47–73, April 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1535-4.

Aksnes:2019:GGI

[APR19] Dag W. Aksnes, Fredrik Niclas Piro, and Kristoffer Rørstad. Gender gaps in international research collaboration: a bibliometric approach. Scientometrics, 120(2):747–774, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03155-3;   http://
link.springer.com/content/pdf/10.1007/s11192-019-
03155-3.pdf.

Ahlgren:2013:GDB

[APT13] Per Ahlgren, Olle Persson, and Robert Tijssen. Geographical distance in bibliometric relations within epistemic communities. Scientometrics, 95(2):771–784, May 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0819-1.

Arora:2013:CND

[APYS13] Sanjay K. Arora, Alan L. Porter, Jan Youtie, and Philip Shapira. Capturing new developments in an emerging technology: an updated search strategy for identifying nanotechnology research outputs. Scientometrics, 95(1):351–370, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0903-6.

Ainsworth:2018:HSD

[AR18] Shirley Ainsworth and Jane M. Russell. Has hosting on science direct improved the visibility of Latin American scholarly journals? A preliminary analysis of data quality. Scientometrics, 115(3):1463–1484, June 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2725-7.

Arbesman:2011:QES

[Arb11] Samuel Arbesman. Quantifying the ease of scientific discovery. Scientometrics, 86(2):245–250, February 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0232-6.

Ardanuy:2012:SCL

[Ard12] Jordi Ardanuy. Scientific collaboration in library and information science viewed through the Web of Knowledge: the Spanish case. Scientometrics, 90(3):877–890, March 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0552-1.

AlRyalat:2018:CER

[ARE^{+}18] Saif Aldeen AlRyalat, Khaled Rawashdeh, Osama El khatib, Abeer Yasin, Fadwa Alqadi, Noor Saleh, Lna Malkawi, Ola Hijjawi, and Mohammad Alessa. The change from an eponym to a representative name: Wegener to granulomatosis with polyangiitis. Scientometrics, 117(3):2077–2089, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2951-z.

Abdullah:2015:EPE

[ARK^{+}15] Noorhidawati Abdullah, Siti Hajar Mohd Roffeei, Yusniza Kamarulzaman, Farrah Dina Yusop, Azian Madun, and Kwan Hoong Ng. Evaluating the performance of electromagnetic fields (EMF) research work (2003–2013). Scientometrics, 105(1):261–278, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1657-8.

Avila-Robinson:2013:EPC

[ÁRM13] Alfonso Ávila-Robinson and Kumiko Miyazaki. Evolutionary paths of change of emerging nanotechnological innovation systems: the case of ZnO nanostructures. Scientometrics, 95(3):829–849, June 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0939-7.

Avila-Robinson:2017:TKB

[ÁRS17] Alfonso Ávila-Robinson and Shintaro Sengoku. Tracing the knowledge-building dynamics in new stem cell technologies through techno-scientific networks. Scientometrics, 112(3):1691–1720, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2436-5.

Alluqmani:2018:WSD

[AS18a] Amnah Alluqmani and Lior Shamir. Writing styles in different scientific disciplines: a data science approach. Scientometrics, 115(2):1071–1085, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2688-8.

Avanesova:2018:CTR

[AS18b] Anna A. Avanesova and Tatyana A. Shamliyan. Comparative trends in research performance of the Russian universities. Scientometrics, 116(3):2019–2052, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2807-6.

Asai:2019:CRS

[Asa19] Sumiko Asai. Changes in revenue structure of a leading open access journal publisher: the case of BMC. Scientometrics, 121(1):53–63, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03200-1.

Asemi:2010:CAI

[Ase10] Asefeh Asemi. A citation analysis of Iranian journals to open access (OA) articles and journals. Scientometrics, 82 (3):487–494, March 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0184-x.

Askeridis:2018:IMC

[Ask18] Johanna M. Askeridis. An h index for Mendeley: comparison of citation-based h indices and a readership-based h_{men} index for 29 authors. Scientometrics, 117(1):615–624, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2882-8.

Alves-Silva:2016:IFO

[ASPF^{+}16] Estevao Alves-Silva, Ana Carolina Figueira Porto, Carine Firmino, Henrique Venancio Silva, Ingrid Becker, Liegy Resende, Livia Borges, Luana Pfeffer, Marcela Silvano, Melina Santos Galdiano, Rafaella Silvestrini, and Renan Moura. Are the impact factor and other variables related to publishing time in ecology journals? Scientometrics, 108(3):1445–1453, September 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2040-0.

Shimada:2017:PSI

[aSS17] Yoshi aki Shimada and Jun Suzuki. Promoting scientodiversity inspired by biodiversity. Scientometrics, 113(3): 1463–1479, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2545-1.

Shimada:2017:PDS

[aSTS17] Yoshi aki Shimada, Naotoshi Tsukada, and Jun Suzuki. Promoting diversity in science in Japan through missionoriented research grants. Scientometrics, 110(3):1415–1435, March 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2224-7.

Asubiaro:2019:HCT

[Asu19] Toluwase Asubiaro. How collaboration type, publication place, funding and author’s role affect citations received by publications from Africa: A bibliometric study of LIS research from 1996 to 2015. Scientometrics, 120(3):1261– 1287, September 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03157-1.

Azeroual:2018:DMR

[ASW18] Otmane Azeroual, Gunter Saake, and Jürgen Wastl. Data measurement in research information systems: metrics for the evaluation of data quality. Scientometrics, 115(3): 1271–1290, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2735-5.

Abdulhayoglu:2017:URG

[AT17] Mehmet Ali Abdulhayoglu and Bart Thijs. Use of Research-Gate and Google CSE for author name disambiguation. Scientometrics, 111(3):1965–1985, June 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Abdulhayoglu:2018:ULS

[AT18] Mehmet Ali Abdulhayoglu and Bart Thijs. Use of locality sensitive hashing (LSH) algorithm to match Web of Science and Scopus. Scientometrics, 116(2):1229–1245, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2569-6.

Aleixandre-Tudo:2019:USR

[ATCCAAB19] Jose Luis Aleixandre-Tudo, Lourdes Castelló-Cogollos, Jose Luis Aleixandre, and Rafael Aleixandre-Benavent. Unravelling the scientific research on grape and wine phenolic compounds: a bibliometric study. Scientometrics, 119 (1):119–147, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03029-8.

Abdulhayoglu:2016:UCG

[ATJ16] Mehmet Ali Abdulhayoglu, Bart Thijs, and Wouter Jeuris. Using character n-grams to match a list of publications to references in bibliographic databases. Scientometrics, 109(3):1525–1546, December 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2066-3.

Aduku:2017:DMR

[ATK17] Kuku Joseph Aduku, Mike Thelwall, and Kayvan Kousha. Do Mendeley reader counts reflect the scholarly impact of conference papers? An investigation of computer science and engineering. Scientometrics, 112(1):573–581, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Andrei:2016:BIF

[ATM16] Tudorel Andrei, Daniel Teodorescu, and Andreea Mirica. Beyond the impact factor: measuring the international visibility of Romanian social sciences journals. Scientometrics, 108(1):1–20, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1949-7.

Alstott:2017:MTS

[ATYL17] Jeff Alstott, Giorgio Triulzi, Bowen Yan, and Jianxi Luo. Mapping technology space by normalizing patent networks. Scientometrics, 110(1):443–479, January 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2107-y.

Abercrombie:2012:SSM

[AUS12] Robert K. Abercrombie, Akaninyene W. Udoeyop, and Bob G. Schlicher. A study of scientometric methods to identify emerging technologies via modeling of milestones. Scientometrics, 91(2):327–342, May 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0614-4.

Ausloos:2013:SLA

[Aus13] M. Ausloos. A scientometrics law about co-authors and their ranking: the co-author core. Scientometrics, 95(3):895–909, June 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0936-x.

Ausloos:2014:BSS

[Aus14a] Marcel Ausloos. Binary scientific star coauthors core size. Scientometrics, 99(2):331–351, May 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1230-x.

Ausloos:2014:ZMP

[Aus14b] Marcel Ausloos. Zipf–Mandelbrot–Pareto model for coauthorship popularity. Scientometrics, 101(3):1565–1586, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1302-y.

Avkiran:2013:EII

[Avk13] Necmi K. Avkiran. An empirical investigation of the influence of collaboration in finance on article impact. Scientometrics, 95(3):911–925, June 2013. CODEN SCNTDX.

ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0892-5.

Aksnes:2014:EBC

[AvLS14] Dag W. Aksnes, Thed N. van Leeuwen, and Gunnar Sivertsen. The effect of booming countries on changes in the relative specialization index (RSI) on country level. Scientometrics, 101(2):1391–1401, November 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1245-3.

Ajiferuke:2010:CAM

[AW10] Isola Ajiferuke and Dietmar Wolfram. Citer analysis as a measure of research impact: library and information science as a case study. Scientometrics, 83(3):623–638, June 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-009-0127-6.

An:2011:CWA

[AW11] Xin Ying An and Qing Qiang Wu. Co-word analysis of the trends in stem cells field based on subject heading weighting. Scientometrics, 88(1):133–144, July 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0374-1.

Arora:2013:ESE

[AYS^{+}13] Sanjay K. Arora, Jan Youtie, Philip Shapira, Lidan Gao, and TingTing Ma. Entry strategies in an emerging technology: a pilot web-based study of graphene firms. Scientometrics, 95(3):1189–1207, June 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0950-7.

Alves:2014:BLA

[AYS14] Alexandre Donizeti Alves, Horacio Hideki Yanasse, and Nei Yoshihiro Soma. Benford’s Law and articles of scientific journals: comparison of JCR(R) and Scopus data. Scientometrics, 98(1):173–184, January 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-013-1030-8.pdf.

Alves:2016:ABI

[AYS16] Alexandre Donizeti Alves, Horacio Hideki Yanasse, and Nei Yoshihiro Soma. An analysis of bibliometric indicators to JCR according to Benford’s Law. Scientometrics, 107(3):1489–1499, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com//content/
pdf/10.1007/s11192-016-1908-3.pdf;   http://link.
springer.com/content/pdf/10.1007/s11192-016-1908-
3.pdf.

Abrizah:2013:LJS

[AZKR13] A. Abrizah, A. N. Zainab, K. Kiran, and R. G. Raj. LIS journals scientific impact and subject categorization: a comparison between Web of Science and Scopus. Scientometrics, 94(2):721–740, February 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0813-7.

Akhmat:2014:RAE

[AZSA14] Ghulam Akhmat, Khalid Zaman, Tan Shukui, and Tauseef Ahmed. RETRACTED ARTICLE: Educational reforms and internationalization of universities: evidence from major regions of the world. Scientometrics, 98(3):2185–2205, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1130-5.

Akhmat:2016:RNE

[AZSA16] Ghulam Akhmat, Khalid Zaman, Tan Shukui, and Tauseef Ahmed. Retraction note to: Educational reforms and internationalization of universities: evidence from major regions of the world. Scientometrics, 108(2):999, August 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-014-1305-8.pdf.

Brischoux:2015:ANE

[BA15] François Brischoux and Frédéric Angelier. Academia’s never-ending selection for productivity. Scientometrics, 103 (1):333–336, April 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1534-5.

Barirani:2013:DAF

[BAB13] Ahmad Barirani, Bruno Agard, and Catherine Beaudry. Discovering and assessing fields of expertise in nanomedicine: a patent co-citation network perspective. Scientometrics, 94 (3):1111–1136, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0891-6.

Bordons:2013:HCR

[BAC13] María Bordons, Javier Aparicio, and Rodrigo Costas. Heterogeneity of collaboration and its relationship with research impact in a biomedical field. Scientometrics, 96(2):443–466, August 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0890-7.

Bai:2018:GSB

[Bai18] Yang Bai. Has the Global South become a playground for Western scholars in information and communication technologies for development? Evidence from a threejournal analysis. Scientometrics, 116(3):2139–2153, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2839-y.

Bakker:2017:LLR

[Bak17] Jurriën Bakker. The log–linear relation between patent citations and patent value. Scientometrics, 110(2):879–892, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2208-7.

Balaban:2012:PNA

[Bal12] Alexandru T. Balaban. Positive and negative aspects of citation indices and journal impact factors. Scientometrics, 92 (2):241–247, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0637-5.

Bangani:2018:IET

[Ban18] Siviwe Bangani. The impact of electronic theses and dissertations: a study of the institutional repository of a university in South Africa. Scientometrics, 115(1):131–151, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2657-2.

Bartneck:2011:EBR

[Bar11] Christoph Bartneck. The end of the beginning: a reflection on the first five years of the HRI conference. Scientometrics, 86(2):487–504, February 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-010-0281-x.pdf.

Barrot:2017:RIP

[Bar17a] Jessie S. Barrot. Research impact and productivity of Southeast Asian countries in language and linguistics. Scientometrics, 110(1):1–15, January 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2163-3.

Bartneck:2017:RSD

[Bar17b] Christoph Bartneck. Reviewers’ scores do not predict impact: bibliometric analysis of the proceedings of the human–robot interaction conference. Scientometrics, 110(1):179–194, January 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2176-y.

Basu:2010:DCS

[Bas10] Aparna Basu. Does a country’s scientific ‘productivity’ depend critically on the number of country journals indexed? Scientometrics, 82(3):507–516, March 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0186-8.

Baskurt:2011:TSA

[Bas11] Oguz K. Baskurt. Time series analysis of publication counts of a university: what are the implications? Scientometrics, 86(3):645–656, March 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0298-1.

Basu:2013:SDR

[Bas13] Aparna Basu. Some differences in research publications of Indian scientists in India and the diaspora, 1986– 2010. Scientometrics, 94(3):1007–1019, March 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0800-z.

Basu:2014:AME

[Bas14] Aparna Basu. The Albuquerque model and efficiency indicators in national scientific productivity with respect to manpower and funding in science. Scientometrics, 100(2): 531–539, August 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1254-2.

Bertin:2016:LPR

[BASL16] Marc Bertin, Iana Atanassova, Cassidy R. Sugimoto, and Vincent Lariviere. The linguistic patterns and rhetorical structure of citation context: an approach using n-grams. Scientometrics, 109(3):1417–1434, December 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2134-8.

Baccini:2010:IEN

[BB10] Alberto Baccini and Lucio Barabesi. Interlocking editorship. A network analysis of the links between economic journals. Scientometrics, 82(2):365–389, February 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0053-7.

Bornmann:2015:EHC

[BB15] Lutz Bornmann and Johann Bauer. Evaluation of the highly-cited researchers’ database for a country: proposals for meaningful analyses on the example of Germany. Scientometrics, 105(3):1997–2003, December 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1619-1.

Bertoli-Barsotti:2016:NI

[BB16] Lucio Bertoli-Barsotti. Normalizing the g-index. Scientometrics, 106(2):645–655, February 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1794-0.

Bertoli-Barsotti:2017:RCP

[BB17a] Lucio Bertoli-Barsotti. Reply to the comments of Prathap. Scientometrics, 112(2):1137–1140, August 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2420-0. See [BBL17b, Pra17a].

Buttliere:2017:PPU

[BB17b] Brett Buttliere and Jürgen Buder. Personalizing papers using Altmetrics: comparing paper ‘quality’ or ‘impact’ to person ‘intelligence’ or ‘personality’. Scientometrics, 111(1):219–239, April 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-017-2246-9.

Belli:2019:SSP

[BB19] Simone Belli and Joan Baltà. Stocktaking scientific publication on bi-regional collaboration between Europe 28 and Latin America and the Caribbean. Scientometrics, 121(3): 1447–1480, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03266-x.

Baccini:2014:CHD

[BBCP14] A. Baccini, L. Barabesi, M. Cioni, and C. Pisani. Crossing the hurdle: the determinants of individual scientific performance. Scientometrics, 101(3):2035–2062, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-014-1395-3.

Bartol:2014:ARF

[BBDS^{+}14] Tomaz Bartol, Gordana Budimir, Doris Dekleva-Smrekar, Miro Pusnik, and Primoz Juznic. Assessment of research fields in Scopus and Web of Science in the view of national research evaluation in Slovenia. Scientometrics, 98(2): 1491–1504, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1148-8.

Bartol:2016:MCA

[BBJS16] Tomaz Bartol, Gordana Budimir, Primoz Juznic, and Karmen Stopar. Mapping and classification of agriculture in Web of Science: other subject categories and research fields may benefit. Scientometrics, 109(2):979–996, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2071-6.

Bertoli-Barsotti:2017:IAE

[BBL17a] Lucio Bertoli-Barsotti and Tommaso Lando. The hindex as an almost-exact function of some basic statistics. Scientometrics, 113(2):1209–1228, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2508-6;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2508-
6.pdf.

Bertoli-Barsotti:2017:TMR

[BBL17b] Lucio Bertoli-Barsotti and Tommaso Lando. A theoretical model of the relationship between the h-index and other simple citation indicators. Scientometrics, 111(3): 1415–1448, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-017-2351-

9.pdf. See comments [Pra17a] and reply [BB17a].

Bogocz:2014:NFL

[BBP14] Jacek Bogocz, Andrzej Bak, and Jaroslaw Polanski. No free lunches in nature? An analysis of the regional distribution of the affiliations of Nature publications. Scientometrics, 101 (1):547–568, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1252-4.

Bornmann:2017:CHC

[BBS17] Lutz Bornmann, Johann Bauer, and Elisabeth Maria Schlagberger. Characteristics of highly cited researchers 2015 in Germany. Scientometrics, 111(1):543–545, April 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-017-2248-7.

Basu:2016:DCI

[BBSS16a] Aparna Basu, Sumit Kumar Banshal, Khushboo Singhal, and Vivek Kumar Singh. Designing a Composite Index for research performance evaluation at the national or regional level: ranking Central Universities in India. Scientometrics, 107(3):1171–1193, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-016-1935-0. See comments [Pra16b,

BBSS16c] and erratum [BBSS16b].

Basu:2016:EDC

[BBSS16b] Aparna Basu, Sumit Kumar Banshal, Khushboo Singhal, and Vivek Kumar Singh. Erratum to: “Designing a Composite Index for research performance evaluation at the national or regional level: ranking Central Universities in India”. Scientometrics, 108(3):1695–1697, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s11192-016-2009-z.pdf. See [BBSS16a].

Basu:2016:RLE

[BBSS16c] Aparna Basu, Sumit Kumar Banshal, Khushboo Singhal, and Vivek Kumar Singh. Response to the letter to the Editor by Gangan Prathap on the article: Designing a composite index for research performance evaluation at the national or regional level: ranking Central Universities in India. Scientometrics, 108(3):1689–1691, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/

accesspage/article/10.1007/s11192-016-1996-0. See

[BBSS16a, Pra16b].

Billaut:2010:SYB

[BBV10] Jean-Charles Billaut, Denis Bouyssou, and Philippe Vincke. Should you believe in the Shanghai ranking? Scientometrics, 84(1):237–263, July 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
009-0115-x.

Borrego:2010:SOI

[BBVO10] Ángel Borrego, Maite Barrios, Anna Villarroya, and Candela Ollé. Scientific output and impact of postdoctoral scientists: a gender perspective. Scientometrics, 83(1):93–101, April 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0025-y.

Batagelj:2013:BN

[BC13a] Vladimir Batagelj and Monika Cerinsek. On bibliographic networks. Scientometrics, 96(3):845–864, September 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0940-1.

Bosquet:2013:AWP

[BC13b] Clément Bosquet and Pierre-Philippe Combes. Are academics who publish more also more cited? Individual determinants of publication and citation records. Scientometrics, 97(3):831–857, December 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0996-6.

Boudry:2017:ADO

[BC17] Christophe Boudry and Ghislaine Chartron. Availability of Digital Object Identifiers in publications archived by PubMed. Scientometrics, 110(3):1453–1469, March 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2225-6.

Bras:2017:ORL

[BCC^{+}17] Oriana Rainho Brás, Jean-Philippe Cointet, Alberto Cambrosio, Leonor David, João Arriscado Nunes, Fátima Cardoso, and Carmen Jerónimo. Oncology research in late twentieth century and turn of the century Portugal: a scientometric approach to its institutional and semantic dimensions. Scientometrics, 113(2):867–888, November 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2491-y.

Bonaccorsi:2017:ETG

[BCHH17] Andrea Bonaccorsi, Tindaro Cicero, Peter Haddawy, and Saeed-Ul Hassan. Explaining the transatlantic gap in research excellence. Scientometrics, 110(1):217–241, January 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2180-2.

Bollen:2017:ESF

[BCJ^{+}17] Johan Bollen, David Crandall, Damion Junk, Ying Ding, and Katy Börner. An efficient system to fund science: from proposal review to peer-to-peer distributions. Scientometrics, 110(1):521–528, January 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2110-3.

Ba:2019:HAA

[BCML19] Zhichao Ba, Yujie Cao, Jin Mao, and Gang Li. A hierarchical approach to analyzing knowledge integration between two fields — a case study on medical informatics and computer science. Scientometrics, 119(3):1455–1486, June 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03103-1.

Bi:2019:SES

[BCT19] Weilong Bi, Ho Fai Chan, and Benno Torgler. Self-esteem, self-symbolizing, and academic recognition: behavioral evidence from curricula vitae. Scientometrics, 119(1):495–525, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03037-8.

Buela-Casal:2012:WDS

[BCZ12] Gualberto Buela-Casal and Izabela Zych. What do the scientists think about the impact factor? Scientometrics, 92 (2):281–292, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0676-y.

Ben-David:2010:RIE

[BD10a] Dan Ben-David. Ranking Israel’s economists. Scientometrics, 82(2):351–364, February 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0049-3.

Bornmann:2010:VSE

[BD10b] Lutz Bornmann and Hans-Dieter Daniel. The validity of staff editors’ initial evaluations of manuscripts: a case study of Angewandte Chemie International Edition. Scientometrics, 85(3):681–687, December 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0215-7.

Baron:2012:PSV

[BD12a] Justus Baron and Henry Delcamp. The private and social value of patents in discrete and cumulative innovation. Scientometrics, 90(2):581–606, February 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0532-5.

Basu:2012:CMM

[BD12b] Aparna Basu and Roland Wagner Dobler. ‘Cognitive mobility’ or migration of authors between fields used in mapping a network of mathematics. Scientometrics, 91(2):353–368, May 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0613-5.

Bougnol:2013:MMO

[BD13] M. L. Bougnol and J. H. Dulá. A mathematical model to optimize decisions to impact multi-attribute rankings. Scientometrics, 95(2):785–796, May 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0844-0.

Baccini:2016:DTA

[BD16a] Alberto Baccini and Giuseppe De Nicolao. Do they agree? Bibliometric evaluation versus informed peer review in the Italian research assessment exercise. Scientometrics, 108(3):1651–1671, September 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-1929-y. See comment [BGJ^{+}16].

Baccini:2016:RCB

[BD16b] Alberto Baccini and Giuseppe De Nicolao. Reply to the comment of Bertocchi et al. Scientometrics, 108(3):1675– 1684, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-016-
2055-6.

Bucheli:2012:GSP

[BDC^{+}12] Victor Bucheli, Adriana Díaz, Juan Pablo Calderón, Pablo Lemoine, Juan Alejandro Valdivia, José Luis Villaveces, and Roberto Zarama. Growth of scientific production in Colombian universities: an intellectual capital-based approach. Scientometrics, 91(2):369–382, May 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0627-7.

Bouabid:2011:IEV

[BDE11] Hamid Bouabid, Mohamed Dalimi, and Zayer ElMajid. Impact evaluation of the voluntary early retirement policy on research and technology outputs of the faculties of science in Morocco. Scientometrics, 86(1):125–132, January 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0271-z.

Bonaccorsi:2017:DSS

[BDF^{+}17] Andrea Bonaccorsi, Cinzia Daraio, Stefano Fantoni, Viola Folli, Marco Leonetti, and Giancarlo Ruocco. Do social sciences and humanities behave like life and hard sciences? Scientometrics, 112(1):607–653, July 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Belter:2013:BAN

[Bel13] Chris W. Belter. A bibliometric analysis of NOAA’s Office of Ocean Exploration and Research. Scientometrics, 95 (2):629–644, May 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-012-0836-
0.pdf.

Belter:2017:RRM

[Bel17] Christopher W. Belter. A relevance ranking method for citation-based search results. Scientometrics, 112(2):731– 746, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2406-y.

Borner:2017:E

[BEMS17] Katy Börner, Bruce Edmonds, Stasa Milojević, and Andrea Scharnhorst. Editorial. Scientometrics, 110(1):387–390, January 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-016-2104-1.pdf.

Bensman:2011:AWH

[Ben11] Stephen J. Bensman. Anne-wil harzing: The publish or perish book: Your guide to effective and responsible citation analysis. Scientometrics, 88(1):339–342, July 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0388-8.

Bensman:2012:IFP

[Ben12] Stephen J. Bensman. The impact factor: its place in Garfield’s thought, in science evaluation, and in library collection management. Scientometrics, 92(2):263–275, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0601-9.

Bentley:2015:CCD

[Ben15] Peter James Bentley. Cross-country differences in publishing productivity of academics in research universities. Scientometrics, 102(1):865–883, January 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1430-4.

Benjafield:2019:KFA

[Ben19] John G. Benjafield. Keyword frequencies in anglophone psychology. Scientometrics, 118(3):1051–1064, March 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03021-2.

Berker:2018:GRS

[Ber18] Yannick Berker. Golden-ratio as a substitute to geometric and harmonic counting to determine multi-author publication credit. Scientometrics, 114(3):839–857, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2632-3.

Burns:2017:LSP

[BF17] C. Sean Burns and Charles W. Fox. Language and socioeconomics predict geographic variation in peer review outcomes at an ecology journal. Scientometrics, 113(2):1113– 1127, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2517-5.

Basurto-Flores:2018:ERA

[BFGVV^{+}18] R. Basurto-Flores, L. Guzmán-Vargas, S. Velasco, A. Medina, and A. Calvo Hernandez. On entropy research analysis: cross-disciplinary knowledge transfer. Scientometrics, 117 (1):123–139, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2860-1;
http://link.springer.com/content/pdf/10.1007/s11192-
018-2860-1.pdf.

Basu:2018:CRL

[BFHS18] A. Basu, P. Foland, G. Holdridge, and R. D. Shelton. China’s rising leadership in science and technology: quantitative and qualitative indicators. Scientometrics, 117(1): 249–269, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2877-5.

Brunson:2014:EEM

[BFM^{+}14] Jason Cory Brunson, Steve Fassino, Antonio McInnes, Monisha Narayan, Brianna Richardson, Christopher Franck, Patrick Ion, and Reinhard Laubenbacher. Evolutionary events in a mathematical sciences research collaboration network. Scientometrics, 99(3):973–998, June 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1209-z.

Barrios:2019:TCI

[BFMRM19] Candelaria Barrios, Esther Flores, M. Ángeles Martínez, and Marta Ruiz-Martínez. Is there convergence in international research collaboration? an exploration at the country level in the basic and applied science fields. Scientometrics, 120(2):631–659, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03133-9.

Batagelj:2017:EFN

[BFS17] Vladimir Batagelj, Anuska Ferligoj, and Flaminio Squazzoni. The emergence of a field: a network analysis of research on peer review. Scientometrics, 113(1):503–532, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2522-8;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2522-
8.pdf.

Bigdeli:2012:ASI

[BG12] Zahed Bigdeli and Ali Gazni. Authors’ sources of information: a new dimension in information scattering.

Scientometrics, 92(3):505–521, September 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0609-1.

Bornmann:2017:ACM

[BG17] Lutz Bornmann and Wolfgang Glänzel. Applying the CSS method to bibliometric indicators used in (university) rankings. Scientometrics, 110(2):1077–1079, February 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2198-5.

Bornmann:2018:WDC

[BG18] Lutz Bornmann and Wolfgang Glänzel. Which differences can be expected when two universities in the Leiden Ranking are compared? Some benchmarks for institutional research evaluations. Scientometrics, 115(2):1101–1105, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2700-3.

Bordons:2015:IRI

[BGAAM15] María Bordons, Borja González-Albo, Javier Aparicio, and Luz Moreno. The influence of R&D intensity of countries on the impact of international collaborative research: evidence from Spain. Scientometrics, 102(2):1385–1400, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1491-4.

Bianchi:2018:PRG

[BGBS18] Federico Bianchi, Francisco Grimaldo, Giangiacomo Bravo, and Flaminio Squazzoni. The peer review game: an agent-based model of scientists facing resource constraints and institutional pressures. Scientometrics, 116(3):1401– 1420, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2825-4;
http://link.springer.com/content/pdf/10.1007/s11192-
018-2825-4.pdf.

Boyack:2017:TIC

[BGG^{+}17] Kevin Boyack, Wolfgang Glänzel, Jochen Gläser, Frank Havemann, Andrea Scharnhorst, Bart Thijs, Nees Jan van Eck, Theresa Velden, and Ludo Waltmann. Topic identification challenge. Scientometrics, 111(2):1223–1224, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-017-2307-0.pdf.

Bertocchi:2016:CDT

[BGJ^{+}16] Graziella Bertocchi, Alfonso Gambardella, Tullio Jappelli, Carmela Anna Nappi, and Franco Peracchi. Comment to: Do they agree? Bibliometric evaluation versus informed peer review in the Italian research assessment exercise. Scientometrics, 108(1):349–353, July 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1965-7. See [BD16a].

Brizan:2016:PCP

[BGJB16] David Guy Brizan, Kevin Gallagher, Arnab Jahangir, and Theodore Brown. Predicting citation patterns: defining and determining influence. Scientometrics, 108(1):183–200, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1950-1.

Bildosola:2017:AMF

[BGM17] Iñaki Bildosola, Pilar Gonzalez, and Paz Moral. An approach for modelling and forecasting research activity related to an emerging technology. Scientometrics, 112(1): 557–572, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Bernabo:2016:SAR

[BGMB16] Nicola Bernabò, Luana Greco, Mauro Mattioli, and Barbara Barboni. A scientometric analysis of reproductive medicine. Scientometrics, 109(1):103–120, October 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-1969-3.

Bilir:2013:RPT

[BGÖ^{+}13] Selçuk Bilir, Ersin Gögüs, Özgecan Önal, Nazli Derya Öztürkmen, and Talar Yontan. Research performance of Turkish astronomers in the period of 1980–2010. Scientometrics, 97(2):477–489, November 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0922-3.

Benito:2019:FIK

[BGR19] Mónica Benito, Pilar Gil, and Rosario Romera. Funding, is it key for standing out in the university rankings? Scientometrics, 121(2):771–792, November 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03202-z.

Borner:2011:MSS

[BGSvdB11] Katy Börner, Wolfgang Glänzel, Andrea Scharnhorst, and Peter van den Besselaar. Modeling science: studying the structure and dynamics of science. Scientometrics, 89(1): 347–348, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-011-0429-
3.pdf.

Bartneck:2010:FCM

[BH10] Christoph Bartneck and Jun Hu. The fruits of collaboration in a multidisciplinary field. Scientometrics, 85(1):41–52, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-010-0242-4.pdf.

Bornmann:2015:ISC

[BH15] Lutz Bornmann and Robin Haunschild. The interest of the scientific community in expert opinions from journal peer review procedures. Scientometrics, 102(3):2187–2188, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-014-1514-1.

Biesenbender:2016:RCDa

[BH16a] Sophie Biesenbender and Stefan Hornbostel. The Research Core Dataset for the German science system: challenges, processes and principles of a contested standardization project. Scientometrics, 106(2):837–847, February 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1816-y.

Biesenbender:2016:RCDb

[BH16b] Sophie Biesenbender and Stefan Hornbostel. The Research Core Dataset for the German science system: developing standards for an integrated management of research information. Scientometrics, 108(1):401–412, July 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1909-2.

Bornmann:2016:HNT

[BH16c] Lutz Bornmann and Robin Haunschild. How to normalize Twitter counts? A first attempt based on journals in the Twitter Index. Scientometrics, 107(3):1405–1422, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com//content/pdf/10.1007/s11192-016-1893-6.pdf;
http://link.springer.com/content/pdf/10.1007/s11192-
016-1893-6.pdf.

Bornmann:2017:DES

[BH17a] Lutz Bornmann and Robin Haunschild. Does evaluative scientometrics lose its main focus on scientific quality by the new orientation towards societal impact? Scientometrics, 110(2):937–943, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-2200-2.pdf.

Bornmann:2017:QIC

[BH17b] Lutz Bornmann and Robin Haunschild. Quality and impact considerations in bibliometrics: a reply to Ricker (in press). Scientometrics, 111(3):1857–1859, June 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Bornmann:2018:ASM

[BH18a] Lutz Bornmann and Robin Haunschild. Allegation of scientific misconduct increases Twitter attention. Scientometrics, 115(2):1097–1100, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2698-6.

Bornmann:2018:PVP

[BH18b] Lutz Bornmann and Robin Haunschild. Plots for visualizing paper impact and journal impact of single researchers in a single graph. Scientometrics, 115(1):385–394, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2658-1;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2658-
1.pdf.

Bharathi:2011:MES

[Bha11] D. Gnana Bharathi. Methodology for the evaluation of scientific journals: Aggregated citations of cited articles. Scientometrics, 86(3):563–574, March 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0291-8.

Badar:2015:KNC

[BHA15] Kamal Badar, Julie M. Hite, and Naeem Ashraf. Knowledge network centrality, formal rank and research performance: evidence for curvilinear and interaction effects. Scientometrics, 105(3):1553–1576, December 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1652-0.

Bhardwaj:2016:SAD

[Bha16] Rajesh Kumar Bhardwaj. Scientometric analysis and dimensions on international business literature. Scientometrics, 106(1):299–317, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1777-1.

Bhattacharya:2018:EGB

[Bha18] Sujit Bhattacharya. Eugene Garfield: brief reflections. Scientometrics, 114(2):401–407, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2620-7.

Badar:2013:ERC

[BHB13] Kamal Badar, Julie M. Hite, and Yuosre F. Badir. Examining the relationship of co-authorship network centrality and gender on academic research performance: the case of chemistry researchers in Pakistan. Scientometrics, 94(2):755–775, February 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0764-z.

Boja:2018:DWS

[BHDI18] Catalin Emilian Boja, Claudiu Herteliu, Marian Dârdala, and Bogdan Vasile Ileanu. Day of the week submission effect for accepted papers in Physica A, PLOS ONE, Nature and Cell. Scientometrics, 117(2):887–918, November 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2911-7.

Bornmann:2018:VCC

[BHH18] Lutz Bornmann, Robin Haunschild, and Sven E. Hug. Visualizing the context of citations referencing papers published by Eugene Garfield: a new type of keyword cooccurrence analysis. Scientometrics, 114(2):427–437, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-017-2591-8;   https://
link.springer.com/content/pdf/10.1007/s11192-017-
2591-8.pdf.

Bornmann:2012:PPR

[BHJD12] Lutz Bornmann, Hanna Herich, Hanna Joos, and Hans-Dieter Daniel. In public peer review of submitted manuscripts, how do reviewer comments differ from comments written by interested members of the scientific community? A content analysis of comments written for Atmospheric Chemistry and Physics. Scientometrics, 93(3):

915–929, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0731-8.

Barnett:2011:CAC

[BHKP11] George A. Barnett, Catherine Huh, Youngju Kim, and Han Woo Park. Citations among communication journals and other disciplines: a network analysis. Scientometrics, 88 (2):449–469, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0381-2.

Borner:2010:RNR

[BHL^{+}10] Katy Börner, Weixia Huang, Micah Linnemeier, Russell J. Duhon, Patrick Phillips, Nianli Ma, Angela M. Zoss, Hanning Guo, and Mark A. Price. Rete-netzwerk-red: analyzing and visualizing scholarly networks using the Network Workbench Tool. Scientometrics, 83(3):863–876, June 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0149-0.

Bornmann:2018:RPY

[BHL18] Lutz Bornmann, Robin Haunschild, and Loet Leydesdorff. Reference publication year spectroscopy (RPYS) of eugene Garfield’s publications. Scientometrics, 114(2):439–448, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-017-2608-3;   https://
link.springer.com/content/pdf/10.1007/s11192-017-
2608-3.pdf.

Bornmann:2016:PDS

[BHM16] Lutz Bornmann, Robin Haunschild, and Werner Marx. Policy documents as sources for measuring societal impact: how often is climate change research mentioned in policyrelated documents? Scientometrics, 109(3):1477–1495, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-016-2115-y.pdf.

Barahona:2018:IRT

[BHPVdPMR18] Igor Barahona, Daría Micaela Hernández, Héctor Hugo Pérez-Villarreal, and María del Pilar Martínez-Ruíz. Identifying research topics in marketing science along the past decade: a content analysis. Scientometrics, 117(1):293–312, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2851-2.

Barth:2014:LSG

[BHS14] Martin Barth, Stefanie Haustein, and Barbara Scheidt. The life sciences in German–Chinese cooperation: an institutional-level co-publication analysis. Scientometrics, 98(1):99–117, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1147-9.

Bar-Ilan:2010:CII

[BI10a] Judit Bar-Ilan. Citations to the “Introduction to informetrics” indexed by WOS, Scopus and Google Scholar. Scientometrics, 82(3):495–506, March 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0185-9.

Bar-Ilan:2010:WSC

[BI10b] Judit Bar-Ilan. Web of Science with the Conference Proceedings Citation Indexes: the case of computer science. Scientometrics, 83(3):809–824, June 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0145-4.

Bar-Ilan:2012:EJR

[BI12a] Judit Bar-Ilan. Erratum to: Journal report card. Scientometrics, 92(2):261, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-012-

0795-5.pdf. See [BI12b].

Bar-Ilan:2012:JRC

[BI12b] Judit Bar-Ilan. Journal report card. Scientometrics, 92(2):249–260, August 2012. CODEN SCNTDX.

ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-

012-0671-3. See erratum [BI12a].

Bar-Ilan:2014:APA

[BI14] Judit Bar-Ilan. Astrophysics publications on arXiv, Scopus and Mendeley: a case study. Scientometrics, 100(1):217– 225, July 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1215-1.

Bar-Ilan:2018:CLE

[BI18a] Judit Bar-Ilan. Comments on the Letter to the Editor on “Multiple versions of the h-index: cautionary use for formal academic purposes” by Jaime A. Teixera da Silva and Judit Dobránszki. Scientometrics, 115(2):1115–1117, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2681-2. See [dSD18b].

Bar-Ilan:2018:EGW

[BI18b] Judit Bar-Ilan. Eugene Garfield on the Web in 2001. Scientometrics, 114(2):389–399, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2590-9.

Bar-Ilan:2017:PRC

[BIH17] Judit Bar-Ilan and Gali Halevi. Post retraction citations in context: a case study. Scientometrics, 113(1):547–565, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2242-0;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2242-
0.pdf.

Bar-Ilan:2018:TCR

[BIH18] Judit Bar-Ilan and Gali Halevi. Temporal characteristics of retracted articles. Scientometrics, 116(3):1771– 1783, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2802-y.

Bar-Ilan:2015:RIV

[BIL15] Judit Bar-Ilan and Mark Levene. The hw-rank: an hindex variant for ranking web pages. Scientometrics, 102 (3):2247–2253, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1477-2.

Buchan:2016:GNM

[BJIB16] Alison M. J. Buchan, Eva Jurczyk, Ruth Isserlin, and Gary D. Bader. Global neuroscience and mental health research: a bibliometrics case study. Scientometrics, 109 (1):515–531, October 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-016-
2094-z.

Bjork:2019:AWN

[Bjø19] R. Bjørk. The age at which Nobel Prize research is conducted. Scientometrics, 119(2):931–939, May 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03065-4.

Bass:2010:DFI

[BK10] Scott D. Bass and Lukasz A. Kurgan. Discovery of factors influencing patent value based on machine learning in patents in the field of nanotechnology. Scientometrics, 82 (2):217–241, February 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0008-z.

Bartneck:2011:DIM

[BK11] Christoph Bartneck and Servaas Kokkelmans. Detecting hindex manipulation through self-citation analysis. Scientometrics, 87(1):85–98, April 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-010-
0306-5.pdf.

Batistic:2015:OSF

[BK15] Sasa Batistic and Robert Kase. The organizational socialization field fragmentation: a bibliometric review. Scientometrics, 104(1):121–146, July 2015. CODEN SCNTDX.

ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
015-1538-1.

Bellotti:2016:ERC

[BKG16] Elisa Bellotti, Luka Kronegger, and Luigi Guadalupi. The evolution of research collaboration within and across disciplines in Italian Academia. Scientometrics, 109(2):783– 811, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-016-2068-
1.pdf.

Beets:2015:AAJ

[BKL15] S. Douglas Beets, Andrea S. Kelton, and Bruce R. Lewis. An assessment of accounting journal quality based on departmental lists. Scientometrics, 102(1):315–332, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-014-1353-0.

Bigdeli:2013:PAI

[BKRG13] Zahed Bigdeli, Morteza Kokabi, Gholam Reza Rajabi, and Ali Gazni. Patterns of authors’ information scattering: towards a causal explanation of information scattering from a scholarly information-seeking behavior perspective. Scientometrics, 96(1):103–131, July 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0885-4.

Bhattacharya:2015:RBI

[BKSS15] Sujit Bhattacharya, Arshia Kaul, Shilpa, and Praveen Sharma. Role of bilateral institution in influencing collaboration: case study of CEFIPRA — a bilateral S&T institution established by India and France. Scientometrics, 102 (1):169–194, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1454-9.

Baydas:2015:ETR

[BKY^{+}15] Ozlem Baydas, Sevda Kucuk, Rabia Meryem Yilmaz, Melike Aydemir, and Yuksel Goktas. Educational technology research trends from 2002 to 2014. Scientometrics, 105(1): 709–725, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1693-4.

Babic:2016:EQS

[BKZ^{+}16] Dragan Babić, Duro Kutlaca, Lazar Zivković, Dijana Strbac, and Dusica Semencenko. Evaluation of the quality of scientific performance of the selected countries of Southeast Europe. Scientometrics, 106(1):405–434, January 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1649-8.

Bador:2010:CAB

[BL10] Pascal Bador and Thierry Lafouge. Comparative analysis between impact factor and h-index for pharmacology and psychiatry journals. Scientometrics, 84(1):65–79, July 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-009-0058-2.

Behrens:2011:MBA

[BL11a] Heinrich Behrens and Peter Luksch. Mathematics 1868– 2008: a bibliometric analysis. Scientometrics, 86(1):179– 194, January 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0249-x.

Breimer:2011:LCS

[BL11b] Lars H. Breimer and Janeth Leksell. Longitudinal and cross-sectional study of registered nurses in Sweden who undertake a PhD showing that nurses continue to publish in English after their PhD but male nurses are more productive than female nurses. Scientometrics, 87(2):337–345, May 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0325-2.

Bouabid:2013:LPL

[BL13] Hamid Bouabid and Vincent Larivière. The lengthening of papers’ life expectancy: a diachronous analysis. Scientometrics, 97(3):695–717, December 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0995-7.

Bornmann:2015:TCB

[BL15] Lutz Bornmann and Loet Leydesdorff. Topical connections between the institutions within an organisation (institutional co-authorships, direct citation links and cocitations). Scientometrics, 102(1):455–463, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-014-1425-1.

Bakare:2017:CCR

[BL17a] Victoria Bakare and Grant Lewison. Country overcitation ratios. Scientometrics, 113(2):1199–1207, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2490-z;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2490-

z.pdf. See editorial content [Ano18b].

Byrne:2017:SSB

[BL17b] Jennifer A. Byrne and Cyril Labbé. Striking similarities between publications from China describing single gene knockdown experiments in human cancer cell lines. Scientometrics, 110(3):1471–1493, March 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2209-6.

Bornmann:2018:CHC

[BL18] Lutz Bornmann and Loet Leydesdorff. Count highly-cited papers instead of papers with h citations: use normalized citation counts and compare “like with like”! Scientometrics, 115(2):1119–1123, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2682-1;   http://link.springer.com/content/pdf/
10.1007/s11192-018-2682-1.pdf.

Benevenuto:2016:HIP

[BLA16] Fabrício Benevenuto, Alberto H. F. Laender, and Bruno L. Alves. The H-index paradox: your coauthors have a higher H-index than you do. Scientometrics, 106(1):469–474, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1776-2.

Barbosa:2017:AIC

[BLdlCV17] Marcelo Werneck Barbosa, Marcelo Bronzo Ladeira, and Alberto de la Calle Vicente. An analysis of international coauthorship networks in the supply chain analytics research area. Scientometrics, 111(3):1703–1731, June 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Blagus:2015:CBM

[BLS15] Rok Blagus, Brane L. Leskosek, and Janez Stare. Comparison of bibliometric measures for assessing relative importance of researchers. Scientometrics, 105(3):1743–1762, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1622-6.

Breschi:2011:AST

[BM11] Stefano Breschi and Franco Malerba. Assessing the scientific and technological output of EU framework programmes: evidence from the FP6 projects in the ICT field. Scientometrics, 88(1):239–257, July 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0378-x.

Bornmann:2012:ESV

[BM12a] Lutz Bornmann and Werner Marx. The effect of several versions of one and the same manuscript published by a journal on its journal impact factor. Scientometrics, 92(2): 277–279, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0656-2.

Bosnjak:2012:PPA

[BM12b] Lana Bosnjak and Ana Marusić. Prescribed practices of authorship: review of codes of ethics from professional bodies and journal guidelines across disciplines. Scientometrics, 93(3):751–763, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0773-y.

Benson:2013:EHK

[BM13a] Christopher L. Benson and Christopher L. Magee. Erratum to: A hybrid keyword and patent class methodology for selecting relevant sets of patents for a technological field. Scientometrics, 96(1):83, July 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-013-1043-3.pdf. See [BM13b].

Benson:2013:HKP

[BM13b] Christopher L. Benson and Christopher L. Magee. A hybrid keyword and patent class methodology for selecting relevant sets of patents for a technological field. Scientometrics, 96(1):69–82, July 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-

012-0930-3. See erratum [BM13a].

Bornmann:2014:HSS

[BM14a] Lutz Bornmann and Werner Marx. How should the societal impact of research be generated and measured? A proposal for a simple and practicable approach to allow interdisciplinary comparisons. Scientometrics, 98(1):211–219, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1020-x.

Bornmann:2014:HEI

[BM14b] Lutz Bornmann and Werner Marx. How to evaluate individual researchers working in the natural and life sciences meaningfully? A proposal of methods based on percentiles of citations. Scientometrics, 98(1):487–509, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1161-y.

Benson:2015:TSI

[BM15] Christopher L. Benson and Christopher L. Magee. Technology structural implications from the extension of a patent search method. Scientometrics, 102(3):1965–1985, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1493-2.

Bornmann:2019:HCL

[BM19] Lutz Bornmann and Julian N. Marewski. Heuristics as conceptual lens for understanding and studying the usage of bibliometrics in research evaluation. Scientometrics, 120(2):419–459, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03018-x;   http://link.springer.com/content/
pdf/10.1007/s11192-019-03018-x.pdf.

Bu:2018:MSS

[BMD^{+}18] Yi Bu, Dakota S. Murray, Ying Ding, Yong Huang, and Yiming Zhao. Measuring the stability of scientific collaboration. Scientometrics, 114(2):463–479, February 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2599-0.

Baghele:2014:ICI

[BMM14] Om N. Baghele, Abhijeet S. Mohkhedkar, and Pooja S. Malpani. Intellectual contribution of Indian periodontists to world literature: a bibliometric evaluation of Pubmed database till 1st March, 2012. Scientometrics, 99(3): 999–1010, June 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1196-0.

Basu:2017:EIA

[BMM17] Tapasree Basu, Ajoy Mallik, and Nripendranath Mandal. Evolving importance of anticancer research using herbal medicine: a scientometric analysis. Scientometrics, 110(3):1375–1396, March 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2223-8.

Bojovic:2014:OFJ

[BMP^{+}14] Srdan Bojović, Rada Matić, Zorica Popović, Miroslava Smiljanić, Milena Stefanović, and Vera Vidaković. An overview of forestry journals in the period 2006–2010 as basis for ascertaining research trends. Scientometrics, 98(2):1331–1346, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1171-9.

Brouthers:2012:BHI

[BMR12] Keith D. Brouthers, Ram Mudambi, and David M. Reeb. The blockbuster hypothesis: influencing the boundaries of knowledge. Scientometrics, 90(3):959–982, March 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0540-5.

Bonilla:2015:ELA

[BMTA15] Claudio A. Bonilla, José M. Merigó, and Carolina Torres-Abad. Economics in Latin America: a bibliometric analysis. Scientometrics, 105(2):1239–1252, November 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1747-7.

Brunswicker:2017:CID

[BMZ^{+}17] Sabine Brunswicker, Sorin Adam Matei, Michael Zentner, Lynn Zentner, and Gerhard Klimeck. Creating impact in the digital space: digital practice dependency in communities of digital scientific innovations. Scientometrics, 110(1):417–442, January 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2106-z.

Breimer:2010:LCS

[BN10] Lars H. Breimer and Torbjörn K. Nilsson. A longitudinal and cross-sectional study of Swedish biomedical PhD processes 1991–2009 with emphasis on international and gender aspects. Scientometrics, 85(2):401–414, November 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0227-3.

Breimer:2014:CAE

[BN14] Lars H. Breimer and Torbjörn K. Nilsson. Considerations for appointing an external examiner of a PhD in the biomedical sciences in Sweden: a questionnaire-based survey. Scientometrics, 98(3):2039–2049, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1150-1.

Bornmann:2011:ETS

[BND11] Lutz Bornmann, Christoph Neuhaus, and Hans-Dieter Daniel. The effect of a two-stage publication process on the Journal Impact Factor: a case study on the interactive open access journal atmospheric chemistry and physics. Scientometrics, 86(1):93–97, January 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0250-4.

Buter:2011:SCR

[BNV11] Reindert K. Buter, Ed. C. M. Noyons, and Anthony F. J. Van Raan. Searching for converging research using field to field citations. Scientometrics, 86(2):325–338, February 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-010-0246-0.pdf.

Bornmann:2011:MS

[Bor11] Lutz Bornmann. Mimicry in science? Scientometrics, 86 (1):173–177, January 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0222-8.

Bornmann:2012:HEJ

[Bor12] Lutz Bornmann. The Hawthorne effect in journal peer review. Scientometrics, 91(3):857–862, June 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0547-y.

Bornmann:2014:RIH

[Bor14] Lutz Bornmann. Ranking institutions by the handicap principle. Scientometrics, 100(2):603–604, August 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-014-1260-4.

Bornmann:2015:AMS

[Bor15a] Lutz Bornmann. Alternative metrics in scientometrics: a meta-analysis of research into three altmetrics. Scientometrics, 103(3):1123–1144, June 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
015-1565-y.

Bornmann:2015:LEC

[Bor15b] Lutz Bornmann. Letter to the Editor: On the conceptualisation and theorisation of the impact caused by publications. Scientometrics, 103(3):1145–1148, June 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1588-4.

Bornmann:2016:PRR

[Bor16] Lutz Bornmann. Is the promotion of research reflected in bibliometric data? A network analysis of highly cited papers on the Clusters of Excellence supported under the Excellence Initiative in Germany. Scientometrics, 107(3):1041–1061, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com//content/
pdf/10.1007/s11192-016-1925-2.pdf;   http://link.
springer.com/content/pdf/10.1007/s11192-016-1925-
2.pdf.

Bornmann:2017:CIJ

[Bor17] Lutz Bornmann. Confidence intervals for Journal Impact Factors. Scientometrics, 111(3):1869–1871, June 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). See comment [Ste17].

Bornmann:2018:FCP

[Bor18] Lutz Bornmann. Field classification of publications in Dimensions: a first case study testing its reliability and validity. Scientometrics, 117(1):637–640, October 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2855-y. See response [HL18].

Boshoff:2010:SSR

[Bos10] Nelius Boshoff. South–South research collaboration of countries in the Southern African Development Community (SADC). Scientometrics, 84(2):481–503, August 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0120-0.

Bjork:2014:TSC

[BOS14] Samuel Bjork, Avner Offer, and Gabriel Söderberg. Time series citation data: the Nobel Prize in Economics. Scientometrics, 98(1):185–196, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0989-5.

Bouabid:2011:RCA

[Bou11] Hamid Bouabid. Revisiting citation aging: a model for citation distribution and life-cycle prediction. Scientometrics, 88(1):199–211, July 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0370-5.

Bouabid:2014:STM

[Bou14a] Hamid Bouabid. Science and technology metrics for research policy evaluation: some insights from a Moroccan experience. Scientometrics, 101(1):899–915, October 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1407-3.

Bougrine:2014:SEC

[Bou14b] Hassan Bougrine. Subfield effects on the core of coauthors. Scientometrics, 98(2):1047–1064, February 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1066-9.

Boyack:2017:IEG

[Boy17a] Kevin W. Boyack. Investigating the effect of global data on topic detection. Scientometrics, 111(2):999–1015, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Boyack:2017:TBM

[Boy17b] Kevin W. Boyack. Thesaurus-based methods for mapping contents of publication sets. Scientometrics, 111(2):1141– 1155, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Bjurstrom:2011:CCI

[BP11] Andreas Bjurström and Merritt Polk. Climate change and interdisciplinarity: a co-citation analysis of IPCC third assessment report. Scientometrics, 87(3):525–550, June 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0356-3.

Benavent-Perez:2012:DFR

[BPGGdMA12] Maria Benavent-Pérez, Juan Gorraiz, Christian Gumpenberger, and Félix de Moya-Anegón. The different flavors of research collaboration: a case study of their influence on university excellence in four world regions. Scientometrics, 93(1):41–58, October 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0638-4.

Bouabid:2016:SCH

[BPHL16] Hamid Bouabid, Adèle Paul-Hus, and Vincent Larivière. Scientific collaboration and high-technology exchanges among BRICS and G-7 countries. Scientometrics, 106(3): 873–899, March 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1806-0.

Barnett:2014:MLN

[BPJ^{+}14] George A. Barnett, Han Woo Park, Ke Jiang, Chuan Tang, and Isidro F. Aguillo. A multi-level network analysis of webcitations among the world’s universities. Scientometrics, 99(1):5–26, April 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1070-0.

Bolanos-Pizarro:2010:CRS

[BPTG10] Máxima Bolaños-Pizarro, Bart Thijs, and Wolfgang Glänzel. Cardiovascular research in Spain. A comparative scientometric study. Scientometrics, 85(2):509–526, November 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0155-2.

Bosnjak:2011:ANT

[BPVM11] Lana Bosnjak, Livia Puljak, Katarina Vukojević, and Ana Marusić. Analysis of a number and type of publications that editors publish in their own journals: case study of scholarly journals in Croatia. Scientometrics, 86(1):227–233, January 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0207-7.

Benito:2011:IQA

[BR11] M. Benito and R. Romera. Improving quality assessment of composite indicators in university rankings: a case study of French and German universities of excellence. Scientometrics, 89(1):153–176, October 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0419-5.

Benito:2012:MED

[BR12] Mónica Benito and Rosario Romera. Modeling the enrollment demand of masters programs for the Spanish public university system. Scientometrics, 91(1):113–130, April 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0593-5.

Braun:2010:Ea

[Bra10] Tibor Braun. Editorial. Scientometrics, 82(1):1–2, January 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-009-0151-6.pdf.

Braun:2012:EWS

[Bra12a] Jasna Dravec Braun. Effects of war on scientific production: mathematics in Croatia from 1968 to 2008. Scientometrics, 93(3):931–936, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0735-4.

Braun:2012:E

[Bra12b] Tibor Braun. Editorial. Scientometrics, 92(2):207–208, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-012-0754-1.pdf.

Breimer:2010:SBP

[Bre10] Lars H. Breimer. Swedish biomedical PhD examination: an international forum and a proposed procedure for Europe. Scientometrics, 83(2):583–587, May 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0081-3.

Breimer:2013:IFE

[Bre13] Lars H. Breimer. Impact of foreign external PhD examiners on academic collaboration patterns. Scientometrics, 96 (1):315–322, July 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0931-2.

Brody:2013:IFI

[Bro13] Stuart Brody. Impact factor: Imperfect but not yet replaceable. Scientometrics, 96(1):255–257, July 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0863-x.

Bruck:2016:RET

[BRS^{+}16] Péter Bruck, István Réthy, Judit Szente, Jan Tobochnik, and Péter Érdi. Recognition of emerging technology trends: class-selective study of citations in the U.S. Patent Citation Network. Scientometrics, 107(3):1465–1475, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-016-1899-0.

Bruer:2010:CWT

[Bru10] John T. Bruer. Can we talk? How the cognitive neuroscience of attention emerged from neurobiology and psychology, 1980–2005. Scientometrics, 83(3):751–764, June 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-009-0128-5.

Brzezinski:2015:PLC

[Brz15] Michal Brzezinski. Power laws in citation distributions: evidence from Scopus. Scientometrics, 103(1):213–228, April 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-014-1524-z.pdf.

Borsi:2011:ARE

[BS11] Balázs Borsi and András Schubert. Agrifood research in Europe: a global perspective. Scientometrics, 86(1):133–154, January 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0235-3.

Bonitz:2013:RMB

[BS13a] Michael Bonitz and Andrea Scharnhorst. Remembering Manfred Bonitz (7.3.1931–14.8.2012) on the first anniversary of his death. Scientometrics, 97(1):121–128, October 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1056-y.

Brandt:2013:UMO

[BS13b] Tasso Brandt and Torben Schubert. Is the university model an organizational necessity? Scale and agglomeration effects in science. Scientometrics, 94(2):541–565, February 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0834-2.

Bartol:2015:NLD

[BS15a] Tomaz Bartol and Karmen Stopar. Nano language and distribution of article title terms according to power laws. Scientometrics, 103(2):435–451, May 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1546-1.

Bjork:2015:APC

[BS15b] Bo-Christer Björk and David Solomon. Article processing charges in OA journals: relationship between price and quality. Scientometrics, 103(2):373–385, May 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1556-z.

Bolli:2015:JMP

[BS15c] Thomas Bolli and Jörg Schläpfer. Job mobility, peer effects, and research productivity in economics. Scientometrics, 104(3):629–650, September 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1625-3.

Bruns:2016:RAU

[BS16] Stephan B. Bruns and David I. Stern. Research assessment using early citation information. Scientometrics, 108(2): 917–935, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1979-1.

Bonaccorsi:2017:DRP

[BS17] Andrea Bonaccorsi and Luca Secondi. The determinants of research performance in European universities: a large scale multilevel analysis. Scientometrics, 112(3):1147– 1178, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2442-7.

Burmaoglu:2019:EAI

[BS19] Serhat Burmaoglu and Ozcan Saritas. An evolutionary analysis of the innovation policy domain: Is there a paradigm shift? Scientometrics, 118(3):823–847, March 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03014-1.

Bhattacharya:2012:CIT

[BSB12] Sujit Bhattacharya, Shilpa, and Madhulika Bhati. China and India: The two new players in the nanotechnology race. Scientometrics, 93(1):59–87, October 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0651-7.

Borner:2018:SMM

[BSBG18] Katy Börner, Adam H. Simpson, Andreas Bueckle, and Robert L. Goldstone. Science map metaphors: a comparison of network versus hexmap-based visualizations. Scientometrics, 114(2):409–426, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2596-3.

Bakhtin:2017:TML

[BSC^{+}17] Pavel Bakhtin, Ozcan Saritas, Alexander Chulok, Ilya Kuzminov, and Anton Timofeev. Trend monitoring for linking science and strategy. Scientometrics, 111(3):2059–2075, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Bin:2015:WDD

[BSFCC15] Adriana Bin, Sergio Salles-Filho, Luiza Maria Capanema, and Fernando Antonio Basile Colugnati. What difference does it make? Impact of peer-reviewed scholarships on scientific production. Scientometrics, 102(2):1167–1188, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1462-9.

Bookstein:2010:TMN

[BSFW10] Fred L. Bookstein, Horst Seidler, Martin Fieder, and Georg Winckler. Too much noise in the times higher education rankings. Scientometrics, 85(1):295–299, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s11192-010-0189-5.pdf.

Braun:2010:Eb

[BSG10] Tibor Braun, András Schubert, and Wolfgang Glänzel. Editorial. Scientometrics, 82(1):3, January 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-009-0152-5.pdf.

Barbosa:2017:WPM

[BSG17] Simone Diniz Junqueira Barbosa, Milene Selbach Silveira, and Isabela Gasparini. What publications metadata tell us about the evolution of a scientific community: the case of the Brazilian human–computer interaction conference series. Scientometrics, 110(1):275–300, January 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2162-4.

Bhattacharya:2015:ECA

[BSK15] Sujit Bhattacharya, Shilpa, and Arshia Kaul. Emerging countries assertion in the global publication landscape of science: a case study of India. Scientometrics, 103(2):387– 411, May 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1551-4.

Boyack:2018:TPR

[BSK18] Kevin W. Boyack, Caleb Smith, and Richard Klavans. Toward predicting research proposal success. Scientometrics, 114(2):449–461, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
017-2609-2.

Burmaoglu:2017:ECH

[BSKB17] Serhat Burmaoglu, Ozcan Saritas, Levent Bekir Kidak, and Ipek Camuz Berber. Evolution of connected health: a network perspective. Scientometrics, 112(3):1419– 1438, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2431-x.

Bornmann:2011:DIA

[BSMD11] Lutz Bornmann, Hermann Schier, Werner Marx, and Hans-Dieter Daniel. Does the h index for assessing single publications really work? A case study on papers published in chemistry. Scientometrics, 89(3):835–843, December 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0472-0.

Burmaoglu:2019:ATR

[BSPL19] Serhat Burmaoglu, Olivier Sartenaer, Alan Porter, and Munan Li. Analysing the theoretical roots of technology emergence: an evolutionary perspective. Scientometrics, 119 (1):97–118, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03033-y.

Bergiante:2015:BSR

[BSS15] Níssia C. R. Bergiante, Marcio P. S. Santos, and Respício A. Espírito Santo, Jr. Bibliometric study of the relationship between business model and air transport. Scientometrics, 105(2):941–958, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1711-6.

Bauer:2013:HDS

[BSvEK13] Hans P. W. Bauer, Gabriel Schui, Alexander von Eye, and Günter Krampen. How does scientific success relate to individual and organizational characteristics? A scientometric study of psychology researchers in the German-speaking countries. Scientometrics, 94(2):523–539, February 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0760-3.

Breitzman:2015:ITS

[BT15] Anthony Breitzman and Patrick Thomas. Inventor team size as a predictor of the future citation impact of patents. Scientometrics, 103(2):631–647, May 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1550-5.

Bihari:2017:EIN

[BT17] Anand Bihari and Sudhakar Tripathi. EM-index: a new measure to evaluate the scientific impact of scientists. Scientometrics, 112(1):659–677, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). See correction [BT18a].

Bihari:2018:CEI

[BT18a] Anand Bihari and Sudhakar Tripathi. Correction to: EM-index: a new measure to evaluate the scientific impact of scientists. Scientometrics, 117(2):1315, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2868-6;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2868-

6.pdf. See [BT17].

Bihari:2018:CYB

[BT18b] Anand Bihari and Sudhakar Tripathi. Correction to: Year based EM-index: a new approach to evaluate the scientific impact of scholars. Scientometrics, 117(2):1317, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2869-5;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2869-

5.pdf. See [BT18c].

Bihari:2018:YBE

[BT18c] Anand Bihari and Sudhakar Tripathi. Year based EM-index: a new approach to evaluate the scientific impact of scholars. Scientometrics, 114(3):1175–1205, March 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2625-2. See correction [BT18b].

Bornmann:2019:DPP

[BT19a] Lutz Bornmann and Alexander Tekles. Disruptive papers published in Scientometrics. Scientometrics, 120(1):331– 336, July 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03113-z.

Bornmann:2019:PDE

[BT19b] Lutz Bornmann and Alexander Tekles. Productivity does not equal usefulness. Scientometrics, 118(2):705–707, February 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2982-5.

Bornmann:2019:HWD

[BTL19] Lutz Bornmann, Alexander Tekles, and Loet Leydesdorff. How well does I3 perform for impact measurement compared to other bibliometric indicators? The convergent validity of several (field-normalized) indicators. Scientometrics, 119(2):1187–1205, May 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03071-6;   http://link.springer.
com/content/pdf/10.1007/s11192-019-03071-6.pdf.

Bowman:2014:PIF

[BTNS14] Timothy D. Bowman, Andrew Tsou, Chaoqun Ni, and Cassidy R. Sugimoto. Post-interdisciplinary frames of reference: exploring permeability and perceptions of disciplinarity in the social sciences. Scientometrics, 101(3):1695–1714, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1338-z.

Bueno:2015:OLI

[Bue15] Newton Paulo Bueno. Are opinion leaders important to spread information to cope with extreme droughts in (all) irrigation systems? A network analysis. Scientometrics, 105(2):817–824, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1734-z.

Burrell:2012:CPI

[Bur12] Quentin L. Burrell. Comments on “A publication index that is independent of age” by Abt. Scientometrics, 91(3):1059–1060, June 2012. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-011-0605-5. See [Abt12].

Burrell:2014:IAP

[Bur14] Quentin L. Burrell. The individual author’s publicationcitation process: theory and practice. Scientometrics, 98 (1):725–742, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1018-4.

Barrios:2013:SPP

[BVB13] Maite Barrios, Anna Villarroya, and Ángel Borrego. Scientific production in psychology: a gender analysis. Scientometrics, 95(1):15–23, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0816-4.

Braam:2014:IDR

[BvdB14] Robert Braam and Peter van den Besselaar. Indicators for the dynamics of research organizations: a biomedical case study. Scientometrics, 99(3):949–971, June 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1235-5.

Banadkouki:2018:RIU

[BVOL18] Mohammad Reza Zare Banadkouki, Mohammad Ali Vahdatzad, Mohammad Saleh Owlia, and Mohammad Mahdi Lotfi. Ranking Iranian universities: an interpretative structural modeling approach. Scientometrics, 117(3):1493– 1512, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2946-9.

Bakker:2016:PCI

[BVZV16] Jurriën Bakker, Dennis Verhoeven, Lin Zhang, and Bart Van Looy. Patent citation indicators: One size fits all? Scientometrics, 106(1):187–211, January 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1786-0.

Boell:2010:JIF

[BW10] Sebastian K. Boell and Concepción S. Wilson. Journal Impact Factors for evaluating scientific performance: use of hlike indicators. Scientometrics, 82(3):613–626, March 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0175-y.

Bornmann:2019:NCI

[BW19] Lutz Bornmann and Klaus Wohlrabe. Normalisation of citation impact in economics. Scientometrics, 120(2):841–884, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03140-w;   http://
link.springer.com/content/pdf/10.1007/s11192-019-

03140-w.pdf. See correction [BW20].

Bornmann:2020:CNC

[BW20] Lutz Bornmann and Klaus Wohlrabe. Correction to: Normalisation of citation impact in economics. Scientometrics, 123(2):1167, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03255-0;   http://link.springer.com/content/
pdf/10.1007/s11192-019-03255-0.pdf. See [BW19].

Bu:2018:UAC

[BWbH^{+}18] Yi Bu, Binglu Wang, Win bin Huang, Shangkun Che, and Yong Huang. Using the appearance of citations in full text on author co-citation analysis. Scientometrics, 116(1):275– 289, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2757-z.

Bornmann:2010:CAR

[BWD10] Lutz Bornmann, Christophe Weymuth, and Hans-Dieter Daniel. A content analysis of referees’ comments: how do comments on manuscripts rejected by a high-impact journal and later published in either a low- or high-impact journal differ? Scientometrics, 83(2):493–506, May 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0011-4.

Bornmann:2012:CVO

[BWD12] Lutz Bornmann, Markus Wolf, and Hans-Dieter Daniel. Closed versus open reviewing of journal manuscripts: how far do comments differ in language use? Scientometrics, 91(3):843–856, June 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0569-5.

Bornmann:2017:CES

[BWdMA17] Lutz Bornmann, Klaus Wohlrabe, and Felix de Moya Anegon. Calculating the excellence shift: How efficiently do institutions produce highly cited papers? Scientometrics, 112(3):1859–1864, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2446-3.

Bajwa:2013:BAB

[BY13] R. S. Bajwa and K. Yaldram. Bibliometric analysis of biotechnology research in Pakistan. Scientometrics, 95 (2):529–540, May 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0839-x.

Bu:2016:MMA

[ByLbH16] Yi Bu, Tian yi Liu, and Win bin Huang. MACA: a modified author co-citation analysis method combined with general descriptive metadata of citations. Scientometrics, 108 (1):143–166, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1959-5.

Bajwa:2013:SAR

[BYR13] R. S. Bajwa, K. Yaldram, and S. Rafique. A scientometric assessment of research output in nanoscience and nanotechnology: Pakistan perspective. Scientometrics, 94(1): 333–342, January 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0699-4.

Bornmann:2017:SAA

[BYY17] Lutz Bornmann, Adam Y. Ye, and Fred Y. Ye. Sequence analysis of annually normalized citation counts: an empirical analysis based on the characteristic scores and scales (CSS) method. Scientometrics, 113(3):1665–1680, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2521-9;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2521-
9.pdf.

Bornmann:2018:IHP

[BYY18] Lutz Bornmann, Adam Y. Ye, and Fred Y. Ye. Identifying “hot papers” and papers with “delayed recognition” in large-scale datasets by using dynamically normalized citation impact scores. Scientometrics, 116(2):655–674, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2772-0;   http://link.

springer.com/content/pdf/10.1007/s11192-018-2772-
0.pdf.

Bao:2017:DCA

[BZ17] Peng Bao and Chengxiang Zhai. Dynamic credit allocation in scientific literature. Scientometrics, 112(1):595–606, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Boukacem-Zeghmouri:2016:RBC

[BZBLP16] Chérifa Boukacem-Zeghmouri, Pascal Bador, Thierry Lafouge, and Hélène Prost. Relationships between consumption, publication and impact in French universities in a value perspective: a bibliometric analysis. Scientometrics, 106 (1):263–280, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1779-z.

Colliander:2012:ECF

[CA12] Cristian Colliander and Per Ahlgren. Experimental comparison of first and second-order similarities in a scientometric context. Scientometrics, 90(2):675–685, February 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0491-x.

Cansun:2018:PSP

[CA18] Sebnem Cansun and Engin Arik. Political science publications about Turkey. Scientometrics, 115(1):169–188, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2672-3.

Colliander:2019:CPL

[CA19] Cristian Colliander and Per Ahlgren. Comparison of publication-level approaches to ex-post citation normalization. Scientometrics, 120(1):283–300, July 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03121-z;   http://link.springer.
com/content/pdf/10.1007/s11192-019-03121-z.pdf.

Cakir:2015:CAG

[ÇAAÇ15] Murat Perit Çakir, Cengiz Acartürk, Oguzhan Alasehir, and Canan Çilingir. A comparative analysis of global and national university ranking systems. Scientometrics, 103 (3):813–848, June 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1586-6.

Cabanac:2011:AIR

[Cab11] Guillaume Cabanac. Accuracy of inter-researcher similarity measures based on topical and social clues. Scientometrics, 87(3):597–620, June 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0358-1.

Cabanac:2013:EPA

[Cab13] Guillaume Cabanac. Experimenting with the partnership ability ϕ-index on a million computer scientists. Scientometrics, 96(1):1–9, July 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0862-y.

Cabanac:2014:EQE

[Cab14] Guillaume Cabanac. Extracting and quantifying eponyms in full-text articles. Scientometrics, 98(3):1631–1645, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1091-8.

Cabanac:2018:WPR

[Cab18] Guillaume Cabanac. What is the primordial reference for … ? — redux. Scientometrics, 114(2):481–488, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2595-4.

Chen:2015:EIE

[CAGL15] Shiji Chen, Clément Arsenault, Yves Gingras, and Vincent Larivière. Exploring the interdisciplinary evolution of a discipline: the case of biochemistry and molecular biology. Scientometrics, 102(2):1307–1323, February 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1457-6.

Campanario:2010:DCI

[Cam10] Juan Miguel Campanario. Distribution of changes in impact factors over time. Scientometrics, 84(1):35–42, July 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-009-0094-y.

Campanario:2011:ESJ

[Cam11] Juan Miguel Campanario. Empirical study of journal impact factors obtained using the classical two-year citation window versus a five-year citation window. Scientometrics, 87(1):189–204, April 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0334-1.

Campanario:2012:SRI

[Cam12] Juan Miguel Campanario. Some research ideas on Journal Impact Factors as a crucial topic in science dynamics. Scientometrics, 92(2):293–295, August 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0650-8.

Campanario:2014:ECS

[Cam14] Juan Miguel Campanario. The effect of citations on the significance of decimal places in the computation of journal impact factors. Scientometrics, 99(2):289–298, May 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1206-2. See comments [Egg14a].

Campanario:2017:JPU

[Cam17] Juan Miguel Campanario. JIF-Plots: using plots of citations versus citable items as a tool to study journals and subject categories and discover new scientometric relationships. Scientometrics, 113(2):1141–1154, November 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11192-017-2506-8. See editorial content [Ano18b].

Campanario:2018:LRL

[Cam18] Juan Miguel Campanario. Are leaders really leading? Journals that are first in Web of Science subject categories in the context of their groups. Scientometrics, 115(1):111–130, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2660-7.

Carroll:2016:MAR

[Car16] Christopher Carroll. Measuring academic research impact: creating a citation profile using the conceptual framework for implementation fidelity as a case study. Scientometrics, 109(2):1329–1340, November 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2085-0.

Cuccurullo:2016:FTP

[CAS16] Corrado Cuccurullo, Massimo Aria, and Fabrizia Sarto. Foundations and trends in performance management. a twenty-five years bibliometric analysis in business and public administration domains. Scientometrics, 108(2):595–611, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1948-8.

Cavacini:2015:CBN

[Cav15a] Antonio Cavacini. Is the CE/ BCE notation becoming a standard in scholarly literature? Scientometrics, 102(2): 1661–1668, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1352-1.

Cavacini:2015:WBD

[Cav15b] Antonio Cavacini. What is the best database for computer science journal articles? Scientometrics, 102(3):2059–2071, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1506-1.

Cavacini:2016:RTM

[Cav16] Antonio Cavacini. Recent trends in Middle Eastern scientific production. Scientometrics, 109(1):423–432, October 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-1932-3.

Cools:2019:PUA

[CAV^{+}19] Evelien Cools, Julia Ausserer, Marc Van de Velde, Peter Hamm, and Peter Paal. Publications from universityaffiliated anaesthesiology departments: a look at Belgium, France and the Netherlands from 2001 to 2015. Scientometrics, 119(2):863–878, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03075-2.

Costas:2011:DAP

[CB11] Rodrigo Costas and María Bordons. Do age and professional rank influence the order of authorship in scientific publications? Some evidence from a micro-level perspective. Scientometrics, 88(1):145–161, July 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-011-0368-z.pdf.

Cerinsek:2015:NAB

[CB15] Monika Cerinsek and Vladimir Batagelj. Network analysis of Zentralblatt MATH data. Scientometrics, 102(1):977–1001, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1419-z.

Coccia:2016:AMM

[CB16] Mario Coccia and Barry Bozeman. Allometric models to measure and analyze the evolution of international research collaboration. Scientometrics, 108(3):1065–1084, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2027-x.

Copiello:2018:FRR

[CB18] Sergio Copiello and Pietro Bonifaci. A few remarks on ResearchGate score and academic reputation. Scientometrics, 114(1):301–306, January 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2582-9.

Copiello:2019:RSF

[CB19] Sergio Copiello and Pietro Bonifaci. ResearchGate Score, full-text research items, and full-text reads: a follow-up study. Scientometrics, 119(2):1255–1262, May 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03063-6.

Chen:2013:ECN

[CBF13] Yunwei Chen, Katy Börner, and Shu Fang. Evolving collaboration networks in Scientometrics in 1978–2010: a micro-macro analysis. Scientometrics, 95(3):1051–1070, June 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0895-2.

Chung:2013:ACT

[CBKL13] Chung Joo Chung, George A. Barnett, Kitae Kim, and Derek Lackaff. An analysis on communication theory and discipline. Scientometrics, 95(3):985–1002, June 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0869-4.

Calver:2018:QIM

[CBWJ18] Michael Calver, Kate Bryant, and Grant Wardell-Johnson. Quantifying the internationality and multidisciplinarity of authors and journals using ecological statistics. Scientometrics, 115(2):731–748, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2692-z.

Chen:2010:ANE

[CC10a] Yu-Shan Chen and Ke-Chiun Chang. Analyzing the nonlinear effects of firm size, profitability, and employee productivity on patent citations of the US pharmaceutical companies by using artificial neural network. Scientometrics, 82 (1):75–82, January 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0034-x.

Chen:2010:NNR

[CC10b] Yu-Shan Chen and Ke-Chiun Chang. The nonlinear nature of the relationships between the patent traits and corporate performance. Scientometrics, 82(1):201–210, January 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0101-3.

Campanario:2011:BLC

[CC11a] Juan Miguel Campanario and María Angeles Coslado. Benford’s Law and citations, articles and impact factors of scientific journals. Scientometrics, 88(2):421–432, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0387-9.

Claro:2011:MMI

[CC11b] João Claro and Carlos A. V. Costa. A made-to-measure indicator for cross-disciplinary bibliometric ranking of researchers performance. Scientometrics, 86(1):113–123, January 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0241-5.

Chen:2012:IQP

[CC12a] Jui-Kuei Chen and I-Shuo Chen. An Inno–Qual performance system for higher education. Scientometrics, 93(3): 1119–1149, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0759-9.

Chen:2012:UEB

[CC12b] Yu-Shan Chen and Ke-Chiun Chang. Using the entropybased patent measure to explore the influences of related and unrelated technological diversification upon technological competences and firm performance. Scientometrics, 90 (3):825–841, March 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0557-9.

Cho:2013:PPA

[CC13] Young-Don Cho and Hoo-Gon Choi. Principal parameters affecting R&D exploitation of nanotechnology research: a case for Korea. Scientometrics, 96(3):881–899, September 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-0974-z.

Campanario:2014:EAC

[CC14] Juan Miguel Campanario and William Cabos. The effect of additional citations in the stability of Journal Citation Report categories. Scientometrics, 98(2):1113–1130, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1116-3.

Chen:2016:WFP

[CC16] Chi-Tung Chen and Dar-Zen Chen. Who files provisional applications in the United States? Scientometrics, 107 (2):555–568, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1855-z.

Cheang:2014:OMJ

[CCLL14] Brenda Cheang, Samuel Kai Wah Chu, Chongshou Li, and Andrew Lim. OR/MS journals evaluation based on a refined PageRank method: an updated and more comprehensive review. Scientometrics, 100(2):339–361, August 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1272-0.

Campanario:2011:EJI

[CCM^{+}11] Juan Miguel Campanario, Jesús Carretero, Vera Marangon, Antonio Molina, and Germán Ros. Effect on the journal impact factor of the number and document type of citing records: a wide-scale study. Scientometrics, 87(1):75–84, April 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0333-2.

Capellari:2014:UOU

[CD14] Saveria Capellari and Domenico De Stefano. Universityowned and university-invented patents: a network analysis on two Italian universities. Scientometrics, 99(2):313–329, May 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1211-5.

Calma:2016:AMJ

[CD16a] Angelito Calma and Martin Davies. Academy of Management Journal, 1958–2014: a citation analysis. Scientometrics, 108(2):959–975, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
016-1998-y.

Castriotta:2016:DAT

[CD16b] Manuel Castriotta and Maria Chiara Di Guardo. Disentangling the automotive technology structure: a patent co-citation analysis. Scientometrics, 107(2):819–837, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1862-0.

Calma:2017:GIC

[CD17] Angelito Calma and Martin Davies. Geographies of influence: a citation network analysis of higher education 1972– 2014. Scientometrics, 110(3):1579–1599, March 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2228-3.

Cheng:2018:KCS

[CD18] Ni Cheng and Ke Dong. Knowledge communication on social media: a case study of biomedical science on Baidu Baike. Scientometrics, 116(3):1749–1770, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2828-1.

Cho:2013:IRF

[CDCK13] Philip S. Cho, Huy Hoang Nhat Do, Muthu Kumar Chandrasekaran, and Min-Yen Kan. Identifying research facilitators in an emerging Asian research area. Scientometrics, 97(1):75–97, October 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1051-3.

Clermont:2015:RSB

[CDD15] Marcel Clermont, Alexander Dirksen, and Harald Dyckhoff. Returns to scale of business administration research in Germany. Scientometrics, 103(2):583–614, May 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1561-2.

Coutinho:2012:BSP

[CDdS^{+}12] Renato X. Coutinho, Eliziane S. Dávila, Wendel M. dos Santos, João B. T. Rocha, Diogo O. G. Souza, Vanderlei Folmer, and Robson L. Puntel. Brazilian scientific production in science education. Scientometrics, 92(3):697–710, September 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0645-5.

Chimhundu:2015:SCN

[CdJD15] Chipo Chimhundu, Kylie de Jager, and Tania Douglas. Sectoral collaboration networks for cardiovascular medical device development in South Africa. Scientometrics, 105(3): 1721–1741, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1743-y.

Chen:2018:SWS

[CDM18] Baitong Chen, Ying Ding, and Feicheng Ma. Semantic word shifts in a scientific domain. Scientometrics, 117(1):211–226, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2843-2.

Costa:2017:BAA

[CdMCdMMdP17] Daniel Fonseca Costa, Francisval de Melo Carvalho, Bruno César de Melo Moreira, and José Willer do Prado. Bibliometric analysis on the association between behavioral finance and decision making with cognitive biases such as overconfidence, anchoring effect and confirmation bias. Scientometrics, 111(3):1775–1799, June 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Costa:2013:SCB

[CdSPdM13] Benedita Marta Gomes Costa, Edilson da Silva Pedro, and Gorete Ribeiro de Macedo. Scientific collaboration in biotechnology: the case of the northeast region in Brazil. Scientometrics, 95(2):571–592, May 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0924-1.

Chen:2014:MEP

[CF14] Yunwei Chen and Shu Fang. Mapping the evolving patterns of patent assignees’ collaboration networks and identifying the collaboration potential. Scientometrics, 101(2): 1215–1231, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1304-9.

Costas:2018:RAC

[CF18] Rodrigo Costas and Thomas Franssen. Reflections around ‘the cautionary use’ of the h-index: response to Teixeira da Silva and Dobránszki. Scientometrics, 115(2):1125–1130, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2683-0;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2683-

0.pdf. See [dSD18b].

Colugnati:2014:PSA

[CFdC^{+}14] Fernando A. B. Colugnati, Sergio Firpo, Paula F. Drummond de Castro, Juan E. Sepulveda, and Sergio L. M. Salles-Filho. A propensity score approach in the impact evaluation on scientific production in Brazilian biodiversity research: the BIOTA program. Scientometrics, 101(1):85–107, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1397-1.

Chan:2014:DBS

[CFG^{+}14] Ho Fai Chan, Bruno S. Frey, Jana Gallus, Markus Schaffner, Benno Torgler, and Stephen Whyte. Do the best scholars attract the highest speaking fees? An exploration of internal and external influence. Scientometrics, 101(1):793–817, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1379-3.

Cugmas:2016:SCA

[CFK16] Marjan Cugmas, Anuska Ferligoj, and Luka Kronegger. The stability of co-authorship structures. Scientometrics, 106 (1):163–186, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1790-4.

Carayol:2012:RCT

[CFL12] Nicolas Carayol, Ghislaine Filliatreau, and Agenor Lahatte. Reference classes: a tool for benchmarking universities’ research. Scientometrics, 93(2):351–371, November 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0672-2.

Coccia:2015:IHP

[CFM15] Mario Coccia, Greta Falavigna, and Alessandro Manello. The impact of hybrid public and market-oriented financing mechanisms on the scientific portfolio and performances of public research labs: a scientometric analysis. Scientometrics, 102(1):151–168, January 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1427-z.

Cabanac:2018:BEI

[CFM18] Guillaume Cabanac, Ingo Frommholz, and Philipp Mayr. Bibliometric-enhanced information retrieval: preface. Scientometrics, 116(2):1225–1227, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2861-0;   http://link.springer.com/
content/pdf/10.1007/s11192-018-2861-0.pdf.

Cechlarova:2014:AER

[CFP14] Katarína Cechlárová, Tamás Fleiner, and Eva Potpinková. Assigning evaluators to research grant applications: the case of Slovak Research and Development Agency. Scientometrics, 99(2):495–506, May 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
013-1187-1.

Cislak:2018:BAR

[CFS18] Aleksandra Cislak, Magdalena Formanowicz, and Tamar Saguy. Bias against research on gender bias. Scientometrics, 115(1):189–200, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2667-0;
http://link.springer.com/content/pdf/10.1007/s11192-
018-2667-0.pdf.

Calabuig:2016:VVI

[CFSSP16] J. M. Calabuig, A. Ferrer-Sapena, and E. A. Sánchez-Pérez. Vector-valued impact measures and generation of specific indexes for research assessment. Scientometrics, 108(3):1425–1443, September 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2039-6.

Chen:2011:MBP

[CG11] Zifeng Chen and Jiancheng Guan. Mapping of biotechnology patents of China from 1995–2008. Scientometrics, 88(1):73– 89, July 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0380-3.

Cerchiello:2014:SI

[CG14] Paola Cerchiello and Paolo Giudici. On a statistical h index. Scientometrics, 99(2):299–312, May 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1194-2.

Campbell:2015:RDA

[CG15a] Nick Campbell and Michelle Grayson. A response to ‘Discussion about the new Nature Index’. Scientometrics, 102(2):1831–1833, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-

014-1516-z. See [HB15].

Confraria:2015:IAS

[CG15b] Hugo Confraria and Manuel Mira Godinho. The impact of African science: a bibliometric analysis. Scientometrics, 102(2):1241–1268, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1463-8.

Chi:2017:EIA

[CG17] Pei-Shan Chi and Wolfgang Glänzel. An empirical investigation of the associations among usage, scientific collaboration and citation impact. Scientometrics, 112(1):403–412, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Chi:2018:CCU

[CG18a] Pei-Shan Chi and Wolfgang Glänzel. Comparison of citation and usage indicators in research assessment in scientific disciplines and journals. Scientometrics, 116(1):537– 554, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2708-8.

Cunillera:2018:TYS

[CG18b] Toni Cunillera and Georgina Guilera. Twenty years of statistical learning: from language, back to machine learning. Scientometrics, 117(1):1–8, October 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2856-x.

Ceballos:2018:FIF

[CGC18] Hector G. Ceballos, Sara E. Garza, and Francisco J. Cantu. Factors influencing the formation of intra-institutional formal research groups: group prediction from collaboration, organisational, and topical networks. Scientometrics, 114 (1):181–216, January 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2561-1.

Casnici:2017:APR

[CGG^{+}17] Niccolò Casnici, Francisco Grimaldo, Nigel Gilbert, Pierpaolo Dondio, and Flaminio Squazzoni. Assessing peer review by gauging the fate of rejected manuscripts: the case of the Journal of Artificial Societies and Social Simulation. Scientometrics, 113(1):533–546, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2241-1;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2241-
1.pdf.

Chi:2019:CCU

[CGG19] Pei-Shan Chi, Juan Gorraiz, and Wolfgang Glänzel. Comparing capture, usage and citation indicators: an altmetric analysis of journal papers in chemistry disciplines. Scientometrics, 120(3):1461–1473, September 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03168-y.

Celik:2014:MEM

[CGK^{+}14] Embiya Celik, Nuray Gedik, Güler Karaman, Turgay Demirel, and Yuksel Goktas. Mistakes encountered in manuscripts on education and their effects on journal rejections. Scientometrics, 98(3):1837–1853, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1137-y.

Cagliero:2018:DCT

[CGKB18] Luca Cagliero, Paolo Garza, Mohammad Reza Kavoosifar, and Elena Baralis. Discovering cross-topic collaborations among researchers by exploiting weighted association rules. Scientometrics, 116(2):1273–1301, August 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2737-3.

Chan:2015:IQA

[CGPT15] Ho Fai Chan, Malka Guillot, Lionel Page, and Benno Torgler. The inner quality of an article: Will time tell? Scientometrics, 104(1):19–41, July 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1581-y.

Correa:2013:HBE

[CGSS13] Marc Correa, Lucinio González-Sabaté, and Ignacio Serrano. Home bias effect in the management literature. Scientometrics, 95(1):417–433, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0876-5.

Callaert:2012:DSF

[CGV12] Julie Callaert, Joris Grouwels, and Bart Van Looy. Delineating the scientific footprint in technology: Identifying scientific publications within non-patent references. Scientometrics, 91(2):383–398, May 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
011-0573-9.

Cardoso:2010:CER

[CGZ10] Ana Rute Cardoso, Paulo Guimarães, and Klaus F. Zimmermann. Comparing the early research performance of PhD graduates in labor economics in Europe and the USA. Scientometrics, 84(3):621–637, September 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0136-5.

Chuang:2012:CBS

[CH12] Kun-Yang Chuang and Yuh-Shan Ho. Comments on “A bibliometric study of the trend in articles related to eutrophication published in Science Citation Index”. Scientometrics, 91(3):1061–1065, June 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11192-011-0604-6. See [YJ11] and reponse [HW12].

Chang:2013:PII

[CH13a] Han-Wen Chang and Mu-Hsuan Huang. Prominent institutions in international collaboration network in astronomy and astrophysics. Scientometrics, 97(2):443–460, November 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-0976-x.

Choung:2013:EPK

[CH13b] Jae-Yong Choung and Hye-Ran Hwang. The evolutionary patterns of knowledge production in Korea. Scientometrics, 94(2):629–650, February 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0780-z.

Chang:2014:CSI

[CH14] Han-Wen Chang and Mu-Hsuan Huang. Cohesive subgroups in the international collaboration network in astronomy and astrophysics. Scientometrics, 101(3):1587–1607, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1312-9.

Comins:2015:DSR

[CH15] Jordan A. Comins and Thomas W. Hussey. Detecting seminal research contributions to the development and use of the global positioning system by reference publication year spectroscopy. Scientometrics, 104(2):575–580, August 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1598-2.

Chang:2013:CCC

[Cha13] Yu-Wei Chang. A comparison of citation contexts between natural sciences and social sciences and humanities. Scientometrics, 96(2):535–553, August 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0956-1.

Chang:2014:ESA

[Cha14] Yu-Wei Chang. Exploring scientific articles contributed by industries in Taiwan. Scientometrics, 99(2):599–613, May 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1222-2.

Chang:2016:IPL

[Cha16] Yu-Wei Chang. Influence of the principle of least effort across disciplines. Scientometrics, 106(3):1117–1133, March 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1838-0.

Chang:2017:EGE

[Cha17a] Shu-Hao Chang. The evolutionary growth estimation model of international cooperative patent networks. Scientometrics, 112(2):711–729, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
017-2378-y.

Chavalarias:2017:WWS

[Cha17b] David Chavalarias. What’s wrong with science?: Modeling the collective discovery processes with the Nobel game. Scientometrics, 110(1):481–503, January 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2109-9.

Chakraborty:2018:RIC

[Cha18a] Tanmoy Chakraborty. Role of interdisciplinarity in computer sciences: quantification, impact and life trajectory. Scientometrics, 114(3):1011–1029, March 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2628-z.

Chang:2018:PSC

[Cha18b] Shu-Hao Chang. A pilot study on the connection between scientific fields and patent classification systems. Scientometrics, 114(3):951–970, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2613-6.

Chang:2018:EIL

[Cha18c] Yu-Wei Chang. Examining interdisciplinarity of library and information science (LIS) based on LIS articles contributed by non-LIS authors. Scientometrics, 116(3):1589– 1613, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2822-7.

Chandler:2019:IES

[Cha19a] Vincent Chandler. Identifying emerging scholars: seeing through the crystal ball of scholarship selection committees. Scientometrics, 120(1):39–56, July 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03120-0.

Chang:2019:ALI

[Cha19b] Yu-Wei Chang. Are articles in library and information science (LIS) journals primarily contributed to by LIS authors? Scientometrics, 121(1):81–104, October 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03186-w.

Chang:2019:DAS

[Cha19c] Yu-Wei Chang. Definition of authorship in social science journals. Scientometrics, 118(2):563–585, February 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2986-1.

Chen:2013:DFE

[CHC13] Ssu-Han Chen, Mu-Hsuan Huang, and Dar-Zen Chen. Driving factors of external funding and funding effects on academic innovation performance in university-industry-government linkages. Scientometrics, 94(3):1077–1098, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0864-9.

Chapa:2017:CAF

[CHC17] Joaquin Chapa, Zeeshan Haq, and Adam S. Cifu. Comparative analysis of the factors associated with citation and media coverage of clinical research. Scientometrics, 112(3):1271–1283, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2428-5.

Chen:2011:UPA

[Che11] Yu-Shan Chen. Using patent analysis to explore corporate growth. Scientometrics, 88(2):433–448, August 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0396-8.

Chen:2012:DES

[Che12] Tsung Teng Chen. The development and empirical study of a literature review aiding system. Scientometrics, 92 (1):105–116, July 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0728-3.

Chen:2015:QUS

[Che15] Xiaotian Chen. Questionable university-sponsored supplements in high-impact journals. Scientometrics, 105(3):1985– 1995, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1644-0.

Chen:2018:UPC

[Che18a] Bikun Chen. Usage pattern comparison of the same scholarly articles between Web of Science (WoS) and Springer. Scientometrics, 115(1):519–537, April 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2616-3. See correction [Che20].

Chen:2018:EGS

[Che18b] Chaomei Chen. Eugene Garfield’s scholarly impact: a scientometric review. Scientometrics, 114(2):489–516, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2594-5.

Chen:2020:CUP

[Che20] Bikun Chen. Correction to: Usage pattern comparison of the same scholarly articles between Web of Science (WoS) and Springer. Scientometrics, 123(2):1173, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03283-w;   http://
link.springer.com/content/pdf/10.1007/s11192-019-

03283-w.pdf. See [Che18a].

Chi:2014:WRD

[Chi14] Pei-Shan Chi. Which role do non-source items play in the social sciences? A case study in political science in Germany. Scientometrics, 101(2):1195–1213, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-014-1433-1.

Chi:2015:CPC

[Chi15] Pei-Shan Chi. Changing publication and citation patterns in political science in Germany. Scientometrics, 105(3):1833– 1848, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1609-3.

Cho:2010:IPB

[CHL10] Cheng-Chung Cho, Ming-Wen Hu, and Meng-Chun Liu. Improvements in productivity based on co-authorship: a case study of published articles in China. Scientometrics, 85(2): 463–470, November 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0263-z.

Chang:2015:ERS

[CHL15] Yu-Wei Chang, Mu-Hsuan Huang, and Chiao-Wen Lin. Evolution of research subjects in library and information science based on keyword, bibliographical coupling, and cocitation analyses. Scientometrics, 105(3):2071–2087, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1762-8.

Cabanac:2015:ACC

[CHM15] Guillaume Cabanac, Gilles Hubert, and Béatrice Milard. Academic careers in computer science: continuance and transience of lifetime co-authorships. Scientometrics, 102 (1):135–150, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1426-0.

Choi:2012:CPN

[Cho12] Sujin Choi. Core-periphery, new clusters, or rising stars?: international scientific collaboration among ‘advanced’ countries in the era of globalization. Scientometrics, 90(1):25–41, January 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0509-4.

Chung:2014:AST

[Chu14] Chung Joo Chung. An analysis of the status of the triple helix and university-industry-government relationships in Asia. Scientometrics, 99(1):139–149, April 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1100-y.

Cao:2012:RST

[CHWL12] Xiaofeng Cao, Yi Huang, Jie Wang, and Shengji Luan. Research status and trends in limnology journals: a bibliometric analysis based on SCI database. Scientometrics, 92(3): 735–746, September 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0623-y.

Chen:2013:PPJ

[CHY13] Chiang-Ping Chen, Jin-Li Hu, and Chih-Hai Yang. Produce patents or journal articles? A cross-country comparison of R&D productivity change. Scientometrics, 94(3):833–849, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0811-9.

Chang:2016:ACF

[CHY16] Yu-Wei Chang, Mu-Hsuan Huang, and Hsiao-Wen Yang. Analysis of coactivity in the field of fuel cells at institutional and individual levels. Scientometrics, 109(1):143–158, October 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-1957-7.

Cisneros:2018:BSF

[CIK^{+}18] Luis Cisneros, Mihai Ibanescu, Christian Keen, Odette Lobato-Calleros, and Juan Niebla-Zatarain. Bibliometric study of family business succession between 1939 and 2017: mapping and analyzing authors’ networks. Scientometrics, 117(2):919–951, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2889-1.

Cramond:2016:PRC

[CIL^{+}16] Fala Cramond, Cadi Irvine, Jing Liao, David Howells, Emily Sena, Gillian Currie, and Malcolm Macleod. Protocol for a retrospective, controlled cohort study of the impact of a change in Nature journals’ editorial policy for life sciences research on the completeness of reporting study design and execution. Scientometrics, 108(1):315–328, July 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-1964-8.pdf.

Chen:2013:PPI

[CJC13] Jennifer H. Chen, Show-Ling Jang, and Chiao-Hui Chang. The patterns and propensity for international co-invention: the case of China. Scientometrics, 94(2):481–495, February 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0821-7.

Chen:2010:MTD

[CJW10] Jennifer H. Chen, Show-Ling Jang, and Sonya H. Wen. Measuring technological diversification: identifying the effects of patent scale and patent scope. Scientometrics, 84(1):265– 275, July 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0143-6.

Chen:2015:BAW

[CJY^{+}15] Haibin Chen, Wei Jiang, Yan Yang, Xin Man, and Mingyi Tang. A bibliometric analysis of waste management research during the period 1997–2014. Scientometrics, 105(2): 1005–1018, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1714-3.

Chen:2019:CRB

[CjZZ^{+}19] Xi Chen, Huan jing Zhao, Shu Zhao, Jie Chen, and Yan ping Zhang. Citation recommendation based on citation tendency. Scientometrics, 121(2):937–956, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03225-6.

Cho:2014:TIN

[CK14] Yongrae Cho and Wonjoon Kim. Technology-industry networks in technology commercialization: evidence from Korean university patents. Scientometrics, 98(3):1785–1810, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1131-4.

CostaRibeiro:2014:MUG

[CKB^{+}14] Leonardo Costa Ribeiro, Glenda Kruss, Gustavo Britto, Américo Tristão Bernardes, and Eduardo da Motta e Albuquerque. A methodology for unveiling global innovation networks: patent citations as clues to cross border knowledge flows. Scientometrics, 101(1):61–83, October 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1351-2.

Cheng:2010:PDP

[CKCK10] Yin-Hui Cheng, Fu-Yung Kuan, Shih-Chieh Chuang, and Yun Ken. Profitability decided by patent quality? An empirical study of the U.S. semiconductor industry. Scientometrics, 82(1):175–183, January 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
009-0080-4.

Carley:2019:RBF

[CKPY19] Stephen F. Carley, Seokbeom Kwon, Alan L. Porter, and Jan L. Youtie. The relationship between forward and backward diversity in CORE datasets. Scientometrics, 120(3): 961–974, September 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03163-3.

Chatzimichael:2017:MPP

[CKT17] Konstantinos Chatzimichael, Pantelis Kalaitzidakis, and Vangelis Tzouvelekas. Measuring the publishing productivity of economics departments in Europe. Scientometrics, 113(2):889–908, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2509-5.

Chen:2011:UAC

[CL11] Liang-Chu Chen and Yen-Hsuan Lien. Using author cocitation analysis to examine the intellectual structure of elearning: A MIS perspective. Scientometrics, 89(3):867– 886, December 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0458-y.

Czellar:2013:QRW

[CL13] Judith Czellar and Jacques Lanarès. Quality of research: which underlying values? Scientometrics, 95(3):1003–1021, June 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0928-x.

Comins:2016:RSD

[CL16] Jordan A. Comins and Loet Leydesdorff. RPYS i/o: software demonstration of a web-based tool for the historiography and visualization of citation classics, sleeping beauties and research fronts. Scientometrics, 107(3):1509–1517, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1928-z.

Choe:2017:SCR

[CL17a] Hochull Choe and Duk Hee Lee. The structure and change of the research collaboration network in Korea (2000–2011): network analysis of joint patents. Scientometrics, 111(2): 917–939, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Comins:2017:CAI

[CL17b] Jordan A. Comins and Loet Leydesdorff. Citation algorithms for identifying research milestones driving biomedical innovation. Scientometrics, 110(3):1495–1504, March 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2238-1.

Claassen:2015:MUQ

[Cla15] Christopher Claassen. Measuring university quality. Scientometrics, 104(3):793–807, September 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1584-8.

Cuxac:2013:ESS

[CLB13] Pascal Cuxac, Jean-Charles Lamirel, and Valerie Bonvallot. Efficient supervised and semi-supervised approaches for affiliations disambiguation. Scientometrics, 97(1):47–58, October 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1025-5.

Calver:2013:PIB

[CLD13] Michael C. Calver, Maggie Lilith, and Christopher R. Dickman. A ‘perverse incentive’ from bibliometrics: could national research assessment exercises (NRAEs) restrict literature availability for nature conservation? Scientometrics, 95(1):243–255, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0908-1.

Clermont:2016:EER

[Cle16] Marcel Clermont. Effectiveness and efficiency of research in Germany over time: an analysis of German business schools between 2001 and 2009. Scientometrics, 108(3):1347–1381, September 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2013-3.

Chen:2010:CNP

[CLHH10] Dar-Zen Chen, Chang-Pin Lin, Mu-Hsuan Huang, and Chen-Yu Huang. Constructing a new patent bibliometric performance measure by using modified citation rate analyses with dynamic backward citation windows. Scientometrics, 82(1):149–163, January 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0044-8.

Chen:2012:SPE

[CLJH12] Jennifer H. Chen, Shihmin Lo, Show-Ling Jang, and Chi-Cho Huang. Strategic partnership and its effect on external learning of technology descendants. Scientometrics, 92 (1):157–179, July 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0734-5.

Choi:2011:RPS

[CLkS11] Dong Geun Choi, Heesang Lee, and Tae kyung Sung. Research profiling for ‘standardization and innovation’. Scientometrics, 88(1):259–278, July 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
011-0344-7.

Chang:2017:HCA

[CLL^{+}17] Yu-Hsin Chang, Kuei-Kuei Lai, Chien-Yu Lin, Fang-Pei Su, and Ming-Chung Yang. A hybrid clustering approach to identify network positions and roles through social network and multivariate analysis. Scientometrics, 113(3): 1733–1755, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2556-y.

Cheng:2015:MRT

[CLLH15] Qing Cheng, Xin Lu, Zhong Liu, and Jincai Huang. Mining research trends with anomaly detection models: the case of social computing research. Scientometrics, 103(2):453–469, May 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1559-9.

Chiu:2010:TSD

[CLLL10] Yi-Chia Chiu, Hsien-Che Lai, Yi-Ching Liaw, and Tai-Yu Lee. Technological scope: diversified or specialized. Scientometrics, 82(1):37–58, January 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0039-5.

Cheang:2015:IPS

[CLLZ15] Brenda Cheang, Chongshou Li, Andrew Lim, and Zhenzhen Zhang. Identifying patterns and structural influences in the scientific communication of business knowledge. Scientometrics, 103(1):159–189, April 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1518-x.

Castriotta:2019:WNE

[CLMN19] Manuel Castriotta, Michela Loi, Elona Marku, and Luca Naitana. What’s in a name? Exploring the conceptual structure of emerging organizations. Scientometrics, 118(2): 407–437, February 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2977-2.

Chen:2018:UFL

[CLO18] Chuchu Chen, Albert N. Link, and Zachary T. Oliver. U.S. federal laboratories and their research partners: a quantitative case study. Scientometrics, 115(1):501–517, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2665-2.

Chen:2019:EII

[CLSW19] Xiuwen Chen, Jianping Li, Xiaolei Sun, and Dengsheng Wu. Early identification of intellectual structure based on coword analysis from research grants. Scientometrics, 121(1):

349–369, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03187-9.

Chen:2019:RED

[CLW^{+}19] Yu-Shan Chen, Yu-Hsien Lin, Tai-Hsi Wu, Shu-Tzu Hung, Pei-Ju Lucy Ting, and Chen-Han Hsieh. Re-examine the determinants of market value from the perspectives of patent analysis and patent litigation. Scientometrics, 120(1):1–17, July 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03119-7.

Cockriel:2018:IDJ

[CM18] William M. Cockriel and James B. McDonald. The influence of dispersion on journal impact measures. Scientometrics, 116(1):609–622, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2755-1.

Cassi:2014:HED

[CMdT14] Lorenzo Cassi, Wilfriedo Mescheba, and Élisabeth de Turckheim. How to evaluate the degree of interdisciplinarity of an institution? Scientometrics, 101(3):1871–1895, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1280-0.

Ciaramella:2017:TPT

[CMM17] Laurie Ciaramella, Catalina Martínez, and Yann Ménière. Tracking patent transfers in different European countries: methods and a first application to medical technologies. Scientometrics, 112(2):817–850, August 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2411-1.

Chang:2011:WMG

[CMO11] Chia-Lin Chang, Michael McAleer, and Les Oxley. What makes a great journal great in the sciences? Which came first, the chicken or the egg? Scientometrics, 87(1):17–40, April 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0335-0.

Calma:2019:JAM

[CMPD19] Angelito Calma, José Martí-Parreño, and Martin Davies. Journal of the Academy of Marketing Science 1973–2018: an analytical retrospective. Scientometrics, 119(2):879–908, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03080-5.

Calderon-Martinez:2015:LEM

[CMRC15] Aurora Calderón-Martínez and Enar Ruiz-Conde. Leading emerging markets: capturing and diffusing scientific knowledge through research-oriented repositories. Scientometrics, 104(3):907–930, September 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1603-9.

Chan:2018:REC

[CMT18] Ho F. Chan, Franklin G. Mixon, Jr., and Benno Torgler. Relation of early career performance and recognition to the probability of winning the Nobel Prize in Economics. Scientometrics, 114(3):1069–1086, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2614-5.

Chan:2019:FSC

[CMT19] Ho Fai Chan, Franklin G. Mixon, Jr., and Benno Torgler. Fame in the sciences: a culturomics approach. Scientometrics, 118(2):605–615, February 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2975-4.

Cainelli:2015:SST

[CMUdF15] Giulio Cainelli, Mario A. Maggioni, T. Erika Uberti, and Annunziata de Felice. The strength of strong ties: How co-authorship affect productivity of academic economists? Scientometrics, 102(1):673–699, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1421-5.

Cortes:2016:PTR

[CMVP16] Lina M. Cortés, Andrés Mora-Valencia, and Javier Perote. The productivity of top researchers: a semi-nonparametric approach. Scientometrics, 109(2):891–915, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2072-5.

Cristino:2018:EEB

[CNC18] Talita Mariane Cristino, Antonio Faria Neto, and Antonio Fernando Branco Costa. Energy efficiency in buildings: analysis of scientific literature and identification of data analysis techniques from a bibliometric study. Scientometrics, 114(3):1275–1326, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2615-4.

Carley:2017:MSP

[CNPG17] Stephen F. Carley, Nils C. Newman, Alan L. Porter, and Jon G. Garner. A measure of staying power: Is the persistence of emergent concepts more significantly influenced by technical domain or scale? Scientometrics, 111(3):2077– 2087, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Carley:2018:ITE

[CNPG18] Stephen F. Carley, Nils C. Newman, Alan L. Porter, and Jon G. Garner. An indicator of technical emergence. Scientometrics, 115(1):35–49, April 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2654-5.

Chadha:2010:RMV

[CO10] Alka Chadha and Raffaele Oriani. R&D market value under weak intellectual property rights protection: the case of India. Scientometrics, 82(1):59–74, January 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0042-x.

Coccia:2018:GPE

[Coc18] Mario Coccia. General properties of the evolution of research fields: a scientometric study of human microbiome, evolutionary robotics and astrobiology. Scientometrics, 117(2):1265–1283, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2902-8.

Colavizza:2017:SRC

[Col17] Giovanni Colavizza. The structural role of the core literature in history. Scientometrics, 113(3):1787–1809, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2550-4.

Colavizza:2018:DSH

[Col18] Giovanni Colavizza. A diachronic study of historiography. Scientometrics, 117(3):2117–2131, December 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2934-0.

Comins:2015:DMT

[Com15] Jordan A. Comins. Data-mining the technological importance of government-funded patents in the private sector. Scientometrics, 104(2):425–435, August 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1616-4.

Copiello:2018:MVP

[Cop18] Sergio Copiello. On the money value of peer review. Scientometrics, 115(1):613–620, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2664-3.

Copiello:2019:OAC

[Cop19a] Sergio Copiello. The open access citation premium may depend on the openness and inclusiveness of the indexing database, but the relationship is controversial because it is ambiguous where the open access boundary lies. Scientometrics, 121(2):995–1018, November 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03221-w.

Copiello:2019:RIA

[Cop19b] Sergio Copiello. Research interest: another undisclosed (and redundant) algorithm by ResearchGate. Scientometrics, 120 (1):351–360, July 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03124-w.

Canibano:2011:ITM

[COS11a] Carolina Cañibano, F. Javier Otamendi, and Francisco Solís. International temporary mobility of researchers: a cross-discipline study. Scientometrics, 89(2):653–675, November 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0462-2.

Creaser:2011:WDU

[COS11b] Claire Creaser, Charles Oppenheim, and Mark A. C. Summers. What do UK academics cite? An analysis of references cited in UK scholarly outputs. Scientometrics, 86 (3):613–627, March 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0321-6.

Chan:2015:DNL

[CÖT15] Ho Fai Chan, Ali Sina Önder, and Benno Torgler. Do Nobel laureates change their patterns of collaboration following prize reception? Scientometrics, 105(3):2215–2235, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1738-8.

Chan:2016:FCD

[CÖT16a] Ho Fai Chan, Ali Sina Önder, and Benno Torgler. The first cut is the deepest: repeated interactions of coauthorship and academic productivity in Nobel laureate teams. Scientometrics, 106(2):509–524, February 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1796-y.

Conley:2016:AEG

[CÖT16b] John P. Conley, Ali Sina Önder, and Benno Torgler. Are all economics graduate cohorts created equal? Gender, job openings, and research productivity. Scientometrics, 108 (2):937–958, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1987-1.

Carley:2012:FDI

[CP12a] Stephen Carley and Alan L. Porter. A forward diversity index. Scientometrics, 90(2):407–427, February 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0528-1.

Cho:2012:GOI

[CP12b] Seong Eun Cho and Han Woo Park. Government organizations’ innovative use of the Internet: The case of the Twitter activity of South Korea’s Ministry for food, agriculture, forestry and fisheries. Scientometrics, 90(1):9–23, January 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0519-2.

Chung:2012:WVS

[CP12c] Chung Joo Chung and Han Woo Park. Web visibility of scholars in media and communication journals. Scientometrics, 93(1):207–215, October 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0707-8.

Chung:2014:MTH

[CP14] Chung Joo Chung and Han Woo Park. Mapping triple helix innovation in developing and transitional economies: webometrics, scientometrics, and informetrics. Scientometrics, 99(1):1–4, April 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-013-1105-
6.pdf.

Cook:2016:DSF

[CP16] James M. Cook and Dawn Plourde. Do scholars follow Betteridge’s Law? The use of questions in journal article titles. Scientometrics, 108(3):1119–1128, September 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2030-2.

Cova:2013:IUC

[CPF13] Tânia F. G. G. Cova, Alberto A. C. C. Pais, and Sebastião J. Formosinho. Iberian universities: a characterisation from ESI rankings. Scientometrics, 94(3):1239–1251, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0792-8.

Correia:2018:SAS

[CPF18] António Correia, Hugo Paredes, and Benjamim Fonseca. Scientometric analysis of scientific publications in CSCW. Scientometrics, 114(1):31–89, January 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2562-0.

Chamorro-Padial:2019:EEA

[CPRSFVG19] Jorge Chamorro-Padial, Rosa Rodriguez-Sánchez, J. Fdez-Valdivia, and J. A. Garcia. An evolutionary explanation of assassins and zealots in peer review. Scientometrics, 120(3):1373–1385, September 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03171-3.

Callaert:2014:SIM

[CPV14] Julie Callaert, Maikel Pellens, and Bart Van Looy. Sources of inspiration? Making sense of scientific references in patents. Scientometrics, 98(3):1617–1629, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1073-x.

Carley:2013:TMP

[CPY13] Stephen Carley, Alan L. Porter, and Jan Youtie. Toward a more precise definition of self-citation. Scientometrics, 94 (2):777–780, February 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0745-2.

Costa:2016:ECN

[CQB16] Mark R. Costa, Jian Qin, and Sarah Bratt. Emergence of collaboration networks around large scale data repositories: a study of the genomics community using Gen-Bank. Scientometrics, 108(1):21–40, July 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1954-x.

Collazo-Reyes:2014:GNI

[CR14] Francisco Collazo-Reyes. Growth of the number of indexed journals of Latin America and the Caribbean: the effect on the impact of each country. Scientometrics, 98(1):197–209, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1036-2.

Cowan:2018:ESF

[CR18] Robin Cowan and Giulia Rossello. Emergent structures in faculty hiring networks, and the effects of mobility on academic performance. Scientometrics, 117(1):527–562, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2858-8.

Chinchilla-Rodriguez:2015:SPC

[CRAJdMACÁ15] Zaida Chinchilla-Rodríguez, Ricardo Arencibia-Jorge, Félix de Moya-Anegón, and Elena Corera-Álvarez. Somes patterns of Cuban scientific publication in Scopus: the current situation and challenges. Scientometrics, 103(3):779–794, June 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1568-8.

Chinchilla-Rodriguez:2018:TBS

[CRBRG^{+}18] Zaida Chinchilla-Rodríguez, Yi Bu, Nicolás Robinson-García, Rodrigo Costas, and Cassidy R. Sugimoto. Travel bans and scientific mobility: utility of asymmetry and affinity indexes to inform science policy. Scientometrics, 116 (1):569–590, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2738-2;
http://link.springer.com/content/pdf/10.1007/s11192-
018-2738-2.pdf.

Chinchilla-Rodriguez:2012:BCA

[CRFM^{+}12] Zaida Chinchilla-Rodríguez, Anuska Ferligoj, Sandra Miguel, Luka Kronegger, and Félix de Moya-Anegón. Blockmodeling of co-authorship networks in library and information science in Argentina: a case study. Scientometrics, 93(3):699–717, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0794-6.

Collazo-Reyes:2017:CPR

[CRLMLM17] Francisco Collazo-Reyes, Maria Elena Luna-Morales, and Evelia Luna-Morales. Change in the publishing regime in Latin America: from a local to universal journal, Archivos de investigación Médica/Archives of Medical Research (1970–2014). Scientometrics, 110(2):695–709, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2207-8.

Collazo-Reyes:2010:EKP

[CRLMRPA10] Francisco Collazo-Reyes, Ma. Elena Luna-Morales, Jane M. Russell, and Miguel Ángel Pérez-Angón. Enriching knowledge production patterns of Mexican physics in particles and fields. Scientometrics, 85(3):791–802, December 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0229-1.

Collazo-Reyes:2017:EMS

[CRLMRPA17] Francisco Collazo-Reyes, María Elena Luna-Morales, Jane M. Russell, and Miguel Ángel Pérez-Angón. Emergence of modern scientific discourse in the American continent: knowledge claims in the discovery of erythronium/vanadium in Mexico (1802–1832). Scientometrics, 110(3):1505–1521, March 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2220-y.

Chinchilla-Rodriguez:2015:WFA

[CRMdMA15] Zaida Chinchilla-Rodríguez, Sandra Miguel, and Félix de Moya-Anegón. What factors affect the visibility of Argentinean publications in humanities and social sciences in Scopus? Some evidence beyond the geographic realm of research. Scientometrics, 102(1):789–810, January 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1414-4.

Chinchilla-Rodriguez:2018:DAR

[CRMPRS18] Zaida Chinchilla-Rodríguez, Sandra Miguel, Antonio Perianes-Rodríguez, and Cassidy R. Sugimoto. Dependencies and autonomy in research performance: examining nanoscience and nanotechnology in emerging countries. Scientometrics, 115(3):1485–1504, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2652-7;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2652-
7.pdf.

Cheng:2014:DPC

[CRR14] Wei Hong Cheng, Sheng Li Ren, and Ronald Rousseau. Digital publishing and China’s core scientific journals: a position paper. Scientometrics, 98(1):11–22, January 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0873-8.

Caputo:2012:LSR

[CRV12] Carlo Caputo, Jaime Requena, and Domingo Vargas. Life sciences research in Venezuela. Scientometrics, 90(3): 781–805, March 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0548-x.

Chinchilla-Rodriguez:2016:BSP

[CRZGVQdMA16] Zaida Chinchilla-Rodríguez, Grisel Zacca-González, Benjamín Vargas-Quesada, and Félix de Moya-Anegón. Benchmarking scientific performance by decomposing leadership of Cuban and Latin American institutions in public health. Scientometrics, 106(3):1239–1264, March 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1831-z.

Chinchilla-Rodriguez:2015:LAS

[CRZGVQMA15] Zaida Chinchilla-Rodríguez, Grisel Zacca-González, Benjamín Vargas-Quesada, and Félix Moya-Anegón. Latin American scientific output in public health: combined analysis using bibliometric, socioeconomic and health indicators. Scientometrics, 102(1):609–628, January 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1349-9.

Chen:2011:RER

[CS11a] Yu-Shan Chen and Chun-Yu Shih. Re-examine the relationship between patents and Tobin’s q. Scientometrics, 89(3): 781–794, December 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0471-1.

Cho:2011:PCN

[CS11b] Ta-Shun Cho and Hsin-Yu Shih. Patent citation network analysis of core and emerging technologies in Taiwan: 1997– 2008. Scientometrics, 89(3):795–811, December 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0457-z.

Cortes-Sanchez:2019:ILA

[CS19] Julián David Cortés-Sánchez. Innovation in Latin America through the lens of bibliometrics: crammed and fading away. Scientometrics, 121(2):869–895, November 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03201-0;   http://link.springer.
com/content/pdf/10.1007/s11192-019-03201-0.pdf.

Chen:2012:ERU

[CSC12] Yu-Shan Chen, Chun-Yu Shih, and Ching-Hsun Chang. The effects of related and unrelated technological diversification on innovation performance and corporate growth in the Taiwan’s semiconductor industry. Scientometrics, 92(1):117– 134, July 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0720-y.

Chen:2013:PMV

[CSC13] Yu-Shan Chen, Chun-Yu Shih, and Ching-Hsun Chang. Patents and market value in the U.S. pharmaceutical industry: new evidence from panel threshold regression. Scientometrics, 97(2):161–176, November 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0999-3.

Chen:2014:ENR

[CSC14] Yu-Shan Chen, Chun-Yu Shih, and Ching-Hsun Chang. Explore the new relationship between patents and market value: a panel smooth transition regression (PSTR) approach. Scientometrics, 98(2):1145–1159, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-1110-9.

Lo:2010:SLS

[cSL10] Szu chia S. Lo. Scientific linkage of science research and technology development: a case of genetic engineering research. Scientometrics, 82(1):109–120, January 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0036-8.

Carter:2017:GCS

[CSO17] T. Edison Carter, Thomas E. Smith, and Philip J. Osteen. Gender comparisons of social work faculty using H-Index scores. Scientometrics, 111(3):1547–1557, June 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Castro:2018:GLA

[CSR^{+}18] Randelys Molina Castro, Maria Victoria Guzmán Sánchez, Yaidelyn Macías Rivero, Romel Calero Ramos, and Ivet Álvarez Díaz. Global and Latin American scientific production related to pneumococcal vaccines. Scientometrics, 115(3):1549–1559, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2722-x.

Campos:2016:ICB

[CSS^{+}16] Juliana Loureiro Almeida Campos, André Sobral, Josivan Soares Silva, Thiago Antonio Sousa Araújo, Washington Soares Ferreira-Júnior, Flávia Rosa Santoro, Gilney Charll dos Santos, and Ulysses Paulino Albuquerque. Insularity and citation behavior of scientific articles in young fields: the case of ethnobiology. Scientometrics, 109(2):1037–1055, November 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2067-2.

Cocosila:2011:EMI

[CST11] Mihail Cocosila, Alexander Serenko, and Ofir Turel. Exploring the management information systems discipline: a scientometric study of ICIS, PACIS and ASAC. Scientometrics, 87(1):1–16, April 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0331-4.

Chan:2015:DGM

[CT15a] Ho Fai Chan and Benno Torgler. Do great minds appear in batches? Scientometrics, 104(2):475–488, August 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1620-8.

Chan:2015:IEM

[CT15b] Ho Fai Chan and Benno Torgler. The implications of educational and methodological background for the career success of Nobel laureates: an investigation of major awards. Scientometrics, 102(1):847–863, January 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1367-7.

Cai:2019:SIA

[CTL^{+}19] Liwei Cai, Jiahao Tian, Jiaying Liu, Xiaomei Bai, Ivan Lee, Xiangjie Kong, and Feng Xia. Scholarly impact assessment: a survey of citation weighting solutions. Scientometrics, 118 (2):453–478, February 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2973-6.

Tseng:2017:MBL

[cTnHwH17] Hsiang chi Tseng, Wei neng Huang, and Ding wei Huang. Modified Benford’s law for two-exponent distributions. Scientometrics, 110(3):1403–1413, March 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2217-6.

Choudhury:2016:TAL

[CU16] Nazim Choudhury and Shahadat Uddin. Time-aware link prediction to explore network effects on temporal knowledge evolution. Scientometrics, 108(2):745–776, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-016-2003-5.

Coursaris:2014:SAS

[CV14] Constantinos K. Coursaris and Wietske Van Osch. A scientometric analysis of social media research (2004– 2011). Scientometrics, 101(1):357–380, October 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1399-z.

Campani:2015:SIG

[CV15] Marco Campani and Ruggero Vaglio. A simple interpretation of the growth of scientific/technological research impact leading to hype-type evolution curves. Scientometrics, 103(1):75–83, April 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1533-6.

Cavero:2014:CSR

[CVC14] José María Cavero, Belén Vela, and Paloma Cáceres. Computer science research: more production, less productivity. Scientometrics, 98(3):2103–2111, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1178-2.

Cavero:2015:EFA

[CVC^{+}15] José María Cavero, Belén Vela, Paloma Cáceres, Carlos Cuesta, and Almudena Sierra-Alonso. The evolution of female authorship in computing research. Scientometrics, 103 (1):85–100, April 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1520-3.

Cartes-Velasquez:2014:BAA

[CVD14] Ricardo Cartes-Velásquez and Carlos Manterola Delgado. Bibliometric analysis of articles published in ISI dental journals, 2007–2011. Scientometrics, 98(3):2223–2233, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1173-7.

Costas:2010:SCM

[CvLB10] Rodrigo Costas, Thed N. van Leeuwen, and María Bordons. Self-citations at the meso and individual levels: effects of different calculation methods. Scientometrics, 82(3):517–537, March 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-010-0187-7.pdf.

Costas:2011:MSS

[CvLvR11] Rodrigo Costas, Thed N. van Leeuwen, and Anthony F. J. van Raan. The “Mendel syndrome” in science: durability of scientific literature and its effects on bibliometric analysis of individual scientists. Scientometrics, 89(1):177–205, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-011-0436-4.pdf.

Chen:2017:HST

[CW17] Ying Chen and Can Wu. The hot spot transformation in the research evolution of maker. Scientometrics, 113(3): 1307–1324, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2542-4.

Chuang:2011:HIP

[CWH11] Kun-Yang Chuang, Ming-Huang Wang, and Yuh-Shan Ho. High-impact papers presented in the subject category of water resources in the essential science indicators database of the Institute for Scientific Information. Scientometrics, 87 (3):551–562, June 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0365-2.

Calver:2010:WMJ

[CWJBT10] Michael Calver, Grant Wardell-Johnson, Stuart Bradley, and Ross Taplin. What makes a journal international? A case study using conservation biology journals. Scientometrics, 85(2):387–400, November 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0273-x.

Chen:2014:ADR

[CWJC14] Huaqi Chen, Yuehua Wan, Shuian Jiang, and Yanxia Cheng. Alzheimer’s disease research in the future: bibliometric analysis of cholinesterase inhibitors from 1993 to 2012. Scientometrics, 98(3):1865–1877, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1132-3.

Chang:2010:UPA

[CWL10] Pao-Long Chang, Chao-Chan Wu, and Hoang-Jyh Leu. Using patent analyses to monitor the technological trends in an emerging field of technology: a case of carbon nanotube field emission display. Scientometrics, 82(1):5–19, January 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0033-y.

Chen:2015:IRF

[CXpHqZ15] Guo Chen, Lu Xiao, Chang ping Hu, and Xue qin Zhao. Identifying the research focus of library and information science institutions in China with institution-specific keywords. Scientometrics, 103(2):707–724, May 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1545-2.

Chen:2018:RPB

[CXWW18] Wei Chen, Qin-Rui Xing, Hui Wang, and Tao Wang. Retracted publications in the biomedical literature with authors from mainland China. Scientometrics, 114(1):217– 227, January 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2565-x.

Chunlei:2019:RDJ

[CXZ19] Zhou Chunlei, Kong Xiangyi, and Lin Zhipeng. Research on Derek John de Solla Price Medal: Prediction based on academic credit analysis. Scientometrics, 118(1):159–175, January 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2964-7.

Chi:2013:ISR

[CY13] Ruobing Chi and Jonathan Young. The interdisciplinary structure of research on intercultural relations: a co-citation network analysis study. Scientometrics, 96(1):147–171, July 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0894-3.

Chen:2013:KBC

[CYH13] Liang-Chu Chen, Ting-Jung Yu, and Chia-Jung Hsieh. KeyGraph-based chance discovery for exploring the development of e-commerce topics. Scientometrics, 95(1):257–275, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0826-2.

Choi:2011:SNA

[CYK^{+}11] Sungchul Choi, Janghyeok Yoon, Kwangsoo Kim, Jae Yeol Lee, and Cheol-Han Kim. SAO network analysis of patents for technology trends identification: a case study of polymer electrolyte membrane technology in proton exchange membrane fuel cells. Scientometrics, 88(3):863–883, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0420-z.

Choi:2012:USM

[CyPP12] Sujin Choi, Ji young Park, and Han Woo Park. Using social media data to explore communication processes within South Korean online innovation communities. Scientometrics, 90(1):43–56, January 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0514-7.

Chen:2012:AAF

[CYT^{+}12] Chuanfu Chen, Yuan Yu, Qiong Tang, Kuei Chiu, Yan Rao, Xuan Huang, and Kai Sun. Assessing the authority of free online scholarly information. Scientometrics, 90(2):543–560, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0524-5.

Chen:2011:TNH

[CYW^{+}11] Yu-Chun Chen, Hsiao-Yun Yeh, Jau-Ching Wu, Ingo Haschler, Tzeng-Ji Chen, and Thomas Wetter. Taiwan’s national health insurance research database: administrative health care database as study object in bibliometrics. Scientometrics, 86(2):365–380, February 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0289-2.

Cui:2018:BRR

[CZ18] Tiening Cui and Jimei Zhang. Bibliometric and review of the research on circular economy through the evolution of Chinese public policy. Scientometrics, 116(2):1013–1037, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2782-y.

Czerwon:2011:O

[Cze11] Hans-Jürgen Czerwon. Obituary. Scientometrics, 87(3): 655–656, June 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-011-0340-
y.pdf.

Chariker:2017:ISM

[CZPR17] Julia H. Chariker, Yihang Zhang, John R. Pani, and Eric C. Rouchka. Identification of successful mentoring communities using network-based analysis of mentor–mentee relationships across Nobel laureates. Scientometrics, 111(3): 1733–1749, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-017-2364-
4.pdf.

Chen:2010:MSE

[CZV10] Chaomei Chen, Jian Zhang, and Michael S. Vogeley. Making sense of the evolution of a scientific domain: a visual analytic study of the Sloan digital sky survey research. Scientometrics, 83(3):669–688, June 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0123-x.

Cao:2013:BAG

[CZW13] Yang Cao, Sixing Zhou, and Guobin Wang. A bibliometric analysis of global laparoscopy research trends during 1997– 2011. Scientometrics, 96(3):717–730, September 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0942-z.

deArruda:2018:HIT

[dACdFC18] Henrique F. de Arruda, Cesar H. Comin, and Luciano da F. Costa. How integrated are theoretical and applied physics? Scientometrics, 116(2):1113–1121, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2777-8.

deAlmeida:2013:BGP

[dAG13] Elenara Chaves Edler de Almeida and Jorge Almeida Guimarães. Brazil’s growing production of scientific articles — how are we doing with review articles and other qualitative indicators? Scientometrics, 97(2):287–315, November 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-0967-y. See comments

[dCPF14] and response [GdA14].

Daud:2015:UML

[DAMC15] Ali Daud, Muhammad Ahmad, M. S. I. Malik, and Dunren Che. Using machine learning techniques for rising star prediction in co-author network. Scientometrics, 102(2): 1687–1711, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1455-8.

Danell:2014:RIC

[Dan14] Jenny-Ann Brodin Danell. Reception of integrative and complementary medicine (ICM) in scientific journals: a citation and co-word analysis. Scientometrics, 98(2):807–821, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1082-9.

Daniel:2019:LBR

[Dan19] Hans-Dieter Daniel. Lutz Bornmann: Recipient of the 2019 Derek John de Solla Price Medal. Scientometrics, 121(3): 1235–1238, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03251-4.

Nascimento:2015:GTS

[dANR15] Felipe de Araújo Nascimento and Flávia Melo Rodrigues. Growth trend of scientific literature on genetic improvement through the database Scopus. Scientometrics, 105(2): 805–816, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1733-0.

deAraujo:2017:BAS

[dART^{+}17] Alcione Lino de Araújo, Bethânia Ávila Rodrigues, Leomara Battisti Telles, Mônica Cristine S. Vaz, and Juliana Vitória M. Bittencourt. A bibliometric analysis of the Scielo database: a Brazilian portfolio of the solidarity economy. Scientometrics, 112(1):1–20, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Das:2016:MPS

[Das16] Tara Das. Measuring production and scholarly use of National Center for Health Statistics publications: a citation analysis in US government information. Scientometrics, 108(3):1287–1298, September 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2018-y.

Ding:2018:DSN

[DAYY18] Jielan Ding, Per Ahlgren, Liying Yang, and Ting Yue. Disciplinary structures in Nature, Science and PNAS: journal and country levels. Scientometrics, 116(3):1817–1852, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2812-9.

Beaver:2012:QOO

[dB12] Donald deB. Beaver. Quantity is only one of the qualities. Scientometrics, 93(1):33–39, October 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0657-1.

Doslu:2016:CSA

[DB16] Metin Doslu and Haluk O. Bingol. Context sensitive article ranking with citation context analysis. Scientometrics, 108 (2):653–671, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1982-6.

Docampo:2019:NAA

[DB19] Domingo Docampo and Jean-Jacques Bessoule. A new approach to the analysis and evaluation of the research output of countries and institutions. Scientometrics, 119(2):

1207–1225, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03089-w.

Damar:2018:ENA

[DBO^{+}18] Hale Turhan Damar, Ozlem Bilik, Guzin Ozdagoglu, Askin Ozdagoglu, and Muhammet Damar. Evaluating the nursing academicians in Turkey in the scope of Web of Science: scientometrics of original articles. Scientometrics, 115 (1):539–562, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2641-x.

Oliveira:2019:RPS

[dBONM^{+}19] Ivone de Bem Oliveira, Rhewter Nunes, Lucia Mattiello, Stela Barros-Ribeiro, Isabela Pavanelli de Souza, Alexandre Siqueira Guedes Coelho, and Rosane Garcia Collevatti. Research and partnership in studies of sugarcane using molecular markers: a scientometric approach. Scientometrics, 119 (1):335–355, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03047-6.

Docampo:2014:IDS

[DC14] Domingo Docampo and Lawrence Cram. On the internal dynamics of the Shanghai ranking. Scientometrics, 98(2): 1347–1366, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1143-0.

Docampo:2015:EIS

[DC15a] Domingo Docampo and Lawrence Cram. On the effects of institutional size in university classifications: the case of the Shanghai ranking. Scientometrics, 102(2):1325–1346, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1488-z.

Dong:2015:PTC

[DC15b] Dahui Dong and Meng-Lin Chen. Publication trends and co-citation mapping of translation studies between 2000 and 2015. Scientometrics, 105(2):1111–1128, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1769-1.

Docampo:2017:API

[DC17] Domingo Docampo and Lawrence Cram. Academic performance and institutional resources: a cross-country analysis of research universities. Scientometrics, 110(2):739–764, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2189-6.

Docampo:2019:HCR

[DC19] Domingo Docampo and Lawrence Cram. Highly cited researchers: a moving target. Scientometrics, 118(3): 1011–1025, March 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2993-2.

Moreno:2016:EUI

[dCCMAW16a] María del Carmen Calatrava Moreno, Thomas Auzinger, and Hannes Werthner. Erratum to: On the uncertainty of interdisciplinarity measurements due to incomplete bibliographic data. Scientometrics, 107(1):233–234, April 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1902-9;   http://link.
springer.com/content/pdf/10.1007/s11192-016-1902-

9.pdf. See [dCCMAW16b].

Moreno:2016:UIM

[dCCMAW16b] María del Carmen Calatrava Moreno, Thomas Auzinger, and Hannes Werthner. On the uncertainty of interdisciplinarity measurements due to incomplete bibliographic data. Scientometrics, 107(1):213–232, April 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1842-4;   http://link.
springer.com/content/pdf/10.1007/s11192-016-1842-

4.pdf. See erratum [dCCMAW16a].

deCarvalho:2019:IAS

[dCdAMB19] Priscila Pereira Suzart de Carvalho, Ricardo de Araújo Kalid, Jorge Laureano Moya Rodríguez, and Sandro Breval Santiago. Interactions among stakeholders in the processes of city logistics: a systematic review of the literature. Scientometrics, 120(2):567–607, August 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03149-1.

deCarvalho:2015:TYU

[dCdSNB15] Gilson Correia de Carvalho, Raymundo José de Sá-Neto, and Francisco Barros. Thirty years of use of multivariate quantitative methods in benthic community ecology of marine and coastal habitats: looking to the past to planning the future. Scientometrics, 105(1):593–610, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1667-6.

DeFilippo:2012:VIR

[DCGZ^{+}12] Daniela De Filippo, Fernando Casani, Carlos García-Zorita, Preiddy Efraín-García, and Elías Sanz-Casado. Visibility in international rankings. Strategies for enhancing the competitiveness of Spanish universities. Scientometrics, 93(3): 949–966, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0749-y.

DeSordi:2016:BIC

[DCM16] José Osvaldo De Sordi, Marco Antonio Conejero, and Manuel Meireles. Bibliometric indicators in the context of regional repositories: proposing the D-index. Scientometrics, 107(1):235–258, April 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1873-x.

Filho:2014:CAB

[dCPF14] Roberto de Camargo Penteado Filho. Comments on the article “Brazil’s growing production of scientific articles — how are we doing with review articles and other qualitative indicators?” by Elenara Chaves Edler de Almeida and Jorge Almeida Guimarães: Scientometrics (2013) 97:287–315. doi: 10.1007/s11192-013-0967-y. Scientometrics, 101(3):2077–2078, December 2014. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-014-1416-2. See [dAG13] and response

[GdA14].

deCampos:2018:CQA

[dCPRP18] Elaine Aparecida Regiani de Campos, Regina Negri Pagani, Luis Mauricio Resende, and Joseane Pontes. Construction and qualitative assessment of a bibliographic portfolio using the methodology Methodi Ordinatio. Scientometrics, 116 (2):815–842, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2798-3.

DiCaro:2012:IDD

[DCS12] Luigi Di Caro, Mario Cataldi, and Claudio Schifanella. The d-index: Discovering dependences among scientific collaborators from their bibliographic data records. Scientometrics, 93(3):583–607, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0762-1.

Du:2017:EIA

[DCY^{+}17] Wei Du, Xusen Cheng, Chen Yang, Jianshan Sun, and Jian Ma. Establishing interoperability among knowledge organization systems for research management: a social network approach. Scientometrics, 112(3):1489–1506, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2457-0.

Delgado:2018:HIB

[DD18] Alberto Falk Delgado and Anna Falk Delgado. Home institution bias in the New England Journal of Medicine? A noninferiority study on citation rates. Scientometrics, 115 (1):607–611, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2584-7;
http://link.springer.com/content/pdf/10.1007/s11192-
017-2584-7.pdf.

Doloreux:2019:TIM

[DdlPPL^{+}19] David Doloreux, Jose Gaviria de la Puerta, Iker Pastor-López, Igone Porto Gómez, Borja Sanz, and Jon Mikel Zabala-Iturriagagoitia. Territorial innovation models: to be or not to be, that’s the question. Scientometrics, 120(3):1163–1191, September 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03181-1.

dosSantosRubem:2015:CAS

[ddMS15] Ana Paula dos Santos Rubem, Ariane Lima de Moura, and João Carlos Correia Baptista Soares de Mello. Comparative analysis of some individual bibliometric indices when applied to groups of researchers. Scientometrics, 102(1): 1019–1035, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1428-y.

DAgostino:2017:HSI

[DDR17] Marcello D’Agostino, Valentino Dardanoni, and Roberto Ghiselli Ricci. How to standardize (if you must). Scientometrics, 113(2):825–843, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2495-7.

Costa:2019:ITB

[DDS^{+}19a] Matheus Becker Da Costa, Leonardo Moraes Aguiar Lima Dos Santos, Jones Luís Schaefer, Ismael Cristofer Baierle, and Elpidio Oscar Benitez Nara. Industry 4.0 technologies basic network identification. Scientometrics, 121(2): 977–994, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03216-7.

Dobranszki:2019:CFA

[DdS19b] Judit Dobránszki and Jaime A. Teixeira da Silva. Corrective factors for author- and journal-based metrics impacted by citations to accommodate for retractions. Scientometrics, 121(1):387–398, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11192-019-03190-0;   http://link.springer.
com/content/pdf/10.1007/s11192-019-03190-0.pdf.

DeVisscher:2013:NPE

[De 13] Alex De Visscher. A new Price’s estimate on the size of scientific specialties based on scientific community structure. Scientometrics, 96(3):937–940, September 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0897-0.

DeMarchi:2016:FST

[De 16a] Mario De Marchi. First steps towards a consistent classification of innovation. Scientometrics, 108(2):983–985, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1994-2.

DeMarchi:2016:TSI

[De 16b] Mario De Marchi. A taxonomy of S&T indicators. Scientometrics, 106(3):1265–1268, March 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1823-z.

DeLanghe:2017:TDS

[De 17] Rogier De Langhe. Towards the discovery of scientific revolutions in scientometric data. Scientometrics, 110(1): 505–519, January 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-016-
2108-x.

Docampo:2015:EUM

[DEC15] D. Docampo, D. Egret, and L. Cram. The effect of university mergers on the Shanghai ranking. Scientometrics, 104(1):175–191, July 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1587-5.

Demir:2018:PCN

[Dem18] Selcuk Besir Demir. Pros and cons of the new financial support policy for Turkish researchers. Scientometrics, 116(3):2053–2068, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2833-4.

David:2015:LPO

[DF15] Daniel David and Petre Frangopol. The lost paradise, the original sin, and the dodo bird: a scientometrics Sapere Aude manifesto as a reply to the Leiden manifesto on scientometrics. Scientometrics, 105(3):2255–2257, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1634-2.

Daraio:2018:AIB

[DFG^{+}18] Cinzia Daraio, Francesco Fabbri, Giulia Gavazzi, Maria Grazia Izzo, Luca Leuzzi, Giammarco Quaglia, and Giancarlo Ruocco. Assessing the interdependencies between scientific disciplinary profiles. Scientometrics, 116(3):1785– 1803, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2816-5.

Pachi:2012:RBC

[dFPYdCL12] Clarice Gameiro da Fonseca Pachi, Jorge Futoshi Yamamoto, Anna Paula Amadeu da Costa, and Luis Fernandez Lopez. Relationship between connectivity and academic productivity. Scientometrics, 93(2):265–278, November 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0666-0.

Demarest:2014:RME

[DFS14] Bradford Demarest, Guo Freeman, and Cassidy R. Sugimoto. The reviewer in the mirror: examining gendered and ethnicized notions of reciprocity in peer review. Scientometrics, 101(1):717–735, October 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1354-z.

DeBattisti:2015:DRS

[DFS15] Francesca De Battisti, Alfio Ferrara, and Silvia Salini. A decade of research in statistics: a topic model approach. Scientometrics, 103(2):413–433, May 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1554-1.

Veneroso:2019:SRB

[dFVDU^{+}19] João Mateus de Freitas Veneroso, Marlon Dias, Alberto Ueda, Sabir Ribas, Berthier Ribeiro-Neto, Nivio Ziviani, and Edmundo de Souza e Silva. P -score: a reputation bibliographic index that complements citation counts. Scientometrics, 121(3):1269–1291, December 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03247-0.

Daraio:2016:GCD

[DG16] Cinzia Daraio and Wolfgang Glänzel. Grand challenges in data integration-state of the art and future perspectives: an introduction. Scientometrics, 108(1):391–400, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s11192-016-1914-5.pdf.

Delgado:2019:LPR

[DGD19] Alberto Falk Delgado, Gregory Garretson, and Anna Falk Delgado. The language of peer review reports on articles published in the BMJ, 2014–2017: an observational study. Scientometrics, 120(3):1225–1235, September 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03160-6;   http://
link.springer.com/content/pdf/10.1007/s11192-019-
03160-6.pdf.

Dorta-Gonzalez:2011:CIC

[DGDG11] Pablo Dorta-González and María-Isabel Dorta-González. Central indexes to the citation distribution: a complement to the h-index. Scientometrics, 88(3):729–745, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0453-3.

Dorta-Gonzalez:2013:CJD

[DGDG13] P. Dorta-González and M. I. Dorta-González. Comparing journals from different fields of science and social science through a JCR subject categories normalized impact factor. Scientometrics, 95(2):645–672, May 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0929-9.

Dorta-Gonzalez:2015:AAC

[DGDGSV15] Pablo Dorta-González, María Isabel Dorta-González, and Rafael Suárez-Vega. An approach to the author citation potential: measures of scientific performance which are invariant across scientific fields. Scientometrics, 102(2):1467– 1496, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1459-4.

Dunaiski:2017:ESA

[DGF17] Marcel Dunaiski, Gillian J. Greene, and Bernd Fischer. Exploratory search of academic publication and citation data using interactive tag cloud visualizations. Scientometrics, 110(3):1539–1571, March 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2236-3.

Dorta-Gonzalez:2017:RGO

[DGGBDG17] Pablo Dorta-González, Sara M. González-Betancor, and María Isabel Dorta-González. Reconsidering the gold open access citation advantage postulate in a multidisciplinary context: an analysis of the subject categories in the Web of Science database 2009–2014. Scientometrics, 112(2): 877–901, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2422-y.

Diallo:2015:OMS

[DGPL15] Saikou Y. Diallo, Ross J. Gore, Jose J. Padilla, and Christopher J. Lynch. An overview of modeling and simulation using content analysis. Scientometrics, 103(3):977–1002, June 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1578-6.

Ding:2013:BER

[DGWZ13] Zuo-Qi Ding, Jian-Ping Ge, Xiao-Ming Wu, and Xiao-Nan Zheng. Bibliometrics evaluation of research performance in pharmacology/pharmacy: China relative to ten representative countries. Scientometrics, 96(3):829–844, September 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-0968-x.

Danell:2013:CPF

[DH13a] Rickard Danell and Mikael Hjerm. Career prospects for female university researchers have not improved. Scientometrics, 94(3):999–1006, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0840-4.

DeWitte:2013:WAE

[DH13b] Kristof De Witte and Lenka Hudrlikova. What about excellence in teaching? a benevolent ranking of universities. Scientometrics, 96(1):337–364, July 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0971-2.

Ding:2014:MSI

[Din14] Ying Ding. Measuring scholarly impact: methods and practice. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-319-10376-8 (paperback), 3-319-10377-6 (e-book). xiv + 346 pp. LCCN

Z669.8 .M43 2014. URL http://www.loc.gov/catdir/
enhancements/fy1501/2014950682-d.html;   http://
www.loc.gov/catdir/enhancements/fy1501/2014950682-
t.html.

deJesus:2015:IBP

[dJC15] Igor Rosa Dias de Jesus and Helder Gomes Costa. Interfaces between production engineering and the public affairs: evidences from bibliometric analysis. Scientometrics, 105(2): 1183–1193, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1724-1.

Duffy:2011:RPA

[DJWS11] Ryan D. Duffy, Alex Jadidian, Gregory D. Webster, and Kyle J. Sandell. The research productivity of academic psychologists: assessment, trends, and best practice recommendations. Scientometrics, 89(1):207–227, October 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0452-4.

Dranev:2018:DRP

[DKS18] Yury Dranev, Maxim Kotsemir, and Boris Syomin. Diversity of research publications: relation to agricultural productivity and possible implications for STI policy. Scientometrics, 116(3):1565–1587, September 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2799-2.

DeMarchi:2016:MIJ

[DL16a] Mario De Marchi and E. Lorenzetti. Measuring the impact of journals, a reprise. Scientometrics, 108(2):995–997, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1980-8.

DeMarchi:2016:MIS

[DL16b] Mario De Marchi and Edoardo Lorenzetti. Measuring the impact of scholarly journals in the humanities field. Scientometrics, 106(1):253–261, January 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1787-z.

Diallo:2016:IKP

[DLGP16] Saikou Y. Diallo, Christopher J. Lynch, Ross Gore, and Jose J. Padilla. Identifying key papers within a journal via network centrality measures. Scientometrics, 107(3): 1005–1020, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1891-8.

Daraio:2016:DIR

[DLL^{+}16a] Cinzia Daraio, Maurizio Lenzerini, Claudio Leporelli, Henk F. Moed, Paolo Naggar, Andrea Bonaccorsi, and Alessandro Bartolucci. Data integration for research and innovation policy: an Ontology-Based Data Management approach. Scientometrics, 106(2):857–871, February 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1814-0.

Daraio:2016:AOB

[DLL^{+}16b] Cinzia Daraio, Maurizio Lenzerini, Claudio Leporelli, Paolo Naggar, Andrea Bonaccorsi, and Alessandro Bartolucci. The advantages of an ontology–based data management approach: openness, interoperability and data quality. Scientometrics, 108(1):441–455, July 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1913-6.

Dong:2017:ASS

[DLL^{+}17] Hongguang Dong, Menghui Li, Ru Liu, Chensheng Wu, and Jinshan Wu. Allometric scaling in scientific fields. Scientometrics, 112(1):583–594, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Dilger:2015:SPA

[DLM15] Alexander Dilger, Laura Lütkenhöner, and Harry Müller. Scholars’ physical appearance, research performance, and feelings of happiness. Scientometrics, 104(2):555–573, August 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1623-5.

Du:2015:MFM

[DLMX15] Wei Du, Raymond Yiu Keung Lau, Jian Ma, and Wei Xu. A multi-faceted method for science classification schemes (SCSs) mapping in networking scientific resources. Scientometrics, 105(3):2035–2056, December 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1742-z.

deMesnard:2010:HAT

[dM10] Louis de Mesnard. On Hochberg et al.’s “The tragedy of the reviewer commons”. Scientometrics, 84(3):903– 917, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0141-8.

deMoya-Anegon:2014:HIP

[dMALIM14] Félix de Moya-Anegón, Carmen López-Illescas, and Henk F. Moed. How to interpret the position of private sector institutions in bibliometric rankings of research institutions. Scientometrics, 98(1):283–298, January 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1087-4.

Decullier:2017:PRP

[DMB17] E. Decullier, H. Maisonneuve, and J. N. Besson. Publication in 6 rehabilitation professions: a five-year professionalbased bibliometric overview. Scientometrics, 113(2):751– 764, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2460-5.

Dastidar:2013:CSD

[DMM13] Prabir G. Dastidar, Ajoy Mallik, and Nripendranath Mandal. Contribution of shrimp disease research to the development of the shrimp aquaculture industry: an analysis of the research and innovation structure across the countries. Scientometrics, 97(3):659–674, December 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0977-9.

Dechezlepretre:2017:IPF

[DMM17] Antoine Dechezleprêtre, Yann Ménière, and Myra Mohnen. International patent families: from application strategies to statistical indicators. Scientometrics, 111(2):793–828, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-017-2311-4.pdf.

Dehon:2010:UEA

[DMV10] Catherine Dehon, Alice McCathie, and Vincenzo Verardi. Uncovering excellence in academic rankings: a closer look at the Shanghai ranking. Scientometrics, 83(2):515–524, May 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0076-0.

Delbari:2015:MII

[DNAH15] Seyyed Ali Delbari, Siew Imm Ng, Yuhanis Abdul Aziz, and Jo Ann Ho. Measuring the influence and impact of competitiveness research: a Web of Science approach. Scientometrics, 105(2):773–788, November 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1731-2.

Machado:2016:ISS

[dNMVQL16] Raymundo das Neves Machado, Benjamín Vargas-Quesada, and Jacqueline Leta. Intellectual structure in stem cell research: exploring Brazilian scientific articles from 2001 to 2010. Scientometrics, 106(2):525–537, February 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1793-1.

Docampo:2011:EUS

[Doc11a] Domingo Docampo. Erratum to: On using the Shanghai ranking to assess the research performance of university systems. Scientometrics, 86(1):237, January 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-010-0315-4.pdf. See [Doc11b].

Docampo:2011:USR

[Doc11b] Domingo Docampo. On using the Shanghai ranking to assess the research performance of university systems. Scientometrics, 86(1):77–92, January 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0280-y. See erratum [Doc11a].

Docampo:2012:ASI

[Doc12] Domingo Docampo. Adjusted sum of institutional scores as an indicator of the presence of university systems in the ARWU ranking. Scientometrics, 90(2):701–713, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0490-y.

Docampo:2013:RSA

[Doc13] Domingo Docampo. Reproducibility of the Shanghai academic ranking of world universities results. Scientometrics, 94(2):567–587, February 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0801-y.

deOliveira:2016:ISS

[dOM16] Alexandre Rodrigues de Oliveira and Carlos Fernando Mello. Importance and susceptibility of scientific productivity indicators: two sides of the same coin. Scientometrics, 109(2):697–722, November 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2047-6.

Donner:2017:DTA

[Don17] Paul Donner. Document type assignment accuracy in the journal citation index data of Web of Science. Scientometrics, 113(1):219–236, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2483-y.

Dorsch:2017:RVA

[Dor17] Isabelle Dorsch. Relative visibility of authors’ publications in different information services. Scientometrics, 112(2): 917–925, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2416-9.

doPrado:2016:MAC

[dPdCAdMC^{+}16] José Willer do Prado, Valderí de Castro Alcântara, Francisval de Melo Carvalho, Kelly Carvalho Vieira, Luiz Kennedy Cruz Machado, and Dany Flávio Tonelli. Multivariate analysis of credit risk and bankruptcy research data: a bibliometric study involving different knowledge fields (1968–2014). Scientometrics, 106(3):1007–1029, March 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1829-6.

Drew:2016:ARI

[DPF^{+}16] Christina H. Drew, Kristianna G. Pettibone, Fallis Owen Finch III, Douglas Giles, and Paul Jordan. Automated Research Impact Assessment: a new bibliometrics approach. Scientometrics, 106(3):987–1005, March 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1828-7.

dePaulo:2018:MCC

[dPSS18] Alex Fabianne de Paulo, Evandro Marcos Saidel Ribeiro, and Geciane Silveira Porto. Mapping countries cooperation networks in photovoltaic technology development based on patent analysis. Scientometrics, 117(2):667–686, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2892-6.

Ding:2011:AII

[DQ11] Jingda Ding and Junping Qiu. An approach to improve the indicator weights of scientific and technological competitiveness evaluation of Chinese universities. Scientometrics, 86 (2):285–297, February 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0268-7.

DeMarchi:2010:NRE

[DR10a] Mario De Marchi and Maurizio Rocchi. Note on R&D expenditures and fixed capital formation. Scientometrics, 85(2):489–494, November 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
010-0225-5.

DeWitte:2010:PPA

[DR10b] Kristof De Witte and Nicky Rogge. To publish or not to publish? On the aggregation and drivers of research performance. Scientometrics, 85(3):657–680, December 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-010-0286-5.pdf.

Dey:2017:SBC

[DRCG17] Ratnadeep Dey, Anurag Roy, Tanmoy Chakraborty, and Saptarshi Ghosh. Sleeping beauties in computer science: characterization and early identification. Scientometrics, 113(3):1645–1663, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2543-3.

Damaceno:2019:BAG

[DRMMC19] Rafael J. P. Damaceno, Luciano Rossi, Rogério Mugnaini, and Jesús P. Mena-Chalco. The Brazilian academic genealogy: evidence of advisor–advisee relationships through quantitative analysis. Scientometrics, 119 (1):303–333, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03023-0.

Dubois:2014:PMA

[DRS14] Pierre Dubois, Jean-Charles Rochet, and Jean-Marc Schlenker. Productivity and mobility in academic research: evidence from mathematicians. Scientometrics, 98 (3):1669–1701, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1112-7.

Demetrescu:2018:AAN

[DRS18] Camil Demetrescu, Andrea Ribichini, and Marco Schaerf. Accuracy of author names in bibliographic data sources: an Italian case study. Scientometrics, 117(3):1777–1791, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2945-x.

daSilva:2017:DCN

[dS17a] Jaime A. Teixeira da Silva. Does China need to rethink its metrics- and citation-based research rewards policies? Scientometrics, 112(3):1853–1857, September 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2430-y. See comment [Shu17].

daSilva:2017:NSC

[dS17b] Jaime A. Teixeira da Silva. Nested self-citation: the citation of a paper’s least divisible unit. Scientometrics, 111(1):547–552, April 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-017-2249-6.

daSilva:2018:GSI

[dS18] Jaime A. Teixeira da Silva. The Google Scholar h-index: useful but burdensome metric. Scientometrics, 117(1): 631–635, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2859-7.

Almendra:2015:RCS

[dSAEE15] Vinicius da Silva Almendra, Denis Enachescu, and Cornelia Enachescu. Ranking computer science conferences using self-organizing maps with dynamic node splitting. Scientometrics, 102(1):267–283, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1436-y.

daSilva:2017:WDS

[dSBC17] Jaime A. Teixeira da Silva and Helmar Bornemann-Cimenti. Why do some retracted papers continue to be cited? Scientometrics, 110(1):365–370, January 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2178-9.

daSilva:2017:HCR

[dSD17] Jaime A. Teixeira da Silva and Judit Dobránszki. Highly cited retracted papers. Scientometrics, 110(3):1653–1661, March 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2227-4.

daSilva:2018:CIE

[dSD18a] Jaime A. Teixeira da Silva and Judit Dobránszki. Citation inflation: the effect of not correcting the scientific literature sufficiently, a case study in the plant sciences. Scientometrics, 116(2):1213–1222, August 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2759-x.

daSilva:2018:MVI

[dSD18b] Jaime A. Teixeira da Silva and Judit Dobránszki. Multiple versions of the h-index: cautionary use for formal academic purposes. Scientometrics, 115(2):1107–1113, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2680-3. See comments

[BI18a, CF18] and rejoinder [dSD18c].

daSilva:2018:RMV

[dSD18c] Jaime A. Teixeira da Silva and Judit Dobránszki. Rejoinder to “Multiple versions of the h-index: cautionary use for formal academic purposes”. Scientometrics, 115(2):1131–1137, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2684-z. See [BI18a, CF18,

dSD18b].

deStefano:2016:DSM

[dSdSSB16] Ercilia de Stefano, Marcio Peixoto de Sequeira Santos, and Ronaldo Balassiano. Development of a software for metric studies of transportation engineering journals. Scientometrics, 109(3):1579–1591, December 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2152-6.

deSouza:2013:RPC

[dSF13] Cristina Gomes de Souza and Marta Lúcia Azevedo Ferreira. Researchers profile, co-authorship pattern and knowledge organization in information science in Brazil. Scientometrics, 95(2):673–687, May 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0882-7.

deSouza:2015:CNM

[dSFSF15] Luiz Gustavo Antonio de Souza, Márcia Azanha Ferraz Dias de Moraes, Maria Ester Soares Dal Poz, and José Maria Ferreira Jardim da Silveira. Collaborative networks as a measure of the innovation systems in second-generation ethanol. Scientometrics, 103(2):355–372, May 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1553-2.

Dai:2015:BAR

[DSG^{+}15] Yunrong Dai, Yonghui Song, Hongjie Gao, Siyu Wang, and Yu Yuan. Bibliometric analysis of research progress in membrane water treatment technology from 1985 to 2013. Scientometrics, 105(1):577–591, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1669-4.

Derrick:2010:CBT

[DSH^{+}10] G. E. Derrick, H. Sturk, A. S. Haynes, S. Chapman, and W. D. Hall. A cautionary bibliometric tale of two cities. Scientometrics, 84(2):317–320, August 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0118-7.

Delanghe:2011:ERP

[DSM11] Henri Delanghe, Brian Sloan, and Ugur Muldur. European research policy and bibliometric indicators, 1990– 2005. Scientometrics, 87(2):389–398, May 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0308-3.

daSilva:2017:CCS

[dSM17] Jaime A. Teixeira da Silva and Aamir Raoof Memon. CiteScore: A cite for sore eyes, or a valuable, transparent metric? Scientometrics, 111(1):553–556, April 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-017-2250-0.

daSilva:2018:GBS

[dSNV18] Carlos Eduardo M. Viegas da Silva, Rubens Nunes, and Elisabete Maria Macedo Viegas. A genealogy of the Brazilian scientific research on freshwater fish farming by means of the academic supervision linkage. Scientometrics, 117(3): 1535–1553, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2940-2.

SantAnna:2014:NPB

[dSSdMAF14] Leonardo da Silva Sant’Anna, Maria Simone de Menezes Alencar, and Aldo Pacheco Ferreira. Nanomaterials patenting in Brazil: some considerations for the national regulatory framework. Scientometrics, 100(3):675–686, September 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1300-0.

doSul:2018:ECD

[dSTL18] Juliana A. Ivar do Sul, Alexander S. Tagg, and Matthias Labrenz. Exploring the common denominator between microplastics and microbiology: a scientometric approach. Scientometrics, 117(3):2145–2157, December 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2936-y.

Darvish:2016:DNK

[DT16] Hamid Darvish and Yasar Tonta. Diffusion of nanotechnology knowledge in Turkey and its network structure. Scientometrics, 107(2):569–592, May 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1854-0.

DEste:2013:PAE

[DTM^{+}13] Pablo D’Este, Puay Tang, Surya Mahdi, Andy Neely, and Mabel Sánchez-Barrioluengo. The pursuit of academic excellence and business engagement: is it irreconcilable? Scientometrics, 95(2):481–502, May 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0955-2.

Dehdarirad:2014:RTG

[DVB14] Tahereh Dehdarirad, Anna Villarroya, and Maite Barrios. Research trends in gender differences in higher education and science: a co-word analysis. Scientometrics, 101(1): 273–290, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1327-2.

Dehdarirad:2015:RWS

[DVB15] Tahereh Dehdarirad, Anna Villarroya, and Maite Barrios. Research on women in science and higher education: a bibliometric analysis. Scientometrics, 103(3):795– 812, June 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1574-x.

Diez-Vial:2017:RES

[DVMS17] Isabel Diez-Vial and Angeles Montoro-Sanchez. Research evolution in science parks and incubators: foundations and new trends. Scientometrics, 110(3):1243–1272, March 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2218-5.

deWinter:2015:RBT

[dW15] J. C. F. de Winter. The relationship between tweets, citations, and article views for PLOS ONE articles. Scientometrics, 102(2):1773–1779, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1445-x.

Du:2018:PFI

[DW18] Jian Du and Yishan Wu. A parameter-free index for identifying under-cited sleeping beauties in science. Scientometrics, 116(2):959–971, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2780-0.

Dong:2016:LAC

[DWGL16] Yanping Dong, Panzhi Wang, Lan Guo, and Hongqun Liu. ”Listing author contribution” does not alter the author inflation in the publications in basic research in four major gastroenterology journals in 10 years. Scientometrics, 107 (3):1501–1507, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1923-4.

deWinter:2014:EGS

[dWZD14] Joost C. F. de Winter, Amir A. Zadpoor, and Dimitra Dodou. The expansion of Google Scholar versus Web of Science: a longitudinal study. Scientometrics, 98(2):1547–1565, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1089-2.

Duan:2017:DPJ

[DX17] Yufeng Duan and Zequan Xiong. Download patterns of journal papers and their influencing factors. Scientometrics, 112(3):1761–1775, September 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2456-1.

Dong:2012:BAS

[DXL^{+}12] Bensi Dong, Guoqiang Xu, Xiang Luo, Yi Cai, and Wei Gao. A bibliometric analysis of solar power research from 1991 to 2010. Scientometrics, 93(3):1101–1117, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0730-9.

Dong:2018:IMI

[DXL^{+}18] Kun Dong, Haiyun Xu, Rui Luo, Ling Wei, and Shu Fang. An integrated method for interdisciplinary topic identification and prediction: a case study on information science and library science. Scientometrics, 115(2):849–868, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2694-x.

DelRosarioBenavides:2018:AAB

[DY18] Maria Del Rosario Benavides and Marcus Antonius Ynalvez. Academics’ “ambidextrous behavior” and doctoral science mentoring practices. Scientometrics, 115(1):79–109, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2670-5.

Dyachenko:2014:IAJ

[Dya14] Ekaterina L. Dyachenko. Internationalization of academic journals: Is there still a gap between social and natural sciences? Scientometrics, 101(1):241–255, October 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1357-9.

Dyachenko:2017:CIM

[Dya17a] Ekaterina L. Dyachenko. Correction to: Internal migration of scientists in Russia and the USA: the case of physicists. Scientometrics, 113(3):1823, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2553-1;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2553-

1.pdf. See [Dya17b].

Dyachenko:2017:IMS

[Dya17b] Ekaterina L. Dyachenko. Internal migration of scientists in Russia and the USA: the case of physicists. Scientometrics, 113(1):105–122, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11192-017-2478-8. See correction [Dya17a].

Zhu:2015:RPE

[dZLwC^{+}15] Wei dong Zhu, Fang Liu, Yu wang Chen, Jian bo Yang, Dong ling Xu, and Dong peng Wang. Research project evaluation and selection: an evidential reasoning rule-based method for aggregating peer review information with reliabilities. Scientometrics, 105(3):1469–1490, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1770-8.

Ennas:2015:DDJ

[EBD15] Gianfranco Ennas, Battista Biggio, and Maria Chiara Di Guardo. Data-driven journal meta-ranking in business and management. Scientometrics, 105(3):1911–1929, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1751-y.

Elango:2016:DHR

[EBK16] Bakthavachalam Elango, Lutz Bornmann, and Govindaraju Kannan. Detecting the historical roots of tribology research: a bibliometric analysis. Scientometrics, 107(1):305–313, April 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1877-6.

Eito-Brun:2016:YSR

[EBR16] Ricardo Eito-Brun and María Ledesma Rodríguez. 50 years of space research in Europe: a bibliometric profile of the European Space Agency (ESA). Scientometrics, 109(1):551–576, October 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2053-8.

Enger:2016:WGH

[EC16] Simen G. Enger and Fulvio Castellacci. Who gets Horizon 2020 research grants? Propensity to apply and probability to succeed in a two-step analysis. Scientometrics, 109(3):1611–1638, December 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2145-5.

El-Din:2016:BAE

[EDEH16] Hanaa M. H. Alam El-Din, Ahmed Sharaf Eldin, and Amro M. S. A. Hanora. Bibliometric analysis of Egyptian publications on Hepatitis C virus from PubMed using data mining of an in-house developed database (HCVDBegy). Scientometrics, 108(2):895–915, August 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-2007-1.

Erfanmanesh:2019:SOQ

[EdS19] Mohammadamin Erfanmanesh and Jaime A. Teixeira da Silva. Is the soundness-only quality control policy of open access mega journals linked to a higher rate of published errors? Scientometrics, 120(2):917–923, August 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03153-5.

Eslami:2013:ECN

[EES13] Hamidreza Eslami, Ashkan Ebadi, and Andrea Schiffauerova. Effect of collaboration network structure on knowledge creation and technological performance: the case of biotechnology in Canada. Scientometrics, 97(1):99–119, October 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1069-6.

Erren:2016:RMW

[EG16]  Thomas C. Erren and J. Valérie Groß.  Research

metrics: What about weighted citations? Scientometrics, 107(1):315–316, April 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-1841-5;   http://link.springer.
com/article/10.1007/s11192-016-1841-5.

ElAichouchi:2018:DRJ

[EG18] Adil El Aichouchi and Philippe Gorry. Delayed recognition of Judah Folkman’s hypothesis on tumor angiogenesis: when a Prince awakens a Sleeping Beauty by selfcitation. Scientometrics, 116(1):385–399, July 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2752-4.

Egghe:2010:CSS

[Egg10a] L. Egghe. Characteristic scores and scales in a Lotkaian framework. Scientometrics, 83(2):455–462, May 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0009-y.

Egghe:2010:DUF

[Egg10b] L. Egghe. The distribution of the uncitedness factor and its functional relation with the impact factor. Scientometrics, 83(3):689–695, June 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0130-y.

Egghe:2010:MSI

[Egg10c] L. Egghe. A model showing the increase in time of the average and median reference age and the decrease in time of the price index. Scientometrics, 82(2):243–248, February 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0057-3.

Egghe:2010:RBS

[Egg10d] L. Egghe. On the relation between Schubert’s h-index of a single paper and its total number of received citations. Scientometrics, 84(1):115–117, July 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0062-6.

Egghe:2010:LER

[Egg10e] Leo Egghe. Letter to the editor: On Randić’s Hsequence. Scientometrics, 84(3):795–797, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-009-0110-2. See [Ran09].

Egghe:2011:IFR

[Egg11a] L. Egghe. The impact factor rank-order distribution revisited. Scientometrics, 87(3):683–685, June 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0338-5.

Egghe:2011:IRR

[Egg11b] L. Egghe. The influence of random removal of sources and items on the h-index. Scientometrics, 88(2):363–370, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0411-0.

Egghe:2011:MDS

[Egg11c] L. Egghe. Mathematical derivation of the scale-dependence of the h-index and other h-type indices. Scientometrics, 87 (2):287–292, May 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0327-0.

Egghe:2011:PNS

[Egg11d] L. Egghe. Problems with “natural selection of academic papers”. Scientometrics, 88(2):663–667, August 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0395-9.

Egghe:2011:SPIa

[Egg11e] L. Egghe. The single publication H-index and the indirect H-index of a researcher. Scientometrics, 88(3):1003– 1004, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-011-
0417-7.

Egghe:2011:SPIb

[Egg11f] L. Egghe. The single publication H-index of papers in the Hirsch-core of a researcher and the indirect Hindex. Scientometrics, 89(3):727–739, December 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0483-x.

Egghe:2013:CIC

[Egg13a] L. Egghe. On the correction of the h-index for career length. Scientometrics, 96(2):563–571, August 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0926-z.

Egghe:2013:RRB

[Egg13b] L. Egghe. A rationale for the relation between the citer h-index and the classical h-index of a researcher. Scientometrics, 94(3):873–876, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0770-1.

Egghe:2013:TJC

[Egg13c] L. Egghe. Theoretical justification of the central area indices and the central interval indices. Scientometrics, 95(1):25– 34, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0803-9.

Egghe:2014:CPJ

[Egg14a] Leo Egghe. Comments on the paper of J. M. Campanario: The effect of citations on the significance of decimal places in the computation of journal impact factors. Scientometrics, 101(3):2071–2075, December 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1412-6. See [Cam14].

Egghe:2014:CYB

[Egg14b] Leo Egghe. Comments on “Year-based h-type indicators”. Scientometrics, 98(3):2369–2370, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-013-1186-2. See [MR13].

Egghe:2013:MCA

[EGR13] Leo Egghe, Raf Guns, and Ronald Rousseau. Measuring co-authors’ contribution to an article’s visibility. Scientometrics, 95(1):55–67, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0832-4.

Etxebarria:2010:USG

[EGU10] Goio Etxebarria and Mikel Gomez-Uranga. Use of Scopus and Google Scholar to measure social sciences production in four major Spanish universities. Scientometrics, 82(2): 333–349, February 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0043-9.

Etxebarria:2012:TSO

[EGUB12] Goio Etxebarria, Mikel Gomez-Uranga, and Jon Barrutia. Tendencies in scientific output on carbon nanotubes and graphene in global centers of excellence for nanotechnology. Scientometrics, 91(1):253–268, April 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0617-9.

Evans:2012:UPI

[EHK12] T. S. Evans, N. Hopkins, and B. S. Kaube. Universality of performance indicators based on citation and reference counts. Scientometrics, 93(2):473–495, November 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0694-9.

Elango:2019:ARI

[EKR19] Bakthavachalam Elango, Marcin Kozak, and Periyaswamy Rajendran. Analysis of retractions in Indian science. Scientometrics, 119(2):1081–1094, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03079-y.

Eldakar:2019:WRI

[Eld19] Metwaly Ali Mohamed Eldakar. Who reads international Egyptian academic articles? An altmetrics analysis of Mendeley readership categories. Scientometrics, 121(1): 105–135, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03189-7.

Ellegaard:2018:ABA

[Ell18] Ole Ellegaard. The application of bibliometric analysis: disciplinary and user aspects. Scientometrics, 116(1):181–202, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2765-z.

Evans:2011:CSP

[ELP11] T. S. Evans, R. Lambiotte, and P. Panzarasa. Community structure and patterns of scientific collaboration in business and management. Scientometrics, 89(1):381–396, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0439-1.

Elkins:2010:CBJ

[EMH^{+}10] Mark R. Elkins, Christopher G. Maher, Robert D. Herbert, Anne M. Moseley, and Catherine Sherrington. Correlation between the Journal Impact Factor and three other journal citation indices. Scientometrics, 85(1):81–93, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0262-0.

Emmer:2019:CBI

[Emm19] Adam Emmer. The careers behind and the impact of solo author articles in Nature and Science. Scientometrics, 120 (2):825–840, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03145-5.

Erdi:2013:PET

[ÉMS^{+}13] Péter Érdi, Kinga Makovi, Zoltán Somogyvári, Katherine Strandburg, Jan Tobochnik, Péter Volf, and László Zalányi. Prediction of emerging technologies based on analysis of the US patent citation network. Scientometrics, 95(1):225–242, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0796-4.

Ebrahimy:2016:PAR

[EMSH16] Saeideh Ebrahimy, Jafar Mehrad, Fatemeh Setareh, and Massoud Hosseinchari. Path analysis of the relationship between visibility and citation: the mediating roles of save, discussion, and recommendation metrics. Scientometrics, 109(3):1497–1510, December 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2130-z.

Ena:2016:MTT

[EMSS16] Oleg Ena, Nadezhda Mikova, Ozcan Saritas, and Anna Sokolova. A methodology for technology trend monitoring: the case of semantic technologies. Scientometrics, 108(3):1013–1041, September 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2024-0.

Erfanian:2017:SOL

[EN17] Elham Erfanian and Amir B. Ferreira Neto. Scientific output: labor or capital intensive? An analysis for selected countries. Scientometrics, 112(1):461–482, July 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Erfanmanesh:2019:WCB

[ENA19] Mohammadamin Erfanmanesh, A. Noorhidawati, and A. Abrizah. What can Bookmetrix tell us about the impact of Springer Nature’s books. Scientometrics, 121(1):521–536, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03198-6.

Erdt:2016:AAS

[ENST16] Mojisola Erdt, Aarthy Nagarajan, Sei-Ching Joanna Sin, and Yin-Leng Theng. Altmetrics: an analysis of the stateof-the-art in measuring research impact on social media. Scientometrics, 109(2):1117–1166, November 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2077-0.

Ebrahimy:2014:DVR

[EO14] Saeideh Ebrahimy and Farideh Osareh. Design, validation, and reliability determination a citing conformity instrument at three levels: normative, informational, and identification. Scientometrics, 99(2):581–597, May 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1188-0.

Engels:2012:CPP

[EOS12] Tim C. E. Engels, Truyken L. B. Ossenblok, and Eric H. J. Spruyt. Changing publication patterns in the social sciences and humanities, 2000–2009. Scientometrics, 93(2):

373–390, November 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0680-2.

Egghe:2012:TPS

[ER12] Leo Egghe and Ronald Rousseau. Theory and practice of the shifted Lotka function. Scientometrics, 91(1):295–301, April 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0539-y.

Egghe:2019:GRB

[ER19a] Leo Egghe and Ronald Rousseau. A geometric relation between the h-index and the Lorenz curve. Scientometrics, 119 (2):1281–1284, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03083-2.

Egghe:2019:MLT

[ER19b] Leo Egghe and Ronald Rousseau. Measures of linear type lead to a characterization of Zipf functions. Scientometrics, 121(3):1707–1715, December 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03257-y.

Eckmann:2012:RBH

[ERW12] Michael Eckmann, Anderson Rocha, and Jacques Wainer. Relationship between high-quality journals and conferences in computer vision. Scientometrics, 90(2):617–630, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0527-2.

Ebadi:2016:HBS

[ES16a] Ashkan Ebadi and Andrea Schiffauerova. How to boost scientific production? A statistical analysis of research funding and other influencing factors. Scientometrics, 106(3): 1093–1116, March 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1825-x.

Ebadi:2016:IIA

[ES16b] Ashkan Ebadi and Andrea Schiffauerova. iSEER: an intelligent automatic computer system for scientific evaluation of researchers. Scientometrics, 107(2):477–498, May 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1852-2.

Eskrootchi:2018:CMR

[ES18] Rogheyeh Eskrootchi and Nadia Sanee. Comparison of medical research performance by thermodynamic and citation analysis methods. Scientometrics, 117(3):2159–2168, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2930-4.

Echeverria:2015:MTD

[ESB15] Mercedes Echeverria, David Stuart, and Tobias Blanke. Medical theses and derivative articles: dissemination of contents and publication patterns. Scientometrics, 102(1): 559–586, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1442-0.

Erman:2015:EME

[ET15] Nusa Erman and Ljupco Todorovski. The effects of measurement error in case of scientific network analysis. Scientometrics, 104(2):453–473, August 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1615-5.

Eto:2013:ECB

[Eto13] Masaki Eto. Evaluations of context-based co-citation searching. Scientometrics, 94(2):651–673, February 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0756-z.

Etzkowitz:2013:MDD

[Etz13a] Henry Etzkowitz. Mistaking dawn for dusk: quantophrenia and the cult of numerology in technology transfer analysis. Scientometrics, 97(3):913–925, December 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1007-7.

Etzkowitz:2013:PSH

[Etz13b] Henry Etzkowitz. Paula Stephan: How economics shapes science and how science shapes the economy. Scientometrics, 96(3):941–946, September 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
013-0972-1.

Ellegaard:2015:BAS

[EW15] Ole Ellegaard and Johan A. Wallin. The bibliometric analysis of scholarly production: How great is the impact? Scientometrics, 105(3):1809–1831, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com//content/
pdf/10.1007/s11192-015-1645-z.pdf;   http://link.
springer.com/content/pdf/10.1007/s11192-015-1645-
z.pdf.

Fu:2010:UCB

[FA10] Lawrence D. Fu and Constantin F. Aliferis. Using contentbased and bibliometric features for machine learning models to predict citation counts in the biomedical literature. Scientometrics, 85(1):257–270, October 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0160-5.

Fu:2013:CMI

[FAA13] Lawrence D. Fu, Yindalon Aphinyanaphongs, and Constantin F. Aliferis. Computer models for identifying instrumental citations in the biomedical literature. Scientometrics, 97(3):871–882, December 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0983-y.

Farooq:2018:MPC

[FAI^{+}18] Muhammad Farooq, Muhammad Asim, Muhammad Imran, Shahid Imran, Jameel Ahmad, and Muhammad Rizwan Younis. Mapping past, current and future energy research trend in Pakistan: a scientometric assessment. Scientometrics, 117(3):1733–1753, December 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2939-8.

Fang:2011:PRC

[Fan11] Hui Fang. Peer review and over-competitive research funding fostering mainstream opinion to monopoly. Scientometrics, 87(2):293–301, May 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
010-0323-4.

Fanelli:2012:NRD

[Fan12] Daniele Fanelli. Negative results are disappearing from most disciplines and countries. Scientometrics, 90(3):891–904, March 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0494-7.

Fanelli:2013:PBT

[Fan13a] Daniele Fanelli. Any publicity is better than none: newspaper coverage increases citations, in the UK more than in Italy. Scientometrics, 95(3):1167–1177, June 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0925-0.

Fanelli:2013:PRR

[Fan13b] Daniele Fanelli. Positive results receive more citations, but only in some disciplines. Scientometrics, 94(2):701–709, February 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0757-y.

Fan:2015:CIR

[Fan15a] Wenqiang Fan. Contribution of the institutional repositories of the Chinese Academy of Sciences to the webometric indicators of their home institutions. Scientometrics, 105(3): 1889–1909, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1758-4.

Fang:2015:VSE

[Fan15b] Yuqing Fang. Visualizing the structure and the evolving of digital medicine: a scientometrics review. Scientometrics, 105(1):5–21, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1696-1.

Fang:2018:AVT

[Fan18] Hui Fang. Analysing the variation tendencies of the numbers of yearly citations for sleeping beauties in science by using derivative analysis. Scientometrics, 115(2):1051–1070, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2687-9.

Fang:2019:TSC

[Fan19] Hui Fang. A transition stage co-citation criterion for identifying the awakeners of sleeping beauty publications. Scientometrics, 121(1):307–322, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03195-9.

Feeney:2010:WSN

[FB10] Mary K. Feeney and Margarita Bernal. Women in STEM networks: who seeks advice and support from women scientists? Scientometrics, 85(3):767–790, December 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0256-y.

Finardi:2016:SCF

[FB16] Ugo Finardi and Andrea Buratti. Scientific collaboration framework of BRICS countries: an analysis of international coauthorship. Scientometrics, 109(1):433–446, October 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-1927-0.

Fernandez-Cano:2018:QSR

[FCCMTRVR18] Antonio Fernández-Cano, Elvira Curiel-Marin, Manuel Torralbo-Rodríguez, and Mónica Vallejo-Ruiz. Questioning the Shanghai Ranking methodology as a tool for the evaluation of universities: an integrative review. Scientometrics, 116(3):2069–2083, September 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2814-7.

Fernandez-Cano:2017:MME

[FCFG17] Antonio Fernandez-Cano and Inés M. Fernández-Guerrero. A multivariate model for evaluating emergency medicine journals. Scientometrics, 110(2):991–1003, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2197-6.

Fernandez-Cano:2012:TSS

[FCTV12] Antonio Fernández-Cano, Manuel Torralbo, and Mónica Vallejo. Time series of scientific growth in Spanish doctoral theses (1848–2009). Scientometrics, 91(1):15–36, April 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0572-x.

Fu:2011:CRC

[FCWH11] Hui-Zhen Fu, Kun-Yang Chuang, Ming-Huang Wang, and Yuh-Shan Ho. Characteristics of research in China assessed with essential science indicators. Scientometrics, 88(3):841– 862, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0416-8.

Fry:2014:EAA

[FD14] Timothy D. Fry and Joan M. Donohue. Exploring the author affiliation index. Scientometrics, 98(3):1647–1667, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1133-2.

Filser:2017:SRF

[FdSdO17] Lukas D. Filser, Fábio Francisco da Silva, and Otávio José de Oliveira. State of research and future research tendencies in lean healthcare: a bibliometric analysis. Scientometrics, 112(2):799–816, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
017-2409-8.

Fuccella:2016:ICA

[FDVZ16] Vittorio Fuccella, Domenico De Stefano, Maria Prosperina Vitale, and Susanna Zaccarin. Improving co-authorship network structures by combining multiple data sources: evidence from Italian academic statisticians. Scientometrics, 107(1):167–184, April 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1872-y.

Fragkiadaki:2014:RIC

[FE14] Eleni Fragkiadaki and Georgios Evangelidis. Review of the indirect citations paradigm: theory and practice of the assessment of papers, authors and journals. Scientometrics, 99(2):261–288, May 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1175-5.

Fragkiadaki:2016:ETN

[FE16a] Eleni Fragkiadaki and Georgios Evangelidis. Erratum to: Three novel indirect indicators for the assessment of papers and authors based on generations of citations. Scientometrics, 108(2):1011, August 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-1874-9.pdf. See [FE16b].

Fragkiadaki:2016:TNI

[FE16b] Eleni Fragkiadaki and Georgios Evangelidis. Three novel indirect indicators for the assessment of papers and authors based on generations of citations. Scientometrics, 106(2):657–694, February 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-

015-1802-4. See erratum [FE16a].

Fedderke:2013:ONR

[Fed13] J. W. Fedderke. The objectivity of national research foundation peer review in South Africa assessed against bibliometric indexes. Scientometrics, 97(2):177–206, November 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-0981-0.

Frandsen:2019:FPL

[FEHC19] Tove Faber Frandsen, Mette Brandt Eriksen, David Mortan Grøne Hammer, and Janne Buck Christensen. Fragmented publishing: a large-scale study of health science. Scientometrics, 119(3):1729–1743, June 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03109-9.

Fernandes:2014:ATS

[Fer14] João M. Fernandes. Authorship trends in software engineering. Scientometrics, 101(1):257–271, October 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1331-6.

Fragkiadaki:2011:VMA

[FESD11] Eleni Fragkiadaki, Georgios Evangelidis, Nikolaos Samaras, and Dimitris A. Dervos. f-value: measuring an article’s scientific impact. Scientometrics, 86(3):671–686, March 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0302-9.

Fernandez:2016:PDS

[FFL16] A. Fernández, E. Ferrándiz, and M. D. León. Proximity dimensions and scientific collaboration among academic institutions in Europe: The closer, the better? Scientometrics, 106(3):1073–1092, March 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1819-8.

Ferreira:2016:CCB

[FFR16] João José M. Ferreira, Cristina I. Fernandes, and Vanessa Ratten. A co-citation bibliometric analysis of strategic management research. Scientometrics, 109(1):1–32, October 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2008-0.

Fernandes:2017:DCP

[FFR^{+}17] Cristina Fernandes, João J. Ferreira, Mário L. Raposo, Cristina Estevão, Marta Peris-Ortiz, and Carlos Rueda-Armengot. The dynamic capabilities perspective of strategic management: a co-citation analysis. Scientometrics, 112 (1):529–555, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Fay:2015:SSG

[FG15] Stéphane Fay and Sébastien Gautrias. A scientometric study of general relativity and quantum cosmology from 2000 to 2012. Scientometrics, 105(1):471–484, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1674-7.

Franceschini:2012:SIA

[FGMM12] Fiorenzo Franceschini, Maurizio Galetto, Domenico Maisano, and Luca Mastrogiacomo. The success-index: an alternative approach to the h-index for evaluating an individual’s research output. Scientometrics, 92(3):621–641, September 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0570-z.

Fumani:2013:ITI

[FGP13] Mohammad Reza Falahati Qadimi Fumani, Marzieh Goltaji, and Pardis Parto. Inconsistent transliteration of Iranian university names: a hazard to Iran’s ranking in ISI Web of Science. Scientometrics, 95(1):371–384, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0818-2.

Fu:2013:CIR

[FH13] Hui-Zhen Fu and Yuh-Shan Ho. Comparison of independent research of China’s top universities using bibliometric indicators. Scientometrics, 96(1):259–276, July 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0912-5.

Fu:2016:HCA

[FH16] Hui-Zhen Fu and Yuh-Shan Ho. Highly cited Antarctic articles using Science Citation Index Expanded: a bibliometric analysis. Scientometrics, 109(1):337–357, October 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-1992-4.

Fukuzawa:2016:SLB

[FI16] Naomi Fukuzawa and Takanori Ida. Science linkages between scientific articles and patents for leading scientists in the life and medical sciences field: the case of Japan. Scientometrics, 106(2):629–644, February 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1795-z.

Fiala:2011:MCI

[Fia11] Dalibor Fiala. Mining citation information from CiteSeer data. Scientometrics, 86(3):553–562, March 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0326-1.

Fields:2015:CEC

[Fie15a] Chris Fields. Close to the edge: co-authorship proximity of Nobel laureates in Physiology or Medicine, 1991–2010, to cross-disciplinary brokers. Scientometrics, 103(1):267–299, April 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1526-5.

Fields:2015:CAP

[Fie15b] Chris Fields. Co-authorship proximity of A. M. Turing Award and John von Neumann Medal winners to the disciplinary boundaries of computer science. Scientometrics, 104(3):809–825, September 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1575-9.

Fields:2015:HSC

[Fie15c] Chris Fields. How small is the center of science? Short cross-disciplinary cycles in co-authorship graphs. Scientometrics, 102(2):1287–1306, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1468-3.

Finardi:2011:TRB

[Fin11] Ugo Finardi. Time relations between scientific production and patenting of knowledge: the case of nanotechnologies. Scientometrics, 89(1):37–50, October 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0443-5.

Finardi:2015:SCB

[Fin15] Ugo Finardi. Scientific collaboration between BRICS countries. Scientometrics, 102(2):1139–1166, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-014-1490-5.

Fakhree:2011:SAM

[FJ11] Mohammad A. Abolghassemi Fakhree and Abolghasem Jouyban. Scientometric analysis of the major Iranian medical universities. Scientometrics, 87(1):205–220, April 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0336-z.

Fell:2016:TGD

[FK16] Clemens B. Fell and Cornelius J. König. Is there a gender difference in scientific collaboration? A scientometric examination of co-authorships among industrial-organizational psychologists. Scientometrics, 108(1):113–141, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-016-1967-5.

Fursov:2017:HAT

[FK17] Konstantin Fursov and Alina Kadyrova. How the analysis of transitionary references in knowledge networks and their centrality characteristics helps in understanding the genesis of growing technology areas. Scientometrics, 111(3): 1947–1963, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). See correction [FK18].

Fursov:2018:CHA

[FK18] Konstantin Fursov and Alina Kadyrova. Correction to: How the analysis of transitionary references in knowledge networks and their centrality characteristics helps in understanding the genesis of growing technology areas. Scientometrics, 117(2):1313, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2779-6;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2779-

6.pdf. See [FK17].

Ferligoj:2015:SCD

[FKM^{+}15] Anuska Ferligoj, Luka Kronegger, Franc Mali, Tom A. B. Snijders, and Patrick Doreian. Scientific collaboration dynamics in a national scientific system. Scientometrics, 104(3):985–1012, September 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
015-1585-7.

Fink:2014:SKP

[FKRS14] Daniel Fink, Youngsun Kwon, Jae Jeung Rho, and Minho So. S&T knowledge production from 2000 to 2009 in two periphery countries: Brazil and South Korea. Scientometrics, 99(1):37–54, April 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1085-6.

Fan:2016:OSR

[FL16] Wenqiang Fan and Qinghui Liu. Open scholarship ranking of Chinese research universities. Scientometrics, 108 (2):673–691, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1983-5.

Feeley:2011:PFJ

[FLB11] Thomas Hugh Feeley, Katherine Hart LaVail, and George A. Barnett. Predicting faculty job centrality in communication. Scientometrics, 87(2):303–314, May 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0324-3.

Frigyesi:2019:EPI

[FLB19] Veronika Frigyesi, Patrice Laget, and Mark Boden. Exploitation of patent information in R&D output analysis for policymaking. Scientometrics, 121(3):1717–1736, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03236-3.

Fu:2014:CRC

[FLH14] Hui-Zhen Fu, Xiao Long, and Yuh-Shan Ho. China’s research in chemical engineering journals in Science Citation Index expanded: a bibliometric analysis. Scientometrics, 98 (1):119–136, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1047-z.

Ferrara:2016:LBP

[FLM16] M. Ferrara, F. Lamperti, and R. Mavilia. Looking for best performers: a pilot study towards the evaluation of science parks. Scientometrics, 106(2):717–750, February 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1804-2.

Feliciani:2019:SRS

[FLM^{+}19] Thomas Feliciani, Junwen Luo, Lai Ma, Pablo Lucas, Flaminio Squazzoni, Ana Marušić, and Kalpana Shankar. A scoping review of simulation models of peer review. Scientometrics, 121(1):555–594, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03205-w;   http://link.springer.
com/content/pdf/10.1007/s11192-019-03205-w.pdf.

Fan:2017:SLT

[FLZ17] Xia Fan, Wenjie Liu, and Guilong Zhu. Scientific linkage and technological innovation capabilities: international comparisons of patenting in the solar energy industry. Scientometrics, 111(1):117–138, April 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-017-2274-5.

Franceschini:2011:BPS

[FM11a] Fiorenzo Franceschini and Domenico Maisano. Bibliometric positioning of scientific manufacturing journals: a comparative analysis. Scientometrics, 86(2):463–485, February 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0301-x.

Franceschini:2011:CI

[FM11b] Fiorenzo Franceschini and Domenico Maisano. Criticism on the hg-index. Scientometrics, 86(2):339–346, February 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0261-1.

Franceschini:2011:ABE

[FM11c] Fiorenzo Franceschini and Domenico Maisano. On the analogy between the evolution of thermodynamic and bibliometric systems: a breakthrough or just a bubble? Scientometrics, 89(1):315–327, October 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0444-4.

Franceschini:2011:PER

[FM11d] Fiorenzo Franceschini and Domenico Maisano. Proposals for evaluating the regularity of a scientist’s research output. Scientometrics, 88(1):279–295, July 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0371-4. See comments [Pra11c].

Franceschini:2012:PPA

[FM12] Fiorenzo Franceschini and Domenico Maisano. Publication and patent analysis of European researchers in the field of production technology and manufacturing systems. Scientometrics, 93(1):89–100, October 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0648-2.

Fernandes:2017:ENA

[FM17] João M. Fernandes and Miguel P. Monteiro. Evolution in the number of authors of computer science publications. Scientometrics, 110(2):529–539, February 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2214-9.

Franceschini:2013:EDD

[FMM13a] Fiorenzo Franceschini, Domenico Maisano, and Luca Mastrogiacomo. The effect of database dirty data on h-index calculation. Scientometrics, 95(3):1179–1188, June 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0871-x.

Franceschini:2013:ERI

[FMM13b] Fiorenzo Franceschini, Domenico Maisano, and Luca Mastrogiacomo. Evaluating research institutions: the potential of the success-index. Scientometrics, 96(1):85–101, July 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0887-2.

Franceschini:2014:CSI

[FMM14] Fiorenzo Franceschini, Domenico Maisano, and Luca Mastrogiacomo. The citer-success-index: a citer-based indicator to select a subset of elite papers. Scientometrics, 101(2): 963–983, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1250-6.

Franceschini:2015:EDI

[FMM15a] Fiorenzo Franceschini, Domenico Maisano, and Luca Mastrogiacomo. Errors in DOI indexing by bibliometric databases. Scientometrics, 102(3):2181–2186, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-014-1503-4.

Franceschini:2015:IOC

[FMM15b] Fiorenzo Franceschini, Domenico Maisano, and Luca Mastrogiacomo. Influence of omitted citations on the bibliometric statistics of the major manufacturing journals. Scientometrics, 103(3):1083–1122, June 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1583-9.

Franceschini:2016:DSW

[FMM16] Fiorenzo Franceschini, Domenico Maisano, and Luca Mastrogiacomo. Do Scopus and WoS correct “old” omitted citations? Scientometrics, 107(2):321–335, May 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1867-8.

Ferrara:2017:EKP

[FMP17a] Massimiliano Ferrara, Roberto Mavilia, and Bruno Antonio Pansera. Extracting knowledge patterns with a social network analysis approach: an alternative methodology for assessing the impact of power inventors. Scientometrics, 113(3):1593–1625, December 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2536-2.

Figuerola:2017:MEL

[FMP17b] Carlos G. Figuerola, Francisco Javier García Marco, and María Pinto. Mapping the evolution of library and information science (1978–2014) using topic modeling on LISA. Scientometrics, 112(3):1507–1535, September 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2432-9.

Franceschini:2010:AII

[FMPP10] Fiorenzo Franceschini, Domenico Maisano, Anna Perotti, and Andrea Proto. Analysis of the ch-index: an indicator to evaluate the diffusion of scientific research output by citers. Scientometrics, 85(1):203–217, October 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0165-0.

Fiala:2017:RHE

[FMS17] Jaroslav Fiala, Jirí J. Mares, and Jaroslav Sesták. Reflections on how to evaluate the professional value of scientific papers and their corresponding citations. Scientometrics, 112(1):697–709, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Faria:2016:HCC

[FMU16] João R. Faria, Franklin G. Mixon, Jr., and Kamal P. Upadhyaya. Human capital, collegiality, and stardom in economics: empirical analysis. Scientometrics, 106(3):917–943, March 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1835-3.

Fox:2017:DRR

[Fox17] Charles W. Fox. Difficulty of recruiting reviewers predicts review scores and editorial decisions at six journals of ecology and evolution. Scientometrics, 113(1):465–477, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2489-5.

Ferreira:2018:BAC

[FP18] Rafael Henrique Mainardes Ferreira and Claudia Tania Picinin. Bibliometric analysis for characterization of oil production in Brazilian territory. Scientometrics, 116(3):1945– 1974, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2829-0.

Ferreira:2014:TCT

[FPS14] Manuel Portugal Ferreira, Cláudia Frias Pinto, and Fernando Ribeiro Serra. The transaction costs theory in international business research: a bibliometric study over three decades. Scientometrics, 98(3):1899–1922, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1172-8.

Frigotto:2011:FSC

[FR11] M. Laura Frigotto and Massimo Riccaboni. A few special cases: scientific creativity and network dynamics in the field of rare diseases. Scientometrics, 89(1):397–420, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0431-9.

Franceschet:2010:CBI

[Fra10] Massimo Franceschet. A comparison of bibliometric indicators for computer science scholars and journals on Web of Science and Google Scholar. Scientometrics, 83(1):243–258, April 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0021-2.

Franses:2014:TTD

[Fra14] Philip Hans Franses. Trends in three decades of rankings of Dutch economists. Scientometrics, 98(2):1257–1268, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1041-5.

Frandsen:2017:PJU

[Fra17] Tove Faber Frandsen. Are predatory journals undermining the credibility of science? A bibliometric analysis of citers. Scientometrics, 113(3):1513–1528, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2520-x.

Foronda-Robles:2016:IDC

[FRdA16] Concepción Foronda-Robles and Luis Galindo-Pérez de Azpillaga. From initial dissemination to consolidated impact: the concept of crisis in the field of tourism. Scientometrics, 109(1):261–281, October 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2016-0.

Freyer:2014:RR

[Fre14] Leo Freyer. Robust rankings. Scientometrics, 100(2): 391–406, August 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-014-1313-
8.pdf.

Firdaus:2019:RBB

[FRF^{+}19] Ahmad Firdaus, Mohd Faizal Ab Razak, Ali Feizollah, Ibrahim Abaker Targio Hashem, Mohamad Hazim, and Nor Badrul Anuar. The rise of “blockchain”: bibliometric analysis of blockchain study. Scientometrics, 120(3):1289– 1331, September 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03170-4.

Ferreira:2017:SLA

[FRPP17] Manuel Portugal Ferreira, Nuno R. Reis, Roberta M. Paula, and Claudia Frias Pinto. Structural and longitudinal analysis of the knowledge base on spin-off research. Scientometrics, 112(1):289–313, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Frietsch:2010:TPI

[FS10] Rainer Frietsch and Ulrich Schmoch. Transnational patents and international markets. Scientometrics, 82(1):185–200, January 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0082-2.

Ferrara:2012:TCM

[FS12] Alfio Ferrara and Silvia Salini. Ten challenges in modeling bibliographic data for bibliometric analysis. Scientometrics, 93(3):765–785, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0810-x.

Fisher:2010:PVN

[FSAB10] Erik Fisher, Catherine P. Slade, Derrick Anderson, and Barry Bozeman. The public value of nanotechnology?

Scientometrics, 85(1):29–39, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0237-1.

Fu:2014:ASN

[FSC14] Tom Z. J. Fu, Qianqian Song, and Dah Ming Chiu. The academic social network. Scientometrics, 101(1):203–239, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1356-x.

Franzoni:2010:UCA

[FSLR10] Chiara Franzoni, Christopher L. Simpkins, Baoli Li, and Ashwin Ram. Using content analysis to investigate the research paths chosen by scientists over time. Scientometrics, 83(1):321–335, April 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0061-7.

Furukawa:2011:QAC

[FSO11] Takao Furukawa, Nobuyuki Shirakawa, and Kumi Okuwada. Quantitative analysis of collaborative and mobility networks. Scientometrics, 87(3):451–466, June 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0360-7.

Furukawa:2012:IMR

[FSOS12] Takao Furukawa, Nobuyuki Shirakawa, Kumi Okuwada, and Kazuya Sasaki. International mobility of researchers in robotics, computer vision and electron devices: A quantitative and comparative analysis. Scientometrics, 91(1): 185–202, April 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0545-0.

Ferrer-Sapena:2015:MPW

[FSSPG^{+}15] A. Ferrer-Sapena, E. A. Sánchez-Pérez, L. M. González, F. Peset, and R. Aleixandre-Benavent. Mathematical properties of weighted impact factors based on measures of prestige of the citing journals. Scientometrics, 105(3):2089– 2108, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1741-0.

Frijters:2019:IPR

[FT19] Paul Frijters and Benno Torgler. Improving the peer review process: a proposed market system. Scientometrics, 119 (2):1285–1288, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03076-1.

Fukuzawa:2014:EAR

[Fuk14] Naomi Fukuzawa. An empirical analysis of the relationship between individual characteristics and research productivity. Scientometrics, 99(3):785–809, June 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1213-3.

Fukugawa:2016:KCD

[Fuk16] Nobuya Fukugawa. Knowledge creation and dissemination by Kosetsushi in sectoral innovation systems: insights from patent data. Scientometrics, 109(3):2303–2327, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2124-x.

Fukuzawa:2017:CPP

[Fuk17] Naomi Fukuzawa. Characteristics of papers published in journals: an analysis of open access journals, country of publication, and languages used. Scientometrics, 112(2): 1007–1023, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2414-y.

Fukugawa:2019:DIP

[Fuk19] Nobuya Fukugawa. Determinants and impacts of public agricultural research: product-level evidence from agricultural Kohsetsushi in Japan. Scientometrics, 120(3):1475– 1498, September 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03158-0.

Fatt:2010:SCB

[FUR10] Choong Kwai Fatt, Ephrance Abu Ujum, and Kuru Ratnavelu. The structure of collaboration in the Journal of Finance. Scientometrics, 85(3):849–860, December 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0254-0.

Flores-Vargas:2018:DEM

[FVVSGM^{+}18] Xochitl Flores-Vargas, Silvano Habrajam Vitar-Sandoval, Jazmín Ivonne Gutiérrez-Maya, Pavel Collazo-Rodríguez, and Francisco Collazo-Reyes. Determinants of the emergence of modern scientific knowledge in mineralogy (Mexico, 1975–1849): a geohistoriometric approach. Scientometrics, 115(3):1505–1515, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2646-5.

Faria:2018:RIH

[FWFM18] João Ricardo Faria, Peter F. Wanke, João J. Ferreira, and Franklin G. Mixon, Jr. Research and innovation in higher education: empirical evidence from research and patenting in Brazil. Scientometrics, 116(1):487–504, July 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2744-4.

Fan:2015:DRA

[FYC15] Xia Fan, Xiaowan Yang, and Liming Chen. Diversified resources and academic influence: patterns of universityindustry collaboration in Chinese research-oriented universities. Scientometrics, 104(2):489–509, August 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1618-2.

Fang:2017:OCA

[FZQ17] Chencheng Fang, Jiantong Zhang, and Wei Qiu. Online classified advertising: a review and bibliometric analysis. Scientometrics, 113(3):1481–1511, December 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2524-6.

Fu:2011:BAC

[FZZ^{+}11] Jun-Ying Fu, Xu Zhang, Yun-Hua Zhao, Mu-Hsuan Huang, and Dar-Zen Chen. Bibliometric analysis of complementary and alternative medicine research over three decades. Scientometrics, 88(2):617–626, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0391-0.

Fu:2012:GPT

[FZZ^{+}12a] Jun-Ying Fu, Xu Zhang, Yun-Hua Zhao, Dar-Zen Chen, and Mu-Hsuan Huang. Global performance of traditional Chinese medicine over three decades. Scientometrics, 90 (3):945–958, March 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0521-8.

Fu:2012:SPC

[FZZ^{+}12b] Jun-Ying Fu, Xu Zhang, Yun-Hua Zhao, He-Feng Tong, Dar-Zen Chen, and Mu-Hsuan Huang. Scientific production and citation impact: a bibliometric analysis in acupuncture over three decades. Scientometrics, 93(3):1061–1079, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0737-2.

Feng:2017:ICW

[FZZ17] Jia Feng, Yun Qiu Zhang, and Hao Zhang. Improving the coword analysis method based on semantic distance. Scientometrics, 111(3):1521–1531, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Glanzel:2018:GNS

[GA18] Wolfgang Glänzel and Mehmet Ali Abdulhayoğlu. Garfield number: on some characteristics of Eugene Garfield’s first and second order co-authorship networks. Scientometrics, 114(2):533–544, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2623-4.

Goldberg:2015:MCN

[GAE15] S. R. Goldberg, H. Anthony, and T. S. Evans. Modelling citation networks. Scientometrics, 105(3):1577–1604, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1737-9.

Gupta:2015:WCR

[GAGT15] B. M. Gupta, K. K. Mueen Ahmed, Ritu Gupta, and Rishi Tiwari. World camel research: a scientometric assessment, 2003–2012. Scientometrics, 102(1):957–975, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-014-1405-5.

Galam:2011:TBA

[Gal11] Serge Galam. Tailor based allocations for multiple authorship: a fractional gh-index. Scientometrics, 89(1):365–379, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0447-1.

Galvez:2017:AAS

[Gál17] Ramiro H. Gálvez. Assessing author self-citation as a mechanism of relevant knowledge diffusion. Scientometrics, 111 (3):1801–1812, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Gonzalez-Alcaide:2016:BII

[GALR16] Gregorio González-Alcaide, Pedro Llorente, and José M. Ramos. Bibliometric indicators to identify emerging research fields: publications on mass gatherings. Scientometrics, 109(2):1283–1298, November 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2083-2.

Gantman:2012:ELP

[Gan12] Ernesto R. Gantman. Economic, linguistic, and political factors in the scientific productivity of countries. Scientometrics, 93(3):967–985, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0736-3.

Gonzalez-Alcaide:2018:ERL

[GAPP18] Gregorio González-Alcaide and Inés Poveda-Pastor. Emerging roles in library and information science: consolidation in the scientific literature and appropriation by professionals of the discipline. Scientometrics, 116(1):319–337, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2766-y.

Garfield:2014:FE

[Gar14] Eugene Garfield. Farewell editorial. Scientometrics, 98 (1):1–2, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-013-
1153-y.

Garousi:2015:BAT

[Gar15] Vahid Garousi. A bibliometric analysis of the Turkish software engineering research community. Scientometrics, 105 (1):23–49, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1663-x.

Gautam:2017:OWS

[Gau17] Pitambar Gautam. An overview of the Web of Science record of scientific publications (2004–2013) from Nepal: focus on disciplinary diversity and international collaboration. Scientometrics, 113(3):1245–1267, December 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2538-0.

Gonzalez-Alcaide:2012:IFN

[GAVZAB12] Gregorio González-Alcaide, Juan Carlos Valderrama-Zurián, and Rafael Aleixandre-Benavent. The Impact Factor in non-English-speaking countries. Scientometrics, 92 (2):297–311, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0692-y.

Gupta:2012:SPO

[GB12] B. M. Gupta and Adarsh Bala. S&T publications output of Nepal: a quantitative analysis, 2001–10. Scientometrics, 93(3):1029–1046, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0778-6.

Gonzalez-Brambila:2014:SCA

[GB14a] Claudia N. Gonzalez-Brambila. Social capital in academia. Scientometrics, 101(3):1609–1625, December 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1424-2.

Gorjiara:2014:NNR

[GB14b] T. Gorjiara and C. Baldock. Nanoscience and nanotechnology research publications: a comparison between Australia and the rest of the world. Scientometrics, 100(1):121–148, July 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1287-6.

Gu:2016:RTA

[GB16] Xin Gu and Karen L. Blackmore. Recent trends in academic journal growth. Scientometrics, 108(2):693–716, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1985-3.

Gu:2017:CAJ

[GB17a] Xin Gu and Karen Blackmore. Characterisation of academic journals in the digital age. Scientometrics, 110(3):1333–1350, March 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2219-4.

Gu:2017:QSA

[GB17b] Xin Gu and Karen Blackmore. Quantitative study on Australian academic science. Scientometrics, 113(2):1009– 1035, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2499-3.

Geraci:2015:GIP

[GBB15] Lisa Geraci, Steve Balsis, and Alexander J. Busch Busch. Gender and the h index in psychology. Scientometrics, 105(3):2023–2034, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
015-1757-5.

Gonzalez-Betancor:2019:PMA

[GBDG19] Sara M. González-Betancor and Pablo Dorta-González. Publication modalities ‘article in press’ and ‘open access’ in relation to journal average citation. Scientometrics, 120(3):1209–1223, September 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03156-2.

Guilera:2013:MAP

[GBGB13] Georgina Guilera, Maite Barrios, and Juana Gómez-Benito. Meta-analysis in psychology: a bibliometric study. Scientometrics, 94(3):943–954, March 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0761-2.

Griffin:2016:CJA

[GBHT16] Darrin J. Griffin, San Bolkan, Jennifer L. Holmgren, and Frank Tutzauer. Central journals and authors in communication using a publication network. Scientometrics, 106 (1):91–104, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1774-4.

Glanzel:2016:DCU

[GBM^{+}16] Wolfgang Glänzel, Raphael Beck, Katrin Milzow, Stig Slipersæter, Gábor Tóth, Michal Kolodziejski, and Pei-Shan Chi. Data collection and use in research funding and performing organisations. General outlines and first results of a project launched by Science Europe. Scientometrics, 106(2):

825–835, February 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1815-z.

Guerrero-Bote:2014:RBD

[GBMA14] Vicente P. Guerrero-Bote and Félix Moya-Anegón. Relationship between downloads and citations at journal and paper levels, and the influence of language. Scientometrics, 101(2):1043–1065, November 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1243-5.

Gentil-Beccot:2010:CRB

[GBMB10] Anne Gentil-Beccot, Salvatore Mele, and Travis C. Brooks. Citing and reading behaviours in high-energy physics. Scientometrics, 84(2):345–355, August 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0111-1.

Glanzel:2015:CC

[GBSZL15] Wolfgang Glänzel, Tibor Braun, András Schubert, and Guido Zosimo-Landolfo. Coping with copying. Scientometrics, 102(1):1–3, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
014-1473-6.

Guan:2010:MMR

[GC10] Jiancheng Guan and Kaihua Chen. Modeling macro–R&D production frontier performance: an application to Chinese province-level R&D. Scientometrics, 82(1):165–173, January 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0030-1.

Garcia-Carpintero:2010:RNE

[GCGP10] Esther García-Carpintero, Begoña Granadino, and Luis M. Plaza. The representation of nationalities on the editorial boards of international journals and the promotion of the scientific output of the same countries. Scientometrics, 84(3):799–811, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0199-3.

Gao:2015:SAP

[GCLcG15] Wei Gao, Yan Chen, Yong Liu, and Huai cheng Guo. Scientometric analysis of phosphorus research in eutrophic lakes. Scientometrics, 102(3):1951–1964, March 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1500-7.

Gazni:2011:IDT

[GD11] Ali Gazni and Fereshteh Didegah. Investigating different types of research collaboration and citation impact: a case study of Harvard University’s publications. Scientometrics, 87(2):251–265, May 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0343-8.

Gazni:2016:RBA

[GD16] Ali Gazni and Fereshteh Didegah. The relationship between authors’ bibliographic coupling and citation exchange: analyzing disciplinary differences. Scientometrics, 107(2):609– 626, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1856-y.

Gomes:2017:DDS

[GD17] Janaína Gomes and Homero Dewes. Disciplinary dimensions and social relevance in the scientific communications on biofuels. Scientometrics, 110(3):1173–1189, March 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2233-6.

Guimaraes:2014:RCA

[GdA14] Jorge Almeida Guimarães and Elenara Chaves Edler de Almeida. Response to the comments on the article: Brazil’s growing production of scientific articles-how are we doing with review articles and other qualitative indicators? Scientometrics, 101(3):2079–2080, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-014-1413-5.

Gracio:2013:DSA

[GdOdAG^{+}13] Maria Cláudia Cabrini Gracio, Ely Francina Tannuri de Oliveira, Júlio de Araujo Gurgel, Maria Isabel Escalona, and Antonio Pulgarin Guerrero. Dentistry scientometric analysis: a comparative study between Brazil and other most productive countries in the area. Scientometrics, 95 (2):753–769, May 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0847-x.

Gore:2016:CMS

[GDP16] Ross Gore, Saikou Diallo, and Jose Padilla. Classifying modeling and simulation as a scientific discipline. Scientometrics, 109(2):615–628, November 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2050-y.

Geraci:2011:WDI

[GE11] Marco Geraci and M. Degli Esposti. Where do Italian universities stand? An in-depth statistical analysis of national and international rankings. Scientometrics, 87(3):667–681, June 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0350-9.

Georges:2017:WCM

[Geo17] Patrick Georges. Western classical music development: a statistical analysis of composers similarity, differentiation and evolution. Scientometrics, 112(1):21–53, July 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-017-2387-x.pdf.

Groh:2011:MMS

[GF11] Georg Groh and Christoph Fuchs. Multi-modal social networks for modeling scientific fields. Scientometrics, 89(2): 569–590, November 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0475-x.

Garousi:2017:QVI

[GF17] Vahid Garousi and João M. Fernandes. Quantity versus impact of software engineering papers: a quantitative study. Scientometrics, 112(2):963–1006, August 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2419-6.

Gadd:2018:IJP

[GFC18] Elizabeth Gadd, Jenny Fry, and Claire Creaser. The influence of journal publisher characteristics on open access policy trends. Scientometrics, 115(3):1371–1393, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2716-8;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2716-
8.pdf.

Groneberg:2018:SBO

[GFK^{+}18] David A. Groneberg, Axel Fischer, Doris Klingelhöfer, Michael H. K. Bendels, David Quarcoo, and Dörthe Brüggmann. The story behind Oncotarget? A bibliometric analysis. Scientometrics, 117(3):2195–2205, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2949-6.

Gao:2012:NMK

[GG12] Xia Gao and Jiancheng Guan. Network model of knowledge diffusion. Scientometrics, 90(3):749–762, March 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0554-z.

Gowanlock:2013:ARI

[GG13] Michael Gowanlock and Rich Gazan. Assessing researcher interdisciplinarity: a case study of the University of Hawaii NASA Astrobiology Institute. Scientometrics, 94(1):133– 161, January 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0765-y.

Gao:2014:NRW

[GG14] Wei Gao and Huai-Cheng Guo. Nitrogen research at watershed scale: a bibliometric analysis during 1959– 2011. Scientometrics, 99(3):737–753, June 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1240-8.

Glanzel:2015:UMV

[GG15a] Wolfgang Glänzel and Juan Gorraiz. Usage metrics versus altmetrics: confusing terminology? Scientometrics, 102(3): 2161–2164, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-014-1472-
7.pdf.

Gorraiz:2015:FBA

[GG15b] Juan Gorraiz and Christian Gumpenberger. A flexible bibliometric approach for the assessment of professorial appointments. Scientometrics, 105(3):1699–1719, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1703-6.

Gazni:2019:IDS

[GG19] Ali Gazni and Zahra Ghaseminik. The increasing dominance of science in the economy: Which nations are successful? Scientometrics, 120(3):1411–1426, September 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03161-5.

Gorraiz:2011:ERI

[GGG^{+}11] Juan Gorraiz, Christian Gumpenberger, Wolfgang Glänzel, Koenraad Debackere, Stefan Hornbostel, and Sybille Hinze. ESSS 2010: A review of the inaugurational European Summer School for Scientometrics in Berlin. Scientometrics, 86(1):235–236, January 2011. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-010-0279-4.

Gumpenberger:2012:ERE

[GGG^{+}12] Christian Gumpenberger, Juan Gorraiz, Wolfgang Glänzel, Koenrad Debackere, Stefan Hornbostel, and Sybille Hinze. Event report: esss 2011 —- scientometric education in Indian summer at the University of Vienna. Scientometrics, 91(1):311–313, April 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-011-0578-
4.pdf.

Gao:2014:APA

[GGG14] Xia Gao, Xi Guo, and Jiancheng Guan. An analysis of the patenting activities and collaboration among industry–university-research institutes in the Chinese ICT sector. Scientometrics, 98(1):247–263, January 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1048-y.

Gorraiz:2016:BCF

[GGG16a] Juan Gorraiz, Christian Gumpenberger, and Thomas Glade. On the bibliometric coordinates of four different research fields in geography. Scientometrics, 107(2):873–897, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1864-y;   http://link.
springer.com/content/pdf/10.1007/s11192-016-1864-
y.pdf.

Gumpenberger:2016:EAA

[GGG16b] Christian Gumpenberger, Wolfgang Glänzel, and Juan Gorraiz. The ecstasy and the agony of the Altmetric score. Scientometrics, 108(2):977–982, August 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-1991-5.pdf.

Gorraiz:2010:ESS

[GGH^{+}10] Juan Gorraiz, Christian Gumpenberger, Stefan Hornbostel, Sybille Hinze, Wolfgang Glänzel, and Koenraad Debackere. European Summer School for Scientometrics (ESSS) to be launched. Scientometrics, 83(2):601–602, May 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-010-0206-8.

Gorraiz:2014:ICI

[GGH^{+}14] Juan Gorraiz, Christian Gumpenberger, Marianne Hörlesberger, Henk Moed, and Edgar Schiebel. The 14th International Conference of the International Society for Scientometrics and Informetrics. Scientometrics, 101(2):937–938, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-014-1438-9.pdf.

Gorraiz:2014:PBR

[GGP14] Juan Gorraiz, Christian Gumpenberger, and Philip J. Purnell. The power of book reviews: a simple and transparent enhancement approach for book citation indexes. Scientometrics, 98(2):841–852, February 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1176-4.

Gao:2011:MCK

[GGR11] Xia Gao, Jiancheng Guan, and Ronald Rousseau. Mapping collaborative knowledge production in China using patent co-inventorships. Scientometrics, 88(2):343–362, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0404-z.

Gorraiz:2014:UVC

[GGS14] Juan Gorraiz, Christian Gumpenberger, and Christian Schlögl. Usage versus citation behaviours in four subject areas. Scientometrics, 101(2):1077–1095, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-014-1271-1.

Glaser:2017:SDD

[GGS17] Jochen Gläser, Wolfgang Glänzel, and Andrea Scharnhorst. Same data-different results? Towards a comparative approach to the identification of thematic structures in science. Scientometrics, 111(2):981–998, May 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-017-2296-z.pdf.

Gorraiz:2011:GRB

[GGW11] Juan Gorraiz, Christian Gumpenberger, and Martin Wieland. Galton 2011 revisited: a bibliometric journey in the footprints of a universal genius. Scientometrics, 88 (2):627–652, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0393-y.

Gumpenberger:2013:EBS

[GGW^{+}13] Christian Gumpenberger, Juan Gorraiz, Martin Wieland, Ivana Roche, Edgar Schiebel, Dominique Besagni, and Claire François. Exploring the bibliometric and semantic nature of negative results. Scientometrics, 95(1):277–297, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0829-z.

Gzoyan:2015:CAS

[GHA^{+}15] Edita G. Gzoyan, Lusine A. Hovhannisyan, Sofya A. Aleksanyan, Narine A. Ghazaryan, Simon R. Hunanyan, Ahmed Bourghida, and Shushanik A. Sargsyan. Comparative analysis of the scientific output of Armenia, Azerbaijan and Georgia. Scientometrics, 102(1):195–212, January 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1452-y.

Guevara:2016:RSU

[GHA^{+}16] Miguel R. Guevara, Dominik Hartmann, Manuel Aristarán, Marcelo Mendoza, and César A. Hidalgo. The research space: using career paths to predict the evolution of the research output of individuals, institutions, and nations. Scientometrics, 109(3):1695–1709, December 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2125-9.

Ghiasi:2018:ICP

[GHS18] Gita Ghiasi, Matthew Harsh, and Andrea Schiffauerova. Inequality and collaboration patterns in Canadian nanotechnology: implications for pro-poor and gender-inclusive policy. Scientometrics, 115(2):785–815, May 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2701-2.

Glanzel:2016:TMP

[GHT16] Wolfgang Glänzel, Sarah Heeffer, and Bart Thijs. A triangular model for publication and citation statistics of individual authors. Scientometrics, 107(2):857–872, May 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1870-0.

Glanzel:2017:LAS

[GHT17] Wolfgang Glänzel, Sarah Heeffer, and Bart Thijs. Lexical analysis of scientific publications for nano-level scientometrics. Scientometrics, 111(3):1897–1906, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Gurney:2012:ADU

[GHvdB12] Thomas Gurney, Edwin Horlings, and Peter van den Besselaar. Author disambiguation using multi-aspect similarity indicators. Scientometrics, 91(2):435–449, May 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-011-0589-1.pdf.

Grauwin:2011:MSI

[GJ11] Sebastian Grauwin and Pablo Jensen. Mapping scientific institutions. Scientometrics, 89(3):943–954, December 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0482-y.

Garg:2014:SPI

[GK14] K. C. Garg and S. Kumar. Scientometric profile of Indian scientific output in life sciences with a focus on the contributions of women scientists. Scientometrics, 98(3):1771–1783, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1107-4.

Gingras:2018:AEU

[GK18] Yves Gingras and Mahdi Khelfaoui. Assessing the effect of the United States’ “citation advantage” on other countries’ scientific impact as measured in the Web of Science (WoS) database. Scientometrics, 114(2):517–532, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2593-6.

Grega:2019:HAP

[GK19] Dominik Grega and Jozef Kolár. Historical analysis of pharmacoeconomic terms. Scientometrics, 119(3):1643–1654, June 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03093-0.

Groneberg:2019:NQQ

[GKB^{+}19] David A. Groneberg, Doris Klingelhöfer, Dörthe Brüggmann, Cristian Scutaru, Axel Fischer, and David Quarcoo. New quality and quantity indices in science (NewQIS): results of the first decade-project progress review. Scientometrics, 121 (1):451–478, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03188-8.

Gomide:2017:NUP

[GKF17] Janaína Gomide, Hugo Kling, and Daniel Figueiredo. Name usage pattern in the synonym ambiguity problem in bibliographic data. Scientometrics, 112(2):747–766, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2410-2.

Ghosh:2015:FIC

[GKK15] Jaideep Ghosh, Avinash Kshitij, and Sandeep Kadyan. Functional information characteristics of large-scale research collaboration: network measures and implications. Scientometrics, 102(2):1207–1239, February 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1475-4.

Guskov:2016:SRR

[GKS16] Andrey Guskov, Denis Kosyakov, and Irina Selivanova. Scientometric research in Russia: impact of science policy changes. Scientometrics, 107(1):287–303, April 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1876-7.

Guskov:2018:BRP

[GKS18] Andrey E. Guskov, Denis V. Kosyakov, and Irina V. Selivanova. Boosting research productivity in top Russian universities: the circumstances of breakthrough. Scientometrics, 117(2):1053–1080, November 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2890-8.

Gupta:2011:MIC

[GKV11] B. M. Gupta, Avinash Kshitij, and Charu Verma. Mapping of Indian computer science research output, 1999– 2008. Scientometrics, 86(2):261–283, February 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0272-y.

Ghiasi:2015:SSI

[GL15] Gita Ghiasi and Vincent Larivière. Sectoral systems of innovation: the case of robotics research activities. Scientometrics, 104(2):407–424, August 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1611-9.

Glanzel:2010:RIC

[Glä10] Wolfgang Glänzel. The role of the h-index and the characteristic scores and scales in testing the tail properties of scientometric distributions. Scientometrics, 83(3):697–709, June 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0124-9.

Glanzel:2012:RCD

[Glä12] Wolfgang Glänzel. The role of core documents in bibliometric network analysis and their relation with h-type indices. Scientometrics, 93(1):113–123, October 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0639-3.

Glanzel:2013:HEP

[Glä13] Wolfgang Glänzel. High-end performance or outlier? Evaluating the tail of scientometric distributions. Scientometrics, 97(1):13–23, October 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1022-8.

Glanzel:2014:GNE

[Glä14] Wolfgang Glänzel. Greetings from the new Editor-in-Chief. Scientometrics, 98(1):3–4, January 2014. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-013-1197-z.

Glanzel:2015:BAR

[Glä15] Wolfgang Glänzel. Bibliometrics-aided retrieval: where information retrieval meets scientometrics. Scientometrics, 102(3):2215–2222, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
014-1480-7.

Glanzel:2018:ECB

[Glä18] Wolfgang Glänzel. Expression of concern: Bibliometric study of Electronic Commerce Research in Information Systems & MIS Journals, Scientometrics, 2016, 109(3), 1455–1476 (https://doi.org/10.1007/s11192-016-2142-8). Scientometrics, 114(3):1423, March 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2575-8;   https://link.springer.
com/content/pdf/10.1007/s11192-017-2575-8.pdf. See

[LHC16].

Gazni:2016:ECI

[GLD16] Ali Gazni, Vincent Larivière, and Fereshteh Didegah. The effect of collaborators on institutions’ scientific impact. Scientometrics, 109(2):1209–1230, November 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2101-4.

Guns:2011:QMB

[GLM11] Raf Guns, Yu Xian Liu, and Dilruba Mahbuba. Q-measures and betweenness centrality in a collaboration network: a case study of the field of informetrics. Scientometrics, 87 (1):133–147, April 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0332-3.

Gasko:2016:NNM

[GLS16] Noémi Gaskó, Rodica Ioana Lung, and Mihai Alexandru Suciu. A new network model for the study of scientific collaborations: Romanian computer science and mathematics co-authorship networks. Scientometrics, 108(2):613–632, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1968-4.

Garcia-Lillo:2016:OAE

[GLÚGML16] Francisco García-Lillo, Mercedes Úbeda-García, and Bartolomé Marco-Lajara. Organizational ambidexterity: exploring the knowledge base. Scientometrics, 107(3):1021– 1040, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1897-2.

Gerken:2012:NIT

[GM12] Jan M. Gerken and Martin G. Moehrle. A new instrument for technology monitoring: novelty in patents measured by semantic patent analysis. Scientometrics, 91(3):645–670, June 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0635-7.

Glanzel:2013:OPT

[GM13] Wolfgang Glänzel and Henk F. Moed. Opinion paper: thoughts and facts on bibliometric indicators. Scientometrics, 96(1):381–394, July 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0898-z.

Gautret:2017:DSA

[GMJ^{+}17] Marjolaine Gautret, Stefano Messori, André Jestin, Marina Bagni, and Alain Boissy. Development of a semi-automatic bibliometric system for publications on animal health and welfare: a methodological study. Scientometrics, 113(2): 803–823, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2494-8.

Guccio:2016:EAC

[GMM16] Calogero Guccio, Marco Ferdinando Martorana, and Isidoro Mazza. Efficiency assessment and convergence in teaching and research in Italian public universities. Scientometrics, 107(3):1063–1094, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
016-1903-8.

Gorraiz:2010:P

[GMS10] Juan Gorraiz, Henk Moed, and Edgar Schiebel. Preface. Scientometrics, 82(3):459–460, March 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-010-0181-0.pdf.

Guo:2018:SEI

[GMSZ18] Feng Guo, Chao Ma, Qingling Shi, and Qingqing Zong. Succinct effect or informative effect: the relationship between title length and the number of citations. Scientometrics, 116(3):1531–1539, September 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2805-8.

Gherardini:2017:YGI

[GN17] Alberto Gherardini and Alberto Nucciotti. Yesterday’s giants and invisible colleges of today. A study on the ‘knowledge transfer’ scientific domain. Scientometrics, 112(1):255– 271, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Georges:2019:VMS

[GN19] Patrick Georges and Ngoc Nguyen. Visualizing music similarity: clustering and mapping 500 classical music composers. Scientometrics, 120(3):975–1003, September 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03166-0;   http://
link.springer.com/content/pdf/10.1007/s11192-019-
03166-0.pdf.

Gyorffy:2018:FIS

[GNHT18] Balázs Győrffy, Andrea Magda Nagy, Péter Herman, and Ádám Török. Factors influencing the scientific performance of Momentum grant holders: an evaluation of the first 117 research groups. Scientometrics, 117(1):409–426, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2852-1;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2852-
1.pdf.

Gul:2015:MER

[GNS^{+}15] Sumeer Gul, Nahida Tun Nisa, Tariq Ahmad Shah, Sangita Gupta, Asifa Jan, and Suhail Ahmad. Middle East: research productivity and performance across nations. Scientometrics, 105(2):1157–1166, November 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1722-3.

Gomez-Nunez:2011:ISJ

[GNVQdMAG11] Antonio J. Gómez-Núñez, Benjamín Vargas-Quesada, Félix de Moya-Anegón, and Wolfgang Glänzel. Improving SCImago journal & country rank (SJR) subject classification through reference analysis. Scientometrics, 89(3):

741–758, December 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0485-8.

Gomez:2019:AED

[Gom19] Juan Carlos Gomez. Analysis of the effect of data properties in automated patent classification. Scientometrics, 121(3): 1239–1268, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03246-1.

Gumpenberger:2013:IGO

[GOPG13] Christian Gumpenberger, María-Antonia Ovalle-Perandones, and Juan Gorraiz. On the impact of gold open access journals. Scientometrics, 96(1):221–238, July 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0902-7.

Garcia-Perez:2013:LVE

[GP13] Miguel A. García-Pérez. Limited validity of equations to predict the future h index. Scientometrics, 96(3):901– 909, September 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-0979-7.

Garcia-Perez:2015:OSI

[GP15] Miguel A. García-Pérez. Online supplemental information: a sizeable black hole for citations. Scientometrics, 102(2): 1655–1659, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1348-x.

Gong:2018:EPP

[GP18a] Hong Gong and Shan Peng. Effects of patent policy on innovation outputs and commercialization: evidence from universities in China. Scientometrics, 117(2):687–703, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2893-5.

Guan:2018:BRB

[GP18b] Jiancheng Guan and Lanxin Pang. Bidirectional relationship between network position and knowledge creation in Scientometrics. Scientometrics, 115(1):201–222, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2673-2.

Giuliani:2010:ASC

[GPN10] Francesco Giuliani, Michele Pio De Petris, and Giovanni Nico. Assessing scientific collaboration through coauthorship and content sharing. Scientometrics, 85(1):13–28, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0264-y.

Garner:2014:DVM

[GPN14] Jon Garner, Alan L. Porter, and Nils C. Newman. Distance and velocity measures: using citations to determine breadth and speed of research impact. Scientometrics, 100(3):687– 703, September 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1316-5.

Grimaldo:2018:RPR

[GPSM18] Francisco Grimaldo, Mario Paolucci, and Jordi Sabater-Mir. Reputation or peer review? The role of outliers. Scientometrics, 116(3):1421–1438, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2826-3;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2826-
3.pdf.

Ghani:2019:CEI

[GQAM19] Rizwan Ghani, Faiza Qayyum, Muhammad Tanvir Afzal, and Hermann Maurer. Comprehensive evaluation of hindex and its extensions in the domain of mathematics. Scientometrics, 118(3):809–822, March 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03007-0.

Guns:2014:RRC

[GR14] Raf Guns and Ronald Rousseau. Recommending research collaborations using link prediction and random forest classifiers. Scientometrics, 101(2):1461–1473, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-1228-9.

Gorry:2016:SBH

[GR16] Philippe Gorry and Pascal Ragouet. ”Sleeping Beauty” and her restless sleep: Charles Dotter and the birth of interventional radiology. Scientometrics, 107(2):773–784, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1859-8.

Garechana:2017:EIM

[GRBBS17] Gaizka Garechana, Rosa Río-Belver, Iñaki Bildosola, and Marisela Rodríguez Salvador. Effects of innovation management system standardization on firms: evidence from text mining annual reports. Scientometrics, 111(3):1987–1999, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Garcia-Romero:2014:BAP

[GREL14] Antonio García-Romero and José Manuel Estrada-Lorenzo. A bibliometric analysis of plagiarism and self-plagiarism through Déjà vu. Scientometrics, 101(1):381–396, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1387-3.

Gorraiz:2012:KFC

[GRG12] Juan Gorraiz, Ralph Reimann, and Christian Gumpenberger. Key factors and considerations in the assessment of international collaboration: a case study for Austria and six countries. Scientometrics, 91(2):417–433, May 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0579-3.

Garcia:2011:OPJ

[GRSFV11] J. A. García, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. Overall prestige of journals with ranking score above a given threshold. Scientometrics, 89(1):229–243, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0442-6.

Garcia:2012:CTE

[GRSFV12a] J. A. García, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. A comparison of top economics departments in the US and EU on the basis of the multidimensional prestige of influential articles in 2010. Scientometrics, 93(3): 681–698, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0708-7.

Garcia:2012:RRO

[GRSFV^{+}12b] J. A. García, Rosa Rodriguez-Sánchez, J. Fdez-Valdivia, Daniel Torres-Salinas, and Francisco Herrera. Ranking of research output of universities on the basis of the multidimensional prestige of influential fields: Spanish universities as a case of study. Scientometrics, 93(3):1081–1099, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0740-7.

Garcia:2013:BRP

[GRSFV^{+}13] J. A. García, Rosa Rodriguez-Sánchez, J. Fdez-Valdivia, Nicolas Robinson-García, and Daniel Torres-Salinas. Benchmarking research performance at the university level with information theoretic measures. Scientometrics, 95(1): 435–452, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0854-y.

Garcia:2014:HSO

[GRSFV14a] J. A. García, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. How the same organizational structures can arise across seemingly unrelated domains of human activities: the example of academic publishing and stock market. Scientometrics, 99(2):447–461, May 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1184-4.

Garcia:2014:SHQ

[GRSFV14b] J. A. García, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. The selection of high-quality manuscripts. Scientometrics, 98(1):299–313, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
013-1034-4.

Garcia:2014:BCS

[GRSFV^{+}14c] J. A. García, Rosa Rodriguez-Sánchez, J. Fdez-Valdivia, Nicolas Robinson-García, and Daniel Torres-Salinas. Bestin-class and strategic benchmarking of scientific subject categories of Web of Science in 2010. Scientometrics, 99(3):615– 630, June 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1000-1.

Garcia:2015:AEG

[GRSFV15] J. A. García, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. The author-editor game. Scientometrics, 104 (1):361–380, July 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1566-x.

Garcia:2016:ARW

[GRSFV16a] J. A. García, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. Authors and reviewers who suffer from confirmatory bias. Scientometrics, 109(2):1377–1395, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2079-y.

Garcia:2016:WRR

[GRSFV16b] J. A. García, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. Why the referees’ reports I receive as an editor are so much better than the reports I receive as an author? Scientometrics, 106(3):967–986, March 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1827-8.

Garcia:2017:POP

[GRSFV17a] J. A. García, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. Problems with open participation in peer review. Scientometrics, 112(3):1881–1885, September 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2445-4.

Garcia:2017:STF

[GRSFV17b] J. A. García, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. STRATEGY: a tool for the formulation of peer-review strategies. Scientometrics, 113(1):45–60, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2470-3.

Garcia:2018:CBA

[GRSFV18] J. A. García, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. Competition between academic journals for scholars’ attention: the ‘Nature effect’ in scholarly communication. Scientometrics, 115(3):1413–1432, June 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2723-9.

Garcia:2019:DBP

[GRSFV19a] J. A. García, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. Do the best papers have the highest probability of being cited? Scientometrics, 118(3):885–890, March 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03008-z.

Garcia:2019:OAI

[GRSFV19b] J. A. Garcia, Rosa Rodriguez-Sánchez, and J. Fdez-Valdivia. The optimal amount of information to provide in an academic manuscript. Scientometrics, 121(3):1685– 1705, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03270-1.

Garcia:2019:AIP

[GRSFVCP19] J. A. García, Rosa Rodriguez-Sánchez, J. Fdez-Valdivia, and Jorge Chamorro-Padial. The author’s ignorance on the publication fees is a source of power for publishers. Scientometrics, 121(3):1435–1445, December 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03231-8.

Garcia:2014:WAA

[GRSFVdMA14] J. A. García, Rosa Rodriguez-Sánchez, J. Fdez-Valdivia, and F. de Moya-Anegón. A web application for aggregating conflicting reviewers’ preferences. Scientometrics, 99 (2):523–539, May 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1198-y.

Garcia:2012:FQJ

[GRSFVMB12] J. A. García, Rosa Rodriguez-Sánchez, J. Fdez-Valdivia, and J. Martinez-Baena. On first quartile journals which are not of highest impact. Scientometrics, 90(3):925–943, March 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0534-3.

Garcia-Romero:2016:ABW

[GRSS16] Antonio García-Romero, Daniel Santín, and Gabriela Sicilia. Another brick in the wall: a new ranking of academic journals in economics using FDH. Scientometrics, 107(1):91–101, April 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1843-3.

Galban-Rodriguez:2019:MCS

[GRTPMLAJ19] Ernesto Galbán-Rodríguez, Déborah Torres-Ponjuán, Yohannis Martí-Lahera, and Ricardo Arencibia-Jorge. Measuring the Cuban scientific output in scholarly journals through a comprehensive coverage approach. Scientometrics, 121(2): 1019–1043, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03233-6.

Guan:2012:TCT

[GS12] Jiancheng Guan and Yuan Shi. Transnational citation, technological diversity and small world in global nanotechnology patenting. Scientometrics, 93(3):609–633, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0706-9.

Glanzel:2015:FCS

[GS15] Wolfgang Glänzel and András Schubert. Foreword to the “Case Studies in Scientometrics” special issues. Scientometrics, 105(1):1–3, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
015-1755-7;   http://link.springer.com/content/pdf/
10.1007/s11192-015-1755-7.pdf.

Glanzel:2018:EPE

[GSB18] W. Glänzel, A. Schubert, and T. Braun. Editorial preface to the Eugene Garfield Memorial Issue. Scientometrics, 114(2):371–372, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2643-8;   https://link.springer.com/content/
pdf/10.1007/s11192-018-2643-8.pdf.

Guns:2018:CCO

[GSE^{+}18] Raf Guns, Linda Sile, Joshua Eykens, Frederik T. Verleysen, and Tim C. E. Engels. A comparison of cognitive and organizational classification of publications in the social sciences and humanities. Scientometrics, 116(2):1093–1111, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2775-x.

Gogoglou:2017:FDC

[GSKM17] Antonia Gogoglou, Antonis Sidiropoulos, Dimitrios Katsaros, and Yannis Manolopoulos. The fractal dimension of a citation curve: quantifying an individual’s scientific output using the geometry of the entire curve. Scientometrics, 111(3):1751–1774, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Ginde:2016:SFM

[GSM^{+}16] Gouri Ginde, Snehanshu Saha, Archana Mathur, Sukrit Venkatagiri, Sujith Vadakkepat, Anand Narasimhamurthy, and B. S. Daya Sagar. ScientoBASE: a framework and model for computing scholastic indicators of non-local influence of journals via native data acquisition algorithms. Scientometrics, 108(3):1479–1529, September 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2006-2.

Gomez-Sancho:2010:NAM

[GSMT10] José María Gómez-Sancho and María Jesús Mancebón-Torrubia. A new approach to measuring scientific production in JCR journals and its application to Spanish public universities. Scientometrics, 85(1):271–293, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0217-5.

Gonzalez-Sala:2019:JAS

[GSOLHO19] Francisco González-Sala, Julia Osca-Lluch, and Julia Haba-Osca. Are journal and author self-citations a visibility strategy? Scientometrics, 119(3):1345–1364, June 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03101-3.

Guzman-Sanchez:2018:CCB

[GSPLVG^{+}18] Maria Victoria Guzmán-Sánchez, Maybel Piñón-Lora, Elio Atenógenes Villaseñor-García, José Luis Jiménez-Andrade, and Humberto Carrillo-Calvet. Characterization of the Cuban biopharmaceutical industry from collaborative networks. Scientometrics, 115(3):1533–1548, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2719-5.

Glanzel:2011:PVP

[GSTD11] Wolfgang Glänzel, András Schubert, Bart Thijs, and Koenraad Debackere. A priori vs. a posteriori normalisation of citation indicators. The case of journal ranking. Scientometrics, 87(2):415–424, May 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0345-6.

Glanzel:2011:UCD

[GT11] Wolfgang Glänzel and Bart Thijs. Using ‘core documents’ for the representation of clusters and topics. Scientometrics, 88(1):297–309, July 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0347-4.

Glanzel:2012:UCD

[GT12] Wolfgang Glänzel and Bart Thijs. Using ‘core documents’ for detecting and labelling new emerging topics. Scientometrics, 91(2):399–416, May 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0591-7.

Glanzel:2017:UHM

[GT17] Wolfgang Glänzel and Bart Thijs. Using hybrid methods and ‘core documents’ for the representation of clusters and topics: the astronomy dataset. Scientometrics, 111(2): 1071–1087, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Glanzel:2018:RBG

[GT18] Wolfgang Glänzel and Bart Thijs. The role of baseline granularity for benchmarking citation impact. the case of CSS profiles. Scientometrics, 116(1):521–536, July 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2747-1.

Gonzalez-Teruel:2018:IEC

[GTAG18] Aurora González-Teruel and Francisca Abad-García. The influence of Elfreda Chatman’s theories: a citation context analysis. Scientometrics, 117(3):1793–1819, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2915-3.

Glanzel:2016:CEB

[GTC16] Wolfgang Glänzel, Bart Thijs, and Pei-Shan Chi. The challenges to expand bibliometric studies from periodical literature to monographic literature with a new data source: the book citation index. Scientometrics, 109(3):2165–2179, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2046-7.

Glanzel:2014:ACB

[GTD14] Wolfgang Glänzel, Bart Thijs, and Koenraad Debackere. The application of citation-based performance classes to the disciplinary and multidisciplinary assessment in national comparison and institutional research assessment. Scientometrics, 101(2):939–952, November 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1247-1.

Gonzalez-Teruel:2015:MRI

[GTGABAG15] Aurora González-Teruel, Gregorio González-Alcaide, Maite Barrios, and María-Francisca Abad-García. Mapping recent information behavior research: an analysis of co-authorship and co-citation networks. Scientometrics, 103(2):687–705, May 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1548-z.

Gimenez-Toledo:2016:TSB

[GTMRE^{+}16] Elea Giménez-Toledo, Jorge Mañana-Rodríguez, Tim C. E. Engels, Peter Ingwersen, Janne Pölönen, Gunnar Sivertsen, Frederik T. Verleysen, and Alesia A. Zuccala. Taking scholarly books into account: current developments in five European countries. Scientometrics, 107(2):685–699, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1886-5.

Gimenez-Toledo:2019:TSB

[GTMRE^{+}19] Elea Giménez-Toledo, Jorge Mañana-Rodríguez, Tim C. E. Engels, Raf Guns, Emanuel Kulczycki, Michael Ochsner, Janne Pölönen, Gunnar Sivertsen, and Alesia A. Zuccala.

Taking scholarly books into account, part II: a comparison of 19 European countries in evaluation and funding. Scientometrics, 118(1):233–251, January 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2956-7.

Gusenbauer:2019:GSO

[Gus19] Michael Gusenbauer. Google Scholar to overshadow them all? Comparing the sizes of 12 academic search engines and bibliographic databases. Scientometrics, 118(1):177–214, January 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2958-5;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2958-
5.pdf.

Gomez-Verjan:2015:TCP

[GVGSEPRC15] J. Gómez-Verjan, I. Gonzalez-Sanchez, E. Estrella-Parra, and R. Reyes-Chilpa. Trends in the chemical and pharmacological research on the tropical trees Calophyllum brasiliense and Calophyllum inophyllum, a global context. Scientometrics, 105(2):1019–1030, November 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1715-2.

Grancay:2017:PPH

[GVS17] Martin Grancay, Jolita Vveinhardt, and Erika Sumilo. Publish or perish: how Central and Eastern European economists have dealt with the ever-increasing academic publishing requirements 2000-2015. Scientometrics, 111(3): 1813–1837, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Gingras:2010:WIB

[GW10a] Yves Gingras and Matthew L. Wallace. Why it has become more difficult to predict Nobel Prize winners: a bibliometric analysis of nominees and winners of the chemistry and physics prizes (1901–2007). Scientometrics, 82(2):401–412, February 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0035-9.

Guan:2010:CSR

[GW10b] Jiancheng Guan and Gangbo Wang. A comparative study of research performance in nanotechnology for China’s inventor-authors and their non-inventing peers. Scientometrics, 84(2):331–343, August 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
009-0140-9.

Gan:2015:RCS

[GW15a] Chunmei Gan and Weijun Wang. Research characteristics and status on social media in China: A bibliometric and coword analysis. Scientometrics, 105(2):1167–1182, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1723-2.

Guan:2015:BCR

[GW15b] Jiancheng Guan and He Wei. A bilateral comparison of research performance at an institutional level. Scientometrics, 104(1):147–173, July 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1599-1.

Glanzel:2016:P

[GW16] Wolfgang Glänzel and Hans Willems. Preface. Scientometrics, 106(2):821–823, February 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
015-1813-1;   http://link.springer.com/content/pdf/
10.1007/s11192-015-1813-1.pdf.

Gnewuch:2017:TCC

[GW17] Matthias Gnewuch and Klaus Wohlrabe. Title characteristics and citations in economics. Scientometrics, 110(3):1573–1578, March 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2216-7.

Graziotin:2014:FSA

[GWA14] Daniel Graziotin, Xiaofeng Wang, and Pekka Abrahamsson. A framework for systematic analysis of open access journals and its application in software engineering and information systems. Scientometrics, 101(3):1627–1656, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1278-7.

Guo:2011:MIM

[GWB11] Hanning Guo, Scott Weingart, and Katy Börner. Mixedindicators model for identifying emerging research areas. Scientometrics, 89(1):421–435, October 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0433-7.

Guardiola-Wanden-Berghe:2013:MSH

[GWBSVWB13] Rocío Guardiola-Wanden-Berghe, Javier Sanz-Valero, and Carmina Wanden-Berghe. Medical subject headings versus American psychological association index terms: indexing eating disorders. Scientometrics, 94(1):305–311, January 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0866-7.

Gunashekar:2017:HDN

[GWG17] Salil Gunashekar, Steven Wooding, and Susan Guthrie. How do NIHR peer review panels use bibliometric information to support their decisions? Scientometrics, 112(3):1813– 1835, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2417-8;
http://link.springer.com/content/pdf/10.1007/s11192-
017-2417-8.pdf.

Guler:2016:SWB

[GWP16] Arzu Tugce Guler, Cathelijn J. F. Waaijer, and Magnus Palmblad. Scientific workflows for bibliometrics. Scientometrics, 107(2):385–398, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
016-1885-6;   http://link.springer.com/content/pdf/
10.1007/s11192-016-1885-6.pdf.

Gok:2015:UWM

[GWS15] Abdullah Gök, Alec Waterworth, and Philip Shapira. Use of web mining in studying innovation. Scientometrics, 102 (1):653–671, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-014-1434-
0.pdf.

Gong:2019:CAF

[GXC^{+}19] Kaile Gong, Juan Xie, Ying Cheng, Vincent Larivière, and Cassidy R. Sugimoto. The citation advantage of foreign language references for Chinese social science papers. Scientometrics, 120(3):1439–1460, September 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03146-4.

Gautam:2012:RCD

[GY12] Pitambar Gautam and Ryuichi Yanagiya. Reflection of cross-disciplinary research at Creative Research Institution (Hokkaido University) in the Web of Science database: appraisal and visualization using bibliometry. Scientometrics, 93(1):101–111, October 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0655-3.

Guan:2015:HDC

[GYZ15] Jiancheng Guan, Yan Yan, and Jingjing Zhang. How do collaborative features affect scientific output? Evidences from wind power field. Scientometrics, 102(1):333–355, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1311-x.

Glanzel:2011:PAC

[GZ11] Wolfgang Glänzel and Ping Zhou. Publication activity, citation impact and bi-directional links between publications and patents in biotechnology. Scientometrics, 86(2): 505–525, February 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0269-6.

Guan:2014:HKD

[GZ14a] Jiancheng Guan and Wenjia Zhu. How knowledge diffuses across countries: a case study in the field of management. Scientometrics, 98(3):2129–2144, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1134-1.

Guan:2014:CCC

[GZ14b] Jiancheng Guan and Kairui Zuo. A cross-country comparison of innovation efficiency. Scientometrics, 100(2):541–575, August 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1288-5.

Guo:2017:ACC

[GZ17] Shesen Guo and Ganzhou Zhang. Analyzing concept complexity, knowledge ageing and diffusion pattern of MOOC. Scientometrics, 112(1):413–430, July 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Glanzel:2018:SRA

[GZ18] Wolfgang Glänzel and Lin Zhang. Scientometric research assessment in the developing world: A tribute to Michael J. Moravcsik from the perspective of the twenty-first century. Scientometrics, 115(3):1517–1532, June 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2647-4.

Ghasemian:2016:TBS

[GZGAC16] Fahimeh Ghasemian, Kamran Zamanifar, Nasser Ghasem-Aqaee, and Noshir Contractor. Toward a better scientific collaboration success prediction model through the feature space expansion. Scientometrics, 108(2):777–801, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-016-1999-x. See erratum

[GZGAC17].

Ghasemian:2017:ETB

[GZGAC17] Fahimeh Ghasemian, Kamran Zamanifar, Nasser Ghasem-Aghaee, and Noshir Contractor. Erratum to: Toward a better scientific collaboration success prediction model through the feature space expansion. Scientometrics, 113 (3):1819, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-2170-4;
http://link.springer.com/content/pdf/10.1007/s11192-
016-2170-4.pdf. See [GZGAC16].

Guo:2015:ECD

[GZJ^{+}15] Shesen Guo, Ganzhou Zhang, Qiuhong Ju, Yu Chen, Qianfeng Chen, and Lulu Li. The evolution of conceptual diversity in economics titles from 1890 to 2012. Scientometrics, 102(3):2073–2088, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1501-6.

Godoy:2015:PPA

[GZM15] Daniela Godoy, Alejandro Zunino, and Cristian Mateos. Publication practices in the Argentinian Computer Science community: a bibliometric perspective. Scientometrics, 102(2):1795–1814, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1450-0.

Harzing:2016:GSS

[HA16] Anne-Wil Harzing and Satu Alakangas. Google Scholar, Scopus and the Web of Science: a longitudinal and crossdisciplinary comparison. Scientometrics, 106(2):787–804, February 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1798-9.

Harzing:2017:MAO

[HA17a] Anne-Wil Harzing and Satu Alakangas. Microsoft Academic is one year old: the Phoenix is ready to leave the nest. Scientometrics, 112(3):1887–1894, September 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2454-3.

Harzing:2017:MAP

[HA17b] Anne-Wil Harzing and Satu Alakangas. Microsoft Academic: is the phoenix getting wings? Scientometrics, 110(1):371–383, January 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2185-x.

Habib:2019:SBB

[HA19] Raja Habib and Muhammad Tanvir Afzal. Sectionsbased bibliographic coupling for research paper recommendation. Scientometrics, 119(2):643–656, May 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03053-8.

Harzing:2014:HIA

[HAA14] Anne-Wil Harzing, Satu Alakangas, and David Adams. hIa: an individual annual h-index to accommodate disciplinary and career length differences. Scientometrics, 99(3):811– 821, June 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1208-0.

Hagen:2010:DDD

[Hag10a] Nils T. Hagen. Deconstructing doctoral dissertations: how many papers does it take to make a PhD? Scientometrics, 85(2):567–579, November 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-010-0214-8.pdf.

Hagen:2010:HPC

[Hag10b] Nils T. Hagen. Harmonic publication and citation counting: sharing authorship credit equitably — not equally, geometrically or arithmetically. Scientometrics, 84(3):785– 793, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-009-0129-
4.pdf.

Hashem:2016:MRO

[HAG^{+}16] Ibrahim Abaker Targio Hashem, Nor Badrul Anuar, Abdullah Gani, Ibrar Yaqoob, Feng Xia, and Samee Ullah Khan. MapReduce: Review and open challenges. Scientometrics, 109(1):389–422, October 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-1945-y.

Hildrun:2012:REP

[HAJ12] Kretschmer Hildrun, Pudovkin Alexander, and Stegmann Johannes. Research evaluation. Part II: gender effects of evaluation: are men more productive and more cited than women? Scientometrics, 93(1):17–30, October 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0658-0. See erratum [KPS12].

Halilem:2011:AIT

[HAL11] Norrin Halilem, Nabil Amara, and Réjean Landry. Is the academic ivory tower becoming a managed structure? A nested analysis of the variance in activities of researchers from natural sciences and engineering in Canada. Scientometrics, 86(2):431–448, February 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0278-5.

Hallonsten:2013:IFF

[Hal13] Olof Hallonsten. Introducing ‘facilitymetrics’: a first review and analysis of commonly used measures of scientific leadership among synchrotron radiation facilities worldwide. Scientometrics, 96(2):497–513, August 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0945-9.

Hallonsten:2014:HEB

[Hal14] Olof Hallonsten. How expensive is big science? Consequences of using simple publication counts in performance assessment of large scientific facilities. Scientometrics, 100 (2):483–496, August 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1249-z.

Hammarfelt:2011:IIB

[Ham11] Björn Hammarfelt. Interdisciplinarity and the intellectual base of literature studies: citation analysis of highly cited monographs. Scientometrics, 86(3):705–725, March 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0314-5.

Hammarfelt:2014:UAA

[Ham14] Björn Hammarfelt. Using altmetrics for assessing research impact in the humanities. Scientometrics, 101(2):1419– 1430, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1261-3.

Han:2011:DCU

[Han11] Chung-Souk Han. On the demographical changes of U.S. research doctorate awardees and corresponding trends in research fields. Scientometrics, 89(3):845–865, December 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0487-6.

Hanazaki:2015:WWA

[Han15] Natalia Hanazaki. Why are we so attached to the “ethno” prefix in Brazil? Scientometrics, 103(2):545–554, May 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1540-7.

Hartley:2012:CCA

[Har12] James Hartley. To cite or not to cite: author self-citations and the impact factor. Scientometrics, 92(2):313–317, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0568-6.

Hardeman:2013:OLR

[Har13a] Sjoerd Hardeman. Organization level research in scientometrics: a plea for an explicit pragmatic approach.

Scientometrics, 94(3):1175–1194, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-012-0806-6.pdf.

Hartley:2013:NWS

[Har13b] James Hartley. New ways of sending Christmas greetings. Scientometrics, 97(3):911–912, December 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-013-1064-y.

Harzing:2013:DCI

[Har13c] Anne-Wil Harzing. Document categories in the ISI Web of Knowledge: Misunderstanding the social sciences? Scientometrics, 94(1):23–34, January 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0738-1.

Harzing:2013:PTG

[Har13d] Anne-Wil Harzing. A preliminary test of Google Scholar as a source for citation data: a longitudinal study of Nobel Prize winners. Scientometrics, 94(3):1057–1075, March 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0777-7.

Hartley:2014:NTC

[Har14a] James Hartley. New technology and Christmas greetings. Scientometrics, 101(3):2081–2082, December 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-014-1341-4.

Harzing:2014:LSG

[Har14b] Anne-Wil Harzing. A longitudinal study of Google Scholar coverage between 2012 and 2013. Scientometrics, 98(1): 565–575, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-0975-y.

Hartley:2015:IT

[Har15a] James Hartley. Inaccuracies in titles. Scientometrics, 103 (1):329–330, April 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-015-
1530-9.

Harzing:2015:HWM

[Har15b] Anne-Wil Harzing. Health warning: might contain multiple personalities-the problem of homonyms in Thomson Reuters Essential Science Indicators. Scientometrics, 105(3):2259– 2270, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1699-y.

Hartley:2016:ETF

[Har16a] James Hartley. Erratum to: Is time up for the Flesch measure of reading ease? Scientometrics, 109(3):2329, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11192-016-2123-y.pdf. See

[Har16c].

Hartley:2016:ITP

[Har16b] James Hartley. Is it true that papers written by jointauthors are cited more than papers written by single ones? What else matters? Scientometrics, 106(2):817–818, February 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/accesspage/article/10.1007/s11192-015-1799-8;
 http://link.springer.com/article/10.1007/s11192-
015-1799-8.

Hartley:2016:TFM

[Har16c] James Hartley. Is time up for the Flesch measure of reading ease? Scientometrics, 107(3):1523–1526, June 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1920-7. See erratum [Har16a].

Harzing:2016:MAS

[Har16d] Anne-Wil Harzing. Microsoft Academic (Search): a Phoenix arisen from the ashes? Scientometrics, 108(3):1637– 1647, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-016-
2026-y.

Hartley:2017:ATC

[Har17] James Hartley. Authors and their citations: a point of view. Scientometrics, 110(2):1081–1084, February 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2211-z.

Hartley:2019:SRB

[Har19a] James Hartley. Some reflections on being cited 10,000 times. Scientometrics, 118(1):375–381, January 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2966-5.

Harzing:2019:TNK

[Har19b] Anne-Wil Harzing. Two new kids on the block: How do Crossref and Dimensions compare with Google Scholar, Microsoft Academic, Scopus and the Web of Science? Scientometrics, 120(1):341–349, July 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03114-y.

Haustein:2016:GCA

[Hau16] Stefanie Haustein. Grand challenges in altmetrics: heterogeneity, data quality and dependencies. Scientometrics, 108 (1):413–423, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1910-9.

Haunschild:2015:DAN

[HB15] Robin Haunschild and Lutz Bornmann. Discussion about the new Nature Index. Scientometrics, 102(2):1829–1830, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-014-1505-2. See

response [CG15a].

Haunschild:2017:HMS

[HB17a] Robin Haunschild and Lutz Bornmann. How many scientific papers are mentioned in policy-related documents? an empirical investigation using Web of Science and Altmetric data. Scientometrics, 110(3):1209–1216, March 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-2237-2.pdf.

Hug:2017:CMA

[HB17b] Sven E. Hug and Martin P. Brändle. The coverage of Microsoft Academic: analyzing the publication output of a university. Scientometrics, 113(3):1551–1571, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2535-3.

Haunschild:2018:FTN

[HB18a] Robin Haunschild and Lutz Bornmann. Field- and timenormalization of data with many zeros: an empirical analysis using citation and Twitter data. Scientometrics, 116 (2):997–1012, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2771-1;
http://link.springer.com/content/pdf/10.1007/s11192-
018-2771-1.pdf.

Heinisch:2018:NGP

[HB18b] Dominik P. Heinisch and Guido Buenstorf. The next generation (plus one): an analysis of doctoral students’ academic fecundity based on a novel approach to advisor identification. Scientometrics, 117(1):351–380, October 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2840-5.

Haunschild:2019:RPP

[HBA19] Robin Haunschild, Lutz Bornmann, and Jonathan Adams. R package for producing beamplots as a preferred alternative to the h index when assessing single researchers (based on downloads from Web of Science). Scientometrics, 120 (2):925–927, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03147-3.

Huang:2018:NVS

[HBDL18] Yong Huang, Yi Bu, Ying Ding, and Wei Lu. Number versus structure: towards citing cascades. Scientometrics, 117(3): 2177–2193, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2952-y.

Hassan:2019:ITR

[HBS^{+}19] Saeed-Ul Hassan, Timothy D. Bowman, Mudassir Shabbir, Aqsa Akhtar, Mubashir Imran, and Naif Radi Aljohani. Influential tweeters in relation to highly cited articles in altmetric big data. Scientometrics, 119(1):481–493, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03044-9.

Hoyla:2016:CCS

[HBT16] Tuomas Höylä, Christoph Bartneck, and Timo Tiihonen. The consequences of competition: simulating the effects of research grant allocation strategies. Scientometrics, 108 (1):263–288, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1940-3.

Huang:2012:CSI

[HC12] Mu-Hsuan Huang and Yu-Wei Chang. A comparative study of interdisciplinary changes between information science and library science. Scientometrics, 91(3):789–803, June 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0619-7.

Hartley:2014:DMW

[HC14a] James Hartley and Guillaume Cabanac. Do men and women differ in their use of tables and graphs in academic publications? Scientometrics, 98(2):1161–1172, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-1096-3.

Huang:2014:DRF

[HC14b] Mu-Hsuan Huang and Chia-Pin Chang. Detecting research fronts in OLED field using bibliographic coupling with sliding window. Scientometrics, 98(3):1721–1744, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-1126-1.

Hartley:2015:AOW

[HC15a] James Hartley and Guillaume Cabanac. An academic odyssey: writing over time. Scientometrics, 103(3):1073– 1082, June 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1562-1.

Hsu:2015:BSS

[HC15b] Chien-Lung Hsu and Chun-Hao Chiang. A bibliometric study of SSME in information systems research. Scientometrics, 102(3):1835–1865, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1334-3.

Hsu:2015:FCR

[HC15c] Chien-Lung Hsu and Chun-Hao Chiang. The financial crisis research: a bibliometric analysis. Scientometrics, 105(1):161–177, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-

015-1698-z. See retraction notice [HC17].

Hartley:2016:TAB

[HC16a] James Hartley and Guillaume Cabanac. Are two authors better than one? Can writing in pairs affect the readability of academic blogs? Scientometrics, 109(3):2119–2122, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2116-x.

Huai:2016:BAP

[HC16b] Cuiqian Huai and Lihe Chai. A bibliometric analysis on the performance and underlying dynamic patterns of water security research. Scientometrics, 108(3):1531–1551, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2019-x.

Huang:2016:CST

[HC16c] Mu-Hsuan Huang and Chia-Pin Chang. A comparative study on three citation windows for detecting research fronts. Scientometrics, 109(3):1835–1853, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2133-9.

Hsu:2017:RNF

[HC17] Chien-Lung Hsu and Chun-Hao Chiang. Retraction note to: The financial crisis research: a bibliometric analysis. Scientometrics, 113(3):1827–1828, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2487-7;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2487-

7.pdf. See [HC15c].

Huang:2019:EIS

[HC19] Jia-Yen Huang and Rong-Chang Chen. Exploring the intellectual structure of cloud patents using non-exhaustive overlaps. Scientometrics, 121(2):739–769, November 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03219-4.

Chen:2015:EAM

[hCcTmWH15] Kuang hua Chen, Muh chyun Tang, Chun mei Wang, and Jieh Hsiang. Exploring alternative metrics of scholarly performance in the social sciences and humanities in Taiwan. Scientometrics, 102(1):97–112, January 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1420-6.

Hartley:2016:BRT

[HCDT16] James Hartley, John Cowan, Cynthia Deeson, and Peter Thomas. Book reviews in time. Scientometrics, 109(3):2123–2128, December 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2114-z.

Hu:2014:HCP

[HCL14] Zhigang Hu, Chaomei Chen, and Zeyuan Liu. How are collaboration and productivity correlated at various career stages of scientists? Scientometrics, 101(2):1553– 1564, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1323-6.

Huang:2014:ETR

[HCLC14] Mu-Hsuan Huang, Ssu-Han Chen, Chia-Ying Lin, and Dar-Zen Chen. Exploring temporal relationships between scientific and technical fronts: a case of biotechnology field. Scientometrics, 98(2):1085–1100, February 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1054-0.

Huang:2015:MTP

[HCS^{+}15] Mu-Hsuan Huang, Dar-Zen Chen, Danqi Shen, Mona S. Wang, and Fred Y. Ye. Measuring technological performance of assignees using trace metrics in three fields. Scientometrics, 104(1):61–86, July 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
015-1604-8.

Chen:2012:CSW

[hCyL12] Kuang hua Chen and Pei yu Liao. A comparative study on world university rankings: a bibliometric survey. Scientometrics, 92(1):89–103, July 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0724-7.

Hayat:2017:ADH

[HD17] Malik Khizar Hayat and Ali Daud. Anomaly detection in heterogeneous bibliographic information networks using coevolution pattern mining. Scientometrics, 113(1):149–175, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2467-y.

Huang:2013:UPR

[HDC13] Mu-Hsuan Huang, Huei-Ru Dong, and Dar-Zen Chen. The unbalanced performance and regional differences in scientific and technological collaboration in the field of solar cells. Scientometrics, 94(1):423–438, January 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0755-0.

Hoppen:2016:NBB

[HdSV16] Natascha Helena Franz Hoppen and Samile Andréa de Souza Vanz. Neurosciences in Brazil: a bibliometric study of main characteristics, collaboration and citations. Scientometrics, 109(1):121–141, October 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-1919-0.

Hung:2015:ECU

[HDW^{+}15] Wen-Chi Hung, Cherng G. Ding, Hung-Jui Wang, Meng-Che Lee, and Chieh-Peng Lin. Evaluating and comparing the university performance in knowledge utilization for patented inventions. Scientometrics, 102(2):1269–1286, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1470-9.

He:2013:RGS

[He13] Tianwei He. Retraction of global scientific publications from 2001 to 2010. Scientometrics, 96(2):555–561, August 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0906-3.

Hsiehchen:2016:DEB

[HE16] David Hsiehchen and Magdalena Espinoza. Detecting editorial bias in medical publishing. Scientometrics, 106(1): 453–456, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1753-9.

Hsiehchen:2017:DBA

[HEH17] David Hsiehchen, Magdalena Espinoza, and Antony Hsieh. Disease burden and the advancement of biomedical knowledge. Scientometrics, 110(1):321–333, January 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2169-x.

Hsiehchen:2018:ECO

[HEH18] David Hsiehchen, Magdalena Espinoza, and Antony Hsieh. Evolution of collaboration and optimization of impact: selforganization in multinational research. Scientometrics, 117 (1):391–407, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2886-4.

Heinze:2013:CAS

[Hei13] Thomas Heinze. Creative accomplishments in science: definition, theoretical considerations, examples from science history, and bibliometric findings. Scientometrics, 95(3):927– 940, June 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0848-9.

Heikkila:2019:IGG

[Hei19] Jussi Heikkilä. IPR gender gaps: a first look at utility model, design right and trademark filings. Scientometrics, 118(3): 869–883, March 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2979-0;
http://link.springer.com/content/pdf/10.1007/s11192-
018-2979-0.pdf.

Henriksen:2016:RCA

[Hen16] Dorte Henriksen. The rise in co-authorship in the social sciences (1980–2013). Scientometrics, 107(2):455–476, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1849-x.

Henriksen:2018:WFA

[Hen18] Dorte Henriksen. What factors are associated with increasing co-authorship in the social sciences? A case study of Danish economics and political science. Scientometrics, 114 (3):1395–1421, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2635-0.

Heneberg:2019:THP

[Hen19] Petr Heneberg. The troubles of high-profile open access megajournals. Scientometrics, 120(2):733–746, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03144-6. See correction

[Hen20].

Heneberg:2020:CTH

[Hen20] Petr Heneberg. Correction to: The troubles of highprofile open access megajournals. Scientometrics, 123(2): 1169–1171, May 2020. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03281-y;
 http://link.springer.com/content/pdf/10.1007/s11192-
019-03281-y.pdf.

Harrigan:2019:FIT

[HF19] Kathryn Rudie Harrigan and Yunzhe Fang. Financial implications of technology-class code popularity and usage among industry competitors. Scientometrics, 121(1):25–51, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03185-x.

Huang:2011:TRB

[HFC11] Ming-Chao Huang, Shih-Chieh Fang, and Shao-Chi Chang. Tracking R&D behavior: bibliometric analysis of drug patents in the orange book. Scientometrics, 88(3):805– 818, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0400-3.

Hu:2014:ESC

[HFL14] Zhengyin Hu, Shu Fang, and Tian Liang. Empirical study of constructing a knowledge organization system of patent documents using topic modeling. Scientometrics, 100(3):

787–799, September 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1328-1.

Henderson:2014:GDC

[HFW^{+}14] Michael T. Henderson, Natalia Fijalkowski, Sean K. Wang, Mitch Maltenfort, Luo Luo Zheng, John Ratliff, Andrew A. Moshfeghi, and Darius M. Moshfeghi. Gender differences in compensation in academic medicine: the results from four neurological specialties within the University of California Healthcare System. Scientometrics, 100(1):297–306, July 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1266-y.

Haddow:2010:CAP

[HG10] Gaby Haddow and Paul Genoni. Citation analysis and peer ranking of Australian social science journals. Scientometrics, 85(2):471–487, November 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0198-4.

Horlings:2013:SSA

[HG13] Edwin Horlings and Thomas Gurney. Search strategies along the academic lifecycle. Scientometrics, 94(3):1137– 1160, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-012-0789-
3.pdf.

Hilario:2017:SCB

[HG17] Carla Mara Hilário and Maria Cláudia Cabrini Grácio. Scientific collaboration in Brazilian researches: a comparative study in the information science, mathematics and dentistry fields. Scientometrics, 113(2):929–950, November 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2498-4.

Havemann:2017:MSO

[HGH17a] Frank Havemann, Jochen Gläser, and Michael Heinz. Memetic search for overlapping topics based on a local evaluation of link communities. Scientometrics, 111(2):1089– 1118, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Hu:2017:MRS

[HGH17b] Zhigang Hu, Fangqi Guo, and Haiyan Hou. Mapping research spotlights for different regions in China. Scientometrics, 110(2):779–790, February 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2175-z.

Hodder:2010:RCN

[HH10] A. Peter W. Hodder and Catherine Hodder. Research culture and new Zealand’s performance-based research fund: some insights from bibliographic compilations of research outputs. Scientometrics, 84(3):887–901, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0201-0.

Hassan:2013:MIK

[HH13] Saeed-Ul Hassan and Peter Haddawy. Measuring international knowledge flows and scholarly impact of scientific research. Scientometrics, 94(1):163–179, January 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0786-6.

Hassan:2015:AKF

[HH15a] Saeed-Ul Hassan and Peter Haddawy. Analyzing knowledge flows of scientific literature through semantic links: a case study in the field of energy. Scientometrics, 103(1):33–46, April 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1528-3.

Heidler:2015:QPE

[HH15b] Richard Heidler and Olof Hallonsten. Qualifying the performance evaluation of big science beyond productivity, impact and costs. Scientometrics, 104(1):295–312, July 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1577-7.

Ho:2015:PDF

[HH15c] Huei-Chen Ho and Yuh-Shan Ho. Publications in dance field in Arts & Humanities Citation Index: a bibliometric analysis. Scientometrics, 105(2):1031–1040, November 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1716-1.

Hartley:2017:DFB

[HH17a] James Hartley and Yuh-Shan Ho. The decline and fall of book reviews in psychology: a bibliometric analysis. Scientometrics, 112(1):655–657, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Hartley:2017:WWS

[HH17b] James Hartley and Yuh-Shan Ho. Who woke the sleeping beauties in psychology? Scientometrics, 112(2):1065–1068, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2326-x.

Ho:2017:HCP

[HH17c] Yuh-Shan Ho and James Hartley. Highly cited publications in World War II: a bibliometric analysis. Scientometrics, 110(2):1065–1075, February 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2199-4.

Ho:2017:SBP

[HH17d] Yuh-Shan Ho and James Hartley. Sleeping beauties in psychology. Scientometrics, 110(1):301–305, January 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2174-0.

Huang:2017:TFM

[HH17e] Jia-Yen Huang and Hung-Tu Hsu. Technology-function matrix based network analysis of cloud computing. Scientometrics, 113(1):17–44, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
017-2469-9.

Huang:2018:AGR

[HH18] Mu-Hsuan Huang and Mei-Jhen Huang. An analysis of global research funding from subject field and funding agencies perspectives in the G9 countries. Scientometrics, 115 (2):833–847, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2677-y.

Horbach:2019:ADP

[HH19] S. P. J. M. Horbach and W. Halffman. The ability of different peer review procedures to flag problematic publications. Scientometrics, 118(1):339–373, January 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2969-2;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2969-
2.pdf.

Huang:2016:RUV

[HHA^{+}16] Youliang Huang, Qian Huang, Sajid Ali, Xing Zhai, Xiaoming Bi, and Renquan Liu. Rehabilitation using virtual reality technology: a bibliometric analysis, 1996–2015. Scientometrics, 109(3):1547–1559, December 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2117-9.

Haunschild:2018:NLR

[HHBB18] Robin Haunschild, Sven E. Hug, Martin P. Brändle, and Lutz Bornmann. The number of linked references of publications in Microsoft Academic in comparison with the Web of Science. Scientometrics, 114(1):367–370, January 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2567-8.

Huang:2015:CSD

[hHC15] Mu hsuan Huang and Chia-Pin Chang. A comparative study on detecting research fronts in the organic light-emitting diode (OLED) field using bibliographic coupling and cocitation. Scientometrics, 102(3):2041–2057, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-014-1494-1.

Hu:2013:CWA

[HHDL13] Chang-Ping Hu, Ji-Ming Hu, Sheng-Li Deng, and Yong Liu. A co-word analysis of library and information science in China. Scientometrics, 97(2):369–382, November 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1076-7.

Hu:2011:JCC

[HHGZ11] Chang-Ping Hu, Ji-Ming Hu, Yan Gao, and Yao-Kun Zhang. A journal co-citation analysis of library and information science in China. Scientometrics, 86(3):657–670, March 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0313-6.

Hassan:2012:BSR

[HHK^{+}12] Saeed-Ul Hassan, Peter Haddawy, Pratikshya Kuinkel, Alexander Degelsegger, and Cosima Blasy. A bibliometric study of research activity in ASEAN related to the EU in FP7 priority areas. Scientometrics, 91(3):1035–1051, June 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0665-1.

Huang:2019:OCT

[hHSL19] Mu hsuan Huang, Wang-Ching Shaw, and Chi-Shiou Lin. One category, two communities: subfield differences in “information science and library science” in Journal Citation Reports. Scientometrics, 119(2):1059–1079, May 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03074-3.

Hassan:2014:BSW

[HHZ14] Saeed-Ul Hassan, Peter Haddawy, and Jia Zhu. A bibliometric study of the world’s research activity in sustainable development and its sub-areas using scientific literature. Scientometrics, 99(2):549–579, May 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1193-3.

Hou:2012:RAP

[HIC12] Angela Yung-Chi Hou, Martin Ince, and Chung-Lin Chiang. A reassessment of Asian Pacific excellence programs in higher education: the Taiwan experience. Scientometrics, 92(1):23–42, July 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0727-4.

Hassan:2017:MSM

[HIG^{+}17] Saeed-Ul Hassan, Mubashir Imran, Uzair Gillani, Naif Radi Aljohani, Timothy D. Bowman, and Fereshteh Didegah. Measuring social media activity of scientific literature: an exhaustive comparison of Scopus and novel altmetrics big data. Scientometrics, 113(2):1037–1057, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2512-x.

Hassan:2018:DCC

[HII^{+}18] Saeed-Ul Hassan, Mubashir Imran, Sehrish Iqbal, Naif Radi Aljohani, and Raheel Nawaz. Deep context of citations using machine-learning models in scholarly full-text articles. Scientometrics, 117(3):1645–1662, December 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2944-y.

Hirsch:2005:IQI

[Hir05] Jorge E. Hirsch. An index to quantify an individual’s scientific research output. Proceedings of the National Academy of Sciences of the United States of America, 102(56):16569– 16572, November 15, 2005. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.

pnas.org/content/102/46/16569.abstract.

Hirsch:2010:IQI

[Hir10] J. E. Hirsch. An index to quantify an individual’s scientific research output that takes into account the effect of multiple coauthorship. Scientometrics, 85(3):741–754, December 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0193-9.

Hirsch:2019:IQI

[Hir19a] J. E. Hirsch. h_{α}: An index to quantify an individual’s scientific leadership. Scientometrics, 118(2):673–686, February 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2994-1.

Hirsch:2019:RCS

[Hir19b] J. E. Hirsch. Response to comment “h_{α}: the scientist as chimpanzee or bonobo”, by Leydesdorff, Bornmann and Opthof. Scientometrics, 118(3):1167–1172, March 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03019-w.

Hanssen:2018:RBQ

[HJL18] Thor-Erik Sandberg Hanssen, Finn Jørgensen, and Berner Larsen. The relation between the quality of research, researchers’ experience, and their academic environment. Scientometrics, 114(3):933–950, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2580-y.

Hopkins:2013:DPP

[HJM^{+}13] Allison L. Hopkins, James W. Jawitz, Christopher McCarty, Alex Goldman, and Nandita B. Basu. Disparities in publication patterns by gender, race and ethnicity based on a survey of a random sample of authors. Scientometrics, 96 (2):515–534, August 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0893-4.

Honekopp:2012:FPS

[HK12] Johannes Hönekopp and Julie Khan. Future publication success in science is better predicted by traditional measures than by the h index. Scientometrics, 90(3):843–853, March 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0551-2.

Halim:2019:DSB

[HK19] Zahid Halim and Shafaq Khan. A data science-based framework to categorize academic journals. Scientometrics, 119 (1):393–423, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03035-w.

Hossain:2015:ELN

[HKWC15] Liaquat Hossain, Faezeh Karimi, Rolf T. Wigand, and John W. Crawford. Evolutionary longitudinal network dynamics of global zoonotic research. Scientometrics, 103 (2):337–353, May 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1557-y.

Ho:2013:MKT

[HL13] Mei Hsiu-Ching Ho and John S. Liu. The motivations for knowledge transfer across borders: the diffusion of data envelopment analysis (DEA) methodology. Scientometrics, 94 (1):397–421, January 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0705-x.

Havemann:2015:BIY

[HL15] Frank Havemann and Birger Larsen. Bibliometric indicators of young authors in astrophysics: Can later stars be predicted? Scientometrics, 102(2):1413–1434, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1476-3.

Hottenrott:2017:FLM

[HL17] Hanna Hottenrott and Cornelia Lawson. A first look at multiple institutional affiliations: a study of authors in Germany, Japan and the UK. Scientometrics, 111(1):285–295, April 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-017-2257-6.pdf.

Herzog:2018:RLF

[HL18] Christian Herzog and Brian Kierkegaard Lunn. Response to the letter ‘field classification of publications in Dimensions: a first case study testing its reliability and validity’. Scientometrics, 117(1):641–645, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2854-z;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2854-

z.pdf. See [Bor18].

Ho:2017:IRR

[HLC17] Mei Hsiu-Ching Ho, John S. Liu, and Kerr C.-T. Chang. To include or not: the role of review papers in citation-based analysis. Scientometrics, 110(1):65–76, January 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2158-0.

Han:2010:TPS

[HLE10] Chung-Souk Han, Su Kyung Lee, and Mark England. Transition to postmodern science-related scientometric data. Scientometrics, 84(2):391–401, August 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0119-6.

Ho:2014:EKD

[HLL14] Mei Hsiu-Ching Ho, Vincent H. Lin, and John S. Liu. Exploring knowledge diffusion among nations: a study of core technologies in fuel cells. Scientometrics, 100(1):149–171, July 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1265-z.

Hung:2014:TCL

[HLLT14] Shih-Chang Hung, John S. Liu, Louis Y. Y. Lu, and Yu-Chiang Tseng. Technological change in lithium iron phosphate battery: the key-route main path analysis. Scientometrics, 100(1):97–120, July 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
014-1276-9.

Hernandez-Linares:2018:IAH

[HLSC18] Remedios Hernández-Linares, Soumodip Sarkar, and Manuel J. Cobo. Inspecting the Achilles heel: a quantitative analysis of 50 years of family business definitions. Scientometrics, 115(2):929–951, May 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2702-1.

Hu:2018:EUM

[HLSW18] Zhigang Hu, Gege Lin, Taian Sun, and Xianwen Wang. An EU without the UK: mapping the UK’s changing roles in the EU scientific research. Scientometrics, 115(3):1185–1198, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2740-8.

Hu:2019:IRC

[HLW19] Zewen Hu, Angela Lin, and Peter Willett. Identification of research communities in cited and uncited publications using a co-authorship network. Scientometrics, 118(1):1–19, January 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2954-9.

Hua:2014:QAA

[HLY14] Weina Hua, Yu Li, and Shunbo Yuan. A quantitative analysis of Antarctic related articles in humanities and social sciences appearing in the world core journals. Scientometrics, 100(1):273–286, July 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1190-6.

Harzing:2015:EPP

[HM15a] Anne-Wil Harzing and Wilfred Mijnhardt. Erratum to: Proof over promise: towards a more inclusive ranking of Dutch academics in Economics & Business. Scientometrics, 102(1):751–752, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-014-1511-4.pdf. See [HM15b].

Harzing:2015:PPT

[HM15b] Anne-Wil Harzing and Wilfred Mijnhardt. Proof over promise: towards a more inclusive ranking of Dutch academics in economics & business. Scientometrics, 102(1):727–749, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-

014-1370-z. See erratum [HM15a].

Haucap:2015:WDR

[HM15c] Justus Haucap and Johannes Muck. What drives the relevance and reputation of economics journals? An update from a survey among economists. Scientometrics, 103(3): 849–877, June 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1542-5.

Han:2018:TST

[HM18] Fang Han and Christopher L. Magee. Testing the science/ technology relationship by analysis of patent citations of scientific papers after decomposition of both science and technology. Scientometrics, 116(2):767–796, August 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2774-y.

Hernandez-Martin:2019:PWD

[HMCD^{+}19] Estela Hernández-Martín, Fernando Calle, Juan C. Dueñas, Miguel Holgado, and Asunción Gómez-Pérez. Participation of women in doctorate, research, innovation, and management activities at Universidad Politécnica de Madrid: analysis of the decade 2006–2016. Scientometrics, 120(3):1059– 1089, September 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03179-9.

Herron:2016:RDI

[HMCL16] Patrick Herron, Aashish Mehta, Cong Cao, and Timothy Lenoir. Research diversification and impact: the case of national nanoscience development. Scientometrics, 109(2):629–659, November 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2062-7.

Hicks:2019:CPM

[HMI19] Diana Hicks, Julia Melkers, and Kimberley R. Isett. A characterization of professional media and its links to research. Scientometrics, 119(2):827–843, May 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03072-5.

Hossain:2012:MDK

[HMK^{+}12] Md. Dulal Hossain, Junghoon Moon, Hyoung Goo Kang, Sung Chul Lee, and Young Chan Choe. Mapping the dynamics of knowledge base of innovations of R&D in Bangladesh: triple helix perspective. Scientometrics, 90 (1):57–83, January 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0507-6.

Hosotsubo:2016:RBA

[HN16] Moritaka Hosotsubo and Ryuei Nishii. Relation between awarding of grants-in-aid for scientific research and characteristics of applicants in Japanese universities. Scientometrics, 109(2):1097–1116, November 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2074-3.

Hoorani:2019:DIE

[HNG19] Bareerah Hafeez Hoorani, Lakshmi Balachandran Nair, and Michael Gibbert. Designing for impact: the effect of rigor and case study design on citations of qualitative case studies in management. Scientometrics, 121(1):285–306, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03178-w.

Ho:2013:CBS

[Ho13a] Yuh-Shan Ho. Comments on “A bibliometric study of earthquake research: 1900–2010”. Scientometrics, 96(3):929–931, September 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-

012-0915-2. See [LZH^{+}12, LZH^{+}13].

Ho:2013:TCR

[Ho13b] Yuh-Shan Ho. The top-cited research works in the science citation index expanded. Scientometrics, 94(3):1297–1312, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0837-z.

Ho:2014:CAS

[Ho14] Yuh-Shan Ho. Classic articles on social work field in social Science Citation Index: a bibliometric analysis. Scientometrics, 98(1):137–155, January 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1014-8.

Ho:2016:RLA

[Ho16] Yuh-Shan Ho. Rebuttal to: Liu et al. “Progress in global parallel computing research: a bibliometric approach”, vol. 95, pp 967–983. Scientometrics, 108(3):1693–1694, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-1993-3. See

[LLGW13].

Haugen:2019:CRP

[HO19] Kjetil K. Haugen and Kai A. Olsen. Could requiring a presentation of the paper and adding a formalized contributor list solve academia’s credibility problem? Scientometrics, 121(2):1229–1233, November 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03229-2.

Hug:2017:CAM

[HOB17] Sven E. Hug, Michael Ochsner, and Martin P. Brändle. Citation analysis with Microsoft Academic. Scientometrics, 111(1):371–378, April 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-017-2247-8.

Holmberg:2010:CIM

[Hol10] Kim Holmberg. Co-inlinking to a municipal Web space: a webometric and content analysis. Scientometrics, 83 (3):851–862, June 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0148-1.

Horta:2018:DSW

[Hor18] Hugo Horta. The declining scientific wealth of Hong Kong and Singapore. Scientometrics, 117(1):427–447, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2845-0.

Hosotsubo:2011:SST

[Hos11] Moritaka Hosotsubo. A statistical study of transferral and promotion mechanisms relating to the appointment of professors at Japanese national universities based on cross tabulation and log-linear model analysis. Scientometrics, 86(2): 405–430, February 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0276-7.

Hou:2017:EEH

[Hou17] Jianhua Hou. Exploration into the evolution and historical roots of citation analysis by referenced publication year spectroscopy. Scientometrics, 110(3):1437–1452, March 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2206-9.

Hypponen:2010:CSR

[HP10] Konstantin Hyppönen and Vivian Michael Paganuzzi. Computer science research articles: the locations of different section types, and a proposal for standardization in the structure. Scientometrics, 84(1):199–220, July 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0089-8.

Holmberg:2018:AIO

[HP18] Kim Holmberg and Han Woo Park. An altmetric investigation of the online visibility of South Korea-based scientific journals. Scientometrics, 117(1):603–613, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2874-8.

Haustein:2014:CAA

[HPBI^{+}14] Stefanie Haustein, Isabella Peters, Judit Bar-Ilan, Jason Priem, Hadas Shema, and Jens Terliesner. Coverage and adoption of altmetrics sources in the bibliometric community. Scientometrics, 101(2):1145–1163, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-1221-3.

Herrmannova:2018:DCR

[HPKS18] Drahomira Herrmannova, Robert M. Patton, Petr Knoth, and Christopher G. Stahl. Do citations and readership identify seminal publications? Scientometrics, 115(1):239–262, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2669-y.

Heibi:2019:SRC

[HPS19] Ivan Heibi, Silvio Peroni, and David Shotton. Software review: COCI, the OpenCitations Index of Crossref open DOI-to-DOI citations. Scientometrics, 121 (2):1213–1228, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03217-6;   http://link.springer.com/content/
pdf/10.1007/s11192-019-03217-6.pdf.

Hu:2018:IGC

[HQY^{+}18] Kai Hu, Kunlun Qi, Siluo Yang, Shengyu Shen, Xiaoqiang Cheng, Huayi Wu, Jie Zheng, Stephen McClure, and Tianxing Yu. Identifying the “Ghost City” of domain topics in a keyword semantic space combining citations. Scientometrics, 114(3):1141–1157, March 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2604-7.

Helene:2011:BSP

[HR11] André Frazão Helene and Pedro Leite Ribeiro. Brazilian scientific production, financial support, established investigators and doctoral graduates. Scientometrics, 89(2): 677–686, November 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0470-2.

Hu:2015:SAD

[HR15] Xiaojun Hu and Ronald Rousseau. A simple approach to describe a company’s innovative activities and their technological breadth. Scientometrics, 102(2):1401–1411, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1464-7.

Hu:2017:NPW

[HR17] Xiaojun Hu and Ronald Rousseau. Nobel Prize winners 2016: Igniting or sparking foundational publications? Scientometrics, 110(2):1053–1063, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2205-x.

Horlesberger:2013:CIF

[HRB^{+}13] Marianne Hörlesberger, Ivana Roche, Dominique Besagni, Thomas Scherngell, Claire François, Pascal Cuxac, Edgar Schiebel, Michel Zitt, and Dirk Holste. A concept for inferring ‘frontier research’ in grant proposals. Scientometrics, 97(2):129–148, November 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1008-6. See erratum [HRB^{+}14].

Horlesberger:2014:ECI

[HRB^{+}14] Marianne Hörlesberger, Ivana Roche, Dominique Besagni, Thomas Scherngell, Claire François, Pascal Cuxac, Edgar Schiebel, Michel Zitt, and Dirk Holste. Erratum to: A concept for inferring ‘frontier research’ in grant proposals. Scientometrics, 100(2):609–610, August 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-014-1277-8.pdf. See [HRB^{+}13].

Hu:2012:SIC

[HRC12] Xiaojun Hu, Ronald Rousseau, and Jin Chen. Structural indicators in citation networks. Scientometrics, 91(2):451– 460, May 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0587-3.

Herranz:2013:EEP

[HRC13] Neus Herranz and Javier Ruiz-Castillo. The end of the “European Paradox”. Scientometrics, 95(1):453–464, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0865-8.

Hayashi:2010:SKD

[HRH10] Maria Cristina Piumbato Innocentini Hayashi, Danilo Rothberg, and Carlos Roberto Massao Hayashi. Scientific knowledge and digital democracy in Brazil: how to assess public health policy debate with applied scientometrics. Scientometrics, 83(3):825–833, June 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0125-8.

Haugen:2016:NNI

[HS16a] Kjetil K. Haugen and Frode E. Sandnes. The new Norwegian incentive system for publication: from bad to worse. Scientometrics, 109(2):1299–1306, November 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2075-2.

Horta:2016:IMI

[HS16b] Hugo Horta and João M. Santos. An instrument to measure individuals’ research agenda setting: the multidimensional research agendas inventory. Scientometrics, 108(3):1243–1265, September 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2012-4.

Huisman:2017:DQP

[HS17] Janine Huisman and Jeroen Smits. Duration and quality of the peer review process: the author’s perspective. Scientometrics, 113(1):633–650, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2310-5;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2310-
5.pdf.

Hassan:2018:NML

[HSAK18] Saeed-Ul Hassan, Iqra Safder, Anam Akram, and Faisal Kamiran. A novel machine-learning approach to measuring scientific knowledge flows using citation context analysis. Scientometrics, 116(2):973–996, August 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2767-x.

Hofer:2010:CPM

[HSBW10] Katharina Maria Hofer, Angela Elisabeth Smejkal, F. Zeynep Bilgin, and Gerhard A. Wuehrer. Conference proceedings as a matter of bibliometric studies: the Academy of International Business 2006–2008. Scientometrics, 84(3):845– 862, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0216-6.

Hsieh:2011:ESU

[Hsi11] Chihmao Hsieh. Explicitly searching for useful inventions: dynamic relatedness and the costs of connecting versus synthesizing. Scientometrics, 86(2):381–404, February 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s11192-010-0290-9.pdf.

Huang:2017:SPT

[HSK17] Hung-Chun Huang, Hsin-Yu Shih, and Tsung-Han Ke. Structure of a patent transaction network. Scientometrics, 111(1):25–45, April 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-017-2258-5.

Henrique:2018:BDC

[HSK18] Bruno Miranda Henrique, Vinicius Amorim Sobreiro, and Herbert Kimura. Building direct citation networks. Scientometrics, 115(2):817–832, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2676-z.

Han:2014:ICL

[HSL^{+}14] Pu Han, Jin Shi, Xiaoyan Li, Dongbo Wang, Si Shen, and Xinning Su. International collaboration in LIS: global trends and networks at the country and institution level. Scientometrics, 98(1):53–72, January 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1146-x.

Hu:2014:SSG

[HSLP14] Yue Hu, Jun Sun, Weimin Li, and Yunlong Pan. A scientometric study of global electric vehicle research. Scientometrics, 98(2):1269–1282, February 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1067-8.

Huang:2015:SMC

[HSPY15] Ying Huang, Jannik Schuehle, Alan L. Porter, and Jan Youtie. A systematic method to create search strategies for emerging technologies based on the Web of Science: illustrated for ‘big data’. Scientometrics, 105(3):2005–2022, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1638-y.

Henzinger:2010:SI

[HSW10] Monika Henzinger, Jacob Suñol, and Ingmar Weber. The stability of the h-index. Scientometrics, 84(2):465–479, August 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0098-7.

Huang:2013:EPP

[HSWC13] Mu-Hsuan Huang, Hui-Yun Sung, Chun-Chieh Wang, and Dar-Zen Chen. Exploring patent performance and technology interactions of universities, industries, governments and individuals. Scientometrics, 96(1):11–26, July 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0878-3.

Huang:2015:BSC

[HSX^{+}15] Cui Huang, Jun Su, Xiang Xie, Xuanting Ye, Zhang Li, Alan Porter, and Jiang Li. A bibliometric study of China’s science and technology policies: 1949–2010. Scientometrics, 102(2): 1521–1539, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1406-4.

Huang:2014:BRO

[HSXL14] Cui Huang, Jun Su, Xiang Xie, and Jiang Li. Basic research is overshadowed by applied research in China: a policy perspective. Scientometrics, 99(3):689–694, June 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1199-x.

Halkos:2011:MEJ

[HT11] George Emm Halkos and Nickolaos G. Tzeremes. Measuring economic journals’ citation efficiency: a data envelopment analysis approach. Scientometrics, 88(3):979–1001, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0421-y.

Holmberg:2014:DDT

[HT14] Kim Holmberg and Mike Thelwall. Disciplinary differences in Twitter scholarly communication. Scientometrics, 101(2): 1027–1042, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1229-3.

Heffernan:2018:IPS

[HT18] Kevin Heffernan and Simone Teufel. Identifying problems and solutions in scientific text. Scientometrics, 116(2):

1367–1382, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2718-6;
http://link.springer.com/content/pdf/10.1007/s11192-
018-2718-6.pdf.

Horstman:2019:CHM

[HT19] Katelyn Horstman and Virginia Trimble. A citation history of measurements of Newton’s constant of gravity. Scientometrics, 119(1):527–541, April 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03031-0;   http://link.springer.
com/content/pdf/10.1007/s11192-019-03031-0.pdf.

Haustein:2011:RDI

[HTHB11] Stefanie Haustein, Dirk Tunger, Gerold Heinrichs, and Gesa Baelz. Reasons for and developments in international scientific collaboration: does an Asia–Pacific research area exist from a bibliometric point of view? Scientometrics, 86 (3):727–746, March 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0295-4.

Hsiao:2015:CBA

[HTL15] Chun-Hua Hsiao, Kai-Yu Tang, and John S. Liu. Citationbased analysis of literature: a case study of technology acceptance research. Scientometrics, 105(2):1091– 1110, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1749-5.

Hu:2011:EKC

[Hu11] Mei-Chih Hu. Evolution of knowledge creation and diffusion: the revisit of Taiwan’s Hsinchu Science Park. Scientometrics, 88(3):949–977, September 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0427-5.

Hudson:2016:ATP

[Hud16] John Hudson. An analysis of the titles of papers submitted to the UK REF in 2014: authors, disciplines, and stylistic details. Scientometrics, 109(2):871–889, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-016-2081-4.pdf.

Hudson:2017:IEP

[Hud17] John Hudson. Identifying economics’ place amongst academic disciplines: a science or a social science? Scientometrics, 113(2):735–750, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
017-2519-3;   http://link.springer.com/content/pdf/
10.1007/s11192-017-2519-3.pdf.

Hung:2012:MUP

[Hun12] Wen Chi Hung. Measuring the use of public research in firm R&D in the Hsinchu Science Park. Scientometrics, 92(1):63– 73, July 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0726-5.

Hur:2017:PKS

[Hur17] Wonchang Hur. The patterns of knowledge spillovers across technology sectors evidenced in patent citation networks. Scientometrics, 111(2):595–619, May 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Heikkila:2018:RUM

[HV18a] Jussi Heikkilä and Michael Verba. The role of utility models in patent filing strategies: evidence from European countries. Scientometrics, 116(2):689–719, August 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2773-z.

Holmberg:2018:WDS

[HV18b] Kim Holmberg and Julia Vainio. Why do some research articles receive more online attention and higher altmetrics? Reasons for online success according to the authors. Scientometrics, 116(1):435–447, July 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2710-1.

Hung:2010:ESW

[HW10] Shiu-Wan Hung and An-Pang Wang. Examining the small world phenomenon in the patent citation network: a case study of the radio frequency identification (RFID) network. Scientometrics, 82(1):121–134, January 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0032-z.

Huang:2012:RCH

[HW12] Yi Huang and Jie Wang. Response to Chuang and Ho’s comments on “A bibliometric study of the trend in articles related to eutrophication published in Science Citation Index”. Scientometrics, 91(3):1067–1071, June 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0607-3. See [YJ11, CH12].

Hennemann:2011:MRS

[HWL11] S. Hennemann, T. Wang, and I. Liefner. Measuring regional science networks in China: a comparison of international and domestic bibliographic data sources. Scientometrics, 88 (2):535–554, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0410-1.

Hu:2014:IKS

[HWLL14] Mei-Chih Hu, Ching-Yan Wu, Jung Hoon Lee, and Yun-Chu Lu. The influence of knowledge source and ambidexterity in the thin film transistor and liquid crystal display industry: evidence from Japan, Korea, and Taiwan. Scientometrics, 99(2):233–260, May 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1226-y.

Hu:2018:DKA

[HWQ^{+}18] Kai Hu, Huayi Wu, Kunlun Qi, Jingmin Yu, Siluo Yang, Tianxing Yu, Jie Zheng, and Bo Liu. A domain keyword analysis approach extending Term Frequency– Keyword Active Index with Google Word2Vec model. Scientometrics, 114(3):1031–1068, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2574-9.

Hu:2018:QAD

[HWS18] Zewen Hu, Yishan Wu, and Jianjun Sun. A quantitative analysis of determinants of non-citation using a panel data model. Scientometrics, 116(2):843–861, August 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2791-x.

Hou:2019:PSB

[HY19] Jianhua Hou and Xiucai Yang. Patent sleeping beauties: evolutionary trajectories and identification methods. Scientometrics, 120(1):187–215, July 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03123-x.

Huang:2015:IAC

[HYC15] Mu-Hsuan Huang, Hsiao-Wen Yang, and Dar-Zen Chen. Industry-academia collaboration in fuel cells: a perspective from paper and patent analysis. Scientometrics, 105(2): 1301–1318, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1748-6.

Hou:2018:ETN

[HYC18] Jianhua Hou, Xiucai Yang, and Chaomei Chen. Emerging trends and new developments in information science: a document co-citation analysis (2009–2016). Scientometrics, 115(2):869–892, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2695-9.

Han:2017:SFB

[HYF^{+}17] Hongqi Han, Changqing Yao, Yuan Fu, Yongsheng Yu, Yunliang Zhang, and Shuo Xu. Semantic fingerprints-based author name disambiguation in Chinese documents. Scientometrics, 111(3):1879–1896, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Huang:2018:PCA

[HYS18] Cui Huang, Chao Yang, and Jun Su. Policy change analysis based on “policy target-policy instrument” patterns: a case study of China’s nuclear energy policy. Scientometrics, 117(2):1081–1114, November 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2899-z.

Hua:2012:QAA

[HYYL12] Weina Hua, Shunbo Yuan, Miaomiao Yan, and Yu Li. A quantitative analysis of Arctic related articles in the humanities and social sciences appearing in the world core journals. Scientometrics, 91(3):703–718, June 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0690-0.

Huang:2014:IND

[HYYR14] Shuiqing Huang, Bo Yang, Sulan Yan, and Ronald Rousseau. Institution name disambiguation for research assessment. Scientometrics, 99(3):823–838, June 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1214-2.

Hu:2017:DIN

[HZ17] Jiming Hu and Yin Zhang. Discovering the interdisciplinary nature of Big Data research through social network analysis and visualization. Scientometrics, 112(1):91–109, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Huang:2015:BAL

[HZD^{+}15] Youliang Huang, Mingquan Zhou, Qingqiong Deng, Juan Zhang, Pengbo Zhou, and XianGang Shang. Bibliometric analysis for the literature of traditional Chinese medicine in PubMed. Scientometrics, 105(1):557–566, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1686-3.

Huang:2017:EIE

[HZL^{+}17] Ying Huang, Donghua Zhu, Qi Lv, Alan L. Porter, Douglas K. R. Robinson, and Xuefeng Wang. Early insights on the Emerging Sources Citation Index (ESCI): an overlay mapbased bibliometric study. Scientometrics, 111(3):2041–2057, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Huang:2017:HMT

[HZQ^{+}17] Ying Huang, Donghua Zhu, Yue Qian, Yi Zhang, Alan L. Porter, Yuqin Liu, and Ying Guo. A hybrid method to trace technology evolution pathways: a case study of 3D printing. Scientometrics, 111(1):185–204, April 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-017-2271-8.

Ikram:2019:ABC

[IA19] Muhammad Touseef Ikram and Muhammad Tanvir Afzal. Aspect based citation sentiment analysis using linguistic patterns for better comprehension of scientific knowledge. Scientometrics, 119(1):73–95, April 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03028-9.

Iyengar:2015:RJU

[IB15] Kishen Iyengar and Venugopal Balijepally. Ranking journals using the dominance hierarchy procedure: an illustration with IS journals. Scientometrics, 102(1):5–23, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1444-y.

Ibanez:2013:RAR

[IBL13] Alfonso Ibáñez, Concha Bielza, and Pedro Larrañaga. Relationship among research collaboration, number of documents and number of citations: a case study in Spanish computer science production in 2000–2009. Scientometrics, 95(2):689–716, May 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0883-6.

Ibrahim:2018:ASE

[Ibr18] Bahaa Ibrahim. Arab Spring’s effect on scientific productivity and research performance in Arab countries. Scientometrics, 117(3):1555–1586, December 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2935-z.

Izquierdo:2016:ACA

[ICC16] Javier Luis Cánovas Izquierdo, Valerio Cosentino, and Jordi Cabot. Analysis of co-authorship graphs of CORE-ranked software conferences. Scientometrics, 109(3):1665–1693, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2136-6.

Iglic:2017:WDR

[IDKF17] Hajdeja Iglic, Patrick Doreian, Luka Kronegger, and Anuska Ferligoj. With whom do researchers collaborate and why? Scientometrics, 112(1):153–174, July 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-017-2386-y.pdf.

Ida:2013:ELS

[IF13] Takanori Ida and Naomi Fukuzawa. Effects of largescale research funding programs: a Japanese case study. Scientometrics, 94(3):1253–1273, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0841-3.

Ivanovic:2015:PSS

[IFH15] Dragan Ivanović, Hui-Zhen Fu, and Yuh-Shan Ho. Publications from Serbia in the Science Citation Index Expanded: a bibliometric analysis. Scientometrics, 105(1):145–160, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1664-9.

Ibarra:2018:CPA

[IFT^{+}18] Mariano Esteban Ibarra, Juan Pablo Ferreira, Milagros Torrents, Magalí Hamui, Fernando Torres, Paula Dominguez, María Fabiana Ossorio, and Fernando Ferrero. Changes in PubMed affiliation indexing improved publication identification by country. Scientometrics, 115(3):1365–1370, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2714-x.

Ivanovic:2014:IPS

[IH14] Dragan Ivanović and Yuh-Shan Ho. Independent publications from Serbia in the science citation index expanded: a bibliometric analysis. Scientometrics, 101(1):603–622, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1396-2.

Ivanovic:2016:ASP

[IJF16] Dragan Ivanović, Milos Jovanović, and Frank Fritsche. Analysis of scientific productivity and cooperation in the republics of former Yugoslavia before, during and after the Yugoslav wars. Scientometrics, 107(2):499–519, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-016-1853-1.

Ingwersen:2014:IPI

[IL14a] Peter Ingwersen and Birger Larsen. Influence of a performance indicator on Danish research production and citation impact 2000–12. Scientometrics, 101(2):1325–1344, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1291-x.

Ivanova:2014:SMT

[IL14b] Inga A. Ivanova and Loet Leydesdorff. A simulation model of the triple helix of university-industry-government relations and the decomposition of the redundancy. Scientometrics, 99(3):927–948, June 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
014-1241-7.

Ibanez:2011:UBN

[ILB11] Alfonso Ibáñez, Pedro Larrañaga, and Concha Bielza. Using Bayesian networks to discover relationships between bibliometric indices. A case study of computer science and artificial intelligence journals. Scientometrics, 89(2):523– 551, November 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0486-7.

Ibanez:2013:CMA

[ILB13] Alfonso Ibáñez, Pedro Larrañaga, and Concha Bielza. Cluster methods for assessing research performance: exploring Spanish computer science. Scientometrics, 97(3):571– 600, December 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-0985-9.

Inglesi-Lotz:2014:TVC

[ILBG14] Roula Inglesi-Lotz, Mehmet Balcilar, and Rangan Gupta. Time-varying causality between research output and economic growth in US. Scientometrics, 100(1):203–216, July 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1257-z.

Ingwersen:2014:IPP

[ILGZ^{+}14] Peter Ingwersen, Birger Larsen, J. Carlos Garcia-Zorita, Antonio Eleazar Serrano-López, and Elias Sanz-Casado. Influence of proceedings papers on citation impact in seven sub-fields of sustainable energy research 2005–2011. Scientometrics, 101(2):1273–1292, November 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1335-2.

Inglesi-Lotz:2011:SIA

[ILP11] Roula Inglesi-Lotz and Anastassios Pouris. Scientometric impact assessment of a research policy instrument: the case of rating researchers on scientific outputs in South Africa. Scientometrics, 88(3):747–760, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0440-8.

Inglesi-Lotz:2013:ISR

[ILP13] R. Inglesi-Lotz and A. Pouris. The influence of scientific research output of academics on economic growth in South Africa: an autoregressive distributed lag (ARDL) application. Scientometrics, 95(1):129–139, April 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0817-3.

Isfandyari-Moghaddam:2013:SFI

[IMH13] Alireza Isfandyari-Moghaddam and Mohammad Hasanzadeh. A study of factors inhibiting research productivity of Iranian women in ISI. Scientometrics, 95(2):797–815, May 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-0980-1.

Isfandyari-Moghaddam:2012:SFA

[IMHG12] Alireza Isfandyari-Moghaddam, Mohammad Hasanzadeh, and Zainab Ghayoori. A study of factors affecting research productivity of Iranian women in ISI. Scientometrics, 91 (1):159–172, April 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0585-5.

Iwami:2014:DME

[IMSK14] Shino Iwami, Junichiro Mori, Ichiro Sakata, and Yuya Kajikawa. Detection method of emerging leading papers using time transition. Scientometrics, 101(2):1515–1533, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1380-x.

Ingwersen:2012:PDJ

[Ing12] Peter Ingwersen. The pragmatics of a diachronic journal impact factor. Scientometrics, 92(2):319–324, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0701-1.

Ioannidis:2006:CMC

[Ioa06] John P. A. Ioannidis. Concentration of the most-cited papers in the scientific literature: Analysis of journal ecosystems. PLoS One, 1(1):e5:1–e5:7, 12 2006. CODEN POL-

NCL. ISSN 1932-6203. URL http://dx.plos.org/10.
1371/journal.pone.0000005.

Inigo:2013:EPC

[IPIU13] Jesus Iñigo, Jose-Alberto Palma, Jorge Iriarte, and Elena Urrestarazu. Evolution of the publications in clinical neurology: scientific impact of different countries during the 2000– 2009 period. Scientometrics, 95(3):941–952, June 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0880-9.

Iqbal:2019:BAP

[IQT^{+}19] Waleed Iqbal, Junaid Qadir, Gareth Tyson, Adnan Noor Mian, Saeed ul Hassan, and Jon Crowcroft. A bibliometric analysis of publications in computer networking research. Scientometrics, 119(2):1121–1155, May 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03086-z.

Igami:2016:DDJ

[IS16] Masatsura Igami and Ayaka Saka. Decreasing diversity in Japanese science, evidence from in-depth analyses of science maps. Scientometrics, 106(1):383–403, January 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1648-9.

Ingwersen:2018:SCR

[ISL18] Peter Ingwersen and Antonio Eleazar Serrano-López. Smart city research 1990–2016. Scientometrics, 117(2):1205– 1236, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2901-9.

Ivanovic:2011:CDM

[ISR11] Dragan Ivanović, Dusan Surla, and Milos Racković. A CERIF data model extension for evaluation and quantitative expression of scientific research results. Scientometrics, 86(1):155–172, January 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
010-0228-2.

Inanc:2011:EAI

[IT11] Ozlem Inanc and Onur Tuncer. The effect of academic inbreeding on scientific effectiveness. Scientometrics, 88(3): 885–898, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0415-9.

Iwami:2017:SDR

[Iwa17] Shino Iwami. Study on the destination of research via knowledge flows. Scientometrics, 112(1):273–288, July 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Iefremova:2018:BAS

[IWK18] Olesia Iefremova, Kamil Wais, and Marcin Kozak. Biographical articles in scientific literature: analysis of articles indexed in Web of Science. Scientometrics, 117(3):1695– 1719, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2923-3;
http://link.springer.com/content/pdf/10.1007/s11192-
018-2923-3.pdf.

Jamali:2018:IRC

[JAAA18] Hamid R. Jamali, Ghasem Azadi-Ahmadabadi, and Saeid Asadi. Interdisciplinary relations of converging technologies: Nano-Bio-Info-Cogno (NBIC). Scientometrics, 116(2): 1055–1073, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2776-9.

Jacso:2012:GTA

[Jac12] Peter Jacso. Grim tales about the impact factor and the h-index in the Web of Science and the Journal Citation Reports databases: reflections on Vanclay’s criticism. Scientometrics, 92(2):325–354, August 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0769-7.

Jacso:2018:SPE

[Jac18] Peter Jacso. The scientometric portrait of Eugene Garfield through the free ResearcherID service from the Web of Science Core Collection of 67 million master records and 1.3 billion references. Scientometrics, 114(2):545–555, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2624-3.

Jamali:2017:CCI

[Jam17] Hamid R. Jamali. Copyright compliance and infringement in ResearchGate full-text journal articles. Scientometrics, 112(1):241–254, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Jarneving:2010:RRF

[Jar10] Bo Jarneving. Regional research and foreign collaboration. Scientometrics, 83(1):295–320, April 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0064-4.

Jaric:2016:HTC

[Jar16] Ivan Jarić. High time for a common plagiarism detection system. Scientometrics, 106(1):457–459, January 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1756-6.

Ji:2019:GNG

[JBC19] Jiaojiao Ji, George A. Barnett, and Jianxun Chu. Global networks of genetically modified crops technology: a patent citation network analysis. Scientometrics, 118(3):737–762, March 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03006-1.

Jeremic:2011:FAE

[JBMR11] Veljko Jeremic, Milica Bulajic, Milan Martic, and Zoran Radojicic. A fresh approach to evaluating the academic ranking of world universities. Scientometrics, 87(3):587– 596, June 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0361-6.

Jang:2011:WDH

[JC11] Show-Ling Jang and Jennifer H. Chen. What determines how long an innovative spell will last? Scientometrics, 86 (1):65–76, January 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0247-z.

Jeong:2012:TRC

[JC12] Seongkyoon Jeong and Jae Young Choi. The taxonomy of research collaboration in science and technology: evidence from mechanical research through probabilistic clustering analysis. Scientometrics, 91(3):719–735, June 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0686-9.

Jiang:2019:EPS

[JC19] Shan Jiang and Hsinchun Chen. Examining patterns of scientific knowledge diffusion based on knowledge cyber infrastructure: a multi-dimensional network approach. Scientometrics, 121(3):1599–1617, December 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03242-5.

Jang:2013:IPG

[JCCC13] Show-Ling Jang, Li-Ju Chen, Jennifer H. Chen, and Yu-Chieh Chiu. Innovation and production in the global solar photovoltaic industry. Scientometrics, 94(3):1021–1036, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0907-2.

Jeong:2011:DRC

[JCK11] Seongkyoon Jeong, Jae Young Choi, and Jaeyun Kim. The determinants of research collaboration modes: exploring the effects of research and researcher characteristics on coauthorship. Scientometrics, 89(3):967–983, December 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0474-y.

Jia:2014:CEU

[JDG14] Xiaofeng Jia, Tao Dai, and Xinbiao Guo. Comprehensive exploration of urban health by bibliometric analysis: 35 years and 11,299 articles. Scientometrics, 99(3):881–894, June 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1220-4.

Jones:2012:TWI

[JDH12] Teresa H. Jones, Claire Donovan, and Steve Hanney. Tracing the wider impacts of biomedical research: a literature search to develop a novel citation categorisation technique. Scientometrics, 93(1):125–134, October 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-012-0642-8.pdf.

Jonkers:2014:MSI

[JDLIV14] K. Jonkers, G. E. Derrick, C. Lopez-Illescas, and P. Van den Besselaar. Measuring the scientific impact of eresearch infrastructures: a citation based approach? Scientometrics, 101(2):1179–1194, November 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1411-7.

Jaric:2012:APS

[JG12] I. Jarić and J. Gessner. Analysis of publications on sturgeon research between 1996 and 2010. Scientometrics, 90(2):715–735, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-

011-0583-7. See erratum [JG14].

Jaric:2014:EAP

[JG14] I. Jarić and J. Gessner. Erratum to: Analysis of publications on sturgeon research between 1996 and 2010. Scientometrics, 98(1):775, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-012-

0710-0.pdf. See [JG12].

Jones:2016:TIS

[JH16] Teresa H. Jones and Steve Hanney. Tracing the indirect societal impacts of biomedical research: development and piloting of a technique based on citations. Scientometrics, 107(3):975–1003, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com//content/
pdf/10.1007/s11192-016-1895-4.pdf;   http://link.
springer.com/content/pdf/10.1007/s11192-016-1895-
4.pdf.

Jovanovic:2010:ECW

[JJR10] Milos M. Jovanović, Marcus John, and Stefan Reschke. Effects of civil war: scientific cooperation in the republics of the former Yugoslavia and the province of Kosovo. Scientometrics, 82(3):627–645, March 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0176-x.

Jovanovic:2012:HDN

[JJS^{+}12] Milica Jovanovic, Veljko Jeremic, Gordana Savic, Milica Bulajic, and Milan Martic. How does the normalization of data affect the ARWU ranking? Scientometrics, 93(2): 319–327, November 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0674-0.

Jeong:2010:ISB

[JK10a] Senator Jeong and Hong-Gee Kim. Intellectual structure of biomedical informatics reflected in scholarly events. Scientometrics, 85(2):541–551, November 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0166-z.

Joo:2010:MRB

[JK10b] Si Hyung Joo and Yeonbae Kim. Measuring relatedness between technological fields. Scientometrics, 83(2):435–454, May 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0108-9.

Jang:2019:HLC

[JK19] Young-Sun Jang and Young Joo Ko. How latecomers catch up to leaders in high-energy physics as Big Science: transition from national system to international collaboration. Scientometrics, 119(1):437–480, April 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03030-1.

Jeong:2015:TCW

[JKC15] Seongkyoon Jeong, Jong-Chan Kim, and Jae Young Choi. Technology convergence: What developmental stage are we in? Scientometrics, 104(3):841–871, September 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1606-6.

Jaric:2014:RAR

[JKJL14] Ivan Jarić, Jelena Knezević-Jarić, and Mirjana Lenhardt. Relative age of references as a tool to identify emerging research fields with an application to the field of ecology and environmental sciences. Scientometrics, 100(2):519–529, August 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1268-9.

Jurajda:2017:SPP

[JKMS17] Stepán Jurajda, Stanislav Kozubek, Daniel Münich, and Samuel Skoda. Scientific publication performance in postcommunist countries: still lagging far behind. Scientometrics, 112(1):315–328, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Jaidka:2019:CHS

[JKN19] Kokil Jaidka, Christopher S. G. Khoo, and Jin-Cheon Na. Characterizing human summarization strategies for text reuse and transformation in literature review writing. Scientometrics, 121(3):1563–1582, December 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03250-5.

Jang:2018:PDI

[JKPL18] Wooseok Jang, Heeyeul Kwon, Yongtae Park, and Hakyeon Lee. Predicting the degree of interdisciplinarity in academic fields: the case of nanotechnology. Scientometrics, 116(1):231–254, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2749-z.

Jung:2015:SRB

[JKSK15] Youngim Jung, Jayhoon Kim, Minho So, and Hwanmin Kim. Statistical relationships between journal use and research output at academic institutions in South Korea. Scientometrics, 103(3):751–777, June 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-015-1563-0.pdf.

Jensen:2014:MDL

[JL14] Pablo Jensen and Katsiaryna Lutkouskaya. The many dimensions of laboratories’ interdisciplinarity. Scientometrics, 98(1):619–631, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
013-1129-y.

Javed:2018:PRO

[JL18a] Saad Ahmed Javed and Sifeng Liu. Predicting the research output/growth of selected countries: application of Even GM(1, 1) and NDGM models. Scientometrics, 115(1):395–413, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-

017-2586-5. See correction [JL19].

Jiang:2018:HSI

[JL18b] Fan Jiang and Niancai Liu. The hierarchical status of international academic awards in social sciences. Scientometrics, 117(3):2091–2115, December 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2928-y.

Javed:2019:CPR

[JL19] Saad Ahmed Javed and Sifeng Liu. Correction to: Predicting the research output/growth of selected countries: application of Even GM (1, 1) and NDGM models. Scientometrics, 120(3):1505, September 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-03003-w;   http://link.springer.com/content/
pdf/10.1007/s11192-018-03003-w.pdf. See [JL18a].

Jokic:2018:SPE

[JMM18] Maja Jokić, Andrea Mervar, and Stjepan Mateljan. Scientific potential of European fully open access journals. Scientometrics, 114(3):1373–1394, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2629-y.

Jokic:2019:CAB

[JMM19] Maja Jokić, Andrea Mervar, and Stjepan Mateljan. Comparative analysis of book citations in social science journals by Central and Eastern European authors. Scientometrics, 120(3):1005–1029, September 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03176-y.

Jamali:2011:ATT

[JN11] Hamid R. Jamali and Mahsa Nikzad. Article title type and its relation with the number of downloads and citations. Scientometrics, 88(2):653–661, August 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0412-z.

Jamali:2015:OAS

[JN15] Hamid R. Jamali and Majid Nabavi. Open access and sources of full-text articles in Google Scholar in different subject fields. Scientometrics, 105(3):1635–1651, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1642-2.

Jamali:2018:HVA

[JNA18] Hamid R. Jamali, Majid Nabavi, and Saeid Asadi. How video articles are cited, the case of JoVE: Journal of Visualized Experiments. Scientometrics, 117(3):1821–1839, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2957-6.

Jirschitzka:2017:IRR

[JOGC17] Jens Jirschitzka, Aileen Oeberst, Richard Göllner, and Ulrike Cress. Inter-rater reliability and validity of peer reviews in an interdisciplinary field. Scientometrics, 113(2): 1059–1092, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2516-6.

Johnes:2018:URW

[Joh18] Jill Johnes. University rankings: What do they really show? Scientometrics, 115(1):585–606, April 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2666-1.

Jonkers:2010:MOC

[Jon10] Koen Jonkers. Models and orphans; concentration of the plant molecular life science research agenda. Scientometrics, 83(1):167–179, April 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0024-z.

Johnston:2013:CST

[JPT13] David W. Johnston, Marco Piatti, and Benno Torgler. Citation success over time: theory or empirics? Scientometrics, 95(3):1023–1029, June 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0910-7.

Juznic:2010:SIP

[JPZ^{+}10] Primoz Juznic, Stojan Peclin, Matjaz Zaucer, Tilen Mandelj, Miro Pusnik, and Franci Demsar. Scientometric indicators: peer-review, bibliometric methods and conflict of interests. Scientometrics, 85(2):429–441, November 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0230-8.

Ji:2014:BAR

[JPZ14] Qing Ji, Xiaoping Pang, and Xi Zhao. A bibliometric analysis of research on Antarctica during 1993–2012. Scientometrics, 101(3):1925–1939, December 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1332-5.

Ju:2015:IPR

[JS15] Yonghan Ju and So Young Sohn. Identifying patterns in rare earth element patents based on text and data mining. Scientometrics, 102(1):389–410, January 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1382-8.

Jiang:2013:GBA

[JSZ13] Xiaorui Jiang, Xiaoping Sun, and Hai Zhuge. Graphbased algorithms for ranking researchers: not all swans are white! Scientometrics, 96(3):743–759, September 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0943-y.

Jimenez-Saez:2013:WLR

[JSZIZ13] Fernando Jiménez-Sáez, Jon Mikel Zabala-Iturriagagoitia, and Jose Luis Zofío. Who leads research productivity growth? Guidelines for R&D policy-makers. Scientometrics, 94(1):273–303, January 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0763-0.

Jia:2014:RMS

[JTZ14] Xiaoyan Jia, Xufei Tan, and Yuehong Zhang. Replication of the methods section in biosciences papers: is it plagiarism? Scientometrics, 98(1):337–345, January 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1033-5.

Jun:2012:ESU

[Jun12] Seung-Pyo Jun. An empirical study of users’ hype cycle based on search traffic: the case study on hybrid cars. Scientometrics, 91(1):81–99, April 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0550-3.

Jansen:2010:KPS

[JvGH10] Dorothea Jansen, Regina von Görtz, and Richard Heidler. Knowledge production and the structure of collaboration networks in two scientific fields. Scientometrics, 83 (1):219–241, April 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0022-1.

Janko:2017:CCC

[JVM17] Ferenc Jankó, Judit Papp Vancsó, and Norbert Móricz. Is climate change controversy good for science? IPCC and contrarian reports in the light of bibliometrics. Scientometrics, 112(3):1745–1759, September 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2440-9.

Joob:2018:PRC

[JW18] Beuy Joob and Viroj Wiwanitkit. Post retraction citations in context: a comment. Scientometrics, 115(3):1291–1292, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2713-y.

Jian:2013:PAO

[JX13] Du Jian and Tang Xiaoli. Perceptions of author order versus contribution among researchers with different professional ranks and the potential of harmonic counts for encouraging ethical co-authorship practices. Scientometrics, 96(1):277– 295, July 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0905-4.

Jiang:2016:SVA

[JYM^{+}16] Hongbing Jiang, Chen Yang, Jian Ma, Thushari Silva, and Huaping Chen. A social voting approach for scientific domain vocabularies construction. Scientometrics, 108(2): 803–820, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1990-6.

Jang:2011:EFE

[JYW11] Show-Ling Jang, Yun-Chen Yu, and Tzu-Ya Wang. Emerging firms in an emerging field: an analysis of patent citations in electronic-paper display technology. Scientometrics, 89 (1):259–272, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0448-0.

Zhang:2015:GGS

[jZhLY15] Shao jie Zhang, Peng hui Lyu, and Yan Yan. Global geographical and scientometric analysis of tourism-themed research. Scientometrics, 105(1):385–401, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1678-3.

Jokic:2010:CSP

[JZL10] Maja Jokić, Kresimir Zauder, and Srebrenka Letina. Croatian scholarly productivity 1991–2005 measured by journals indexed in Web of Science. Scientometrics, 83(2):375–395, May 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0071-5.

Kato:2013:RBR

[KA13] Maki Kato and Asao Ando. The relationship between research performance and international collaboration in chemistry. Scientometrics, 97(3):535–553, December 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1011-y.

Kato:2017:NTI

[KA17] Maki Kato and Asao Ando. National ties of international scientific collaboration and researcher mobility found in Nature and Science. Scientometrics, 110(2):673–694, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2183-z.

Karamourzov:2012:DTS

[Kar12] Renat Karamourzov. The development trends of science in the CIS countries on the basis of some scientometric indicators. Scientometrics, 91(1):1–14, April 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0592-6.

Kazakis:2014:BER

[Kaz14] Nikolaos A. Kazakis. Bibliometric evaluation of the research performance of the Greek civil engineering departments in national and European context. Scientometrics, 101(1): 505–525, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1326-3.

Kazakis:2015:RAC

[Kaz15] Nikolaos A. Kazakis. The research activity of the current faculty of the Greek chemical engineering departments: a bibliometric study in national and international context. Scientometrics, 103(1):229–250, April 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1523-0.

Klavans:2010:TOR

[KB10] Richard Klavans and Kevin W. Boyack. Toward an objective, reliable and accurate method for measuring research leadership. Scientometrics, 82(3):539–553, March 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0188-6.

Kenna:2011:CMD

[KB11a] R. Kenna and B. Berche. Critical mass and the dependency of research quality on group size. Scientometrics, 86(2): 527–540, February 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0282-9.

Kissin:2011:TJS

[KB11b] Igor Kissin and Edwin L. Bradley, Jr. Top journals selectivity index: is it acceptable for drugs beyond the field of analgesia? Scientometrics, 88(2):589–597, August 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0403-0.

Kliegl:2011:ICP

[KB11c] Reinhold Kliegl and Douglas Bates. International collaboration in psychology is on the rise. Scientometrics, 87 (1):149–158, April 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0299-0.

Kissin:2012:TJS

[KB12] Igor Kissin and Edwin L. Bradley, Jr. Top journals selectivity index and “me-too” drugs. Scientometrics, 91 (1):131–142, April 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0556-x.

Kissin:2013:SBP

[KB13] Igor Kissin and Edwin L. Bradley, Jr. A surname-based patent-related indicator: the contribution of Jewish inventors to US patents. Scientometrics, 97(2):357–368, November 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1005-9.

Klarenbeek:2018:MMH

[KB18] Tracy Klarenbeek and Nelius Boshoff. Measuring multidisciplinary health research at South African universities: a comparative analysis based on co-authorships and journal subject categories. Scientometrics, 116(3):1461–1485, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2813-8.

Kallmes:2017:DFM

[KBAK17] Kevin M. Kallmes, Waleed Brinjikji, Ahmed T. Ahmed, and David F. Kallmes. Difficulty in finding manuscript reviewers is not associated with manuscript acceptance rates: a study of the peer-review process at the journal Radiology. Scientometrics, 111(2):971–978, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Kozak:2015:HEE

[KBL15] Marcin Kozak, Lutz Bornmann, and Loet Leydesdorff. How have the Eastern European countries of the former Warsaw Pact developed since 1990? A bibliometric study. Scientometrics, 102(2):1101–1117, February 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-014-1439-8.pdf.

Kenekayoro:2014:ACA

[KBT14] Patrick Kenekayoro, Kevan Buckley, and Mike Thelwall. Automatic classification of academic web page types. Scientometrics, 101(2):1015–1026, November 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1292-9.

Kenekayoro:2015:CRG

[KBT15] Patrick Kenekayoro, Kevan Buckley, and Mike Thelwall. Clustering research group website homepages. Scientometrics, 102(3):2023–2039, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1497-y.

Kutlaca:2015:AQQ

[KBZS15] Djuro Kutlaca, Dragan Babić, Lazar Zivković, and Dijana Strbac. Analysis of quantitative and qualitative indicators of SEE countries scientific output. Scientometrics, 102(1): 247–265, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1290-y.

Ke:2012:SPS

[KC12] Hao-Ren Ke and Ya-Ning Chen. Structure and pattern of social tags for keyword selection behaviors. Scientometrics, 92(1):43–62, July 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0718-5.

Kim:2015:SRE

[KC15] Meen Chul Kim and Chaomei Chen. A scientometric review of emerging trends and new developments in recommendation systems. Scientometrics, 104(1):239–263, July 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1595-5.

Kim:2014:DPT

[KCK14] Euiseok Kim, Yongrae Cho, and Wonjoon Kim. Dynamic patterns of technological convergence in printed electronics technologies: patent citation network. Scientometrics, 98 (2):975–998, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1104-7.

Kozma:2019:RSA

[KCM19] Csaba Kozma and Clara Calero-Medina. The role of South African researchers in intercontinental collaboration. Scientometrics, 121(3):1293–1321, December 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03230-9;   http://link.springer.
com/content/pdf/10.1007/s11192-019-03230-9.pdf.

Ko:2011:SOK

[KCP11] Young Man Ko, Soo-Ryun Cho, and Yong Seok Park. A study on the optimization of KCI-based index (kor-factor) in evaluating Korean journals. Scientometrics, 88(1):61–71, July 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0384-z.

Khan:2012:CDE

[KCP12] Gohar Feroz Khan, Seong Eun Cho, and Han Woo Park. A comparison of the Daegu and Edinburgh musical industries: a triple helix approach. Scientometrics, 90(1):85–99, January 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0504-9.

Kudlow:2017:ODC

[KCT^{+}17] Paul Kudlow, Matthew Cockerill, Danielle Toccalino, Devin Bissky Dziadyk, Alan Rutledge, Aviv Shachak, Roger S. McIntyre, Arun Ravindran, and Gunther Eysenbach. Online distribution channel increases article usage on Mendeley: a randomized controlled trial. Scientometrics, 112(3):1537–1556, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
017-2438-3;   http://link.springer.com/content/pdf/
10.1007/s11192-017-2438-3.pdf.

Khan:2019:FRF

[KCU19] Arif Khan, Nazim Choudhury, and Shahadat Uddin. Few research fields play major role in interdisciplinary grant success. Scientometrics, 119(1):237–246, April 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03043-w.

Kim:2014:NBA

[KD14] Jinseok Kim and Jana Diesner. A network-based approach to coauthorship credit allocation. Scientometrics, 101(1): 587–602, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1253-3.

Kallens:2018:EMT

[KD18] Pablo Contreras Kallens and Rick Dale. Exploratory mapping of theoretical landscapes through word use in abstracts. Scientometrics, 116(3):1641–1674, September 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2811-x.

Kim:2019:FBL

[KD19] Jinseok Kim and Jana Diesner. Formational bounds of link prediction in collaboration networks. Scientometrics, 119 (2):687–706, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03055-6.

Kretschmer:2015:TDV

[KdBBK15] Hildrun Kretschmer, Donald de B. Beaver, and Theo Kretschmer. Three-dimensional visualization and animation of emerging patterns by the process of self-organization in collaboration networks. Scientometrics, 104(1):87–120, July 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1579-5.

Kazakis:2014:ERP

[KDFL14] Nikolaos A. Kazakis, Anastasios D. Diamantidis, Leonidas L. Fragidis, and Miltos K. Lazarides. Evaluating the research performance of the Greek medical schools using bibliometrics. Scientometrics, 98(2):1367–1384, February 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1049-x.

Ke:2013:FMS

[Ke13] Weimao Ke. A fitness model for scholarly impact analysis. Scientometrics, 94(3):981–998, March 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0787-5.

Kegen:2015:CSA

[Keg15] Nadine V. Kegen. Cohesive subgroups in academic networks: unveiling clique integration of top-level female and male researchers. Scientometrics, 103(3):897–922, June 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1572-z.

Kenekayoro:2018:INE

[Ken18] Patrick Kenekayoro. Identifying named entities in academic biographies with supervised learning. Scientometrics, 116 (2):751–765, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2797-4.

Kulczycki:2018:PPS

[KEP^{+}18] Emanuel Kulczycki, Tim C. E. Engels, Janne Pölönen, Kasper Bruun, Marta Dusková, Raf Guns, Robert Nowotniak, Michal Petr, Gunnar Sivertsen, Andreja Istenic Starcic, and Alesia Zuccala. Publication patterns in the social sciences and humanities: evidence from eight European countries. Scientometrics, 116(1):463–486, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2711-0;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2711-
0.pdf.

Kolesnikov:2018:RRS

[KFB18] Sergey Kolesnikov, Eriko Fukumoto, and Barry Bozeman. Researchers’ risk-smoothing publication strategies: Is productivity the enemy of impact? Scientometrics, 116(3):1995–2017, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2793-8.

Konig:2015:TGD

[KFKS15] Cornelius J. König, Clemens B. Fell, Linus Kellnhofer, and Gabriel Schui. Are there gender differences among researchers from industrial/organizational psychology? Scientometrics, 105(3):1931–1952, December 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1646-y.

Kaur:2010:MDS

[KG10a] Har Kaur and B. M. Gupta. Mapping of dental science research in India: a scientometric analysis of India’s research output, 1999–2008. Scientometrics, 85(1):361–376, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0213-9.

Klitkou:2010:RBA

[KG10b] Antje Klitkou and Magnus Gulbrandsen. The relationship between academic patenting and scientific publishing in Norway. Scientometrics, 82(1):93–108, January 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0050-x.

Kretschmer:2013:ICW

[KG13] Hildrun Kretschmer and Wolfgang Glänzel. The 8th International Conference on Webometrics, Informetrics and Scientometrics & 13th COLLNET Meeting. Scientometrics, 97(1):1, October 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-013-
1078-5.pdf.

Kufenko:2016:BCE

[KG16] Vadim Kufenko and Niels Geiger. Business cycles in the economy and in economics: an econometric analysis. Scientometrics, 107(1):43–69, April 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1866-9.

Klingelhofer:2018:FYA

[KGB^{+}18] Doris Klingelhöfer, David A. Groneberg, Markus Braun, Dörthe Brüggmann, and Jenny Jaque. Fifteen years after September 11: Where is the medical research heading? A scientometric analysis. Scientometrics, 117(1):45–60, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2878-4.

Kshitij:2015:EIS

[KGG15] Avinash Kshitij, Jaideep Ghosh, and Brij Mohan Gupta. Embedded information structures and functions of coauthorship networks: evidence from cancer research collaboration in India. Scientometrics, 102(1):285–306, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1310-y.

Kim:2014:ICC

[KGL^{+}14] Byunghoon Kim, Gianluca Gazzola, Jae-Min Lee, Dohyun Kim, Kanghoe Kim, and Myong K. Jeong. Inter-cluster connectivity analysis for technology opportunity discovery. Scientometrics, 98(3):1811–1825, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1097-2.

Karpagam:2011:MNN

[KGNB11] R. Karpagam, S. Gopalakrishnan, M. Natarajan, and B. Ramesh Babu. Mapping of nanoscience and nanotechnology research in India: a scientometric analysis, 1990– 2009. Scientometrics, 89(2):501–522, November 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0477-8.

Karaulova:2016:SSP

[KGSS16] Maria Karaulova, Abdullah Gök, Oliver Shackleton, and Philip Shapira. Science system path-dependencies and their influences: nanotechnology research in Russia. Scientometrics, 107(2):645–670, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
016-1916-3;   http://link.springer.com/content/pdf/
10.1007/s11192-016-1916-3.pdf.

Kim:2017:TPE

[KGY^{+}17] Byunghoon Kim, Gianluca Gazzola, Jaekyung Yang, Jae-Min Lee, Byoung-Youl Coh, Myong K. Jeong, and Young-Seon Jeong. Two-phase edge outlier detection method for technology opportunity discovery. Scientometrics, 113 (1):1–16, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2472-1.

Kocak:2019:MCA

[KGZML^{+}19] Murat Kocak, Carlos García-Zorita, Sergio Marugán-Lázaro, Murat Perit Çakır, and Elías Sanz-Casado. Mapping and clustering analysis on neuroscience literature in Turkey: a bibliometric analysis from 2000 to 2017. Scientometrics, 121(3):1339–1366, December 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03259-w.

Kniffin:2017:ANT

[KH17] Kevin M. Kniffin and Andrew S. Hanks. Antecedents and near-term consequences for interdisciplinary dissertators. Scientometrics, 111(3):1225–1250, June 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-017-2317-y.pdf.

Khan:2013:ESM

[Kha13a] Gohar Feroz Khan. Erratum to: Social media-based systems: an emerging area of information systems research and practice. Scientometrics, 95(1):181–182, April 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-012-0901-8.pdf. See [Kha13b].

Khan:2013:SMB

[Kha13b] Gohar Feroz Khan. Social media-based systems: an emerging area of information systems research and practice. Scientometrics, 95(1):159–180, April 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0831-5. See erratum [Kha13a].

Kivinen:2017:SPG

[KHA17] Osmo Kivinen, Juha Hedman, and Kalle Artukka. Scientific publishing and global university rankings. How well are top publishing universities recognized? Scientometrics, 112(1): 679–695, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Kim:2018:CRA

[KHH18] Sung Kim, Derek Hansen, and Richard Helps. Computing research in the academy: insights from theses and dissertations. Scientometrics, 114(1):135–158, January 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2572-y.

Kim:2012:THA

[KHJ^{+}12] Haneul Kim, Minghao Huang, Furong Jin, David Bodoff, Junghoon Moon, and Young Chan Choe. Triple helix in the agricultural sector of northeast Asian countries: a comparative study between Korea and China. Scientometrics, 90 (1):101–120, January 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0517-4.

Kivinen:2013:PAR

[KHK13] Osmo Kivinen, Juha Hedman, and Päivi Kaipainen. Productivity analysis of research in natural sciences, technology and clinical medicine: an input-output model applied in comparison of top 300 ranked universities of 4 North European and 4 East Asian countries. Scientometrics, 94(2): 683–699, February 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0808-4.

Kibanov:2019:SSS

[KHR^{+}19] Mark Kibanov, Raphael H. Heiberger, Simone Rödder, Martin Atzmueller, and Gerd Stumme. Social studies of scholarly life with sensor-based ethnographic observations. Scientometrics, 119(3):1387–1428, June 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03097-w.

Karlsson:2015:MUB

[KHS^{+}15] Alexander Karlsson, Björn Hammarfelt, H. Joe Steinhauer, Göran Falkman, Nasrine Olson, Gustaf Nelhans, and Jan Nolin. Modeling uncertainty in bibliometrics and information retrieval: an information fusion approach. Scientometrics, 102(3):2255–2274, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1481-6.

Keathley-Herring:2016:AMR

[KHVGA^{+}16] Heather Keathley-Herring, Eileen Van Aken, Fernando Gonzalez-Aleu, Fernando Deschamps, Geert Letens, and Pablo Cardenas Orlandini. Assessing the maturity of a research area: bibliometric review and proposed framework. Scientometrics, 109(2):927–951, November 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2096-x.

Kim:2010:VKS

[Kim10] Mee-Jean Kim. Visibility of Korean science journals: an analysis between citation measures among international composition of editorial board and foreign authorship. Scientometrics, 84(2):505–522, August 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0168-x.

Kim:2014:BAP

[Kim14] Mee-Jean Kim. A bibliometric analysis of publications by the School of Biological Sciences, Seoul National University, South Korea. Scientometrics, 98(2):999–1019, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1084-7.

Kim:2018:EAN

[Kim18] Jinseok Kim. Evaluating author name disambiguation for digital libraries: a case of DBLP. Scientometrics, 116(3):1867–1886, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-

018-2824-5. See correction [Kim19a].

Kim:2019:CEA

[Kim19a] Jinseok Kim. Correction to: Evaluating author name disambiguation for digital libraries: a case of DBLP. Scientometrics, 118(1):383, January 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2960-y;   http://link.springer.com/content/pdf/
10.1007/s11192-018-2960-y.pdf. See [Kim18].

Kim:2019:FIA

[Kim19b] Jinseok Kim. A fast and integrative algorithm for clustering performance evaluation in author name disambiguation. Scientometrics, 120(2):661–681, August 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03143-7.

Kissin:2011:CBI

[Kis11a] Igor Kissin. Can a bibliometric indicator predict the success of an analgesic? Scientometrics, 86(3):785–795, March 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0320-7.

Kissin:2011:SBB

[Kis11b] Igor Kissin. A surname-based bibliometric indicator: publications in biomedical journal. Scientometrics, 89(1):273– 280, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0437-3.

Kumar:2013:MRC

[KJ13] Sameer Kumar and Jariah Mohd. Jan. Mapping research collaborations in the business and management field in Malaysia, 1980–2010. Scientometrics, 97(3):491–517, December 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-0994-8.

Kumar:2014:RCN

[KJ14] Sameer Kumar and Jariah Mohd. Jan. Research collaboration networks of two OIC nations: comparative study between Turkey and Malaysia in the field of ‘energy fuels’, 2009–2011. Scientometrics, 98(1):387–414, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-1059-8.

Klimek:2016:SFG

[KJES16] Peter Klimek, Aleksandar S. Jovanovic, Rainer Egloff, and Reto Schneider. Successful fish go with the flow: citation impact prediction based on centrality measures for termdocument networks. Scientometrics, 107(3):1265–1282, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1926-1.

Kim:2014:IIL

[KJS14] Meen Chul Kim, Yoo Kyung Jeong, and Min Song. Investigating the integrated landscape of the intellectual topology of bioinformatics. Scientometrics, 101(1):309–335, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1417-1.

Kong:2017:EDR

[KJW^{+}17] Xiangjie Kong, Huizhen Jiang, Wei Wang, Teshome Megersa Bekele, Zhenzhen Xu, and Meng Wang. Exploring dynamic research interest and academic influence for scientific collaborator recommendation. Scientometrics, 113(1):369–385, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2485-9.

Kretschmer:2013:GBE

[KK13] Hildrun Kretschmer and Theo Kretschmer. Gender bias and explanation models for the phenomenon of women’s discriminations in research careers. Scientometrics, 97(1):25–36, October 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1023-7.

kumar:2015:SAM

[KK15] R. Santha Kumar and K. Kaliyaperumal. A scientometric analysis of mobile technology publications. Scientometrics, 105(2):921–939, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1710-7.

Kreuchauff:2017:PSS

[KK17] Florian Kreuchauff and Vladimir Korzinov. A patent search strategy based on machine learning for the emerging field of service robotics. Scientometrics, 111(2):743–772, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Kim:2018:IIT

[KK18] Jinseok Kim and Jenna Kim. The impact of imbalanced training data on machine learning for author name disambiguation. Scientometrics, 117(1):511–526, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2865-9.

Korytkowski:2019:EHC

[KK19] Przemysław Korytkowski and Emanuel Kulczycki. Examining how country-level science policy shapes publication patterns: the case of Poland. Scientometrics, 119(3):1519–1543, June 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03092-1;   http://
link.springer.com/content/pdf/10.1007/s11192-019-
03092-1.pdf.

Kirchner:2017:MFA

[KKBW17] Marlene K. Kirchner, Lubor Kostál, Boris Bilcík, and Christoph Winckler. Mapping farm animal welfare research in an enlarged Europe: international collaboration, bibliometric output, research resources and relation to economic indices. Scientometrics, 113(2):909–922, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2505-9.

Karsli:2018:CDS

[KKCG18] Meva Bayrak Karsli, Sinem Karabey, Nergiz Ercil Cagiltay, and Yuksel Goktas. Comparison of the discussion sections of PhD dissertations in educational technology: the case of Turkey and the USA. Scientometrics, 117(3):1381–1403, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2955-8.

Kretschmer:2012:GBJ

[KKdBK12] Hildrun Kretschmer, Ramesh Kundra, Donald deB. Beaver, and Theo Kretschmer. Gender bias in journals of gender studies. Scientometrics, 93(1):135–150, October 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0661-5.

Kutlar:2013:CTA

[KKE13] Aziz Kutlar, Ali Kabasakal, and Mehmet Sena Ekici. Contributions of Turkish academicians supervising PhD dissertations and their universities to economics: an evaluation of the 1990–2011 period. Scientometrics, 97(3):639–658, December 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-0973-0.

Ko:2014:ATI

[KKK^{+}14] Sung-Seok Ko, Namuk Ko, Doyeon Kim, Hyunseok Park, and Janghyeok Yoon. Analyzing technology impact networks for R&D planning using patents: combined application of network approaches. Scientometrics, 101(1):917–936, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1343-2.

Koczkodaj:2014:QES

[KKL14] Waldemar W. Koczkodaj, Konrad Kulakowski, and Antoni Ligeza. On the quality evaluation of scientific entities in Poland supported by consistency-driven pairwise comparisons method. Scientometrics, 99(3):911–926, June 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s11192-014-1258-y.pdf.

Kastrin:2017:ASR

[KKLP17] Andrej Kastrin, Jelena Klisara, Borut Luzar, and Janez Povh. Analysis of Slovenian research community through bibliographic networks. Scientometrics, 110(2):791–813, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2203-z.

Kastrin:2018:SDP

[KKLP18] Andrej Kastrin, Jelena Klisara, Borut Luzar, and Janez Povh. Is science driven by principal investigators? Scientometrics, 117(2):1157–1182, November 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2900-x.

Kim:2019:GAL

[KKOS19] Jinseok Kim, Jinmo Kim, and Jason Owen-Smith. Generating automatically labeled data for author name disambiguation: an iterative clustering method. Scientometrics, 118 (1):253–280, January 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2968-3.

Kawaguchi:2016:RCT

[KKS16] Daiji Kawaguchi, Ayako Kondo, and Keiji Saito. Researchers’ career transitions over the life cycle. Scientometrics, 109(3):1435–1454, December 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2131-y.

Koczkodaj:2017:HRN

[KKS^{+}17] W. W. Koczkodaj, T. Kakiashvili, A. Szymańska, J. Montero-Marin, R. Araya, J. Garcia-Campayo, K. Rutkowski, and D. Strzalka. How to reduce the number of rating scale items without predictability loss? Scientometrics, 111(2):581– 593, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-017-2283-4.pdf.

Kyriakidou:2018:BRP

[KKT^{+}18] Margarita Kyriakidou, Aigli Kyriakoudi, Nikolaos A. Triarides, Konstantinos Z. Vardakas, and Matthew E. Falagas. Biomedical research productivity and economic crisis in Greece: a 22-year study. Scientometrics, 116(3):1559– 1564, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2827-2.

Kungas:2013:REC

[KKV^{+}13] Peep Küngas, Siim Karus, Svitlana Vakulenko, Marlon Dumas, Cristhian Parra, and Fabio Casati. Reverseengineering conference rankings: what does it take to make a reputable conference? Scientometrics, 96(2):651–665, August 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0938-8.

Kousha:2016:MTW

[KL16] Kayvan Kousha and Jonathan Levitt. Michael Thelwall wins the 2015 Derek John de Solla Price Medal. Scientometrics, 108(2):485–488, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-015-1647-x.pdf.

Kim:2017:FIM

[KL17] Jeeeun Kim and Sungjoo Lee. Forecasting and identifying multi-technology convergence based on patent data: the case of IT and BT industries in 2020. Scientometrics, 111(1):47–65, April 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-017-2275-4.

Kim:2014:ECN

[KLCS14] Hee Dae Kim, Duk Hee Lee, Hochull Choe, and Il Won Seo. The evolution of cluster network structure and firm growth: a study of industrial software clusters. Scientometrics, 99 (1):77–95, April 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1094-5.

Klincewicz:2016:EDT

[Kli16] Krzysztof Klincewicz. The emergent dynamics of a technological research topic: the case of graphene. Scientometrics, 106(1):319–345, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1780-6.

Kim:2014:ARC

[KLL14] Yangson Kim, Hee Jin Lim, and Soo Jeung Lee. Applying research collaboration as a new way of measuring research performance in Korean universities. Scientometrics, 99(1):97–115, April 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1095-4.

Karlovcec:2016:CPD

[KLM16] Mario Karlovcec, Borut Luzar, and Dunja Mladenić. Coreperiphery dynamics in collaboration networks: the case study of Slovenia. Scientometrics, 109(3):1561–1578, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2154-4. See

[KM15a].

Kao:2012:AIM

[KLP12] Chiang Kao, Shiang-Tai Liu, and Hwei-Lan Pao. Assessing improvement in management research in Taiwan. Scientometrics, 92(1):75–87, July 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0721-x.

Karunan:2017:DII

[KLP17] Kavitha Karunan, Hiran H. Lathabai, and Thara Prabhakaran. Discovering interdisciplinary interactions between two research fields using citation networks. Scientometrics, 113(1):335–367, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2481-0.

Khan:2016:TCS

[KLPP16] Gohar Feroz Khan, Sungjoon Lee, Ji Young Park, and Han Woo Park. Theories in communication science: a structural analysis using webometrics and social network approach. Scientometrics, 108(2):531–557, August 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1822-0.

Kostoff:2011:SII

[KM11] Ronald N. Kostoff and Stephen A. Morse. Structure and infrastructure of infectious agent research literature: SARS. Scientometrics, 86(1):195–209, January 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0240-6.

Kwon:2012:SSM

[KM12] Ki-Seok Kwon and Ben R. Martin. Synergy or separation mode: the relationship between the academic research and the knowledge-transfer activities of Korean academics. Scientometrics, 90(1):177–200, January 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0513-8.

Karlovcec:2015:ISF

[KM15a] Mario Karlovcec and Dunja Mladenić. Interdisciplinarity of scientific fields and its evolution based on graph of project collaboration and co-authoring. Scientometrics, 102(1):433–454, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-

014-1355-y. See comments [Rod16].

Kaur:2015:CMR

[KM15b] Har Kaur and Preeti Mahajan. Collaboration in medical research: a case study of India. Scientometrics, 105(1): 683–690, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1691-6.

Kaur:2015:NVI

[KM15c] Har Kaur and Preeti Mahajan. National versus international growth of medical research of India: a case study. Scientometrics, 105(2):973–989, November 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1746-8.

Kaur:2015:RMI

[KM15d] Har Kaur and Preeti Mahajan. Ranking of medical institutes of India for quality and quantity: a case study. Scientometrics, 105(2):1129–1139, November 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1720-5.

Kumar:2016:BCH

[KM16] Sameer Kumar and Bernd Markscheffel. Bonded-communities in HantaVirus research: a research collaboration network (RCN) analysis. Scientometrics, 109(1):533–550, October 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-1942-1.

Katchanov:2017:SPJ

[KM17a] Yurij L. Katchanov and Yulia V. Markova. The “space of physics journals”: topological structure and the Journal Impact Factor. Scientometrics, 113(1):313–333, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2471-2.

Khachatryan:2017:DSP

[KM17b] Davit Khachatryan and Brigitte Muehlmann. Determinants of successful patent applications to combat financial fraud. Scientometrics, 111(3):1353–1383, June 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Kacem:2018:ASS

[KM18a] Ameni Kacem and Philipp Mayr. Analysis of search stratagem utilisation. Scientometrics, 116(2):1383–1400, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2821-8.

Kenworthy:2018:CET

[KM18b] Thomas P. Kenworthy and W. Edward McMullan. In consideration of entrepreneurship theory. Scientometrics, 115 (2):767–783, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2699-5.

Kumar:2018:CII

[KM18c] K. Reji Kumar and Shibu Manuel. Collaborations of Indian institutions which conduct mathematical research: A study from the perspective of social network analysis. Scientometrics, 117(2):1041–1051, November 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2898-0.

Karimi:2018:CBR

[KMD^{+}18] Samaneh Karimi, Luis Moraes, Avisha Das, Azadeh Shakery, and Rakesh Verma. Citance-based retrieval and summarization using IR and machine learning. Scientometrics, 116(2):1331–1366, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2785-8.

Kronegger:2012:CSS

[KMFD12] Luka Kronegger, Franc Mali, Anuska Ferligoj, and Patrick Doreian. Collaboration structures in Slovenian scientific communities. Scientometrics, 90(2):631–647, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0493-8.

Khan:2011:NCM

[KMP11a] Gohar Feroz Khan, Junghoon Moon, and Han Woo Park. Network of the core: mapping and visualizing the core of scientific domains. Scientometrics, 89(3):759–779, December 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0489-4.

Khan:2011:STP

[KMP^{+}11b] Gohar Feroz Khan, Junghoon Moon, Han Woo Park, Bobby Swar, and Jae Jeung Rho. A socio-technical perspective on e-government issues in developing countries: a scientometrics approach. Scientometrics, 87(2):267–286, May 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0322-5.

Katchanov:2016:HPW

[KMS16] Yurij L. Katchanov, Yulia V. Markova, and Natalia A. Shmatko. How physics works: scientific capital in the space of physics institutions. Scientometrics, 108(2):875–893, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-2005-3.

Kullenberg:2015:HTM

[KN15] Christopher Kullenberg and Gustaf Nelhans. The happiness turn? Mapping the emergence of “happiness studies” using cited references. Scientometrics, 103(2):615–630, May 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1536-3.

Kaur:2019:KGW

[KNK^{+}19] Kiran Kaur, Kwan Hoong Ng, Ray Kemp, Yin Yee Ong, Zaharah Ramly, and Ai Peng Koh. Knowledge generation in the wake of the Fukushima Daiichi nuclear power plant disaster. Scientometrics, 119(1):149–169, April 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03024-z.

Kuld:2018:RMA

[KO18] Lukas Kuld and John O’Hagan. Rise of multi-authored papers in economics: Demise of the ‘lone star’ and why? Scientometrics, 114(3):1207–1225, March 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2588-3.

Katsurai:2019:TME

[KO19] Marie Katsurai and Shunsuke Ono. TrendNets: mapping emerging research trends from dynamic co-word networks via sparse representation. Scientometrics, 121 (3):1583–1598, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03241-6;   http://link.springer.com/content/
pdf/10.1007/s11192-019-03241-6.pdf.

Kretschmer:2012:ICW

[KÖG12] Hildrun Kretschmer, Bülent Özel, and Wolfgang Glänzel. The 7th International Conference on Webometrics, Informetrics and Scientometrics & 12th COLLNET Meeting. Scientometrics, 93(1):1–2, October 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-012-0664-2.pdf.

Kolasa:2012:SCC

[Kol12] Wladysław Marek Kolasa. Specific character of citations in historiography (using the example of Polish history). Scientometrics, 90(3):905–923, March 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-011-0553-0.pdf.

Kondo:2012:PRI

[Kon12] Masayuki Kondo. A public research institute that created and led a large industrial group in Japan. Scientometrics, 90 (1):141–162, January 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0505-8.

Korom:2018:DSE

[Kor18] Philipp Korom. Does scientific eminence endure? making sense of the most cited economists, psychologists and sociologists in textbooks (1970–2010). Scientometrics, 116(2): 909–939, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2781-z.

Korom:2019:BVE

[Kor19] Philipp Korom. A bibliometric visualization of the economics and sociology of wealth inequality: a world apart? Scientometrics, 118(3):849–868, March 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-03000-z;   http://link.springer.
com/content/pdf/10.1007/s11192-018-03000-z.pdf.

Kostoff:2014:LRD

[Kos14] Ronald N. Kostoff. Literature-related discovery: common factors for Parkinson’s disease and Crohn’s disease. Scientometrics, 100(3):623–657, September 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1298-3.

Kosmulski:2015:CYP

[Kos15] Marek Kosmulski. Careers of young Polish chemists. Scientometrics, 102(2):1455–1465, February 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-014-1461-x.pdf.

Koseoglu:2016:MIC

[Kos16a] Mehmet Ali Koseoglu. Mapping the institutional collaboration network of strategic management research: 1980– 2014. Scientometrics, 109(1):203–226, October 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-1894-5.

Kosten:2016:CUR

[Kos16b] Joost Kosten. A classification of the use of research indicators. Scientometrics, 108(1):457–464, July 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-1904-7.pdf.

Kosmulski:2018:YTO

[Kos18a] Marek Kosmulski. Are you in top 1% (1o/oo)? Scientometrics, 114(2):557–565, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2526-4;   https://link.springer.
com/content/pdf/10.1007/s11192-017-2526-4.pdf.

Kosteas:2018:PLR

[Kos18b] Vasilios D. Kosteas. Predicting long-run citation counts for articles in top economics journals. Scientometrics, 115(3): 1395–1412, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2703-0.

Kozlowski:2015:IIN

[Koz15] Jan Kozlowski. Innovation indices: the need for positioning them where they properly belong. Scientometrics, 104 (3):609–628, September 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-015-
1632-4.pdf.

Khan:2012:ETH

[KP12a] Gohar Feroz Khan and Han Woo Park. Editorial: Triple helix and innovation in Asia using scientometrics, webometrics, and informetrics. Scientometrics, 90(1):1–7, January 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0506-7.

Kim:2012:MTB

[KP12b] Minjeong Kim and Han Woo Park. Measuring Twitterbased political participation and deliberation in the South Korean context by using social network and triple helix indicators. Scientometrics, 90(1):121–140, January 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0508-5.

Kazi:2016:TNP

[KPJ16] ParvezAhamad Kazi, Manasi Patwardhan, and Pushkar Joglekar. Towards a new perspective on context based citation index of research articles. Scientometrics, 107(1): 103–121, April 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1844-2.

Koler-Povh:2014:IOA

[KPJT14] Teja Koler-Povh, Primoz Juznic, and Goran Turk. Impact of open access on citation of scholarly publications in the field of civil engineering. Scientometrics, 98(2):1033–1045, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1101-x.

Kim:2019:ITO

[KPL19] Kyuwoong Kim, Kyeongmin Park, and Sungjoo Lee. Investigating technology opportunities: the use of SAOx analysis. Scientometrics, 118(1):45–70, January 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2962-9.

Kovanis:2016:CSA

[KPRT16] Michail Kovanis, Raphaël Porcher, Philippe Ravaud, and Ludovic Trinquart. Complex systems approach to scientific publication and peer-review system: development of an agent-based model calibrated with empirical journal data. Scientometrics, 106(2):695–715, February 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1800-6;   http://link.

springer.com/content/pdf/10.1007/s11192-015-1800-
6.pdf.

Kretschmer:2012:ERE

[KPS12] Hildrun Kretschmer, Alexander Pudovkin, and Johannes Stegmann. Erratum to: Research evaluation. Part II: gender effects of evaluation: are men more productive and more cited than women? Scientometrics, 93(1):31, October 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s11192-012-0739-0.pdf. See [HAJ12].

Kwon:2012:GSS

[KPSL12] Ki-Seok Kwon, Han Woo Park, Minho So, and Loet Leydesdorff. Has globalization strengthened South Korea’s national research system? National and international dynamics of the triple helix of scientific co-authorship relationships in South Korea. Scientometrics, 90(1):163–176, January 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0512-9.

Kwon:2016:NIT

[KPY16] Seokbeom Kwon, Alan Porter, and Jan Youtie. Navigating the innovation trajectories of technology by combining specialization score analyses for publications and patents: graphene and nano-enabled drug delivery. Scientometrics, 106(3):1057–1071, March 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1826-9.

Kulczycki:2017:DEB

[KR17a] Emanuel Kulczycki and Ewa A. Rozkosz. Does an expertbased evaluation allow us to go beyond the Impact Factor? Experiences from building a ranking of national journals in Poland. Scientometrics, 111(1):417–442, April 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-017-2261-x.pdf.

Kyvik:2017:RCG

[KR17b] Svein Kyvik and Ingvild Reymert. Research collaboration in groups and networks: differences across academic fields. Scientometrics, 113(2):951–967, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2497-5;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2497-
5.pdf.

Krampen:2010:ACB

[Kra10] Günter Krampen. Acceleration of citing behavior after the millennium? Exemplary bibliometric reference analyses for psychology journals. Scientometrics, 83(2):507–513, May 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0093-z.

Krauskopf:2013:SIA

[Kra13] E. Krauskopf. Standardization of the institutional address. Scientometrics, 94(3):1313–1315, March 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0852-0.

Krampen:2016:STA

[Kra16] Günter Krampen. Scientometric trend analyses of publications on the history of psychology: Is psychology becoming an unhistorical science? Scientometrics, 106(3):1217–1238, March 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com//content/pdf/10.1007/s11192-016-1834-4.pdf;
http://link.springer.com/content/pdf/10.1007/s11192-
016-1834-4.pdf.

Krawczyk:2017:ARM

[Kra17] Michal Krawczyk. Are all researchers male? Gender misattributions in citations. Scientometrics, 110(3):1397–1402, March 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-016-2192-y.pdf.

Krauskopf:2019:MDS

[Kra19] Erwin Krauskopf. Missing documents in Scopus: the case of the journal Enfermeria Nefrologica. Scientometrics, 119 (1):543–547, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03040-z.

Katz:2019:CSI

[KRP19] J. Sylvan Katz and Guillermo Armando Ronda-Pupo. Cooperation, scale-invariance and complex innovation systems: a generalization. Scientometrics, 121(2):1045–1065, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03215-8.

Kumar:2014:IRC

[KRR14] Sameer Kumar, Vala Ali Rohani, and Kuru Ratnavelu. International research collaborations of ASEAN nations in economics, 1979–2010. Scientometrics, 101(1):847–867, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1363-y.

Kotsemir:2017:MAV

[KS17] Maxim Kotsemir and Sergey Shashnov. Measuring, analysis and visualization of research capacity of university at the level of departments and staff members. Scientometrics, 112(3):1659–1689, September 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2450-7.

Kim:2018:MET

[KS18] Junmo Kim and Juneseuk Shin. Mapping extended technological trajectories: integration of main path, derivative paths, and technology junctures. Scientometrics, 116(3):1439–1459, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2834-3.

Kosecki:2011:SCU

[KSB11] Stanislaw Kosecki, Robbin Shoemaker, and Charlotte Kirk Baer. Scope, characteristics, and use of the U.S. Department of Agriculture’s intramural research. Scientometrics, 88(3): 707–728, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0359-0.

Kademani:2013:PTM

[KSSB13] B. S. Kademani, Anil Sagar, Ganesh Surwase, and K. Bhanumurthy. Publication trends in materials science: a global perspective. Scientometrics, 94(3):1275–1295, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0835-1.

Kawashima:2015:AES

[KT15] Hirotaka Kawashima and Hiroyuki Tomizawa. Accuracy evaluation of Scopus Author ID based on the largest funding database in Japan. Scientometrics, 103(3):1061–1071, June 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1580-z.

Kim:2016:ESS

[KTLD16] Jinseok Kim, Liang Tao, Seok-Hyoung Lee, and Jana Diesner. Evolution and structure of scientific co-publishing network in Korea between 1948–2011. Scientometrics, 107 (1):27–41, April 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1878-5.

Kovanis:2017:EAS

[KTRP17] Michail Kovanis, Ludovic Trinquart, Philippe Ravaud, and Raphaël Porcher. Evaluating alternative systems of peer review: a large-scale agent-based modelling approach to scientific publication. Scientometrics, 113(1):651–671, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2375-1;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2375-
1.pdf.

Kumar:2011:CSI

[KTT11] Ranjeet Kumar, R. C. Tripathi, and M. D. Tiwari. A case study of impact of patenting in the current developing economies in Asia. Scientometrics, 88(2):575–587, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0405-y.

Kovacs:2015:ESO

[KVC15] Adrián Kovács, Bart Van Looy, and Bruno Cassiman. Exploring the scope of open innovation: a bibliometric review of a decade of research. Scientometrics, 104(3):951– 983, September 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1628-0.

Krampen:2011:FTD

[KvES11] Günter Krampen, Alexander von Eye, and Gabriel Schui. Forecasting trends of development of psychology from a bibliometric perspective. Scientometrics, 87(3):687–694, June 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0357-2.

Khan:2015:ITM

[KW15] Gohar Feroz Khan and Jacob Wood. Information technology management domain: emerging themes and keyword analysis. Scientometrics, 105(2):959–972, November 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1712-5.

Koopman:2017:MIB

[KW17] Rob Koopman and Shenghui Wang. Mutual information based labelling and comparing clusters. Scientometrics, 111 (2):1157–1167, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Wan:2010:JDI

[kWhHRkS10] Jin kun Wan, Ping huan Hua, Ronald Rousseau, and Xiu kun Sun. The journal download immediacy index (DII): experiences using a Chinese full-text database. Scientometrics, 82(3):555–566, March 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0171-2.

Kwiek:2018:HRP

[Kwi18] Marek Kwiek. High research productivity in vertically undifferentiated higher education systems: Who are the top performers? Scientometrics, 115(1):415–462, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2644-7;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2644-
7.pdf.

Kawamura:2018:FMU

[KWM^{+}18] Takahiro Kawamura, Katsutaro Watanabe, Naoya Matsumoto, Shusaku Egami, and Mari Jibu. Funding map using paragraph embedding based on semantic diversity. Scientometrics, 116(2):941–958, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2783-x;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2783-
x.pdf.

Koopman:2017:CTB

[KWS17] Rob Koopman, Shenghui Wang, and Andrea Scharnhorst. Contextualization of topics: browsing through the universe of bibliographic information. Scientometrics, 111(2):1119– 1139, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Krampen:2015:CSD

[KWW15] Günter Krampen, Peter Weiland, and Jürgen Wiesenhütter. Citation success of different publication types: a case study on all references in psychology publications from the German-speaking countries (D-A-CH-L-L) in 2009, 2010, and 2011. Scientometrics, 104(3):827–840, September 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s11192-015-1573-y.pdf.

Khor:2016:IIC

[KY16] K. A. Khor and L.-G. Yu. Influence of international coauthorship on the research citation impact of young universities. Scientometrics, 107(3):1095–1110, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com//content/
pdf/10.1007/s11192-016-1905-6.pdf;   http://link.
springer.com/content/pdf/10.1007/s11192-016-1905-
6.pdf.

Khan:2017:ARB

[KY17] Muhammad Salman Khan and Muhammad Younas. Analyzing readers behavior in downloading articles from IEEE Digital Library: a study of two selected journals in the field of education. Scientometrics, 110(3):1523–1537, March 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2232-7.

Ketzler:2013:CAE

[KZ13] Rolf Ketzler and Klaus F. Zimmermann. A citation-analysis of economic research institutes. Scientometrics, 95(3): 1095–1112, June 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-012-0850-
2.pdf.

Kim:2016:HTD

[KZC16] Meen Chul Kim, Yongjun Zhu, and Chaomei Chen. How are they different? A quantitative domain comparison of information visualization and data visualization (2000– 2014). Scientometrics, 107(1):123–165, April 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1830-0.

Kang:2019:THS

[KZSZ19] Weimin Kang, Shuliang Zhao, Wei Song, and Tao Zhuang. Triple helix in the science and technology innovation centers of China from the perspective of mutual information: a comparative study between Beijing and Shanghai. Scientometrics, 118(3):921–940, March 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03017-y.

Lovakov:2019:BAP

[LA19] Andrey Lovakov and Elena Agadullina. Bibliometric analysis of publications from post-Soviet countries in psychological journals in 1992–2017. Scientometrics, 119(2): 1157–1171, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03087-y.

Laakso:2014:GOA

[Laa14] Mikael Laakso. Green open access policies of scholarly journal publishers: a study of what, when, and where self-archiving is allowed. Scientometrics, 99(2):475–494, May 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1205-3.

Lortie:2013:DCI

[LABL13] Christopher J. Lortie, Lonnie W. Aarssen, Amber E. Budden, and Roosa Leimu. Do citations and impact factors relate to the real numbers in publications? A case study of citation rates, impact, and effect sizes in ecology and evolutionary biology. Scientometrics, 94(2):675–682, February 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-012-0822-6.pdf.

Luiza-Andrade:2017:FDS

[LAdAMJ17] Ana Luiza-Andrade, Luciano Fogaça de Assis Montag, and Leandro Juen. Functional diversity in studies of aquatic macroinvertebrates community. Scientometrics, 111(3): 1643–1656, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Leydesdorff:2015:PIE

[LAHH15] Loet Leydesdorff, Floortje Alkemade, Gaston Heimeriks, and Rinke Hoekstra. Patents as instruments for exploring innovation dynamics: geographic and technological perspectives on “photovoltaic cells”. Scientometrics, 102(1):

629–651, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1447-8.

Lafouge:2015:DTP

[LAL15] Thierry Lafouge, Abdellatif Agouzal, and Genevieve Lallich. The deconstruction of a text: the permanence of the generalized Zipf law-the inter-textual relationship between entropy and effort amount. Scientometrics, 104(1):193–217, July 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1600-z.

Lamirel:2012:NAA

[Lam12] Jean-Charles Lamirel. A new approach for automatizing the analysis of research topics dynamics: application to optoelectronics research. Scientometrics, 93(1):151–166, October 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0771-0.

Lander:2013:SCB

[Lan13] Bryn Lander. Sectoral collaboration in biomedical research and development. Scientometrics, 94(1):343–357, January 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0776-8.

Lariviere:2012:SSC

[Lar12] Vincent Larivière. On the shoulders of students? The contribution of PhD students to the advancement of knowledge. Scientometrics, 90(2):463–481, February 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0495-6.

Lepori:2014:SWD

[LAS14] Benedetto Lepori, Isidro F. Aguillo, and Marco Seeber. Size of web domains and interlinking behavior of higher education institutions in Europe. Scientometrics, 100(2):497–518, August 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1242-6.

Lu:2014:ECA

[LAW14] Kun Lu, Isola Ajiferuke, and Dietmar Wolfram. Extending citer analysis to journal impact evaluation. Scientometrics, 100(1):245–260, July 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1274-y.

Lazaridis:2010:RUD

[Laz10] Themis Lazaridis. Ranking university departments using the mean h-index. Scientometrics, 82(2):211–216, February 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0048-4.

Leydesdorff:2012:TDS

[LB12] Loet Leydesdorff and Lutz Bornmann. Testing differences statistically with the Leiden ranking. Scientometrics, 92 (3):781–783, September 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-012-
0636-6.pdf.

Leydesdorff:2019:III

[LBA19] Loet Leydesdorff, Lutz Bornmann, and Jonathan Adams. The integrated impact indicator revisited (I3*): a nonparametric alternative to the journal impact factor. Scientometrics, 119(3):1669–1694, June 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03099-8;   http://link.springer.com/content/
pdf/10.1007/s11192-019-03099-8.pdf.

Lancho-Barrantes:2019:SMB

[LBCO19] Bárbara S. Lancho-Barrantes and Francisco J. Cantú-Ortiz. Science in Mexico: a bibliometric analysis. Scientometrics, 118(2):499–517, February 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2985-2.

Lancho-Barrantes:2013:CIB

[LBGBdMA13] Bárbara S. Lancho-Barrantes, Vicente P. Guerrero-Bote, and Félix de Moya-Anegón. Citation increments between collaborating countries. Scientometrics, 94(3):817–831, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0797-3.

Lancho-Barrantes:2010:IHR

[LBGBMA10] Bárbara S. Lancho-Barrantes, Vicente P. Guerrero-Bote, and Félix Moya-Anegón. The Iceberg Hypothesis revisited. Scientometrics, 85(2):443–461, November 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0209-5.

Leydesdorff:2019:SCB

[LBO19] Loet Leydesdorff, Lutz Bornmann, and Tobias Opthof. h_{α}: the scientist as chimpanzee or bonobo. Scientometrics, 118(3):1163–1166, March 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03004-3;   http://link.springer.com/content/
pdf/10.1007/s11192-019-03004-3.pdf.

Leahey:2019:USC

[LBRR19] Erin Leahey, Sondra N. Barringer, and Misty Ring-Ramirez. Universities’ structural commitment to interdisciplinary research. Scientometrics, 118(3):891–919, March 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2992-3.

Leydesdorff:2017:GCJ

[LBW17] Loet Leydesdorff, Lutz Bornmann, and Caroline S. Wagner. Generating clustered journal maps: an automated system for hierarchical classification. Scientometrics, 110(3): 1601–1614, March 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-016-2226-
5.pdf.

Liu:2012:PCC

[LC12] Shengbo Liu and Chaomei Chen. The proximity of cocitation. Scientometrics, 91(2):495–511, May 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0575-7.

Liao:2018:EKP

[LC18] Chien Hsiang Liao and Mu-Yen Chen. Exploring knowledge patterns of library and information science journals within the field: a citation analysis from 2009 to 2016. Scientometrics, 117(3):1991–2008, December 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2925-1.

Liu:2012:CPT

[LCC12] Hsuan-I Liu, Bi-Chun Chang, and Kuan-Chia Chen. Collaboration patterns of Taiwanese scientific publications in various research areas. Scientometrics, 92(1):145–155, July 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0719-4.

Liu:2014:LRB

[LCD^{+}14] Shengbo Liu, Chaomei Chen, Kun Ding, Bo Wang, Kan Xu, and Yuan Lin. Literature retrieval based on citation context. Scientometrics, 101(2):1293–1307, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-014-1233-7.

Liaw:2014:CTI

[LCFC14] Yi-Ching Liaw, Te-Yi Chan, Chin-Yuan Fan, and Cheng-Hsin Chiang. Can the technological impact of academic journals be evaluated? The practice of non-patent reference (NPR) analysis. Scientometrics, 101(1):17–37, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1337-0.

Lozano:2019:CNA

[LCIADG19] S. Lozano, L. Calzada-Infante, B. Adenso-Díaz, and S. García. Complex network analysis of keywords cooccurrence in the recent efficiency analysis literature. Scientometrics, 120(2):609–629, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03132-w.

Liu:2016:MNR

[LCLX16] Ping Liu, Bao-Li Chen, Kan Liu, and Hao Xie. Magnetic nanoparticles research: a scientometric analysis of development trends and research fronts. Scientometrics, 108(3):1591–1602, September 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2017-z.

Leydesdorff:2013:GMS

[LCR13] Loet Leydesdorff, Stephen Carley, and Ismael Rafols. Global maps of science based on the new Web-of-Science categories. Scientometrics, 94(2):589–593, February 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-012-0784-8.pdf.

Leydesdorff:2016:CRM

[LCS^{+}16] Loet Leydesdorff, Jordan A. Comins, Aaron A. Sorensen, Lutz Bornmann, and Iina Hellsten. Cited references and Medical Subject Headings (MeSH) as two different knowledge representations: clustering and mappings at the paper level. Scientometrics, 109(3):2077–2091, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s11192-016-2119-7.pdf.

Liang:2012:RCU

[LCWY12] Liming Liang, Lixin Chen, Yishan Wu, and Junpeng Yuan. The role of Chinese universities in enterprise-university research collaboration. Scientometrics, 90(1):253–269, January 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0503-x.

Lin:2014:ARR

[LCY14] Pin-Hua Lin, Jong-Rong Chen, and Chih-Hai Yang. Academic research resources and academic quality: a crosscountry analysis. Scientometrics, 101(1):109–123, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1362-z.

Li:2017:HDF

[LCZ17] Ming Li, Xiangdong Chen, and Gupeng Zhang. How does firm size affect technology licensing? Empirical evidence from China. Scientometrics, 112(3):1249–1269, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2451-6.

Lindahl:2016:IVE

[LD16] Jonas Lindahl and Rickard Danell. The information value of early career productivity in mathematics: a ROC analysis of prediction errors in bibliometricly informed decision making. Scientometrics, 109(3):2241–2262, December 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-2097-9.pdf.

Liu:2017:BRA

[LDG17] Weishu Liu, Yishan Ding, and Mengdi Gu. Book reviews in academic journals: patterns and dynamics. Scientometrics, 110(1):355–364, January 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2172-2.

Lopes:2017:FER

[LdSdFFNM17] Renato Matos Lopes, Daniel José Garcia dos Santos de Faria, Antonio Augusto Fidalgo-Neto, and Fabio Batista Mota. Facebook in educational research: a bibliometric analysis. Scientometrics, 111(3):1591–1621, June 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Lopez-Duarte:2019:CND

[LDVSGD19] Cristina López-Duarte, Marta M. Vidal-Suárez, and Belén González-Díaz. Cross-national distance and international business: an analysis of the most influential recent models. Scientometrics, 121(1):173–208, October 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03203-y.

Lopez-Duarte:2016:URN

[LDVSGDR16] Cristina López-Duarte, Marta M. Vidal-Suárez, Belén González-Díaz, and Nuno Rosa Reis. Understanding the relevance of national culture in international business research: a quantitative analysis. Scientometrics, 108(3):1553–1590, September 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2044-9.

Lu:2017:UIC

[LDZ17] Chao Lu, Ying Ding, and Chengzhi Zhang. Understanding the impact change of a highly cited article: a contentbased citation analysis. Scientometrics, 112(2):927–945, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2398-7.

Liu:2017:ERS

[LdZwC^{+}17] Fang Liu, Wei dong Zhu, Yu wang Chen, Dong ling Xu, and Jian bo Yang. Evaluation, ranking and selection of R&D projects by multiple experts: an evidential reasoning rule based approach. Scientometrics, 111(3):1501–1519, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Leblond:2012:ASC

[Leb12] Mathieu Leblond. Author self-citations in the field of ecology. Scientometrics, 91(3):943–953, June 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0565-9.

Lee:2010:APP

[Lee10a] Gregory John Lee. Assessing publication performance of research units: extensions through operational research and economic techniques. Scientometrics, 84(3): 717–734, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0210-z.

Lee:2010:SSD

[Lee10b] Yong-Gil Lee. Sectoral strategic differences of technological development between electronics and chemistry: a historical view from analyses of Korean-invented US patents during the period of 1989–1992. Scientometrics, 82(1):83–92, January 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0052-8.

Lee:2012:SCL

[Lee12] Yong-Gil Lee. Strengthening competency linkage to innovation at Korean universities. Scientometrics, 90(1):219–230, January 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0497-4.

Lee:2015:UMN

[Lee15] Hakyeon Lee. Uncovering the multidisciplinary nature of technology management: journal citation network analysis. Scientometrics, 102(1):51–75, January 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1350-3.

Lee:2019:PRP

[Lee19a] Danielle H. Lee. Predicting the research performance of early career scientists. Scientometrics, 121(3):1481– 1504, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03232-7.

Lee:2019:PPC

[Lee19b] Danielle H. Lee. Predictive power of conference-related factors on citation rates of conference papers. Scientometrics, 118(1):281–304, January 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2943-z.

Lei:2016:WSM

[Lei16] Lei Lei. When science meets cluttered writing: adjectives and adverbs in academia revisited. Scientometrics, 107(3):

1361–1372, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1896-3.

Levitt:2015:WON

[Lev15] Jonathan M. Levitt. What is the optimal number of researchers for social science research? Scientometrics, 102 (1):213–225, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1441-1.

Leydesdorff:2011:KBA

[Ley11a] Loet Leydesdorff. Katy Börner: atlas of science: visualizing what we know. Scientometrics, 88(2):675–677, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-011-0409-7.pdf.

Leydesdorff:2011:SIO

[Ley11b] Loet Leydesdorff. “Structuration” by intellectual organization: the configuration of knowledge in relations among structural components in networks of science. Scientometrics, 88(2):499–520, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-011-
0397-7.pdf.

Leydesdorff:2012:AJI

[Ley12] Loet Leydesdorff. Alternatives to the journal impact factor: I3 and the top-10% (or top-25%?) of the most-highly cited papers. Scientometrics, 92(2):355–365, August 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-012-0660-6.pdf.

Leydesdorff:2013:EIN

[Ley13a] Loet Leydesdorff. An evaluation of impacts in “nanoscience & nanotechnology”: steps towards standards for citation analysis. Scientometrics, 94(1):35–55, January 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-012-0750-5.pdf.

Leydesdorff:2013:SDA

[Ley13b] Loet Leydesdorff. Statistics for the dynamic analysis of scientometric data: the evolution of the sciences in terms of trajectories and regimes. Scientometrics, 96(3):731– 741, September 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0917-0.

Leydesdorff:2015:CIP

[Ley15a] Loet Leydesdorff. Can intellectual processes in the sciences also be simulated? The anticipation and visualization of possible future states. Scientometrics, 105(3): 2197–2214, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http:

//link.springer.com//content/pdf/10.1007/s11192-
015-1630-6.pdf;   http://link.springer.com/content/
pdf/10.1007/s11192-015-1630-6.pdf.

Leydesdorff:2015:CTL

[Ley15b] Loet Leydesdorff. Can technology life-cycles be indicated by diversity in patent classifications? The crucial role of variety. Scientometrics, 105(3):1441–1451, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com//content/
pdf/10.1007/s11192-015-1639-x.pdf;   http://link.
springer.com/content/pdf/10.1007/s11192-015-1639-
x.pdf.

Leydesdorff:2018:DIH

[Ley18] Loet Leydesdorff. Diversity and interdisciplinarity: how can one distinguish and recombine disparity, variety, and balance? Scientometrics, 116(3):2113–2121, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2810-y;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2810-
y.pdf.

Liu:2012:FSC

[LF12a] Xuan Zhen Liu and Hui Fang. Fairly sharing the credit of multi-authored papers and its application in the modification of h-index and g-index. Scientometrics, 91(1):37–49, April 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0571-y.

Liu:2012:PRC

[LF12b] Xuan Zhen Liu and Hui Fang. Peer review and overcompetitive research funding fostering mainstream opinion to monopoly. Part II. Scientometrics, 90(2):607–616, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0526-3.

Liu:2014:IPM

[LF14a] Xuan Zhen Liu and Hui Fang. The impact of publications from mainland China on the trends in alphabetical authorship. Scientometrics, 99(3):865–879, June 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1219-x.

Liu:2014:SGL

[LF14b] Xuan Zhen Liu and Hui Fang. Scientific group leaders’ authorship preferences: an empirical investigation. Scientometrics, 98(2):909–925, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1083-8.

Liu:2017:WWC

[LF17] Xuan Zhen Liu and Hui Fang. What we can learn from tweets linking to research papers. Scientometrics, 111(1): 349–369, April 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-017-
2279-0.

Levene:2019:CCI

[LFBI19] Mark Levene, Trevor Fenner, and Judit Bar-Ilan. Characterisation of the χ-index and the rec-index. Scientometrics, 120 (2):885–896, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03151-7.

Liu:2014:CGT

[LFLG14] Ai-Yuan Liu, Hui-Zhen Fu, Shi-Ying Li, and Yu-Qing Guo. Comments on “Global trends of solid waste research from 1997 to 2011 by using bibliometric analysis”. Scientometrics, 98(1):767–774, January 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1086-5. See [YCL^{+}13a].

Lillquist:2010:DDC

[LG10] Eva Lillquist and Sheldon Green. The discipline dependence of citation statistics. Scientometrics, 84(3):749– 762, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0162-3.

Liu:2015:DEC

[LG15] Na Liu and Jiancheng Guan. Dynamic evolution of collaborative networks: evidence from nano-energy research in China. Scientometrics, 102(3):1895–1919, March 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1508-z.

Liu:2016:MIS

[LG16] Chengliang Liu and Qinchang Gui. Mapping intellectual structures and dynamics of transport geography research: a scientometric overview from 1982 to 2014. Scientometrics, 109(1):159–184, October 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2045-8.

Liu:2011:HCM

[LGD11] Xinhai Liu, Wolfgang Glänzel, and Bart De Moor. Hybrid clustering of multi-view data via Tucker-2 model and its application. Scientometrics, 88(3):819–839, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0348-3.

Liu:2012:OHC

[LGD12] Xinhai Liu, Wolfgang Glänzel, and Bart De Moor. Optimal and hierarchical clustering of large-scale hybrid networks for scientific mapping. Scientometrics, 91(2):473–493, May 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0600-x.

Liu:2014:PDB

[LGH^{+}14] Weishu Liu, Mengdi Gu, Guangyuan Hu, Chao Li, Huchang Liao, Li Tang, and Philip Shapira. Profile of developments in biomass-based bioenergy research: a 20-year perspective. Scientometrics, 99(2):507–521, May 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1152-z.

Lang:2010:SCL

[LGL10] Pamela Lang, Fábio C. Gouveia, and Jacqueline Leta. Site co-link analysis applied to small networks: a new methodological approach. Scientometrics, 83(1):157–166, April 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0092-0.

Lathabai:2018:IAP

[LGPC18] Hiran H. Lathabai, Susan George, Thara Prabhakaran, and Manoj Changat. An integrated approach to path analysis for weighted citation networks. Scientometrics, 117(3): 1871–1904, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2917-1.

Lacasa:2017:TCC

[LGR17] Iciar Dominguez Lacasa, Alexander Giebler, and Slavo Radosević. Technological capabilities in Central and Eastern Europe: an analysis based on priority patents. Scientometrics, 111(1):83–102, April 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-017-2277-2.

Lung:2018:HMR

[LGS18] Rodica Ioana Lung, Noémi Gaskó, and Mihai Alexandru Suciu. A hypergraph model for representing scientific output. Scientometrics, 117(3):1361–1379, December 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2908-2.

Li:2013:KMR

[LGZ^{+}13] Changling Li, Fengjiao Guo, Ling Zhi, Zhiping Han, and Feifan Liu. Knowledge management research status in China from 2006 to 2010: based on analysis of the degree theses. Scientometrics, 94(1):95–111, January 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0858-7.

Lin:2012:RBC

[LH12] Wen-Yau Cathy Lin and Mu-Hsuan Huang. The relationship between co-authorship, currency of references and author self-citations. Scientometrics, 90(2):343–360, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0523-6.

Liao:2014:GTA

[LH14] Jingqiu Liao and Yi Huang. Global trend in aquatic ecosystem research from 1992 to 2011. Scientometrics, 98(2): 1203–1219, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1071-z.

Lu:2018:FSI

[LHBC18] Wei Lu, Yong Huang, Yi Bu, and Qikai Cheng. Functional structure identification of scientific documents in computer science. Scientometrics, 115(1):463–486, April 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2640-y.

Lin:2016:BSE

[LHC16] Arthur Jing Lin, Chien-Lung Hsu, and Chun-Hao Chiang. Bibliometric study of electronic commerce research in information systems & MIS journals. Scientometrics, 109(3):1455–1476, December 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2142-8. See concern [Glä18].

Li:2018:DPC

[LHCH18] Shaobo Li, Jie Hu, Yuxin Cui, and Jianjun Hu. DeepPatent: patent classification with convolutional neural networks and word embedding. Scientometrics, 117(2):721–744, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2905-5.

Lepori:2016:SIP

[LHG16a] Benedetto Lepori, Sybille Hinze, and Wolfgang Glänzel. Special issue: papers from the 20th International Conference on Science and Technology Indicators. Scientometrics, 109(3):2161–2164, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-2128-6.pdf.

Liu:2016:PPF

[LHG16b] Weishu Liu, Guangyuan Hu, and Mengdi Gu. The probability of publishing in first-quartile journals. Scientometrics, 106(3):1273–1276, March 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1821-1.

Liang:2019:QTT

[LHLH19] Guoqiang Liang, Haiyan Hou, Xiaodan Lou, and Zhigang Hu. Qualifying threshold of “take-off” stage for successfully disseminated creative ideas. Scientometrics, 120(3):1193– 1208, September 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03154-4.

Liesch:2011:EIB

[LHM^{+}11] Peter W. Liesch, Lars Håkanson, Sara L. McGaughey, Stuart Middleton, and Julia Cretchley. The evolution of the international business field: a scientometric investigation of articles published in its premier journal. Scientometrics, 88(1):17–42, July 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0372-3.

Liu:2018:PCM

[LHTL18] Fang Liu, Guangyuan Hu, Li Tang, and Weishu Liu. The penalty of containing more non-English articles. Scientometrics, 114(1):359–366, January 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2577-6.

Liu:2012:CWA

[LHW12] Gao-Yong Liu, Ji-Ming Hu, and Hui-Ling Wang. A co-word analysis of digital library field in China. Scientometrics, 91 (1):203–217, April 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0586-4.

Lin:2016:CBE

[LHW16] Zhenquan Lin, Shanci Hou, and Jinshan Wu. The correlation between editorial delay and the ratio of highly cited papers in Nature, Science and Physical Review Letters. Scientometrics, 107(3):1457–1464, June 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1936-z.

Luan:2014:SIM

[LHWW14] Chunjuan Luan, Haiyan Hou, Yongtao Wang, and Xianwen Wang. Are significant inventions more diversified? Scientometrics, 100(2):459–470, August 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1303-x.

Li:2014:CCA

[Li14]  Jiang Li.  Citation curves of “all-elements-sleeping-

beauties”: “flash in the pan” first and then “delayed recognition”. Scientometrics, 100(2):595–601, August 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1217-z.

Li:2015:NTD

[Li15] Munan Li. A novel three-dimension perspective to explore technology evolution. Scientometrics, 105(3):1679– 1697, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1591-9.

Li:2016:WWL

[Li16] John T. Li. What we learn from the shifts in highly cited data from 2001 to 2014? Scientometrics, 108(1):57–82, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1958-6.

Li:2017:EPN

[Li17] Ning Li. Evolutionary patterns of national disciplinary profiles in research: 1996–2015. Scientometrics, 111(1): 493–520, April 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-017-
2259-4.

Li:2018:CRT

[Li18] Munan Li. Classifying and ranking topic terms based on a novel approach: role differentiation of author keywords. Scientometrics, 116(1):77–100, July 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2741-7.

Li:2019:VSS

[Li19] Munan Li. Visualizing the studies on smart cities in the past two decades: a two-dimensional perspective. Scientometrics, 120(2):683–705, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03134-8.

Liao:2011:HIR

[Lia11] Chien Hsiang Liao. How to improve research quality? Examining the impacts of collaboration intensity and member diversity in collaboration networks. Scientometrics, 86 (3):747–761, March 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0309-2.

Lopez-Illescas:2011:RUS

[LIdMAM11] Carmen López-Illescas, Félix de Moya-Anegón, and Henk F. Moed. A ranking of universities should account for differences in their disciplinary specialization. Scientometrics, 88 (2):563–574, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0398-6.

Linmans:2010:WBH

[Lin10] A. J. M. Linmans. Why with bibliometrics the humanities does not need to be the weakest link. Scientometrics, 83 (2):337–354, May 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0088-9.

Lin:2011:SPL

[Lin11] Fang Lin. A study on power-law distribution of hostnames in the URL references. Scientometrics, 88(1):191– 198, July 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0377-y.

Lin:2012:RSC

[Lin12] Wen-Yau Cathy Lin. Research status and characteristics of library and information science in Taiwan: a bibliometric analysis. Scientometrics, 92(1):7–21, July 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0725-6.

Lin:2018:ACF

[Lin18] Chi-Shiou Lin. An analysis of citation functions in the humanities and social sciences research from the perspective of problematic citation analysis assumptions. Scientometrics, 116(2):797–813, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2770-2.

Liu:2013:VPP

[Liu13] Guifeng Liu. Visualization of patents and papers in terahertz technology: a comparative study. Scientometrics, 94 (3):1037–1056, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0782-x.

Liu:2016:CCA

[Liu16] Weishu Liu. Comments on “A comparative analysis of scientific publications in management journals by authors from Mainland China, Hong Kong, Taiwan, and Macau: 2003-2012”. Scientometrics, 106(3):1269–1272, March 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1809-x. See [LXL15].

Liu:2017:NBC

[Liu17] Rey-Long Liu. A new bibliographic coupling measure with descriptive capability. Scientometrics, 110(2):915–935, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2196-7.

Liu:2019:DSS

[Liu19] Weishu Liu. The data source of this study is Web of Science Core Collection? Not enough. Scientometrics, 121(3): 1815–1824, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03238-1.

Lin:2010:IMC

[LJ10] Chih-Hao Lin and Show-Ling Jang. The impact of M&As on company innovation: evidence from the US medical device industry. Scientometrics, 84(1):119–131, July 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0096-9.

Light:2016:KME

[Lja16] Ryan Light and jimi adams. Knowledge in motion: the evolution of HIV/AIDS research. Scientometrics, 107(3):

1227–1248, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1933-2.

Liu:2015:MKD

[LJC^{+}15] Xuan Liu, Shan Jiang, Hsinchun Chen, Catherine A. Larson, and Mihail C. Roco. Modeling knowledge diffusion in scientific innovation networks: an institutional comparison between China and US with illustration for nanotechnology. Scientometrics, 105(3):1953–1984, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1761-9.

Luiz:2016:EFR

[LJJ^{+}16] Joao Victor Rojas Luiz, Daniel Jugend, Charbel José Chiappeta Jabbour, Octaviano Rojas Luiz, and Fernando Bernardi de Souza. Ecodesign field of research throughout the world: mapping the territory by using an evolutionary lens. Scientometrics, 109(1):241–259, October 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2043-x.

Li:2015:BAF

[LJKG15] Jie Li, Aleksandar Jovanovic, Peter Klimek, and Xiaohong Guo. Bibliometric analysis of fracking scientific literature. Scientometrics, 105(2):1273–1284, November 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1739-7.

Loudcher:2015:COI

[LJMF15] Sabine Loudcher, Wararat Jakawat, Edmundo Pavel Soriano Morales, and Cécile Favre. Combining OLAP and information networks for bibliographic data analysis: a survey. Scientometrics, 103(2):471–487, May 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1539-0.

Lee:2016:SMT

[LJS16] Keeheon Lee, Hyojung Jung, and Min Song. Subjectmethod topic network analysis in communication studies.

Scientometrics, 109(3):1761–1787, December 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2135-7.

Lee:2017:KND

[LK17] Sanghoon Lee and Wonjoon Kim. The knowledge network dynamics in a mobile ecosystem: a patent citation analysis. Scientometrics, 111(2):717–742, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Lapon-Kandelshein:2011:BRS

[LKP11] Esther Lapon-Kandelshein and Gila Prebor. Bibliographical research in the study of Hebrew printing: a bibliometric analysis. Scientometrics, 88(3):899–913, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0423-9.

Leydesdorff:2014:IOM

[LKR14] Loet Leydesdorff, Duncan Kushnir, and Ismael Rafols. Interactive overlay maps for US patent (USPTO) data based on international patent classification (IPC). Scientometrics, 98(3):1583–1599, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0923-2.

Lee:2014:TOI

[LKS^{+}14] Yongho Lee, So Young Kim, Inseok Song, Yongtae Park, and Juneseuk Shin. Technology opportunity identification customized to the technological capability of SMEs through two-stage patent analysis. Scientometrics, 100(1):227–244, July 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1216-0.

Lee:2015:CAR

[LKSK15] Hoyeop Lee, Jueun Kwak, Min Song, and Chang Ouk Kim. Coherence analysis of research and education using topic modeling. Scientometrics, 102(2):1119–1137, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1453-x.

Lewison:2016:CEG

[LKW^{+}16] Grant Lewison, Sameer Kumar, Chan-Yuan Wong, Philip Roe, and Richard Webber. The contribution of ethnic groups to Malaysian scientific output, 1982–2014, and the effects of the new economic policy. Scientometrics, 109(3):1877–1893, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-2139-3.pdf.

Leydesdorff:2017:MPC

[LKY17] Loet Leydesdorff, Dieter Franz Kogler, and Bowen Yan. Mapping patent classifications: portfolio and statistical analysis, and the comparison of strengths and weaknesses. Scientometrics, 112(3):1573–1591, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2449-0;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2449-
0.pdf.

Lee:2010:DCB

[LL10] Yong-Gil Lee and Ji-Hoon Lee. Different characteristics between auctioned and non-auctioned patents. Scientometrics, 82(1):135–148, January 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
009-0029-7.

Lou:2012:EGR

[LL12] Yu-Chiung Lou and Hsiao-Fang Lin. Estimate of global research trends and performance in family therapy in social Science Citation Index. Scientometrics, 90(3):807–823, March 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0549-9.

Labbe:2013:DFP

[LL13a] Cyril Labbé and Dominique Labbé. Duplicate and fake publications in the scientific literature: how many SCIgen papers in computer science? Scientometrics, 94(1):379–396, January 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0781-y.

Lu:2013:IAI

[LL13b] Louis Y. Y. Lu and John S. Liu. An innovative approach to identify the knowledge diffusion path: the case of resourcebased theory. Scientometrics, 94(1):225–246, January 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0744-3.

Li:2015:PEC

[LL15] Jiang Li and Yueting Li. Patterns and evolution of coauthorship in China’s humanities and social sciences. Scientometrics, 102(3):1997–2010, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1471-8.

Laakso:2016:JCR

[LL16] Mikael Laakso and Juho Lindman. Journal copyright restrictions and actual open access availability: a study of articles published in eight top information systems journals (2010–2014). Scientometrics, 109(2):1167–1189, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2078-z.

Lee:2019:TOA

[LL19] Changyong Lee and Gyumin Lee. Technology opportunity analysis based on recombinant search: patent landscape analysis for idea generation. Scientometrics, 121(2): 603–632, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03224-7.

Lai:2017:SAE

[LLC^{+}17] Kuei-Kuei Lai, Chien-Yu Lin, Yu-Hsin Chang, Ming-Chung Yang, and Wen-Goang Yang. A structured approach to explore technological competencies through R&D portfolio of photovoltaic companies by patent statistics. Scientometrics, 111(3):1327–1351, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Lee:2011:ROE

[LLCL11] Ling-Chu Lee, Pin-Hua Lin, Yun-Wen Chuang, and Yi-Yang Lee. Research output and economic productivity: a Granger causality test. Scientometrics, 89(2):465– 478, November 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0476-9.

Liu:2014:CRB

[LLG14] Ai-Yuan Liu, Shi-Ying Li, and Yu-Qing Guo. Characteristics of research on bioinformatics in China assessed with Science Citation Index expanded. Scientometrics, 99(2):371– 391, May 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1231-9.

Liu:2013:PGP

[LLGW13] Zhongqiu Liu, Yaolin Liu, Yangjie Guo, and Hua Wang. Progress in global parallel computing research: a bibliometric approach. Scientometrics, 95(3):967–983, June 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0927-y. See rebuttal [Ho16].

Ling:2016:GAL

[LLH^{+}16] Xiaoxi Ling, Yu Liu, Zhen Huang, Parantu K. Shah, and Cheng Li. A graphical article-level metric for intuitive comparison of large-scale literatures. Scientometrics, 106(1): 41–50, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1782-4.

Liu:2019:FNM

[LLH19] John S. Liu, Louis Y. Y. Lu, and Mei Hsiu-Ching Ho. A few notes on main path analysis. Scientometrics, 119 (1):379–391, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03034-x.

Lyu:2017:SEI

[LLHN17] Yibo Lyu, Quanshan Liu, Binyuan He, and Jingfei Nie. Structural embeddedness and innovation diffusion: the moderating role of industrial technology grouping. Scientometrics, 111(2):889–916, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Lee:2012:BAD

[LLL12] Ling-Chu Lee, Yi-Yang Lee, and Yi-Ching Liaw. Bibliometric analysis for development of research strategies in agricultural technology: the case of Taiwan. Scientometrics, 93(3): 813–830, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0833-3.

Li:2018:III

[LLLL18] Yang Li, Huajiao Li, Nairong Liu, and Xueyong Liu. Important institutions of interinstitutional scientific collaboration networks in materials science. Scientometrics, 117 (1):85–103, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2837-0.

Lorenzo:2016:BIS

[LLP^{+}16] Gonzalo Lorenzo, Asunción Lledó, Jorge Pomares, Rosabel Roig, and Pilar Arnaiz. Bibliometric indicators in the study of Asperger syndrome between 1990 and 2014. Scientometrics, 109(1):377–388, October 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-1975-5.

Leta:2010:ICS

[LLRG10] Jacqueline Leta, Birger Larsen, Ronald Rousseau, and Wolfgang Glänzel. The 12th International Conference on Scientometrics and Informetrics. Scientometrics, 83(3):603–604, June 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-010-0219-3.pdf.

Luan:2013:DCT

[LLW13] Chunjuan Luan, Zeyuan Liu, and Xianwen Wang. Divergence and convergence: technology-relatedness evolution in solar energy industry. Scientometrics, 97(2):461– 475, November 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1057-x.

Liu:2016:GRT

[LLW^{+}16] Fenglian Liu, Aiwen Lin, Huanhuan Wang, Yuling Peng, and Song Hong. Global research trends of geographical information system from 1961 to 2010: a bibliometric analysis. Scientometrics, 106(2):751–768, February 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1789-x.

Liu:2018:MVN

[LLX^{+}18] Yu Liu, Dan Lin, Xiujuan Xu, Shimin Shan, and Quan Z. Sheng. Multi-views on Nature Index of Chinese academic institutions. Scientometrics, 114(3):823–837, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2581-x.

Luor:2014:TCE

[LLYC14] Tainyi Luor, Hsi-Peng Lu, Hueiju Yu, and Kuoliang Chang. Trends in and contributions to entrepreneurship research: a broad review of literature from 1996 to June 2012. Scientometrics, 99(2):353–369, May 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-013-1203-5.pdf.

Leydesdorff:2010:DUP

[LM10] Loet Leydesdorff and Martin Meyer. The decline of university patenting and the end of the Bayh–Dole effect. Scientometrics, 83(2):355–362, May 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0001-6. See reply [LM13a].

Lewison:2011:FRR

[LM11] Grant Lewison and Valentina Markusova. Female researchers in Russia: have they become more visible? Scientometrics, 89(1):139–152, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0435-5.

Leydesdorff:2013:REC

[LM13a] Loet Leydesdorff and Martin Meyer. A reply to Etzkowitz’ comments to Leydesdorff and Martin (2010): technology transfer and the end of the Bayh–Dole effect. Scientometrics, 97(3):927–934, December 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0997-5. See [LM10].

Liu:2013:TDK

[LM13b] Xiang Liu and Feicheng Ma. Transfer and distribution of knowledge creation activities of bio-scientists in knowledge space. Scientometrics, 95(1):299–310, April 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0827-1.

Liu:2015:HCI

[LM15] Linqing Liu and Shiye Mei. How can an indigenous concept enter the international academic circle: the case of guanxi. Scientometrics, 105(1):645–663, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1687-2.

Liu:2016:VGR

[LM16] Linqing Liu and Shiye Mei. Visualizing the GVC research: a co-occurrence network based bibliometric analysis. Scientometrics, 109(2):953–977, November 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2100-5.

Lamba:2019:MTD

[LM19] Manika Lamba and Margam Madhusudhan. Mapping of topics in DESIDOC Journal of Library and Information Technology, India: a study. Scientometrics, 120(2):477–505, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03137-5.

Luque-Martinez:2016:CSI

[LMdBG16] Teodoro Luque-Martínez and Salvador del Barrio-García. Constructing a synthetic indicator of research activity. Scientometrics, 108(3):1049–1064, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2037-8.

Lang:2019:ACS

[LMKG19] Raynell Lang, Marcy Mintz, Hartmut B. Krentz, and M. John Gill. An approach to conference selection and evaluation: advice to avoid “predatory” conferences. Scientometrics, 118(2):687–698, February 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2981-6.

Leite:2011:NII

[LML11] Paula Leite, Rogério Mugnaini, and Jacqueline Leta. A new indicator for international visibility: exploring Brazilian scientific community. Scientometrics, 88(1):311–319, July 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0379-9.

Landini:2015:SDN

[LMM15] Fabio Landini, Franco Malerba, and Roberto Mavilia. The structure and dynamics of networks of scientific collaborations in Northern Africa. Scientometrics, 105(3):1787– 1807, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1635-1.

Loos:2016:MSA

[LMR16] Mauricio Johnny Loos, Eugenio Merino, and Carlos Manuel Taboada Rodriguez. Mapping the state of the art of ergonomics within logistics. Scientometrics, 109(1):85–101, October 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-1960-z.

Low:2014:ETI

[LNK^{+}14a] Wah Yun Low, Kwan Hoong Ng, M. A. Kabir, Ai Peng Koh, and Janaki Sinnasamy. Erratum to: Trend and impact of international collaboration in clinical medicine papers published in Malaysia. Scientometrics, 100(2):607, August 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-014-1270-2.pdf.

Low:2014:TII

[LNK^{+}14b] Wah Yun Low, Kwan Hoong Ng, M. A. Kabir, Ai Peng Koh, and Janaki Sinnasamy. Trend and impact of international collaboration in clinical medicine papers published in Malaysia. Scientometrics, 98(2):1521–1533, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s11192-013-1121-6.pdf.

Lopez-Navarro:2015:WDP

[LNMQRR15] Irene López-Navarro, Ana I. Moreno, Miguel Ángel Quintanilla, and Jesús Rey-Rocha. Why do I publish research articles in English instead of my own language? Differences in Spanish researchers’ motivations across scientific domains. Scientometrics, 103(3):939–976, June 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1570-1.

Lezama-Nicolas:2018:BMA

[LNRSRBB18] René Lezama-Nicolás, Marisela Rodríguez-Salvador, Rosa Río-Belver, and Iñaki Bildosola. A bibliometric method for assessing technological maturity: the case of additive manufacturing. Scientometrics, 117(3):1425–1452, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2941-1;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2941-
1.pdf.

Leydesdorff:2012:REV

[LO12] Loet Leydesdorff and Tobias Opthof. A rejoinder on energy versus impact indicators. Scientometrics, 90(2):745–748, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-011-0502-y.pdf.

Lopez-Olmedo:2017:RPP

[LOMLPA^{+}17] R. Lopez-Olmedo, R. Marmolejo-Leyva, M. A. Perez-Angon, L. L. Villa-Vazquez, and E. Zayago-Lau. The role of public policies in the decentralization process of Mexican science and the formation of new researchers in institutions outside the Mexico City area. Scientometrics, 112(3):1343– 1366, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2423-x.

Lopresti:2010:CAE

[Lop10] Robert Lopresti. Citation accuracy in environmental science journals. Scientometrics, 85(3):647–655, December 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0293-6.

Lopez-Olmedo:2019:PMC

[LOPAGS19] Roberto Lopez-Olmedo, Miguel Angel Perez-Angon, and Norma Georgina Gutiérrez-Serrano. Participation of Mexican civil society organizations in scientific publications. Scientometrics, 119(1):55–72, April 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03027-w.

Lortie:2010:LEG

[Lor10] Christopher J. Lortie. Letter to the Editor: A global comment on scientific publications, productivity, people, and beer. Scientometrics, 84(2):539–541, August 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0077-z.

Lorentzen:2014:WBW

[Lor14] David Gunnarsson Lorentzen. Webometrics benefitting from web mining? An investigation of methods and applications of two research fields. Scientometrics, 99(2):409–445, May 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1227-x.

Lim:2010:ITK

[LP10] Hyojeong Lim and Yongtae Park. Identification of technological knowledge intermediaries. Scientometrics, 84(3): 543–561, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0133-8.

Lee:2012:EWV

[LP12] Moosung Lee and Han Woo Park. Exploring the web visibility of world-class universities. Scientometrics, 90(1): 201–218, January 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0515-6.

Laakso:2018:OAE

[LP18a] Mikael Laakso and Andrea Polonioli. Open access in ethics research: an analysis of open access availability and author self-archiving behaviour in light of journal copyright restrictions. Scientometrics, 116(1):291–317, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2751-5;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2751-
5.pdf.

Li:2018:FDR

[LP18b] Munan Li and Alan L. Porter. Facilitating the discovery of relevant studies on risk analysis for three-dimensional printing based on an integrated framework. Scientometrics, 114 (1):277–300, January 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2570-0.

Light:2014:ODO

[LPB14] Robert P. Light, David E. Polley, and Katy Börner. Open data and open code for big science of science studies. Scientometrics, 101(2):1535–1551, November 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1238-2.

Lathabai:2017:CPA

[LPC17] Hiran H. Lathabai, Thara Prabhakaran, and Manoj Changat. Contextual productivity assessment of authors and journals: a network scientometric approach. Scientometrics, 110(2):711–737, February 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2202-0.

Leydesdorff:2014:RMS

[LPL14] Loet Leydesdorff, Han Woo Park, and Balazs Lengyel. A routine for measuring synergy in university-industry-government relations: mutual information as a triple-helix and quadruple-helix indicator. Scientometrics, 99(1):27–35, April 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1079-4.

Li:2016:TCB

[LPL16] Zhi Li, Qinke Peng, and Che Liu. Two citation-based indicators to measure latent referential value of papers. Scientometrics, 108(3):1299–1313, September 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2000-8.

Lee:2017:TMT

[LPMK17] Chul Lee, Gunno Park, Klaus Marhold, and Jina Kang. Top management team’s innovation-related characteristics and the firm’s explorative R&D: an analysis based on patent data. Scientometrics, 111(2):639–663, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Lyu:2017:BAS

[LPZ17] Qiu-Ju Lyu, Qiang-Hong Pu, and Jin Zhang. Bibliometric analysis of scientific publications in endocrinology and metabolism from China, Japan, and South Korea. Scientometrics, 110(1):105–112, January 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2179-8.

Li:2017:EEE

[LQW17] Xuerong Li, Han Qiao, and Shouyang Wang. Exploring evolution and emerging trends in business model study: a co-citation analysis. Scientometrics, 111(2):869–887, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

LeMoigne:2012:SIC

[LR12a] Philippe Le Moigne and Pascal Ragouet. Science as instrumentation. The case for psychiatric rating scales. Scientometrics, 93(2):329–349, November 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0673-1.

Lewison:2012:EIC

[LR12b] Grant Lewison and Philip Roe. The evaluation of Indian cancer research, 1990–2010. Scientometrics, 93(1):167–181, October 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0633-9.

Lozano:2014:AES

[LRA14] Sergi Lozano, Xosé-Pedro Rodríguez, and Alex Arenas. Atapuerca: evolution of scientific collaboration in an emergent large-scale research infrastructure. Scientometrics, 98(2): 1505–1520, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1162-x.

Laverde-Rojas:2019:CSP

[LRC19] Henry Laverde-Rojas and Juan C. Correa. Can scientific productivity impact the economic complexity of countries? Scientometrics, 120(1):267–282, July 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03118-8;   http://link.springer.
com/content/pdf/10.1007/s11192-019-03118-8.pdf.

Li:2018:ICR

[LRS^{+}18] Xiancheng Li, Wenge Rong, Haoran Shi, Jie Tang, and Zhang Xiong. The impact of conference ranking systems in computer science: a comparative regression analysis. Scientometrics, 116(2):879–907, August 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2763-1.

Lewison:2016:LCR

[LRWS16] Grant Lewison, Philip Roe, Richard Webber, and Richard Sullivan. Lung cancer researchers, 2008–2013: their sex and ethnicity. Scientometrics, 106(1):105–117, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1785-1;   http://link.
springer.com/content/pdf/10.1007/s11192-015-1785-
1.pdf.

Li:2018:WSU

[LRY18] Kai Li, Jason Rollins, and Erjia Yan. Web of Science use in published research and review papers 1997-2017: a selective, dynamic, cross-domain, content-based analysis. Scientometrics, 115(1):1–20, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
017-2622-5;   http://link.springer.com/content/pdf/
10.1007/s11192-017-2622-5.pdf.

Liang:2013:NEJ

[LRZ13] Liming Liang, Ronald Rousseau, and Zhen Zhong. Non-English journals and papers in physics and chemistry: bias in citations? Scientometrics, 95(1):333–350, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0828-0.

Lewison:2015:CIS

[LS15] Grant Lewison and Richard Sullivan. Conflicts of interest statements on biomedical papers. Scientometrics, 102(3): 2151–2159, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1507-0.

Lercher:2016:PVO

[LS16] Aaron Lercher and Lawrence Smolinsky. Persistent value of older scientific journal articles. Scientometrics, 108(3):1205–1220, September 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2011-5.

Lee:2017:EED

[LS17a] Bo Kyeong Lee and So Young Sohn. Exploring the effect of dual use on the value of military technology patents based on the renewal decision. Scientometrics, 112(3):1203– 1227, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2443-6.

Lee:2017:WMF

[LS17b] Jungpyo Lee and So Young Sohn. What makes the first forward citation of a patent occur earlier? Scientometrics, 113 (1):279–298, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2480-1.

Lobo:2019:SIN

[LS19a] José Lobo and Deborah Strumsky. Sources of inventive novelty: two patent classification schemas, same story. Scientometrics, 120(1):19–37, July 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03102-2.

Lyu:2019:ABJ

[LS19b] Guoliang Lyu and Ganwei Shi. On an approach to boosting a journal’s citation potential. Scientometrics, 120(3):1387– 1409, September 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03172-2.

Lee:2010:AOB

[LSC10] Pei-Chun Lee, Hsin-Ning Su, and Te-Yi Chan. Assessment of ontology-based knowledge network formation by vector-space model. Scientometrics, 85(3):689–703, December 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0267-8.

Lee:2012:CNP

[LSCK12] Duk Hee Lee, Il Won Seo, Ho Chull Choe, and Hee Dae Kim. Collaboration network patterns and research performance: the case of Korean public research institutions. Scientometrics, 91(3):925–942, June 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0602-8.

Luo:2018:EPC

[LSE^{+}18] Feiheng Luo, Aixin Sun, Mojisola Erdt, Aravind Sesagiri Raamkumar, and Yin-Leng Theng. Exploring prestigious citations sourced from top universities in bibliometrics and altmetrics: a case study in the computer science discipline. Scientometrics, 114(1):1–17, January 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2571-z.

Lundgren:2015:RBD

[LSL15] Silje Lundgren, Margrit Shildrick, and David Lawrence. Rethinking bibliometric data concerning gender studies: a response to Söderlund and Madison. Scientometrics, 105(3):1389–1398, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-

015-1767-3. See [MS16a, MS16a].

Lima:2015:APT

[LSM^{+}15] Harlley Lima, Thiago H. P. Silva, Mirella M. Moro, Rodrygo L. T. Santos, Wagner Meira, Jr., and Alberto H. F. Laender. Assessing the profile of top Brazilian computer science researchers. Scientometrics, 103(3):879–896, June 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1569-7.

Lobo:2013:SPU

[LSR13] J. Lobo, D. Strumsky, and J. Rothwell. Scaling of patenting with urban population size: evidence from global metropolitan areas. Scientometrics, 96(3):819–828, September 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-0970-3.

Li:2015:NMA

[LSS15] Shuqing Li, Ying Sun, and Dagobert Soergel. A new method for automatically constructing domain-oriented term taxonomy based on weighted word co-occurrence analysis. Scientometrics, 103(3):1023–1042, June 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1571-0. See erratum [LSS16].

Li:2016:ENM

[LSS16] Shuqing Li, Ying Sun, and Dagobert Soergel. Erratum to: A new method for automatically constructing domain-oriented term taxonomy based on weighted word co-occurrence analysis. Scientometrics, 108(2):1005, August 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-1832-6.pdf. See [LSS15].

Liu:2011:SIK

[LSY11] Chen Liu, Wei Shan, and Jing Yu. Shaping the interdisciplinary knowledge network of China: a network analysis based on citation data from 1981 to 2010. Scientometrics, 89 (1):89–106, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0450-6.

Levitt:2010:DHC

[LT10a] Jonathan M. Levitt and Mike Thelwall. Does the higher citation of collaborative research differ from region to region? A case study of economics. Scientometrics, 85(1):171–183, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0197-5.

Lewison:2010:NBF

[LT10b] Grant Lewison and Thomas Turnbull. News in brief and features in new scientist magazine and the biomedical research papers that they cite, August 2008 to July 2009. Scientometrics, 85(1):345–359, October 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0221-9.

Levitt:2016:LTP

[LT16] Jonathan M. Levitt and Mike Thelwall. Long term productivity and collaboration in information science. Scientometrics, 108(3):1103–1117, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2061-8.

Li:2012:VOR

[LTG12] Xuemei Li, Mike Thelwall, and Dean Giustini. Validating online reference managers for scholarly impact measurement. Scientometrics, 91(2):461–471, May 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0580-x.

Liu:2015:FRC

[LTGH15] Weishu Liu, Li Tang, Mengdi Gu, and Guangyuan Hu. Feature report on China: a bibliometric analysis of Chinarelated articles. Scientometrics, 102(1):503–517, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1371-y.

Liu:2018:UAA

[LTK^{+}18] Jiaying Liu, Tao Tang, Xiangjie Kong, Amr Tolba, Zafer Al-Makhadmeh, and Feng Xia. Understanding the advisoradvisee relationship via scholarly data analysis. Scientometrics, 116(1):161–180, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2762-2.

Liu:2019:TDI

[LTK^{+}19] Jiaying Liu, Jiahao Tian, Xiangjie Kong, Ivan Lee, and Feng Xia. Two decades of information systems: a bibliometric review. Scientometrics, 118(2):617–643, February 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2974-5.

Ladle:2012:AIG

[LTM12] Richard J. Ladle, Peter A. Todd, and Ana. C. M. Malhado. Assessing insularity in global science. Scientometrics, 93(3): 745–750, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0703-z.

Lund:2019:ECH

[Lun19] Brady Lund. Examination of correlates of H-index as a measure of research productivity for library and information science faculty in the United States and Canada. Scientometrics, 120(2):897–915, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03152-6.

Luo:2012:ODR

[Luo12] Lieh-Ming Luo. Optimal diversification for R&D project portfolios. Scientometrics, 91(1):219–229, April 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0537-0.

Larcombe:2011:SCC

[LV11] Alexander N. Larcombe and Sasha C. Voss. Self-citation: comparison between radiology, European radiology and radiology for 1997–1998. Scientometrics, 87(2):347–356, May 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-010-0328-z. See erratum

[LV12].

Larcombe:2012:ESC

[LV12] Alexander N. Larcombe and Sasha C. Voss. Erratum to: Self-citation in American Physiological Society Journals. Scientometrics, 90(3):1027, March 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-010-0337-y.pdf. See [LV11].

Lariviere:2011:SDR

[LVGV^{+}11] Vincent Larivière, Etienne Vignola-Gagné, Christian Villeneuve, Pascal Gélinas, and Yves Gingras. Sex differences in research funding, productivity and impact: an analysis of Québec university professors. Scientometrics, 87(3):483– 498, June 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0369-y.

Lepori:2015:PEF

[LVHS^{+}15] Benedetto Lepori, Valerio Veglio, Barbara Heller-Schuh, Thomas Scherngell, and Michael Barber. Participations to European Framework Programs of higher education institutions and their association with organizational characteristics. Scientometrics, 105(3):2149–2178, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1768-2.

Larsen:2010:RGS

[LvI10] Peder Olesen Larsen and Markus von Ins. The rate of growth in scientific publication and the decline in coverage provided by Science Citation Index. Scientometrics, 84 (3):575–603, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-010-
0202-z.pdf.

Laurens:2018:APP

[LVSL18] Patricia Laurens, Lionel Villard, Antoine Schoen, and Philippe Larédo. The artificial patents in the PATSTAT database: How much do they matter when computing indicators of internationalisation based on worldwide priority patents? Scientometrics, 114(1):91–112, January 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2578-5.

Liu:2010:FDB

[LW10] Chen-Yuan Liu and Jhen-Cheng Wang. Forecasting the development of the biped robot walking technique in Japan through S-curve model analysis. Scientometrics, 82(1): 21–36, January 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0055-5.

Liberman:2015:ISD

[LW15] Sofia Liberman and Kurt Bernardo Wolf. Independent simultaneous discoveries visualized through network analysis: the case of linear canonical transforms. Scientometrics, 104(3):715–735, September 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1602-x.

Leydesdorff:2016:PCB

[LWB16] Loet Leydesdorff, Paul Wouters, and Lutz Bornmann. Professional and citizen bibliometrics: complementarities and ambivalences in the development and use of indicators — a state-of-the-art report. Scientometrics, 109(3):2129– 2150, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-016-2150-
8.pdf.

Leydesdorff:2018:BDJ

[LWB18a] Loet Leydesdorff, Caroline S. Wagner, and Lutz Bornmann. Betweenness and diversity in journal citation networks as measures of interdisciplinarity — a tribute to Eugene Garfield. Scientometrics, 114(2):567–592, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-017-2528-2;   https://
link.springer.com/content/pdf/10.1007/s11192-017-
2528-2.pdf.

Leydesdorff:2018:DCR

[LWB18b] Loet Leydesdorff, Caroline S. Wagner, and Lutz Bornmann. Discontinuities in citation relations among journals: self-organized criticality as a model of scientific revolutions and change. Scientometrics, 116(1):623–644, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2734-6;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2734-
6.pdf.

Lepori:2016:DUU

[LWIB16] Benedetto Lepori, Michael Wise, Diana Ingenhoff, and Alexander Buhmann. The dynamics of university units as a multi-level process. Credibility cycles and resource dependencies. Scientometrics, 109(3):2279–2301, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2080-5.

Li:2017:QAR

[LWL17] Xingchen Li, Qiang Wu, and Yuanyuan Liu. A quantitative analysis of researcher citation personal display considering disciplinary differences and influence factors. Scientometrics, 113(2):1093–1112, November 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2501-0.

Liu:2015:DIS

[LWM^{+}15] Ping Liu, Qiong Wu, Xiangming Mu, Kaipeng Yu, and Yiting Guo. Detecting the intellectual structure of library and information science based on formal concept analysis. Scientometrics, 104(3):737–762, September 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1629-z.

Li:2017:SAP

[LWR^{+}17] Yixi Li, Yuan Wang, Xue Rui, Yaxiu Li, Yang Li, Huanzhi Wang, Jian Zuo, and Yindong Tong. Sources of atmospheric pollution: a bibliometric analysis. Scientometrics, 112(2): 1025–1045, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2421-z.

Low:2016:SSM

[LWT16] Wan Jing Low, Paul Wilson, and Mike Thelwall. Stopped sum models and proposed variants for citation data. Scientometrics, 107(2):369–384, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
016-1847-z.

Lv:2011:BTA

[LWW^{+}11] Peng Hui Lv, Gui-Fang Wang, Yong Wan, Jia Liu, Qing Liu, and Fei cheng Ma. Bibliometric trend analysis on global graphene research. Scientometrics, 88(2):399–419, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0386-x.

Liu:2015:SEC

[LX15] Peng Liu and Haoxiang Xia. Structure and evolution of co-authorship network in an interdisciplinary research field. Scientometrics, 103(1):101–134, April 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1525-y.

Zhijun:2019:ERA

[LX19] Zhijun LI and Jinfen XU. The evolution of research article titles: the case of Journal of Pragmatics 1978–2018. Scientometrics, 121(3):1619–1634, December 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03244-3.

Lin:2013:FTL

[LXDL13] Lili Lin, Zhuoming Xu, Ying Ding, and Xiaozhong Liu. Finding topic-level experts in scholarly networks. Scientometrics, 97(3):797–819, December 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0988-6.

Luo:2018:DDH

[LXH^{+}18] Qingzhou Luo, Jianhong Cecilia Xia, Gaby Haddow, Michele Willson, and Jun Yang. Does distance hinder the collaboration between Australian universities in the humanities, arts and social sciences? Scientometrics, 115(2):695– 715, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2686-x.

Liu:2015:CAS

[LXL15] Xiaodong Liu, Qi Xu, and Meina Li. A comparative analysis of scientific publications in management journals by authors from Mainland China, Hong Kong, Taiwan, and Macau: 2003–2012. Scientometrics, 105(1):135–143, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1665-8. See comments [Liu16] and

erratum [LXL16].

Liu:2016:ECA

[LXL16] Xiaodong Liu, Qi Xu, and Meina Li. Erratum to: A comparative analysis of scientific publications in management journals by authors from Mainland China, Hong Kong, Taiwan, and Macau: 2003–2012. Scientometrics, 108(2):1007–1009, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11192-016-1840-6.pdf. See

[LXL15].

Li:2017:UOD

[LXWC17] Jianping Li, Yongjia Xie, Dengsheng Wu, and Yuanping Chen. Underestimating or overestimating the distribution inequality of research funding? The influence of funding sources and subdivision. Scientometrics, 112(1):55–74, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Li:2012:PAE

[LY12] Jiang Li and Fred Y. Ye. The phenomenon of allelements-sleeping-beauties in scientific literature. Scientometrics, 92(3):795–799, September 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0643-7.

Lei:2016:RCI

[LY16a] Lei Lei and Sheng Yan. Readability and citations in information science: evidence from abstracts and articles of four journals (2003–2012). Scientometrics, 108(3):1155– 1169, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-016-
2036-9.

Li:2016:DSB

[LY16b] Jiang Li and Fred Y. Ye. Distinguishing sleeping beauties in science. Scientometrics, 108(2):821–828, August 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1977-3.

Li:2012:PER

[LYGQ12] Feng Li, Yong Yi, Xiaolong Guo, and Wei Qi. Performance evaluation of research universities in mainland China, Hong Kong and Taiwan: based on a two-dimensional approach. Scientometrics, 90(2):531–542, February 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0544-1.

Lykke:2018:CBA

[Lyk18] Nina Lykke. Can’t bibliometric analysts do better? How quality assessment without field expertise does not work. Scientometrics, 117(1):655–666, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2872-x;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2872-
x.pdf.

Liu:2015:VIS

[LYLD15] Zhigao Liu, Yimei Yin, Weidong Liu, and Michael Dunford. Visualizing the intellectual structure and evolution of innovation systems research: a bibliometric analysis. Scientometrics, 103(1):135–158, April 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1517-y.

Lee:2017:MTS

[LYS^{+}17] Mi Kyung Lee, Ho Young Yoon, Marc Smith, Hye Jin Park, and Han Woo Park. Mapping a Twitter scholarly communication network: a case of the association of Internet researchers’ conference. Scientometrics, 112(2):767–797, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2413-z.

Li:2019:VTA

[LYW19] Pin Li, Guoli Yang, and Chuanqi Wang. Visual topical analysis of library and information science. Scientometrics, 121(3):1753–1791, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03239-0.

Li-Ying:2013:CFL

[LYWSV13] Jason Li-Ying, Yuandi Wang, Søren Salomo, and Wim Vanhaverbeke. Have Chinese firms learned from their prior technology in-licensing? An analysis based on patent citations. Scientometrics, 95(1):183–195, April 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0802-x.

Leydesdorff:2014:MKB

[LZ14] Loet Leydesdorff and Ping Zhou. Measuring the knowledgebased economy of China in terms of synergy among technological, organizational, and geographic attributes of firms. Scientometrics, 98(3):1703–1719, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1179-1.

Laurens:2010:DGF

[LZB10] Patricia Laurens, Michel Zitt, and Elise Bassecoulard. Delineation of the genomics field by hybrid citation-lexical methods: interaction with experts and validation process. Scientometrics, 82(3):647–662, March 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0177-9.

Liu:2017:EPG

[LZC17] Fengchao Liu, Na Zhang, and Cong Cao. An evolutionary process of global nanotechnology collaboration: a social network analysis of patents at USPTO. Scientometrics, 111(3): 1449–1465, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Lou:2018:RMI

[LZCZ18] Wen Lou, Yuehua Zhao, Yuchen Chen, and Jin Zhang. Research or management? An investigation of the impact of leadership roles on the research performance of academic administrators. Scientometrics, 117(1):191–209, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2842-3.

Li:2015:EBS

[LZFW15] Yongli Li, Guijie Zhang, Yuqiang Feng, and Chong Wu. An entropy-based social network community detecting method and its application to scientometrics. Scientometrics, 102(1):1003–1017, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1377-5.

Liu:2013:CGI

[LZGQ13] Yu Liu, Wei Zuo, Ying Gao, and Yanhong Qiao. Comprehensive geometrical interpretation of h-type indices. Scientometrics, 96(2):605–615, August 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0916-1.

Liu:2012:BSE

[LZH^{+}12] Xingjian Liu, F. Benjamin Zhan, Song Hong, Beibei Niu, and Yaolin Liu. A bibliometric study of earthquake research: 1900–2010. Scientometrics, 92(3):747–765, September 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0599-z. See comments [Ho13a] and

replies [LZH^{+}13].

Liu:2013:RCB

[LZH^{+}13] Xingjian Liu, F. Benjamin Zhan, Song Hong, Beibei Niu, and Yaolin Liu. Replies to comments on “a bibliometric study of earthquake research: 1900–2010”. Scientometrics, 96(3):933–936, September 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0914-3. See [LZH^{+}12, Ho13a].

Luan:2010:PSC

[LZL10] Chunjuan Luan, Chunyan Zhou, and Aiyun Liu. Patent strategy in Chinese universities: a comparative perspective. Scientometrics, 84(1):53–63, July 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0194-8.

Liang:2014:SRM

[LZR14] Liming Liang, Zhen Zhong, and Ronald Rousseau. Scientists’ referencing (mis)behavior revealed by the dissemination network of referencing errors. Scientometrics, 101(3): 1973–1986, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1275-x.

Lei:2012:IAC

[LZZ^{+}12] Xiao-Ping Lei, Zhi-Yun Zhao, Xu Zhang, Dar-Zen Chen, Mu-Hsuan Huang, and Yun-Hua Zhao. The inventive activities and collaboration pattern of universityindustry-government in China based on patent analysis. Scientometrics, 90(1):231–251, January 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0510-y.

Lei:2013:TCP

[LZZ^{+}13] Xiao-Ping Lei, Zhi-Yun Zhao, Xu Zhang, Dar-Zen Chen, Mu-Hsuan Huang, Jia Zheng, Run-Sheng Liu, Jing Zhang, and Yun-Hua Zhao. Technological collaboration patterns in solar cell industry based on patent inventors and assignees analysis. Scientometrics, 96(2):427–441, August 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0944-x.

Ma:2012:DAI

[Ma12] Ruimin Ma. Discovering and analyzing the intellectual structure and its evolution of LIS in China, 1998– 2007. Scientometrics, 93(3):645–659, December 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0702-0.

McGillivray:2019:RBU

[MA19] Barbara McGillivray and Mathias Astell. The relationship between usage and citations in an open access mega-journal. Scientometrics, 121(2):817–838, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03228-3;   http://
link.springer.com/content/pdf/10.1007/s11192-019-
03228-3.pdf.

Maghrebi:2011:CAL

[MAA^{+}11] Morteza Maghrebi, Ali Abbasi, Saeid Amiri, Reza Monsefi, and Ahad Harati. A collective and abridged lexical query for delineation of nanotechnology publications. Scientometrics, 86(1):15–25, January 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0304-7.

Mubin:2017:HUR

[MAA17] Omar Mubin, Abdullah Al Mahmud, and Muneeb Ahmad. HCI down under: reflecting on a decade of the OzCHI conference. Scientometrics, 112(1):367–382, July 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Mubin:2018:TFW

[MAA18] Omar Mubin, Mudassar Arsalan, and Abdullah Al Mahmud. Tracking the follow-up of work in progress papers. Scientometrics, 114(3):1159–1174, March 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2631-4;   https://link.springer.
com/content/pdf/10.1007/s11192-017-2631-4.pdf.

Morillo:2018:HAI

[MÁB18] Fernanda Morillo and Belén Álvarez-Bornstein. How to automatically identify major research sponsors selecting keywords from the WoS Funding Agency field. Scientometrics, 117(3):1755–1770, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2947-8.

Madani:2015:TMB

[Mad15] Farshad Madani. ‘Technology Mining’ bibliometrics analysis: applying network analysis and cluster analysis. Scientometrics, 105(1):323–335, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1685-4.

Magnone:2014:ECT

[Mag14a] Edoardo Magnone. The extreme case of terrorism: a scientometric analysis. Scientometrics, 101(1):179–201, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1378-4.

Magnone:2014:NGR

[Mag14b] Edoardo Magnone. A novel graphical representation of sentence complexity: the description and its application. Scientometrics, 98(2):1301–1329, February 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1074-9.

Morillo:2013:TAA

[MAGAM13] Fernanda Morillo, Javier Aparicio, Borja González-Albo, and Luz Moreno. Towards the automation of address identification. Scientometrics, 94(1):207–224, January 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0733-6.

Moya-Anegon:2013:RGS

[MAGBBM13] Félix Moya-Anegón, Vicente P. Guerrero-Bote, Lutz Bornmann, and Henk F. Moed. The research guarantors of scientific papers and the output counting: a promising new approach. Scientometrics, 97(2):421–434, November 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-1046-0.

Murgado-Armenteros:2015:ACE

[MAGSTRC15] E. M. Murgado-Armenteros, M. Gutiérrez-Salcedo, F. J. Torres-Ruiz, and M. J. Cobo. Analysing the conceptual evolution of qualitative marketing research through science mapping analysis. Scientometrics, 102(1):519–557, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1443-z.

Manh:2015:SPV

[Man15] Ho Dung Manh. Scientific publications in Vietnam as seen from Scopus during 1996–2013. Scientometrics, 105 (1):83–95, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1655-x.

Moed:2013:SSM

[MAP13] Henk F. Moed, M’hamed Aisati, and Andrew Plume. Studying scientific migration in Scopus. Scientometrics, 94(3): 929–942, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0783-9.

Martinez:2011:PFW

[Mar11] Catalina Martínez. Patent families: When do different definitions really matter? Scientometrics, 86(1):39–63, January 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0251-3.

Martin-Alcazar:2019:ASC

[MARMSG19] Fernando Martín-Alcázar, Marta Ruiz-Martínez, and Gonzalo Sánchez-Gardey. Assessing social capital in academic research teams: a measurement instrument proposal. Scientometrics, 121(2):917–935, November 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03212-x.

Mendez:2014:TEM

[MASM14] David I. Méndez, M. Ángeles Alcaraz, and Françoise Salager-Meyer. Titles in English-medium astrophysics research articles. Scientometrics, 98(3):2331–2351, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-013-1174-6. See erratum

[MASM16].

Mendez:2016:ETE

[MASM16] David I. Méndez, M. Ángeles Alcaraz, and Françoise Salager-Meyer. Erratum to: Titles in English-medium astrophysics research articles. Scientometrics, 108(2):1003, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s11192-015-1555-0.pdf. See

[MASM14].

Matthews:2012:SAU

[Mat12] Alan Peter Matthews. South African universities in world rankings. Scientometrics, 92(3):675–695, September 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0611-7.

Matthews:2013:PPP

[Mat13] Alan Peter Matthews. Physics publication productivity in South African universities. Scientometrics, 95(1):69–86, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0842-2.

Marx:2010:HAD

[MB10a] Werner Marx and Lutz Bornmann. How accurately does Thomas Kuhn’s model of paradigm change describe the transition from the static view of the universe to the big bang theory in cosmology? Scientometrics, 84(2):441–464, August 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0107-x.

Mauleon:2010:MFI

[MB10b] Elba Mauleón and María Bordons. Male and female involvement in patenting activity in Spain. Scientometrics, 83(3):605–621, June 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0131-x.

Marx:2013:EPT

[MB13] Werner Marx and Lutz Bornmann. The emergence of plate tectonics and the Kuhnian model of paradigm shift: a bibliometric case study based on the Anna Karenina principle. Scientometrics, 94(2):595–614, February 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0741-6.

Marx:2014:TOS

[MB14] Werner Marx and Lutz Bornmann. Tracing the origin of a scientific legend by reference publication year spectroscopy (RPYS): the legend of the Darwin finches. Scientometrics, 99(3):839–844, June 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1200-8.

Marx:2015:CSS

[MB15] Werner Marx and Lutz Bornmann. On the causes of subject-specific citation rates in Web of Science. Scientometrics, 102(2):1823–1827, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1499-9.

Marx:2016:CPB

[MB16a] Werner Marx and Lutz Bornmann. Change of perspective: bibliometrics from the point of view of cited references — a literature overview on approaches to the evaluation of cited references in bibliometrics. Scientometrics, 109(2):1397–1415, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-2111-2.pdf.

Mirnezami:2016:EHR

[MB16b] Seyed Reza Mirnezami and Catherine Beaudry. The effect of holding a research chair on scientists’ productivity. Scientometrics, 107(2):399–454, May 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1848-y.

Maltseva:2019:SNA

[MB19] Daria Maltseva and Vladimir Batagelj. Social network analysis as a field of invasions: bibliographic approach to study SNA development. Scientometrics, 121(2):1085– 1128, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03193-x.

Mas-Bleda:2013:CPW

[MBA13] Amalia Más-Bleda and Isidro F. Aguillo. Can a personal website be useful as an information source to assess individual scientists? The case of European highly cited researchers. Scientometrics, 96(1):51–67, July 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0952-5.

Markusova:2018:UMV

[MBL18] Valentina Markusova, Valentin Bogorov, and Alexander Libkind. Usage metrics vs classical metrics: analysis of Russia’s research output. Scientometrics, 114(2):593–603, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2597-2.

Maciel:2019:TRI

[MBP19] R. Fileto Maciel, P. Saskia Bayerl, and Marta Macedo Kerr Pinheiro. Technical research innovations of the US national security system. Scientometrics, 120(2):539–565, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03148-2.

Mavros:2013:CNC

[MBR^{+}13] Michael N. Mavros, Vangelis Bardakas, Petros I. Rafailidis, Thalia A. Sardi, Elena Demetriou, and Matthew E. Falagas. Comparison of number of citations to full original articles versus brief reports. Scientometrics, 94(1):203–206, January 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0752-3.

Meyer-Brotz:2017:EEP

[MBSB17] Fabian Meyer-Brötz, Edgar Schiebel, and Leo Brecht. Experimental evaluation of parameter settings in calculation of hybrid similarities: effects of first- and second-order similarity, edge cutting, and weighting factors. Scientometrics, 111(3):1307–1325, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Mas-Bleda:2016:CAI

[MBT16] Amalia Mas-Bleda and Mike Thelwall. Can alternative indicators overcome language biases in citation counts? A comparison of Spanish and UK research. Scientometrics, 109(3):2007–2030, December 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2118-8.

Mas-Bleda:2014:DHC

[MBTKA14] Amalia Mas-Bleda, Mike Thelwall, Kayvan Kousha, and Isidro F. Aguillo. Do highly cited researchers successfully use the social web? Scientometrics, 101(1):337–356, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1345-0.

Macias-Chapula:2010:ILR

[MC10] C. A. Macías-Chapula. Influence of local and regional publications in the production of public health research papers in Latin America. Scientometrics, 84(3):703–716, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0153-4.

Miyairi:2012:BCH

[MC12] Nobuko Miyairi and Han-Wen Chang. Bibliometric characteristics of highly cited papers from Taiwan, 2000– 2009. Scientometrics, 92(1):197–205, July 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0722-9.

Macias-Chapula:2013:CAH

[MC13] César A. Macías-Chapula. Comparative analysis of health public policy research results among Mexico, Chile and Argentina. Scientometrics, 95(2):615–628, May 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0855-x.

Marroquin:2015:EWT

[MC15] Andrés Marroquín and Julio H. Cole. Economical writing (or, “Think Hemingway”). Scientometrics, 103(1):251–259, April 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1522-1.

Morillo:2015:HCG

[MCB15] Fernanda Morillo, Rodrigo Costas, and María Bordons. How is credit given to networking centres in their publications? A case study of the Spanish CIBER research structures. Scientometrics, 103(3):923–938, June 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1564-z.

McCain:2014:AOI

[McC14] Katherine W. McCain. Assessing obliteration by incorporation in a full-text database: JSTOR, economics, and the concept of “bounded rationality”. Scientometrics, 101(2): 1445–1459, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1237-3.

McCain:2018:BGC

[McC18] Katherine W. McCain. Beyond Garfield’s Citation Index: an assessment of some issues in building a personal name Acknowledgments Index. Scientometrics, 114(2):605–631, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2598-1.

Merigo:2016:ARI

[MCCU16] José M. Merigó, Christian A. Cancino, Freddy Coronado, and David Urbano. Academic research in innovation: a country analysis. Scientometrics, 108(2):559–593, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1984-4.

Mendlowicz:2011:TGG

[MCL^{+}11] Mauro Vitor Mendlowicz, Evandro Silva Freire Coutinho, Jerson Laks, Leonardo Franklin Fontenelle, Alexandre Martins Valença, William Berger, Ivan Figueira, and Gláucia Azambuja de Aguiar. Is there a ‘gender gap’ in authorship of the main Brazilian psychiatric journals at the beginning of the 21st century? Scientometrics, 86(1):27–37, January 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0296-3.

Miller:2013:ETF

[MCL13] J. Corey Miller, Keith H. Coble, and Jayson L. Lusk. Evaluating top faculty researchers and the incentives that motivate them. Scientometrics, 97(3):519–533, December 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-0987-7.

Mao:2017:TSC

[MCLL17] Jin Mao, Yujie Cao, Kun Lu, and Gang Li. Topic scientific community in science: a combined perspective of scientific collaboration and topics. Scientometrics, 112(2):851–875, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2418-7.

Moed:2012:CBM

[MCR^{+}12] Henk F. Moed, Lisa Colledge, Jan Reedijk, Felix Moya-Anegon, Vicente Guerrero-Bote, Andrew Plume, and Mayur Amin. Citation-based metrics are appropriate tools in journal assessment provided that they are accurate and used in an informed way. Scientometrics, 92(2):367–376, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0679-8.

Mehmood:2016:PGS

[MCvFP16] Arif Mehmood, Gyu Sang Choi, Otto F. von Feigenblatt, and Han Woo Park. Proving ground for social network analysis in the emerging research area “Internet of Things” (IoT). Scientometrics, 109(1):185–201, October 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-1931-4.

Mutz:2012:GPS

[MD12] Rüdiger Mutz and Hans-Dieter Daniel. The generalized propensity score methodology for estimating unbiased journal impact factors. Scientometrics, 92(2):377–390, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0670-4.

Monroy:2018:TSB

[MD18] Sonia E. Monroy and Hernando Diaz. Time series-based bibliometric analysis of the dynamics of scientific production. Scientometrics, 115(3):1139–1159, June 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2728-4.

Maia:2019:MLC

[MdBdP^{+}19] Saulo Cardoso Maia, Gideon Carvalho de Benedicto, José Willer do Prado, David Alastair Robb, Oscar Neto de Almeida Bispo, and Mozar José de Brito. Mapping the literature on credit unions: a bibliometric investigation grounded in Scopus and Web of Science. Scientometrics, 120(3):929–960, September 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03165-1.

Marzi:2017:PPI

[MDDG17] Giacomo Marzi, Marina Dabić, Tugrul Daim, and Edwin Garces. Product and process innovation in manufacturing firms: a 30-year bibliometric analysis. Scientometrics, 113(2):673–704, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2500-1.

Milanez:2014:PNA

[MdFdA^{+}14] Douglas Henrique Milanez, Leandro Innocentini Lopes de Faria, Roniberto Morato do Amaral, Daniel Rodrigo Leiva, and José Angelo Rodrigues Gregolin. Patents in nanotechnology: an analysis using macro-indicators and forecasting curves. Scientometrics, 101(2):1097–1112, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1244-4.

Morillo:2014:DNC

[MDFGAM14] Fernanda Morillo, Adrián A. Díaz-Faes, Borja González-Albo, and Luz Moreno. Do networking centres perform better? An exploratory analysis in psychiatry and Gastroenterology/ hepatology in Spain. Scientometrics, 98(2): 1401–1416, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1183-5.

Meyer:2010:CAS

[MDG10] M. Meyer, K. Debackere, and W. Glänzel. Can applied science be ‘good science’? Exploring the relationship between patent citations and citation impact in nanoscience. Scientometrics, 85(2):527–539, November 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0154-3.

Montefusco:2019:IIA

[MdNS^{+}19] Adilson Marcos Montefusco, Felipe Parra do Nascimento, Luiz Ubirajara Sennes, Ricardo Ferreira Bento, and Rui Imamura. Influence of international authorship on citations in Brazilian medical journals: a bibliometric analysis. Scientometrics, 119(3):1487–1496, June 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03104-0.

Medina:2015:ORM

[Med15] Fida Medina. The output of researchers in Morocco compared to some North African countries from 1996 to 2012, and its relationship to governmental major decisions on higher education and scientific research. Scientometrics, 105 (1):367–384, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1701-8.

Medina:2018:WDE

[Med18] Anderson Matos Medina. Why do ecologists search for coauthorships? Patterns of co-authorship networks in ecology (1977–2016). Scientometrics, 116(3):1853–1865, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2835-2.

Megnigbeto:2013:ICS

[Mêg13a] Eustache Mêgnigbêto. International collaboration in scientific publishing: the case of West Africa (2001–2010). Scientometrics, 96(3):761–783, September 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0963-2.

Megnigbeto:2013:SPB

[Mêg13b] Eustache Mêgnigbêto. Scientific publishing in Benin as seen from Scopus. Scientometrics, 94(3):911–928, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0843-1.

Megnigbeto:2013:SPW

[Mêg13c] Eustache Mêgnigbêto. Scientific publishing in West Africa: comparing Benin with Ghana and Senegal. Scientometrics, 95(3):1113–1139, June 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0948-6.

Morillo:2015:BAT

[MEG15] Fernanda Morillo and Preiddy Efrain-Garcia. A bibliometric analysis of technology centres. Scientometrics, 104(3): 685–713, September 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1631-5.

Meho:2007:RRC

[Meh07] Lokman I. Meho. The rise and rise of citation analysis. Physics World, 20(1):32–36, January 2007. CO-DEN PHWOEW. ISSN 0953-8585. URL http://

physicsworldarchive.iop.org/full/pwa-pdf/20/1/phwv20i1a33.
pdf.

Meng:2018:GDP

[Men18] Yu Meng. Gender distinctions in patenting: Does nanotechnology make a difference? Scientometrics, 114(3):971–992, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2607-4.

Messinis:2011:TCC

[Mes11] George Messinis. Triadic citations, country biases and patent value: the case of pharmaceuticals. Scientometrics, 89(3):813–833, December 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
011-0473-z.

Michels:2014:SAC

[MF14] Carolin Michels and Jun-Ying Fu. Systematic analysis of coverage and usage of conference proceedings in Web of Science. Scientometrics, 100(2):307–327, August 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1309-4.

Mrowinski:2016:RTP

[MFF^{+}16] Maciej J. Mrowinski, Agata Fronczak, Piotr Fronczak, Olgica Nedic, and Marcel Ausloos. Review time in peer review: quantitative analysis and modelling of editorial workflows. Scientometrics, 107(1):271–286, April 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1871-z;   http://link.
springer.com/content/pdf/10.1007/s11192-016-1871-
z.pdf.

Moehrle:2012:MTP

[MG12] Martin G. Moehrle and Jan M. Gerken. Measuring textual patent similarity on the basis of combined concepts: design decisions and their consequences. Scientometrics, 91 (3):805–826, June 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0682-0.

Miramontes:2016:EEC

[MGB16] Jorge Rodriguez Miramontes and C. N. Gonzalez-Brambila. The effects of external collaboration on research output in engineering. Scientometrics, 109(2):661–675, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2054-7.

Miranda:2019:CSD

[MGC19] Ruben Miranda and Esther Garcia-Carpintero. Comparison of the share of documents and citations from different quartile journals in 25 research areas. Scientometrics, 121 (1):479–501, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03210-z.

Moliner:2017:USC

[MGGdP17] Liliana Arroyo Moliner, Eva Gallardo-Gallardo, and Pedro Gallo de Puelles. Understanding scientific communities: a social network approach to collaborations in talent management research. Scientometrics, 113(3):1439–1462, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2537-1.

Martins:2010:AQS

[MGLZ10] Waister Silva Martins, Marcos André Gonçalves, Alberto H. F. Laender, and Nivio Ziviani. Assessing the quality of scientific conferences based on bibliographic citations. Scientometrics, 83(1):133–155, April 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0078-y.

Maisonobe:2017:GGS

[MGM^{+}17] Marion Maisonobe, Michel Grossetti, Béatrice Milard, Laurent Jégou, and Denis Eckert. The global geography of scientific visibility: a deconcentration process (1999– 2011). Scientometrics, 113(1):479–493, October 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2463-2.

Meyer:2014:THI

[MGMW14] Martin Meyer, Kevin Grant, Piera Morlacchi, and Dagmara Weckowska. Triple helix indicators as an emergent area of enquiry: a bibliometric perspective. Scientometrics, 99 (1):151–174, April 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1103-8.

McLevey:2018:IID

[MGMY^{+}18] John McLevey, Alexander V. Graham, Reid McIlroy-Young, Pierson Browne, and Kathryn S. Plaisance. Interdisciplinarity and insularity in the diffusion of knowledge: an analysis of disciplinary boundaries between philosophy of science and the sciences. Scientometrics, 117(1):331–349, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2866-8.

Machado:2014:DFC

[MGT14] J. A. Tenreiro Machado, Alexandra M. S. F. Galhano, and Juan J. Trujillo. On development of fractional calculus during the last fifty years. Scientometrics, 98(1):577–582, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1032-6.

Moed:2014:BAT

[MH14] Henk F. Moed and Gali Halevi. A bibliometric approach to tracking international scientific migration. Scientometrics, 101(3):1987–2001, December 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1307-6.

Maity:2015:RTL

[MH15] Bapan Kumar Maity and Sudip Ranjan Hatua. Research trends of library management in LIS in India since 1950– 2012. Scientometrics, 105(1):337–346, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1673-8.

Ma:2016:CEL

[MH16a] Rongzhen Ma and Yuh-Shan Ho. Comparison of environmental laws publications in Science Citation Index Expanded and Social Science Index: a bibliometric analysis. Scientometrics, 109(1):227–239, October 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2010-6.

Maity:2016:DME

[MH16b] Bapan Kumar Maity and Sudip Ranjan Hatua. Designing a model to evaluate scholarly publications with special reference to social sciences in India. Scientometrics, 109(3):2031–2048, December 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2137-5.

Moksony:2014:RRS

[MHC14] Ferenc Moksony, Rita Hegedűs, and Melinda Császár. Rankings, research styles, and publication cultures: a study of American sociology departments. Scientometrics, 101(3): 1715–1729, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1218-y.

Martinez:2015:CHC

[MHC^{+}15] M. A. Martínez, M. Herrera, E. Contreras, A. Ruíz, and E. Herrera-Viedma. Characterizing highly cited papers in social work through H-Classics. Scientometrics, 102(2): 1713–1729, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1460-y.

Marx:2017:SRU

[MHFB17] Werner Marx, Robin Haunschild, Bernie French, and Lutz Bornmann. Slow reception and under-citedness in climate change research: A case study of Charles David Keeling, discoverer of the risk of global warming. Scientometrics, 112(2):1079–1092, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2405-z;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2405-
z.pdf.

Mryglod:2016:QES

[MHKB16] O. Mryglod, Yu. Holovatch, R. Kenna, and B. Berche. Quantifying the evolution of a scientific topic: reaction of the academic community to the Chornobyl disaster. Scientometrics, 106(3):1151–1166, March 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1820-2.

Martinez:2014:CCC

[MHLGHV14] M. A. Martínez, M. Herrera, J. López-Gijón, and E. Herrera-Viedma. H-Classics: characterizing the concept of citation classics through H-index. Scientometrics, 98(3): 1971–1983, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1155-9.

Mehta:2012:GGN

[MHM^{+}12a] Aashish Mehta, Patrick Herron, Yasuyuki Motoyama, Richard Appelbaum, and Timothy Lenoir. Globalization and de-globalization in nanotechnology research: the role of China. Scientometrics, 93(2):439–458, November 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0687-8.

Mryglod:2012:EPS

[MHM12b] O. Mryglod, Yu. Holovatch, and I. Mryglod. Editorial process in scientific journals: analysis and modeling. Scientometrics, 91(1):101–112, April 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0536-1.

Mauleon:2013:AGB

[MHM^{+}13] Elba Mauleón, Laura Hillán, Luz Moreno, Isabel Gómez, and María Bordons. Assessing gender balance among journal authors and editorial board members. Scientometrics, 95(1):87–114, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0824-4.

Marx:2017:WEW

[MHTB17] Werner Marx, Robin Haunschild, Andreas Thor, and Lutz Bornmann. Which early works are cited most frequently in climate change research literature? A bibliometric approach based on Reference Publication Year Spectroscopy. Scientometrics, 110(1):335–353, January 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s11192-016-2177-x.pdf.

Mikki:2010:CGS

[Mik10] Susanne Mikki. Comparing Google Scholar and ISI Web of Science for earth sciences. Scientometrics, 82(2):321–331, February 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0038-6.

Mikki:2017:SPB

[Mik17] Susanne Mikki. Scholarly publications beyond pay-walls: increased citation advantage for open publishing. Scientometrics, 113(3):1529–1538, December 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2554-0.

Mixon:2018:DAS

[Mix18] Franklin G. Mixon, Jr. Do academics swing for the fences after tenure? Analysis of attributions data from economics research. Scientometrics, 116(3):2155–2160, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2819-2.

Martinez:2012:RAC

[MJC12] H. Martinez, A. Jaime, and J. Camacho. Relative absorptive capacity: a research profiling. Scientometrics, 92(3): 657–674, September 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0652-6.

Martinez:2014:BPA

[MJC14] H. Martinez, A. Jaime, and J. Camacho. Biotechnology profile analysis in Colombia. Scientometrics, 101(3):1789– 1804, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1408-2.

McCarty:2013:PAI

[MJHG13] Christopher McCarty, James W. Jawitz, Allison Hopkins, and Alex Goldman. Predicting author h-index using characteristics of the co-author network. Scientometrics, 96(2): 467–483, August 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0933-0.

Mejia:2018:UAD

[MK18] Cristian Mejia and Yuya Kajikawa. Using acknowledgement data to characterize funding organizations by the types of research sponsored: the case of robotics research. Scientometrics, 114(3):883–904, March 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2617-2.

Mizoguchi:2019:CAC

[MK19] Hirokuni Mizoguchi and Shingo Kano. Comparative analysis of correlations of research and development indicators for rare diseases among Japan, the US, and Europe. Scientometrics, 120(2):361–374, August 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03129-5.

Mustafee:2014:EMS

[MKF14] Navonil Mustafee, Korina Katsaliaki, and Paul Fishwick. Exploring the modelling and simulation knowledge base through journal co-citation analysis. Scientometrics, 98(3): 2145–2159, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1136-z.

Mryglod:2013:ASM

[MKHB13a] O. Mryglod, R. Kenna, Yu. Holovatch, and B. Berche. Absolute and specific measures of research group excellence. Scientometrics, 95(1):115–127, April 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0874-7.

Mryglod:2013:CCB

[MKHB13b] O. Mryglod, R. Kenna, Yu. Holovatch, and B. Berche. Comparison of a citation-based indicator and peer review for absolute and specific measures of research-group excellence. Scientometrics, 97(3):767–777, December 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1058-9.

Mryglod:2015:PRRa

[MKHB15a] O. Mryglod, R. Kenna, Yu. Holovatch, and B. Berche. Predicting results of the Research Excellence Framework using departmental h-index. Scientometrics, 102(3):2165–2180, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1512-3. See [MKHB15b].

Mryglod:2015:PRRb

[MKHB15b] O. Mryglod, R. Kenna, Yu. Holovatch, and B. Berche. Predicting results of the research excellence framework using departmental h-index: revisited. Scientometrics, 104(3):1013–1017, September 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-

015-1567-9. See [MKHB15a].

Mugabushaka:2016:BII

[MKP16] Alexis-Michel Mugabushaka, Anthi Kyriakou, and Theo Papazoglou. Bibliometric indicators of interdisciplinarity: the potential of the Leinster–Cobbold diversity indices to study disciplinary diversity. Scientometrics, 107(2):593–607, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1865-x.

Mohabbati-Kalejahi:2017:SSS

[MKYM^{+}17] Nasrin Mohabbati-Kalejahi, Mohammad Ali Alamdar Yazdi, Fadel M. Megahed, Sydney Y. Schaefer, Lara A. Boyd, Catherine E. Lang, and Keith R. Lohse. Streamlining science with structured data archives: insights from stroke rehabilitation. Scientometrics, 113(2):969–983, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2482-z.

Mingers:2010:CCC

[ML10] John Mingers and Evangelia A. E. C. G. Lipitakis. Counting the citations: a comparison of Web of Science and Google Scholar in the field of business and management. Scientometrics, 85(2):613–625, November 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0270-0.

Milojevic:2013:IMI

[ML13] Stasa Milojević and Loet Leydesdorff. Information metrics (iMetrics): a research specialty with a socio-cognitive identity? Scientometrics, 95(1):141–157, April 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0861-z.

Ma:2016:ERF

[ML16] Vincent C. Ma and John S. Liu. Exploring the research fronts and main paths of literature: a case study of shareholder activism research. Scientometrics, 109(1):33–52, October 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2035-x.

Ma:2018:GSC

[ML18] Qian Ma and Wenlan Li. Growing scientific collaboration between Hong Kong and Mainland China since the handover: a 20-year bibliometric analysis. Scientometrics, 117(3):1479–1491, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2916-2.

Ma:2014:SCF

[MLC14] Feicheng Ma, Yating Li, and Baitong Chen. Study of the collaboration in the field of the Chinese humanities and social sciences. Scientometrics, 100(2):439–458, August 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1301-z.

Ma:2018:TBR

[MLOY18] Tingcan Ma, Ruinan Li, Guiyan Ou, and Mingliang Yue. Topic based research competitiveness evaluation. Scientometrics, 117(2):789–803, November 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2891-7.

Meyer:2014:OEE

[MLT^{+}14] M. Meyer, D. Libaers, B. Thijs, K. Grant, W. Glänzel, and K. Debackere. Origin and emergence of entrepreneurship as a research field. Scientometrics, 98(1):473–485, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1021-9.

Meng:2015:MVC

[MLT^{+}15] Xiangfeng Meng, Xinhai Liu, YunHai Tong, Wolfgang Glänzel, and Shaohua Tan. Multi-view clustering with exemplars for scientific mapping. Scientometrics, 105(3):1527– 1552, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1682-7.

Markusova:2012:RPC

[MLVJ12] V. A. Markusova, A. N. Libkind, A. E. Varshavsly, and C. N. M. Jansz. Research performance and collaboration in the Novosibirsk region. Scientometrics, 91(2):513–526, May 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0597-1.

Ma:2014:PTK

[MLY^{+}14] Fei-Cheng Ma, Peng-Hui Lyu, Qiang Yao, Lan Yao, and Shi-Jing Zhang. Publication trends and knowledge maps of global translational medicine research. Scientometrics, 98 (1):221–246, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1003-y.

Mallik:2014:BAG

[MM14a] Ajoy Mallik and Nripendranath Mandal. Bibliometric analysis of global publication output and collaboration structure study in microRNA research. Scientometrics, 98(3): 2011–2037, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1128-z.

Manisha:2014:BCC

[MM14b] Manjari Manisha and G. Mahesh. Bibliometric characteristics of champion works of China and India. Scientometrics, 98(2):1101–1111, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1093-6.

Maraut:2014:IAI

[MM14c] Stéphane Maraut and Catalina Martínez. Identifying author-inventors from Spain: methods and a first insight into results. Scientometrics, 101(1):445–476, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-014-1409-1.

Moeller:2015:CKP

[MM15a] Ansgar Moeller and Martin G. Moehrle. Completing keyword patent search with semantic patent search: introducing a semiautomatic iterative method for patent near search based on semantic similarities. Scientometrics, 102 (1):77–96, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1446-9.

Mutschke:2015:SMS

[MM15b] Peter Mutschke and Philipp Mayr. Science models for search: a study on combining scholarly information retrieval and scientometrics. Scientometrics, 102(3):2323– 2345, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1485-2.

Makkonen:2016:SCB

[MM16] Teemu Makkonen and Timo Mitze. Scientific collaboration between ‘old’ and ‘new’ member states: Did joining the European Union make a difference? Scientometrics, 106(3):1193–1215, March 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com//content/
pdf/10.1007/s11192-015-1824-y.pdf;   http://link.
springer.com/content/pdf/10.1007/s11192-015-1824-
y.pdf.

Mingers:2017:ENG

[MM17a] John Mingers and Martin Meyer. Erratum to: Normalizing Google Scholar data for use in research evaluation. Scientometrics, 112(2):1123–1124, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2424-9;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2424-

9.pdf. See [MM17b].

Mingers:2017:NGS

[MM17b] John Mingers and Martin Meyer. Normalizing Google Scholar data for use in research evaluation. Scientometrics, 112(2):1111–1121, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
017-2415-x;   http://link.springer.com/content/pdf/
10.1007/s11192-017-2415-x.pdf. See erratum [MM17a].

Marini:2018:TWG

[MM18] Giulio Marini and Viviana Meschitti. The trench warfare of gender discrimination: evidence from academic promotions to full professor in Italy. Scientometrics, 115(2):989–1006, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-018-2696-8. See correc-

tion [MM19].

Marini:2019:CTW

[MM19] Giulio Marini and Viviana Meschitti. Correction to: The trench warfare of gender discrimination: evidence from academic promotions to full professor in Italy. Scientometrics, 118(2):719, February 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2983-4;
http://link.springer.com/content/pdf/10.1007/s11192-
018-2983-4.pdf. See [MM18].

Moed:2018:TRR

[MMA18] Henk F. Moed, Valentina Markusova, and Mark Akoev. Trends in Russian research output indexed in Scopus and Web of Science. Scientometrics, 116(2):1153–1180, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2769-8.

Miguel:2010:ISE

[MMAHS10] Sandra Miguel, Félix Moya-Anegón, and Víctor Herrero-Solana. The impact of the socio-economic crisis of 2001 on the scientific system of Argentina from the scientometric perspective. Scientometrics, 85(2):495–507, November 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0266-9.

Martin-Martin:2016:BPS

[MMOMALC16] Alberto Martín-Martín, Enrique Orduna-Malea, Juan M. Ayllón, and Emilio Delgado López-Cózar. Back to the past: on the shoulders of an academic search engine giant. Scientometrics, 107(3):1477–1487, June 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1917-2.

Martin-Martin:2018:CHC

[MMOMLC18a] Alberto Martín-Martín, Enrique Orduna-Malea, and Emilio Delgado López-Cózar. Coverage of highly-cited documents in Google Scholar, Web of Science, and Scopus: a multidisciplinary comparison. Scientometrics, 116(3):2175– 2188, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2820-9.

Martin-Martin:2018:NMD

[MMOMLC18b] Alberto Martín-Martín, Enrique Orduna-Malea, and Emilio Delgado López-Cózar. A novel method for depicting academic disciplines through Google Scholar Citations: The case of bibliometrics. Scientometrics, 114(3):1251–1273, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2587-4.

Mutschke:2011:SMV

[MMSS11] Peter Mutschke, Philipp Mayr, Philipp Schaer, and York Sure. Science models as value-added services for scholarly information systems. Scientometrics, 89(1):349–364, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0430-x.

Maghami:2015:QQA

[MnaeR^{+}15] Mohammad Reza Maghami, Shahin navabi asl, Mohammad esmaeil Rezadad, Nader Ale Ebrahim, and Chandima Gomes. Qualitative and quantitative analysis of solar hydrogen generation literature from 2001 to 2014. Scientometrics, 105(2):759–771, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1730-3;   http://link.
springer.com/content/pdf/10.1007/s11192-015-1730-
3.pdf.

Milanez:2016:DPR

[MNdF16] Douglas Henrique Milanez, Ed Noyons, and Leandro Innocentini Lopes de Faria. A delineating procedure to retrieve relevant publication data in research areas: the case of nanocellulose. Scientometrics, 107(2):627–643, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-016-1922-5.

Maluleka:2016:FIR

[MOA16a] Jan Resenga Maluleka, Omwoyo Bosire Onyancha, and Isola Ajiferuke. Factors influencing research collaboration in LIS schools in South Africa. Scientometrics, 107(2):337–355, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1846-0.

Masterton:2016:LVH

[MOA16b] George Masterton, Erik J. Olsson, and Staffan Angere. Linking as voting: how the Condorcet jury theorem in political science is relevant to webometrics. Scientometrics, 106 (3):945–966, March 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1837-1.

Moehrle:2010:MTP

[Moe10] Martin G. Moehrle. Measures for textual patent similarities: a guided way to select appropriate approaches. Scientometrics, 85(1):95–109, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0243-3.

Moed:2016:CIC

[Moe16a] Henk F. Moed. Comprehensive indicator comparisons intelligible to non-experts: the case of two SNIP versions. Scientometrics, 106(1):51–65, January 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1781-5.

Moed:2016:ISD

[Moe16b] Henk F. Moed. Iran’s scientific dominance and the emergence of South-East Asian countries as scientific collaborators in the Persian Gulf Region. Scientometrics, 108(1):305– 314, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1946-x.

Moed:2017:CCA

[Moe17] Henk F. Moed. A critical comparative analysis of five world university rankings. Scientometrics, 110(2):967–990, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2212-y.

Mohammadi:2012:KMI

[Moh12] Ehsan Mohammadi. Knowledge mapping of the Iranian nanoscience and technology: a text mining approach. Scientometrics, 92(3):593–608, September 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0644-6.

Mooij:2015:EPF

[Moo15] Ton Mooij. Exploring a prototype framework of web-based and peer-reviewed “European Educational Research Quality Indicators” (EERQI). Scientometrics, 102(1):1037–1055, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1429-x.

Mingers:2017:UGS

[MOO17] John Mingers, Jesse R. O’Hanley, and Musbaudeen Okunola. Using Google Scholar institutional level data to evaluate the quality of university research. Scientometrics, 113(3):1627–1643, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
017-2532-6;   http://link.springer.com/content/pdf/
10.1007/s11192-017-2532-6.pdf.

Morillo:2016:PPI

[Mor16] Fernanda Morillo. Public-private interactions reflected through the funding acknowledgements. Scientometrics, 108(3):1193–1204, September 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2032-0.

Morillo:2019:CIR

[Mor19] Fernanda Morillo. Collaboration and impact of research in different disciplines with international funding (from the EU and other foreign sources). Scientometrics, 120(2):807–823, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03150-8.

Moustafa:2015:DIA

[Mou15a] Khaled Moustafa. Don’t infer anything from unavailable data. Scientometrics, 105(3):2271–2272, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-015-1750-z.

Moustafa:2015:TBE

[Mou15b] Khaled Moustafa. Is there bias in editorial choice? Yes. Scientometrics, 105(3):2249–2251, December 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1617-3.

Moustafa:2016:PPO

[Mou16] Khaled Moustafa. A proposal for print-online hybrid publishing system. Scientometrics, 108(3):1649–1650, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-1944-z.

Ma:2015:APT

[MP15] Jing Ma and Alan L. Porter. Analyzing patent topical information to identify technology pathways and potential opportunities. Scientometrics, 102(1):811–827, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-014-1392-6.

Mendonca:2018:GAS

[MPF18] Sandro Mendonça, João Pereira, and Manuel Ennes Ferreira. Gatekeeping African studies: what does “editormetrics” indicate about journal governance? Scientometrics, 117(3):1513–1534, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2909-1.

Mongeon:2016:JCW

[MPH16] Philippe Mongeon and Adèle Paul-Hus. The journal coverage of Web of Science and Scopus: a comparative analysis. Scientometrics, 106(1):213–228, January 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1765-5.

Mistele:2019:PAC

[MPH19] Tobias Mistele, Tom Price, and Sabine Hossenfelder. Predicting authors’ citation counts and h-indices with a neural network. Scientometrics, 120(1):87–104, July 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03110-2.

Molleri:2018:TUR

[MPM18] Jefferson Seide Molléri, Kai Petersen, and Emilia Mendes. Towards understanding the relation between citations and research quality in software engineering studies. Scientometrics, 117(3):1453–1478, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2907-3;   http://link.springer.com/content/pdf/
10.1007/s11192-018-2907-3.pdf.

Moskaleva:2018:RIS

[MPS^{+}18] Olga Moskaleva, Vladimir Pislyakov, Ivan Sterligov, Mark Akoev, and Svetlana Shabanova. Russian Index of Science Citation: Overview and review. Scientometrics, 116(1):449– 462, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2758-y.

Ma:2013:QSE

[MPY^{+}13] Zheng Ma, Yuntao Pan, Zhenglu Yu, Jingting Wang, Jia Jia, and Yishan Wu. A quantitative study on the effectiveness of peer review for academic journals. Scientometrics, 95(1):1– 13, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0879-2.

Mahbuba:2010:SRI

[MR10] Dilruba Mahbuba and Ronald Rousseau. Scientific research in the Indian subcontinent: selected trends and indicators 1973–2007 comparing Bangladesh, Pakistan and Sri Lanka with India, the local giant. Scientometrics, 84(2):403–420, August 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0203-y.

Mahbuba:2013:YBT

[MR13] Dilruba Mahbuba and Ronald Rousseau. Year-based h-type indicators. Scientometrics, 96(3):785–797, September 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0934-z. See comments [Egg14b].

Mamun:2015:TFE

[MR15] Shamsul Arifeen Khan Mamun and Mohammad Mafizur Rahman. Is there any feedback effect between academic research publication and research collaboration? Evidence from an Australian university. Scientometrics, 105(3):2179– 2196, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1759-3.

Mayer:2018:HDR

[MR18a] Sabrina J. Mayer and Justus M. K. Rathmann. How does research productivity relate to gender? Analyzing gender differences for multiple publication dimensions. Scientometrics, 117(3):1663–1693, December 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2933-1.

Migheli:2018:MAA

[MR18b] Matteo Migheli and Giovanni Battista Ramello. The market of academic attention. Scientometrics, 114(1):113–133, January 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2564-y.

Manana-Rodriguez:2013:SPS

[MRGT13] Jorge Mañana-Rodríguez and Elea Giménez-Toledo. Scholarly publishing in social sciences and humanities, associated probabilities of belonging and its spectrum: a quantitative approach for the Spanish case. Scientometrics, 94 (3):893–910, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0838-y.

Mannana-Rodriguez:2018:SMS

[MRGT18] Jorge Mannana-Rodriguez and Elea Giménez-Toledo. Specialization and multidisciplinarity of scholarly book publishers: differences between Spanish University Presses and other scholarly publishers. Scientometrics, 114(1):19–30, January 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2563-z.

Mahian:2015:FBE

[MRLW15] Omid Mahian, Mohammad Mehdi Rashidi, Giulio Lorenzini, and Somchai Wongwises. Fame bias in editorial choice: Yes or no? Scientometrics, 105(3):2253–2254, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-015-1636-0.

Morooka:2014:BAI

[MRN14] Kazuko Morooka, Mila M. Ramos, and Fonseca N. Nathaniel. A bibliometric approach to interdisciplinarity in Japanese rice research and technology development. Scientometrics, 98(1):73–98, January 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1119-0.

Muller:2017:DSA

[MRR17] Mark-Christoph Müller, Florian Reitz, and Nicolas Roy. Data sets for author name disambiguation: an empirical analysis and a new resource. Scientometrics, 111(3): 1467–1500, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-017-2363-
5.pdf.

Mimouni:2016:SCR

[MRS^{+}16] Michael Mimouni, Motti Ratmansky, Yaron Sacher, Sharon Aharoni, and Aviva Mimouni-Bloch. Self-citation rate and impact factor in pediatrics. Scientometrics, 108(3):1455–1460, September 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2025-z.

Michels:2012:GSD

[MS12] Carolin Michels and Ulrich Schmoch. The growth of science and database coverage. Scientometrics, 93(3):831–846, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0732-7.

Mishra:2013:MSA

[MS13] Vinod Mishra and Russell Smyth. Are more senior academics really more research productive than junior academics? Evidence from Australian law schools. Scientometrics, 96(2):411–425, August 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0886-3.

Michels:2014:IBS

[MS14] Carolin Michels and Ulrich Schmoch. Impact of bibliometric studies on the publication behaviour of authors. Scientometrics, 98(1):369–385, January 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1015-7.

Mayr:2015:CBI

[MS15a] Philipp Mayr and Andrea Scharnhorst. Combining bibliometrics and information retrieval: preface. Scientometrics, 102(3):2191–2192, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-015-
1529-2.pdf.

Mayr:2015:SIR

[MS15b] Philipp Mayr and Andrea Scharnhorst. Scientometrics and information retrieval: weak-links revitalized. Scientometrics, 102(3):2193–2199, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-014-1484-3.pdf.

Madison:2016:CGS

[MS16a] Guy Madison and Therese Söderlund. Can gender studies be studied? Reply to comments on Söderlund and Madison. Scientometrics, 108(1):329–335, July 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1963-9. See [LSL15, SM15].

Morichika:2016:UDD

[MS16b] Noriyuki Morichika and Sotaro Shibayama. Use of dissertation data in science policy research. Scientometrics, 108 (1):221–241, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1962-x.

Madison:2018:CCS

[MS18a] Guy Madison and Therese Söderlund. Comparisons of content and scientific quality indicators across peer-reviewed journal articles with more or less gender perspective: gender studies can do better. Scientometrics, 115(3):1161–1183, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2729-3;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2729-
3.pdf.

Meschede:2018:CMC

[MS18b] Christine Meschede and Tobias Siebenlist. Cross-metric compatability and inconsistencies of altmetrics. Scientometrics, 115(1):283–297, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2674-1.

Morillo:2013:ANC

[MSA13] Fernanda Morillo, Ignacio Santabárbara, and Javier Aparicio. The automatic normalisation challenge: detailed addresses identification. Scientometrics, 95(3):953–966, June 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-0965-0.

Minniti:2018:MDO

[MSB18] Sergio Minniti, Valeria Santoro, and Simone Belli. Mapping the development of Open Access in Latin America and Caribbean countries. An analysis of Web of Science Core Collection and SciELO Citation Index (2005–2017). Scientometrics, 117(3):1905–1930, December 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2950-0.

Mukherjee:2018:NSC

[MSC18] Bhaskar Mukherjee, Sinisa Subotić, and Ajay Kumar Chaubey. And now for something completely different: the congruence of the Altmetric Attention Score’s structure between different article groups. Scientometrics, 114 (1):253–275, January 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2559-8.

Manganote:2016:EHE

[MSdBC16] Edmilson J. T. Manganote, Peter A. Schulz, and Carlos Henrique de Brito Cruz. Effect of high energy physics large collaborations on higher education institutions citations and rankings. Scientometrics, 109(2):813–826, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2048-5.

Mansourzadeh:2019:NUR

[MSDJ19] Mohammad Javad Mansourzadeh, Behrooz Shahmoradi, Hossein Dehdarirad, and Elmira Janavi. A note on using revealed comparative advantages in scientometrics studies. Scientometrics, 121(1):595–599, October 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03207-8.

Moller:2016:AEG

[MSH16] Torger Möller, Marion Schmidt, and Stefan Hornbostel. Assessing the effects of the German Excellence Initiative with bibliometric methods. Scientometrics, 109(3):2217– 2239, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-016-
2090-3.

Mattsson:2011:CRA

[MSL11] Pauline Mattsson, Carl Johan Sundberg, and Patrice Laget. Is correspondence reflected in the author position? A bibliometric study of the relation between corresponding author and byline position. Scientometrics, 87(1):99–105, April 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0310-9.

Matic:2015:TRT

[MSP^{+}15] Rada Matić, Srdan Stamenković, Zorica Popović, Milena Stefanović, Vera Vidaković, Miroslava Smiljanić, and Srdan Bojović. Tree responses, tolerance and acclimation to stress: Does current research depend on the cultivation status of studied species? Scientometrics, 105(2):1209–1222, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1726-z.

Ma:2012:PWN

[MSYW12] Caifeng Ma, Cheng Su, Junpeng Yuan, and Yishan Wu. Papers written by Nobel Prize winners in physics before they won the prize: an analysis of their language and journal of publication. Scientometrics, 93(3):1151–1163, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0748-z.

Monteleone:2012:GAA

[MT12a] S. Monteleone and B. Torrisi. Geographical analysis of the academic brain drain in Italy. Scientometrics, 93(2): 413–430, November 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0683-z.

Murad:2012:IIR

[MT12b] Abdulkader A. Murad and Dimiter T. Tomov. Institutionalization and internationalization of research on the applications of the geographical information systems in health planning. Scientometrics, 91(1):143–158, April 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0567-7.

Mohammadi:2013:ANS

[MT13a] Ehsan Mohammadi and Mike Thelwall. Assessing nonstandard article impact using F1000 labels. Scientometrics, 97(2):383–395, November 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
013-0993-9.

Moiwo:2013:CDC

[MT13b] Juana Paul Moiwo and Fulu Tao. The changing dynamics in citation index publication position China in a race with the USA for global leadership. Scientometrics, 95(3):1031–1050, June 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0846-y.

Minguillo:2015:WBI

[MT15] David Minguillo and Mike Thelwall. Which are the best innovation support infrastructures for universities? evidence from R&D output and commercial activities. Scientometrics, 102(1):1057–1081, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1458-5.

Mubin:2018:ACA

[MTA^{+}18] Omar Mubin, Dhaval Tejlavwala, Mudassar Arsalan, Muneeb Ahmad, and Simeon Simoff. An assessment into the characteristics of award winning papers at CHI. Scientometrics, 116(2):1181–1201, August 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2778-7.

Minguillo:2015:DSP

[MTT15] David Minguillo, Robert Tijssen, and Mike Thelwall. Do science parks promote research and technology? A scientometric analysis of the UK. Scientometrics, 102(1):701–725, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1435-z.

Mixon:2017:SIT

[MTU17] Franklin G. Mixon, Jr., Benno Torgler, and Kamal P. Upadhyaya. Scholarly impact and the timing of major awards in economics. Scientometrics, 112(3):1837–1852, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2429-4.

Mueller:2016:AFC

[Mue16] Christoph Emanuel Mueller. Accurate forecast of countries’ research output by macro-level indicators. Scientometrics, 109(2):1307–1328, November 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2084-1.

Mueller:2018:WST

[Mue18] Georg P. Mueller. When the search for truth fails: A computer simulation of the impact of the publication bias on the meta-analysis of scientific literature. Scientometrics, 117(3): 2061–2076, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2942-0.

Must:2012:ATE

[Mus12] Ülle Must. Alone or together: examples from history research. Scientometrics, 91(2):527–537, May 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0596-2.

Monsarrat:2019:PSH

[MV19] Paul Monsarrat and Jean-Noel Vergnes. The progressive substitution of hazard ratios for relative risks in biomedical research. Scientometrics, 119(2):1263–1267, May 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03059-2.

Makkonen:2013:BRI

[MvdH13] Teemu Makkonen and Robert P. van der Have. Benchmarking regional innovative performance: composite measures and direct innovation counts. Scientometrics, 94(1): 247–262, January 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0753-2.

Magerman:2010:EFA

[MVS10] Tom Magerman, Bart Van Looy, and Xiaoyan Song. Exploring the feasibility and accuracy of latent semantic analysis based text mining techniques to detect similarity between patent documents and scientific publications. Scientometrics, 82(2):289–306, February 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0046-6.

Ma:2012:JII

[MWDC12] Tingcan Ma, Gui-Fang Wang, Ke Dong, and Mukun Cao. The Journal’s integrated impact index: a new indicator for journal evaluation. Scientometrics, 90(2):649–658, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0538-z.

Mamtora:2014:ESR

[MWH14] Jayshree Mamtora, Jacqueline K. Wolstenholme, and Gaby Haddow. Environmental sciences research in northern Australia, 2000–2011: a bibliometric analysis within the context of a national research assessment exercise. Scientometrics, 98(1):265–281, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
013-1037-1.

Ma:2018:AIC

[MXZ18] Shutian Ma, Jin Xu, and Chengzhi Zhang. Automatic identification of cited text spans: a multi-classifier approach over imbalanced dataset. Scientometrics, 116(2):1303–1330, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2754-2.

Ma:2016:UIS

[MY16] Ruimin Ma and Erjia Yan. Uncovering inter-specialty knowledge communication using author citation networks. Scientometrics, 109(2):839–854, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2091-2.

Mesdaghinia:2015:AMR

[MYN^{+}15] Alireza Mesdaghinia, Masoud Younesian, Simin Nasseri, Ramin Nabizadeh Nodehi, and Mahdi Hadi. Analysis of the microbial risk assessment studies from 1973 to 2015: a bibliometric case study. Scientometrics, 105(1):691–707, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1692-5.

Mun:2019:SDT

[MYP19] Changbae Mun, Sejun Yoon, and Hyunseok Park. Structural decomposition of technological domain using patent co-classification and classification hierarchy. Scientometrics, 121(2):633–652, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03223-8.

Yan:2015:ACS

[mYqS15] Su mei Yan and Ji qing Sun. Assessing China’s salt lake resources R&D based on bibliometrics analysis. Scientometrics, 105(2):1141–1155, November 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1721-4.

Michayluk:2014:DLA

[MZ14] David Michayluk and Ralf Zurbruegg. Do lead articles signal higher quality in the digital age? Evidence from finance journals. Scientometrics, 98(2):961–973, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-1115-4.

Montazerian:2019:NPN

[MZE19] Maziar Montazerian, Edgar Dutra Zanotto, and Hellmut Eckert. A new parameter for (normalized) evaluation of Hindex: countries as a case study. Scientometrics, 118(3): 1065–1078, March 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2996-z.

Noruzi:2012:MIP

[NA12] Alireza Noruzi and Mohammadhiwa Abdekhoda. Mapping Iranian patents based on international patent classification (IPC), from 1976 to 2011. Scientometrics, 93(3): 847–856, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0743-4.

Navarrete:2014:RPS

[NA14] Ian A. Navarrete and Victor B. Asio. Research productivity in soil science in the Philippines. Scientometrics, 100 (1):261–272, July 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1202-6.

Ni:2018:RBI

[NA18] Ping Ni and Xinying An. Relationship between international collaboration papers and their citations from an economic perspective. Scientometrics, 116(2):863–877, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2784-9.

Narin:2012:DPP

[Nar12] Francis Narin. Decades of progress, or the progress of decades? Scientometrics, 92(2):391–393, August 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0678-9.

Nasir:2011:TAI

[NASR11] Anthony Nasir, Tariq Mahmood Ali, Sheikh Shahdin, and Tariq Ur Rahman. Technology achievement index 2009: ranking and comparative study of nations. Scientometrics, 87(1):41–62, April 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0285-6.

Nikolic:2011:BAD

[NBR^{+}11] N. Nikolic, J.-L. Baglinière, C. Rigaud, C. Gardes, M. L. Masquilier, and C. Taverny. Bibliometric analysis of diadromous fish research from 1970s to 2010: a case study of seven species. Scientometrics, 88(3):929–947, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0422-x.

Nuzzolese:2019:DAW

[NCG^{+}19] Andrea Giovanni Nuzzolese, Paolo Ciancarini, Aldo Gangemi, Silvio Peroni, Francesco Poggi, and Valentina Presutti. Do altmetrics work for assessing research quality? Scientometrics, 118(2):539–562, February 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2988-z.

Nedic:2016:PCP

[ND16] Olgica Nedić and Aleksandar Dekanski. Priority criteria in peer review of scientific articles. Scientometrics, 107(1):15– 26, April 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1869-6.

Neuhausler:2013:PFM

[NF13] Peter Neuhäusler and Rainer Frietsch. Patent families as macro level patent value indicators: applying weights to account for market differences. Scientometrics, 96(1):27– 49, July 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0870-y.

Nicolaisen:2019:ZIL

[NF19] Jeppe Nicolaisen and Tove Faber Frandsen. Zero impact: a large-scale study of uncitedness. Scientometrics, 119(2): 1227–1254, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03064-5.

Natale:2012:MRA

[NFH12] Fabrizio Natale, Gianluca Fiore, and Johann Hofherr. Mapping the research on aquaculture. A bibliometric analysis of aquaculture literature. Scientometrics, 90(3):983–999, March 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0562-z.

Nair:2016:WMG

[NG16] Lakshmi Balachandran Nair and Michael Gibbert. What makes a ‘good’ title and (how) does it matter for citations? A review and general model of article title attributes in management science. Scientometrics, 107(3):1331–1359, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1937-y.

Niazi:2011:ABC

[NH11] Muaz Niazi and Amir Hussain. Agent-based computing from multi-agent systems to agent-based models: a visual survey. Scientometrics, 89(2):479–499, November 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0468-9.

Nourmohammadi:2014:PIW

[NH14] Hamzehali Nourmohammadi and Fateme Hodaei. Perspective of Iranian women’s scientific production in high priority fields of science and technology. Scientometrics, 98(2): 1455–1471, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1098-1.

Nguyen:2017:ICS

[NHLL17] Tuan V. Nguyen, Thao P. Ho-Le, and Ut V. Le. International collaboration in scientific research in Vietnam: an analysis of patterns and impact. Scientometrics, 110(2):1035–1051, February 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2201-1.

Niu:2014:GTS

[NHY^{+}14] Beibei Niu, Song Hong, Jiefei Yuan, Sha Peng, Zhen Wang, and Xu Zhang. Global trends in sediment-related research in earth science during 1992–2011: a bibliometric analysis. Scientometrics, 98(1):511–529, January 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1065-x.

Nichols:2014:TMA

[Nic14] Leah G. Nichols. A topic model approach to measuring interdisciplinarity at the national science foundation. Scientometrics, 100(3):741–754, September 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1319-2.

Niidome:2017:RPD

[Nii17] Yutaka Niidome. The relation of patent description and examination with validity: an empirical study. Scientometrics, 111(1):159–183, April 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-017-2272-7.

Nejati:2010:TDA

[NJ10] Ammar Nejati and Seyyed Mehdi Hosseini Jenab. A two-dimensional approach to evaluate the scientific production of countries (case study: the basic sciences). Scientometrics, 84(2):357–364, August 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0103-1.

Nasar:2018:IES

[NJM18] Zara Nasar, Syed Waqar Jaffry, and Muhammad Kamran Malik. Information extraction from scientific articles: a survey. Scientometrics, 117(3):1931–1990, December 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2921-5.

Norton:2017:MGD

[NLCC17] Wynne E. Norton, Alina Lungeanu, David A. Chambers, and Noshir Contractor. Mapping the growing discipline of dissemination and implementation science in health. Scientometrics, 112(3):1367–1390, September 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2455-2.

Nguyen:2011:SOR

[NP11] Tuan V. Nguyen and Ly T. Pham. Scientific output and its relationship to knowledge economy: an analysis of ASEAN countries. Scientometrics, 89(1):107–117, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0446-2.

Nishy:2012:ICE

[NPP^{+}12] P. Nishy, Yatish Panwar, Suresh Prasad, G. K. Mandal, and Gangan Prathap. An impact-citations-exergy (iCX) trajectory analysis of leading research institutions in India. Scientometrics, 91(1):245–251, April 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0594-4.

Nabout:2015:PGP

[NPT^{+}15] João Carlos Nabout, Micael Rosa Parreira, Fabrício Barreto Teresa, Fernanda Melo Carneiro, Hélida Ferreira da Cunha, Luciana de Souza Ondei, Samantha Salomão Caramori, and Thannya Nascimento Soares. Publish (in a group) or perish (alone): the trend from single- to multi-authorship in biological papers. Scientometrics, 102(1):357–364, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-014-1385-5.

Niu:2014:NSD

[NQ14] Fenggao Niu and Junping Qiu. Network structure, distribution and the growth of Chinese international research collaboration. Scientometrics, 98(2):1221–1233, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-1170-x.

Najmi:2017:RTD

[NRAW17] Ali Najmi, Taha H. Rashidi, Alireza Abbasi, and S. Travis Waller. Reviewing the transport domain: an evolutionary bibliometrics and network analysis. Scientometrics, 110(2):843–865, February 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2171-3.

Ni:2013:VCF

[NSC13] Chaoqun Ni, Cassidy R. Sugimoto, and Blaise Cronin. Visualizing and comparing four facets of scholarly communication: producers, artifacts, concepts, and gatekeepers. Scientometrics, 94(3):1161–1173, March 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0849-8.

Nakamura:2011:CLA

[NSH^{+}11] Hiroko Nakamura, Shinji Suzuki, Tomobe Hironori, Yuya Kajikawa, and Ichiro Sakata. Citation lag analysis in supply chain research. Scientometrics, 87(2):221–232, May 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0341-x.

Nakamura:2015:EPF

[NSKO15] Hiroko Nakamura, Shinji Suzuki, Yuya Kajikawa, and Masataka Osawa. The effect of patent family information in patent citation network analysis: a comparative case study in the drivetrain domain. Scientometrics, 104(2):437–452, August 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1626-2.

Nunes-Silva:2019:GIA

[NSMMDB19] Liária Nunes-Silva, Alan Malacarne, Ricardo Fontes Macedo, and Robelius De-Bortoli. Generation of intangible assets in higher education institutions. Scientometrics, 121(2):957–975, November 2019. CODEN SCNTDX.

ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03226-5.

Nobre:2017:SLA

[NT17] Gustavo Cattelan Nobre and Elaine Tavares. Scientific literature analysis on big data and Internet of Things applications on circular economy: a bibliometric study. Scientometrics, 111(1):463–492, April 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-017-2281-6.

Nabout:2018:DTS

[NTM^{+}18] João Carlos Nabout, Fabrício Barreto Teresa, Karine Borges Machado, Vitor Hugo Mendonça do Prado, Luis Mauricio Bini, and José Alexandre Felizola Diniz-Filho. Do traditional scientometric indicators predict social media activity on scientific knowledge? An analysis of the ecological literature. Scientometrics, 115(2):1007–1015, May 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2678-x.

Nederhof:2010:HCN

[NvLvR10] Anton J. Nederhof, Thed N. van Leeuwen, and Anthony F. J. van Raan. Highly cited non-journal publications in political science, economics and psychology: a first exploration. Scientometrics, 83(2):363–374, May 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0086-y.

Nie:2019:ARS

[NZL^{+}19] Yubing Nie, Yifan Zhu, Qika Lin, Sifan Zhang, Pengfei Shi, and Zhendong Niu. Academic rising star prediction via scholar’s evaluation model and machine learning techniques. Scientometrics, 120(2):461–476, August 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03131-x.

Ortega:2010:DNS

[OA10a] José Luis Ortega and Isidro F. Aguillo. Describing national science and technology systems through a multivariate approach: country participation in the 6th framework programmes. Scientometrics, 84(2):321–330, August 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0109-8.

Ortega:2010:NCF

[OA10b] José Luis Ortega and Isidro F. Aguillo. Network collaboration in the 6th framework programmes: country participation in the health thematic area. Scientometrics, 84(3): 835–844, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0212-x.

Ortega:2010:SER

[OA10c] José Luis Ortega and Isidro F. Aguillo. Shaping the European research collaboration in the 6th Framework Programme health thematic area through network analysis. Scientometrics, 85(1):377–386, October 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0218-4.

Ouimet:2011:ISA

[OBG11] Mathieu Ouimet, Pierre-Olivier Bédard, and François Gélineau. Are the h-index and some of its alternatives discriminatory of epistemological beliefs and methodological preferences of faculty members? The case of social scientists in Quebec. Scientometrics, 88(1):91–106, July 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-011-0364-3.pdf.

Osuna:2011:OSA

[OCCSM11] Carmen Osuna, Laura Cruz-Castro, and Luis Sanz-Menéndez. Overturning some assumptions about the effects of evaluation systems on publication performance. Scientometrics, 86(3):575–592, March 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0312-7.

OLeary:2015:BBB

[OCJB15] James D. O’Leary, Mark W. Crawford, Eva Jurczyk, and Alison Buchan. Benchmarking bibliometrics in biomedical research: research performance of the University of Toronto’s Faculty of Medicine, 2008–2012. Scientometrics, 105(1): 311–321, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1676-5.

Oliveira:2012:CBR

[OCM^{+}12] Eduardo A. Oliveira, Enrico A. Colosimo, Daniella R. Martelli, Isabel G. Quirino, Maria Christina L. Oliveira, Leonardo S. Lima, Ana Cristina Simões e Silva, and Hercílio Martelli-Júnior. Comparison of Brazilian researchers in clinical medicine: are criteria for ranking welladjusted? Scientometrics, 90(2):429–443, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0492-9.

Ossenblok:2015:EBS

[OE15] Truyken L. B. Ossenblok and Tim C. E. Engels. Edited books in the social sciences and humanities: Characteristics and collaboration analysis. Scientometrics, 104(1):219–237, July 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1544-3.

Orosz:2016:QCR

[OFP16] Katalin Orosz, Illés J. Farkas, and Péter Pollner. Quantifying the changing role of past publications. Scientometrics, 108(2):829–853, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
016-1971-9.

Olmeda-Gomez:2017:CWA

[OGOPPR17] Carlos Olmeda-Gómez, Maria-Antonia Ovalle-Perandones, and Antonio Perianes-Rodríguez. Co-word analysis and thematic landscapes in Spanish information science literature, 1985–2014. Scientometrics, 113(1):195–217, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2486-8.

Olmeda-Gomez:2019:OTG

[OGRMOP19] Carlos Olmeda-Gómez, Carlos Romá-Mateo, and Maria-Antonia Ovalle-Perandones. Overview of trends in global epigenetic research (2009–2017). Scientometrics, 119(3): 1545–1574, June 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03095-y.

Ohniwa:2019:GPE

[OH19] Ryosuke L. Ohniwa and Aiko Hibino. Generating process of emerging topics in the life sciences. Scientometrics, 121(3): 1549–1561, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03248-z.

Ohniwa:2010:TRF

[OHT10] Ryosuke L. Ohniwa, Aiko Hibino, and Kunio Takeyasu. Trends in research foci in life science fields over the last 30 years monitored by emerging topics. Scientometrics, 85 (1):111–127, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0252-2.

Ozcan:2017:PIR

[OI17] Sercan Ozcan and Nazrul Islam. Patent information retrieval: approaching a method and analysing nanotechnology patent collaborations. Scientometrics, 111(2):941–970, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-017-2325-y.pdf.

Ocholla:2012:ICS

[OING12] Dennis Ocholla, Peter Ingwersen, Ed C. M. Noyons, and Wolfgang Glänzel. The 13th International Conference on Scientometrics and Informetrics. Scientometrics, 91(2):315– 316, May 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-011-0608-2.

Okulicz-Kozaryn:2013:CWA

[OK13] Adam Okulicz-Kozaryn. Cluttered writing: adjectives and adverbs in academia. Scientometrics, 96(3):679– 681, September 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0937-9.

Oleinik:2017:ARC

[OKCPS17] Anton Oleinik, Svetlana Kirdina-Chandler, Irina Popova, and Tatyana Shatalova. On academic reading: citation patterns and beyond. Scientometrics, 113(1):417–435, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2466-z.

Ozel:2014:CAP

[OKK14] Bulent Ozel, Hildrun Kretschmer, and Theo Kretschmer. Co-authorship pair distribution patterns by gender. Scientometrics, 98(1):703–723, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
013-1145-y.

Opthof:2011:CPW

[OL11] Tobias Opthof and Loet Leydesdorff. A comment to the paper by Waltman et al., Scientometrics, 87, 467–481, 2011. Scientometrics, 88(3):1011–1016, September 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-011-0424-8.pdf. See [WvEvL^{+}11b] and

replies [WvEvL^{+}11a, Pra12a].

Oleinik:2012:PPR

[Ole12] Anton Oleinik. Publication patterns in Russia and the West compared. Scientometrics, 93(2):533–551, November 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0698-5.

Olijnyk:2015:AHE

[Oli15a] Nicholas V. Olijnyk. An algorithmic historiography of the Ebola research specialty: mapping the science behind Ebola. Scientometrics, 105(1):623–643, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1688-1;   http://link.
springer.com/content/pdf/10.1007/s11192-015-1688-
1.pdf.

Olijnyk:2015:QEI

[Oli15b] Nicholas V. Olijnyk. A quantitative examination of the intellectual profile and evolution of information security from 1965 to 2015. Scientometrics, 105(2):883–904, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1708-1.

Onyancha:2011:KPT

[OM11] Omwoyo Bosire Onyancha and Jan Resenga Maluleka. Knowledge production through collaborative research in sub-Saharan Africa: how much do countries contribute to each other’s knowledge output and citation impact? Scientometrics, 87(2):315–336, May 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
010-0330-5.

Orduna-Malea:2015:RON

[OMA15] Enrique Orduna-Malea and Selenay Aytac. Revealing the online network between university and industry: the case of Turkey. Scientometrics, 105(3):1849–1866, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1596-4.

Orduna-Malea:2015:MES

[OMAMMLC15] Enrique Orduna-Malea, Juan M. Ayllón, Alberto Martín-Martín, and Emilio Delgado López-Cózar. Methods for estimating the size of Google Scholar. Scientometrics, 104(3): 931–949, September 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1614-6.

Orduna-Malea:2019:UTE

[OMAT19] Enrique Orduna-Malea, Selenay Aytac, and Clara Y. Tran. Universities through the eyes of bibliographic databases: a retroactive growth comparison of Google Scholar, Scopus and Web of Science. Scientometrics, 121(1):433–450, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03208-7.

Omar:2017:GMA

[OMCP17] Muhammad Omar, Arif Mehmood, Gyu Sang Choi, and Han Woo Park. Global mapping of artificial intelligence in Google and Google Scholar. Scientometrics, 113(3):1269– 1305, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2534-4.

Ospina-Mateus:2019:BAM

[OMJLVSN19] Holman Ospina-Mateus, Leonardo Augusto Quintana Jiménez, Francisco J. Lopez-Valdes, and Katherinne Salas-Navarro. Bibliometric analysis in motorcycle accident research: a global overview. Scientometrics, 121(2):793– 815, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03234-5.

Orduna-Malea:2014:GSM

[OMLC14] Enrique Orduña-Malea and Emilio Delgado López-Cózar. Google Scholar Metrics evolution: an analysis according to languages. Scientometrics, 98(3):2353–2367, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-1164-8.

Orduna-Malea:2015:DSO

[OMLC15] Enrique Orduña-Malea and Emilio Delgado López-Cózar. The dark side of open access in Google and Google Scholar: the case of Latin–American repositories. Scientometrics, 102 (1):829–846, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1369-5.

Orduna-Malea:2017:DRS

[OMMMTLC17] Enrique Orduna-Malea, Alberto Martín-Martín, Mike Thelwall, and Emilio Delgado López-Cózar. Do ResearchGate Scores create ghost academic reputations? Scientometrics, 112(1):443–460, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Orduna-Malea:2013:PMU

[OMOR13a] Enrique Orduña-Malea and José-Antonio Ontalba-Ruipérez. Proposal for a multilevel university cybermetric analysis model. Scientometrics, 95(3):863–884, June 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0868-5.

Orduna-Malea:2013:SLS

[OMOR13b] Enrique Orduña-Malea and José-Antonio Ontalba-Ruipérez. Selective linking from social platforms to university websites: a case study of the Spanish academic system. Scientometrics, 95(2):593–614, May 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0851-1.

Orduna-Malea:2014:UAL

[OMR14] Enrique Orduña-Malea and John J. Regazzi. U.S. academic libraries: understanding their web presence and their relationship with economic indicators. Scientometrics, 98(1): 315–336, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1001-0.

Ohba:2012:SBO

[ON12] Norio Ohba and Kumiko Nakao. Sleeping beauties in ophthalmology. Scientometrics, 93(2):253–264, November 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0667-z.

Owusu-Nimo:2017:RCG

[ONB17] Frederick Owusu-Nimo and Nelius Boshoff. Research collaboration in Ghana: patterns, motives and roles. Scientometrics, 110(3):1099–1121, March 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2221-x.

OLeary:2012:RPA

[OO12] James D. O’Leary and Owen O’Sullivan. Research productivity among trainee anaesthetists in Ireland: a crosssectional study. Scientometrics, 93(2):431–438, November 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0684-y.

Oosterhaven:2015:TMJ

[Oos15] Jan Oosterhaven. Too many journals? Towards a theory of repeated rejections and ultimate acceptance. Scientometrics, 103(1):261–265, April 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-015-1527-4.pdf.

Ovalle-Perandones:2013:IEF

[OPGW^{+}13] María-Antonia Ovalle-Perandones, Juan Gorraiz, Martin Wieland, Christian Gumpenberger, and Carlos Olmeda-Gómez. The influence of European framework programmes on scientific collaboration in nanotechnology. Scientometrics, 97(1):59–74, October 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1028-2.

Orazbayev:2017:SOE

[Ora17] Sultan Orazbayev. Sequential order as an extraneous factor in editorial decision. Scientometrics, 113(3):1573–1592, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2531-7.

Ontalba-Ruiperez:2016:IIR

[OROMAA16] José-Antonio Ontalba-Ruipérez, Enrique Orduna-Malea, and Adolfo Alonso-Arroyo. Identifying institutional relationships in a geographically distributed public health system using interlinking and co-authorship methods. Scientometrics, 106(3):1167–1191, March 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1839-z.

Ortega:2011:CPP

[Ort11] José Luis Ortega. Collaboration patterns in patent networks and their relationship with the transfer of technology: the case study of the CSIC patents. Scientometrics, 87(3):657– 666, June 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0363-4.

Ortega:2015:HAS

[Ort15] José Luis Ortega. How is an academic social site populated? A demographic study of Google Scholar citations population. Scientometrics, 104(1):1–18, July 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1593-7.

Ortega:2016:TRT

[Ort16] José Luis Ortega. To be or not to be on Twitter, and its relationship with the tweeting and citation of research papers. Scientometrics, 109(2):1353–1364, November 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2113-0.

Ortega:2017:PRA

[Ort17] José Luis Ortega. Are peer-review activities related to reviewer bibliometric performance? A scientometric analysis of Publons. Scientometrics, 112(2):947–962, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2399-6.

Ortega:2018:RAA

[Ort18] José Luis Ortega. Reliability and accuracy of altmetric providers: a comparison among Altmetric.com, PlumX and Crossref Event Data. Scientometrics, 116(3):2123– 2138, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2838-z.

Onder:2017:CFB

[ÖS17] Ali Sina Önder and Sascha Schweitzer. Catching up or falling behind? Promising changes and persistent patterns across cohorts of economics PhDs in German-speaking countries from 1991 to 2008. Scientometrics, 110(3):1297–1331, March 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2222-9.

Osorio:2018:IPC

[Osó18] António Osório. On the impossibility of a perfect counting method to allocate the credits of multi-authored publications. Scientometrics, 116(3):2161–2173, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2815-6.

Oswald:2010:SMM

[Osw10] Andrew J. Oswald. A suggested method for the measurement of world-leading research (illustrated with data on economics). Scientometrics, 84(1):99–113, July 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0087-x.

Okraku:2017:IPN

[OVJM17] Therese Kennelly Okraku, Raffaele Vacca, James W. Jawitz, and Christopher McCarty. Identity and publication in nonuniversity settings: academic co-authorship and collaboration. Scientometrics, 111(1):401–416, April 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-017-2280-7.

Ozel:2012:CSK

[Oze12a] Bulent Ozel. Collaboration structure and knowledge diffusion in Turkish management academia. Scientometrics, 93 (1):183–206, October 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0641-9.

Ozel:2012:ICS

[Oze12b] Bulent Ozel. Individual cognitive structures and collaboration patterns in academia. Scientometrics, 91(2):539–555, May 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0624-x.

Onel:2011:SAN

[OZK11] Selen Onel, Abe Zeid, and Sagar Kamarthi. The structure and analysis of nanotechnology co-author and citation networks. Scientometrics, 89(1):119–138, October 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0434-6.

Pendlebury:2012:CCT

[PA12] David A. Pendlebury and Jonathan Adams. Comments on a critique of the Thomson Reuters journal impact factor. Scientometrics, 92(2):395–401, August 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0689-6.

Packalen:2019:EFS

[Pac19] Mikko Packalen. Edge factors: scientific frontier positions of nations. Scientometrics, 118(3):787–808, March 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2991-4.

Pajic:2015:GSS

[Paj15] Dejan Pajić. Globalization of the social sciences in Eastern Europe: genuine breakthrough or a slippery slope of the research evaluation practice? Scientometrics, 102(3): 2131–2150, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1510-5.

Parker:2013:CSE

[PAL13] John N. Parker, Stefano Allesina, and Christopher J. Lortie. Characterizing a scientific elite (b): publication and citation patterns of the most highly cited scientists in environmental science and ecology. Scientometrics, 94(2):469–480, February 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0859-6.

Pal:2015:SDC

[Pal15] Jiban K. Pal. Scientometric dimensions of cryptographic research. Scientometrics, 105(1):179–202, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1661-z.

Panat:2014:DAR

[Pan14] Rahul Panat. On the data and analysis of the research output of India and China: India has significantly fallen behind China. Scientometrics, 100(2):471–481, August 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1236-4.

Park:2014:ILL

[Par14a] Han Woo Park. An interview with Loet Leydesdorff: the past, present, and future of the triple helix in the age of big data. Scientometrics, 99(1):199–202, April 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-013-1123-4.pdf.

Park:2014:MEC

[Par14b] Han Woo Park. Mapping election campaigns through negative entropy: Triple and quadruple helix approach to South Korea’s 2012 presidential election. Scientometrics, 99(1): 187–197, April 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1122-5.

Park:2014:TTH

[Par14c] Han Woo Park. Transition from the triple helix to N-tuple helices? An interview with Elias G. Carayannis and David F. J. Campbell. Scientometrics, 99(1):203–207, April 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/content/
pdf/10.1007/s11192-013-1124-3.pdf.

Park:2015:RLM

[Par15] Sungmin Park. The R&D logic model: Does it really work? An empirical verification using successive binary logistic regression models. Scientometrics, 105(3):1399–1439, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1764-6.

Patriota:2018:NLF

[Pat18] Alexandre Galvão Patriota. Is NHST logically flawed? Commentary on: “NHST is still logically flawed”. Scientometrics, 116(3):2189–2191, September 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2817-4. See [Sch18a, Sch18b].

Pautasso:2010:WFD

[Pau10] Marco Pautasso. Worsening file-drawer problem in the abstracts of natural, medical and social science databases. Scientometrics, 85(1):193–202, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0233-5.

Prozesky:2012:BTM

[PB12] Heidi Prozesky and Nelius Boshoff. Bibliometrics as a tool for measuring gender-specific research performance: an example from South African invasion ecology. Scientometrics, 90(2):383–406, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0478-7.

Ponomariov:2016:WCA

[PB16] Branco Ponomariov and Craig Boardman. What is coauthorship? Scientometrics, 109(3):1939–1963, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2127-7.

Packalen:2017:NRS

[PB17a] Mikko Packalen and Jay Bhattacharya. Neophilia ranking of scientific journals. Scientometrics, 110(1):43–64, January 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2157-1.

Pooladian:2017:MIM

[PB17b] Aida Pooladian and Ángel Borrego. Methodological issues in measuring citations in Wikipedia: a case study in Library and Information Science. Scientometrics, 113(1):455–464, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2474-z.

Pavan:2018:APC

[PB18] Cleusa Pavan and Marcia C. Barbosa. Article processing charge (APC) for publishing open access articles: the Brazilian scenario. Scientometrics, 117(2):805–823, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2896-2.

Porter:2014:ISI

[PC14] Alan Porter and Denise Chiavetta. Introduction to special issue on TechMining. Scientometrics, 100(3):611– 612, September 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-014-
1340-5.

Ping:2018:LIS

[PC18] Qing Ping and Chaomei Chen. LitStoryTeller+: an interactive system for multi-level scientific paper visual storytelling with a supportive text mining toolbox. Scientometrics, 116(3):1887–1944, September 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2803-x.

Palacios-Callender:2018:SCC

[PCR18] Miriam Palacios-Callender and Stephen A. Roberts. Scientific collaboration of Cuban researchers working in Europe: understanding relations between origin and destination countries. Scientometrics, 117(2):745–769, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2888-2;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2888-
2.pdf.

Perez-Cabezas:2018:HCP

[PCRMCB^{+}18] Veronica Perez-Cabezas, Carmen Ruiz-Molinero, Ines Carmona-Barrientos, Enrique Herrera-Viedma, Manuel J. Cobo, and Jose A. Moral-Munoz. Highly cited papers in rheumatology: identification and conceptual analysis. Scientometrics, 116(1):555–568, July 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2712-z.

Pislyakov:2010:CET

[PD10] Vladimir Pislyakov and Ekaterina Dyachenko. Citation expectations: are they realized? Study of the Matthew index for Russian papers published abroad. Scientometrics, 83 (3):739–749, June 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0144-5.

Pluskiewicz:2019:SQI

[PDAN19] W. Pluskiewicz, B. Drozdzowska, P. Adamczyk, and K. Noga. Scientific Quality Index: a composite sizeindependent metric compared with h-index for 480 medical researchers. Scientometrics, 119(2):1009–1016, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03078-z;   http://
link.springer.com/content/pdf/10.1007/s11192-019-
03078-z.pdf.

Pinto:2013:ILS

[PEFP13] María Pinto, María Isabel Escalona-Fernández, and Antonio Pulgarín. Information literacy in social sciences and health sciences: a bibliometric study (1974–2011). Scientometrics, 95(3):1071–1094, June 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0899-y.

Peidu:2019:CAP

[Pei19] Ch Peidu. Can authors’ position in the ascription be a measure of dominance? Scientometrics, 121(3):1527–1547, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03254-1.

Pendlebury:2019:CPB

[Pen19] David A. Pendlebury. Charting a path between the simple and the false and the complex and unusable: Review of Henk F. Moed, Applied Evaluative Informetrics [in the series Qualitative and Quantitative Analysis of Scientific and Scholarly Communication, Wolfgang Glänzel, Andras Schubert (eds.)]. Scientometrics, 119(1):549–560, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03045-8.

Pinto:2015:SPI

[PEPUT15] María Pinto, M. Isabel Escalona, Antonio Pulgarín, and Alejandro Uribe-Tirado. The scientific production of Ibero– American authors on information literacy (1985–2013). Scientometrics, 102(2):1555–1576, February 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1498-x.

Persson:2010:HCP

[Per10] Olle Persson. Are highly cited papers more international? Scientometrics, 83(2):397–401, May 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0007-0.

Peritz:2018:GHI

[Per18] Bluma C. Peritz. Gene and his influence on my research and promotion. Scientometrics, 114(2):633–635, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2600-y.

Petrovich:2018:AKP

[Pet18a] Eugenio Petrovich. Accumulation of knowledge in parascientific areas: the case of analytic philosophy. Scientometrics, 116(2):1123–1151, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2796-5.

Petrovich:2018:RW

[Pet18b]  Eugenio Petrovich.  Reply to Wray.  Scientomet-

rics, 117(1):651–654, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-

018-2871-y. See [Wra18].

Palomo:2017:WPS

[PFDL17] Jesus Palomo, Cristina Figueroa-Domecq, and Pilar Laguna. Women, peace and security state-of-art: a bibliometric analysis in social sciences based on SCOPUS database. Scientometrics, 113(1):123–148, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2484-x.

Pagel:2019:YAG

[PFL19] Paul S. Pagel, Julie K. Freed, and Cynthia A. Lien. A 50-year analysis of gender differences in United States authorship of original research articles in two major anesthesiology journals. Scientometrics, 121(1):371–386, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03192-y.

Pinto:2019:SPM

[PFPCM^{+}19] María Pinto, Rosaura Fernández-Pascual, David Caballero-Mariscal, Dora Sales, David Guerrero, and Alejandro Uribe. Scientific production on mobile information literacy in higher education: a bibliometric analysis (2006– 2017). Scientometrics, 120(1):57–85, July 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03115-x.

Pudovkin:2012:RNI

[PG12] A. I. Pudovkin and Eugene Garfield. Rank normalization of impact factors will resolve Vanclay’s dilemma with TRIF. Scientometrics, 92(2):409–412, August 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0634-8.

Paolucci:2014:MCS

[PG14a] Mario Paolucci and Francisco Grimaldo. Mechanism change in a simulation of peer review: from junk support to elitism. Scientometrics, 99(3):663–688, June 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-014-1239-1.pdf.

Persson:2014:DHA

[PG14b] Olle Persson and Wolfgang Glänzel. Discouraging honorific authorship. Scientometrics, 98(2):1417–1419, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1042-4.

Gao:2012:HDC

[pGDTP12] Ji ping Gao, Kun Ding, Li Teng, and Jie Pang. Hybrid documents co-citation analysis: making sense of the interaction between science and technology in technology diffusion. Scientometrics, 93(2):459–471, November 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0691-z.

Gao:2019:RFE

[pGSyW^{+}19] Ji ping Gao, Cheng Su, Hai yan Wang, Li hua Zhai, and Yun tao Pan. Research fund evaluation based on academic publication output analysis: the case of Chinese research fund evaluation. Scientometrics, 119(2):959–972, May 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03073-4.

Pouris:2014:REC

[PH14] Anastassios Pouris and Yuh-Shan Ho. Research emphasis and collaboration in Africa. Scientometrics, 98(3): 2169–2184, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1156-8.

Paul-Hus:2015:FYG

[PHBN^{+}15] Adèle Paul-Hus, Rébecca L. Bouvier, Chaoqun Ni, Cassidy R. Sugimoto, Vladimir Pislyakov, and Vincent Larivière. Forty years of gender disparities in Russian science: a historical bibliometric analysis. Scientometrics, 102(2): 1541–1553, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1386-4.

PhD:2018:ADJ

[PhD18] Erwin Krauskopf PhD. An analysis of discontinued journals by Scopus. Scientometrics, 116(3):1805–1815, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2808-5.

Paul-Hus:2016:CDC

[PHDC16] Adèle Paul-Hus, Nadine Desrochers, and Rodrigo Costas. Characterization, description, and considerations for the use of funding acknowledgement data in Web of Science. Scientometrics, 108(1):167–182, July 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1953-y.

Park:2017:ERI

[PHL17] Jungkyu Park, Eunnyeong Heo, and Dongjun Lee. Effective R&D investment planning based on technology spillovers: the case of Korea. Scientometrics, 111(1):67–82, April 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-017-2276-3.

Pratt:2012:DIS

[PHS12] Jean A. Pratt, Karina Hauser, and Cassidy R. Sugimoto. Defining the intellectual structure of information systems and related college of business disciplines: a bibliometric analysis. Scientometrics, 93(2):279–304, November 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0668-y.

Petersen:2017:EGJ

[PHV17] Jessica Petersen, Fabian Hattke, and Rick Vogel. Editorial governance and journal impact: a study of management and business journals. Scientometrics, 112(3):1593– 1614, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2434-7.

Perlin:2018:PPR

[PIB18] Marcelo S. Perlin, Takeyoshi Imasato, and Denis Borenstein. Is predatory publishing a real threat? Evidence from a large database study. Scientometrics, 116(1):255– 273, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2750-6.

Pietrucha:2018:CSD

[Pie18] Jacek Pietrucha. Country-specific determinants of world university rankings. Scientometrics, 114(3):1129–1139, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-017-2634-1;   https://
link.springer.com/content/pdf/10.1007/s11192-017-
2634-1.pdf.

Pinto:2015:VES

[Pin15] María Pinto. Viewing and exploring the subject area of information literacy assessment in higher education (2000– 2011). Scientometrics, 102(1):227–245, January 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1440-2.

Piro:2019:RDC

[Pir19] Fredrik Niclas Piro. The R&D composition of European countries: concentrated versus dispersed profiles. Scientometrics, 119(2):1095–1119, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03062-7.

Peclin:2012:EIC

[PJB^{+}12] Stojan Peclin, Primoz Juznic, Rok Blagus, Mojca Cizek Sajko, and Janez Stare. Effects of international collaboration and status of journal on impact of papers. Scientometrics, 93(3):937–948, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0768-8.

Pan:2019:GST

[PJL19] Weiwei Pan, Lirong Jian, and Tao Liu. Grey system theory trends from 1991 to 2018: a bibliometric analysis and visualization. Scientometrics, 121(3):1407–1434, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03256-z.

Park:2017:AVT

[PJY17] Inchae Park, Yujin Jeong, and Byungun Yoon. Analyzing the value of technology based on the differences of patent citations between applicants and examiners. Scientometrics, 111(2):665–691, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Parinov:2014:SLR

[PK14] Sergey Parinov and Mikhail Kogalovsky. Semantic linkages in research information systems as a new data source for scientometric studies. Scientometrics, 98(2):927–943, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1108-3.

Peters:2016:RDE

[PKL^{+}16] Isabella Peters, Peter Kraker, Elisabeth Lex, Christian Gumpenberger, and Juan Gorraiz. Research data explored: an extended analysis of citations and altmetrics. Scientometrics, 107(2):723–744, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/

article/10.1007/s11192-016-1887-4;   http://link.
springer.com/content/pdf/10.1007/s11192-016-1887-
4.pdf.

Pagani:2015:MOP

[PKR15] Regina Negri Pagani, João Luiz Kovaleski, and Luis Mauricio Resende. Methodi Ordinatio: a proposed methodology to select and rank relevant scientific papers encompassing the impact factor, number of citation, and year of publication. Scientometrics, 105(3):2109–2135, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1744-x.

Pudovkin:2012:REP

[PKSG12] A. Pudovkin, H. Kretschmer, J. Stegmann, and E. Garfield. Research evaluation. Part I: productivity and citedness of a German medical research institution. Scientometrics, 93 (1):3–16, October 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0659-z.

Poon:2017:TGG

[PL17] Wai Ching Poon and Gareth D. Leeves. Is there gender gap unequivocally? Evidence from research output 1958–2008. Scientometrics, 111(3):1687–1701, June 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Pohl:2018:RCI

[PL18] Hans Pohl and Jason E. Lane. Research contributions of international branch campuses to the scientific wealth of academically developing countries. Scientometrics, 116(3):1719–1734, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2790-y.

Parker:2010:CSE

[PLA10] John N. Parker, Christopher Lortie, and Stefano Allesina. Characterizing a scientific elite: the social characteristics of the most highly cited scientists in environmental science and ecology. Scientometrics, 85(1):129–143, October 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-010-0234-4.pdf.

Pedrini:2018:AHR

[PLBZ18] Matteo Pedrini, Valentina Langella, Mario Alberto Battaglia, and Paola Zaratin. Assessing the health research’s social impact: a systematic review. Scientometrics, 114(3): 1227–1250, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2585-6.

Prabhakaran:2019:CCC

[PLG19] Thara Prabhakaran, Hiran H. Lathabai, and Susan George. Competing, complementary and co-existing paradigms in techno-scientific literature: A case study of nanotechnology for engineering. Scientometrics, 118(3):941–977, March 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03013-2.

Prabhakaran:2018:TPP

[PLGC18] Thara Prabhakaran, Hiran H. Lathabai, Susan George, and Manoj Changat. Towards prediction of paradigm shifts from scientific literature. Scientometrics, 117(3):1611–1644, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2931-3.

Park:2018:IPU

[PLJ18] Tae-Young Park, Hyungjoo Lim, and Ilyong Ji. Identifying potential users of technology for technology transfer using patent citation analysis: a case analysis of a Korean research institute. Scientometrics, 116(3):1541–1558, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2792-9.

Pezzoni:2014:HKI

[PLT14] Michele Pezzoni, Francesco Lissoni, and Gianluca Tarasconi. How to kill inventors: testing the Massacrator© algorithm for inventor disambiguation. Scientometrics, 101 (1):477–504, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1375-7.

Peng:2015:GTR

[PLW^{+}15] Yuling Peng, Aiwen Lin, Ke Wang, Fenglian Liu, Fang Zeng, and Li Yang. Global trends in DEM-related research from 1994 to 2013: a bibliometric analysis. Scientometrics, 105 (1):347–366, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1666-7.

Ponomarev:2014:BPI

[PLWS14] Ilya V. Ponomarev, Brian K. Lawton, Duane E. Williams, and Joshua D. Schnell. Breakthrough paper indicator 2.0: can geographical diversity and interdisciplinarity improve the accuracy of outstanding papers prediction? Scientometrics, 100(3):755–765, September 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1320-9.

Patra:2018:RIS

[PM18] Swapan Kumar Patra and Mammo Muchie. Research and innovation in South African universities: from the triple helix’s perspective. Scientometrics, 116(1):51–76, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2764-0.

Payumo:2019:MBP

[PMJF19] Jane G. Payumo, Jamie Monson, Amy Jamison, and Bradley W. Fenwick. Metrics-based profiling of university research engagement with Africa: research management, gender, and internationalization perspective. Scientometrics, 121(2):675–698, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03211-y.

Puuska:2014:IDC

[PML14] Hanna-Mari Puuska, Reetta Muhonen, and Yrjö Leino. International and domestic co-publishing and their citation impact in different disciplines. Scientometrics, 98(2): 823–839, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1181-7.

Parreira:2017:RGD

[PML^{+}17] Micael Rosa Parreira, Karine Borges Machado, Ramiro Logares, José Alexandre Felizola Diniz-Filho, and João Carlos Nabout. The roles of geographic distance and socioeconomic factors on international collaboration among ecologists. Scientometrics, 113(3):1539–1550, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2502-z.

Prathap:2016:DHI

[PMN16] Gangan Prathap, S. Mini, and P. Nishy. Does high impact factor successfully predict future citations? An analysis using Peirce’s measure. Scientometrics, 108(3):1043– 1047, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-016-
2034-y.

Piepenbrink:2015:TLT

[PN15] Anke Piepenbrink and Elkin Nurmammadov. Topics in the literature of transition economies and emerging markets. Scientometrics, 102(3):2107–2130, March 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1513-2.

Padial:2010:WED

[PNS^{+}10] André Andrian Padial, João Carlos Nabout, Tadeu Siqueira, Luis Mauricio Bini, and José Alexandre Felizola Diniz-Filho. Weak evidence for determinants of citation frequency in ecological articles. Scientometrics, 85(1):1–12, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0231-7.

Palacios-Nunez:2018:DTA

[PNVCB18] Guadalupe Palacios-Núñez, Gabriel Vélez-Cuartas, and Juan D. Botero. Developmental tendencies in the academic field of intellectual property through the identification of invisible colleges. Scientometrics, 115(3):1561–1574, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2648-3.

Polonioli:2016:DUD

[Pol16a] Andrea Polonioli. Debunking unwarranted defenses of the status quo in the humanities and social sciences. Scientometrics, 107(3):1519–1522, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1906-5.

Polonioli:2016:MFI

[Pol16b] Andrea Polonioli. Metrics, flawed indicators, and the case of philosophy journals. Scientometrics, 108(2):987–994, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1941-2.

Pouris:2010:SAS

[Pou10] Anastassios Pouris. A scientometric assessment of the Southern Africa Development Community: science in the tip of Africa. Scientometrics, 85(1):145–154, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0260-2.

Pouris:2012:SRS

[Pou12] Anastassios Pouris. Scientometric research in South Africa and successful policy instruments. Scientometrics, 91(2): 317–325, May 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0581-9.

Pouris:2011:SPH

[PP11] Anthipi Pouris and Anastassios Pouris. Scientometrics of a pandemic: HIV/AIDS research in South Africa and the World. Scientometrics, 86(2):541–552, February 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0277-6.

Pisoschi:2016:OAS

[PP16] Aurelia Magdalena Pisoschi and Claudia Gabriela Pisoschi. Is open access the solution to increase the impact of scientific journals? Scientometrics, 109(2):1075–1095, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2088-x.

Park:2018:REA

[PP18] Hyejin Park and Han Woo Park. Research evaluation of Asian countries using altmetrics: comparing South Korea, Japan, Taiwan, Singapore, and China. Scientometrics, 117(2):771–788, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2884-6.

Pinto:2014:VIL

[PPE14] María Pinto, Antonio Pulgarín, and M. Isabel Escalona. Viewing information literacy concepts: a comparison of two branches of knowledge. Scientometrics, 98(3):2311–2329, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1166-6.

Polyakov:2017:DAC

[PPI17] Maksym Polyakov, Serhiy Polyakov, and Md Sayed Iftekhar. Does academic collaboration equally benefit impact of research across topics? The case of agricultural, resource, environmental and ecological economics. Scientometrics, 113(3):1385–1405, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2523-7.

Picinin:2016:CPR

[PPK^{+}16] Claudia Tania Picinin, Luiz Alberto Pilatti, João Luiz Kovaleski, Alexandre Reis Graeml, and Bruno Pedroso. Comparison of performance of researchers recipients of CNPq productivity grants in the field of Brazilian production engineering. Scientometrics, 109(2):855–870, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2070-7.

Pradhan:2017:CIP

[PPM^{+}17] Dinesh Pradhan, Partha Sarathi Paul, Umesh Maheswari, Subrata Nandi, and Tanmoy Chakraborty. C^{3}-index: a PageRank based multi-faceted metric for authors’ performance measurement. Scientometrics, 110(1):253–273, January 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2168-y.

Pinto:2014:DIV

[PQG14] María Pinto, David Guerrero Quesada, and Ximo Granell. Dissemination of information and visibility of the European Higher Education Area through the websites of Spanish universities: a longitudinal metric analysis, 2007– 2012. Scientometrics, 98(2):1235–1255, February 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1062-0.

Pepe:2010:CSN

[PR10] Alberto Pepe and Marko A. Rodriguez. Collaboration in sensor network research: an in-depth longitudinal analysis of assortative mixing patterns. Scientometrics, 84(3): 687–701, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-009-0147-
2.pdf.

Plotnikova:2014:CPR

[PR14] Tatiana Plotnikova and Bastian Rake. Collaboration in pharmaceutical research: exploration of country-level determinants. Scientometrics, 98(2):1173–1202, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-1182-6.

Poelmans:2015:FDA

[PR15] Eline Poelmans and Sandra Rousseau. Factors determining authors’ willingness to wait for editorial decisions from economic history journals. Scientometrics, 102(2):1347–1374, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1469-2.

Prathap:2010:MPE

[Pra10a] Gangan Prathap. The 100 most prolific economists using the p-index. Scientometrics, 84(1):167–172, July 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0068-0.

Prathap:2010:GMB

[Pra10b] Gangan Prathap. Going much beyond the Durfee square: enhancing the h_{T} index. Scientometrics, 84(1):149–152, July 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0067-1.

Prathap:2010:IAJ

[Pra10c] Gangan Prathap. The iCE approach for journal evaluation. Scientometrics, 85(2):561–565, November 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0239-z.

Prathap:2010:IMA

[Pra10d] Gangan Prathap. An iCE map approach to evaluate performance and efficiency of scientific production of countries. Scientometrics, 85(1):185–191, October 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0192-x.

Prathap:2010:TPM

[Pra10e] Gangan Prathap. Is there a place for a mock hindex? Scientometrics, 84(1):153–165, July 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0066-2.

Prathap:2011:EEE

[Pra11a] Gangan Prathap. The energy–exergy–entropy (or EEE) sequences in bibliometric assessment. Scientometrics, 87 (3):515–524, June 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0367-0.

Prathap:2011:FHI

[Pra11b] Gangan Prathap. The fractional and harmonic p-indices for multiple authorship. Scientometrics, 86(2):239–244, February 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0257-x.

Prathap:2011:LEC

[Pra11c] Gangan Prathap. Letter to the Editor: Comments on the paper of Franceschini and Maisano: Proposals for evaluating the regularity of a scientist’s research output. Scientometrics, 88(3):1005–1010, September 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0413-y. See [FM11d].

Prathap:2011:QWQ

[Pra11d] Gangan Prathap. Quasity, when quantity has a quality all of its own-toward a theory of performance. Scientometrics, 88 (2):555–562, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0401-2.

Prathap:2012:CPO

[Pra12a] Gangan Prathap. A comment to the papers by Opthof and Leydesdorff, Scientometrics, 88, 1011–1016, 2011 and Waltman et al., Scientometrics, 88, 1017–1022, 2011. Scientometrics, 90(2):737–743, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0500-0. See [OL11, WvEvL^{+}11a].

Prathap:2012:EIP

[Pra12b] Gangan Prathap. Energy indicators and percentile ranking normalization. Scientometrics, 91(3):997–1003, June 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0620-1.

Prathap:2012:EJP

[Pra12c] Gangan Prathap. Evaluating journal performance metrics. Scientometrics, 92(2):403–408, August 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0746-1.

Prathap:2012:QQQ

[Pra12d] Gangan Prathap. The quality–quantity–quasity and energy–exergy–entropy exegesis of expected value calculation of citation performance. Scientometrics, 91(1):269–275, April 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0516-5.

Prathap:2013:SOI

[Pra13] Gangan Prathap. Second order indicators for evaluating international scientific collaboration. Scientometrics, 95 (2):563–570, May 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0804-8.

Prathap:2014:BDF

[Pra14a] Gangan Prathap. Big data and false discovery: analyses of bibliometric indicators from large data sets. Scientometrics, 98(2):1421–1422, February 2014. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-013-1063-z.

Prathap:2014:SPI

[Pra14b] Gangan Prathap. Single parameter indices and bibliometric outliers. Scientometrics, 101(3):1781–1787, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-1225-z.

Prathap:2014:TDB

[Pra14c] Gangan Prathap. A three-dimensional bibliometric evaluation of research in polymer solar cells. Scientometrics, 101 (1):889–898, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1346-z.

Piro:2016:HDP

[PRA16a] Fredrik Niclas Piro, Kristoffer Rørstad, and Dag W. Aksnes. How does prolific professors influence on the citation impact of their university departments? Scientometrics, 107(3):941–961, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1900-y.

Prathap:2016:LEC

[Pra16b] Gangan Prathap. Letter to the editor: comments on the paper of Aparna Basu et al.: Designing a composite index for research performance evaluation at the national or regional level: ranking central universities in India. Scientometrics, 108(3):1685–1687, September 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2020-4. See [BBSS16a].

Prathap:2017:LEC

[Pra17a] Gangan Prathap. Letter to the editor: Comments on the paper of Lucio Bertoli-Barsotti and Tommaso Lando: A theoretical model of the relationship between the h-index and other simple citation indicators. Scientometrics, 112(2):1133–1136, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2404-0. See [BBL17b, BB17a].

Prathap:2017:SWI

[Pra17b] Gangan Prathap. Scientific wealth and inequality within nations. Scientometrics, 113(2):923–928, November 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2511-y.

Prathap:2017:TDB

[Pra17c]  Gangan Prathap.  A three-dimensional bibliometric

evaluation of recent research in India. Scientometrics, 110(3):1085–1097, March 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2181-1.

Prathap:2018:BTT

[Pra18a] Gangan Prathap. A bibliometric tale of two cities: Hong Kong and Singapore. Scientometrics, 117(3):2169–2175, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2927-z.

Prathap:2018:EGM

[Pra18b] Gangan Prathap. Eugene Garfield: from the metrics of science to the science of metrics. Scientometrics, 114(2): 637–650, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2525-5.

Prathap:2018:LED

[Pra18c] Gangan Prathap. Letter to the editor: Dimensionless citation indicators. Scientometrics, 115(3):1433–1435, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2736-4.

Prathap:2018:PRU

[Pra18d] Gangan Prathap. Performance of research universities in post-communist countries. Scientometrics, 117(3):2037– 2039, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2938-9.

Prathap:2018:TIO

[Pra18e] Gangan Prathap. Totalized input-output assessment of research productivity of nations using multi-dimensional input and output. Scientometrics, 115(1):577–583, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2605-6.

Prathap:2019:BTP

[Pra19a] Gangan Prathap. Balance: a thermodynamic perspective. Scientometrics, 119(1):247–255, April 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03020-3.

Prathap:2019:EOR

[Pra19b] Gangan Prathap. Expected, observed and relative paper scores from heterogeneous author-paper-citation networks. Scientometrics, 119(2):1275–1279, May 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03039-6.

Prathap:2019:LER

[Pra19c] Gangan Prathap. Letter to the editor: Revisiting the h-index and the p-index. Scientometrics, 121(3):1829– 1833, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03240-7.

Prathap:2019:LES

[Pra19d] Gangan Prathap. Letter to the editor: Second-order htype indicators. Scientometrics, 121(3):1825–1827, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03249-y.

Prathap:2019:PNI

[Pra19e] Gangan Prathap. The Pinski–Narin influence weight and the Ramanujacharyulu power-weakness ratio indicators revisited. Scientometrics, 119(2):1173–1185, May 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03046-7.

Prathap:2019:SDS

[Pra19f] Gangan Prathap. Scale-dependent stratification: a skylineshoreline scatter plot. Scientometrics, 119(2):1269–1273, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03038-7.

Praus:2019:HRC

[Pra19g] Petr Praus. High-ranked citations percentage as an indicator of publications quality. Scientometrics, 120(1):319–329, July 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03128-6.

Popova:2017:CBC

[PRDG17] Olga Popova, Dmitry Romanov, Alexander Drozdov, and Alexander Gerashchenko. Citation-based criteria of the significance of the research activity of scientific teams. Scientometrics, 112(3):1179–1202, September 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2427-6.

Pritychenko:2015:ITN

[Pri15] B. Pritychenko. Intriguing trends in nuclear physics authorship. Scientometrics, 105(3):1781–1786, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1605-7.

Pritychenko:2016:ELL

[Pri16a] B. Pritychenko. Evolving landscape of low-energy nuclear physics publications. Scientometrics, 109(3):2067–2076, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2149-1.

Pritychenko:2016:FAN

[Pri16b] B. Pritychenko. Fractional authorship in nuclear physics. Scientometrics, 106(1):461–468, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1766-4.

Perianes-Rodriguez:2019:EJC

[PROG19] Antonio Perianes-Rodríguez and Carlos Olmeda-Gómez. Effects of journal choice on the visibility of scientific publications: a comparison between subscription-based and full Open Access models. Scientometrics, 121(3):1737–1752, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03265-y.

Perianes-Rodriguez:2010:DIV

[PROGMA10] Antonio Perianes-Rodríguez, Carlos Olmeda-Gómez, and Félix Moya-Anegón. Detecting, identifying and visualizing research groups in co-authorship networks. Scientometrics, 82(2):307–319, February 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0040-z.

Perianes-Rodriguez:2015:WBD

[PRRC15] Antonio Perianes-Rodriguez and Javier Ruiz-Castillo. Within- and between-department variability in individual productivity: the case of economics. Scientometrics, 102(2): 1497–1520, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1449-6.

Perianes-Rodriguez:2016:CTW

[PRRC16] Antonio Perianes-Rodriguez and Javier Ruiz-Castillo. A comparison of two ways of evaluating research units working in different scientific fields. Scientometrics, 106(2):539–561, February 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1801-5.

Perovic:2016:ORT

[PRSB16] Slobodan Perović, Sandro Radovanović, Vlasta Sikimić, and Andrea Berber. Optimal research team composition: data envelopment analysis of Fermilab experiments. Scientometrics, 108(1):83–111, July 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1947-9.

Pautasso:2010:PRD

[PS10] Marco Pautasso and Hanno Schäfer. Peer review delay and selectivity in ecology journals. Scientometrics, 84(2): 307–315, August 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0105-z.

Piergiovanni:2013:MYS

[PS13] Roberta Piergiovanni and Enrico Santarelli. The more you spend, the more you get? The effects of R&D and capital expenditures on the patenting activities of biotechnology firms. Scientometrics, 94(2):497–521, February 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0711-z.

Payumo:2015:BAA

[PS15] Jane G. Payumo and Taurean C. Sutton. A bibliometric assessment of ASEAN collaboration in plant biotechnology. Scientometrics, 103(3):1043–1059, June 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1582-x.

Pastor:2016:DRO

[PS16a] José Manuel Pastor and Lorenzo Serrano. The determinants of the research output of universities: specialization, quality and inefficiencies. Scientometrics, 109(2):1255– 1281, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-016-
2102-3.

Piro:2016:HCD

[PS16b] Fredrik Niclas Piro and Gunnar Sivertsen. How can differences in international university rankings be explained? Scientometrics, 109(3):2263–2278, December 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2056-5.

Payumo:2017:IOA

[PSB^{+}17] Jane Payumo, Taurean Sutton, Derek Brown, Dan Nordquist, Marc Evans, Danna Moore, and Prema Arasu. Input-output analysis of international research collaborations: a case study of five U.S. universities. Scientometrics, 111(3):1657–1671, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Porter:2019:LAL

[PSY^{+}19] Alan L. Porter, David J. Schoeneck, Jan Youtie, Gregg E. A. Solomon, Seokbeom Kwon, and Stephen F. Carley. Learning about learning: patterns of sharing of research knowledge among education, border, and cognitive science fields. Scientometrics, 118(3):1093–1117, March 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03012-3.

Pastor:2015:ROE

[PSZ15] José Manuel Pastor, Lorenzo Serrano, and Irene Zaera. The research output of European higher education institutions. Scientometrics, 102(3):1867–1893, March 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1509-y.

Pan:2017:ACF

[PT17] Fan Pan and Guoxiao Tao. Alex Chengyu, Fang and Jing, Cao: Text genres and registers: The computation of linguistic features. Scientometrics, 113(3):1815–1818, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2546-0.

Perakakis:2010:NSA

[PTMT10] Pandelis Perakakis, Michael Taylor, Marco Mazza, and Varvara Trachana. Natural selection of academic papers. Scientometrics, 85(2):553–559, November 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0253-1.

Perakakis:2011:URO

[PTMT11] Pandelis Perakakis, Michael Taylor, Marco G. Mazza, and Varvara Trachana. Understanding the role of open peer review and dynamic academic articles. Scientometrics, 88 (2):669–673, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0402-1.

Puuska:2010:ESG

[Puu10] Hanna-Mari Puuska. Effects of scholar’s gender and professional position on publishing productivity in different publication types. Analysis of a Finnish university. Scientometrics, 82(2):419–437, February 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0037-7.

Pecha:2015:ACH

[PV15] Ondrej Pecha and Jirí Vanecek. Analysis of the Czech and Hungarian physiology publications 1994–2011. Scientometrics, 105(2):991–1003, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1713-4.

Peng:2013:NCB

[PW13] Tai Quan Peng and Zhen-Zhen Wang. Network closure, brokerage, and structural influence of journals: a longitudinal study of journal citation network in Internet research (2000– 2010). Scientometrics, 97(3):675–693, December 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1012-x.

Park:2017:ERD

[PW17] Hyoungjoo Park and Dietmar Wolfram. An examination of research data sharing and re-use: implications for data citation practice. Scientometrics, 111(1):443–461, April 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-017-2240-2.

Park:2014:ACI

[PY14] Hyunseok Park and Janghyeok Yoon. Assessing coreness and intermediarity of technology sectors using patent co-classification analysis: the case of Korean national R&D. Scientometrics, 98(2):853–890, February 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1109-2.

Park:2019:SCI

[PY19] Han Woo Park and Jungwon Yoon. Structural characteristics of institutional collaboration in North Korea analyzed through domestic publications. Scientometrics, 119(2):771– 787, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03056-5.

Pan:2016:DDS

[PYH16] Xuelian Pan, Erjia Yan, and Weina Hua. Disciplinary differences of software use and impact in scientific literature. Scientometrics, 109(3):1593–1610, December 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2138-4.

Park:2012:IPI

[PYK12] Hyunseok Park, Janghyeok Yoon, and Kwangsoo Kim. Identifying patent infringement using SAO based semantic technological similarities. Scientometrics, 90(2):515–529, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0522-7.

Park:2013:IEC

[PYK13] Hyunseok Park, Janghyeok Yoon, and Kwangsoo Kim. Identification and evaluation of corporations for merger and acquisition strategies using patent information and text mining. Scientometrics, 97(3):883–909, December 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1010-z.

Park:2016:NCA

[PYL16] Han Woo Park, Jungwon Yoon, and Loet Leydesdorff. The normalization of co-authorship networks in the bibliometric evaluation: the government stimulation programs of China and Korea. Scientometrics, 109(2):1017–1036, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-1978-2.

Pak:2018:SCS

[PYW18] CholMyong Pak, Guang Yu, and Weibin Wang. A study on the citation situation within the citing paper: citation distribution of references according to mention frequency. Scientometrics, 114(3):905–918, March 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2627-0.

Potthoff:2017:TGB

[PZ17] Matthias Potthoff and Fabian Zimmermann. Is there a gender-based fragmentation of communication science? An investigation of the reasons for the apparent gender homophily in citations. Scientometrics, 112(2):1047–1063, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2392-0.

Quayle:2018:RRA

[QA18] Michael Quayle and Maura Adshead. The resilience of regional African HIV/AIDS research networks to the withdrawal of international authors in the subfield of public administration and governance: lessons for funders and collaborators. Scientometrics, 117(1):163–173, October 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2863-y.

Qayyum:2019:IIC

[QA19] Faiza Qayyum and Muhammad Tanvir Afzal. Identification of important citations by exploiting research articles’ metadata and cue-terms from content. Scientometrics, 118 (1):21–43, January 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2961-x.

Qin:2019:SSV

[QDK19] Xionghe Qin, Debin Du, and Mei-Po Kwan. Spatial spillovers and value chain spillovers: evaluating regional R&D efficiency and its spillover effects in China. Scientometrics, 119(2):721–747, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03054-7.

Qiu:2014:CSS

[QDY14] Jun-Ping Qiu, Ke Dong, and Hou-Qiang Yu. Comparative study on structure and correlation among author co-occurrence networks in bibliometrics. Scientometrics, 101(2):1345–1360, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1315-6.

Jiang:2017:MFA

[qJnShPL17] Sheng qiang Jiang, An na Shi, Zhi hang Peng, and Xin Li. Major factors affecting cross-city R&D collaborations in China: evidence from cross-sectional co-patent data between 224 cities. Scientometrics, 111(3):1251–1266, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Qureshi:2014:IHE

[QJZ^{+}14] Muhammad Imran Qureshi, Saquib Yusaf Janjua, Khalid Zaman, Mohammad Saeed Lodhi, and Yasir Bin Tariq. Internationalization of higher education institutions: implementation of DMAIC cycle. Scientometrics, 98(3): 2295–2310, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1163-9.

Quental:2012:RAJ

[QL12] Nuno Quental and Júlia M. Lourenço. References, authors, journals and scientific disciplines underlying the sustainable development literature: a citation analysis. Scientometrics, 90(2):361–381, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0533-4.

Quan:2019:DCL

[QMSM^{+}19] Wei Quan, Philippe Mongeon, Maxime Sainte-Marie, Rongying Zhao, and Vincent Larivière. On the development of China’s leadership in international collaborations. Scientometrics, 120(2):707–721, August 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03111-1.

Qian:2017:CRA

[QRJ^{+}17] Yifan Qian, Wenge Rong, Nan Jiang, Jie Tang, and Zhang Xiong. Citation regression analysis of computer science publications in different ranking categories and subfields. Scientometrics, 110(3):1351–1374, March 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2235-4.

Quesada:2010:MAH

[Que10] Antonio Quesada. More axiomatics for the Hirsch index. Scientometrics, 82(2):413–418, February 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0026-x.

Quesada:2011:FCH

[Que11] Antonio Quesada. Further characterizations of the Hirsch index. Scientometrics, 87(1):107–114, April 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0307-4.

Qi:2017:SSG

[QZL^{+}17] Mengjiao Qi, An Zeng, Menghui Li, Ying Fan, and Zengru Di. Standing on the shoulders of giants: the effect of outstanding scientists on young collaborators’ careers. Scientometrics, 111(3):1839–1850, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Qu:2017:PRF

[QZZ17] Zhao Qu, Shanshan Zhang, and Chunbo Zhang. Patent research in the field of library and information science: Less useful or difficult to explore? Scientometrics, 111(1):205–217, April 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-017-2269-2.

Qi:2018:MBR

[QZZD18] Yashuang Qi, Na Zhu, Yujia Zhai, and Ying Ding. The mutually beneficial relationship of patents and scientific literature: topic evolution in nanoscience. Scientometrics, 115 (2):893–911, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2693-y.

Raheel:2018:EIV

[RAA18] Muhammad Raheel, Samreen Ayaz, and Muhammad Tanvir Afzal. Evaluation of h-index, its variants and extensions based on publication age & citation intensity in civil engineering. Scientometrics, 114(3):1107–1127, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2633-2.

Raimbault:2019:EIS

[Rai19] Juste Raimbault. Exploration of an interdisciplinary scientific landscape. Scientometrics, 119(2):617–641, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03090-3.

Repiso:2018:PEW

[RAM18] Rafael Repiso, Josu Ahedo, and Julio Montero. The presence of the encyclicals in Web of Science: a bibliometric approach. Scientometrics, 115(1):487–500, April 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2636-z.

Randic:2009:CVL

[Ran09] Milan Randić. Citations versus limitations of citations: beyond Hirsch index. Scientometrics, 80(3):809–818, September 2009. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-008-2128-2. See letter

[Egg10e].

Ravikumar:2015:MIS

[RAS15] S. Ravikumar, Ashutosh Agrahari, and S. N. Singh. Mapping the intellectual structure of scientometrics: a coword analysis of the journal Scientometrics (2005–2010). Scientometrics, 102(1):929–955, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1402-8.

Rajagopal:2013:ROP

[RASP13] Thangavel Rajagopal, Govindaraju Archunan, Muthuraj Surulinathi, and Ponnirul Ponmanickam. Research output in pheromone biology: a case study of India. Scientometrics, 94(2):711–719, February 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0788-4.

Rossetto:2018:SEI

[RBBG18] Dennys Eduardo Rossetto, Roberto Carlos Bernardes, Felipe Mendes Borini, and Cristiane Chaves Gattaz. Structure and evolution of innovation research in the last 60 years: review and future trends in the field of business through the citations and co-citations analysis. Scientometrics, 115 (3):1329–1363, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2709-7.

Ruiz:2010:EMR

[RBC^{+}10] Cristhian Fabián Ruiz, Ricardo Bonilla, Diego Chavarro, Luis Antonio Orozco, Roberto Zarama, and Xavier Polanco. Efficiency measurement of research groups using data envelopment analysis and Bayesian networks. Scientometrics, 83 (3):711–721, June 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0122-y.

Roche:2010:ICT

[RBF^{+}10] Ivana Roche, Dominique Besagni, Claire François, Marianne Hörlesberger, and Edgar Schiebel. Identification and characterisation of technological topics in the field of molecular biology. Scientometrics, 82(3):663–676, March 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0178-8.

Radicchi:2013:ABI

[RC13a] Filippo Radicchi and Claudio Castellano. Analysis of bibliometric indicators for individual scholars in a large data set. Scientometrics, 97(3):627–637, December 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1027-3.

Ruiz-Castillo:2013:RSE

[RC13b] Javier Ruiz-Castillo. The role of statistics in establishing the similarity of citation distributions in a static and a dynamic context. Scientometrics, 96(1):173–181, July 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0954-3.

Ruiz-Conde:2014:UIR

[RCCM14] Enar Ruiz-Conde and Aurora Calderón-Martínez. University institutional repositories: competitive environment and their role as communication media of scientific knowledge. Scientometrics, 98(2):1283–1299, February 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1159-5.

Roos:2014:BSP

[RCdJ^{+}14] Daniel Henrique Roos, Luciana Calabró, Sandra Lopes de Jesus, Diogo Onofre Souza, Nilda Vargas Barbosa, and João Batista Teixeira da Rocha. Brazilian scientific production in areas of biological sciences: a comparative study on the modalities of full doctorate in Brazil or abroad. Scientometrics, 98(1):415–427, January 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1017-5.

Repiso:2019:AAI

[RCETS19] Rafael Repiso, Antonio Castillo-Esparcia, and Daniel Torres-Salinas. Altmetrics, alternative indicators for Web of Science communication studies journals. Scientometrics, 119(2):941–958, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03070-7.

Rigby:2018:JPR

[RCJ18] J. Rigby, D. Cox, and K. Julian. Journal peer review: a bar or bridge? an analysis of a paper’s revision history and turnaround time, and the effect on citation. Scientometrics, 114(3):1087–1105, March 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2630-5;   https://link.springer.
com/content/pdf/10.1007/s11192-017-2630-5.pdf.

Reijnhoudt:2014:SEG

[RCN^{+}14] Linda Reijnhoudt, Rodrigo Costas, Ed Noyons, Katy Börner, and Andrea Scharnhorst. ‘Seed + expand’: a general methodology for detecting publication oeuvres of individual researchers. Scientometrics, 101(2):1403– 1417, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-014-1256-
0.pdf.

Ruocco:2013:EAC

[RD13] Giancarlo Ruocco and Cinzia Daraio. An empirical approach to compare the performance of heterogeneous academic fields. Scientometrics, 97(3):601–625, December 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1038-0.

Rotolo:2019:WDS

[RF19] Thomas Rotolo and Scott Frickel. When disasters strike environmental science: a case-control study of changes in scientific collaboration networks. Scientometrics, 120(1):301– 317, July 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03122-y;   http://
link.springer.com/content/pdf/10.1007/s11192-019-
03122-y.pdf.

Romo-Fernandez:2013:CWB

[RFGBMA13] Luz M. Romo-Fernández, Vicente P. Guerrero-Bote, and Félix Moya-Anegón. Co-word based thematic analysis of renewable energy (1990–2010). Scientometrics, 97(3):743– 765, December 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1009-5.

Ren:2012:EIS

[RG12] Quan’e Ren and Xuemei Gong. Evaluation index system for academic papers of humanities and social sciences. Scientometrics, 93(3):1047–1060, December 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0790-x.

Raban:2015:ETL

[RG15] Daphne R. Raban and Avishag Gordon. The effect of technology on learning research trends: a bibliometric analysis over five decades. Scientometrics, 105(1):665–681, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1690-7.

Russell:2018:CEL

[RG18] Jane M. Russell and María Victoria Guzmán. Connecting and empowering the Latin American research community in the global sphere: an introduction to the special issue on the metric study of regional science and technology. Scientometrics, 115(3):1439–1441, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2745-3;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2745-
3.pdf.

Robinson-Garcia:2014:WDU

[RGCM14] Nicolás Robinson-García and Clara Calero-Medina. What do university rankings by fields rank? Exploring discrepancies between the organizational structure of universities and bibliometric classifications. Scientometrics, 98(3): 1955–1970, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1157-7.

Rodrigues:2017:OSP

[RGdCMM17] Lucas Oliveira Rodrigues, Marcos Martins Gouvêa, Flávia Ferreira de Carvalho Marques, and Samanta Cardozo Mourão. Overview of the scientific production in the pharmacy area in Brazil: profile and productivity of researchers granted with fellowships by the National Council for Scientific and Technological Development. Scientometrics, 110(3):1157–1171, March 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2210-0.

Reyes-Gonzalez:2016:UCA

[RGGBV16] Leonardo Reyes-Gonzalez, Claudia N. Gonzalez-Brambila, and Francisco Veloso. Using co-authorship and citation analysis to identify research groups: a new way to assess performance. Scientometrics, 108(3):1171–1191, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2029-8.

Rahman:2016:MMB

[RGLE16] A. I. M. Jakaria Rahman, Raf Guns, Loet Leydesdorff, and Tim C. E. Engels. Measuring the match between evaluators and evaluees: cognitive distances between panel members and research groups at the journal level. Scientometrics, 109(3):1639–1663, December 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2132-x.

Robinson-Garcia:2014:IIN

[RGTSLCH14] Nicolás Robinson-García, Daniel Torres-Salinas, Emilio Delgado López-Cózar, and Francisco Herrera. An insight into the importance of national university rankings in an international context: the case of the I-UGR rankings of Spanish universities. Scientometrics, 101(2):1309–1324, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1263-1.

Rousseau:2018:UCI

[RH18] Ronald Rousseau and Xiaojun Hu. Under-cited influential work by Eugene Garfield. Scientometrics, 114(2):651–657, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2530-8.

Rhaiem:2017:MDA

[Rha17] Mehdi Rhaiem. Measurement and determinants of academic research efficiency: a systematic review of the evidence. Scientometrics, 110(2):581–615, February 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2173-1.

Russell:2016:CDM

[RHGKD16] Jane M. Russell, Yoscelina Hernández-García, and Mina Kleiche-Dray. Collaboration dynamics of Mexican research in chemistry and its relationship with communication patterns. Scientometrics, 109(1):283–316, October 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2069-0.

Rabiei:2017:UTM

[RHMH17] Mohammad Rabiei, Seyyed-Mahdi Hosseini-Motlagh, and Abdorrahman Haeri. Using text mining techniques for identifying research gaps and priorities: a case study of the environmental science in Iran. Scientometrics, 110(2):815–842, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2195-8.

Ricker:2015:NAP

[Ric15] Martin Ricker. A numerical algorithm with preference statements to evaluate the performance of scientists. Scientometrics, 103(1):191–212, April 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-014-
1521-2.pdf.

Ricker:2017:LEA

[Ric17] Martin Ricker. Letter to the Editor: About the quality and impact of scientific articles. Scientometrics, 111(3): 1851–1855, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-017-2374-
2.pdf.

Rigby:2013:LIP

[Rig13] John Rigby. Looking for the impact of peer review: does count of funding acknowledgements really predict research impact? Scientometrics, 94(1):57–73, January 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-012-0779-5.pdf.

Rigby:2014:HDD

[RJ14] J. Rigby and K. Julian. On the horns of a dilemma: does more funding for research lead to more research or a waste of resources that calls for optimization of researcher portfolios? An analysis using funding acknowledgement data. Scientometrics, 101(2):1067–1075, November 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1259-x.

Rodriguez:2015:NMS

[RKT^{+}15] Andrew Rodriguez, Byunghoon Kim, Mehmet Turkoz, Jae-Min Lee, Byoung-Youl Coh, and Myong K. Jeong. New multi-stage similarity measure for calculation of pairwise patent similarity in a patent citation network. Scientometrics, 103(2):565–581, May 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
015-1531-8.

Rexha:2018:AID

[RKZK18] Andi Rexha, Mark Kröll, Hermann Ziak, and Roman Kern. Authorship identification of documents with high content similarity. Scientometrics, 115(1):223–237, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2661-6;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2661-
6.pdf.

Roe:2014:SEN

[RLW14] Philip Roe, Grant Lewison, and Richard Webber. The sex and ethnicity or national origins of researchers in astronomy and oncology in four countries, 2006–2007 and 2011–2012. Scientometrics, 100(1):287–296, July 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1223-1.

Rafols:2010:DNC

[RM10] Ismael Rafols and Martin Meyer. Diversity and network coherence as indicators of interdisciplinarity: case studies in bionanoscience. Scientometrics, 82(2):263–287, February 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0041-y.

Rashidi:2018:ISA

[RM18] Naser Rashidi and Hussein Meihami. Informetrics of Scientometrics abstracts: a rhetorical move analysis of the research abstracts published in Scientometrics journal. Scientometrics, 116(3):1975–1994, September 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2795-6.

Ramos:2012:CBP

[RMA12] Marcelo Alves Ramos, Joabe Gomes Melo, and Ulysses Paulino Albuquerque. Citation behavior in popular scientific papers: what is behind obscure citations? The case of ethnobotany. Scientometrics, 92(3):711–719, September 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0662-4.

Ragone:2013:PRC

[RMCM13] Azzurra Ragone, Katsiaryna Mirylenka, Fabio Casati, and Maurizio Marchese. On peer review in computer science: analysis of its effectiveness and suggestions for improvement. Scientometrics, 97(2):317–356, November 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1002-z.

Reis:2017:MMI

[RMdO17] Thalita Laua Reis, Maria Augusta Siqueira Mathias, and Otavio Jose de Oliveira. Maturity models: identifying the state-of-the-art and the scientific gaps from a bibliometric study. Scientometrics, 110(2):643–672, February 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2182-0.

Rostami:2014:ECT

[RMH14] Fatemeh Rostami, Asghar Mohammadpoorasl, and Mohammad Hajizadeh. The effect of characteristics of title on citation rates of articles. Scientometrics, 98(3):2007–2010, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1118-1.

Rodriguez-Navarro:2019:PEF

[RNB19] Alonso Rodríguez-Navarro and Ricardo Brito. Probability and expected frequency of breakthroughs: basis and use of a robust method of research assessment. Scientometrics, 119 (1):213–235, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03022-1.

Richter:2019:BTM

[RNF19] Amy Richter, Kelvin Tsun Wai Ng, and Bahareh Fallah. Bibliometric and text mining approaches to evaluate landfill design standards. Scientometrics, 118(3):1027–1049, March 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03011-4.

Ruiz-Navas:2018:CLQ

[RNM18] Santiago Ruiz-Navas and Kumiko Miyazaki. A complement to lexical query’s search-term selection for emerging technologies: the case of “big data”. Scientometrics, 117(1): 141–162, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2857-9.

Rodela:2016:UDA

[Rod16] Romina Rodela. On the use of databases about research performance: comments on Karlovcec and Mladenić (2015) and others using the SICRIS database. Scientometrics, 109(3):2151–2157, December 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2073-4.

Rodriguez:2017:DIC

[Rod17] Jorge Mañana Rodríguez. Disciplinarity and interdisciplinarity in citation and reference dimensions: knowledge importation and exportation taxonomy of journals. Scientometrics, 110(2):617–642, February 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2190-0.

Rousseau:2011:CMC

[Rou11] Ronald Rousseau. Comments on the modified collaborative coefficient. Scientometrics, 87(1):171–174, April 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0300-y.

Rousseau:2012:CHT

[Rou12a] Ronald Rousseau. Comments on “A Hirsch-type index of coauthor partnership ability”. Scientometrics, 91(1):309–310, April 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-011-0606-4. See

[Sch12b].

Rousseau:2012:UJI

[Rou12b] Ronald Rousseau. Updating the journal impact factor or total overhaul? Scientometrics, 92(2):413–417, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0649-1.

Rousseau:2018:RRH

[Rou18] Ronald Rousseau. The repeat rate: from Hirschman to Stirling. Scientometrics, 116(1):645–653, July 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2724-8.

Rousseau:2019:BRC

[Rou19] Ronald Rousseau. Balassa = revealed competitive advantage = activity. Scientometrics, 121(3):1835–1836, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03273-y.

Ronda-Pupo:2017:CBI

[RP17a] Guillermo Armando Ronda-Pupo. The citation-based impact of complex innovation systems scales with the size of the system. Scientometrics, 112(1):141–151, July 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Ronda-Pupo:2017:EDT

[RP17b] Guillermo Armando Ronda-Pupo. The effect of document types and sizes on the scaling relationship between citations and co-authorship patterns in management journals. Scientometrics, 110(3):1191–1207, March 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2231-8.

Rodriguez-Prieto:2019:DRS

[RPAMR19] Oscar Rodriguez-Prieto, Lourdes Araujo, and Juan Martinez-Romo. Discovering related scientific literature beyond semantic similarity: a new co-citation approach. Scientometrics, 120(1):105–127, July 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03125-9.

Ronda-Pupo:2015:RAC

[RPDCRVRP15] Guillermo Armando Ronda-Pupo, Carlos Díaz-Contreras, Guillermo Ronda-Velázquez, and Jorge Carlos Ronda-Pupo. The role of academic collaboration in the impact of Latin– American research on management. Scientometrics, 102(2): 1435–1454, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1486-1.

Ronda-Pupo:2010:DSC

[RPGM10] Guillermo Armando Ronda-Pupo and Luis Ángel Guerras-Martín. Dynamics of the scientific community network within the strategic management field through the strategic management journal 1980–2009: the role of cooperation. Scientometrics, 85(3):821–848, December 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0287-4.

Ronda-Pupo:2016:CNK

[RPGM16] Guillermo Armando Ronda-Pupo and Luis Ángel Guerras-Martín. Collaboration network of knowledge creation and dissemination on management research: ranking the leading institutions. Scientometrics, 107(3):917–939, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-016-1924-3.

Ronda-Pupo:2016:SRB

[RPK16] Guillermo Armando Ronda-Pupo and J. Sylvan Katz. The scaling relationship between citation-based performance and international collaboration of Cuban articles in natural sciences. Scientometrics, 107(3):1423–1434, June 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1939-9.

Ronda-Pupo:2017:SRB

[RPK17] Guillermo Armando Ronda-Pupo and J. Sylvan Katz. The scaling relationship between degree centrality of countries and their citation-based performance on Management Information Systems. Scientometrics, 112(3):1285– 1299, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2459-y.

Ronda-Pupo:2018:PLR

[RPK18] Guillermo Armando Ronda-Pupo and J. Sylvan Katz. The power law relationship between citation impact and multiauthorship patterns in articles in Information Science & Library Science journals. Scientometrics, 114(3):919–932, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2612-7.

Roessner:2013:VII

[RPNC13] David Roessner, Alan L. Porter, Nancy J. Nersessian, and Stephen Carley. Validating indicators of interdisciplinarity: linking bibliometric measures to studies of engineering research labs. Scientometrics, 94(2):439–468, February 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0872-9.

Ronda-Pupo:2018:ERG

[RPP18] Guillermo Armando Ronda-Pupo and Thong Pham. The evolutions of the rich get richer and the fit get richer phenomena in scholarly networks: the case of the strategic management journal. Scientometrics, 116(1):363–383, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2761-3.

Ramos-Pardo:2017:PET

[RPSA17] Francisco Javier Ramos-Pardo and Pablo Sánchez-Antolín. Production of educational theory doctoral theses in Spain (2001–2015). Scientometrics, 112(3):1615–1630, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2435-6.

Reinstaller:2017:UPA

[RR17] Andreas Reinstaller and Peter Reschenhofer. Using PageRank in the analysis of technological progress through patents: an illustration for biotechnological inventions. Scientometrics, 113(3):1407–1438, December 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2549-x.

Ribeiro:2010:MST

[RRBA10] Leonardo Costa Ribeiro, Ricardo Machado Ruiz, Américo Tristão Bernardes, and Eduardo Motta Albuquerque. Matrices of science and technology interactions and patterns of structured growth: implications for development. Scientometrics, 83(1):55–75, April 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
009-0020-3.

Reiners:2016:DUE

[RRL16] William A. Reiners, Derek S. Reiners, and Jeffrey A. Lockwood. Differentiation of U.S. ecologists into professional guilds based on professional traits. Scientometrics, 106(1): 281–298, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1771-7.

Rey-Rocha:2015:ODB

[RRLNAG15] Jesús Rey-Rocha, Irene López-Navarro, and M. Teresa Antonio-García. Opening doors to basic-clinical collaboration and translational research will improve researchers’ performance. Scientometrics, 105(3):2057–2069, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1754-8.

Ruiz-Rosero:2019:SSS

[RRRGVD19] Juan Ruiz-Rosero, Gustavo Ramirez-Gonzalez, and Jesus Viveros-Delgado. Software survey: ScientoPy, a scientometric tool for topics trend analysis in scientific publications. Scientometrics, 121(2):1165–1188, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03213-w.

Ribeiro:2018:GPN

[RRSA18] Leonardo Costa Ribeiro, Márcia Siqueira Rapini, Leandro Alves Silva, and Eduardo Motta Albuquerque. Growth patterns of the network of international collaboration in science. Scientometrics, 114(1):159–179, January 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2573-x.

Rodriguez:2012:ABT

[RS12] V. Rodriguez and A. Soeparwata. ASEAN benchmarking in terms of science, technology, and innovation from 1999 to 2009. Scientometrics, 92(3):549–573, September 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-011-0603-7.pdf.

Rodriguez-Sanchez:2014:EGB

[RSGFV14] Rosa Rodriguez-Sánchez, J. A. García, and J. Fdez-Valdivia. Evolutionary games between subject categories. Scientometrics, 101(1):869–888, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1255-1.

Rodriguez-Sanchez:2018:EDI

[RSGFV18] Rosa Rodriguez-Sánchez, J. A. García, and J. Fdez-Valdivia. Editorial decisions with informed and uninformed reviewers. Scientometrics, 117(1):25–43, October 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2875-7.

Righi:2017:MPR

[RT17] Simone Righi and Károly Takács. The miracle of peer review and development in science: an agentbased model. Scientometrics, 113(1):587–607, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2244-y;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2244-
y.pdf.

Rost:2017:SNA

[RTP17] Katja Rost, Thorsten Teichert, and Alan Pilkington. Social network analytics for advanced bibliometrics: referring to actor roles of management journals instead of journal rankings. Scientometrics, 112(3):1631–1657, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2441-8.

Ribeiro:2018:CRC

[RV18a] M. D. Ribeiro and S. M. R. Vasconcelos. Correction to: Retractions covered by Retraction Watch in the 2013– 2015 period: prevalence for the most productive countries. Scientometrics, 114(2):735, February 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2653-6;   https://link.springer.

com/content/pdf/10.1007/s11192-018-2653-6.pdf. See

[RV18b].

Ribeiro:2018:RCR

[RV18b] M. D. Ribeiro and S. M. R. Vasconcelos. Retractions covered by Retraction Watch in the 2013–2015 period: prevalence for the most productive countries. Scientometrics, 114(2):719–734, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2621-6. See correction [RV18a].

Ramos-Vielba:2010:MUI

[RVFEdlM10] Irene Ramos-Vielba, Manuel Fernández-Esquinas, and Elena Espinosa de-los Monteros. Measuring universityindustry collaboration in a regional innovation system. Scientometrics, 84(3):649–667, September 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0113-z.

Ravallion:2011:MSI

[RW11] Martin Ravallion and Adam Wagstaff. On measuring scholarly influence by citations. Scientometrics, 88(1):321–337, July 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0375-0.

Rochon:2015:PEA

[RWG^{+}15] Paula A. Rochon, Wei Wu, Jerry H. Gurwitz, Sunila R. Kalkar, Joel Thomson, and Sudeep S. Gill. Prospective evaluation of the accessibility of Internet references in leading general medical journals. Scientometrics, 102(2):1375–1384, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1489-y.

Radosevic:2014:TGS

[RY14] Slavo Radosevic and Esin Yoruk. Are there global shifts in the world science base? Analysing the catching up and falling behind of world regions. Scientometrics, 101(3):1897–1924, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-014-1344-1.pdf.

Ryan:2016:VIA

[Rya16] James C. Ryan. A validation of the individual annual h-index (hIa): application of the hIa to a qualitatively and quantitatively different sample. Scientometrics, 109(1):577–590, October 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-1972-8.

Raj:2012:RMI

[RZ12] R. G. Raj and A. N. Zainab. Relative measure index: a metric to measure the quality of journals. Scientometrics, 93(2): 305–317, November 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0675-z.

Sakr:2011:DDR

[SA11] Sherif Sakr and Mohammad Alomari. A decade of database research publications: a look inside. Scientometrics, 88(2): 521–533, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0385-y.

Sakr:2012:DDC

[SA12] Sherif Sakr and Mohammad Alomari. A decade of database conferences: a look inside the program committees. Scientometrics, 91(1):173–184, April 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0530-7.

Shideler:2016:MSJ

[SA16] Geoffrey S. Shideler and Rafael J. Araújo. Measures of scholarly journal quality are not universally applicable to determining value of advertised annual subscription price. Scientometrics, 107(3):963–973, June 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1943-0.

Shideler:2017:RIM

[SA17] Geoffrey S. Shideler and Rafael J. Araújo. Reviewer interest in a manuscript may predict its future citation potential. Scientometrics, 113(2):1171–1176, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2492-x.

Saad:2010:AIE

[Saa10] Gad Saad. Applying the h-index in exploring bibliometric properties of elite marketing scholars. Scientometrics, 83 (2):423–433, May 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0069-z.

Safon:2013:WDG

[Saf13] Vicente Safón. What do global university rankings really measure? The search for the X factor and the X entity. Scientometrics, 97(2):223–244, November 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0986-8.

Safon:2019:IRR

[Saf19] Vicente Safón. Inter-ranking reputational effects: an analysis of the Academic Ranking of World Universities (ARWU) and the Times Higher Education World University Rankings (THE) reputational relationship. Scientometrics, 121(2): 897–915, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03214-9.

Sahoo:2016:ARP

[Sah16] Shaon Sahoo. Analyzing research performance: proposition of a new complementary index. Scientometrics, 108 (2):489–504, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1988-0.

Sakai:2019:WPR

[Sak19] Daisuke Sakai. Who is peer reviewed? Comparing publication patterns of peer-reviewed and non-peer-reviewed papers in Japanese political science. Scientometrics, 121(1):65–80, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03197-7.

Salimi:2017:QAS

[Sal17] Negin Salimi. Quality assessment of scientific outputs using the BWM. Scientometrics, 112(1):195–213, July 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-017-2284-3.pdf.

Sangwal:2012:APN

[San12a] Keshra Sangwal. Application of progressive nucleation mechanism for the citation behavior of individual papers of different authors. Scientometrics, 92(3):643–655, September 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-011-0564-x.pdf.

Sangwal:2012:AIP

[San12b] Keshra Sangwal. On the age-independent publication index. Scientometrics, 91(3):1053–1058, June 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-012-0628-6.pdf.

Sangwal:2012:RBC

[San12c] Keshra Sangwal. On the relationship between citations of publication output and Hirsch index h of authors: conceptualization of tapered Hirsch index h_{T}, circular citation area radius R and citation acceleration a. Scientometrics, 93(3):987–1004, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-012-0805-7.pdf.

Sangwal:2012:PNM

[San12d] Keshra Sangwal. Progressive nucleation mechanism for the growth behavior of items and its application to cumulative papers and citations of individual authors. Scientometrics, 92(3):575–591, September 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-011-0610-8.pdf.

Sangwal:2013:SCR

[San13] Keshra Sangwal. Some citation-related characteristics of scientific journals published in individual countries. Scientometrics, 97(3):719–741, December 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-013-1053-1.pdf.

Sandnes:2018:DNA

[San18] Frode Eika Sandnes. Do Norwegian academics who publish more earn higher salaries? Scientometrics, 115(1):263–281, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2639-4.

Sangam:2018:GHL

[SAPR18] S. L. Sangam, Uma B. Arali, C. G. Patil, and Ronald Rousseau. Growth of the hepatitis literature over the period 1976–2015: What can the relative priority index teach us? Scientometrics, 115(1):351–368, April 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2668-z.

Shehatta:2019:ICS

[SAR19] Ibrahim Shehatta and Abdullah M. Al-Rubaish. Impact of country self-citations on bibliometric indicators and ranking of most productive countries. Scientometrics, 120(2): 775–791, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03139-3.

Soriano:2018:BAI

[SÁV18] Alba Santa Soriano, Carolina Lorenzo Álvarez, and Rosa María Torres Valdés. Bibliometric analysis to identify an emerging research area: Public relations intelligence — a challenge to strengthen technological observatories in the network society. Scientometrics, 115(3):1591–1614, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2651-8.

Shrivats:2014:FTI

[SB14] S. Varun Shrivats and Sujit Bhattacharya. Forecasting the trend of international scientific collaboration. Scientometrics, 101(3):1941–1954, December 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1364-x.

Saritas:2015:EUF

[SB15] Ozcan Saritas and Serhat Burmaoglu. The evolution of the use of foresight methods: a scientometric analysis of global FTA research output. Scientometrics, 105(1):497–508, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1671-x.

Seeber:2017:DSB

[SB17] Marco Seeber and Alberto Bacchelli. Does single blind peer review hinder newcomers? Scientometrics, 113(1):567–585, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2264-7;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2264-
7.pdf.

Stopar:2019:DCC

[SB19] Karmen Stopar and Tomaz Bartol. Digital competences, computer skills and information literacy in secondary education: mapping and visualization of trends and concepts. Scientometrics, 118(2):479–498, February 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2990-5.

Said:2019:MNL

[SBA^{+}19] Anwar Said, Timothy D. Bowman, Rabeeh Ayaz Abbasi, Naif Radi Aljohani, Saeed-Ul Hassan, and Raheel Nawaz. Mining network-level properties of Twitter altmetrics data. Scientometrics, 120(1):217–235, July 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03112-0.

Schlagberger:2016:WID

[SBB16] Elisabeth Maria Schlagberger, Lutz Bornmann, and Johann Bauer. At what institutions did Nobel laureates do their prize-winning work? An analysis of biographical information on Nobel laureates from 1994 to 2014. Scientometrics, 109(2):723–767, November 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-2059-2.pdf.

Sanyal:2019:EAS

[SBD^{+}19] Debarshi Kumar Sanyal, Plaban Kumar Bhowmick, Partha Pratim Das, Samiran Chattopadhyay, and T. Y. S. S. Santosh. Enhancing access to scholarly publications with surrogate resources. Scientometrics, 121(2):1129–1164, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03227-4.

Small:2019:CCN

[SBK19] Henry Small, Kevin W. Boyack, and Richard Klavans. Citations and certainty: a new interpretation of citation counts. Scientometrics, 118(3):1079–1092, March 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03016-z.

Squazzoni:2017:SPR

[SBM17] Flaminio Squazzoni, Elise Brezis, and Ana Marusić. Scientometrics of peer review. Scientometrics, 113(1):501–502, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2518-4;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2518-
4.pdf.

Schmoch:2019:IAS

[SBSR19] Ulrich Schmoch, Bernd Beckert, and Petra Schaper-Rinkel. Impact assessment of a support programme of science-based emerging technologies. Scientometrics, 118(3):1141–1161, March 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-03002-x.

Singh:2015:SMR

[SBSU15] Vivek Kumar Singh, Sumit Kumar Banshal, Khushboo Singhal, and Ashraf Uddin. Scientometric mapping of research on ‘Big Data’. Scientometrics, 105(2):727– 741, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1729-9.

Scarpa:2018:INA

[SBT18] Federico Scarpa, Vincenzo Bianco, and Luca A. Tagliafico. The impact of the national assessment exercises on self-citation rate and publication venue: an empirical investigation on the engineering academic sector in Italy. Scientometrics, 117(2):997–1022, November 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2913-5.

Shin:2010:MAA

[SC10] Jung Cheol Shin and William K. Cummings. Multilevel analysis of academic publishing across disciplines: research preference, collaboration, and time on research. Scientometrics, 85(2):581–594, November 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0236-2.

Sugimoto:2013:CGT

[SC13] Cassidy R. Sugimoto and Blaise Cronin. Citation gamesmanship: testing for evidence of ego bias in peer review. Scientometrics, 95(3):851–862, June 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0845-z.

Sahin:2018:SPC

[SC18] Köksal Sahin and Gökçe Candan. Scientific productivity and cooperation in Turkic world: a bibliometric analysis. Scientometrics, 115(3):1199–1229, June 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2730-x.

Sanz-Casado:2016:UCS

[SCGZR16] Elias Sanz-Casado, Carlos García-Zorita, and Ronald Rousseau. Using h-cores to study the most-cited articles of the twenty-first century. Scientometrics, 108(1):243–261, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1956-8.

Sanz-Casado:2013:RER

[SCGZSL^{+}13] Elias Sanz-Casado, J. Carlos Garcia-Zorita, Antonio Eleazar Serrano-López, Birger Larsen, and Peter Ingwersen. Renewable energy research 1995–2009: a case study of wind power research in EU, Spain, Germany and Denmark. Scientometrics, 95(1):197–224, April 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0825-3.

Schubert:2010:RBH

[Sch10a] András Schubert. A reference-based Hirschian similarity measure for journals. Scientometrics, 84(1):133–147, July 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0072-4.

Schultz:2010:THB

[Sch10b] David M. Schultz. Are three heads better than two? How the number of reviewers and editor behavior affect the rejection rate. Scientometrics, 84(2):277–292, August 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0084-0.

Schubert:2011:AVP

[Sch11a] Torben Schubert. Assessing the value of patent portfolios: an international country comparison. Scientometrics, 88(3): 787–804, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0454-2.

Schultz:2011:RRM

[Sch11b] David M. Schultz. Rejection rates for multiple-part manuscripts. Scientometrics, 86(2):251–259, February 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0258-9.

Schiebel:2012:VRF

[Sch12a] Edgar Schiebel. Visualization of research fronts and knowledge bases by three-dimensional areal densities of bibliographically coupled publications and co-citations. Scientometrics, 91(2):557–566, May 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0626-8.

Schubert:2012:HTI

[Sch12b] András Schubert. A Hirsch-type index of co-author partnership ability. Scientometrics, 91(1):303–308, April 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0559-7. See comments [Rou12a].

Schreiber:2013:HMD

[Sch13a] Michael Schreiber. How much do different ways of calculating percentiles influence the derived performance indicators? A case study. Scientometrics, 97(3):821–829, December 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-0984-x.

Schubert:2013:MSB

[Sch13b] András Schubert. Measuring the similarity between the reference and citation distributions of journals. Scientometrics, 96(1):305–313, July 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0889-0.

Schubert:2014:HRI

[Sch14a] András Schubert. How to rank if you must: two useful guiding books. Scientometrics, 98(1):763–765, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-0998-4.

Schubert:2014:SRF

[Sch14b] András Schubert. Sentences to remember from the first 100 volumes of the journal Scientometrics. Scientometrics, 100(1):1–13, July 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-014-1282-
y.pdf.

Schubert:2014:TSE

[Sch14c] Torben Schubert. Are there scale economies in scientific production? On the topic of locally increasing returns to scale. Scientometrics, 99(2):393–408, May 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1207-1.

Schneider:2015:NHS

[Sch15a] Jesper W. Schneider. Null hypothesis significance tests. A mix-up of two different theories: the basis for widespread confusion and numerous misinterpretations. Scientometrics, 102(1):411–432, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-

014-1251-5. See commentary [Wu18].

Schubert:2015:RI

[Sch15b] András Schubert. Rescaling the h-index. Scientometrics, 102(2):1647–1653, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
014-1373-9.

Schubert:2015:XCH

[Sch15c] András Schubert. X-centage: a Hirsch-inspired indicator for distributions of percentage-valued variables and its use for measuring heterodisciplinarity. Scientometrics, 102(1): 307–313, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1281-z.

Schulz:2016:UMC

[Sch16] Jan Schulz. Using Monte Carlo simulations to assess the impact of author name disambiguation quality on different bibliometric analyses. Scientometrics, 107(3):1283–1298, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1892-7.

Schubert:2017:PPC

[Sch17a] András Schubert. Power positions in cardiology publications. Scientometrics, 112(3):1721–1743, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2448-1.

Schubert:2017:SDP

[Sch17b] András Schubert. Science dynamics: from production to evaluation — two recent books. Scientometrics, 112(2): 1141–1145, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2408-9.

Schneider:2018:NSL

[Sch18a] Jesper W. Schneider. NHST is still logically flawed. Scientometrics, 115(1):627–635, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11192-018-2655-4. See commentary [Pat18] and response [Sch18b].

Schneider:2018:RCN

[Sch18b] Jesper W. Schneider. Response to commentary on “Is NHST logically flawed”. Scientometrics, 116(3):2193– 2194, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2818-3. See

[Pat18, Sch18a].

Su:2012:PLP

[SCL12] Hsin-Ning Su, Carey Ming-Li Chen, and Pei-Chun Lee. Patent litigation precaution method: analyzing characteristics of US litigated and non-litigated patents from 1976 to 2010. Scientometrics, 92(1):181–195, July 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0716-7.

Sui:2015:BAR

[SCLC15] Xiaoyun Sui, Yongxia Chen, Zhi Lu, and Yifeng Chen. A bibliometric analysis of research papers related to the Mekong River. Scientometrics, 105(1):419–434, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1683-6.

Shiau:2013:CCC

[SD13] Wen-Lung Shiau and Yogesh K. Dwivedi. Citation and co-citation analysis to identify core and emerging knowledge in electronic commerce research. Scientometrics, 94 (3):1317–1337, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0807-5.

Sun:2018:ITS

[SD18] Xiaoling Sun and Kun Ding. Identifying and tracking scientific and technological knowledge memes from citation networks of publications and patents. Scientometrics, 116(3):1735–1748, September 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2836-1.

Stopar:2016:CAM

[SDEB16] Karmen Stopar, Damjana Drobne, Klemen Eler, and Tomaz Bartol. Citation analysis and mapping of nanoscience and nanotechnology: identifying the scope and interdisciplinarity of research. Scientometrics, 106(2):563–581, February 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1797-x.

Salie:2019:SBO

[SdJDD19] Faatiema Salie, Kylie de Jager, Carsten Dreher, and Tania S. Douglas. The scientific base for orthopaedic device development in South Africa: spatial and sectoral evolution of knowledge development. Scientometrics, 119(1):31–54, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03041-y.

Stevic:2019:NHM

[SDP^{+}19] Zeljko Stević, Irena Dalić, Dragan Pamucar, Zdravko Nunić, Slavko Vesković, Marko Vasiljević, and Ilija Tanackov. A new hybrid model for quality assessment of scientific conferences based on Rough BWM and SERVQUAL. Scientometrics, 119(1):1–30, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03032-z.

Silaghi-Dumitrescu:2014:SAR

[SDS14a] Radu Silaghi-Dumitrescu and Augusta Sabau. Scientometric analysis of relative performance in a key university in Romania. Scientometrics, 99(2):463–474, May 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1232-8.

Su:2014:DAV

[SDS14b] Xinning Su, Sanhong Deng, and Si Shen. The design and application value of the Chinese social Science Citation Index. Scientometrics, 98(3):1567–1582, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0921-4.

Sonne:2019:RPD

[SDS19] James W. H. Sonne, Nicole T. Dawson, and Gerald V. Smith. Research productivity of Doctor of Physical Therapy faculty promoted in the Western United States. Scientometrics, 119(2):707–719, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03042-x;   http://link.springer.com/content/
pdf/10.1007/s11192-019-03042-x.pdf.

Shiau:2015:SCM

[SDT15] Wen-Lung Shiau, Yogesh K. Dwivedi, and Chia-Han Tsai. Supply chain management: exploring the intellectual structure. Scientometrics, 105(1):215–230, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1680-9.

Sotudeh:2018:SOA

[SE18] Hajar Sotudeh and Zohreh Estakhr. Sustainability of open access citation advantage: the case of Elsevier’s authorpays hybrid open access journals. Scientometrics, 115(1): 563–576, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2663-4.

Sampaio:2017:NAS

[SFBS17] Ricardo B. Sampaio, Bruna P. F. Fonseca, Ashwin Bahulkar, and Boleslaw K. Szymanski. Network analysis to support public health: evolution of collaboration among leishmaniasis researchers. Scientometrics, 111(3):2001–2021, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Schmoch:2016:EWC

[SFM16] Ulrich Schmoch, Habib M. Fardoun, and Abdulfattah S. Mashat. Establishing a world–class university in Saudi Arabia: intended and unintended effects. Scientometrics, 109(2):1191–1207, November 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2089-9.

Shirakawa:2012:GCT

[SFNO12] Nobuyuki Shirakawa, Takao Furukawa, Minoru Nomura, and Kumi Okuwada. Global competition and technological transition in electrical, electronic, information and communication engineering: quantitative analysis of periodicals and conference proceedings of the IEEE. Scientometrics, 91 (3):895–910, June 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0566-8.

Sebo:2019:FAP

[SFR^{+}19] Paul Sebo, Jean Pascal Fournier, Claire Ragot, Pierre-Henri Gorioux François R. Herrmann, and Hubert Maisonneuve. Factors associated with publication speed in general medical journals: a retrospective study of bibliometric data. Scientometrics, 119(2):1037–1058, May 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03061-8;   http://link.springer.
com/content/pdf/10.1007/s11192-019-03061-8.pdf.

Schloegl:2010:CCU

[SG10] Christian Schloegl and Juan Gorraiz. Comparison of citation and usage indicators: the case of oncology journals. Scientometrics, 82(3):567–580, March 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0172-1.

Sun:2016:EDS

[SG16] Yutao Sun and Seamus Grimes. The emerging dynamic structure of national innovation studies: a bibliometric analysis. Scientometrics, 106(1):17–40, January 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1778-0.

Schlogl:2014:CDC

[SGG^{+}14] Christian Schlögl, Juan Gorraiz, Christian Gumpenberger, Kris Jack, and Peter Kraker. Comparison of downloads, citations and readership data for two information systems journals. Scientometrics, 101(2):1113–1128, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1365-9.

Safaei:2016:SUM

[SGM^{+}16] Mohammad Reza Safaei, Marjan Goodarzi, Omid Mahian, Mahidzal Dahari, and Somchai Wongwises. A survey of using multiple affiliations by scholars in scientific articles. Scientometrics, 107(1):317–318, April 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-1875-8;   http://link.
springer.com/article/10.1007/s11192-016-1875-8.

Smith:2015:STR

[SGN15] Charles H. Smith, Patrick Georges, and Ngoc Nguyen. Statistical tests for ‘related records’ search results. Scientometrics, 105(3):1665–1677, December 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1610-x.

Sarigol:2017:QEE

[SGSS17] Emre Sarigöl, David Garcia, Ingo Scholtes, and Frank Schweitzer. Quantifying the effect of editor–author relations on manuscript handling times. Scientometrics, 113 (1):609–631, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2309-y;
http://link.springer.com/content/pdf/10.1007/s11192-
017-2309-y.pdf.

Sotudeh:2015:CAA

[SGY15] Hajar Sotudeh, Zahra Ghasempour, and Maryam Yaghtin. The citation advantage of author-pays model: the case of Springer and Elsevier OA journals. Scientometrics, 104(2): 581–608, August 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1607-5.

Santos:2015:GGS

[SH15a] João M. Santos and Hugo Horta. The generational gap of science: a dynamic cluster analysis of doctorates in an evolving scientific system. Scientometrics, 104(1):381–406, July 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1558-x.

Sarwar:2015:BAS

[SH15b] Raheem Sarwar and Saeed-Ul Hassan. A bibliometric assessment of scientific productivity and international collaboration of the Islamic World in science and technology (S&T) areas. Scientometrics, 105(2):1059–1077, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1718-z.

Sun:2015:CSO

[SH15c] Zhu-Mei Sun and Wei-Na Hua. A comparative study of ocean engineering research between China and the world. Scientometrics, 105(1):51–63, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1670-y.

Sotudeh:2018:CDP

[SH18] Hajar Sotudeh and Mojgan Houshyar. Comparing discrimination powers of text and citation-based context types. Scientometrics, 114(1):229–251, January 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2566-9.

Safder:2019:BEI

[SH19] Iqra Safder and Saeed-Ul Hassan. Bibliometric-enhanced information retrieval: a novel deep feature engineering approach for algorithm searching from full-text publications. Scientometrics, 119(1):257–277, April 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03025-y.

Shapiro:2012:RIK

[Sha12] Matthew A. Shapiro. Receiving information at Korean and Taiwanese universities, industry, and GRIs. Scientometrics, 90(1):289–309, January 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0501-z.

Schoen:2014:PNG

[SHB14] Anja Schoen, Dominik Heinisch, and Guido Buenstorf. Playing the ‘name game’ to identify academic patents in Germany. Scientometrics, 101(1):527–545, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-014-1400-x.

Shibayama:2011:DAR

[Shi11] Sotaro Shibayama. Distribution of academic research funds: a case of Japanese national research grant. Scientometrics, 88(1):43–60, July 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0392-z.

Shirabe:2014:ISC

[Shi14] Masashi Shirabe. Identifying SCI covered publications within non-patent references in U.S. utility patents. Scientometrics, 101(2):999–1014, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1293-8.

Song:2014:ATE

[SHK14] Min Song, Go Eun Heo, and Su Yeon Kim. Analyzing topic evolution in bioinformatics: investigation of dynamics of the field with conference data in DBLP. Scientometrics, 101 (1):397–428, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1246-2.

Song:2015:ILA

[SHL15] Min Song, Go Eun Heo, and Dahee Lee. Identifying the landscape of Alzheimer’s disease research with network and content analysis. Scientometrics, 102(1):905–927, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1372-x.

Schiebel:2010:ADM

[SHR^{+}10] Edgar Schiebel, Marianne Hörlesberger, Ivana Roche, Claire François, and Dominique Besagni. An advanced diffusion model to identify emergent research issues: the case of optoelectronic devices. Scientometrics, 83(3):765–781, June 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0137-4.

Seo:2015:TAA

[SHS15] Su Jin Seo, Eun Jin Han, and So Young Sohn. Trend analysis of academic research and technical development pertaining to gas hydrates. Scientometrics, 105(2):905– 920, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1709-0.

Shu:2017:CDC

[Shu17] Fei Shu. Comment to: Does China need to rethink its metrics- and citation-based research rewards policies? Scientometrics, 113(2):1229–1231, November 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2504-x. See [dS17a].

Sohrabi:2017:EKR

[SI17] Babak Sohrabi and Hamideh Iraj. The effect of keyword repetition in abstract and keyword frequency per journal in predicting citation counts. Scientometrics, 110(1):243–251, January 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2161-5.

Siler:2013:CCI

[Sil13] Kyle Siler. Citation choice and innovation in science studies. Scientometrics, 95(1):385–415, April 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0881-8.

Singh:2018:CRP

[Sin18] Varsha Singh. Comparing research productivity of returnee-PhDs in science, engineering, and the social sciences. Scientometrics, 115(3):1241–1252, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2706-x.

Savic:2014:SES

[SIR^{+}14] Milos Savić, Mirjana Ivanović, Milos Radovanović, Zoran Ognjanović, Aleksandar Pejović, and Tatjana Jaksić Krüger. The structure and evolution of scientific collaboration in Serbian mathematical journals. Scientometrics, 101(3):1805–1830, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1295-6.

Savic:2017:AII

[SIS17] Milos Savić, Mirjana Ivanović, and Bojana Dimić Surla. Analysis of intra-institutional research collaboration: a case of a Serbian faculty of sciences. Scientometrics, 110(1): 195–216, January 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-016-
2167-z.

Sivertsen:2016:DIS

[Siv16a] Gunnar Sivertsen. Data integration in Scandinavia. Scientometrics, 106(2):849–855, February 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1817-x.

Sivertsen:2016:PIC

[Siv16b] Gunnar Sivertsen. Patterns of internationalization and criteria for research assessment in the social sciences and humanities. Scientometrics, 107(2):357–368, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1845-1;   http://link.
springer.com/content/pdf/10.1007/s11192-016-1845-
1.pdf.

Schultz:2010:MIE

[SJ10] Laura I. Schultz and Frederick L. Joutz. Methods for identifying emerging general purpose technologies: a case study of nanotechnologies. Scientometrics, 85(1):155–170, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0244-2.

Singh:2019:SEI

[SJ19] Chakresh Kumar Singh and Shivakumar Jolad. Structure and evolution of Indian physics co-authorship networks. Scientometrics, 118(2):385–406, February 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-02999-5.

Smith:2018:CRP

[SJOC18] Thomas E. Smith, Kat S. Jacobs, Philip J. Osteen, and T. Edison Carter. Comparing the research productivity of social work doctoral programs using the h-index. Scientometrics, 116(3):1513–1530, September 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2832-5.

Soos:2011:TTR

[SK11] Sándor Soós and George Kampis. Towards a typology of research performance diversity: the case of top Hungarian players. Scientometrics, 87(2):357–371, May 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0351-8.

Soos:2012:BBS

[SK12] Sándor Soós and George Kampis. Beyond the basemap of science: mapping multiple structures in research portfolios: evidence from Hungary. Scientometrics, 93(3):869–891, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0713-x.

Song:2013:DKS

[SK13] Min Song and Su Yeon Kim. Detecting the knowledge structure of bioinformatics by mining full-text collections. Scientometrics, 96(1):183–201, July 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0900-9.

Sotudeh:2014:GDS

[SK14a] Hajar Sotudeh and Nahid Khoshian. Gender differences in science: the case of scientific productivity in nano science & technology during 2005–2007. Scientometrics, 98 (1):457–472, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1031-7.

Sotudeh:2014:GWP

[SK14b] Hajar Sotudeh and Nahid Khoshian. Gender, web presence and scientific productivity in nanoscience and nanotechnology. Scientometrics, 99(3):717–736, June 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1234-6.

Swar:2014:MIK

[SK14c] Bobby Swar and Gohar Feroz Khan. Mapping ICT knowledge infrastructure in South Asia. Scientometrics, 99(1): 117–137, April 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1099-0.

Saetnan:2016:EEK

[SK16] Eli Rudinow Saetnan and Richard Philip Kipling. Evaluating a European knowledge hub on climate change in agriculture: Are we building a better connected community? Scientometrics, 109(2):1057–1074, November 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-2064-5.pdf.

Shibayama:2017:IPD

[SK17] Sotaro Shibayama and Yoshie Kobayashi. Impact of Ph.D. training: a comprehensive analysis based on a Japanese national doctoral survey. Scientometrics, 113(1):387–415, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2479-7;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2479-
7.pdf.

Shashnov:2018:RLB

[SK18] Sergey Shashnov and Maxim Kotsemir. Research landscape of the BRICS countries: current trends in research output, thematic structures of publications, and the relative influence of partners. Scientometrics, 117(2):1115–1155, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2883-7.

Shin:2014:AND

[SKCK14] Dongwook Shin, Taehwan Kim, Joongmin Choi, and Jungsun Kim. Author name disambiguation using a graph model with node splitting and merging based on bibliographic information. Scientometrics, 100(1):15–50, July 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1289-4.

Sidiropoulos:2015:RII

[SKM15] Antonis Sidiropoulos, Dimitrios Katsaros, and Yannis Manolopoulos. Ranking and identifying influential scientists versus mass producers by the perfectionism index. Scientometrics, 103(1):1–31, April 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1515-0.

Skoric:2014:IBD

[Sko14] Marko M. Skoric. The implications of big data for developing and transitional economies: Extending the triple helix? Scientometrics, 99(1):175–186, April 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1106-5.

Skute:2019:OBB

[Sku19] Igors Skute. Opening the black box of academic entrepreneurship: a bibliometric analysis. Scientometrics, 120(1):237–265, July 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03116-w;   http://link.springer.com/content/
pdf/10.1007/s11192-019-03116-w.pdf.

Shapira:2017:TES

[SKY17] Philip Shapira, Seokbeom Kwon, and Jan Youtie. Tracking the emergence of synthetic biology. Scientometrics, 112(3):1439–1469, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
017-2452-5;   http://link.springer.com/content/pdf/
10.1007/s11192-017-2452-5.pdf.

Su:2010:MKS

[SL10] Hsin-Ning Su and Pei-Chun Lee. Mapping knowledge structure by keyword co-occurrence: a first look at journal papers in technology foresight. Scientometrics, 85(1):65–79, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0259-8.

Sivertsen:2012:CBC

[SL12a] Gunnar Sivertsen and Birger Larsen. Comprehensive bibliographic coverage of the social sciences and humanities in a citation index: an empirical analysis of the potential. Scientometrics, 91(2):567–575, May 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0615-3.

Smolinsky:2012:CRM

[SL12b] Lawrence Smolinsky and Aaron Lercher. Citation rates in mathematics: a study of variation by subdiscipline. Scientometrics, 91(3):911–924, June 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0647-3.

Shelton:2013:SCW

[SL13] R. D. Shelton and Grant Lewison. Scientific collaboration as a window and a door into North Korea. Scientometrics, 97(1):3–11, October 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0946-8.

Schymura:2014:IEC

[SL14] Michael Schymura and Andreas Löschel. Incidence and extent of co-authorship in environmental and resource economics: evidence from the Journal of Environmental Economics and Management. Scientometrics, 99(3):631–661, June 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1248-0.

Sun:2016:PEP

[SL16] Yutao Sun and Kai Liu. Proximity effect, preferential attachment and path dependence in inter-regional network: a case of China’s technology transaction. Scientometrics, 108 (1):201–220, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1951-0.

Schaefer:2017:OVD

[SL17] Kerstin J. Schaefer and Ingo Liefner. Offshore versus domestic: Can EM MNCs reach higher R&D quality abroad? Scientometrics, 113(3):1349–1370, December 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2533-5.

Silva:2017:PAT

[SLD^{+}17] Thiago H. P. Silva, Alberto H. F. Laender, Clodoveu A. Davis, Jr., Ana Paula Couto da Silva, and Mirella M. Moro. A profile analysis of the top Brazilian computer science graduate programs. Scientometrics, 113(1):237–255, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2462-3.

Sangam:2010:MGI

[SLG10] S. L. Sangam, Liang Liming, and Gireesh A. Ganjihal. Modeling the growth of Indian and Chinese liquid crystals literature as reflected in Science Citation Index (1997– 2006). Scientometrics, 84(1):49–52, July 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0079-x.

Sala:2017:CMS

[SLGO17] Francisco González Sala, Julia Osca Lluch, Francisco Tortosa Gil, and María Peñaranda Ortega. Characteristics of monographic special issues in Ibero–American psychology journals: visibility and relevance for authors and publishers. Scientometrics, 112(2):1069–1077, August 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2372-4.

Shu:2018:CTI

[SLH18] Fei Shu, Wen Lou, and Stefanie Haustein. Can Twitter increase the visibility of Chinese publications? Scientometrics, 116(1):505–519, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2732-8.

Serrano-Lopez:2017:WPR

[SLISC17] Antonio Eleazar Serrano-López, Peter Ingwersen, and Elias Sanz-Casado. Wind power research in Wikipedia: Does Wikipedia demonstrate direct influence of research publications and can it be used as adequate source in research evaluation? Scientometrics, 112(3):1471–1488, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2447-2.

Shin:2012:KBI

[SLK12] Jung Cheol Shin, Soo Jeung Lee, and Yangson Kim. Knowledge-based innovation and collaboration: a triplehelix approach in Saudi Arabia. Scientometrics, 90(1): 311–326, January 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0518-3.

Sun:2015:HWC

[SLXD15] Xiaoling Sun, Hongfei Lin, Kan Xu, and Kun Ding. How we collaborate: characterizing, modeling and predicting scientific collaborations. Scientometrics, 104(1):43–60, July 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1597-3.

Schulz:2012:RCR

[SM12] Peter A. Schulz and Edmilson J. T. Manganote. Revisiting country research profiles: learning about the scientific cultures. Scientometrics, 93(2):517–531, November 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0696-7.

Smyth:2014:AIR

[SM14] Russell Smyth and Vinod Mishra. Academic inbreeding and research productivity and impact in Australian law schools. Scientometrics, 98(1):583–618, January 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1052-2.

Soderlund:2015:CGS

[SM15] Therese Söderlund and Guy Madison. Characteristics of gender studies publications: a bibliometric analysis based on a Swedish population database. Scientometrics, 105(3):1347–1387, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11192-015-1702-7. See comments [LSL15] and reply [MS16a]..

Sahragard:2016:DSI

[SM16a] Rahman Sahragard and Hussein Meihami. A diachronic study on the information provided by the research titles of applied linguistics journals. Scientometrics, 108(3):1315– 1331, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-016-
2049-4.

Shehatta:2016:CAT

[SM16b] Ibrahim Shehatta and Khalid Mahmood. Correlation among top 100 universities in the major six global rankings: policy implications. Scientometrics, 109(2):1231–1254, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2065-4.

Soderlund:2017:ORE

[SM17] Therese Söderlund and Guy Madison. Objectivity and realms of explanation in academic journal articles concerning sex/gender: a comparison of gender studies and the other social sciences. Scientometrics, 112(2):1093–1109, August 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2407-x;   http://link.

springer.com/content/pdf/10.1007/s11192-017-2407-
x.pdf.

Small:2010:MSI

[Sma10] Henry Small. Maps of science as interdisciplinary discourse: co-citation contexts and the role of analogy. Scientometrics, 83(3):835–849, June 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0121-z.

Small:2011:IMS

[Sma11] Henry Small. Interpreting maps of science using citation context sentiments: a preliminary investigation. Scientometrics, 87(2):373–388, May 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
011-0349-2.

Salager-Meyer:2011:SGF

[SMAABJ11] Françoise Salager-Meyer, María Ángeles Alcaraz-Ariza, Marianela Luzardo Briceño, and Georges Jabbour. Scholarly gratitude in five geographical contexts: a diachronic and cross-generic approach of the acknowledgment paratext in medical discourse (1950–2010). Scientometrics, 86 (3):763–784, March 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0329-y.

Sassetti:2018:ECS

[SMCC18] Sara Sassetti, Giacomo Marzi, Vincenzo Cavaliere, and Cristiano Ciappei. Entrepreneurial cognition and socially situated approach: a systematic and bibliometric analysis. Scientometrics, 116(3):1675–1718, September 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2809-4.

Santos:2018:LBP

[SMF18] Gina Santos, Carla Susana Marques, and João J. Ferreira. A look back over the past 40 years of female entrepreneurship: mapping knowledge networks. Scientometrics, 115(2):953– 987, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2705-y.

Smith:2012:IFS

[Smi12] Derek R. Smith. Impact factors, scientometrics and the history of citation-based research. Scientometrics, 92(2): 419–427, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0685-x.

Sun:2016:VMO

[SML16] Jianjun Sun, Chao Min, and Jiang Li. A vector for measuring obsolescence of scientific articles. Scientometrics, 107 (2):745–757, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1884-7.

Saez-Martin:2017:API

[SMLHCP17] Alejandro Sáez-Martín, Antonio M. López-Hernandez, and Carmen Caba-Perez. Access to public information: a scientometric study of legal versus voluntary transparency in the public sector. Scientometrics, 113(3):1697–1720, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2541-5.

Sotudeh:2015:CBC

[SMM15] Hajar Sotudeh, Zahra Mazarei, and Mahdieh Mirzabeigi. CiteULike bookmarks are correlated to citations at journal and author levels in library and information science. Scientometrics, 105(3):2237–2248, December 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1745-9.

Sarkar:2019:ROS

[SMM^{+}19] Biplab Sarkar, Anusheel Munshi, Arjunan Manikandan, Tharmarnadar Ganesh, and Bidhu Kalyan Mohanti. Radiation oncology and social media platforms — use, benefits, pitfalls. Scientometrics, 118(2):699–703, February 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2976-3.

Song:2015:CSO

[SMY15] Yanhui Song, Feng Ma, and Siluo Yang. Comparative study on the obsolescence of humanities and social sciences in China: under the new situation of web. Scientometrics, 102 (1):365–388, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1410-8.

Sun:2010:MRA

[SN10] Yuan Sun and Masamitsu Negishi. Measuring the relationships among university, industry and other sectors in Japan’s national innovation system: a comparison of new approaches with mutual information indicators. Scientometrics, 82(3):677–685, March 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0179-7.

Sasvari:2019:EIS

[SND19] Péter Sasvári, András Nemeslaki, and László Duma. Exploring the influence of scientific journal ranking on publication performance in the Hungarian social sciences: the case of law and economics. Scientometrics, 119(2):595–616, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03081-4;   http://
link.springer.com/content/pdf/10.1007/s11192-019-
03081-4.pdf.

Snijder:2016:ROA

[Sni16] Ronald Snijder. Revisiting an open access monograph experiment: measuring citations and tweets 5 years later. Scientometrics, 109(3):1855–1875, December 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-2160-6.pdf.

Sobkowicz:2011:SOC

[Sob11] Paweł Sobkowicz. Simulations of opinion changes in scientific communities. Scientometrics, 87(2):233–250, May 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0339-4.

Sabaghinejad:2016:EPA

[SOBM16] Zivar Sabaghinejad, Farideh Osareh, Fatima Baji, and Parastou Parsaei Mohammadi. Estimating the partnership ability of Scientometrics journal authors based on WoS from 2001 to 2013 according to φ-index. Scientometrics, 109(1):73–84, October 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2001-7.

Soler:2006:RIS

[Sol06] José M. Soler. A rational indicator of scientific creativity. arxiv.org, August 1, 2006. URL http://adsabs.harvard.

edu/abs/2006physics...8006S;   http://arxiv.org/
abs/physics/0608006.

Sooryamoorthy:2010:MRS

[Soo10a] Radhamany Sooryamoorthy. Medical research in south Africa: a scientometric analysis of trends, patterns, productivity and partnership. Scientometrics, 84(3):863– 885, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0169-9.

Sooryamoorthy:2010:SSC

[Soo10b] Radhamany Sooryamoorthy. Science and scientific collaboration in South Africa: apartheid and after. Scientometrics, 84(2):373–390, August 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
009-0106-y.

Sooryamoorthy:2011:SPE

[Soo11a] Radhamany Sooryamoorthy. Scientific publications of engineers in South Africa, 1975–2005. Scientometrics, 86(1): 211–226, January 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0288-3.

Soos:2011:FAS

[Soó11b] Sándor Soós. The functional anatomy of science mapping. Scientometrics, 89(2):723–726, November 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0480-0.

Sooryamoorthy:2014:PPC

[Soo14a] Radhamany Sooryamoorthy. Publication productivity and collaboration of researchers in south Africa: new empirical evidence. Scientometrics, 98(1):531–545, January 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0990-z.

Soos:2014:ASB

[Soó14b] Sándor Soós. Age-sensitive bibliographic coupling reflecting the history of science: The case of the species problem. Scientometrics, 98(1):23–51, January 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1080-y.

Sooryamoorthy:2017:DTC

[Soo17] Radhamany Sooryamoorthy. Do types of collaboration change citation? A scientometric analysis of social science publications in South Africa. Scientometrics, 111(1):379–400, April 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-017-2265-6.

Sooryamoorthy:2018:PSA

[Soo18] Radhamany Sooryamoorthy. The production of science in Africa: an analysis of publications in the science disciplines, 2000–2015. Scientometrics, 115(1):317–349, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2675-0.

Sooryamoorthy:2019:SKS

[Soo19] Radhamany Sooryamoorthy. Scientific knowledge in South Africa: information trends, patterns and collaboration. Scientometrics, 119(3):1365–1386, June 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03096-x.

Sotudeh:2010:ISR

[Sot10] Hajar Sotudeh. Are Iranian scientists recognized as their productivity enhances? A comparison of Iran’s impact to global norms in different subfields of Science Citation Index during 2002–2005. Scientometrics, 83(1):39–54, April 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-009-0018-x.

Sotudeh:2012:HSS

[Sot12] Hajar Sotudeh. How sustainable a scientifically developing country could be in its specialties? The case of Iran’s publications in SCI in the 21st century compared to 1980s. Scientometrics, 91(1):231–243, April 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0542-3.

Sahu:2012:DAS

[SP12a] Satya Ranjan Sahu and Krushna Chandra Panda. A deductive approach to select or rank journals in multifaceted subject, oceanography. Scientometrics, 92(3):609–619, September 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0631-y.

Shapiro:2012:RDS

[SP12b] Matthew A. Shapiro and Han Woo Park. Regional development in South Korea: accounting for research area in centrality and networks. Scientometrics, 90(1):271–287, January 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0498-3.

Sahu:2014:DMA

[SP14] Satya Ranjan Sahu and Krushna Chandra Panda. Does the multi-authorship trend influence the quality of an article? Scientometrics, 98(3):2161–2168, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1127-0.

Sanchez:2018:THE

[SPB18] Gerardo Urrutia Sánchez, Lilian Prado, and Wolfgang Bietenholz. Theoretical high energy physics in Latin America from 1990 to 2012: a statistical study. Scientometrics, 116 (1):125–146, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2739-1.

Silva:2016:PIA

[SPdSM16] Thiago H. P. Silva, Gustavo Penha, Ana Paula Couto da Silva, and Mirella M. Moro. A performance indicator for academic communities based on external publication profiles. Scientometrics, 107(3):1389–1403, June 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1901-x.

Sepulveda:2014:PAS

[SPS14] Juan Sepúlveda, Adriana Paternina, and Andrés Suarez. Patent applications as source for measuring technological performance. Scientometrics, 98(2):1385–1395, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1050-4.

Sachithanantham:2015:SAR

[SR15] Shanmugam Sachithanantham and Selvaraju Raja. Scientometric analysis of rabies research literature in India: 1950– 2014. Scientometrics, 105(1):567–575, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1694-3.

Salimi:2016:MEU

[SR16] Negin Salimi and Jafar Rezaei. Measuring efficiency of university-industry Ph.D. projects using best worst method. Scientometrics, 109(3):1911–1938, December 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-2121-0.pdf.

Slade:2016:FRF

[SRF16] Catherine P. Slade, Saundra J. Ribando, and C. Kevin Fortner. Faculty research following merger: a job stress and social identity theory perspective. Scientometrics, 107(1):71– 89, April 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1881-x.

Sanchez-Riofrio:2015:BPR

[SRGMF15] Angélica María Sánchez-Riofrío, Luis Ángel Guerras-Martín, and Francisco Javier Forcadell. Business portfolio restructuring: a comprehensive bibliometric review. Scientometrics, 102(3):1921–1950, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1495-0.

Stefenon:2013:TYB

[SRP13] V. M. Stefenon, L. F. W. Roesch, and A. B. Pereira. Thirty years of Brazilian research in Antarctica: ups, downs and perspectives. Scientometrics, 95(1):325–331, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0809-3.

Shen:2015:EDR

[SRW^{+}15] Si Shen, Ronald Rousseau, Dongbo Wang, Danhao Zhu, Huoyu Liu, and Ruilun Liu. Editorial delay and its relation to subsequent citations: the journals Nature, Science and Cell. Scientometrics, 105(3):1867–1873, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1592-8.

Shen:2018:DPI

[SRW18] Si Shen, Ronald Rousseau, and Dongbo Wang. Do papers with an institutional e-mail address receive more citations than those with a non-institutional one? Scientometrics, 115(2):1039–1050, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2691-0.

Savanur:2010:MCC

[SS10a] Kiran Savanur and R. Srikanth. Modified collaborative coefficient: a new measure for quantifying the degree of research collaboration. Scientometrics, 84(2):365–371, August 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0100-4.

Schubert:2010:MSJ

[SS10b] András Schubert and Sándor Soós. Mapping of science journals based on h-similarity. Scientometrics, 83(2):589–600, May 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0167-y.

Schubert:2010:CCP

[SS10c] Torben Schubert and Radhamany Sooryamoorthy. Can the centre-periphery model explain patterns of international scientific collaboration among threshold and industrialised countries? The case of South Africa and Germany. Scientometrics, 83(1):181–203, April 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0074-2.

Selek:2014:UII

[SS14] Salih Selek and Ayman Saleh. Use of h index and g index for American academic psychiatry. Scientometrics, 99(2):541– 548, May 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1204-4.

Sriwannawit:2015:LSB

[SS15] Pranpreya Sriwannawit and Ulf Sandström. Large-scale bibliometric review of diffusion research. Scientometrics, 102(2):1615–1645, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
014-1448-7.

Simon:2016:TDM

[SS16] H. Simon and N. Sick. Technological distance measures: new perspectives on nearby and far away. Scientometrics, 107 (3):1299–1320, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1888-3.

Schubert:2018:WHG

[SS18] András Schubert and Gábor Schubert. Whatever happened to Garfield’s constant? Scientometrics, 114(2):659–667, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2527-3.

Salah:2016:SPI

[SSAG16] Almila Akdag Salah, Cassidy Sugimoto, Umut Al, and Wolfgang Glänzel. Selected papers of the 15th International Conference of the International Society for Scientometrics and Informetrics (ISSI), Bogaziçi University, Istanbul, Turkey, 29 June–4 July 2015. Scientometrics, 107(2):319– 320, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1934-1;   http://link.
springer.com/content/pdf/10.1007/s11192-016-1934-
1.pdf.

Siqueira:2017:RMV

[SSdOS17] Bruno V. L. Siqueira, Bruno E. Soares, Danilo R. de Oliveira, and Cássia M. Sakuragui. The regionalization of medicalized vernacular names of medicinal plants in Brazil. Scientometrics, 110(2):945–966, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2184-y.

Stoupas:2018:RRA

[SSG^{+}18] Georgios Stoupas, Antonis Sidiropoulos, Antonia Gogoglou, Dimitrios Katsaros, and Yannis Manolopoulos. Rainbow ranking: an adaptable, multidimensional ranking method for publication sets. Scientometrics, 116(1):147– 160, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2731-9.

Silva:2019:CAN

[SSN19] Fabio S. V. Silva, Peter A. Schulz, and Everard C. M. Noyons. Co-authorship networks and research impact in large research facilities: benchmarking internal reports and bibliometric databases. Scientometrics, 118(1):93–108, January 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2967-4.

Slyder:2011:CPL

[SSS^{+}11] Jacob B. Slyder, Beth R. Stein, Brent S. Sams, David M. Walker, B. Jacob Beale, Jeffrey J. Feldhaus, and Carolyn A. Copenheaver. Citation pattern and lifespan: a comparison of discipline, institution, and individual. Scientometrics, 89(3):955–966, December 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0467-x.

Sugimoto:2016:ASC

[SST^{+}16] Cassidy R. Sugimoto, Thomas J. Sugimoto, Andrew Tsou, Stasa Milojević, and Vincent Larivière. Age stratification and cohort effects in scholarly communication: a study of social sciences. Scientometrics, 109(2):997–1016, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2087-y.

Song:2018:PWS

[SSZL18] You Song, Fangling Situ, Hongjun Zhu, and Jinzhi Lei. To be the Prince to wake up Sleeping Beauty: the rediscovery of the delayed recognition studies. Scientometrics, 117 (1):9–24, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2830-7.

Schubert:2014:NJC

[ST14a] András Schubert and András Telcs. A note on the Jaccardized Czekanowski similarity index. Scientometrics, 98(2):1397–1399, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
013-1044-2.

Sud:2014:EA

[ST14b] Pardeep Sud and Mike Thelwall. Evaluating altmetrics. Scientometrics, 98(2):1131–1143, February 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1117-2.

Sud:2014:LTM

[ST14c] Pardeep Sud and Mike Thelwall. Linked title mentions: a new automated link search candidate. Scientometrics, 101(3):1831–1849, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1374-8.

Suarez-Tamayo:2018:ERR

[STCRPA18] Marcela Suárez-Tamayo, Francisco Collazo-Reyes, and Miguel Ángel Pérez-Angón. Emerging roles of regional journals in the accreditation of knowledge in tropical medicine: Biomédica and Memorias do Instituto Oswaldo Cruz, 2007– 2015. Scientometrics, 115(3):1615–1625, June 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2649-2.

Stern:2017:CBC

[Ste17] David I. Stern. Comment on Bornmann (2017): Confidence intervals for Journal Impact Factors. Scientometrics, 113(3):1811–1813, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2507-7. See [Bor17].

Steinbruchel:2019:CIP

[Ste19] Christoph Steinbrüchel. A citation index for principal investigators. Scientometrics, 118(1):305–320, January 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2971-8.

Sugimoto:2011:LAC

[Sug11] Cassidy R. Sugimoto. Looking across communicative genres: a call for inclusive indicators of interdisciplinarity. Scientometrics, 86(2):449–461, February 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0275-8.

Suominen:2014:PGG

[Suo14] Arho Suominen. Phases of growth in a green tech research network: a bibliometric evaluation of fuel cell technology from 1991 to 2010. Scientometrics, 100(1):51–72, July 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1285-8.

Singh:2015:CSR

[SUP15] Vivek Kumar Singh, Ashraf Uddin, and David Pinto. Computer science research: the top 100 institutions in India and in the world. Scientometrics, 104(2):529–553, August 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1612-8.

Sile:2019:MCP

[SV19] Linda Sile and Raf Vanderstraeten. Measuring changes in publication patterns in a context of performance-based research funding systems: the case of educational research in the University of Gothenburg (2005–2014). Scientometrics, 118(1):71–91, January 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2963-8.

Sorzano:2014:CSP

[SVCFI14] C. O. S. Sorzano, J. Vargas, G. Caffarena-Fernández, and A. Iriarte. Comparing scientific performance among equals. Scientometrics, 101(3):1731–1745, December 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1368-6.

Schneider:2019:ENC

[SvLVA19] Jesper W. Schneider, Thed van Leeuwen, Martijn Visser, and Kaare Aagaard. Examining national citation impact by comparing developments in a fixed and a dynamic journal set. Scientometrics, 119(2):973–985, May 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03082-3.

Soos:2018:LTT

[SVS18] Sándor Soós, Zsófia Vida, and András Schubert. Longterm trends in the multidisciplinarity of some typical natural and social sciences, and its implications on the SSH versus STM distinction. Scientometrics, 114(3):795–822, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2589-2.

Scarazzati:2019:ECS

[SW19a] Stefano Scarazzati and Lili Wang. The effect of collaborations on scientific research output: the case of nanoscience in Chinese regions. Scientometrics, 121(2):839–868, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03220-x;   http://
link.springer.com/content/pdf/10.1007/s11192-019-
03220-x.pdf.

Stockemer:2019:PEA

[SW19b] Daniel Stockemer and Michael J. Wigginton. Publishing in English or another language: An inclusive study of scholar’s language publication preferences in the natural, social and interdisciplinary sciences. Scientometrics, 118(2):645–652, February 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2987-0.

Sung:2014:CSP

[SWCH14] Hui-Yun Sung, Chun-Chieh Wang, Dar-Zen Chen, and Mu-Hsuan Huang. A comparative study of patent counts by the inventor country and the assignee country. Scientometrics, 100(2):577–593, August 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
013-1192-4.

Shiu:2014:DEK

[SWH14a] Jyh-Wen Shiu, Chan-Yuan Wong, and Mei-Chih Hu. The dynamic effect of knowledge capitals in the public research institute: insights from patenting analysis of ITRI (Taiwan) and ETRI (Korea). Scientometrics, 98(3):2051–2068, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-013-1158-6. See erratum

[SWH14b].

Shiu:2014:EDE

[SWH14b] Jyh-Wen Shiu, Chan-Yuan Wong, and Mei-Chih Hu. Erratum to: The dynamic effect of knowledge capitals in the public research institute: insights from patenting analysis of ITRI (Taiwan) and ETRI (Korea). Scientometrics, 100(2):605, August 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-014-1325-

4.pdf. See [SWH14a].

Sun:2016:SCE

[SX16] Yutao Sun and Belle Selene Xia. The scholarly communication of economic knowledge: a citation analysis of Google Scholar. Scientometrics, 109(3):1965–1978, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2140-x.

Schubert:2016:ICO

[SY16a] Torben Schubert and Guoliang Yang. Institutional change and the optimal size of universities. Scientometrics, 108(3):1129–1153, September 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2015-1.

Solarin:2016:GAI

[SY16b] Sakiru Adebola Solarin and Yuen Yee Yen. A global analysis of the impact of research output on economic growth. Scientometrics, 108(2):855–874, August 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-2002-6.

Shen:2019:LES

[SYDW19] Zhesi Shen, Liying Yang, Zengru Di, and Jinshan Wu. Large enough sample size to rank two groups of data reliably according to their means. Scientometrics, 118(2):653–671, February 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2995-0.

Sung:2017:VTP

[SYLC17] Hui-Yun Sung, Hsi-Yin Yeh, Jin-Kwan Lin, and Ssu-Han Chen. A visualization tool of patent topic evolution using a growing cell structure neural network. Scientometrics, 111 (3):1267–1285, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Shapira:2010:ESS

[SYP10] Philip Shapira, Jan Youtie, and Alan L. Porter. The emergence of social science research on nanotechnology. Scientometrics, 85(2):595–611, November 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0204-x.

Sun:2012:CGG

[SZ12] Fengjun Sun and Lijun Zhu. Citation genetic genealogy: a novel insight for citation analysis in scientific literature. Scientometrics, 91(2):577–589, May 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0588-2.

Stevenson:2015:TAI

[SZ15] Jennifer Ann Stevenson and Jin Zhang. A temporal analysis of institutional repository research. Scientometrics, 105(3): 1491–1525, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1728-x.

Sun:2018:MKD

[SZ18] Yaowu Sun and Yi Zhai. Mapping the knowledge domain and the theme evolution of appropriability research between 1986 and 2016: a scientometric review. Scientometrics, 116 (1):203–230, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2748-0.

Sweileh:2014:BAD

[SZAJS14] Waleed M. Sweileh, Sa’ed H. Zyoud, Samah W. Al-Jabi, and Ansam F. Sawalha. Bibliometric analysis of diabetes mellitus research output from Middle Eastern Arab countries during the period (1996–2012). Scientometrics, 101 (1):819–832, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1361-0.

Shakiba:2016:EAT

[SZAS16] Masoud Shakiba, Azam Zavvari, Nader Aleebrahim, and Mandeep Jit Singh. Evaluating the academic trend of RFID technology based on SCI and SSCI publications from 2001 to 2014. Scientometrics, 109(1):591–614, October 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2095-y.

Song:2016:RET

[SZD16] Jinbo Song, Honglian Zhang, and Wanli Dong. A review of emerging trends in global PPP research: analysis and visualization. Scientometrics, 107(3):1111–1147, June 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1918-1.

Saquib:2017:CDR

[SZMS17] Nazmus Saquib, Mohammed Saddik Zaghloul, AbdulRahman Mazrou, and Juliann Saquib. Cardiovascular disease research in Saudi Arabia: a bibliometric analysis. Scientometrics, 112(1):111–140, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Saquib:2018:QAC

[SZMS18] Juliann Saquib, Mohamed Saddik Zaghloul, AbdulRahman Mazrou, and Nazmus Saquib. A quality assessment of clinical research on type 2 diabetes in Saudi Arabia. Scientometrics, 116(3):2085–2096, September 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2823-6.

Shi:2018:DPD

[SZZC18] Shunshun Shi, Wenyu Zhang, Shuai Zhang, and Jie Chen. Does prestige dimension influence the interdisciplinary performance of scientific entities in knowledge flow? Evidence from the e-government field. Scientometrics, 117(2):1237– 1264, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2914-4.

Teodorescu:2011:GIC

[TA11] Daniel Teodorescu and Tudorel Andrei. The growth of international collaboration in East European scholarly communities: a bibliometric analysis of journal articles published between 1989 and 2009. Scientometrics, 89(2):711– 722, November 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0466-y.

Taskin:2014:SPA

[TA14a] Zehra Taskin and Umut Al. Standardization problem of author affiliations in citation indexes. Scientometrics, 98 (1):347–368, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1004-x.

Teodorescu:2014:ECC

[TA14b] Daniel Teodorescu and Tudorel Andrei. An examination of “citation circles” for social sciences journals in Eastern European countries. Scientometrics, 99(2):209–231, May 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-1210-6.

Taskin:2015:CIA

[TA15] Zehra Taskin and Arsev U. Aydinoglu. Collaborative interdisciplinary astrobiology research: a bibliometric study of the NASA Astrobiology Institute. Scientometrics, 103 (3):1003–1022, June 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1576-8.

Tong:2017:ETN

[TA17] Sichao Tong and Per Ahlgren. Evolution of three Nobel Prize themes and a Nobel snub theme in chemistry: a bibliometric study with focus on international collaboration. Scientometrics, 112(1):75–90, July 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-017-2377-z.pdf.

Taskin:2018:CBC

[TA18] Zehra Taskin and Umut Al. A content-based citation analysis study based on text categorization. Scientometrics, 114 (1):335–357, January 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2560-2.

Tahamtan:2016:FAN

[TAA16] Iman Tahamtan, Askar Safipour Afshar, and Khadijeh Ahamdzadeh. Factors affecting number of citations: a comprehensive review of the literature. Scientometrics, 107(3): 1195–1225, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1889-2.

Tahira:2013:SAE

[TAB13] Muzammil Tahira, Rose Alinda Alias, and Aryati Bakri. Scientometric assessment of engineering in Malaysian universities. Scientometrics, 96(3):865–879, September 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0961-4.

Tahira:2016:MLI

[TABA16] Muzammil Tahira, Rose Alinda Alias, Aryati Bakri, and A. Abrizah. Meso-level institutional and journal related indices for Malaysian engineering research. Scientometrics, 107(2):521–535, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1850-4.

Tahamtan:2019:WDC

[TB19a] Iman Tahamtan and Lutz Bornmann. What do citation counts measure? An updated review of studies on citations in scientific documents published between 2006 and 2018. Scientometrics, 121(3):1635–1684, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03243-4.

Tahmooresnejad:2019:CEV

[TB19b] Leila Tahmooresnejad and Catherine Beaudry. Capturing the economic value of triadic patents. Scientometrics, 118 (1):127–157, January 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2959-4.

Turko:2016:IPT

[TBB^{+}16] Tamara Turko, Gennady Bakhturin, Vitaly Bagan, Stanislav Poloskov, and Dmitry Gudym. Influence of the program “5top 100” on the publication activity of Russian universities. Scientometrics, 109(2):769–782, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2060-9.

Thor:2018:ISI

[TBMM18] Andreas Thor, Lutz Bornmann, Werner Marx, and Rüdiger Mutz. Identifying single influential publications in a research field: new analysis opportunities of the CRExplorer. Scientometrics, 116(1):591–608, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2733-7;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2733-
7.pdf.

Tahmooresnejad:2015:RPF

[TBS15] Leila Tahmooresnejad, Catherine Beaudry, and Andrea Schiffauerova. The role of public funding in nanotechnology scientific production: Where Canada stands in comparison to the United States. Scientometrics, 102(1):753–787, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1432-2.

Turki:2019:DAS

[TBT19] Houcemeddine Turki, Mohamed Ben Aouicha, and Mohamed Ali Hadj Taieb. Discussing Arab Spring’s effect on scientific productivity and research performance in Arab countries. Scientometrics, 120(1):337–339, July 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03127-7.

Tao:2012:ECC

[TBW^{+}12] Tianzhu Tao, Lulong Bo, Fei Wang, Jinbao Li, and Xiaoming Deng. Equal contributions and credit given to authors in anesthesiology journals during a 10-year period. Scientometrics, 91(3):1005–1010, June 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0558-8.

Trimble:2011:AAP

[TC11] Virgina Trimble and Jose A. Ceja. Are American astrophysics papers accepted more quickly than others? Scientometrics, 89(1):281–289, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0438-2.

Trimble:2013:AAP

[TC13] Virginia Trimble and Jose A. Ceja. Are American astrophysics papers accepted more quickly than others? Part II: correlations with citation rates, subdisciplines, and author numbers. Scientometrics, 95(1):45–54, April 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0793-7.

Tsai:2016:CEA

[TCB16] Chin-Chang Tsai, Elizabeth A. Corley, and Barry Bozeman. Collaboration experiences across scientific disciplines and cohorts. Scientometrics, 108(2):505–529, August 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1997-z.

Tang:2017:LSI

[TCC17] Muh-Chyun Tang, Yun Jen Cheng, and Kuang Hua Chen. A longitudinal study of intellectual cohesion in digital humanities using bibliometric analyses. Scientometrics, 113(2): 985–1008, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2496-6.

Tao:2015:TPC

[TCH^{+}15] Juan Tao, Rongxiao Che, Dekui He, Yunzhi Yan, Xiaoyun Sui, and Yifeng Chen. Trends and potential cautions in food web research from a bibliometric analysis. Scientometrics, 105(1):435–447, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1679-2.

Taramasco:2010:ATF

[TCR10] Carla Taramasco, Jean-Philippe Cointet, and Camille Roth. Academic team formation as evolving hypergraphs. Scientometrics, 85(3):721–740, December 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0226-4.

Tseng:2013:SAE

[TCT^{+}13] Yuen-Hsien Tseng, Chun-Yen Chang, M. Shane Tutwiler, Ming-Chao Lin, and James P. Barufaldi. A scientometric analysis of the effectiveness of Taiwan’s educational research projects. Scientometrics, 95(3):1141–1166, June 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0966-z.

Thijs:2017:IAP

[TDG17] Bart Thijs, Koenraad Debackere, and Wolfgang Glänzel. Improved author profiling through the use of citation classes. Scientometrics, 111(2):829–839, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Tunger:2018:BAC

[TE18] Dirk Tunger and Marc Eulerich. Bibliometric analysis of corporate governance research in German-speaking countries: applying bibliometrics to business research using a custom-made database. Scientometrics, 117(3):2041– 2059, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2919-z.

Teixeira:2011:MVC

[Tei11] Aurora A. C. Teixeira. Mapping the (in)visible college(s) in the field of entrepreneurship. Scientometrics, 89(1):1–36, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0445-3.

Terekhov:2017:BSR

[Ter17] Alexander I. Terekhov. Bibliometric spectroscopy of Russia’s nanotechnology: 2000–2014. Scientometrics, 110(3): 1217–1242, March 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-016-
2234-5.

Tan:2014:BAR

[TFH14] Jiang Tan, Hui-Zhen Fu, and Yuh-Shan Ho. A bibliometric analysis of research on proteomics in science citation index expanded. Scientometrics, 98(2):1473–1490, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1125-2.

Tatry:2014:EUL

[TFJD14] Marie-Violaine Tatry, Dominique Fournier, Benoît Jeannequin, and Françoise Dosba. EU27 and USA leadership in fruit and vegetable research: a bibliometric study from 2000 to 2009. Scientometrics, 98(3):2207–2222, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1160-z.

Tal:2016:LPC

[TG16] Diana Tal and Avishag Gordon. Leadership of the present, current theories of multiple involvements: a bibliometric analysis. Scientometrics, 107(1):259–269, April 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1880-y.

Tal:2017:PAL

[TG17] Diana Tal and Avishag Gordon. Publication attributes of leadership: what do they mean? Scientometrics, 112(3):1391–1402, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
017-2425-8.

Tal:2018:AIW

[TG18a] Diana Tal and Avishag Gordon. Antisemitism and Islamophobia: what does a bibliometric study reveal? Scientometrics, 117(3):1349–1359, December 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2953-x.

Thijs:2018:CLC

[TG18b] Bart Thijs and Wolfgang Glänzel. The contribution of the lexical component in hybrid clustering, the case of four decades of “Scientometrics”. Scientometrics, 115(1):21–33, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2659-0.

Tsai:2013:PRT

[TH13] Chih-Fong Tsai and Chihli Hung. Popular research topics in multimedia. Scientometrics, 95(1):465–479, April 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0932-1.

Tietze:2019:IMA

[TH19] Anna Tietze and Philip Hofmann. The h-index and multiauthor h_{m}-index for individual researchers in condensed matter physics. Scientometrics, 119(1):171–185, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03051-w.

Taba:2015:TUL

[THAL15] Seyedamir Tavakoli Taba, Liaquat Hossain, Simon Reay Atkinson, and Sarah Lewis. Towards understanding longitudinal collaboration networks: a case of mammography performance research. Scientometrics, 103(2):531–544, May 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1560-3.

Turki:2018:VLE

[THB18] Houcemeddine Turki, Mohamed Ali Hadj Taieb, and Mohamed Ben Aouicha. The value of letters to the editor. Scientometrics, 117(2):1285–1287, November 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2906-4.

Thelwall:2012:JIE

[The12] Mike Thelwall. Journal impact evaluation: a webometric perspective. Scientometrics, 92(2):429–441, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0669-x.

Thelwall:2016:ICB

[The16] Mike Thelwall. Interpreting correlations between citation counts and other indicators. Scientometrics, 108(1):337– 347, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1973-7.

Thelwall:2017:MRC

[The17a] Mike Thelwall. Are Mendeley reader counts useful impact indicators in all fields? Scientometrics, 113(3):1721– 1731, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2557-x.

Thelwall:2017:AOT

[The17b] Mike Thelwall. Avoiding obscure topics and generalising findings produces higher impact research. Scientometrics, 110(1):307–320, January 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2159-z.

Thelwall:2017:JBI

[The17c] Mike Thelwall. Judit Bar-Ilan: information scientist, computer scientist, scientometrician. Scientometrics, 113(3): 1235–1244, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2551-3.

Thelwall:2018:CMA

[The18a] Michael Thelwall. Can Microsoft Academic be used for citation analysis of preprint archives? the case of the Social Science Research Network. Scientometrics, 115(2):913–928, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2704-z.

Thelwall:2018:DGR

[The18b] Mike Thelwall. A decade of Garfield readers. Scientometrics, 114(2):669–674, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2601-x.

Thelwall:2018:DBJ

[The18c] Mike Thelwall. Differences between journals and years in the proportions of students, researchers and faculty registering Mendeley articles. Scientometrics, 115(2):717–729, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2689-7.

Thelwall:2018:DGC

[The18d] Mike Thelwall. Do gendered citation advantages influence field participation? Four unusual fields in the USA 1996– 2017. Scientometrics, 117(3):2133–2144, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2926-0.

Thelwall:2018:DMA

[The18e] Mike Thelwall. Does Microsoft Academic find early citations? Scientometrics, 114(1):325–334, January 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2558-9.

Thelwall:2018:EMR

[The18f] Mike Thelwall. Early Mendeley readers correlate with later citation counts. Scientometrics, 115(3):1231–1240, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2715-9.

Thelwall:2019:CRC

[The19a] Mike Thelwall. Are classic references cited first? An analysis of citation order within article sections. Scientometrics, 120 (2):723–731, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03142-8.

Thelwall:2019:IHC

[The19b] Mike Thelwall. The influence of highly cited papers on field normalised indicators. Scientometrics, 118(2):519–537, February 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-03001-y.

Tibana-Herrera:2018:GAL

[THFBdMA18] Gerardo Tibaná-Herrera, María Teresa Fernández-Bajón, and Félix de Moya-Anegón. Global analysis of the e-learning scientific domain: a declining category? Scientometrics, 114 (2):675–685, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2592-7.

Takeda:2010:TMC

[TK10] Yoshiyuki Takeda and Yuya Kajikawa. Tracking modularity in citation networks. Scientometrics, 83(3):783–792, June 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0158-z.

Thelwall:2016:CCT

[TK16] Mike Thelwall and Kayvan Kousha. Are citations from clinical trials evidence of higher impact research? An analysis of ClinicalTrials.gov. Scientometrics, 109(2):1341– 1351, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-016-
2112-1.

Thelwall:2017:RVG

[TK17] Mike Thelwall and Kayvan Kousha. ResearchGate versus Google Scholar: Which finds more early citations? Scientometrics, 112(2):1125–1131, August 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2400-4.

Thelwall:2017:MRI

[TKA17] Mike Thelwall, Kayvan Kousha, and Mahshid Abdoli. Is medical research informing professional practice more highly cited? Evidence from AHFS DI essentials in drugs.com. Scientometrics, 112(1):509–527, July 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Tien:2018:DAG

[TL18] Nguyen Minh Tien and Cyril Labbé. Detecting automatically generated sentences with grammatical structure similarity. Scientometrics, 116(2):1247–1271, August 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2789-4.

Tang:2014:RBS

[TLSH14] Erzi Tang, Fengchao Liu, Jia Sun, and Zulfqar Haider. The relationship between scientists and science: knowledgebased innovation output. Scientometrics, 98(3):1827–1835, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1138-x.

Teixeira:2012:BPE

[TM12] Aurora A. C. Teixeira and Luisa Mota. A bibliometric portrait of the evolution, scientific roots and influence of the literature on university-industry links. Scientometrics, 93(3): 719–743, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0823-5.

Thor:2016:NFC

[TMLB16] Andreas Thor, Werner Marx, Loet Leydesdorff, and Lutz Bornmann. New features of CitedReferencesExplorer (CR-Explorer). Scientometrics, 109(3):2049–2051, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2082-3.

Thelwall:2019:NEC

[TN19] Mike Thelwall and Tamara Nevill. No evidence of citation bias as a determinant of STEM gender disparities in US biochemistry, genetics and molecular biology research. Scientometrics, 121(3):1793–1801, December 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03271-0.

Teixeira:2018:EBI

[TO18] Eduardo Kunzel Teixeira and Mirian Oliveira. Editorial board interlocking in knowledge management and intellectual capital research field. Scientometrics, 117(3):1853– 1869, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2937-x.

Todeschini:2011:INB

[Tod11] Roberto Todeschini. The j-index: a new bibliometric index and multivariate comparisons between other common indices. Scientometrics, 87(3):621–639, June 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0346-5.

Todorovsky:2014:FSW

[Tod14] Dimitar Todorovsky. Follow-up study: on the working time budget of a university teacher. 45 years self-observation. Scientometrics, 101(3):2063–2070, December 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1284-9.

Tol:2011:CWC

[Tol11] Richard S. J. Tol. Credit where credit’s due: accounting for co-authorship in citation counts. Scientometrics, 89(1): 291–299, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-011-0451-
5.pdf.

Tol:2012:SVA

[Tol12] Richard S. J. Tol. Shapley values for assessing research production and impact of schools and scholars. Scientometrics, 90(3):763–780, March 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-011-0555-y.pdf.

Tomaszewski:2017:CCR

[Tom17] Robert Tomaszewski. Citations to chemical resources in scholarly articles: CRC Handbook of Chemistry and Physics and The Merck Index. Scientometrics, 112(3):1865– 1879, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2437-4.

Tomaszewski:2018:CSC

[Tom18] Robert Tomaszewski. A comparative study of citations to chemical encyclopedias in scholarly articles: Kirk–Othmer Encyclopedia of Chemical Technology and Ullmann’s Encyclopedia of Industrial Chemistry. Scientometrics, 117(1): 175–189, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2844-1.

Torrisi:2013:APC

[Tor13] Benedetto Torrisi. Academic productivity correlated with well-being at work. Scientometrics, 94(2):801–815, February 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0709-6.

Torrisi:2014:MAA

[Tor14] Benedetto Torrisi. A multidimensional approach to academic productivity. Scientometrics, 99(3):755–783, June 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1149-7.

Toivanen:2011:ARI

[TP11] Hannes Toivanen and Branco Ponomariov. African regional innovation systems: bibliometric analysis of research collaboration patterns 2005–2009. Scientometrics, 88(2):471–493, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0390-1.

Tseng:2014:ATT

[TR14] Ampere A. Tseng and Miroslav Raudensky. Assessments of technology transfer activities of US universities and associated impact of Bayh–Dole Act. Scientometrics, 101(3): 1851–1869, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1404-6.

Trimble:2010:GAT

[Tri10] Virginia Trimble. A generation of astronomical telescopes, their users, and publications. Scientometrics, 84(1):21–34, July 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0070-6.

Tang:2011:CUS

[TS11a] Li Tang and Philip Shapira. China–US scientific collaboration in nanotechnology: patterns and dynamics. Scientometrics, 88(1):1–16, July 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
011-0376-z.

Tang:2011:RDI

[TS11b] Li Tang and Philip Shapira. Regional development and interregional collaboration in the growth of nanotechnology research in China. Scientometrics, 86(2):299–315, February 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0274-9.

Tsai:2011:RTA

[Tsa11] Hsu-Hao Tsai. Research trends analysis by comparing data mining and customer relationship management through bibliometric methodology. Scientometrics, 87(3):425–450, June 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0353-6.

Tsay:2012:P

[Tsa12] Ming-Yueh Tsay. Preface. Scientometrics, 92(1):1–5, July 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0723-8.

Tsai:2015:RTF

[Tsa15] Hsu-Hao Tsai. The research trends forecasted by bibliometric methodology: a case study in e-commerce from 1996 to July 2015. Scientometrics, 105(2):1079–1089, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1719-y.

Sun:2013:MIT

[tScL13] Yu tao Sun and Feng chao Liu. Measuring international trade-related technology spillover: a composite approach of network analysis and information theory. Scientometrics, 94 (3):963–979, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0860-0.

Thijs:2013:DSO

[TSG13] Bart Thijs, Edgar Schiebel, and Wolfgang Glänzel. Do second-order similarities provide added-value in a hybrid approach? Scientometrics, 96(3):667–677, September 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0896-1.

Torres-Salinas:2011:MIF

[TSMTDLCH11] Daniel Torres-Salinas, Jose G. Moreno-Torres, Emilio Delgado-López-Cózar, and Francisco Herrera. A methodology for institution–field ranking based on a bidimensional analysis: the IFQ^{2}A index. Scientometrics, 88(3):

771–786, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0418-6.

Torres-Salinas:2014:ACC

[TSRGCCJC14] Daniel Torres-Salinas, Nicolás Robinson-García, Álvaro Cabezas-Clavijo, and Evaristo Jiménez-Contreras. Analyzing the citation characteristics of books: edited books, book series and publisher types in the book citation index. Scientometrics, 98(3):2113–2127, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1168-4.

Torres-Salinas:2017:FCG

[TSRGG17] Daniel Torres-Salinas, Nicolás Robinson-Garcia, and Juan Gorraiz. Filling the citation gap: measuring the multidimensional impact of the academic book at institutional level with PlumX. Scientometrics, 113(3):1371–1384, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2539-z.

Tseng:2013:JCL

[TT13] Yuen-Hsien Tseng and Ming-Yueh Tsay. Journal clustering of library and information science for subfield delineation using the bibliometric analysis toolkit: CATAR. Scientometrics, 95(2):503–528, May 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
013-0964-1.

Trappey:2017:IPE

[TTC17] Amy J. C. Trappey, Charles V. Trappey, and Curry L. S. Chung. IP portfolios and evolution of biomedical additive manufacturing applications. Scientometrics, 111(1): 139–157, April 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/accesspage/article/10.1007/s11192-017-
2273-6.

Tonta:2010:DUL

[TÜ10] Yasar Tonta and Yurdagül Ünal. Does Urquhart’s law hold for consortial use of electronic journals? Scientometrics, 83 (3):793–808, June 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0156-1.

Tu:2019:WCL

[Tu19] Jing Tu. What connections lead to good scientific performance? Scientometrics, 118(2):587–604, February 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-02997-7.

Tan:2015:IAD

[TUCR15] Hui Xuan Tan, Ephrance Abu Ujum, Kwai Fatt Choong, and Kuru Ratnavelu. Impact analysis of domestic and international research collaborations: a Malaysian case study. Scientometrics, 102(1):885–904, January 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1393-5.

Turki:2016:USP

[Tur16] Houcemeddine Turki. Usefulness of the single publication hindex. Scientometrics, 106(3):1281–1282, March 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-015-1808-y.

Tohmo:2017:DIL

[TV17] Timo Tohmo and Jutta Viinikainen. Does intersectoral labour mobility pay for academics? Scientometrics, 113 (1):83–103, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2477-9.

Teixeira:2017:SBT

[TVA17] Aurora A. C. Teixeira, Pedro Cosme Vieira, and Ana Patrícia Abreu. Sleeping Beauties and their princes in innovation studies. Scientometrics, 110(2):541–580, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2186-9.

Tang:2010:BFN

[TW10] Li Tang and John P. Walsh. Bibliometric fingerprints: name disambiguation based on approximate structure equivalence of cognitive maps. Scientometrics, 84(3):763–784, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0196-6.

Tijssen:2016:TFC

[TW16] Robert J. W. Tijssen and Jos Winnink. Twenty-first century macro-trends in the institutional fabric of science: bibliometric monitoring and analysis. Scientometrics, 109(3):2181–2194, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-2041-z.pdf.

Tijssen:2018:CRE

[TW18] Robert J. W. Tijssen and Jos J. Winnink. Capturing ‘R&D excellence’: indicators, international statistics, and innovative universities. Scientometrics, 114(2):687–699, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-017-2602-9;   https://
link.springer.com/content/pdf/10.1007/s11192-017-
2602-9.pdf.

Tian:2012:CTQ

[TYWZ12] Deqiao Tian, Yunzhou Yu, Yumin Wang, and Tao Zheng. Comparison of trends in the quantity and variety of science citation index (SCI) literature on human pathogens between China and the United States. Scientometrics, 93 (3):1019–1027, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-012-
0772-z.pdf.

Tijssen:2016:UIR

[TYYW16] Robert J. W. Tijssen, Alfredo Yegros-Yegros, and Jos J. Winnink. University-industry R&D linkage metrics: validity and applicability in world university rankings. Scientometrics, 109(2):677–696, November 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-2098-8.pdf.

Tian:2015:EID

[TZ15] Deqiao Tian and Tao Zheng. Emerging infectious disease: trends in the literature on SARS and H7N9 influenza. Scientometrics, 105(1):485–495, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1681-8;   http://link.
springer.com/content/pdf/10.1007/s11192-015-1681-
8.pdf.

Thijs:2015:BCH

[TZG15] Bart Thijs, Lin Zhang, and Wolfgang Glänzel. Bibliographic coupling and hierarchical clustering for the validation and improvement of subject-classification schemes. Scientometrics, 105(3):1453–1467, December 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1641-3.

Ain:2019:EIC

[uARA19] Qurat ul Ain, Hira Riaz, and Muhammad Tanvir Afzal. Evaluation of h-index and its citation intensity based variants in the field of mathematics. Scientometrics, 119(1): 187–211, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03009-y.

Uddin:2016:STF

[UBTS16] Ashraf Uddin, Jaideep Bhoosreddy, Marisha Tiwari, and Vivek Kumar Singh. A sciento-text framework to characterize research strength of institutions at fine-grained thematic area level. Scientometrics, 106(3):1135–1150, March 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-016-1836-2.

Uddin:2019:RFE

[UCH19] Shahadat Uddin, Nazim Choudhury, and Md Ekramul Hossain. A research framework to explore knowledge evolution and scholarly quantification of collaborative research. Scientometrics, 119(2):789–803, May 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03057-4.

Uddin:2012:TEA

[UHAR12] Shahadat Uddin, Liaquat Hossain, Alireza Abbasi, and Kim Rasmussen. Trend and efficiency analysis of co-authorship network. Scientometrics, 90(2):687–699, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0511-x.

Bajwa:2019:HMR

[uHBK19] Nida ul Habib Bajwa and Cornelius J. König. How much is research in the top journals of industrial/organizational psychology dominated by authors from the U.S.? Scientometrics, 120(3):1147–1161, September 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03180-2.

Bajwa:2019:WMG

[uHBLKH19] Nida ul Habib Bajwa, Markus Langer, Cornelius J. König, and Hannah Honecker. What might get published in management and applied psychology? Experimentally manipulating implicit expectations of reviewers regarding hedges. Scientometrics, 120(3):1351–1371, September 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03164-2.

Ucar:2014:GNR

[ULFRU^{+}14] Iñaki Ucar, Felipe López-Fernandino, Pablo Rodriguez-Ulibarri, Laura Sesma-Sanchez, Veronica Urrea-Micó, and Joaquín Sevilla. Growth in the number of references in engineering journal papers during the 1972–2013 period. Scientometrics, 98(3):1855–1864, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1113-6.

Uriona-Maldonado:2012:SAS

[UMdSV12] Mauricio Uriona-Maldonado, Raimundo N. M. dos Santos, and Gregorio Varvakis. State of the art on the systems of innovation research: a bibliometrics study up to 2009. Scientometrics, 91(3):977–996, June 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0653-5.

Ukrainski:2014:CPS

[UMK14] Kadri Ukrainski, Jaan Masso, and Hanna Kanep. Cooperation patterns in science within Europe: the standpoint of small countries. Scientometrics, 99(3):845–863, June 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-1224-0.

Upham:2010:IKC

[URU10a] S. Phineas Upham, Lori Rosenkopf, and Lyle H. Ungar. Innovating knowledge communities. Scientometrics, 83(2): 525–554, May 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0102-2.

Upham:2010:PKS

[URU10b] S. Phineas Upham, Lori Rosenkopf, and Lyle H. Ungar. Positioning knowledge: schools of thought and new knowledge creation. Scientometrics, 83(2):555–581, May 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0097-8.

Upham:2010:ERF

[US10] S. Phineas Upham and Henry Small. Emerging research fronts in science and technology: patterns of new knowledge development. Scientometrics, 83(1):15–38, April 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0051-9.

Uddin:2015:SMC

[USPO15] Ashraf Uddin, Vivek Kumar Singh, David Pinto, and Ivan Olmos. Scientometric mapping of computer science research in Mexico. Scientometrics, 105(1):97–114, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1654-y.

Vega-Almeida:2018:DVY

[VACCAJ18] Rosa Lidia Vega-Almeida, Humberto Carrillo-Calvet, and Ricardo Arencibia-Jorge. Diseases and vector: a 10 years view of scientific literature on Aedes aegypti. Scientometrics, 115(3):1627–1634, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2650-9.

Villasenor:2017:MCS

[VAJCC17] Elio Atenógenes Villaseñor, Ricardo Arencibia-Jorge, and Humberto Carrillo-Calvet. Multiparametric characterization of scientometric performance profiles assisted by neural networks: a study of Mexican higher education institutions. Scientometrics, 110(1):77–104, January 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2166-0.

Vannucci:2010:DDC

[Van10] Stefano Vannucci. Dominance dimension: a common parametric formulation for integer-valued scientific impact indices. Scientometrics, 84(1):43–48, July 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0091-1.

Vanclay:2012:IFO

[Van12] Jerome K. Vanclay. Impact factor: outdated artefact or stepping-stone to journal certification? Scientometrics, 92 (2):211–238, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0561-0.

Vanecek:2014:EPB

[Van14] Jiri Vanecek. The effect of performance-based research funding on output of R&D results in the Czech Republic. Scientometrics, 98(1):657–681, January 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1061-1.

Varga:2011:MSI

[Var11] Attila V. Varga. Measuring the semantic integrity of scientific fields: a method and a study of sociology, economics and biophysics. Scientometrics, 88(1):163–177, July 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0342-9.

Varshney:2012:GED

[Var12] Lav R. Varshney. The Google effect in doctoral theses. Scientometrics, 92(3):785–793, September 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0654-4.

Vega-Arce:2019:RPT

[VASNU^{+}19] Maribel Vega-Arce, Gonzalo Salas, Gastón Núñez-Ulloa, Cristián Pinto-Cortez, Ivelisse Torres Fernandez, and Yuh-Shan Ho. Research performance and trends in child sexual abuse research: a Science Citation Index Expandedbased analysis. Scientometrics, 121(3):1505–1525, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03267-w.

vanArensbergen:2012:GDS

[vAvdWvdB12] Pleun van Arensbergen, Inge van der Weijden, and Peter van den Besselaar. Gender differences in scientific productivity: a persisting phenomenon? Scientometrics, 93(3):857–868, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-012-0712-y.pdf.

Vanclay:2012:MER

[VB12] Jerome K. Vanclay and Lutz Bornmann. Metrics to evaluate research performance in academic institutions: a critique of ERA 2010 as applied in forestry and the indirect H_{2} index as a possible alternative. Scientometrics, 91(3):751– 771, June 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0618-8.

vanBochove:2013:ESS

[vB13] Cornelis A. van Bochove. Economic statistics and scientometrics. Scientometrics, 96(3):799–818, September 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-0960-5.

Velden:2017:CTE

[VBG^{+}17] Theresa Velden, Kevin W. Boyack, Jochen Gläser, Rob Koopman, Andrea Scharnhorst, and Shenghui Wang. Comparison of topic extraction approaches and their results. Scientometrics, 111(2):1169–1221, May 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Vera-Baceta:2019:WSS

[VBTK19] Miguel-Angel Vera-Baceta, Michael Thelwall, and Kayvan Kousha. Web of Science and Scopus language coverage. Scientometrics, 121(3):1803–1813, December 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03264-z.

Vela:2012:PWS

[VCC12] Belén Vela, Paloma Cáceres, and José María Cavero. Participation of women in software engineering publications. Scientometrics, 93(3):661–679, December 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0774-x.

vandenBosch:2016:ESE

[vdBBdK16] Antal van den Bosch, Toine Bogers, and Maurice de Kunder. Estimating search engine index size variability: a 9year longitudinal study. Scientometrics, 107(2):839–856, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1863-z;   http://link.
springer.com/content/pdf/10.1007/s11192-016-1863-
z.pdf.

vandenBesselaar:2016:GDR

[vdBS16] Peter van den Besselaar and Ulf Sandström. Gender differences in research performance and its impact on careers: a longitudinal case study. Scientometrics, 106(1):143–162, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1775-3;   http://link.
springer.com/content/pdf/10.1007/s11192-015-1775-
3.pdf.

vandenBesselaar:2018:SGD

[vdBSS18] Peter van den Besselaar, Ulf Sandström, and Hélène Schiffbaenker. Studying grant decision-making: a linguistic analysis of review reports. Scientometrics, 117(1):313–329, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2848-x;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2848-
x.pdf.

vanderPol:2018:CEK

[vdPR18] Johannes van der Pol and Jean-Paul Rameshkoumar. The co-evolution of knowledge and collaboration networks: the role of the technology life-cycle. Scientometrics, 114(1): 307–323, January 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2579-4.

Vancauwenbergh:2016:RIC

[VDV16] Sadia Vancauwenbergh, Pieter De Leenheer, and Geert Van Grootel. On research information and classification governance in an inter-organizational context: the Flanders Research Information Space. Scientometrics, 108(1):425–439, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1912-7.

Verleysen:2014:IPR

[VE14] Frederik T. Verleysen and Tim C. E. Engels. Internationalization of peer reviewed and non-peer reviewed book publications in the social sciences and humanities.

Scientometrics, 101(2):1431–1444, November 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1267-x.

Valderrama:2018:PSM

[VEJC^{+}18a] Pilar Valderrama, Manuel Escabias, Evaristo Jiménez-Contreras, Alberto Rodríguez-Archilla, and Mariano J. Valderrama. Proposal of a stochastic model to determine the bibliometric variables influencing the quality of a journal: application to the field of dentistry. Scientometrics, 115 (2):1087–1095, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2707-9.

Valderrama:2018:MLC

[VEJC^{+}18b] Pilar Valderrama, Manuel Escabias, Evaristo Jiménez-Contreras, Mariano J. Valderrama, and Pilar Baca. A mixed longitudinal and cross-sectional model to forecast the journal impact factor in the field of dentistry. Scientometrics, 116(2):1203–1212, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2801-z.

Velema:2012:CNB

[Vel12] Thijs A. Velema. The contingent nature of brain gain and brain circulation: their foreign context and the impact of return scientists on the scientific community in their country of origin. Scientometrics, 93(3):893–913, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0751-4.

Veugelers:2010:TMS

[Veu10] Reinhilde Veugelers. Towards a multipolar science world: trends and impact. Scientometrics, 82(2):439–456, February 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0045-7.

vanEck:2010:SSV

[vEW10] Nees Jan van Eck and Ludo Waltman. Software survey: VOSviewer, a computer program for bibliometric mapping. Scientometrics, 84(2):523–538, August 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-009-0146-3.pdf.

vanEck:2017:CBC

[vEW17] Nees Jan van Eck and Ludo Waltman. Citation-based clustering of publications using CitNetExplorer and VOSviewer. Scientometrics, 111(2):1053–1070, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-017-2300-7.pdf.

vanEck:2010:ATI

[vEWNB10] Nees Jan van Eck, Ludo Waltman, Ed C. M. Noyons, and Reindert K. Buter. Automatic term identification for bibliometric mapping. Scientometrics, 82(3):581–596, March 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-010-0173-0.pdf.

Vanecek:2010:BEF

[VFA10] Jiri Vanecek, Martin Fatun, and Vladimir Albrecht. Bibliometric evaluation of the FP-5 and FP-6 results in the Czech Republic. Scientometrics, 83(1):103–114, April 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0028-8.

Vieira:2011:IIR

[VG11a] Elizabeth S. Vieira and José A. N. F. Gomes. An impact indicator for researchers. Scientometrics, 89(2):607– 629, November 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0464-0.

Vieira:2011:JRI

[VG11b] Elizabeth S. Vieira and José A. N. F. Gomes. The journal relative impact: an indicator for journal assessment. Scientometrics, 89(2):631–651, November 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0469-8.

Vicente-Gomila:2014:CSS

[VG14] Jose M. Vicente-Gomila. The contribution of syntacticsemantic approach to the search for complementary literatures for scientific or technical discovery. Scientometrics, 100(3):659–673, September 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
014-1299-2.

Vilkins:2017:TEC

[VG17] Samantha Vilkins and Will J. Grant. Types of evidence cited in Australian Government publications. Scientometrics, 113(3):1681–1695, December 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2544-2.

Vicente-Gomila:2017:DSC

[VGPdlC^{+}17] Jose M. Vicente-Gomila, Anna Palli, Begoña de la Calle, Miguel A. Artacho, and Sara Jimenez. Discovering shifts in competitive strategies in probiotics, accelerated with Tech-Mining. Scientometrics, 111(3):1907–1923, June 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Vainio:2017:HTS

[VH17] Julia Vainio and Kim Holmberg. Highly tweeted science articles: who tweets them? An analysis of Twitter user profile descriptions. Scientometrics, 112(1):345–366, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

derStocken:2016:ACB

[VHD^{+}16] Tom Van der Stocken, Jean Hugé, Evelien Deboelpaep, Maarten P. M. Vanhove, Luc Janssens de Bisthoven, and Nico Koedam. Academic capacity building: holding up a mirror. Scientometrics, 106(3):1277–1280, March 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1811-3.

Verbree:2015:OFI

[VHG^{+}15] Maaike Verbree, Edwin Horlings, Peter Groenewegen, Inge Van der Weijden, and Peter van den Besselaar. Organizational factors influencing scholarly performance: a multivariate study of biomedical research groups. Scientometrics, 102(1):25–49, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1437-x.

Vana:2016:CJM

[VHH16] Laura Vana, Ronald Hochreiter, and Kurt Hornik. Computing a journal meta-ranking using paired comparisons and adaptive lasso estimators. Scientometrics, 106(1):229–251, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1772-6.

Vilibic:2010:HMS

[Vil10] Ivica Vilibić. How much the shared ocean or lake basins connect the researchers in neighbouring countries? Scientometrics, 83(2):463–470, May 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
009-0010-5.

Vinkler:2010:IES

[Vin10a] P. Vinkler. Indicators are the essence of scientometrics and bibliometrics. Scientometrics, 85(3):861–866, December 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0159-y.

Vinkler:2010:INI

[Vin10b] Peter Vinkler. The π_{v}-index: a new indicator to characterize the impact of journals. Scientometrics, 82(3):461–475, March 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0182-z.

Vinkler:2012:GIF

[Vin12a] Péter Vinkler. The Garfield Impact Factor, one of the fundamental indicators in scientometrics. Scientometrics, 92 (2):471–483, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0688-7.

Vinluan:2012:RPE

[Vin12b] Lorelei R. Vinluan. Research productivity in education and psychology in the Philippines and comparison with ASEAN countries. Scientometrics, 91(1):277–294, April 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0496-5.

Vinkler:2017:SIE

[Vin17] Péter Vinkler. The size and impact of the elite set of publications in scientometric assessments. Scientometrics, 110(1):163–177, January 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2165-1.

Vinkler:2018:SSR

[Vin18] Peter Vinkler. Structure of the scientific research and science policy. Scientometrics, 114(2):737–756, February 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2568-7.

Vinkler:2019:CJE

[Vin19] Péter Vinkler. Core journals and elite subsets in scientometrics. Scientometrics, 121(1):241–259, October 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03199-5;   http://link.springer.
com/content/pdf/10.1007/s11192-019-03199-5.pdf.

vanLeeuwen:2012:DSB

[vL12] Thed van Leeuwen. Discussing some basic critique on Journal Impact Factors: revision of earlier comments. Scientometrics, 92(2):443–455, August 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-012-0677-x.pdf.

vanLeeuwen:2013:REM

[vLCCMV13] Thed van Leeuwen, Rodrigo Costas, Clara Calero-Medina, and Martijn Visser. The role of editorial material in bibliometric research performance assessments. Scientometrics, 95(2):817–828, May 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0904-5.

Vugteveen:2014:DIR

[VLV14] Pim Vugteveen, Rob Lenders, and Peter Van den Besselaar. The dynamics of interdisciplinary research fields: the case of river research. Scientometrics, 100(1):73–96, July 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1286-7.

vanLeeuwen:2016:BAO

[vLvWW16] Thed N. van Leeuwen, Erik van Wijk, and Paul F. Wouters. Bibliometric analysis of output and impact based on CRIS data: a case study on the registered output of a Dutch university. Scientometrics, 106(1):1–16, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1788-y;   http://link.
springer.com/content/pdf/10.1007/s11192-015-1788-
y.pdf.

Vakilian:2015:BAL

[VMM15] Mohammadmahdi Vakilian, Burhanuddin Yeop Majlis, and Maryam Mousavi. A bibliometric analysis of lab-on-a-chip research from 2001 to 2013. Scientometrics, 105(2):789– 804, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1732-1.

Valentin:2016:OOI

[VNA16] Finn Valentin, Maria Theresa Norn, and Lars Alkaersig. Orientations and outcome of interdisciplinary research: the case of research behaviour in translational medical science. Scientometrics, 106(1):67–90, January 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1784-2.

Verleysen:2017:PMA

[VO17] Frederik T. Verleysen and Truyken L. B. Ossenblok. Profiles of monograph authors in the social sciences and humanities: an analysis of productivity, career stage, co-authorship, disciplinary affiliation and gender, based on a regional bibliographic database. Scientometrics, 111(3):1673–1686, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

vonProff:2013:ICD

[vPD13] Sidonia von Proff and Anja Dettmann. Inventor collaboration over distance: a comparison of academic and corporate patents. Scientometrics, 94(3):1217–1238, March 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0812-8.

Viiu:2016:RDC

[VPM16] Gabriel-Alexandru Vîiu, Mihai Paunescu, and Adrian Miroiu. Research-driven classification and ranking in higher education: an empirical appraisal of a Romanian policy experience. Scientometrics, 107(2):785–805, May 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1860-2.

vanRaan:2012:PJI

[vR12] Anthony F. J. van Raan. Properties of journal impact in relation to bibliometric research group performance indicators. Scientometrics, 92(2):457–469, August 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-012-0747-0.pdf.

vanRaan:2017:SBC

[vR17] Anthony F. J. van Raan. Sleeping beauties cited in patents: Is there also a dormitory of inventions? Scientometrics, 110(3):1123–1156, March 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-2215-8.pdf.

Vaughan:2012:EWK

[VRF12] Liwen Vaughan and Esteban Romero-Frías. Exploring Web keyword analysis as an alternative to link analysis: a multiindustry case. Scientometrics, 93(1):217–232, October 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0640-x.

vanRaan:2011:SLE

[vRvLV11] Anthony F. J. van Raan, Thed N. van Leeuwen, and Martijn S. Visser. Severe language effect in university rankings: particularly Germany and France are wronged in citation-based rankings. Scientometrics, 88(2):495–498, August 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-011-0382-1.pdf.

vanRaan:2018:DYS

[vRW18] Anthony F. J. van Raan and Jos J. Winnink. Do younger Sleeping Beauties prefer a technological prince? Scientometrics, 114(2):701–717, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2603-8;   https://link.springer.
com/content/pdf/10.1007/s11192-017-2603-8.pdf.

Vogl:2018:PBA

[VSK18] Sebastian Vogl, Thomas Scherndl, and Anton Kühberger. #Psychology: a bibliometric analysis of psychological literature in the online media. Scientometrics, 115(3):1253–1269, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2727-5;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2727-
5.pdf.

Vanderelst:2012:PIP

[VSS12] Dieter Vanderelst, Sara Speybroeck, and Niko Speybroeck. The perceived impact of publications on neglected tropical zoonoses as measured by their impact factor. Scientometrics, 90(2):331–342, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0488-5.

Vanga:2015:GFA

[VSVR15] Sai Kranthi Vanga, Ashutosh Singh, Brinda Harish Vagadia, and Vijaya Raghavan. Global food allergy research trend: a bibliometric analysis. Scientometrics, 105(1):203–213, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1660-0.

Vieira:2010:FMM

[VT10] Pedro Cosme Vieira and Aurora A. C. Teixeira. Are finance, management, and marketing autonomous fields of scientific research? An analysis based on journal citations. Scientometrics, 85(3):627–646, December 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0292-7.

Vaughan:2017:IDD

[VTY17] Liwen Vaughan, Juan Tang, and Rongbin Yang. Investigating disciplinary differences in the relationships between citations and downloads. Scientometrics, 111(3):1533–1545, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Velden:2010:NAA

[VuHL10] Theresa Velden, Asif ul Haque, and Carl Lagoze. A new approach to analyzing patterns of collaboration in coauthorship networks: mesoscopic analysis and interpretation. Scientometrics, 85(1):219–242, October 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0224-6.

Vahanikkila:2016:HDS

[VVN16] Hannu Vähänikkilä, Jorma I. Virtanen, and Pentti Nieminen. How do statistics in dental articles differ from those articles published in highly visible medical journals? Scientometrics, 108(3):1417–1424, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2028-9.

vanWilgen:2016:BAI

[vWBS^{+}16] Brian W. van Wilgen, Nelius Boshoff, Izak P. J. Smit, Sofia Solano-Fernandez, and Luanita van der Walt. A bibliometric analysis to illustrate the role of an embedded research capability in South African National Parks. Scientometrics, 107(1):185–212, April 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1879-4.

vanWesel:2014:WDC

[vWWtH14] Maarten van Wesel, Sally Wyatt, and Jeroen ten Haaf. What a difference a colon makes: how superficial factors influence subsequent citation. Scientometrics, 98(3): 1601–1615, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1154-x.

Velden:2017:MCS

[VYL17] Theresa Velden, Shiyan Yan, and Carl Lagoze. Mapping the cognitive structure of astrophysics by infomap clustering of the citation network and topic affinity analysis. Scientometrics, 111(2):1033–1051, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

vanZyl:2013:GPD

[vZ13] J. Martin van Zyl. The generalized Pareto distribution fitted to research outputs of countries. Scientometrics, 94(3): 1099–1109, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0798-2.

Valderrama-Zurian:2019:BNF

[VZAMG19] Juan-Carlos Valderrama-Zurián, Remedios Aguilar-Moya, and Juan Gorraiz. On the bibliometric nature of a foreseeable relationship: open access and education. Scientometrics, 120(3):1031–1057, September 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03175-z.

Woltmann:2018:TUI

[WA18a] Sabrina L. Woltmann and Lars Alkærsig. Tracing university-industry knowledge transfer through a text mining approach. Scientometrics, 117(1):449–472, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2849-9.

Wray:2018:RS

[WA18b] K. Brad Wray and Line Edslev Andersen. Retractions in science. Scientometrics, 117(3):2009–2019, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2922-4.

Waaijer:2013:CSP

[Waa13] Cathelijn J. F. Waaijer. Careers in science: policy issues according to Nature and Science editorials. Scientometrics, 96 (2):485–495, August 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-0958-z.

Wada:2016:OPA

[Wad16] Tetsuo Wada. Obstacles to prior art searching by the trilateral patent offices: empirical evidence from International Search Reports. Scientometrics, 107(2):701–722, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1858-9;   http://link.
springer.com/content/pdf/10.1007/s11192-016-1858-

9.pdf. See erratum [Wad17].

Wada:2017:EOP

[Wad17] Tetsuo Wada. Erratum to: Obstacles to prior art searching by the trilateral patent offices: empirical evidence from International Search Reports. Scientometrics, 113(3): 1821, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1966-6;
http://link.springer.com/content/pdf/10.1007/s11192-
016-1966-6.pdf. See [Wad16].

Wada:2018:CEP

[Wad18] Tetsuo Wada. The choice of examiner patent citations for refusals: evidence from the trilateral offices. Scientometrics, 117(2):825–843, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2885-5.

Walters:2015:MQQ

[Wal15] Glenn D. Walters. Measuring the quantity and quality of scholarly productivity in criminology and criminal justice: a test of three integrated models. Scientometrics, 102(3): 2011–2022, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1496-z.

Walters:2016:AAO

[Wal16] Glenn D. Walters. Adding authorship order to the quantity and quality dimensions of scholarly productivity: evidence from group- and individual-level analyses. Scientometrics, 106(2):769–785, February 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1803-3.

Wang:2013:CTW

[Wan13] Jian Wang. Citation time window choice for research impact evaluation. Scientometrics, 94(3):851–872, March 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0775-9.

Wang:2016:SCAa

[Wan16] Lili Wang. The structure and comparative advantages of China’s scientific research: quantitative and qualitative perspectives. Scientometrics, 106(1):435–452, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1650-2.

Winarko:2016:AQT

[WAT16] Bambang Winarko, A. Abrizah, and Muzammil Tahira. An assessment of quality, trustworthiness and usability of Indonesian agricultural science journals: stated preference versus revealed preference study. Scientometrics, 108(1): 289–304, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1970-x.

Wray:2015:PSV

[WB15] K. Brad Wray and Lutz Bornmann. Philosophy of science viewed through the lense of “referenced publication years spectroscopy” (RPYS). Scientometrics, 102(3):1987–1996, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1465-6.

Wang:2012:BTM

[WBH^{+}12] Jian Wang, Kaspars Berzins, Diana Hicks, Julia Melkers, Fang Xiao, and Diogo Pinheiro. A boosted-trees method for name disambiguation. Scientometrics, 93(2): 391–411, November 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0681-1.

Wang:2017:TCC

[WBX^{+}17] Wei Wang, Xiaomei Bai, Feng Xia, Teshome Megersa Bekele, Xiaoyan Su, and Amr Tolba. From triadic closure to conference closure: the role of academic conferences in promoting scientific collaborations. Scientometrics, 113(1): 177–193, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2468-x.

Wang:2018:QER

[WBX18] Binglu Wang, Yi Bu, and Yang Xu. A quantitative exploration on reasons for citing articles from the perspective of cited authors. Scientometrics, 116(2):675–687, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2787-6.

Wang:2018:TNB

[WC18] Mengyang Wang and Lihe Chai. Three new bibliometric indicators/approaches derived from keyword analysis. Scientometrics, 116(2):721–750, August 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2768-9.

Wang:2015:BAR

[WCB^{+}15] Lixian Wang, Xi Chen, Anming Bao, Xiaoyun Zhang, Miao Wu, Yun Hao, and Jingjing He. A bibliometric analysis of research on Central Asia during 1990–2014. Scientometrics, 105(2):1223–1237, November 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1727-y.

Wang:2012:SPA

[WCK^{+}12] Ju-O Wang, Tzeng-Ji Chen, Senyeong Kao, Te-Chun Yeh, Li-Fang Chou, and Shung-Tai Ho. Scientific publications by anesthesia departments in East Asia. Scientometrics, 92 (1):135–143, July 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0717-6.

Wang:2014:AER

[WCL14] Xiaoguang Wang, Qikai Cheng, and Wei Lu. Analyzing evolution of research topics with NEViewer: a new method based on dynamic co-word networks. Scientometrics, 101(2):1253–1271, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1347-y.

Wu:2013:AND

[WD13] Jiang Wu and Xiu-Hao Ding. Author name disambiguation in scientific collaboration and mobility cases. Scientometrics, 96(3):683–697, September 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0978-8.

Wien:2017:CIR

[WDL17] Charlotte Wien, Bertil F. Dorch, and Asger Væring Larsen. Contradicting incentives for research collaboration. Scientometrics, 112(2):903–915, August 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2412-0.

Wittek:2017:RSL

[WDN17] Peter Wittek, Sándor Darányi, and Gustaf Nelhans. Ruling out static latent homophily in citation networks. Scientometrics, 110(2):765–777, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-2194-9.pdf.

Wolszczak-Derlacz:2011:EEP

[WDP11] Joanna Wolszczak-Derlacz and Aleksandra Parteka. Efficiency of European public higher education institutions: a two-stage multicountry approach. Scientometrics, 89(3): 887–917, December 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-011-0484-
9.pdf.

Wang:2016:SCAb

[WDS16] Hao Wang, Sanhong Deng, and Xinning Su. A study on construction and analysis of discipline knowledge structure of Chinese LIS based on CSSCI. Scientometrics, 109(3):1725–1759, December 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2146-4.

Webb:2016:TUP

[Web16] Stephen Webb. Twitter use in physics conferences. Scientometrics, 108(3):1267–1286, September 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-016-2031-1.pdf.

Wong:2017:STI

[WF17] Chan-Yuan Wong and Hon-Ngen Fung. Science-technology-industry correlative indicators for policy targeting on emerging technologies: exploring the core competencies and promising industries of aspirant economies. Scientometrics, 111(2):841–867, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Wei:2018:SBR

[WF18] Yaoyu Wei and Weiwei Fan. A study of book reviews in SCI–Expanded, SSCI, and A&HCI journals by researchers from five countries: 2006–2015. Scientometrics, 115(2):637– 654, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2679-9.

Wang:2016:TDF

[WFG16] Xianwen Wang, Zhichao Fang, and Xinhui Guo. Tracking the digital footprints to scholarly articles from social media. Scientometrics, 109(2):1365–1376, November 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2086-z.

Wolcott:2016:MTD

[WFH^{+}16] Holly N. Wolcott, Matthew J. Fouch, Elizabeth R. Hsu, Leo G. DiJoseph, Catherine A. Bernaciak, James G. Corrigan, and Duane E. Williams. Modeling time-dependent andindependent indicators to facilitate identification of breakthrough research papers. Scientometrics, 107(2):807–817, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1861-1.

Wang:2016:UPS

[WFS16] Xianwen Wang, Zhichao Fang, and Xiaoling Sun. Usage patterns of scholarly articles on Web of Science: a study on Web of Science usage count. Scientometrics, 109(2):917–926, November 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2093-0.

Wang:2019:CWP

[WFZD19a] Fenghua Wang, Ying Fan, An Zeng, and Zengru Di. Can we predict ESI highly cited publications? Scientometrics, 118 (1):109–125, January 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2965-6.

Wang:2019:NCC

[WFZD19b] Fenghua Wang, Ying Fan, An Zeng, and Zengru Di. A nonlinear collective credit allocation in scientific publications. Scientometrics, 119(3):1655–1668, June 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03107-x.

Wong:2010:MBS

[WG10] Chan-Yuan Wong and Kim-Leng Goh. Modeling the behaviour of science and technology: self-propagating growth in the diffusion process. Scientometrics, 84(3):669– 686, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0220-x.

Watts:2011:DCA

[WG11] Christopher Watts and Nigel Gilbert. Does cumulative advantage affect collective learning in science? An agent-based simulation. Scientometrics, 89(1):437–463, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0432-8.

Wong:2012:PDS

[WG12] Chan-Yuan Wong and Kim-Leng Goh. The pathway of development: science and technology of NIEs and selected Asian emerging economies. Scientometrics, 92(3): 523–548, September 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0622-z.

Wen:2012:TPO

[WH12] Hang Wen and Yi Huang. Trends and performance of oxidative stress research from 1991 to 2010. Scientometrics, 91(1):51–63, April 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0535-2.

Huang:2015:TEM

[wH15] Ding wei Huang. Temporal evolution of multi-author papers in basic sciences from 1960 to 2010. Scientometrics, 105(3):

2137–2147, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1760-x.

Wang:2016:RTM

[WH16] Chong-Chen Wang and Yuh-Shan Ho. Research trend of metal-organic frameworks: a bibliometric analysis. Scientometrics, 109(1):481–513, October 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-1986-2.

Wang:2013:CUE

[WHC^{+}13] Yuandi Wang, Jiashun Huang, Yantai Chen, Xiongfeng Pan, and Jin Chen. Have Chinese universities embraced their third mission? New insight from a business perspective. Scientometrics, 97(2):207–222, November 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1055-z.

Wang:2010:NSI

[WhCL10] Jyun-Cheng Wang, Cheng hsin Chiang, and Shu-Wei Lin. Network structure of innovation: can brokerage or closure predict patent quality? Scientometrics, 84(3):735– 748, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0211-y.

Wylie:2018:FDJ

[WHH^{+}18] Lindsey E. Wylie, Katherine P. Hazen, Lori A. Hoetger, Joshua A. Haby, and Eve M. Brank. Four decades of the journal Law and Human Behavior: a content analysis. Scientometrics, 115(2):655–693, May 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2685-y.

White:2010:SNT

[Whi10] Howard D. White. Some new tests of relevance theory in information science. Scientometrics, 83(3):653–667, June 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0138-3.

White:2015:CCA

[Whi15] Howard D. White. Co-cited author retrieval and relevance theory: examples from the humanities. Scientometrics, 102 (3):2275–2299, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1483-4.

White:2018:PGB

[Whi18] Howard D. White. Pennants for Garfield: bibliometrics and document retrieval. Scientometrics, 114(2):757–778, February 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2610-9.

Wang:2015:CSE

[WHL^{+}15] Yuandi Wang, Die Hu, Weiping Li, Yiwei Li, and Qiang Li. Collaboration strategies and effects on university research: evidence from Chinese universities. Scientometrics, 103(2):725–749, May 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1552-3.

Wang:2016:DTB

[WHLP16] Yuandi Wang, Ruifeng Hu, Weiping Li, and Xiongfeng Pan. Does teaching benefit from university-industry collaboration? Investigating the role of academic commercialization and engagement. Scientometrics, 106(3):1037–1055, March 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1818-9.

Wang:2019:PRI

[WHS19] Hei-Chia Wang, Tzu-Ting Hsu, and Yunita Sari. Personal research idea recommendation using research trends and a hierarchical topic model. Scientometrics, 121(3):1385– 1406, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03258-x.

Wang:2019:WDG

[WHW^{+}19] Yanwen Wang, Song Hong, Yifei Wang, Xi Gong, Chao He, Zhendong Lu, and F. Benjamin Zhan. What is the difference in global research on Central Asia before and after the collapse of the USSR: a bibliometric analysis. Scientometrics, 119(2):909–930, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03069-0.

Hsu:2011:CBI

[wHwH11] Jiann wien Hsu and Ding wei Huang. Correlation between impact and collaboration. Scientometrics, 86(2):317–324, February 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0265-x.

Hsu:2016:IFD

[wHwH16] Jiann wien Hsu and Ding wei Huang. Impact factor distribution revisited with graphical representation. Scientometrics, 107(3):1321–1329, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1921-6.

Wu:2014:DIM

[WHZ14] Shue Wu, Zhenlei Huang, and Weizhou Zhong. Does inertia matter for parts manufacturers’ innovation? Scientometrics, 101(1):705–716, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1191-5.

Wildgaard:2015:CAL

[Wil15] Lorna Wildgaard. A comparison of 17 author-level bibliometric indicators for researchers in astronomy, environmental science, philosophy and public health in Web of Science and Google Scholar. Scientometrics, 104(3):873– 906, September 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1608-4.

Wang:2019:BJL

[WJCC19] Mingyang Wang, Shijia Jiao, Kah-Hin Chai, and Guangsheng Chen. Building journal’s long-term impact: using indicators detected from the sustained active articles. Scientometrics, 121(1):261–283, October 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03196-8.

Wu:2015:DIR

[WJD15] Jiang Wu, Miao Jin, and Xiu-Hao Ding. Diversity of individual research disciplines in scientific funding. Scientometrics, 103(2):669–686, May 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1549-y.

Wood:2015:ITN

[WK15] Jacob Wood and Gohar Feroz Khan. International trade negotiation analysis: network and semantic knowledge infrastructure. Scientometrics, 105(1):537–556, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1651-1.

Wang:2017:CAB

[WK17] Shenghui Wang and Rob Koopman. Clustering articles based on semantic similarity. Scientometrics, 111(2):1017– 1031, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Wu:2019:FKD

[WKHS19] Qingqiang Wu, Yichen Kuang, Qingqi Hong, and Yingying She. Frontier knowledge discovery and visualization in cancer field based on KOS and LDA. Scientometrics, 118 (3):979–1010, March 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2989-y.

Wood:2016:WEO

[WKK16] Jacob Wood, Woocheol Kim, and Gohar Feroz Khan. Work engagement in organizations: a social network analysis of the domain. Scientometrics, 109(1):317–336, October 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-1974-6.

Wang:2014:BAR

[WL14] Jieyong Wang and Zhigao Liu. A bibliometric analysis on rural studies in human geography and related disciplines. Scientometrics, 101(1):39–59, October 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1388-2.

Wei:2018:IBN

[WL18] Yaoyu Wei and Lei Lei. Institution bias in the New England Journal of Medicine? A bibliometric analysis of publications (1997–2016). Scientometrics, 117(3):1771–1775, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2948-7.

Wang:2017:DLR

[WLC17] Mingyang Wang, Shi Li, and Guangsheng Chen. Detecting latent referential articles based on their vitality performance in the latest 2 years. Scientometrics, 112(3):1557– 1571, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2433-8.

Wang:2014:ITT

[WLD^{+}14] Bo Wang, Shengbo Liu, Kun Ding, Zeyuan Liu, and Jing Xu. Identifying technological topics and institutiontopic distribution probability for patent competitive intelligence analysis: a case study in LTE technology. Scientometrics, 101(1):685–704, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1342-3.

Wang:2012:SFR

[WLDW12] Xianwen Wang, Di Liu, Kun Ding, and Xinran Wang. Science funding and research output: a study on 10 countries. Scientometrics, 91(2):591–599, May 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0576-6.

Wang:2015:IDW

[WLF15] Ling Ling Wang, Xuan Zhen Liu, and Hui Fang. Investigation of the degree to which articles supported by research grants are published in open access health and life sciences journals. Scientometrics, 104(2):511–528, August 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1624-4.

Wu:2017:TEA

[WLH^{+}17] Feifei Wu, Ruiyu Li, Lucheng Huang, Hong Miao, and Xin Li. Theme evolution analysis of electrochemical energy storage research based on CitNetExplorer. Scientometrics, 110(1):113–139, January 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2164-2.

Wang:2013:HRB

[WLHZ13] Haijun Wang, Minyan Liu, Song Hong, and Yanhua Zhuang. A historical review and bibliometric analysis of GPS research from 1991–2010. Scientometrics, 95(1):35–44, April 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0853-z.

Wang:2012:RSB

[WLLL12] Zhong-Yi Wang, Gang Li, Chun-Ya Li, and Ang Li. Research on the semantic-based co-word analysis. Scientometrics, 90(3):855–875, March 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
011-0563-y.

Wang:2015:DPB

[WLM15] Xianwen Wang, Chen Liu, and Wenli Mao. Does a paper being featured on the cover of a journal guarantee more attention and greater impact? Scientometrics, 102(2):1815–1821, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1456-7.

Wang:2015:EOA

[WLMF15a] Xianwen Wang, Chen Liu, Wenli Mao, and Zhichao Fang. Erratum to: The open access advantage considering citation, article usage and social media attention. Scientometrics, 103(3):1149, June 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-015-1589-3.pdf. See [WLMF15b].

Wang:2015:OAA

[WLMF15b] Xianwen Wang, Chen Liu, Wenli Mao, and Zhichao Fang. The open access advantage considering citation, article usage and social media attention. Scientometrics, 103(2):555– 564, May 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-015-1547-0. See erratum

[WLMF15a].

Wang:2014:DPT

[WLN^{+}14] Yuandi Wang, Jian Li, Lutao Ning, Deming Zeng, and Xin Gu. Dynamic patterns of technology collaboration: a case study of the Chinese automobile industry, 1985– 2010. Scientometrics, 101(1):663–683, October 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1366-8.

Wu:2014:UAD

[WLPH14] Hao Wu, Bo Li, Yijian Pei, and Jun He. Unsupervised author disambiguation using Dempster–Shafer theory. Scientometrics, 101(3):1955–1972, December 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1283-x.

Wang:2014:CNP

[WLR^{+}14] Xuefeng Wang, Rongrong Li, Shiming Ren, Donghua Zhu, Meng Huang, and Pengjun Qiu. Collaboration network and pattern analysis: case study of dye-sensitized solar cells. Scientometrics, 98(3):1745–1762, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1180-8.

Wang:2014:HDB

[WLY14] Yuandi Wang and Jason Li-Ying. How do the BRIC countries play their roles in the global innovation arena? A study based on USPTO patents during 1990–2009. Scientometrics, 98(2):1065–1083, February 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1141-2.

Wu:2015:RRP

[WLZ^{+}15] Dengsheng Wu, Minglu Li, Xiaoqian Zhu, Hongfang Song, and Jianping Li. Ranking the research productivity of business and management institutions in Asia–Pacific region: empirical research in leading ABS journals. Scientometrics, 105(2):1253–1272, November 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1752-x.

Wang:2019:CTS

[WLZ^{+}19] Pancheng Wang, Shasha Li, Haifang Zhou, Jintao Tang, and Ting Wang. Cited text spans identification with an improved balanced ensemble model. Scientometrics, 120(3):1111–1145, September 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03167-z.

Wang:2017:LIR

[WM17] Xuemei Wang and Mingguo Ma. The luminous intensity of regional ‘night-light’ output can predict the growing volume of published scientific research by ‘luminaries’ in developing countries. Scientometrics, 110(2):1005–1010, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2188-7.

Walters:2019:DFJ

[WM19] William H. Walters and Susanne Markgren. Do faculty journal selections correspond to objective indicators of citation impact? Results for 20 academic departments at Manhattan College. Scientometrics, 118(1):321–337, January 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2972-7.

Wang:2017:CSS

[WMH^{+}17] Xuefeng Wang, Pingping Ma, Ying Huang, Junfang Guo, Donghua Zhu, Alan L. Porter, and Zhinan Wang. Combining SAO semantic analysis and morphology analysis to identify technology opportunities. Scientometrics, 111(1):3– 24, April 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-017-2260-y.

Wang:2012:CCE

[WMT^{+}12] Yuan Wang, Ruimin Ma, Tianhao Tang, Xia Liu, Ping Xie, Jianxin Wang, Jianxiang Liu, Haideng Zhou, and Siwei Zhang. The comprehensive competitiveness evaluation of American universities in bridge engineering. Scientometrics, 91(3):693–701, June 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0616-x.

Wang:2013:CEB

[WMW^{+}13] Xianwen Wang, Wenli Mao, Chuanli Wang, Lian Peng, and Haiyan Hou. Chinese elite brain drain to USA: an investigation of 100 United States national universities. Scientometrics, 97(1):37–46, October 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1024-6.

Wang:2014:UHS

[WMXZ14] Xianwen Wang, Wenli Mao, Shenmeng Xu, and Chunbo Zhang. Usage history of scientific literature: Nature metrics and metrics of Nature publications. Scientometrics, 98(3): 1923–1933, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1167-5.

Wang:2013:INR

[WNS13] Lili Wang, Ad Notten, and Alexandru Surpatean. Interdisciplinarity of nano research fields: a keyword mining approach. Scientometrics, 94(3):877–892, March 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0856-9.

Wolfram:2015:SRB

[Wol15] Dietmar Wolfram. The symbiotic relationship between information retrieval and informetrics. Scientometrics, 102(3): 2201–2214, March 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1479-0.

Wong:2013:PCD

[Won13] Chan-Yuan Wong. On a path to creative destruction: science, technology and science-based technological trajectories of Japan and south Korea. Scientometrics, 96(1):323– 336, July 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0941-0.

Wong:2019:CSP

[Won19] Chan-Yuan Wong. A century of scientific publication: towards a theorization of growth behavior and researchorientation. Scientometrics, 119(1):357–377, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-019-03048-5.

Wiles:2010:EBQ

[WOW10] Louise Wiles, Timothy Olds, and Marie Williams. Evidence base, quantitation and collaboration: three novel indices for bibliometric content analysis. Scientometrics, 85 (1):317–328, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0163-2.

Wiles:2013:TFY

[WOW13] Louise Wiles, Timothy Olds, and Marie Williams. Twentyfive years of Australian nursing and allied health professional journals: bibliometric analysis from 1985 through 2010. Scientometrics, 94(1):359–378, January 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0704-y.

Weismayer:2017:IER

[WP17] Christian Weismayer and Ilona Pezenka. Identifying emerging research fields: a longitudinal latent semantic keyword analysis. Scientometrics, 113(3):1757–1785, December 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2555-z.

Wallace:2018:EBB

[WP18] Frederick H. Wallace and Timothy J. Perri. Economists behaving badly: publications in predatory journals. Scientometrics, 115(2):749–766, May 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2690-1.

Wang:2013:DRT

[WPCG13] Yuandi Wang, Xiongfeng Pan, Yantai Chen, and Xin Gu. Do references in transferred patent documents signal learning opportunities for the receiving firms? Scientometrics, 95 (2):731–752, May 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-0962-3.

Wang:2014:VKS

[WPW^{+}14] Yuandi Wang, Xin Pan, Xinyu Wang, Jin Chen, Lutao Ning, and Ying Qin. Visualizing knowledge space: a case study of Chinese licensed technology, 2000–2012. Scientometrics, 98 (3):1935–1954, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1135-0.

Wang:2012:RCC

[WQY12] Feifei Wang, Junping Qiu, and Houqiang Yu. Research on the cross-citation relationship of core authors in scientometrics. Scientometrics, 91(3):1011–1033, June 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0621-0.

Wray:2010:RSS

[Wra10] K. Brad Wray. Rethinking the size of scientific specialties: correcting Price’s estimate. Scientometrics, 83(2):471–476, May 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0060-8.

Wray:2014:SPP

[Wra14] K. Brad Wray. Specialization in philosophy: a preliminary study. Scientometrics, 98(3):1763–1769, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1102-9.

Wray:2016:NNE

[Wra16a] K. Brad Wray. No new evidence for a citation benefit for Author-Pay Open Access Publications in the social sciences and humanities. Scientometrics, 106(3):1031–1035, March 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1833-5.

Wray:2016:SNN

[Wra16b] K. Brad Wray. Still no new evidence: Author-Pay Open Access in the social sciences and humanities. Scientometrics, 107(3):1527–1529, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
016-1907-4.

Wray:2018:NMN

[Wra18] K. Brad Wray. A note on measuring normal science. Scientometrics, 117(1):647–650, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2870-z. See reply [Pet18b].

Wang:2019:MPS

[WRC^{+}19] Xuefeng Wang, Huichao Ren, Yun Chen, Yuqin Liu, Yali Qiao, and Ying Huang. Measuring patent similarity with SAO semantic analysis. Scientometrics, 121(1):1–23, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03191-z.

Walter:2017:BBB

[WRM17] Lothar Walter, Alfred Radauer, and Martin G. Moehrle. The beauty of brimstone butterfly: novelty of patents identified by near environment analysis based on text mining. Scientometrics, 111(1):103–115, April 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-017-2267-4.

Wang:2014:HFD

[WRV14] Yuandi Wang, Nadine Roijakkers, and Wim Vanhaverbeke. How fast do Chinese firms learn and catch up? Evidence from patent citations. Scientometrics, 98(1):743–761, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1016-6.

Wong:2010:UPA

[WS10] Poh Kam Wong and Annette Singh. University patenting activities and their link to the quantity and quality of scientific publications. Scientometrics, 83(1):271–294, April 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0003-4.

Wang:2011:FAA

[WS11] Jue Wang and Philip Shapira. Funding acknowledgement analysis: an enhanced tool to investigate research sponsorship impacts: the case of nanotechnology. Scientometrics, 87(3):563–586, June 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0362-5.

Wong:2013:DCP

[WS13a] Poh Kam Wong and Annette Singh. Do co-publications with industry lead to higher levels of university technology commercialization activity? Scientometrics, 97(2): 245–265, November 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1029-1.

Wuehrer:2013:KDA

[WS13b] Gerhard A. Wuehrer and Angela Elisabeth Smejkal. The knowledge domain of the Academy of International Business Studies (AIB) conferences: a longitudinal scientometric perspective for the years 2006–2011. Scientometrics, 95 (2):541–561, May 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0909-0.

Wang:2016:SCD

[WSC16] Hao Wang, Hua-Wei Shen, and Xue-Qi Cheng. Scientific credit diffusion: Researcher level or paper level? Scientometrics, 109(2):827–837, November 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2057-4.

Wang:2016:TES

[WSH16] Chun-Chieh Wang, Hui-Yun Sung, and Mu-Hsuan Huang. Technological evolution seen from the USPC reclassifications. Scientometrics, 107(2):537–553, May 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1851-3.

Wildgaard:2014:RCA

[WSL14] Lorna Wildgaard, Jesper W. Schneider, and Birger Larsen. A review of the characteristics of 108 author-level bibliometric indicators. Scientometrics, 101(1):125–158, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1423-3.

Wilkinson:2014:SCE

[WST14] David Wilkinson, Pardeep Sud, and Mike Thelwall. Substance without citation: evaluating the online impact of grey literature. Scientometrics, 98(2):797–806, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-1068-7.

Winnink:2014:RDS

[WT14] J. J. Winnink and Robert J. W. Tijssen. R&D dynamics and scientific breakthroughs in HIV/AIDS drugs development: the case of integrase inhibitors. Scientometrics, 101(1):1–16, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1330-7.

Winnink:2015:ESI

[WT15] J. J. Winnink and Robert J. W. Tijssen. Early stage identification of breakthroughs at the interface of science and technology: lessons drawn from a landmark publication. Scientometrics, 102(1):113–134, January 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1451-z.

Wang:2015:CIC

[WTG15] Lei Wang, Bart Thijs, and Wolfgang Glänzel. Characteristics of international collaboration in sport sciences publications and its influence on citation impact. Scientometrics, 105(2):843–862, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1735-y.

Wang:2016:ABA

[WTM^{+}16] Yunfei Wang, Siming Tan, Yanyan Ma, Xia Zhao, Zhiling Wang, Zhiyong Chu, and Honghua Qin. Application of bibliometrics in analysis of output differences among countries under International Ocean Discovery Program. Scientometrics, 109(1):447–462, October 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2052-9.

Wu:2013:GKD

[Wu13] Jiang Wu. Geographical knowledge diffusion and spatial diversity citation rank. Scientometrics, 94(1):181–201, January 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0715-8.

Wu:2014:CTI

[Wu14] Ching-Yan Wu. Comparisons of technological innovation capabilities in the solar photovoltaic industries of Taiwan, China, and Korea. Scientometrics, 98(1):429–446, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1120-7.

Wu:2018:TIL

[Wu18] Jinshan Wu. Is there an intrinsic logical error in null hypothesis significance tests? Commentary on: “Null hypothesis significance tests. A mix-up of two different theories: the basis for widespread confusion and numerous misinterpretations”. Scientometrics, 115(1):621–625, April 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2656-3. See [Sch15a].

Wainer:2013:CBB

[WV13] Jacques Wainer and Paula Vieira. Correlations between bibliometrics and peer evaluation for all disciplines: the evaluation of Brazilian scientists. Scientometrics, 96(2):395–410, August 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-0969-9.

Waaijer:2011:MBB

[WvBvE11] Cathelijn J. F. Waaijer, Cornelis A. van Bochove, and Nees Jan van Eck. On the map: Nature and Science editorials. Scientometrics, 86(1):99–112, January 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-010-0205-9.pdf.

Waltman:2013:SNI

[WvE13] Ludo Waltman and Nees Jan van Eck. Source normalized indicators of citation impact: an overview of different approaches and an empirical comparison. Scientometrics, 96(3):699–716, September 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0913-4.

Waltman:2011:CBB

[WvEvL^{+}11a] Ludo Waltman, Nees Jan van Eck, Thed N. van Leeuwen, Martijn S. Visser, and Anthony F. J. van Raan. On the correlation between bibliometric indicators and peer review: reply to Opthof and Leydesdorff. Scientometrics, 88(3): 1017–1022, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-011-

0425-7.pdf. See [OL11, Pra12a].

Waltman:2011:TNC

[WvEvL^{+}11b] Ludo Waltman, Nees Jan van Eck, Thed N. van Leeuwen, Martijn S. Visser, and Anthony F. J. van Raan. Towards a new crown indicator: an empirical analysis. Scientometrics, 87(3):467–481, June 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-011-

0354-5.pdf. See comment [OL11] and reply [WvEvL^{+}11a].

Wu:2011:IEP

[WW11] Qiang Wu and Dietmar Wolfram. The influence of effects and phenomena on citations: a comparative analysis of four citation perspectives. Scientometrics, 89(1):245–258, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0456-0.

Wagner:2012:USR

[WW12] Caroline S. Wagner and Shing Kit Wong. Unseen science? Representation of BRICs in global science. Scientometrics, 90(3):1001–1013, March 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
011-0481-z.

Walters:2015:WCL

[WW15] William H. Walters and Esther Isabelle Wilder. Worldwide contributors to the literature of library and information science: top authors, 2007–2012. Scientometrics, 103 (1):301–327, April 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1519-9.

Wang:2019:WCB

[WWC19] Mingyang Wang, Zhenyu Wang, and Guangsheng Chen. Which can better predict the future success of articles? Bibliometric indices or alternative metrics. Scientometrics, 119 (3):1575–1595, June 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03052-9.

Wang:2017:MIR

[WWH^{+}17] Xuefeng Wang, Zhinan Wang, Ying Huang, Yun Chen, Yi Zhang, Huichao Ren, Rongrong Li, and Jinhui Pang. Measuring interdisciplinarity of a research system: detecting distinction between publication categories and citation categories. Scientometrics, 111(3):2023–2039, June 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Wagner:2017:GIC

[WWL17] Caroline S. Wagner, Travis A. Whetsell, and Loet Leydesdorff. Growth of international collaboration in science: revisiting six specialties. Scientometrics, 110(3):1633–1652, March 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2230-9.

Wang:2014:ICB

[WWP14] Weichao Wang, Yishan Wu, and Yuntao Pan. An investigation of collaborations between top Chinese universities: a new quantitative approach. Scientometrics, 98(2):1535– 1545, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1072-y.

Wang:2017:NSS

[WWP17] Lili Wang, Xianwen Wang, and Niels J. Philipsen. Network structure of scientific collaborations between China and the EU member states. Scientometrics, 113(2):765–781, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2488-6;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2488-
6.pdf.

Wang:2013:TSR

[WWX13] Xianwen Wang, Zhi Wang, and Shenmeng Xu. Tracing scientist’s research trends realtimely. Scientometrics, 95 (2):717–729, May 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0884-5.

Wang:2012:RCA

[WXLL12] Xianwen Wang, Shenmeng Xu, Di Liu, and Yongxia Liang. The role of Chinese–American scientists in China–US scientific collaboration: a study in nanotechnology. Scientometrics, 91(3):737–749, June 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
012-0693-x.

Wang:2013:ISC

[WXW^{+}13] Xianwen Wang, Shenmeng Xu, Zhi Wang, Lian Peng, and Chuanli Wang. International scientific collaboration of China: collaborating countries, institutions and individuals. Scientometrics, 95(3):885–894, June 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0877-4.

Wang:2016:BAR

[WXZ^{+}16] Yuan Wang, Cuiyun Xiang, Peng Zhao, Guozhu Mao, and Huibin Du. A bibliometric analysis for the research on river water quality assessment and simulation during 2000–2014. Scientometrics, 108(3):1333–1346, September 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2014-2.

Wong:2012:MTI

[WY12] Chan-Yuan Wong and Xiao-Shan Yap. Mapping technological innovations through patent analysis: a case study of foreign multinationals and indigenous firms in China. Scientometrics, 91(3):773–787, June 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0595-3.

Wang:2019:DCN

[WY19] Jingbei Wang and Naiding Yang. Dynamics of collaboration network community and exploratory innovation: the moderation of knowledge networks. Scientometrics, 121(2): 1067–1084, November 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03235-4.

Wang:2012:DFI

[WYAY12] Mingyang Wang, Guang Yu, Shuang An, and Daren Yu. Discovery of factors influencing citation impact based on a soft fuzzy rough set model. Scientometrics, 93(3):635– 644, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0766-x.

Wang:2017:SCP

[WYB^{+}17] Wei Wang, Shuo Yu, Teshome Megersa Bekele, Xiangjie Kong, and Feng Xia. Scientific collaboration patterns vary with scholars’ academic ages. Scientometrics, 112(1):329– 343, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Wang:2010:BAP

[WYH10] Ming-Huang Wang, Te-Chen Yu, and Yuh-Shan Ho. A bibliometric analysis of the performance of water research. Scientometrics, 84(3):813–820, September 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0112-0.

Waltman:2011:RFN

[WYvE11] Ludo Waltman, Erjia Yan, and Nees Jan van Eck. A recursive field-normalized bibliometric performance indicator: an application to the field of library and information science. Scientometrics, 89(1):301–314, October 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/content/pdf/
10.1007/s11192-011-0449-z.pdf.

Wang:2011:MTF

[WYY11] Mingyang Wang, Guang Yu, and Daren Yu. Mining typical features for highly cited papers. Scientometrics, 87(3):695– 706, June 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0366-1.

Wu:2017:SIV

[WZ17] Qiang Wu and Peng Zhang. Some indices violating the basic domination relation. Scientometrics, 113(1):495–500, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2475-y.

Wang:2019:EME

[WZ19a] Chongfeng Wang and Gupeng Zhang. Examining the moderating effect of technology spillovers embedded in the intra- and inter-regional collaborative innovation networks of China. Scientometrics, 119(2):561–593, May 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03084-1.

Wu:2019:FHI

[WZ19b] Xiaolan Wu and Chengzhi Zhang. Finding high-impact interdisciplinary users based on friend discipline distribution in academic social networking sites. Scientometrics, 119 (2):1017–1035, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03067-2.

Wang:2019:EIO

[WZCC19] Mingyang Wang, Jiaqi Zhang, Guangsheng Chen, and Kah-Hin Chai. Examining the influence of open access on journals’ citation obsolescence by modeling the actual citation process. Scientometrics, 119(3):1621–1641, June 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03085-0.

Wang:2019:RSP

[WZFD19] Yanan Wang, An Zeng, Ying Fan, and Zengru Di. Ranking scientific publications considering the aging characteristics of citations. Scientometrics, 120(1):155–166, July 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03117-9.

Wang:2013:KTA

[WZLZ13] Xuezhao Wang, Yajuan Zhao, Rui Liu, and Jing Zhang. Knowledge-transfer analysis based on co-citation clustering. Scientometrics, 97(3):859–869, December 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1077-6.

Wang:2012:LPW

[WZS12] Dongbo Wang, Danhao Zhu, and Xinning Su. Lotka phenomenon in the words’ syntactic distribution complexity. Scientometrics, 90(2):483–498, February 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0546-z.

Wang:2015:SNM

[WZW15] Zongshui Wang, Hong Zhao, and Yan Wang. Social networks in marketing research 2001–2014: a co-word analysis. Scientometrics, 105(1):65–82, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1672-9.

Wang:2011:PCC

[WZX11] Xianwen Wang, Xi Zhang, and Shenmeng Xu. Patent co-citation networks of fortune 500 companies. Scientometrics, 88(3):761–770, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0414-x.

Xu:2015:CIS

[XA15] Ziyun Xu and Éric Archambault. Chinese interpreting studies: structural determinants of MA students’ career choices. Scientometrics, 105(2):1041–1058, November 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1717-0.

Xu:2018:UFI

[XBD^{+}18] Jian Xu, Yi Bu, Ying Ding, Sinan Yang, Hongli Zhang, Chen Yu, and Lin Sun. Understanding the formation of interdisciplinary research from the perspective of keyword evolution: a case study on joint attention. Scientometrics, 117(2):973–995, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2897-1.

Xiao:2016:ETH

[XCS^{+}16] Lu Xiao, Guo Chen, Jianjun Sun, Shuguang Han, and Chengzhi Zhang. Exploring the topic hierarchy of digital library research in China using keyword networks: a K-core decomposition approach. Scientometrics, 108(3):1085–1101, September 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2051-x.

Xu:2019:ISC

[XDB^{+}19] Jian Xu, Ying Ding, Yi Bu, Shuqing Deng, Chen Yu, Yimin Zou, and Andrew Madden. Interdisciplinary scholarly communication: an exploratory study for the field of joint attention. Scientometrics, 119(3):1597–1619, June 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03106-y.

Xu:2018:NBS

[XG18] Ran Xu and Navid Ghaffarzadegan. Neuroscience bridging scientific disciplines in health: Who builds the bridge, who pays for it? Scientometrics, 117(2):1183–1204, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2887-3.

Xie:2019:CBP

[XGCK19] Juan Xie, Kaile Gong, Ying Cheng, and Qing Ke. The correlation between paper length and citations: a metaanalysis. Scientometrics, 118(3):763–786, March 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03015-0.

Xie:2019:PFW

[XGL^{+}19] Juan Xie, Kaile Gong, Jiang Li, Qing Ke, Hyonchol Kang, and Ying Cheng. A probe into 66 factors which are possibly associated with the number of citations an article received. Scientometrics, 119(3):1429–1454, June 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03094-z.

Xu:2016:ITI

[XGY^{+}16] Haiyun Xu, Ting Guo, Zenghui Yue, Lijie Ru, and Shu Fang. Interdisciplinary topics of information science: a study based on the terms interdisciplinarity index series. Scientometrics, 106(2):583–601, February 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1792-2.

Xu:2019:TDE

[XHA^{+}19] Shuo Xu, Liyuan Hao, Xin An, Dongsheng Zhai, and Hongshen Pang. Types of DOI errors of cited references in Web of Science with a cleaning method. Scientometrics, 120(3):1427–1437, September 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03162-4.

Xie:2015:SIA

[Xie15] Ping Xie. Study of international anticancer research trends via co-word and document co-citation visualization analysis. Scientometrics, 105(1):611–622, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1689-0.

Xie:2019:CGM

[Xie19] Zheng Xie. A cooperative game model for the multimodality of coauthorship networks. Scientometrics, 121(1):503–519, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03183-z.

Xu:2015:ISI

[XLR15] Fang Xu, Wenbin Liu, and Ronald Rousseau. Introducing sub-impact factor (SIF-) sequences and an aggregated SIF-indicator for journal ranking. Scientometrics, 102(2): 1577–1593, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1401-9.

Xu:2018:OTS

[XLZ^{+}18] Shuo Xu, Junwan Liu, Dongsheng Zhai, Xin An, Zheng Wang, and Hongshen Pang. Overlapping thematic structures extraction with mixed-membership stochastic blockmodel. Scientometrics, 117(1):61–84, October 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2841-4.

Xu:2013:RTB

[XM13] Li Xu and Dora Marinova. Resilience thinking: a bibliometric analysis of socio-ecological research. Scientometrics, 96(3):911–927, September 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-0957-0.

Su:2015:SCE

[xShLY^{+}15] Liang xing Su, Peng hui Lyu, Zheng Yang, Shuai Ding, and Kai le Zhou. Scientometric cognitive and evaluation on smart city related construction and building journals data. Scientometrics, 105(1):449–470, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1697-0.

Xu:2015:FRS

[XTZ15] Xin Xu, Alice M. Tan, and Star X. Zhao. Funding ratios in social science: the perspective of countries/ territories level and comparison with natural sciences. Scientometrics, 104(3):673–684, September 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1633-3.

Xie:2019:ETS

[XWL19] Yundong Xie, Qiang Wu, and Xingchen Li. Editorial team scholarly index (ETSI): an alternative indicator for evaluating academic journal reputation. Scientometrics, 120(3):1333–1349, September 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03177-x.

Xie:2017:MCB

[XXL^{+}17] Zheng Xie, Zonglin Xie, Miao Li, Jianping Li, and Dongyun Yi. Modeling the coevolution between citations and coauthorship of scientific papers. Scientometrics, 112(1):483– 507, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Xu:2018:WWW

[XYHD18] Shenmeng Xu, Houqiang Yu, Bradley M. Hemminger, and Xie Dong. Who, what, why? An exploration of JoVE scientific video publications in tweets. Scientometrics, 117(2): 845–856, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2880-x.

Xu:2017:MSD

[XYW^{+}17] Hai-Yun Xu, Zeng-Hui Yue, Chao Wang, Kun Dong, Hong-Shen Pang, and Zhengbiao Han. Multi-source data fusion study in scientometrics. Scientometrics, 111(2):773–792, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Xing:2019:SNE

[XZFD19] Yanmeng Xing, An Zeng, Ying Fan, and Zengru Di. The strong nonlinear effect in academic dropout. Scientometrics, 120(2):793–805, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
019-03135-7.

Xiang:2015:WER

[XZZ15] Huimin Xiang, Jiaen Zhang, and Qiandong Zhu. Worldwide earthworm research: a scientometric analysis, 2000– 2015. Scientometrics, 105(2):1195–1207, November 2015.

CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1725-0.

Yin:2010:CEI

[YAC10] Chun-Yang Yin, Mohd Jindra Aris, and Xi Chen. Combination of Eigenfactor^{T M} and h-index to evaluate scientific journals. Scientometrics, 84(3):639–648, September 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-009-0116-9.

Yan:2014:TBP

[Yan14] Erjia Yan. Topic-based PageRank: toward a topic-level scientific evaluation. Scientometrics, 100(2):407–437, August 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1308-5.

Youtie:2014:SDR

[YB14] Jan Youtie and Barry Bozeman. Social dynamics of research collaboration: norms, practices, and ethical issues in determining co-authorship rights. Scientometrics, 101(2): 953–962, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1391-7.

Yang:2010:ARC

[YC10] Phil Yihsing Yang and Yuan-Chieh Chang. Academic research commercialization and knowledge production and diffusion: the moderating effects of entrepreneurial commitment. Scientometrics, 83(2):403–421, May 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0075-1.

Yi:2012:OSK

[YC12] Sangyoon Yi and Jinho Choi. The organization of scientific knowledge: the structural characteristics of keyword networks. Scientometrics, 90(3):1015–1026, March 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0560-1.

Yoon:2011:IPF

[YCK11] Janghyeok Yoon, Sungchul Choi, and Kwangsoo Kim. Invention property-function network analysis of patents: a case of silicon-based thin film solar cells. Scientometrics, 86 (3):687–703, March 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0303-8.

Yang:2013:GTS

[YCL^{+}13a] Lie Yang, Zhulei Chen, Ting Liu, Zhe Gong, Yingjian Yu, and Jia Wang. Global trends of solid waste research from 1997 to 2011 by using bibliometric analysis. Scientometrics, 96(1):133–146, July 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0911-6. See comments [LFLG14].

Yang:2013:ROA

[YCL^{+}13b] Lie Yang, Zhulei Chen, Ting Liu, Rui Wan, Jia Wang, and Wengang Xie. Research output analysis of municipal solid waste: a case study of China. Scientometrics, 96(2): 641–650, August 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-0982-z.

Youtie:2017:TRT

[YCPS17] Jan Youtie, Stephen Carley, Alan L. Porter, and Philip Shapira. Tracking researchers and their outputs: new insights from ORCIDs. Scientometrics, 113(1):437–453, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2473-0.

Yan:2012:OCT

[YDJ12] Erjia Yan, Ying Ding, and Elin K. Jacob. Overlaying communities and topics: an analysis on publication networks. Scientometrics, 90(2):499–513, February 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0531-6.

Yan:2010:MLI

[YDZ10] Erjia Yan, Ying Ding, and Qinghua Zhu. Mapping library and information science in China: a coauthorship network analysis. Scientometrics, 83(1):115–131, April 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0027-9.

Ye:2014:DSB

[Ye14] Jiyuan Ye. Development, significance and background information about the “Chinese Book Citation Index” (CBkCI) demonstration database. Scientometrics, 98(1):557–564, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1092-7.

Yan:2018:HMN

[YG18] Yan Yan and Jiancheng Guan. How multiple networks help in creating knowledge: evidence from alternative energy patents. Scientometrics, 115(1):51–77, April 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2638-5.

Yair:2017:LLE

[YGD17] Gad Yair, Nofar Gueta, and Nitza Davidovitch. The law of limited excellence: publication productivity of Israel Prize laureates in the life and exact sciences. Scientometrics, 113 (1):299–311, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2465-0.

Yang:2015:MEC

[YGW^{+}15] Xue Yang, Xin Gu, Yuandi Wang, Guangyuan Hu, and Li Tang. The Matthew effect in China’s science: evidence from academicians of Chinese Academy of Sciences. Scientometrics, 102(3):2089–2105, March 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1502-5.

Yang:2015:GTC

[YHC^{+}15] Lie Yang, Lizi Han, Zhulei Chen, Jiabin Zhou, and Jia Wang. Growing trend of China’s contribution to haze research. Scientometrics, 105(1):525–535, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1668-5.

Yuan:2018:WIR

[YHL^{+}18] Lili Yuan, Yanni Hao, Minglu Li, Chunbing Bao, Jianping Li, and Dengsheng Wu. Who are the international research collaboration partners for China? A novel data perspective based on NSFC grants. Scientometrics, 116(1):401–422, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2753-3.

Yang:2019:NAJ

[YHL19] Lijun Yang, Liangxiu Han, and Naxin Liu. A new approach to journal co-citation matrix construction based on the number of co-cited articles in journals. Scientometrics, 120(2): 507–517, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03141-9.

Yanagisawa:2010:ONS

[YIK^{+}10] Kazuaki Yanagisawa, Keishiro Ito, Chisato Katsuki, Kei Kawashima, and Masashi Shirabe. An outcome of nuclear safety research in JAERI: case study for LOCA. Scientometrics, 84(3):563–573, September 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0195-7.

Yi:2011:BST

[YJ11] Huang Yi and Wang Jie. A bibliometric study of the trend in articles related to eutrophication published in Science Citation Index. Scientometrics, 89(3):919–927, December 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0479-6. See comments [CH12] and

response [HW12].

Yoon:2011:IRE

[YK11] Janghyeok Yoon and Kwangsoo Kim. Identifying rapidly evolving technological trends for R&D planning using SAObased semantic patent networks. Scientometrics, 88(1):213– 228, July 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0383-0.

Yoon:2012:DSN

[YK12] Janghyeok Yoon and Kwangsoo Kim. Detecting signals of new technological opportunities using semantic patent analysis and outlier detection. Scientometrics, 90(2):445–461, February 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0543-2.

Yasukawa:2014:VUE

[YK14] Satoshi Yasukawa and Shingo Kano. Validating the usefulness of examiners’ forward citations from the viewpoint of applicants’ self-selection during the patent application procedure. Scientometrics, 99(3):895–909, June 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1195-1.

Yasukawa:2015:CEF

[YK15] Satoshi Yasukawa and Shingo Kano. Comparison of examiners’ forward citations in the United States and Japan with pairs of equivalent patent applications. Scientometrics, 102(2):1189–1205, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1466-5.

Yeo:2013:QAR

[YKCK13] Woondong Yeo, Seonho Kim, Byoung-Youl Coh, and Jaewoo Kang. A quantitative approach to recommend promising technologies for SME innovation: a case study on knowledge arbitrage from LCD to solar cell. Scientometrics, 96 (2):589–604, August 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0935-y.

Yeo:2014:ASM

[YKLK14] Woondong Yeo, Seonho Kim, Jae-Min Lee, and Jaewoo Kang. Aggregative and stochastic model of main path identification: a case study on graphene. Scientometrics, 98(1): 633–655, January 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1140-3.

Yu:2010:IRC

[YL10] Guang Yu and Yi-Jun Li. Identification of referencing and citation processes of scientific journals based on the citation distribution model. Scientometrics, 82(2):249–261, February 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0085-z.

Yang:2012:APP

[YL12] Kiduk Yang and Jongwook Lee. Analysis of publication patterns in Korean library and information science research. Scientometrics, 93(2):233–251, November 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0663-3.

Yan:2018:FIS

[YLC18] Yan Yan, Zhewen Liao, and Xiaosong Chen. Fixedincome securities: bibliometric review with network analysis. Scientometrics, 116(3):1615–1640, September 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2800-0.

You:2017:DTF

[YLH^{+}17] Hanlin You, Mengjun Li, Keith W. Hipel, Jiang Jiang, Bingfeng Ge, and Hante Duan. Development trend forecasting for coherent light generator technology based on patent citation network analysis. Scientometrics, 111(1):297–315, April 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-017-2252-y.

Yang:2018:NNM

[YLH18] Liu Yang, Keping Li, and Hangfei Huang. A new network model for extracting text keywords. Scientometrics, 116 (1):339–361, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2743-5.

You:2017:EMI

[YLJ^{+}17] Hanlin You, Mengjun Li, Jiang Jiang, Bingfeng Ge, and Xueting Zhang. Evolution monitoring for innovation sources using patent cluster analysis. Scientometrics, 111(2):693– 715, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Yoon:2010:DAK

[YLL10] Byungun Yoon, Sungjoo Lee, and Gwanghee Lee. Development and application of a keyword-based knowledge map for effective R&D planning. Scientometrics, 85(3):803–820, December 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0294-5.

Yan:2015:ARP

[YLL15a] Bei-Ni Yan, Tian-Shyug Lee, and Tsung-Pei Lee. Analysis of research papers on e-commerce (2000–2013): based on a text mining approach. Scientometrics, 105(1):403–417, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1675-6.

Yan:2015:MIS

[YLL15b] Bei-Ni Yan, Tian-Shyug Lee, and Tsung-Pei Lee. Mapping the intellectual structure of the Internet of Things (IoT) field (2000–2014): a co-word analysis. Scientometrics, 105(2): 1285–1300, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1740-1.

Yang:2015:UCP

[YLL^{+}15c] Guan-Can Yang, Gang Li, Chun-Ya Li, Yun-Hua Zhao, Jing Zhang, Tong Liu, Dar-Zen Chen, and Mu-Hsuan Huang. Using the Comprehensive Patent Citation network (CPC) to evaluate patent value. Scientometrics, 105(3):1319– 1346, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1763-7.

Yang:2016:DIF

[YLSW16] Dong-Hui Yang, Xin Li, Xiaoxia Sun, and Jie Wan. Detecting impact factor manipulation with data mining techniques. Scientometrics, 109(3):1989–2005, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2144-6.

Yao:2014:CPF

[YLY^{+}14] Qiang Yao, Peng-Hui Lyu, Lian-Ping Yang, Lan Yao, and Zhi-Yong Liu. Current performance and future trends in health care sciences and services research. Scientometrics, 101(1):751–779, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1383-7.

Yang:2010:LAC

[YMSQ10] Siluo Yang, Feng Ma, Yanhui Song, and Junping Qiu. A longitudinal analysis of citation distribution breadth for Chinese scholars. Scientometrics, 85(3):755–765, December 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0245-1.

Yoon:2015:ESK

[Yoo15] Jungwon Yoon. The evolution of South Korea’s innovation system: moving towards the triple helix model? Scientometrics, 104(1):265–293, July 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
015-1541-6.

Yoshikane:2013:MRA

[Yos13] Fuyuki Yoshikane. Multiple regression analysis of a patent’s citation frequency and quantitative characteristics: the case of Japanese patents. Scientometrics, 96(1):365–379, July 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-0953-4.

Youtie:2014:UCS

[You14] Jan Youtie. The use of citation speed to understand the effects of a multi-institutional science center. Scientometrics, 100(3):613–621, September 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1297-4.

Youk:2019:WWD

[YP19] Sungbin Youk and Hee Sun Park. Where and what do they publish? Editors’ and editorial board members’ affiliated institutions and the citation counts of their endogenous publications in the field of communication. Scientometrics, 120(3):1237–1260, September 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03169-x.

Yang:2010:NAI

[YPH10] Chang Hoon Yang, Han Woo Park, and Jungeun Heo. A network analysis of interdisciplinary research relationships: the Korean government’s R&D grant program. Scientometrics, 83(1):77–92, April 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0157-0.

Yoon:2013:ITC

[YPK13] Janghyeok Yoon, Hyunseok Park, and Kwangsoo Kim. Identifying technological competition trends for R&D planning using dynamic patent maps: SAO-based content analysis. Scientometrics, 94(1):313–331, January 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0830-6.

Yau:2014:CSD

[YPNS14] Chyi-Kwei Yau, Alan Porter, Nils Newman, and Arho Suominen. Clustering scientific documents with topic modeling. Scientometrics, 100(3):767–786, September 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1321-8.

Yi:2013:CNB

[YQW13] Yong Yi, Wei Qi, and Dandan Wu. Are CIVETS the next BRICs? A comparative analysis from scientometrics perspective. Scientometrics, 94(2):615–628, February 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0791-9.

Yang:2010:ESU

[YQX10] Siluo Yang, Junping Qiu, and Zunyan Xiong. An empirical study on the utilization of web academic resources in humanities and social sciences based on web citations. Scientometrics, 84(1):1–19, July 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0142-7.

Ye:2010:PCI

[YR10] Fred Y. Ye and Ronald Rousseau. Probing the h-core: an investigation of the tail-core ratio for rank distributions. Scientometrics, 84(2):431–439, August 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0099-6.

Yang:2013:ELB

[YS13] Bo Yang and Ying Sun. An exploration of link-based knowledge map in academic web space. Scientometrics, 96(1):239– 253, July 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0919-y.

Yoshikane:2014:DFP

[YS14] Fuyuki Yoshikane and Takafumi Suzuki. Diversity of fields in patent citations: synchronic and diachronic changes. Scientometrics, 98(3):1879–1897, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1165-7.

Yu:2011:RGO

[YSD11] Qi Yu, Hongfang Shao, and Zhiguang Duan. Research groups of oncology co-authorship network in China. Scientometrics, 89(2):553–567, November 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0465-z.

Yaghtin:2019:QND

[YSM^{+}19] Maryam Yaghtin, Hajar Sotudeh, Mahdieh Mirzabeigi, Seyed Mostafa Fakhrahmad, and Mehdi Mohammadi. In quest of new document relations: evaluating co-opinion relations between co-citations and its impact on information retrieval effectiveness. Scientometrics, 119(2):987–1008, May 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03058-3.

Yi:2017:RIC

[YSND17] Nannan Yi, Nicolas Standaert, Benoit Nemery, and Kris Dierickx. Research integrity in China: precautions when searching the Chinese literature. Scientometrics, 110(2):1011–1016, February 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2191-z.

Yoshikane:2012:ARB

[YST12] Fuyuki Yoshikane, Yutaka Suzuki, and Keita Tsuji. Analysis of the relationship between citation frequency of patents and diversity of their backward citations for Japanese patents. Scientometrics, 92(3):721–733, September 2012. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0612-6.

Yeh:2013:BCA

[YSY^{+}13] Hsi-Yin Yeh, Yi-Shan Sung, Hsiao-Wen Yang, Wan-Chu Tsai, and Dar-Zen Chen. The bibliographic coupling approach to filter the cited and uncited patent citations: a case of electric vehicle technology. Scientometrics, 94(1):75–93, January 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0820-8.

Tsay:2011:BAC

[yT11] Ming yueh Tsay. A bibliometric analysis and comparison on three information science journals: JASIST, IPM, JOD, 1998–2008. Scientometrics, 89(2):591–606, November 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0460-4.

Tsay:2015:KFD

[yT15] Ming yueh Tsay. Knowledge flow out of the domain of information science: a bibliometric and citation analysis study. Scientometrics, 102(1):487–502, January 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1339-y.

Tsay:2016:CCD

[yTmShL16] Ming yueh Tsay, Tung mei Shen, and Ming hsin Liang. A comparison of citation distributions of journals and books on the topic “information society”. Scientometrics, 106(2): 475–508, February 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1791-3.

Tsay:2017:BAJ

[yTnL17] Ming yueh Tsay and Chia ning Li. Bibliometric analysis of the journal literature on women’s studies. Scientometrics, 113(2):705–734, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2493-9.

Tsay:2019:CUC

[yTwTlW19] Ming yueh Tsay, Yu wei Tseng, and Tai luan Wu. Comprehensiveness and uniqueness of commercial databases and open access systems. Scientometrics, 121(3):1323–1338, December 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03252-3.

Yu:2015:SRA

[Yu15] Dejian Yu. A scientometrics review on aggregation operator research. Scientometrics, 105(1):115–133, October 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-015-1695-2.

Yu:2017:CAD

[Yu17] Houqiang Yu. Context of altmetrics data matters: an investigation of count type and user category. Scientometrics, 111(1):267–283, April 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-017-2251-z.

Yuret:2015:ICA

[Yur15] Tolga Yuret. Interfield comparison of academic output by using department level data. Scientometrics, 105(3):1653– 1664, December 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1621-7.

Yuret:2016:DAS

[Yur16a] Tolga Yuret. Does alphabetization significantly affect academic careers? Scientometrics, 108(3):1603–1619, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2058-3.

Yuret:2016:ITI

[Yur16b] Tolga Yuret. International trade in ideas. Scientometrics, 107(3):899–916, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1883-8.

Yuret:2018:AWI

[Yur18a] Tolga Yuret. Author-weighted impact factor and reference return ratio: can we attain more equality among fields? Scientometrics, 116(3):2097–2111, September 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2806-7.

Yuret:2018:PSA

[Yur18b] Tolga Yuret. Path to success: an analysis of US educated elite academics in the United States. Scientometrics, 117 (1):105–121, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2850-3.

Yuret:2018:TTA

[Yur18c] Tolga Yuret. Tenure and turnover of academics in six undergraduate programs in the United States. Scientometrics, 116(1):101–124, July 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2742-6.

Yang:2012:ITV

[YWC12] Ying Yang, Mingzhi Wu, and Lei Cui. Integration of three visualization methods based on co-word analysis. Scientometrics, 90(2):659–673, February 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0541-4.

Yarwood:2014:LTB

[YWG14] Maree R. Yarwood, Michael A. Weston, and Stephen T. Garnett. From little things, big things grow; trends and fads in 110 years of Australian ornithology. Scientometrics, 98 (3):2235–2254, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1144-z.

Yan:2016:DHT

[YWL16] Zhenbin Yan, Qiang Wu, and Xingchen Li. Do Hirschtype indices behave the same in assessing single publications? An empirical study of 29 bibliometric indicators. Scientometrics, 109(3):1815–1833, December 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2147-3.

Yan:2018:FFC

[YWS18] Erjia Yan, Chaojiang Wu, and Min Song. The funding factor: a cross-disciplinary examination of the association between research funding and citation impact. Scientometrics, 115(1):369–384, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2583-8.

Yang:2017:RBA

[YWW17] Siluo Yang, Dietmar Wolfram, and Feifei Wang. The relationship between the author byline and contribution lists: a comparison of three general medical journals. Scientometrics, 110(3):1273–1296, March 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2239-0.

Yu:2010:CKD

[YWY10] Guang Yu, Ming-Yang Wang, and Da-Ren Yu. Characterizing knowledge diffusion of nanoscience & nanotechnology by citation analysis. Scientometrics, 84(1):81–97, July 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-009-0090-2.

Yu:2017:MLM

[YWZ^{+}17] Dejian Yu, Wanru Wang, Shuai Zhang, Wenyu Zhang, and Rongyu Liu. A multiple-link, mutually reinforced journal-ranking model to measure the prestige of journals. Scientometrics, 111(1):521–542, April 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-017-2262-9.

Yang:2018:DIO

[YXW18] Siluo Yang, Xin Xing, and Dietmar Wolfram. Difference in the impact of open-access papers published by China and the USA. Scientometrics, 115(2):1017–1037, May 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2697-7.

Yamashita:2014:IRI

[YY14] Yasuhiro Yamashita and Daisuke Yoshinaga. Influence of researchers’ international mobilities on publication: a comparison of highly cited and uncited papers. Scientometrics, 101(2):1475–1489, November 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1384-6.

Yu:2016:DAJ

[YY16] Liping Yu and Houqiang Yu. Does the average JIF percentile make a difference? Scientometrics, 109(3):1979–1987, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2156-2.

Yang:2012:CDS

[YYDH12] Li Ying Yang, Ting Yue, Jie Lan Ding, and Tao Han. A comparison of disciplinary structure in science between the G7 and the BRIC countries by bibliometric methods. Scientometrics, 93(2):497–516, November 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0695-8.

Yu:2010:RBC

[YYL10] Guang Yu, Dong-Hui Yang, and Wang Liang. Reliabilitybased citation impact factor and the manipulation of impact factor. Scientometrics, 83(1):259–270, April 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0083-1.

Yu:2014:CIP

[YYLW14] Tian Yu, Guang Yu, Peng-Yu Li, and Liang Wang. Citation impact prediction for scientific papers using stepwise regression analysis. Scientometrics, 101(2):1233–1252, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1279-6.

Yoon:2017:QHS

[YYP17] Jungwon Yoon, Joshua SungWoo Yang, and Han Woo Park. Quintuple helix structure of Sino–Korean research collaboration in science. Scientometrics, 113(1):61–81, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2476-x.

Yuan:2010:PAW

[YYS^{+}10] Jun Peng Yuan, Wei Ping Yue, Cheng Su, Zheng Wu, Zheng Ma, Yun Tao Pan, Nan Ma, Zhi Yu Hu, Fei Shi, Zheng Lu Yu, and Yi Shan Wu. Patent activity on water pollution and treatment in China-a scientometric perspective. Scientometrics, 83(3):639–651, June 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-009-0126-7.

Yin:2017:DAE

[YZ17] Zhifeng Yin and Qiang Zhi. Dancing with the academic elite: a promotion or hindrance of research production? Scientometrics, 110(1):17–41, January 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2151-7.

Yan:2018:AFF

[YZB18] Weiwei Yan, Yin Zhang, and Wendy Bromfield. Analyzing the follower-followee ratio to determine user characteristics and institutional participation differences among research universities on ResearchGate. Scientometrics, 115 (1):299–316, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2637-6.

Yang:2017:ROC

[YZW^{+}17] Chao Yang, Donghua Zhu, Xuefeng Wang, Yi Zhang, Guangquan Zhang, and Jie Lu. Requirement-oriented core technological components’ identification based on SAO analysis. Scientometrics, 112(3):1229–1248, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2444-5.

Zanjirchi:2019:FDF

[ZAJ19] Seyed Mahmoud Zanjirchi, Mina Rezaeian Abrishami, and Negar Jalilian. Four decades of fuzzy sets theory in operations management: application of life-cycle, bibliometrics and content analysis. Scientometrics, 119(3):1289–1309, June 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03077-0.

Zoss:2012:MIW

[ZB12] Angela M. Zoss and Katy Börner. Mapping interactions within the evolving science of science and innovation policy community. Scientometrics, 91(2):631–644, May 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0574-8.

Zhou:2015:OAP

[ZB15] Ping Zhou and Lutz Bornmann. An overview of academic publishing and collaboration between China and Germany. Scientometrics, 102(2):1781–1793, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-014-1418-0.

Zhao:2014:AAC

[ZC14] Rongying Zhao and Bikun Chen. Applying author cocitation analysis to user interaction analysis: a case study on instant messaging groups. Scientometrics, 101(2):985– 997, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1314-7.

Zuccala:2016:DMI

[ZC16] Alesia Zuccala and Roberto Cornacchia. Data matching, integration, and interoperability for a metric assessment of monographs. Scientometrics, 108(1):465–484, July 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-016-1911-8.

Zdravkovic:2016:EPS

[ZCKZ16] Marta Zdravkovic, Linley Chiwona-Karltun, and Eren Zink. Experiences and perceptions of South–South and North– South scientific collaboration of mathematicians, physicists and chemists from five southern African universities. Scientometrics, 108(2):717–743, August 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1989-z.

Zhang:2014:CEM

[ZCL14] Zhihui Zhang, Ying Cheng, and Nian Cai Liu. Comparison of the effect of mean-based method and z-score for field normalization of citations at the level of Web of Science subject categories. Scientometrics, 101(3):1679–1693, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1294-7.

Zhang:2015:INE

[ZCL15] Zhihui Zhang, Ying Cheng, and Nian Cai Liu. Improving the normalization effect of mean-based method from the perspective of optimization: optimization-based linear methods and their performance. Scientometrics, 102(1):587–607, January 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1398-0.

Zulueta:2011:RIW

[ZCMVQS11] M. Angeles Zulueta, Gisela Cantos-Mateos, Benjamín Vargas-Quesada, and Carmen Sánchez. Research involving women and health in the Medline database, 1965– 2005: co-term analysis and visualization of main lines of research. Scientometrics, 88(3):679–706, September 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-011-0455-1.

Zahedi:2014:HWD

[ZCW14] Zohreh Zahedi, Rodrigo Costas, and Paul Wouters. How well developed are altmetrics? A cross-disciplinary analysis of the presence of ‘alternative metrics’ in scientific publications. Scientometrics, 101(2):1491–1513, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-014-1264-0.

Zhang:2016:IOS

[ZCZ^{+}16] Yi Zhang, Kaihua Chen, Guilong Zhu, Richard C. M. Yam, and Jiancheng Guan. Inter-organizational scientific collaborations and policy effects: an ego-network evolutionary perspective of the Chinese Academy of Sciences.

Scientometrics, 108(3):1383–1415, September 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-2022-2.

Zhuang:2015:RTH

[ZDZ^{+}15] Yanhua Zhuang, Chao Du, Liang Zhang, Yun Du, and Sisi Li. Research trends and hotspots in soil erosion from 1932 to 2013: a literature review. Scientometrics, 105(2): 743–758, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1706-3.

Zelnio:2012:IGC

[Zel12] Ryan Zelnio. Identifying the global core-periphery structure of science. Scientometrics, 91(2):601–615, May 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0598-0.

Zhang:2017:ATD

[ZFY^{+}17] Guijie Zhang, Yuqiang Feng, Guang Yu, Luning Liu, and Yanqiqi Hao. Analyzing the time delay between scientific research and technology patents based on the citation distribution model. Scientometrics, 111(3):1287–1306, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Zhou:2010:DAC

[ZG10] Ping Zhou and Wolfgang Glänzel. In-depth analysis on China’s international cooperation in science. Scientometrics, 82(3):597–612, March 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
010-0174-z.

Zhao:2011:ICT

[ZG11] Qingjun Zhao and Jiancheng Guan. International collaboration of three ‘giants’ with the G7 countries in emerging nanobiopharmaceuticals. Scientometrics, 87(1):159–170, April 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0311-8.

Zhang:2012:WDM

[ZG12a] Lin Zhang and Wolfgang Glänzel. Where demographics meets scientometrics: towards a dynamic career analysis. Scientometrics, 91(2):617–630, May 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0590-8.

Zhao:2012:MDR

[ZG12b] Qingjun Zhao and Jiancheng Guan. Modeling the dynamic relation between science and technology in nanotechnology. Scientometrics, 90(2):561–579, February 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0520-9.

Zhang:2013:IMP

[ZG13a] Minghua Zhang and Michael L. Grieneisen. The impact of misconduct on the published medical and non-medical literature, and the news media. Scientometrics, 96(2):573–587, August 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0920-5.

Zhao:2013:LDB

[ZG13b] Qingjun Zhao and Jiancheng Guan. Love dynamics between science and technology: some evidences in nanoscience and nanotechnology. Scientometrics, 94(1):113–132, January 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0785-7.

Zhu:2013:BSS

[ZG13c] Wenjia Zhu and Jiancheng Guan. A bibliometric study of service innovation research: based on complex network analysis. Scientometrics, 94(3):1195–1216, March 2013. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0888-1.

Zhang:2017:HIM

[ZG17a] Chao Zhang and Jiancheng Guan. How to identify metaknowledge trends and features in a certain research field?

Evidences from innovation and entrepreneurial ecosystem. Scientometrics, 113(2):1177–1197, November 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2503-y.

Zhang:2017:CIR

[ZG17b] Cui Zhang and Jing Guo. China’s international research collaboration: evidence from a panel gravity model. Scientometrics, 113(2):1129–1139, November 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2513-9.

Zhang:2017:SRI

[ZG17c] JingJing Zhang and Jiancheng Guan. Scientific relatedness and intellectual base: a citation analysis of un-cited and highly-cited papers in the solar energy field. Scientometrics, 110(1):141–162, January 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2155-3.

Zhang:2017:CBCa

[ZG17d] Lin Zhang and Wolfgang Glänzel. A citation-based crossdisciplinary study on literature ageing: part II — diachronous aspects. Scientometrics, 111(3):1559–1572, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Zhang:2017:CBCb

[ZG17e] Lin Zhang and Wolfgang Glänzel. A citation-based crossdisciplinary study on literature aging: part I — the synchronous approach. Scientometrics, 111(3):1573–1589, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Zuo:2017:MRE

[ZG17f] Kairui Zuo and Jiancheng Guan. Measuring the R&D efficiency of regions by a parallel DEA game model. Scientometrics, 112(1):175–194, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Zacca-Gonzalez:2018:MSO

[ZGCRVQ18] Grisel Zacca-González, Zaida Chinchilla-Rodríguez, and Benjamín Vargas-Quesada. Medical scientific output and specialization in Latin American countries. Scientometrics, 115(3):1635–1650, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

http://link.springer.com/article/10.1007/s11192-
018-2717-7.

Zhou:2018:IBL

[ZGJ18] Wen Zhou, Jiayi Gu, and Yifan Jia. h-index-based link prediction methods in citation network. Scientometrics, 117 (1):381–390, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2867-7.

Zhang:2014:ISW

[ZGL14] Gupeng Zhang, Jiancheng Guan, and Xielin Liu. The impact of small world on patent productivity in China. Scientometrics, 98(2):945–960, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1142-1.

Zheng:2017:CEA

[ZGL^{+}17] Jie Zheng, Jianya Gong, Rui Li, Kai Hu, Huayi Wu, and Siluo Yang. Community evolution analysis based on coauthor network: a case study of academic communities of the journal of “Annals of the Association of American Geographers”. Scientometrics, 113(2):845–865, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2515-7.

Zhang:2016:DEC

[ZGY16] Lin Zhang, Wolfgang Glänzel, and Fred Y. Ye. The dynamic evolution of core documents: an experimental study based on h-related literature (2005–2013). Scientometrics, 106(1): 369–381, January 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1705-4.

Zhu:2017:VKD

[ZH17] Jie Zhu and Weijian Hua. Visualizing the knowledge domain of sustainable development research between 1987 and 2015: a bibliometric analysis. Scientometrics, 110(2):893–914, February 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2187-8.

Zhao:2010:CIG

[Zha10] Dangzhi Zhao. Characteristics and impact of grant-funded research: a case study of the library and information science field. Scientometrics, 84(2):293–306, August 2010. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-010-0191-y.

Zhang:2014:GTC

[Zha14] Li Zhang. Growing trend of China’s contribution to tissue engineering. Scientometrics, 98(2):1423–1433, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1111-8.

Zhang:2017:EJI

[Zha17] Fuli Zhang. Evaluating journal impact based on weighted citations. Scientometrics, 113(2):1155–1169, November 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2510-z.

Zhao:2018:ERG

[Zha18] Qu Zhao. Electromobility research in Germany and China: structural differences. Scientometrics, 117(1):473–493, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2873-9.

Zheng:2019:UMI

[Zhe19] Lukun Zheng. Using mutual information as a cocitation similarity measure. Scientometrics, 119(3):1695–1713, June 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03098-9.

Zanotto:2016:UIC

[ZHG16] Sonia R. Zanotto, Cristina Haeffner, and Jorge A. Guimarães. Unbalanced international collaboration affects adversely the usefulness of countries’ scientific output as well as their technological and social impact. Scientometrics, 109 (3):1789–1814, December 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-016-
2126-8.pdf.

Zhu:2019:DEP

[ZHL19] Junwen Zhu, Guangyuan Hu, and Weishu Liu. DOI errors and possible solutions for Web of Science. Scientometrics, 118(2):709–718, February 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2980-7.

Zhu:2014:MRR

[ZHMX14] Jia Zhu, Saeed-Ul Hassan, Hamid Turab Mirza, and Qing Xie. Measuring recent research performance for Chinese universities using bibliometric methods. Scientometrics, 101 (1):429–443, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1389-1.

Zhu:2017:WSO

[Zhu17] Yimei Zhu. Who support open access publishing? Gender, discipline, seniority and other factors associated with academics’ OA practice. Scientometrics, 111(2):557–579, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-017-2316-z.pdf.

Zhou:2019:HAD

[ZHZY19] Xiao Zhou, Lu Huang, Yi Zhang, and Miaomiao Yu. A hybrid approach to detecting technological recombination based on text mining and patent network analysis. Scientometrics, 121(2):699–737, November 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03218-5.

Zinilli:2016:CPF

[Zin16] Antonio Zinilli. Competitive project funding and dynamic complex networks: evidence from Projects of National Interest (PRIN). Scientometrics, 108(2):633–652, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-016-1976-4.

Zitt:2011:BCS

[Zit11] M. Zitt. Behind citing-side normalization of citations: some properties of the journal impact factor. Scientometrics, 89 (1):329–344, October 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-011-0441-7.

Zitt:2012:JIF

[Zit12] Michel Zitt. The journal impact factor: angel, devil, or scapegoat? A comment on J. K. Vanclay’s article 2011. Scientometrics, 92(2):485–503, August 2012. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-012-0697-6.

Zitt:2015:MLR

[Zit15] Michel Zitt. Meso-level retrieval: IR-bibliometrics interplay and hybrid citation-words methods in scientific fields delineation. Scientometrics, 102(3):2223–2245, March 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1482-5.

Zhang:2010:JCC

[ZJLG10] Lin Zhang, Frizo Janssens, Liming Liang, and Wolfgang Glänzel. Journal cross-citation analysis for validation and improvement of journal-based subject classification in bibliometric research. Scientometrics, 82(3):687–706, March 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0180-1.

Zemtsov:2019:ARI

[ZK19] Stepan Zemtsov and Maxim Kotsemir. An assessment of regional innovation system efficiency in Russia: the application of the DEA approach. Scientometrics, 120(2):375–404, August 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03130-y.

Zhang:2016:MBR

[ZKC^{+}16] Yi Zhang, Mingting Kou, Kaihua Chen, Jiancheng Guan, and Yuchen Li. Modelling the Basic Research Competitiveness Index (BR-CI) with an application to the biomass energy field. Scientometrics, 108(3):1221–1241, September 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
accesspage/article/10.1007/s11192-016-2042-y.

Zavadskas:2011:SPR

[ZKD11] Edmundas Kazimieras Zavadskas, Raimundas Kirvaitis, and Eleonora Dagiene. Scientific publications released in the Baltic states. Scientometrics, 88(1):179–190, July 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0394-x.

Zheng:2015:MII

[ZL15a] Juntao Zheng and Niancai Liu. Mapping of important international academic awards. Scientometrics, 104(3):763– 791, September 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1613-7.

Zhou:2015:APC

[ZL15b] Ping Zhou and Xiaozan Lv. Academic publishing and collaboration between China and Germany in physics. Scientometrics, 105(3):1875–1887, December 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1643-1.

Zou:2017:MCC

[ZL17] Yawen Zou and Manfred D. Laubichler. Measuring the contributions of Chinese scholars to the research field of systems biology from 2005 to 2013. Scientometrics, 110(3):1615–1631, March 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2213-x.

Zavale:2018:UIL

[ZL18a] Nelson Casimiro Zavale and Patrício Vitorino Langa. University-industry linkages’ literature on Sub-Saharan Africa: systematic literature review and bibliometric account. Scientometrics, 116(1):1–49, July 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2760-4.

Zdenek:2018:AEB

[ZL18b] Radek Zdenek and Jana Lososová. An analysis of editorial board members’ publication output in agricultural economics and policy journals. Scientometrics, 117(1):563–578, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2881-9.

Zhang:2014:CIN

[ZLF^{+}14] Guijie Zhang, Luning Liu, Yuqiang Feng, Zhen Shao, and Yongli Li. Cext-N index: a network node centrality measure for collaborative relationship distribution. Scientometrics, 101(1):291–307, October 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1358-8.

Zong:2018:SBN

[ZLF18] ZhangJian Zong, XuanZhen Liu, and Hui Fang. Sleeping beauties with no prince based on the co-citation criterion. Scientometrics, 117(3):1841–1852, December 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2932-2.

Zhai:2015:RST

[ZLG^{+}15] Xing Zhai, Zhihong Li, Kuo Gao, Youliang Huang, Lin Lin, and Le Wang. Research status and trend analysis of global biomedical text mining studies in recent 10 years. Scientometrics, 105(1):509–523, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1700-9.

Zhu:2015:KCO

[ZLH^{+}15] Lin Zhu, Xiantao Liu, Sha He, Jun Shi, and Ming Pang. Keywords co-occurrence mapping knowledge domain research base on the theory of Big Data in oil and gas industry. Scientometrics, 105(1):249–260, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1658-7.

Zhao:2016:CRC

[ZLH^{+}16] Yong Zhao, Dong Li, Mingjie Han, Chenying Li, and Dongmei Li. Characteristics of research collaboration in biotechnology in China: evidence from publications indexed in the SCIE. Scientometrics, 107(3):1373–1387, June 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-016-1898-1.

Zhang:2015:OCI

[ZLL^{+}15] Liang Zhang, Sisi Li, Hugo A. Loáiciga, Yanhua Zhuang, and Yun Du. Opportunities and challenges of interbasin water transfers: a literature review with bibliometric analysis. Scientometrics, 105(1):279–294, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1656-9.

Zhang:2017:DCW

[ZLL^{+}17] Qian-Ru Zhang, Yue Li, Jia-Shu Liu, Yi-Dan Chen, and Li-He Chai. A dynamic co-word network-related approach on the evolution of China’s urbanization research. Scientometrics, 111(3):1623–1642, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Zhu:2019:SBW

[ZLL19] Junwen Zhu, Fang Liu, and Weishu Liu. The secrets behind Web of Science’s DOI search. Scientometrics, 119(3): 1745–1753, June 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03091-2.

Zhou:2019:NMI

[ZLLD19] Yuan Zhou, Heng Lin, Yufei Liu, and Wei Ding. A novel method to identify emerging technologies using a semisupervised topic clustering model: a case of 3D printing industry. Scientometrics, 120(1):167–185, July 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03126-8;   http://link.springer.
com/content/pdf/10.1007/s11192-019-03126-8.pdf.

Zhao:2019:DSC

[ZLLL19] Rongying Zhao, Xinlai Li, Zhisen Liang, and Danyang Li. Development strategy and collaboration preference in S&T of enterprises based on funded papers: a case study of Google. Scientometrics, 121(1):323–347, October 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03182-0.

Zhuang:2013:GRS

[ZLN^{+}13] Yanhua Zhuang, Xingjian Liu, Thuminh Nguyen, Qingqing He, and Song Hong. Global remote sensing research trends during 1991–2010: a bibliometric analysis. Scientometrics, 96(1):203–219, July 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-012-0918-z.

Zhang:2014:BAR

[ZLT^{+}14] Baogang Zhang, Ye Liu, Caixing Tian, Zhijun Wang, Ming Cheng, Nan Chen, and Chuanping Feng. A bibliometric analysis of research on upflow anaerobic sludge blanket (UASB) from 1983 to 2012. Scientometrics, 100(1):189–202, July 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1189-z.

Zhao:2018:DFP

[ZLTY18] Star X. Zhao, Wen Lou, Alice M. Tan, and Shuang Yu. Do funded papers attract more usage? Scientometrics, 115 (1):153–168, April 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2662-5.

Zhang:2016:SSC

[ZLW16] Jing Zhang, Xiaomin Liu, and Lili Wu. The study of subjectclassification based on journal coupling and expert subjectclassification system. Scientometrics, 107(3):1149–1170, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-016-1890-9.

Zhang:2019:KNM

[ZLW19] Guijie Zhang, Luning Liu, and Fangfang Wei. Key nodes mining in the inventor–author knowledge diffusion network. Scientometrics, 118(3):721–735, March 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03005-2.

Zhai:2014:EAC

[ZLYF14] Li Zhai, Xiujuan Li, Xiangbin Yan, and Weiguo Fan. Evolutionary analysis of collaboration networks in the field of information systems. Scientometrics, 101(3):1657–1677, December 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1360-1.

Zhang:2019:SMP

[ZLZ19] Tingting Zhang, Baozhen Lee, and Qinghua Zhu. Semantic measure of plagiarism using a hierarchical graph model. Scientometrics, 121(1):209–239, October 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03204-x.

Zhi:2016:FAI

[ZM16] Qiang Zhi and Tianguang Meng. Funding allocation, inequality, and scientific research output: an empirical study based on the life science sector of Natural Science Foundation of China. Scientometrics, 106(2):603–628, February 2016. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-015-1773-5.

Zhang:2019:STR

[ZML19] Qi Zhang, Rui Mao, and Rui Li. Spatial-temporal restricted supervised learning for collaboration recommendation. Scientometrics, 119(3):1497–1517, June 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03100-4.

Zhang:2018:CTP

[ZMW^{+}18] Yongjun Zhang, Jialin Ma, Zijian Wang, Bolun Chen, and Yongtao Yu. Collective topical PageRank: a model to evaluate the topic-dependent academic impact of scientific papers. Scientometrics, 114(3):1345–1372, March 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2626-1.

Zhang:2017:ETF

[ZNB^{+}17] Jun Zhang, Zhaolong Ning, Xiaomei Bai, Xiangjie Kong, Jinmeng Zhou, and Feng Xia. Exploring time factors in measuring the scientific impact of scholars. Scientometrics, 112(3):1301–1321, September 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2458-z.

Zong:2019:RDC

[Zon19] Qianjin Zong. Response to Dr. Copiello’s comments on “The impact of video abstract on citation counts”. Scientometrics, 120(3):1499–1504, September 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03174-0. See [ZXT^{+}19].

Zhou:2015:CAP

[ZP15] Ping Zhou and Youneng Pan. A comparative analysis of publication portfolios of selected economies. Scientometrics, 105(2):825–842, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1707-2.

Zou:2016:QSO

[ZP16] Christopher Zou and Jordan B. Peterson. Quantifying the scientific output of new researchers using the zpindex. Scientometrics, 106(3):901–916, March 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1807-z.

Zhang:2017:STM

[ZPC17] Yi Zhang, Alan L. Porter, and Denise Chiavetta. Scientometrics for tech mining. Scientometrics, 111(3):1875–1878, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s11192-017-2343-9.pdf.

Zhai:2014:ICS

[ZPG^{+}14] Lihua Zhai, Yuntao Pan, Yu Guo, Zheng Ma, and Fei Bi. International comparative study on nanofiltration membrane technology based on relevant publications and patents. Scientometrics, 101(2):1361–1374, November 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1394-4.

Zhang:2017:EBI

[ZQH^{+}17] Yi Zhang, Yue Qian, Ying Huang, Ying Guo, Guangquan Zhang, and Jie Lu. An entropy-based indicator system for measuring the potential of patents in technological innovation: rejecting moderation. Scientometrics, 111(3):1925– 1946, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Zhang:2018:CEC

[ZRL18] Zhihui Zhang, Jason E. Rollins, and Evangelia Lipitakis. China’s emerging centrality in the contemporary international scientific collaboration network. Scientometrics, 116(2):1075–1091, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2788-5.

Zhou:2012:GFD

[ZRY^{+}12] Qiuju Zhou, Ronald Rousseau, Liying Yang, Ting Yue, and Guoliang Yang. A general framework for describing diversity within systems and similarity between systems with applications in informetrics. Scientometrics, 93(3):787–812, December 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0767-9.

Zhao:2011:ISS

[ZS11] Dangzhi Zhao and Andreas Strotmann. Intellectual structure of stem cell research: a comprehensive author cocitation analysis of a highly collaborative and multidisciplinary field. Scientometrics, 87(1):115–131, April 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-010-0317-2.

Zamzami:2017:IIC

[ZS17] Nuha Zamzami and Andrea Schiffauerova. The impact of individual collaborative activities on knowledge creation and transmission. Scientometrics, 111(3):1385–1413, June 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Zaharie:2018:NMR

[ZS18] Monica Aniela Zaharie and Marco Seeber. Are nonmonetary rewards effective in attracting peer reviewers? A natural experiment. Scientometrics, 117(3):1587–1609, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2912-6;   http://link.
springer.com/content/pdf/10.1007/s11192-018-2912-
6.pdf.

Zhang:2018:SBK

[ZSC18] Yu Zhang, Morteza Saberi, and Elizabeth Chang. A semantic-based knowledge fusion model for solutionoriented information network development: a case study in intrusion detection field. Scientometrics, 117(2):857– 886, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2904-6.

Zhang:2018:ICR

[ZSCR^{+}18] Lin Zhang, Beibei Sun, Zaida Chinchilla-Rodríguez, Lixin Chen, and Ying Huang. Interdisciplinarity and collaboration: on the relationship between disciplinary diversity in departmental affiliations and reference lists. Scientometrics, 117(1):271–291, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2853-0.

Zong:2013:DDL

[ZSY^{+}13] Qian-Jin Zong, Hong-Zhou Shen, Qin-Jian Yuan, Xiao-Wei Hu, Zhi-Ping Hou, and Shun-Guo Deng. Doctoral dissertations of library and information science in China: A co-word analysis. Scientometrics, 94(2):781–799, February 2013. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-012-0799-1.

Zhai:2014:FCC

[ZSY14] Qinghua Zhai, Jing Su, and Minghai Ye. Focus on China: the current status of entrepreneurship research in China. Scientometrics, 98(3):1985–2006, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1114-5.

Zhou:2014:FCR

[ZT14] Ping Zhou and Huibao Tian. Funded collaboration research in mathematics in China. Scientometrics, 99(3):695–715, June 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1212-4.

Zhang:2018:HRP

[ZT18] Guiyang Zhang and Chaoying Tang. How R&D partner diversity influences innovation performance: an empirical study in the nano-biopharmaceutical field. Scientometrics, 116(3):1487–1512, September 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-018-2831-6. See correction [ZT19].

Zhang:2019:CHR

[ZT19] Guiyang Zhang and Chaoying Tang. Correction to: How R&D partner diversity influences innovation performance: an empirical study in the nano-biopharmaceutical field. Scientometrics, 120(3):1507, September 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03050-x;   http://link.springer.
com/content/pdf/10.1007/s11192-019-03050-x.pdf. See

[ZT18].

Zopiatis:2015:PPF

[ZTC15] Anastasios Zopiatis, Antonis L. Theocharous, and Panayiotis Constanti. ‘The past is prologue to the future’: an introspective view of hospitality and tourism research. Scientometrics, 102(2):1731–1753, February 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1431-3.

Zou:2018:CPT

[ZTP18a] Christopher Zou, Julia Tsui, and Jordan B. Peterson. Correction to: The publication trajectory of graduate students, post-doctoral fellows, and new professors in psychology. Scientometrics, 117(2):1311, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-

2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2576-7;   http://link.
springer.com/content/pdf/10.1007/s11192-017-2576-

7.pdf. See [ZTP18b].

Zou:2018:PTG

[ZTP18b] Christopher Zou, Julia Tsui, and Jordan B. Peterson. The publication trajectory of graduate students, postdoctoral fellows, and new professors in psychology. Scientometrics, 117(2):1289–1310, November 2018. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.

1007/s11192-017-2540-6. See correction [ZTP18a].

Zharova:2018:HMP

[ZTRH18] Alona Zharova, Janine Tellinger-Rice, and Wolfgang Karl Härdle. How to measure the performance of a Collaborative Research Center. Scientometrics, 117(2):1023– 1040, November 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2910-8.

Zuccala:2010:MRS

[Zuc10] Alesia Zuccala. The mathematical review system: does reviewer status play a role in the citation process? Scientometrics, 84(1):221–235, July 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s11192-010-
0161-4.pdf.

Zhang:2011:SBS

[ZVC11] Jian Zhang, Michael S. Vogeley, and Chaomei Chen. Scientometrics of big science: a case study of research in the Sloan Digital Sky Survey. Scientometrics, 86(1):1–14, January 2011. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0318-1.

Zhao:2011:VRP

[ZW11] Rongying Zhao and Ju Wang. Visualizing the research on pervasive and ubiquitous computing. Scientometrics, 86 (3):593–612, March 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0283-8.

Zhao:2014:STA

[ZW14] Rongying Zhao and Shengnan Wu. Study on themes and authors’ influence of open access in China. Scientometrics, 101(2):1165–1177, November 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1415-3.

Zhang:2017:ESI

[ZW17a] Ben Zhang and Xiaohong Wang. Empirical study on influence of university–industry collaboration on research performance and moderating effect of social capital: evidence from engineering academics in China. Scientometrics, 113 (1):257–277, October 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-017-2464-1.

Zhao:2017:AIE

[ZW17b] Rongying Zhao and Mingkun Wei. Academic impact evaluation of Wechat in view of social media perspective. Scientometrics, 112(3):1777–1791, September 2017. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-017-2439-2.

Zhao:2017:IEO

[ZW17c] Rongying Zhao and Mingkun Wei. Impact evaluation of open source software: an Altmetrics perspective. Scientometrics, 110(2):1017–1033, February 2017. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2204-y.

Zhang:2018:PGG

[ZW18a] Li Zhang and Erin Watson. The prevalence of green and grey open access: Where do physical science researchers archive their publications? Scientometrics, 117(3):2021–2035, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-018-2924-2.

Zhang:2018:WHC

[ZW18b] Liwei Zhang and Jue Wang. Why highly cited articles are not highly tweeted? A biology case. Scientometrics, 117 (1):495–509, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2876-6.

Zhao:2019:ECI

[ZW19] Rongying Zhao and Xu Wang. Evaluation and comparison of influence in international Open Access journals between China and USA. Scientometrics, 120(3):1091–1110, September 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03159-z.

Zhu:2013:SWP

[ZWHH13] Danhao Zhu, Dongbo Wang, Saeed-Ul Hassan, and Peter Haddawy. Small-world phenomenon of keywords network based on complex network. Scientometrics, 97(2):

435–442, November 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1019-3.

Zhu:2018:NML

[ZWL^{+}18] Jia Zhu, Xingcheng Wu, Xueqin Lin, Changqin Huang, Gabriel Pui Cheong Fung, and Yong Tang. A novel multiple layers name disambiguation framework for digital libraries using dynamic clustering. Scientometrics, 114(3):781–794, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-017-2611-8.

Zheng:2015:BAI

[ZWW^{+}15] Tianlong Zheng, Juan Wang, Qunhui Wang, Chunhong Nie, Nicholas Smale, Zhining Shi, and Xiaona Wang. A bibliometric analysis of industrial wastewater research: current trends and future prospects. Scientometrics, 105(2): 863–882, November 2015. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-015-1736-x.

Zheng:2016:BAM

[ZWW^{+}16] Tianlong Zheng, Juan Wang, Qunhui Wang, Chunhong Nie, Zhining Shi, Xiaona Wang, and Zhen Gao. A bibliometric analysis of micro/nano-bubble related research: current trends, present application, and future prospects. Scientometrics, 109(1):53–71, October 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2004-4.

Zhang:2018:BAH

[ZWW^{+}18] Nan Zhang, Shanshan Wan, Peiling Wang, Peng Zhang, and Qiang Wu. A bibliometric analysis of highly cited papers in the field of economics and business based on the Essential Science Indicators database. Scientometrics, 116(2): 1039–1053, August 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2786-7.

Zhang:2022:CEA

[ZWX22] Xi Zhang, Xianhai Wang, and Hui Xiong. Correction to: An effectiveness analysis of altmetrics indices for different levels of artificial intelligence publications. Scientometrics, 127(3):1657, March 2022. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL

https://link.springer.com/article/10.1007/s11192-

021-04146-z. See [ZWZ^{+}19].

Zhang:2019:EAA

[ZWZ^{+}19] Xi Zhang, Xianhai Wang, Hongke Zhao, Patricia Ordóñez de Pablos, Yongqiang Sun, and Hui Xiong. An effectiveness analysis of altmetrics indices for different levels of artificial intelligence publications. Scientometrics, 119(3):1311–1344, June 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03088-x. See correction

[ZWX22].

Zhang:2010:BAW

[ZXH10] Gangfeng Zhang, Shaodong Xie, and Yuh-Shan Ho. A bibliometric analysis of world volatile organic compounds research trends. Scientometrics, 83(2):477–492, May 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-009-0065-3.

Zhou:2014:LPW

[ZXLX14] Zhiwei Zhou, Rui Xing, Jing Liu, and Feiyue Xing. Landmark papers written by the Nobelists in physics from 1901 to 2012: a bibliometric analysis of their citations and journals. Scientometrics, 100(2):329–338, August 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-014-1306-7.

Zhang:2016:SDB

[ZXM^{+}16] Wei Zhang, Xiaolin Xu, Chenghan Ming, Zijun Mao, Jing Shi, and Yaqian Xiang. Surviving in the dispute: A bibliometric analysis of global GMF-related research, 1995– 2014. Scientometrics, 109(1):359–375, October 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-1995-1.

Zong:2019:IVA

[ZXT^{+}19] Qianjin Zong, Yafen Xie, Rongchan Tuo, Jingshi Huang, and Yang Yang. The impact of video abstract on citation counts: evidence from a retrospective cohort study of New Journal of Physics. Scientometrics, 119(3):1715–1727, June 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-019-03108-w. See response

[Zon19].

Zhang:2017:SBM

[ZXZ17] Leihan Zhang, Ke Xu, and Jichang Zhao. Sleeping beauties in meme diffusion. Scientometrics, 112(1):383–402, July 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Zhu:2015:DSA

[ZY15] Yongjun Zhu and Erjia Yan. Dynamic subfield analysis of disciplines: an examination of the trading impact and knowledge diffusion patterns of computer science. Scientometrics, 104(1):335–359, July 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1594-6.

Zhang:2019:RIP

[ZY19] Xiaoqian Zhang and Feng Yang. Rural informatization policy evolution in China: a bibliometric study. Scientometrics, 120(1):129–153, July 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03105-z.

Zyczkowski:2010:CGW

[Zyc10] Karol Zyczkowski. Citation graph, weighted impact factors and performance indices. Scientometrics, 85(1):301–315, October 2010. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-010-0208-6.

Zhang:2017:IPD

[ZYF^{+}17] Guijie Zhang, Guang Yu, Yuqiang Feng, Luning Liu, and Zhenhua Yang. Improving the publication delay model to characterize the patent granting process. Scientometrics, 111(2):621–637, May 2017. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

Zhang:2015:SRS

[ZYG15] Jingjing Zhang, Yan Yan, and Jiancheng Guan. Scientific relatedness in solar energy: a comparative study between the USA and China. Scientometrics, 102(2):1595–1613, February 2015. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-014-1487-0.

Zhang:2018:JMA

[ZYNZ18] Fengqing Zhang, Erjia Yan, Xin Niu, and Yongjun Zhu. Joint modeling of the association between NIH funding and its three primary outcomes: patents, publications, and citation impact. Scientometrics, 117(1):591–602, October 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-018-2846-z.

Zhu:2016:UEA

[ZYS16] Yongjun Zhu, Erjia Yan, and Min Song. Understanding the evolving academic landscape of library and information science through faculty hiring data. Scientometrics, 108(3):1461–1478, September 2016. CODEN SC-NTDX. ISSN 0138-9130 (print), 1588-2861 (electronic).

URL http://link.springer.com/accesspage/article/
10.1007/s11192-016-2033-z.

Zhai:2014:EAI

[ZYSS14] Li Zhai, Xiangbin Yan, Joshana Shibchurn, and Xiaohong Song. Evolutionary analysis of international collaboration network of Chinese scholars in management research. Scientometrics, 98(2):1435–1454, February 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-013-1040-6.

Zhao:2016:GPS

[ZYT^{+}16] Star X. Zhao, Shuang Yu, Alice M. Tan, Xin Xu, and Haiyan Yu. Global pattern of science funding in economics. Scientometrics, 109(1):463–479, October 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/accesspage/
article/10.1007/s11192-016-1961-y.

Zhu:2014:RHN

[ZYX^{+}14] Jia Zhu, Yi Yang, Qing Xie, Liwei Wang, and Saeed-Ul Hassan. Robust hybrid name disambiguation framework for large databases. Scientometrics, 98(3):2255–2274, March 2014. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-013-1151-0.

Zhai:2014:III

[ZYZ14] Li Zhai, Xiangbin Yan, and Bin Zhu. The H_{l}-index: improvement of H-index based on quality of citing papers. Scientometrics, 98(2):1021–1031, February 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1039-z.

Zheng:2011:IEK

[ZyZZ^{+}11] Jia Zheng, Zhi yun Zhao, Xu Zhang, Dar zen Chen, Mu hsuan Huang, Xiao ping Lei, Ze yu Zhang, Yun hua Zhao, and Run sheng Liu. Industry evolution and key technologies in China based on patent analysis. Scientometrics, 87(1):175–188, April 2011. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-010-0316-3.

Zheng:2014:ICD

[ZyZZ^{+}14] Jia Zheng, Zhi yun Zhao, Xu Zhang, Dar zen Chen, and Mu hsuan Huang. International collaboration development in nanotechnology: a perspective of patent network analysis. Scientometrics, 98(1):683–702, January 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1081-x.

Zhao:2011:MKD

[ZZ11] Limei Zhao and Qingpu Zhang. Mapping knowledge domains of Chinese digital library research output, 1994– 2010. Scientometrics, 89(1):51–87, October 2011. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-011-0428-4.

Zhang:2014:PAA

[ZZ14] F. Zhang and X. Zhang. Patent activity analysis of vibration-reduction control technology in high-speed railway vehicle systems in China. Scientometrics, 100(3):723– 740, September 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1318-3.

Zhou:2015:GLR

[ZZ15] Xiaodan Zhou and Guohui Zhao. Global liposome research in the period of 1995–2014: a bibliometric analysis. Scientometrics, 105(1):231–248, October 2015. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1659-6.

Zhao:2016:EAC

[ZZ16] Yuehua Zhao and Rongying Zhao. An evolutionary analysis of collaboration networks in scientometrics. Scientometrics, 107(2):759–772, May 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1857-x.

Zhao:2018:NCM

[ZZD^{+}18] Shu Zhao, Dong Zhang, Zhen Duan, Jie Chen, Yan ping Zhang, and Jie Tang. A novel classification method for paper-reviewer recommendation. Scientometrics, 115(3): 1293–1313, June 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2726-6.

Zhou:2016:RSP

[ZZFD16] Jianlin Zhou, An Zeng, Ying Fan, and Zengru Di. Ranking scientific publications with similarity-preferential mechanism. Scientometrics, 106(2):805–816, February 2016. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-015-1805-1.

Zhou:2018:IIS

[ZZFD18a] Jianlin Zhou, An Zeng, Ying Fan, and Zengru Di. Identifying important scholars via directed scientific collaboration networks. Scientometrics, 114(3):1327–1343, March 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-017-2619-0.

Zhou:2018:RWS

[ZZFD18b] Jianlin Zhou, An Zeng, Ying Fan, and Zengru Di. The representative works of scientists. Scientometrics, 117(3): 1721–1732, December 2018. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-018-2918-0.

Zhang:2010:EPA

[ZZL^{+}10] Ling Zhang, Huan Zhao, Qiushi Li, Juan Wang, and Xin Tan. Establishment of paper assessment system based on academic disciplinary benchmarks. Scientometrics, 84(2): 421–429, August 2010. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-009-0132-9.

Zhang:2019:PLS

[ZZL19] Yi Zhang, Fen Zhao, and Jianguo Lu. P2V: large-scale academic paper embedding. Scientometrics, 121(1):399–432, October 2019. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-019-03206-9.

Zhao:2019:MAI

[ZZLS19] Fen Zhao, Yi Zhang, Jianguo Lu, and Ofer Shai. Measuring academic influence using heterogeneous author-citation networks. Scientometrics, 118(3):1119–1140, March 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03010-5.

Zhang:2014:THI

[ZZP^{+}14a] Yi Zhang, Xiao Zhou, Alan L. Porter, Jose M. Vicente Gomila, and An Yan. Triple helix innovation in China’s dye-sensitized solar cell industry: hybrid methods with semantic TRIZ and technology roadmapping. Scientometrics, 99(1):55–75, April 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-013-1090-9.

Zhou:2014:PAM

[ZZP^{+}14b] Xiao Zhou, Yi Zhang, Alan L. Porter, Ying Guo, and Donghua Zhu. A patent analysis method to trace technology evolutionary pathways. Scientometrics, 100(3):705– 721, September 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1317-4.

Zhang:2014:HCT

[ZZPG14] Yi Zhang, Xiao Zhou, Alan L. Porter, and Jose M. Vicente Gomila. How to combine term clumping and technology roadmapping for newly emerging science & technology competitive intelligence: “problem & solution” pattern based semantic TRIZ tool and case study. Scientometrics, 101(2): 1375–1389, November 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-014-1262-2.

Zhao:2014:BDM

[ZZW14] Limei Zhao, Qingpu Zhang, and Liang Wang. Benefit distribution mechanism in the team members’ scientific research collaboration network. Scientometrics, 100(2):363–389, August 2014. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.springer.

com/article/10.1007/s11192-014-1322-7. See erratum

[ZZW16].

Zhao:2016:EBD

[ZZW16] Limei Zhao, Qingpu Zhang, and Liang Wang. Erratum to: Benefit distribution mechanism in the team members’ scientific research collaboration network. Scientometrics, 108 (2):1001, August 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s11192-014-1403-

7.pdf. See [ZZW14].

Zhao:2019:CKC

[ZZW19a] Liming Zhao, Haihong Zhang, and Wenqing Wu. Cooperative knowledge creation in an uncertain network environment based on a dynamic knowledge supernetwork. Scientometrics, 119(2):657–685, May 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03049-4.

Zhu:2019:ISD

[ZZW^{+}19b] Lin Zhu, Donghua Zhu, Xuefeng Wang, Scott W. Cunningham, and Zhinan Wang. An integrated solution for detecting rising technology stars in co-inventor networks. Scientometrics, 121(1):137–172, October 2019. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-019-03194-w.

Zhou:2013:BIC

[ZZY13] Ping Zhou, Yongfeng Zhong, and Meigen Yu. A bibliometric investigation on China–UK collaboration in food and agriculture. Scientometrics, 97(2):267–285, November 2013. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861

(electronic). URL http://link.springer.com/article/
10.1007/s11192-012-0947-7.

Zhang:2019:TSG

[ZZY19] Helena H. Zhang, Alesia A. Zuccala, and Fred Y. Ye. Tracing the ‘swan groups’ of physics and economics in the key publications of Nobel Laureates. Scientometrics, 119(1): 425–436, April 2019. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-019-03036-9.

Zheng:2012:IST

[ZZZ^{+}12] Jia Zheng, Zhi-Yun Zhao, Xu Zhang, Dar-Zen Chen, Mu-Hsuan Huang, Xiao-Ping Lei, Ze-Yu Zhang, and Yun-Hua Zhao. International scientific and technological collaboration of China from 2004 to 2008: a perspective from paper and patent analysis. Scientometrics, 91(1):65–80, April 2012. CODEN SCNTDX. ISSN 0138-9130 (print),

1588-2861 (electronic). URL http://link.springer.com/
article/10.1007/s11192-011-0529-0.

Zheng:2014:ICM

[ZZZ^{+}14] Jia Zheng, Zhiyun Zhao, Xu Zhang, Mu hsuan Huang, and Dar zen Chen. Influences of counting methods on country rankings: a perspective from patent analysis. Scientometrics, 98(3):2087–2102, March 2014. CO-DEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (elec-

tronic). URL http://link.springer.com/article/10.
1007/s11192-013-1139-9.

Zhou:2016:MBI

[ZZZC16] Qingqing Zhou, Chengzhi Zhang, Star X. Zhao, and Bikun Chen. Measuring book impact based on the multigranularity online review mining. Scientometrics, 107(3): 1435–1455, June 2016. CODEN SCNTDX. ISSN 0138-9130 (print), 1588-2861 (electronic). URL http://link.

springer.com/article/10.1007/s11192-016-1930-5.